A Complete Bibliography of Publications in Journal of Computational and Applied Mathematics (1990–1999)

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

04 September 2023 Version 1.05

Title word cross-reference

(−∞, ∞) [LW91]. (P ± i0) [TB99]. (P ± i0)^{λ} [TB99]. (ψ(x) − ψ(y >))^{2} [De 91a]. 1 [Dor99a, RS99]. 1/π [BB93]. 1/\sqrt{1 − s^{2}} [RP96]. $107.25 [Bul90b]. $162.75 [Ver94b]. $168.50 [Ver94b]. 2 [DPS98, DV99, DM97a, HS92a, Lay93, Sak99, VDM96, WF97, dS97b]. $24.95 [Eme94e]. 2^{α} [EHGNT90]. 2 × 2 [Leó99]. 3 [CM96d, Eme94c, HKW99, Kle99]. $43.95 [Ano94j]. 7 [KN90]. $92.00 [Roo90b]. ^{(1)} [Sid99]. ^{1} [Ber94, CM96b, DOS94, Jun92, Man96, VM94]. ^{2} [Ber93b, HS94b]. ^{3} [SK96, Zha95]. ^{4} [SS91b]. _{1} [CMR99, Zei92]. 10 [MN96]. _{2} [Köl95a, LGR96, Lew97a, Zei92].

{_{2}Φ_{0}((A, α_{j}), (1, 1); z)}_{j}^{n} = 1 [DDL93]. _{3} [LGR96, Lew97a]. _{3}F_{2} [Ext99, Mil06]. _{4} [Sch94b]. _{8}Ø_{7} [GM95]. _{9} [MN96]. _{q+1}F_{q} [KK97]. _{t} [SS91b]. _{xx} [SS91b]. A [Bar99b]. A_{0} [CMR93]. B [Ano91s]. C [Ber94, CLR97, CM96b, DOS94, HS94b, Jun92, Man96, Pas92, SK96, Zha95, Li96]. C^{1} [Pfe91]. C^{2} [Lai96]. c_{0} [Rha92]. {C_{n}^{(−N+12)}}_{n} ≥ 0 [dMPP98]. D [CS97c, Sid97, Sid99, CRD94, Dou96]. D_{ω} [FHR98]. δ [BCCM91]. δ^{(k)}(P ± i0 − m^{2}) [Tél98]. E [Gra96, Vek97]. e^{z} [BPV93]. E_{n}^{r} [Spa99]. E_{n}^{r^{2}} [Spa99]. ϵ [LC97, Sal93]. F [LGR96, Let98, Lew97a, Zei92]. F_{1}(a, b, b′; c; x, y) [CDTV99]. F_{n}(z) = F_{n−1}(f_{n}(z)), f_{n} → f [Gil90]. G [Ada95, Lev90b, Lev94]. G^{2} [FK96, PTP99].

G_{2} [Ekh91]. g_{i}(x) = x^{α_{i}}(− ln x)^{β_{i}} [Gra96]. γδ(r − a) [Goo91]. H [Ber93b, KNO97, Sab96, CZS96, DOAG91, GC97, LM97, SS91b, Sur91]. H/h [KV95]. h^{4} [Moh97]. H^{∞} [Ano97m]. H_{μ}(z) [KSK92]. θ̂ [Fdi99]. hp [Mon94b]. I [Daa96, DL94a, GM90b, Hür97a, Leo98b, MBP93, Mas92a, Sch94c]. j [Hom96]. J^{′_{μ}(z)} [KSK92, KSK92]. J_{0}(x) [Let96]. j_{pq} [FFK97]. J_{v}(x) [LW96b]. k [HY98, KW99, LC97, TB99]. K(1/b_{n}) [Lan99a]. L [FLM92, Sch94b, VM94, dS97a]. L^{2} [GWC96, Joh98, Zin99]. L_{2} [BI96, RS92, Wen97]. λ [MPSM93]. λ_{n} ↘ 0 [Zho94]. ln Γ(x) [Köl94b]. log_{e}^{2}(1 + x) [MM90]. M [Bec90, Bec92, BM95b, DTB93, DM97b, He94, dS97a]. m_{α}(λ) [WGK99]. N [LM90, Mey94, Spu99, Tok96a, Che98a, CH94b, DDL93, Dub98, GK96c, HY98, SA97, TH94]. n → ∞ [Voi95]. O(τ^{2} + h^{4}) [Sty96]. P [AS96, CAZ98, CMR99, FP90, IP99, VVVD99b, VB93, Che97a, CLL98, DOAG91, GMdLSG99, Gri99, GC97, HN95, HNT96, HNT98, LM97, LM95, MP95, Med98a, Mon94b, Sur91]. P_{1}P_{0} [Joh98]. P_{n}(x) [LW95]. π [LM90]. ψ(x) − ψ(y >) [De 91a]. ψ^{(k)}(x) [Köl96b]. Q [FGKP95, ÁNQR99, ARS96, AAVB98, CSE95, CIM94, DDL93, FV95, FV96, Flo96, FRK99, GH96, GIM96, KM96, Koo93, Sch93, Zen95]. QR [BI92, Dub97, FFH97]. q → 1^{−} [Elb92]. R [AH95, CS95b, BBH90]. R^{n} [Bul90b]. ρ [Car90]. S [GM97, dS97a]. T_{+m} [SG94]. θ [HJ95b, Liu96, Lu91]. T QR [GG97b]. U [GK96c, SS91b]. U(n) [MN96]. U(n)||U(n − 1) [Voi95]. V [Axe96]. W [Sha97b]. X [CLR97, De 91a, Köl95a, Let98, Zei92]. x(s) = q^{2s} [Lew98]. x = 1/4 [Köl96b]. x = 3/4 [Köl96b]. x^{d} = 0 [Yam97]. {x_{n}^{λ}} [Zho94]. X_{t} = X_{t−3p} ⊕ X_{t−3q} [Fus90]. y [De 91a]. y′(t) + y(qt) = 0 [Elb92]. y′′ = f(t, y, y′) [RA96]. y′′ = f(x, y)

[IR97, XZ98, IP99]. Z [KNO96, Rös95b]. z^{3} [Yam96]. ζ(2) [Pré96]. ζ(3) [Pré96].

-1 [Dam99]. -adaptive [CZS96]. -algebra [KM96]. -algorithm [Car90, Fdi99, Sal93, Gra96]. -analogue [CSE95, Koo93]. -analogues [ÁNQR99]. -approximation [BI96]. -beta [FV96]. -body [Mey94, Spu99]. -Charlier [Zen95]. -classical [FRK99]. -component [TH94]. -continued [Lev90b, Lev94]. -control [Ano97m]. -convergence [CS97c]. -curved [Ber94]. -cycle [Axe96, Sha97b]. -cyclic [HN95, HNT96, HNT98]. -D [Eme94c, CM96d, Kle99, Lay93, Sak99]. -difference [FRK99]. -dimensional [CH94b, LM90]. -disk [Flo96]. -divergence [MPSM93]. -elementary [FFK97]. -error [Joh98]. -extension [Mon94b]. -finite [Joh98]. -fold [He94]. -forms [VB93]. -Fourier [ARS96]. -fractions [BCCM91, Li96]. -function [Ada95, FGKP95]. -gamma [FV96]. -Hahn [GIM96]. -harmonic [CLL98]. -Hermite [AAVB98]. -hierarchical [DOS94]. -hypergeometric [DDL93]. -independent [KV95]. -ISBN [VR91]. -Laguerre [CIM94, Zen95]. -Laplacian [GMdLSG99, Med98a]. -Legendre [Sch93]. -matrices [KNO96, KNO97]. -method [Liu96]. -methods [HJ95b, Lu91]. -mutual [MPSM93]. -operators [BM95b]. -order [AH95]. -orthogonal [Dou96]. -orthogonality [GM97]. -Padé [Bec90, Bec92]. -point [SA97, Tok96a]. -prediction [Vek97]. -reproducing [CMR99]. -semigroups [Rha92]. -sphere [KN90]. -spline [HS94b, SK96, Pfe91]. -splines [dS97a, Sab96]. -stability [CMR93, Zin99]. -stable [VVVD99b, AS96, Bar99b, CAZ98, FLM92, FP90, IP99]. -stage [dS97b]. -step [HS92a]. -Stirling [Zen95]. -table [Pas92]. -th [Dub98, Gri99, LM95, TB99]. -th-order [DM97b]. -transformation [CRD94, Sid99, Sid97]. -tuple [DDL93]. -uniform [KW99]. -version [Che97a, GH96, MP95]. -versions [Sur91].

0 [Ano93k, Bul90b, Die94a, Eme94a, Eme94d, Eme94b, Eme94f, Mee96, Roo90b, Uyt96]. 0-13-190075-7 [Uyt96]. 0-19-850024-6 [Ano98q]. 0-262-04142-1 [Ano94j]. 0-262-05048-X [Ano94m]. 0-262-14054-3 [Ano94k]. 0-262-51073-1 [Eme94a]. 0-262-61094-9 [Eme94e]. 0-262-72026-4 [Ano98r]. 0-444-70366-7 [Bul90b]. 0-444-87367-8 [Roo90b]. 0-444-89596-5 [Ver94b]. 0-444-89597-5 [Ver94b]. 0-444-89598-1 [Ver94b]. 0-521-40407-X [Eme94b]. 0-582-27654-3 [Mee96]. 0-7923-4521-5 [Ano98p]. 0-8247-8676-9 [Ano93k]. 0-8493-4417-4 [Ano94l]. 0-8493-7721-8 [Eme94f].

1 [Ano91r, Ano93u, Eme94a, Eme94d, Kri92, Lev90a, Lev91]. 1-56032-104-0 [Eme94d]. 107 [KO01]. 130 [DP94b]. 14 [Ano97-41]. 146 [Ano98p]. 14th [Ano97-41, Ano93a]. 15th [Ano97a]. 16th [Ano93b]. 1991 [Ano91-32]. 1st [Lev91].

2 [Bul94b]. 200A [Ano90k]. 29.95 [Ano94j]. 2d [SR91]. 2ET [Eme94d]. 2nd [Lev91]. 2RU [Eme94b].

3 [Bul94a, Bul94b, Die94a, Eme94c, Lev90a, Lev91, Mee96, VR91]. 3-11-013784-4 [Bul94a]. 3-411-16891-9 [Kra96]. 3-528-07636-4 [Eme94c]. 3-540-12760-7 [VR91]. 3-540-54812-2 [Ver96]. 3-540-55560-9 [Bul94b]. 3-540-57554-5 [Ver96]. 3-540-94047-2 [Pie96]. 3-540-94101-0 [Die94a]. 3-540-97040-1 [Lev90a]. 3-540-97151-3 [Lev91]. 3-540-97152-1 [Lev91]. 3-7643-5091-1 [Bul96a, Bul96c]. 3-7643-5097-0 [Bul96b].

4 [Bul94a, Eme94c]. 48 [Bul96a, Bul96c]. 4th [Ano97r].

51 [DP94b].

60 [Ano98q]. 61 [Kre11]. 68 [Kra96, Ver96]. 6SF [Ano94l, Eme94f].

7 [Bul90b, Uyt96, VR91]. 78 [Ano97-31, Bul96b]. 7ET [Ano94k, Ano94m, Eme94a, Eme94e]. 7th [Ano91-36].

8 [Ano97-46, Eme94f, Roo90b]. 80 [Pie96]. 87 [Alz98]. 89 [Die02].

9 [Ano93k, Bul94b]. ’94 [Ano93l, Ano93c]. 97 [Ano97-47, Ano97-51]. 99 [KR99]. 9th [Ano93j].

= [CLR97, CB92, SS91b, Yam96].

Abel [Abd90, AE91, BCRV91, Cap90]. abscissae [Not97, Not98]. Absolute [BEM98, DP95]. Absorbing [KKS96a, DT96, RP96, Ren97b]. abstract [Mas99]. Academic [Ano98p]. Accelerable [Osa90]. Accelerated [Del95, KNO96, Bai98, BKF^{+}96, GS91a]. accelerating [Kor94b, Sad90, Thu99]. Acceleration [Ben94b, Osa91, Pré94, Ver97, Bel90a, BR92a, BMP96, EH98, Fdi99, GB90a, Gra96, Kza95, Le 95, LB99, LG95, Liu95, Nie95, Osa96, Sed90, SG94, Zha98]. accelerations [CF93]. accuracy [DSH95, GSSV92, LJQ95, Mig95, MJ96, Moh98, Sto97, Tsa96, XZ98, vN93]. Accurate [Lau99, Moo95, Sim98, SS97a, Sun96]. Acknowledgement [Cam91]. Acknowledgment [Rad97]. Acoustic [WRW92, CM92, Ked93, MET94]. acoustics [DOAG91, SB96]. across [vdHSvdV95]. actualization [BCP99]. Adams [CW99, HU99, vdHM99]. adaptations [SR91]. adapted [GH97a]. Adaptive [CR96, CRD94, Mit91, RZ99, Ste99, Bän91, BES91, BH95, BT95, CZS96, CP96, FP97b, Hig93, HF95, KL99, Kle99, Li92, LLM91, RD96, Riv91, TK98, Ver94a]. adaptively [QS99]. Addenda [Ano97-31]. addition [Koe96]. Additional [LM93a, RA97, RW97, Thr99, VV95]. Additive [RA96]. adequate [Bab99]. ADI [Sta93]. adjacent [Hoc97]. adjoint [BS94, DM94, Müh96]. adjustment [Seg95]. Advance [Ano98p]. Advanced [Ano91z, Ano91y, Ano97-30, Li99b]. Advances [Bul96b, Bul96c, Bul96a, Pry94, Eme94b]. advection [AP96, ATV98, FP99, LV99, Ram99, Sac96, Sak99, Wan97]. advection-dominated [Sac96]. aerodynamic [vdVNTV97]. aerodynamics [SB96]. affected [eM91]. affine [AC95, CK91, DG92, KK95]. Agarwal [Ano93k, Ano98p]. age [Cus94]. age-structured [Cus94]. aggregation [Lit92, Lit95]. ahead [Hoc97, VB97]. Aided [Ano93-30]. AIDS [GY95]. air [LV99]. airfoil [BHS92, HS94a]. Airy [Mac94]. Albert [Ano94k]. Algebra [Ano97-48, Ano97-49, Ano98r, FV95, FV96, KM96, Bul94a]. Algebraic [Ano97o, Ano97-46, Bak99, Sha97a, AM97, AP92b, AL98, Cas96, DR90, KM94b, Lew91, Lop90, Ren97a, RS99, SD98, Smi91, Ver96, YC90, BGVO95]. algebraically [BMP96]. algebras [Fay97, HV91]. Algorithm [Vek97, AP98b, AR92a, AGR93, AC95, AD92b, Ari95, BCCM91, BHS92, BM96b, Bog98, BI92, CCJ97, Car90, CZ90, CK91, CDM91, CM96d, Cri97, DBFV96, DF91, Dub97, DRW95, Fdi99, Fuk90, GM99, GGF99, Gra96, HGL96, HZ96, HR99, HL95, Koo93, Koo98, LGT90, LG95, LR96a, MH94, McC98, Mes96a, Mes96b, MKK96, Miy93, Moh90, PWT99, Raf94, RF99b, RF99a, Rat94, Ren96, RW97, Sal93, SK90, SW95b, Sto95, SS97d, TPA95, VJ96, Van90a, Van90b, Ven93, Wol94, YM99, ZP95, vG95, vdFvGdRdV95, vdVNTV97, Bul94c]. Algorithmic [BBD^{+}96, Hür97a]. Algorithms [Ano96-29, Ano97-48, Eme94a, FV91, ATV98, AN95, BES91, BF93, Bog97, BM96c, DHP94, DT93, Ell95, Fay97, FRR94, GG97b, Heu92, HF95, Kem90b, Kem90a, KKNU97, Kra96, Lab97, Lab98, LSTZ94, MO96b, Mat90, May95, PP91, Pop96, Pry94, RH99, SÁFV96, TG91, Wan94, YR99, Yey92, ZZ91, dR99, Roo90b]. Allee [Cus94]. Allgower [VR91]. alloys [VV98]. almost [BM93a, BC94, CAZ98, GP93, VVVD99b]. Alternating [FM97, Han91a, JW91, YK96]. alternating-type [YK96]. alternative [GWC96, IS95a]. aluminium [DJR91]. ambient [CSS93]. American [ZFV98]. Ami [Eme94a]. amplification [BM95a, BM95b]. Amsterdam [Bul90b]. analogue [AGR93, CSE95, Esp99, Koe96, Koo93]. analogues [ÁNQR99, MMST96]. analogy [Ros95a]. analyses [Lis95]. Analysis [lCHcH97a, CB92, CSvD98, DGH95, LV99, Lev97, Lev91, QS99, RVY97, Rob98, eMB91, AC95, And95, Ano97-34, Ano97-35, Bab96, BOS97, BS93a, BK95b, BW91, Bul90b, Cha96, Chr90, CK91, CH94a, CRV90, CRV97, De 99, DT93, DKKV98, Dun97, DS96b, Eir93, Ell94, FPR92, GG97a, GMdLSG99, GJ95, Gje97, HLFC99, JP91, JTNW93, Jun94, KT92, LB98, LWZ98, LR96c, MT98, MET94, MJG98, Mon93, Ned94, NS95, Ned95b, NL94, Pet99a, Pry94, SV92a, SZ97, SF90, SR91, TF91, Tic90, WGK99]. Analytic [JNC99, MBMF99, Ata94, BL93, Elb92, GL90a, GSSV92, JS92, Kza95, PH92, SL99, Sch96b, Sza93]. Analytic-numerical [JNC99]. Analytical [BK90, Hom96, Jód92, KSFAH98, PB98, ADG99, Chr98, TD97]. analyzing [AV96]. and/or [Ask93, Riv91]. Andreas [Kra96]. Andrews [EHK^{+}95]. Angewandte [Ano91r, Ano93u]. angular [KM99]. anharmonic [Wit95]. Animating [Ano97b]. annealing [TPA95]. annihilation [YM99]. announcement [Ano97q]. Announcements [Ano94a, Ano90a]. annoy [Ask93]. answer [BB95]. Anti [APH94, Han92]. Anti-periodic [APH94]. any [ZRG93]. aperiodic [Zie90]. Apéry [All97, Rad95]. APL [Ano93c]. Appell [Bel93a, CDTV99, Dou96]. Appendix [Ano94b]. Appl [Alz98, Ano97-31, Die02, DP94b, KO01, Kre11, KR99]. Applicability [ZV95]. Application [BGN^{+}90, BH95, HT91, JV96, ADH98b, ATBB99, AEG98, BPV93, BHD93, CH98, Die96, DK94, För93, GMBW96, HS97, Ioa90, LW99, LL99a, Mat90, NS91, NL94, PS94, Pas91, Sid99, Sim92, TA95, XZ98, YC90, BK95a]. Applications [Ano91g, Ano93c, Ano93z, Ano96-29, Bul96b, Bul96c, DL94a, DL94b, Drm90, Dup94, HB92b, HKW99, Hür97b, AG98, ADG99, And96, Arg98, Bar99a, Ben94a, Ber94, BW91, Bre98, Bul90a, Bul96a, CZ94, CZ95, DJ98, DP97, DF95, EHK^{+}95, Erd93, Fro91, Hig92, Ito94, Li92, Mil93, Miy93, Müh90, NV94, SD98, Sch98, SWW96, SDL99, Shi99, SV96, Sto99, TE97, Tic90, Tri93, XM98, Ano91y, Ano91-35, Ano91-36, Ano93-28, Ano97-30, Eme94d, Pie96]. Applied [Ano90m, Ano91-27, Ano93a, Ano93-27, Ano97a, Bul91, Die94a, Ano90o, Ano90n, Ano91-28, Ano91-29, Ano92j, Ano92k, Ano92l, Ano94-27, Ano94-28, Ano95z, Ano95-27, Ano96-30, Ano96-31, Ano96-32, Ano97-43, Ano97-44, Ano97-45, Ano98-32, Ano98-33, Ano99z, BD90, CY91, CH94a, FS90, GR92a, Hyv97, IO91, JTNW93, OO96, Pen99, Sch94c, Sch94d, SFGF^{+}99, Yam96, Yam97]. Approach [Ano97-49, Abd90, AG98, Ano97m, Bab98, BBD^{+}96, BLG91, BJM97, CS93a, CDL96b, DR90, DDG98, Ext91, Ext95, Faz94, FS96a, FV95, FR99, GMS94, HGC99, Ise93, JS92, Kos94, LYFW95, Lit95, NR90, NM92, Ozi91a, Sta90, TG91, TD97, TE97, VB90a, VB90b, Vog94, Wit96a, lX97, YR99, Zei90]. approached [Pre90]. approaches [Zel95]. Approx [VR91]. approximant [BGM93a, Cuy99, Pan94]. approximants [AP92a, AW95b, BGM90, BGM93b, BGM94, BGM95, Bak99, BSW99, Bec92, BGVHN95a, BGVO95, BDMGVO97a, BDMGVO97b, BGVHN97, BL93, CLB92, CDL96b, CDL96a, DDL93, GM93, GP97, GP98, GP99, GVN90a, GVO94, GM90b, GMS91, GMR94, GMBW96, Gui97, HRGL98, IL99, KW90, Le 95, Mat97, Pre90, Pré96, Rob96, RW92, Sid90, Sta97, Thu99, Tok96a, Tok96b, Voe95, Wim90, dlCYL98, Zho97]. Approximate [A.94, eM91, CNP99, NS91, RS95, TH94, Buo96, GK97, GPCGV91, GB90b, Han91b, Jo’90, Jód92, Leó99, Ren97a, vdHS98]. Approximating [MW95, BC91, PH92, Rab90, Ste95, YD96, Yey92]. Approximation [AK97, AW95b, Ano91z, Ano91y, Ano93w, ADT95, BC98b, CP99, DNZ98, Fun92, GlZ94, LP98, LN95, Pre90, WM94, WM96a, AC90, Aco97a, AY94, AKV92, Anh99, Ano91u, BG95, Bec90, BI96, Bre90c, BV90, CS95a, CDM91, DL96, DC95, DVB96, DG93, DF91, DMGVO99, Dri95, DGH95, Ern96, ELW93, EH97b, FS91, FW99, Fra94, FP99, GVN90b, GMR97, Ill98, Ito94, JS92, Jer91, Jia94, KT93, Lad97, Let96, Lio96, Liu97, LL99a, LSSZ90, MRT95, Mas93, ME93, Mat98, MW90b, McC93, MW99b, MR92a, MR92b, Mon96b, Ned97b, Ors90, Ors97, Ozi92, Pas90, Per92, Ree94, Rin94, RT91, SSS96, SB99, SS94c, Sta98, TD90, VB94, VVD94, VM94, VMV94, Wen97, Zho94, Bul90a]. Approximations [DE90, BE91, Bes96, Ehr94, Jód90, KO99, KO01, Kju95, Let95b, LW95, Let98, LT93, MJG98, OF96, QS99, RP96, SV92b, SL96, VD91, VZ94, Wan94, Wan97]. Approximative [Nee94]. Arbel [Eme94a]. arbitrary [Ad94, Bec92, CLM95, CNS97, Col98, DNZ98, KS99, LS95, Luc95, Mit91, Pan95b, PP96a, Sid97]. arc [De 90a, How93, HSW91, MW95, MW99b, MR94, Mön96a, WM94]. arch [lCHcH97b]. architectures [DT93, Ano94l]. arcs [AK97, WM96a, YD96]. Area [TRBS99, HMT90, Maa94, PP96a]. argument [Cah90, Ext99, Mil06, Zha96b]. arguments [Köl94a, MA98, NPB92]. arising [ABD^{+}96b, BG95, BV99, CM97, Ext98a, JMT94, Ned97b, Thr90]. Aristide [Bul96c]. Arithmetic [Ano91-32, Ano97-51, Sch94a, Sch94f]. arithmetical [Rad95]. ARMA [Tre91]. Arnold [Kra96]. Arnoldi [BGF97, Gra93, Hyv97]. Arov [FFK97]. Arov-normalized [FFK97]. arrangements [Orl99]. Art [Ano94m]. Artificial [Dub96, DD95]. Askey [Ban99, Rah96, RZAG98]. aspect [Ush91]. Aspects [Ano91q, Eir93, ADG99, BGVO95, Mas92a, Mas92b, NNR93, Sid90]. assessment [Aff90]. asset [Pan96c]. Associated [Hen90a, Val95, Bel90b, BM93a, BGVHN95a, BS94, CGV97, CJT90, DLMT99, DT98, Flo96, FKR98, FRK99, GMS94, Hen90b, HV91, JNS90, Let93, Let95a, LRV96, Let99, Lew93, Lew95, Lew97b, Lin98, MR92a, Pan93a, PS95a, Rah96, Rán92, RGZA98, RV98, Sab96, SA97, Sif98, SC95b, Tri93, Van91a, ZRG93, GIM96]. Association [Bul90a]. assumption [GW91]. assumptions [DGC95, Thr99]. astrophysical [SR99, Sto99]. astrophysics [BGM99, IM99, RW99, SS99a]. Asymptotic [And92, BMS98, DP93b, DM97b, DS96b, Ehr95b, ElM94, Elb92, GL94, Guo95a, Hwa96, JP91, Jia94, JMMW90, Kit91, KT92, Lóp99, MF98, Mas99, MW90a, Med95, Med96, NW97, Pao98, PSG91, Rov92, SL98, SFB94, Tem92, Tre92a, WL92, WZ97, BS95, BPP97, Ber92a, BE91, BP96a, BA98, BCG98, Dun97, ES97, GR92a, GK96b, How90, IY97, Kul92, Lim96, MMPP97, Mat90, Mat97, Med97, Med98b, SV95b, SLK91, SLK92, TD90, Tem95, Waz91, YMW98, ZLQ96]. asymptotic-numerical [How90]. asymptotically [BHS^{+}98, Hug94]. Asymptotics [CTV96, CIM94, CI97, DT98, Dun98, MMFMB97, MP99b, Pan93a, PKB98, BR92a, Boe91, BG98, Deh94, IS98, MFMBPC97, PW92, Wim91, dlCYL98]. Asynchronous [BH98, FS95, Bai98, BEC99, Fro91, SBK98, Xu99]. athletic [DNS98]. atmosphere [HB99, WD99]. atomic [SFGF^{+}99]. attenuation [MET93]. Austria [Ano91s]. Author [Ano90b, Ano90c, Ano90d, Ano90e, Ano90f, Ano91a, Ano91b, Ano91c, Ano91d, Ano91e, Ano92a, Ano92b, Ano92c, Ano93d, Ano93e, Ano93f, Ano93g, Ano93h, Ano94c, Ano94d, Ano94e, Ano94f, Ano94g, Ano94h, Ano94i, Ano95a, Ano95b, Ano95c, Ano95d, Ano95e, Ano95f, Ano95g, Ano95h, Ano95i, Ano96a, Ano96b, Ano96c, Ano96d, Ano96e, Ano96f, Ano96g, Ano96h, Ano96i, Ano96j, Ano96k, Ano97c, Ano97d, Ano97e, Ano97f, Ano97g, Ano97h, Ano97i, Ano97j, Ano97k, Ano97l, Ano98d, Ano98e, Ano98f, Ano98a, Ano98g, Ano98h, Ano98i, Ano98j, Ano98k, Ano98l, Ano98m, Ano98n, Ano98o, Ano98c, Ano98b, Ano99a, Ano99b, Ano99c, Ano99d, Ano99e]. automated [PWT99]. Automatic [Bec94, Fis91, MP92, BF94a, Leo98b]. Automation [Ano93l]. automorphisms [KN99]. autoregressive [Mig95, SV95b]. auxiliary [Gra96]. average [LM92, LR96b]. Avoiding [EH98, Sca95, Ste90a]. AXB [CLR97]. Axelsson [Mar95a]. axis [RF99b, RF99a]. Aylesford [Ano94l, Eme94f].

B [Ano90k, Eme94e, Ver94b, PWT99, Str94]. B-3030 [Ano90k]. B-spline [PWT99]. B-splines [Str94]. B.I [Kra96]. back [KV99]. backpacker [Was95]. backprojection [Rat94]. backward [Eir93, Ste90b]. Balanced [Miy92, LR92b]. Ball [ZZ91]. Ballore [Cuy90, Cuy92]. balls [HX99]. Banach [Aco97c, ACQ94, Lad97]. Band [Jin96a]. Banded [Mic92]. Barnes [Par92]. barycentric [Ber97, BM97]. base [Roo90a]. based [Bar90, BP96b, BM96c, BHGVN94, CR96, CH96b, CH97, DL96, DPS98, DNN92, DEP97, DLP98, Die98, Die02, Dos95, EM98a, EK98, Epp91, GS97, GC97, KO99, KO01, Mat90, MKK96, Rab90, Sak99, SP96, San96, Sch90, Seg95, VVD94, VDV90, VVD92, dS96a, dS97b, Fie90]. Basel [Bul91, Bul96a, Bul96b, Bul96c]. bases [Alw96, DOS94, LW96a, Mit91, Mon96b, Sab93, SS94b]. Bashforth [CW99]. Basic [Die95, Eme94d, CIM94, GIM96, GK96c, KKZN97, Koe96, Mar96, Pas90, Ver96]. basis [AE91, FI96, YH95, YA95]. Batchelor [AKM97]. Bayesian [Zel95]. BCG [Cel98]. BDF [CMR93, JC95, Ste90a]. BDM [Che94]. be [Cuy99, Zei92, CZS96]. beach [Ehr94]. beam [lCHcH97b, ESEHEK93, Fen94, JM94, LM94b]. beams [Vrb95, Wes90]. behavior [BA98, DP93b, DM97b, GK96b, Hom96, Med95, Med97, Med98b, Pao98, Tre92a, YMW98]. behaviour [BPP97, BMS98, Ber99, Ehr95b, ElM94, Elb92, FB96, HD93, Hol95, LMM94, PSG91, PS93a, Rov92, vdVV93]. Belgium [Ano90k]. Bell [Köl94a, KS96b]. BEM [DM97a, LN95]. benchmark [Sim96]. Bengt [Van91b]. Benin [Ano93i]. bent [ASWO91]. bent-nose [ASWO91]. Bergman [BW96, PS93a, RNM97]. Berlin [Bul94a, Bul96a, Bul96b, Bul96c, Die94a, Pie96]. Bernstein [GlZ94, JS93, Not90, Raf94, Sab93]. Bernstein-type [JS93]. Besov [Aco97a]. Bessel [AS97, AS99, Ada95, Boe91, CIM94, EEKL93, EM94b, GMS94, IS90, IKK91, IS92, IS93, IKF91, Köl95b, Köl96a, Laf91, LM93a, LW96b, Lew91, LU96, LS95, Luc95, Mar95b, Mav95, MW90a, Mil97c, Mul95, Sid97, SR92, SFB94, WZ97, Zha96b]. Bessel-function [Boe91]. Best [Aco97a, ESS99, Hür97a, Hür97b, JS92, BSW99, Wen97, LW96b, LU96]. Bestness [Lor92]. beta [CQRZ97, FV96, Mil98, Tem92]. BETECH [Ano93j]. between [AD92a, AGRZ97, AGRZ98, Gil99b, GRZA97, GRZA98b, HY98, IY95, Lev90b, LB99, RZG95b, SL97, TB99]. Beyer [EHG90]. beyond [MPPP95]. Bézier [AK97, FK96, LS96, Loe96, PTP99, Raf94, WM94, WM96a, WM96b]. bi [BHS^{+}98]. bi-p [BHS^{+}98]. bias [SV95b]. bibliography [McN93, McN97, McN99]. BICGSTAB [WS97]. biconfluent [AZD91]. biconic [ASWO91]. bidiagonal [Wri93]. bidimensional [ADT97, DJR91]. BIE [Gan93]. Bieberbach [PSG91]. bifurcation [ASW97, FW99, Gov94, LST94, LR96c, RS99, SFW99, Smi96]. bifurcations [JV96, Pel94]. Biharmonic [Li99c, AEG98, CZ90, Chr98, FT98, Kar94, Sca94]. biharmonic-type [Kar94]. bilinear [DE99, Gab92]. bimetal [Jek91]. Binding [BB95]. Bingham [Ned95a]. Bini [Bul94c]. biofilm [CC99]. biology [BB95]. Biorthogonal [RV98, Zhe98, dR99]. bipartite [Pal92]. bipartitioning [Han96]. bipartitionnaires [Han96]. biquartic [Kob95]. Birkhäuser [Bul94c, Bul96a, Bul96c, Bul96b]. Birkhoff [GM90a, Zhu97]. birth [BV95, Val93, Val96]. Bivariate [AY94, Chu97a, CDM91, DNZ98, GM90a, LL91, Zha95]. blackbox [De 90b]. BLAS [GP93]. Bleimann [JS93]. blended [Loe96]. blending [Li99c, LC99a]. Block [SSY98, VB93, Yun98, AM96, AB97b, BEC99, Bur93, BC94, CT99, FKM97, GP93, Jbi99, JC95, Jin96a, KM99, SG96, SBK98, VV96, VB97, Wri93, Xu99, ZYCL97]. block-angular [KM99]. block-bidiagonal [Wri93]. block-iterative [BC94]. block-tridiagonal [Yun98]. blow [BCG98, DKKV98, HMV98, Lac98, MS98b]. blow-up [BCG98, DKKV98, HMV98, Lac98, MS98b]. Blowup [BB98a]. Board [Ano97s, Ano92e, Ano92i, Ano94q, Ano94u, Ano90g, Ano90h, Ano90i, Ano90j, Ano91j, Ano91k, Ano91l, Ano91m, Ano91n, Ano92d, Ano92f, Ano92g, Ano92h, Ano93n, Ano93o, Ano93p, Ano93q, Ano93r, Ano94p, Ano94r, Ano94s, Ano94t, Ano94v, Ano95j, Ano95k, Ano95l, Ano95n, Ano95o, Ano95p, Ano95q, Ano95r, Ano95m, Ano96n, Ano96o, Ano96p, Ano96q, Ano96r, Ano96s, Ano96t, Ano96u, Ano96v, Ano96w, Ano96x, Ano97t, Ano97u, Ano97v, Ano97w, Ano97x, Ano97y, Ano97z, Ano97-27, Ano97-28, Ano97-29, Ano98s, Ano98t, Ano98u, Ano98v, Ano98w, Ano98x, Ano98y, Ano98z, Ano98-27, Ano98-28, Ano98-29, Ano98-30, Ano98-31, Ano99m, Ano99n, Ano99o, Ano99p]. Bochner [GH96]. bodies [Mar97, Sch94b]. body [Gre93, Krí95, Mey94, Spu99]. Bogdanov [JP91]. Boltzmann [BGN^{+}90, Bab99, CL97, Gör99, TK98]. bombardment [BBD^{+}96]. Bonn [Ano91r, Ano93u]. Book [Ano93k, Ano94j, Ano94k, Ano94l, Ano94m, Ano98p, Ano98q, Ano98r, Bul90a, Bul90b, Bul91, Bul94a, Bul94b, Bul94c, Bul96a, Bul96b, Bul96c, Die94a, Eme94a, Eme94d, Eme94e, Eme94b, Eme94c, Eme94f, Kra96, Lev90a, Lev91, Mee96, Pie96, Roo90b, Uyt96, Van91b, VR91, Ver94b, Ver96]. Boolean [GlZ94]. Boor [Die94a]. bordered [Gov94]. Boston [Bul96a, Bul96b, Bul96c]. bottom [eMM97a]. bound [CGR99, DRW95, LS98, RS94]. boundaries [DD95, PP96a, RF99b, RF99a, Ren97b]. Boundary [Ano93b, Ano93j, CG94, DLR92, EK94, Ira94, Jód90, Li99c, LC99a, LN96, MR92a, Waa91, eMB91, ADG99, AKM97, AO98, AGW91, AM96, AM97, AB97b, AAP98, ADT96, ADT97, Anh99, Bab99, BZ93, BV99, BR98, BCV98, BS93b, BF93, BT96, CL99, CS93b, Che97a, Chr90, CH94a, CS96b, Chr98, CLM95, DD95, Deh99, DOAG91, DTB93, Dos95, DT96, Dub96, EGBE93, ES97, Elo98, Ern96, Fea91, FZ93, FW99, GL90b, Gre93, GL98, HB92a, HY98, HKW99, IO91, JK98a, Jód91a, Jon93, JL99, KKS96a, LW96a, LST94, Lim94, LG95, LC99b, LT93, LOH92, LR93, Liu97, Liu99, LP98, Lov95, MS98a, MW90b, MVMP96, MET94, Moh97, MS99b, OF96, Ozi91b, Pao98, Pen99, PACH90, Plu95, QS99, QA97, RP96, Roo90a, Ruo94, SJAM98, Say97, SH96, SA97, Shu96, ST97]. boundary [Sin92, SIC92, Ste99, Ste95, SS97d, Sun96, Tak92, TB97, TM92, Usm92, VVD94, VVD92, Voe95, WA98, Wri93]. boundary-layer [eMB91]. boundary-value [Anh99, DTB93, Elo98, Gre93, LC99b, Roo90a, ST97, Sin92, TM92, VVD92]. bounded [GH97b, Lan99a, LS99, TG93]. Boundedness [Aco97b, GPRV93, Hai90, LD98]. bounding [LR95]. Bounds [HJ95a, Laf91, LL98a, Mat95, AW95a, AH95, Bal93, Ber99, BGF97, Dei95, Die97, Die98, Die02, EHG90, Gil96, Han91b, Hür97a, Hür97b, Jód90, Jo’90, Jód92, JNC99, LW96b, LU96, MR95b, Mat97, Moe95, Moh95, NYW98, Ozi92, SF91, SL99, Thr99, ZT98, ZLQ96, ma92]. Box [ADT92, Die94a]. Bramble [ZV95]. branch-and-bound [DRW95]. Braunschweig [Eme94c]. breaking [ABM96, LL99b]. Bressoud [Lev90a]. bridge [FSZ94, PP92]. Brjuno [MCM99]. Brownian [PP92]. Bruaset [Mee96]. bubble [Sca95]. Buchholz [AS99]. Building [Eme94b]. Burgers [KBÖ99, OO96]. butterfly [SDL99]. Butzer [JS93]. BVP [Gan93]. BVPs [AW98, SSS96, Smi96].

C [Ano91h, Bul90a, Die94a, Eme94d, PZ98, Ano94l, Eme94f]. C-differentiable [PZ98]. c/o [Ano94l, Eme94f]. cable [CHM94]. CAGD [Ano91f]. Calculating [Fra99, ZL96]. calculation [vdFvGdRdV95]. calculations [CP96, Cha95, Ols99, RL99]. Calculus [Lev91, Ano97b, BC98b, FS96a]. Cambridge [Ano98r, Eme94b]. can [Cuy99]. cannibalism [FSW94]. cannot [Zei92]. Canonical [KM94b, OS94, Lan95]. cantilever [ESEHEK93]. caps [BR98]. capturing [IB99, Yan95]. Carathéodory [Pan94]. Carlitz [Val95]. Carlo [Ano96-29, Ano97-42, DT93, DDG98]. Cartesian [Cha95]. cascade [BM95b]. case [Alv99, AGRZ97, AGRZ98, BDMGVO97b, Cuy99, Dib98, GP97, GRZA97, LW96b, Nak91, Ned94, ND95, Pan97, TG93, VB90b, VZ94]. cases [BY99, KK97]. catalytic [BN92]. catastrophe [SV92a]. category [LM94c]. Cauchy [Ber92b, BK99b, CR97, CL93, Cri97, DS90, DS91, DL96, DPS98, DMGVO99, Die94b, Die95, Die96, Die97, HB92a, JL97, KL98, Leó99, Mac99, Mas99, Müh96, Ors90, San96]. Cauchy-type [Ber92b]. cavitation [BCV98]. Cayley [DLP98, GMBW96]. CB2 [Eme94b]. CEEDE [Ano93l]. Celestijnenland [Ano90k]. cell [JMT94, MR96, Swa96, Swa97, Syd99]. cells [DJR91]. cellular [Mie91, RZ99]. centered [LR92a]. Central [Bla95, GR92a]. centre [LP92]. certain [Aco97c, ÁN98, And90, Arg98, AL97b, BBM96, BGM94, Bav95b, Bav98b, Ben97, CS97b, CH94a, DPS98, De 91a, DMGVO99, Dzi99, Ehr98, GVGPSL93, Grü98, HKL96, IRZ96, Koe93, Köl94a, LP92, Mil97b, MM98, Moo95, Nik96, Rin94, SH96, Sty92, SS97d]. CG [Chr92]. CG-like [Chr92]. CGS [Krz96]. CGSTAB [CF93]. Chain [Ano97m, vD95]. Chain-scattering [Ano97m]. chain-sequence [vD95]. chains [Lin98, Mai93, TG91]. Champaign [Pie96]. chance [Gel95]. change [GR95]. change-of-variance [GR95]. channel [Xia96]. channels [LLM91]. chaos [dMM99]. Chaotic [BE98, Bai97, Ber99, VJ96]. characteristic [Ill98, SSS96, Yan92]. Characteristics [BCV98]. Characterization [Bar99b, Cop94, Han92, PS95b, Dam98, De 91b, KYY97]. characterizations [KKZN97]. characters [CS95b]. charge [Ama94]. Charles [Eme94e]. Charlier [Bav98c, LRV96, Let99, Zen95]. Chawla [Men91]. Chebychev [Ren97b]. Chebyshev [Bav95b, CB92, CB95, CPK96, EK96, Eln97, ES92, HT94, HLCW93, HN98, Kju95, Köl95a, Kui95a, Kui95c, Kui95b, Lap93, Mak92, Mas93, ME93, McC93, Not97, Not98, Pas91, RV94, SV95a, SS97a]. Chebyshev-type [Kui95c, Kui95b, RV94]. chemical [BV99]. Chenies [Ano94k, Ano94m, Eme94a, Eme94e]. Cherenkov [Bab96]. choice [DRW95, Ish99, Lah92, Lah96, SC95a]. choose [BB98b]. Christoffel [Bre90b, Det95, FZ93, Van93a]. Ciarlet [Bul90b]. circle [ACM98, BC98a, BC99, BT92a, BHGVN94, BGVHN97, CM95, CS96a, CS97a, Gra93, Li99a, MG91, Mig95, Pan93b, Pan93c, Pan95a, Pan96a, Pan96b, SL99, VB94]. circuit [DP97, GH97a]. circuits [SD98, SZ97]. Circular [LR92a, eMB91, AK97, AK91, lCHcH97b, De 90a, Dor99b, How93, LR96a, MR94, TF91, WM96a, YD96]. circularity [McC93]. Cl [Köl95a]. Class [BB93, Arg98, Bai98, BEC99, BZ93, Bru94, CNP99, CS97a, CS93b, CM96c, EGBE93, Fit91, FHR98, FM98b, HN90, HZL94, IS93, IS95b, IY95, IY97, JN91, JNC99, LT93, MS98a, McC93, Mil97a, Moh97, Moh98, Moh90, Nie94, QA97, RZQ99, Sch97, Sch94f, Smi91, SV92b, SLK91, SLK92, Sun96, VDV90, YK96, YMW98, Zha92, dO90]. Classes [Dzi99, Ehr95a, Fdi99, Kor94a, KL90, Sab95]. Classical [DGC95, Pry94, ÁNGM95, AGRZ97, AGRZ98, BM98, Ber92b, FKR98, FHR98, FRK99, GMS95, GRZA97, GRZA98b, KR96, KBÖ97, Lew95, Lew96b, Mag95a, Mar93b, RBDM90, RZG95a, RGZA98, SV95a, SS94a, Sla93, TD97, TE97, WD99, ZRG93, Zhe98]. Classification [ZCG99]. Clausen [Köl95a]. Clebsch [ÁNYD98]. Clenshaw [AGR93]. Close [Pon98a]. Close-to-convexity [Pon98a]. Closed [Jód91a, KO99, KO01, SV95b, Ata94, CTV96, Jód91b, SS94a, Wim90]. Closed-form [Jód91a, KO99, KO01, Jód91b, Wim90]. closely [DE97]. clouds [Kle99]. clustering [RTK95]. CMFT97 [Ano97n]. Co [Bul90b, Bul94a, Let95a, Let93, Let99, RBDM90, RZG95a]. co-modified [RBDM90]. Co-recursive [Let95a, Let93, Let99, RZG95a]. coarse [Cha95]. coarse-mesh [Cha95]. coastal [LL99b]. code [Sto99, Wam99, SR99]. codes [HU99, SW95b, Zha94]. coefficient [BK90, LR96b]. Coefficients [CL91, A.94, ABM97, AGRZ97, AGRZ98, ADMFSR98, Bav98b, Ber93a, Bla95, CZ93, Doh98, FHR98, GL99, GRZA97, IS95b, Köl95a, KPP97, KM94b, Lau99, Lew92, Lew96a, Lew96b, Lew98, Lov95, Mag95a, MJ96, MJG96a, MJG96b, Moh98, Odl92, RZG95b, SSS96, SS94b, Tre91, lX97]. Coercive [ATV98, Ned94, ND95]. coherent [MPPP95, MMFMB97, MP99b, Pan97]. collapse [CK91, Kle99, MM99c]. collected [Bul94b]. collection [CRD94]. Collector [BP96a, PKB98]. collision [Bab99, Gör99]. Collocation [Bru94, ES92, Lad97, AP92b, Bar99b, Ber92b, BW96, BY96a, BY96b, CH94a, CPK96, ES97, ES96, Fun92, FP99, Guo95b, HS92a, HS94a, IO91, KNP97, Lev97, Lop90, PACH90, QA97, Ruo94, SK96, SSS96, Shu96].

Collocation-approximation [Lad97]. collocation-type [Guo95b]. Colloquium [Ano91-33]. column [McC99]. combination [GR92c]. Combinational [Ano97o]. combinations [Bav95b, Li97, Li99b, LW99, LLY99]. combinatorial [AZ97]. Combined [CDF95, FFLM95, LB90, Lyn92, Pet90]. combustion [SC95a, Eme94f]. Comet [Mey94]. Comment [Mil06, VVVD99b]. Comments [Anh99, Sez92]. common [GS99, MKK96]. Comonotone [CM96b, Man96]. compact [MMFMB97]. compactly [FI96]. Companion [Ran96, JN92]. comparing [AV96, Atk90]. Comparison [AP98a, BKM95, GBP97, Sab95, ST99, AE90, CF93, CM92, EW99, GWC96, Mel97, MKK96, MdCVR99, Par97]. compensation [Axe91, Lu95]. competing [HLS94]. competitive [Zit94]. compilation [CR93, Coo99]. complementarity [Bai98, BE98, Ben97, MFI95]. complements [Lan99a]. complete [CK99, Köl94a, KS96b, MP93, Sch94e]. completely [PdA99]. completeness [RM95a]. complètes [Bul94b]. completion [FKM97]. Complex [Ano91t, Gil90, Ata94, BPV93, CGV97, GM93, Gre93, Ked93, KSK92, LR92a, LR96a, ME93, McC93, Mei93, Miy92, Miy93, NPB92, PC96, Pet90, Ver96, VIA96, Zha96b]. component [DV99, TH94]. components [CSS93]. composed [SE99]. composite [Car93, FK96, LSSZ90, PTP99]. composition [Bal93]. Compositions [Lor90]. compound [Köh93, Rol99]. compressible [FFLM95, SDP97]. Comput [Alz98, Ano97-31, Die02, DP94b, KO01, Kre11, KR99]. Computable [Han94a, Moh95, CGR99, JN91]. Computation [Ano93a, Ano97a, CDW90, CK91, DEHG92, KM95, LSSZ90, LSZ96, Mac94, Sid97, WMC97, AM97, ACD92, BFG96, BCV98, BY99, Bou90, CLB92, Car96, CK92, CH96c, EK98, Eme94e, EH98, GBP97, Hem96, HS97, LW91, LR96c, Mie91, MZ95, Mor92, Pet99b, PB92, TA95, WGK99, Xia96, Zha96b, Ano91-32]. Computational [Ano90m, Ano91g, Ano91-27, Ano93v, Ano93-27, Ano97n, ABD^{+}96b, FM98b, GY95, HT99, PCC95, RW99, Ush91, VR91, Ano90o, Ano90n, Ano91-28, Ano91-29, Ano92j, Ano92k, Ano92l, Ano94-27, Ano94-28, Ano95z, Ano95-27, Ano96-30, Ano96-31, Ano96-32, Ano97-32, Ano97-43, Ano97-44, Ano97-45, Ano98-32, Ano98-33, Ano99z, ADT92, BT92b, DSH95, Eme94b, FT96, MOT91, NW94, SS99a, TG91]. computations [BKF^{+}96, Bul94c, ER96, VDM96, YM99]. compute [Jon93, VV98]. Computer [Ano90k, Ano91i, Ano91-32, Ano93m, Ano93-30, Ano97-48, Ano97-51, Bul94c, GJ95, Gen90, Van91b, DT93, KNST91, Kor96, Sch97, Sez92]. computerized [Rat94]. Computers [Bul90a, FMT91, Hei95b, KC92]. Computing [Ano94j, Ano97-47, Ano97-51, Ano97-52, CZ93, IKF91, Kob95, LP92, MB93, SS94b, AKM97, Ano91-30, Ano93i, Ano97-42, Ari95, Bec92, BMX97, Bul90a, Ehr94, Hai92, Let96, LMJ93, LJQ95, MKK96, PRS90, PS94, PP96a, PM92, Ren96, Smi96, TM92, VVD92, Wit95]. concave [BS93a]. concentration [BEM97a, BEM97b, BEM98]. concentrically [AK91]. concept [Cuy99, ZYCL97]. Concepts [Ano96-29, Eme94d]. Concerning [Arg97]. Concurrent [Ano93l]. condition [AL97b, Bat98, BCP99, CH94a, CL98, Deh99, MB92, MdCVR99, Nie95, Sen99, Wei96]. Conditional [EH90, LSSZ90]. conditionally [IP99]. conditioned [NS91, Sha93]. conditioning [CS96b, FS91, MP95]. conditions [AM97, Arg98, BS93b, DV99, DD95, Dub96, EH97b, EHS98, FFRV92, GL98, HB92a, KKS96a, Kla94, LOH92, Liu99, LP92, Pao98, RP96, Ste95, SBK98, Sun96, Tak92, Tre92b]. conduction [GL98, HZ95]. conductivity [BP97b, HS92c, XK92]. conductor [Wic90]. cone [Pro94]. Conference [Ano91p, Ano91q, Ano91t, Ano91-34, Ano91-35, Ano93l, Ano93v, Ano93s, Ano93-28, Ano93-30, Ano97-30, Ano97-47, Ano99f, Ano99g, Ano99h, Ano93i, Ano93j, Ano95x, Ano91f, Ano91g, Ano93b, Ano96l, Ano97r]. Conferences [Ano94n, Ano94o, Ano96m, Ano97p, Ano90a, Ano94a]. confluent [AS97, AS99, NPB92]. conformal [Ama94, Cha95, DP93a, DE93, DEP97, FPS99, FZ93, Fra99, HLCW93, How93, PS93a, PS94, RNM97, Weg90b, Weg96]. conforming [KP90, Kar94, Mon94b]. Congress [Ano93a, Ano97a, Ano91w, Ano94y, Ano96-27, Ano91u, Ano97q]. congruences [Rad91]. congruential [DEHG92, EHT90, EHG90, EHN91, EH92a, EH92b, EH94, EH95, EH97a, Emm97b, Weg98]. Conic [KW99]. conjecture [dBS99, Sta97]. Conjugate [ACGW91, CLW95, Heu92, LLP97, BMP96, ESS99, FSV96, Fuk90, HY98, LGT90, SS90, YI96]. connected [Ama94, DE97, Weg96, Mar93b]. Connection [GRZA98a, Aco97c, AGRZ97, AGRZ98, ADMFSR98, Gil99b, GRZA97, Lew96a, Lew98, RZG95b]. connections [SL97]. conquer [EFL97, FB90]. conservation [Leó99, SdO99]. conservative [AT99, DG93, PdA99]. conserving [OP94]. Considerations [Die93]. consistency [WB94]. consistent [AM97]. consistently [HN90]. Constant [PCV99, Ber93a, Ekh91, ESS99, LM90]. constants [Ehr96b, Hwa96]. constrained [AU96, AY94, Eln97, EK98, KT93, Lit92, Sch96a, Wol94, ZP95]. constraint [Moh95]. constraints [BI96, Cri94, FPR92, FT96, RW97, SFI95]. Constructing [HX99, SFW99, ZC99, Ozi92, ST91, ZZ91, Zhe99]. Construction [EH92a, Lab98, BH91, BMP96, BJM97, CL93, GPS92, GM97, JM99, JR93, JL92, Kin90, Lew92, MLD96, VC94, YH95, Zho97, dSS97]. constructions [FM98b]. constructive [Bab98, Faz94, Ise93, NYW98, Sid90]. constructively [Nee94]. contact [BG95, BK95a, BR92b, GK95, Ned94, ND95, Ned97b, Wei96]. contact-two-phase [ND95]. containing [De 91a, KS96a, Lóp98]. contaminant [ABD^{+}96b]. contamination [GR95]. Contiguous [GM95, GIM96, Bus99]. continua [Plu95]. Continuation [VR91, ACGW91, CLW95, Gov94, MM91, VJ96, WM92]. Continued [Hov90, MCM99, AP92a, BCH^{+}99, BF94b, CJT90, Gil96, Gil99a, Gil99b, GM95, Kal95a, Kuc99, Lan99a, Lev90b, Lev94, LB99, Lor92, Lor93, Lor95, MM90, Mat90, Rob99, Sha97a, SG94, Thr90, Thr99, VV95, Waa91, Zen95]. continuity [BEM98, DP95, WB94]. Continuous [AVW99, BJ99, Con99, GRZA97, PT96, Bru94, CFR96, FK96, Gil99a, GRZA98b, GIM96, Ill98, JL97, LMP97, SU99]. continuous-time [CFR96]. continuously [Tho90]. Continuum [Ano97-41]. contour [Ata94]. contractions [Lor90, SA94]. contractive [BP91]. contractivity [dO90]. contrast [BP97b, RTK95, SF92a]. Control [Ano97r, SW95a, APH94, Ano97m, Bel90a, BVZ90, DGH95, Ehr96a, Eln97, EM98b, HLS94, KSFAH98, KKNU97, Lit92, Lit95, Lov95, Mau98, SJAM98, Tho90, TRBS99, VS90, WB94, YY98]. controlled [RE94]. convection [AM95, Bre90a, Epp91, GS97, Lin99, Pen99, Sty95, Swa97, Yan92]. convection-dominated [Yan92]. Convective [CSS93, eMB91, GMdLSG99]. Convergence [AD97, BT96, BGVHN95a, BDMGVO97a, Cha96, CFR96, CRV97, DPS98, DDL93, FD93, Fie90, GG97a, GB90a, GVN90a, HN90, Ham94, HD93, HRGL98, Hyv97, KNO97, Kza95, Lan99a, Leó99, LG95, Pan96a, Sed90, Voe95, Xu99, ZM97, dS97a, Arg91, Arg92a, Arg97, Arg98, AH95, AFK93, Bai97, BGM95, Bar97, Bat98, Bel90a, Ben94a, Ben94b, BR92a, BGVHN97, BL93, BC94, Cap90, Che97b, CS97c, CLM95, Cuy90, Cuy92, CDL96b, Cve99, DS91, Die94b, DP94a, DP94b, ES97, EH91, EH97b, EHS98, Fer97, Gau92, Gra96, GMS91, Gut97, HN95, HNT96, HS94a, HMT90, Hom96, HK94, IY97, Jód90, JK98b, KV95, Le 95, LB99, LL98b, Lin98, Lor93, Lor95, LM93b, MN91, Mas92c, Miy92, Osa96, Osa98, Pan93c, Pan94, Pan96b, PK97, PH98]. convergence [Rob96, Ruo94, Sch94c, Sch94d, SG94, Sid99, Son97, Sta97, SBK98, SV92c, SA94, Thu99, Tok96a, Tse95, Waa91, Wan94, Yam96, Zha92, vdVV93]. convergent [Cop94, Roo90a, Thr90, Tre99]. converges [Sto95]. converging [SdO99]. Convex [LR92b, BC91, EH98, Jüt97, Lan99a, Pro94, Sch94c, Sch94d, TRBS99, Ver94b, Weg93b, WD94]. convex-cone-valued [Pro94]. Convexity [Ano97o, BT92a, CGEP96, GR92b, KW99, Pon98a, Sch94e]. CONVODE [Mou93]. Convolution [VV94, Ber92a, FB96, Kan98, Mal90, Rat94, Rös95b]. convolutions [Str94]. cooling [CSS93]. cooperation [FSW94]. coordinates [HB92b, IG90, MJG98]. Copson [BEL93b]. core [MM99c]. corners [ES96, GL90b]. Cornu [ATBB99]. Corput [Pag92]. corrected [LW91]. correction [Dic91, DLR92, HD93, Paq94, SP96, VVVD99a, VD91]. corrections [BDV94, Bur93, SS91a, YZY96]. corrector [Bur93, Kha94, Sim95, TSM96, TS97, UH98, vdHSdS96]. correlation [SK90]. Correspondence [Ano99i]. corresponding [BCCM91, CS97a, Mor92]. Corrigendum [Alz98, KO01, Kre11, KR99]. coset [Voi95]. cosmic [Cou99]. Cosnard [Roo90b]. Cost [Ano94k, Ano94m, Zha96a]. cost-effective [Zha96a]. Cotes [BDV94, Ehr96a, VDV90]. counting [Kam91]. Coupled [KT93, BJ99, Han91b, Jo’90, Jód91a, Jód91b, JN91, JN92, JNC99, MLD96, ND95, Ned97a, Ned97b, Xia96]. coupling [CZS96, MVMP96]. Coupon [BP96a, PKB98]. Courant [Lev91]. cover [Bul90b]. Cowell [FP97b, Fra95c]. CP [IDV98]. cracks [Alv99]. CRAY [GP93]. credibility [BLG91]. Criteria [Aff90, CS97b, XC96]. criterion [Die96, MR96]. critical [DEHG92, Dib98, HAMeM96, Hol95]. cross [FPR92, SWY98]. cross-entropy [FPR92]. crosscuts [FPS99]. crude [SB94]. CTAC [Ano91g]. Cubature [Sto97, BH91, BP97a, CRD94, CR93, CH94b, Coo99, HX99, HR99]. cube [Spa99, Tic90]. cubic [BWS99, Hig92, Jun92, LP92, MS93, PTP99, PWT99, SU99, WM96a, WM96b, ZM97]. cubics [Man96, MW97]. curl [Mon94b]. curl-conforming [Mon94b]. Current [Wim91]. Curriculum [Ano91h]. curvature [FG91, SU99]. Curve [Böc96, HZ96, McC93, PWT99]. Curved [KK95, Ber94, Han94b, HBL96, RF99b, RF99a]. Curves [Ano91f, AK97, CGEP96, ES96, FK96, MW95, PTP99, PWT99, SS94a, Shu96, SS94c, SS99b, Sya99, WM94, YD96, ZC99]. Curvilinear [FPS99]. cusps [Mor92]. cut [DL94a, DL94b, SWY98]. cutting [KKS96b]. cycle [Axe96, Sha97b]. cycles [ABM96]. Cyclic [BBH90, Osb97, Che98a, HN95, HNT96, HNT98]. cylinder [eMB91, Cha97, GR92b, TF91]. cylindrical [IG90, SLK92, Zha91].

D [Ano91r, Ano93u, Eme94b, Eme94c, CM96d, DPS98, DV99, DM97a, Dor99a, HKW99, Kle99, Lay93, Sak99, VDM96, WF97]. D-3500 [Ano91r]. D-5300 [Ano93u]. D-problems [Dor99a]. D-transport [WF97]. d2lri [HR99]. DAE [BK90, JC95, RH99]. DAEs [CM97, DP97, GR97]. DAEs. [HGC99]. damped [GK96b]. Damping [Ext95]. Darboux [Bre90b]. Dario [Bul94c]. Darrell [Eme94d]. data [ANS96, ADT95, Bab96, CL93, CM96b, DLR92, FI96, Fra99, HMS96, LST96, LR96b, PdST96, RS94]. data-dependent [DLR92]. David [Eme94e, Lev90a]. Dawson [Let98]. death [BV95, Val93, Val96]. decaying [BEM97b, BEM98]. decision [Zel95]. decision-making [Zel95]. Decomposing [VBK91]. Decomposition [Che98a, ATV98, BF93, Bog97, Bog98, CZ95, Dos95, FPS99, FR99, GC97, Hei95b, KP90, Kin90, KV95, LDV91, Lit95, PS94, RVY97, Rod99, Sch94f, TK98, Van90a, vdPBW96]. decomposition-aggregation [Lit95]. decompositions [Yan95]. decoupled [Sty96]. decoupling [Die93, Sin92]. Deducing [Gro91b]. defect [Dic91, HD93]. defect-correction [Dic91]. Deferred [VVVD99a]. deficient [SSY98]. defined [DLR92, Dzi99, Li96, SS99b, Sya99]. definite [Gra93, Nik96, Sha94b, Wen99]. definiteness [Die96]. Definition [BGM90]. deflation [EBP96]. Deflection [AK91]. Deformable [Yán99]. deformation [JSWW95, Ned97b]. degenerate [Del95]. degenerate-kernel [Del95]. degree [BHS^{+}98, Bar99b, FV91, MS98b, Sto97]. Dekker [Ano93k, Bul91]. delay [BMS98, BS98a, BD90, Bru94, BY96b, Cah90, CS97b, CS97c, Dib98, FW99, GS92b, HF95, HJ95b, HHL97, HC99, HLFC99, Liu96, Lu91, LR96c, Mas99, Tho90, WB94]. delay-differential [HHL97, WB94]. delays [CL98, HC99, MT98]. delta [TB99]. deltas [RM95a]. démonstration [Han96]. dense [LR92b, Mag95b]. Density [DW94, BGVHN99, Tag98]. Department [Ano91i, Ano93m, Ano90k]. dependence [Bla95]. dependent [AR92a, Bän91, BB95, De 90b, DLR92, KB91, Köh93, Plu95, Rin94, WB94, XM98]. depending [BMS98, BS98a]. Derivation [CH96a, CH98, Chr98, HM97, BM96c, Sla93, YA95]. derivative [Aco97b, Doh98, ES90, EH98, GH97b, IKK91, Köl94b, Pon98b, TB99]. Derivatives [MA98, Ury94, AW95a, CP93, Ell95, Gil96, Hai92, KO99, KO01, LMF99, MN91, Mas92c, MJ96, TA95, ZZ96, Zha92]. derived [DD95, Rös95b]. Descent [SS97c, SLK91]. description [BCG98, RF99a]. Design [Ano93l, BR92b, But93, KK91, YR99, Ano93-30]. designs [DNN92]. Detailed [RF99a, MET94]. detection [PdST96]. detectors [Bab96]. determinant [Wit96a]. Determinants [BGM93a, Van93a]. Determination [GR92c, OO94, Ata94, BM97, CP93, DF91, HNT98, Lan95, LR96b, RS94, STS91]. determining [Gem96, Pop96]. Development [Ano93z, SK90]. device [FLM92, Hem90a, LM92]. Deville [Eme94c]. Dfl [Ver94b]. Dfl.220.00 [Bul90b]. Dfl.220.00/US [Bul90b]. diagonal [Axe91, Fro91, GP93, Lio96, Lu95, Sta97, VVD92, ZYCL97]. Diagonalization [IRZ96]. diagonally [Cha98b]. diagram [RF99b, RF99a]. diagrams [SFW99]. diamond [Kan98]. dichotomic [Die93]. Dichotomy [vVBZ93, Osb97]. dictionary [SC95a]. Difference [DNS98, MS98b, PP97, ÁN98, Ano97-31, Ano98p, BS95, BDS96, Bav97, Bav98a, BT91, Bul90b, CDF95, DP93b, DM97b, Els99, FM98a, FKR98, FRK99, GRZA98a, Hus94, JV96, Jód91a, JA92, KO99, KO01, KPP97, KL90, KBÖ99, Lay93, LRV96, Lew97b, Li97, LT93, Lin99, Med95, Med96, Med97, Med98b, MJ96, MJG96a, MJG96b, Moh97, Moh98, PdA99, QS99, RGZA98, SV92a, SH96, SA97, SF90, SW97, Sim98, SV92b, Sty92, Sty96, Sun96, SZ98, Tre92a, VVD92, Vol90, Wit96b, WL92, XC96, ZLQ96, ZCG99, dMM99, Ano93k]. differences [Bav95a, De 91b, Deh99]. differencing [AM95]. Different [Ano97-38, Yan95]. differentiability [EHS98]. Differentiable [Ito94, Doh98, PZ98]. Differential [Ano91z, Ano97-46, Ano97-52, Bav98a, BK99a, IS90, JKLL95, JKY97, KLL97, AG98, ADH98b, APH94, AL97a, AR92a, Alw96, AM97, And92, Ano91-33, Ano97-31, ABM97, AZD91, AÁNMP98, AP92b, BS95, BDS96, BKM96, BPP97, BMS98, BS98a, Bav97, Bav98b, BJZ93, Ber93a, BJ99, BT96, But92a, CCJ97, Cas96, CS98, CS97c, CS99, CKP94, CPK96, DZY^{+}94, DL99, Dib98, Die93, DK94, Dzi99, Elb92, Elo98, EK94, EM94b, ELW95, Fit91, FW99, GG97a, GPS92, Gel95, GRZA98a, GPCGV91, GMS94, GS92b, HW99, Han91b, Han92, HF95, HJ95b, HHL97, HC99, HLFC99, IS92, JL92, Jód90, Jo’90, Jód91a, Jód92, JA92, JN92, JNC99, KNST91, Kha94, KKL97, KKL98, KT92, Koe93, Kul92, KM94b, KY97, Lem94, Lim94, Liu96, LN96, Lop90, Lu91, LR96c, Mag95a, Mas99, MLD96, Mau98]. differential [McC93, Mic99, MJ96, Mou93, MS92, MV92, Osb97, Ozi92, PP97, RA97, Ren97a, RS99, RZ99, RVY97, RBDM90, Ron90, RZG95a, Rov92, Rov96, SD98, Sla93, Som93, Tho90, Tra96, Tri93, Val95, Van93c, Waz91, WB94, Wri93, YMW98, ZRG93, ZD94]. Differential-Algebraic [Ano97-46, AP92b, KM94b, Lop90, Ren97a, RS99, SD98]. differential-difference [BS95, BDS96, JA92]. differential-functional [BKM96, PP97]. differential-integral [Dzi99]. differentiation [Bec94, BF94a, Fis91, Ste90b, SS99b]. Diffraction [EEAES99, AHEM97, DS98, Dor99b]. Diffusion [Lac98, MP97, MPS96, AP96, AM95, ATV98, BB98a, Bog98, Bre90a, CNP99, Cha95, Cho92, Deh99, Dub96, FP99, GS97, HMV98, JMT94, Kha94, LD98, LV99, Le 98, Lin99, Pao98, PC96, Pen99, Ram99, Sak99, Sty95, Swa97, Wan97, Yan92]. diffusion-reaction [LV99]. diffusions [LL99a]. digital [Li99b, Uyt96]. digitized [YD96]. Dijk [Bul94b]. dilation [Ano96-34]. dilogarithms [SL97]. dimension [Bou90, Bre90a, Car93, GV96]. dimension-reducing [GV96]. dimensional [Abr91, AR92a, Bän91, BH91, CH94b, De 99, Deh99, EIM91, GM99, Goo91, GS99, Grü92, GB90b, Hem90a, HR99, Hug94, Ira94, IG90, Jia91, KvdB92, KKS96b, KZ92, Kuc99, KBÖ99, LDV91, LM90, Med98a, Moh98, Nak91, NRSS99, OO98, Ozi91a, Pag92, PP92, RIG98, SJAM98, Sim92, TLL91, WS97, Wit95, Wit96a, Wit96c, XZ98, XK92, XM98, YO98, ZL90]. dimensions [BF93, HS94b, Kos94, Sac96]. diminishing [CMR95]. DIMSIMs [JM99]. diode [Hem90a]. dipolar [Mar97, vdFvGdRdV95]. dipole [YO98]. Dirac [RM95a, TB99, TD90]. Direct [Spu99, BS95, BM97, BCP99, Bre90b, CP96, Cha99, Chr98, CDL96b, Ehr95a, Epp92, KP90, Mön96a, OO96, SF92a, hC93, Mes96a]. direction [JW91, LGT90, Sab96]. directions [MAR93a, YZY96]. Dirichlet [AB96, BGH98, BELN92, CL91, DV99, HSW91, Ish91, Klu92, LN95, NV94]. Dirichlet-to-Neumann [BGH98]. DIRK [hC94a]. disc [Weg93b]. discontinuities [CG94, LWZ98]. discontinuity [Faz90]. discontinuous [A.94, CL99, LP98, ZM97]. discovering [VJ96]. discrepancy [EHN91, EH92b, EH94, Fau90, KT97b, Str95]. Discrete [AGRZ97, AGRZ98, Cel98, CRV97, Esp99, Gör99, Sab93, SFGF^{+}99, SWW96, VB94, ÁNGM95, ÁNYD98, ÁN98, AAP98, AVW99, Bar90, BP96b, BEM97a, Bul96c, ES96, Flo96, FKR98, FG93, GMS95, GRZA98b, GS97, Guo95b, HKL96, JDV95, KR98, KR99, Lew96a, Lew98, Mag95b, Mei93, Rat94, RGZA98, Say97, Sta90, Weg93a, YH95, YA95]. Discrete-time [CRV97]. Discretisation [BTCC99, FLM92]. discretisations [Ram99]. discretised [FB96]. Discretization [MTW96, Bab99, Ber93a, CG94, CDF95, Gel95, GR97, Heu92, Joh98, Lep94, Mon93, NKS98, Ned97b, Shi99, Swa96, Swa97]. Discretizations [CS98, Mic99, FD93, Sau91, SR91]. Discretized [Weg90b, CPK96, Wei96]. Discriminating [AD92a]. Disequilibrium [NZ90]. disfocal [Elo98]. disk [BZ93, BG95, Flo96, KN99, LL98b, Tri93]. disks [GH99a]. dispersion [PM92]. dispersive [DPTW95]. displacement [KKZ95, Yan90]. dissipation [BDKM96]. dissipative [Cha98b, HS92c, Tsi98]. dissolution [MP97, MPS96, VV98]. dissolution/precipitation [MP97]. distance [DP93a]. Distances [Mah92, DEHG92]. distortion [MM99a]. Distributed [Ano93s, WF97, BGN^{+}90, CM96d, CFBT99, Drm90, Hac97, LSTZ94, MT98, Roo90b, SL99]. Distribution [EH97a, DT98, EH95, GH99b, Kan97, KKNU97, MG91, RZG95a, Tél98, ZRG93]. distributional [GMS95]. distributions [BMR99, CJT90, DMGVO99, KR98, KR99, Mai93, Par97, Rol99, TG91, TB99]. disturbances [Zin99]. divergence [BGM95, MPSM93, YH95, YA95]. divergent [Wen90]. diverse [And96]. divide [EFL97, FB90]. divide-and-conquer [EFL97]. Division [PP96b]. divisor [MKK96]. DM [Bul94a, Bul94b, Bul96a, Bul96b, Bul96c, Die94a, Kra96, Lev90a, Lev91, Pie96, VR91, Ver96]. doctoral [Ano90k, Ano91r, Ano91s, Ano93u, Ano94w, Ano94x, Ano96z, Ano97-37]. does [Oku93]. Doets [Ano94j]. Domain [KV95, LDV91, ATV98, Ano93-29, Bea99, Bog97, Bog98, DV99, Dos95, FPS99, GN96, GC97, HS92c, Hei95b, HSW91, KP90, PS94, RVY97, Rod99, Sch94f, TK98, Tok96b, Yan95]. domain-oriented [GN96]. domains [Ama94, DM97a, HN90, HN95, HNT96, He94, Kar94, LN95, RF99b, RF99a, Sch94f]. dominance [Fro91, ZYCL97]. dominant [Han92]. dominated [Sac96, Yan92]. Dordrecht [Ano98p]. Dormand [CMR90]. Double [Dub98, AE90, ASW97, CGV97, GL91, OM91, OM99, SX99, VPR97, Voi95, Wit96a, Wit96c]. double-well [Wit96a, Wit96c]. doubling [JV96]. doubly [Ama94]. doubly-connected [Ama94]. draining [TBF99]. drilling [Wer95]. droplet [EBV^{+}99]. dual [AC95, DS98, Fay97, GIM96, RW97, Tri93]. Duality [Bre90c, DVB96, Dos95]. ducts [Abr91, MET93]. Dudenstraße [Kra96]. due [AHEM97, EEAES99]. Duffing [RE94, VJ96]. d’un [Han96]. Durand [Bat98, DF91, Yam96, Yam97]. dusty [HB90]. Dynamic [Was95, DRW95, Yan90, Yan92, Pan95c]. dynamical [LMP97]. dynamics [Eme94b, Gil90, Kor94a, LMSSD96]. Dyson [Rha92].

E-algorithm [LG95]. earth [XK92, vVBZ93, GGF99]. ECC97 [Ano97r]. eccentrically [AK91]. ecological [Kor94a]. ed [Ano94l, Ano94m, Bul94b, Lev91]. edge [LN99]. Edinburgh [Eme94b]. edition [Ano97-40, RW99, Ver96]. Editional [Ano97s]. Editorial [AR92b, Ano92e, Ano92i, Ano94q, Ano94u, Ano99j, Ano99k, Ano99l, BGPW94a, RR94a, Ano90g, Ano90h, Ano90i, Ano90j, Ano91j, Ano91k, Ano91l, Ano91m, Ano91n, Ano92d, Ano92f, Ano92g, Ano92h, Ano93n, Ano93o, Ano93p, Ano93q, Ano93r, Ano94p, Ano94r, Ano94s, Ano94t, Ano94v, Ano95j, Ano95k, Ano95l, Ano95n, Ano95o, Ano95p, Ano95q, Ano95r, Ano95m, Ano96n, Ano96o, Ano96p, Ano96q, Ano96r, Ano96s, Ano96t, Ano96u, Ano96v, Ano96w, Ano96x, Ano97t, Ano97u, Ano97v, Ano97w, Ano97x, Ano97y, Ano97z, Ano97-27, Ano97-28, Ano97-29, Ano98s, Ano98t, Ano98u, Ano98v, Ano98w, Ano98x, Ano98y, Ano98z, Ano98-27, Ano98-28, Ano98-29, Ano98-30, Ano98-31, Ano99m, Ano99n, Ano99o, Ano99p]. Eds [Bul94a, Roo90b, VR91, Bul90a, Bul96a, Bul96b, Ver94b]. educational [AH98]. Effect [MET93, BM95b, BDKM96, Cus94, Ned95a]. effective [And95, Bou90, CH94a, Zha96a]. effects [BC93, Ext91, Ext95, KKZ95, LL99b, Pan95c]. efficiency [Fos97, Fos98, GPMR96, IY95, Pet90, SOMF91, VBK91]. Efficient [Bur93, PC96, Sid91, AB97a, FT96, KP90, KC92, Moo95, Ozi91a, RW97, SS90, Sch97]. Eiermann [RW92]. eigenfunctions [Ano97-31, Bav97, Bav98a, BK99a, VZ94, WA98]. eigenproblem [Sha94b]. eigenproblems [HS97, HL95]. Eigensolutions [Fea91]. eigenstress [Pro95]. eigensystems [TA95]. Eigenvalue [GMdLSG99, AKV92, BS94, DB99, DV99, DS99, EFL97, Lar94, Lau91, Lui98, Sha94a, SSF95, VD91, WA98, ZT98, ma92]. Eigenvalues [And90, Cha98a, AL97b, CT99, Dei95, DM94, Dun98, FP97a, GGF99, Jon93, Sha95, TM92, VVD92, VD94, VZ94, Wit95]. eigenvector [Lau91]. eigenvectors [Dei95]. eighth [Sim99]. eighth-order [Sim99]. Einstein [Fra99]. Eisenstein [AP92a]. Ekkehart [Kra96]. Elastic [GK95, DOAG91, LWZ98]. Elastic-plastic [GK95]. Elasticity [Ano96l, ADG99, AN95, CH96c, Hol95, HSW91, JK98a, Mar97, ND95, YL97]. elasto [CZS96]. elasto-plastic [CZS96]. elastoplastic [BK95b]. elastostatics [Mar96]. electric [AR97, GH97a]. electrical [Pas91]. Electricity [Lis95]. electrochemical [But93, Chr90, Ozi91a]. Electrohydrodynamic [AR97, MEES95]. electrolytic [DJR91]. electromagnetic [HS92c, Hei95b, HKW99, Syd99]. Electronic [Ano93l, CSS93]. electrophysiology [Fun92]. electrostatic [Grü98]. Electrostatics [Mul95]. elegant [BE91]. Element [Ano93b, Ano93j, AP96, AGW91, AKV92, AL98, BAG^{+}96, Bän91, BKF^{+}96, BS98b, Che97a, lCHcH97b, DC95, DV99, DOAG91, DGH95, Eme94d, FFRV92, FFLM95, Fit91, GL90b, GC97, Heu92, Ish91, IO91, JK98a, Joh98, JL99, Jun92, KZ94, LM97, Li99c, LC99a, LM92, Liu97, LJQ95, LN99, MP95, MRT95, Mat98, MVMP96, Mit91, NYW98, Ned94, Ned97b, OF96, Riv91, SSS96, Sau91, Sca94, SFGF^{+}99, SB99, Smi96, SR91, Ste99, Ste95, Sur91, TF91, TB97, VD94, VZ94, Wan94, WS97, Wan97, Wer95, Wu90, XM98, YN95, Yan90, Yan92, YL97, ZZ97, vN93]. element-boundary [Ste99]. element-by-element [WS97]. element-finite [FFLM95]. elementary [FFK97, Jer95]. elements [BHS^{+}98, BGH98, Ber94, But93, lCHcH97a, CM96c, KK95, Mon94b, Sca95, Thr99, ZV95]. elimination [Fos97, Fos98]. ellipse [Weg96]. ellipsoid [Maa94]. Elliptic [Ano91p, A.94, BF93, Bog97, Cha97, Che94, CK92, CS99, CK99, DV99, GBP97, GK97, GN96, Hug94, Ish91, JL92, LW96a, LDV91, LM97, LR96b, LN99, Maa94, MP95, Mat98, MS98a, McC93, Moh97, OF96, RVY97, Sch94f, Smi96, Tre92b, Wei96, YN95, vN93]. Elsevier [Bul90b, Ver94b]. Embedded [HH90, AS96, SS92]. Embedding [TPA95]. Emden [Lim96]. Emmer [Ano94m]. Empirical [IY97, HK94]. EMS’91 [Ano91o]. enclosing [FS95, Hei95a]. enclosure [Plu95]. end [GL90a, GL91, SS91a]. Endpoint [GPRV93, Ver97]. endpoints [SLK91]. Energy [Wit96a, Wit96b, Wit96c, But92b, Ewi96, HT99, Pan95c, Rab94]. enforced [VDM96]. Engineering [Ano93l, Ano97-39, Eme94e, Pas91]. engineers [Ano96y]. England [Ano94l, Eme94e, Eme94b, Eme94f]. ensembles [Hac97]. entire [Ber95, FT98, Le 95, Thr90]. entries [Axe91]. entropy [FPR92, FT96, Mas92a, Mas92b]. entropy-type [FPR92]. ENUMATH [Ano97-30]. enumeration [GR92a]. environment [CFBT99]. environmental [Ewi96]. Environments [Ano93s]. epidemic [GY95]. EPLAN [Lis95]. Epsilon [May95]. Epsilon-inflation [May95]. equation [Abd96a, AP96, AM95, Alw96, AZD91, AÁNMP98, AS96, BGN^{+}90, BS98a, BHS92, BDKM96, BELN92, CMR95, CCJ97, Cha99, CH96c, CL97, DNS98, DSH95, Dem95, DT96, Elb92, ES96, EIM91, Epp92, ES92, ELW95, FKR98, FRK99, GPCGV91, GMS94, GS92b, Guo95a, Ham99, HS94a, HLCW93, HKW99, IY97, Ira94, Ise93, IO91, IDV98, IG90, KNST91, Kan98, Kha96, Kre95b, Kre11, KBÖ99, LRV96, Lew97b, LB90, LOH92, Liu99, LSTZ94, LR96b, Lub93, Mal90, MTW96, MdCVR99, OO94, OO96, PB98, PdA99, Rán92, Reg95, Ren97b, RBDM90, RGZA98, Rov92, Rov96, SX99, Sim90, Sim92, SR92, Sim95, SW97, Sim98, Sim99, SS97a, Slo91, SS91b, SZ98, TSM96, TS97, Tre92a, Tre92b, Val95, Waz91, Wic90, XZ98, Yam96, Yam97, YO98, ZT98, ZRG93, ZCG99, vN93]. Equations [Ano91z, Ano97-38, Ano97-46, Ano97-52, Ned95a, AS99, Abd90, AG98, AE91, ADH98b, APH94, AL97a, AU96, AR92a, ATV98, ÁN98, AM97, AEG98, Ano91-33, Ano93k, Ano97-40, ABM97, Arg92b, AP92b, AN95, Bab98, BOS97, BH98, BEC99, BKM95, BS95, BBM96, BDS96, BKM96, BM96a, BPP97, BMS98, BD90, Bak99, BB98a, Bän91, Bar97, BGM96a, BJZ93, Ber93a, Ber92b, BR92a, Boc95, BW96, BP96b, Bre90a, BT96, BCRV91, Bru94, BY96a, BY96b, BCG98, Bul90b, Buo96, BF94b, But92a, Cah90, CS97b, CMGP99, Cap90, Cas96, CY91, CY94, Che97b, CS98, CS97c, CLL98, Cho92, CS99, CS96b, Chr98, CKP94, CRV90, CFR96, CL98, CPK96, DR90, DBFV96, DZY^{+}94, Del95, Dib98, Dic91, DR94, DR95, Die93, DM97a, DPTW95, DP93b, DM97b, DK92, DK94]. equations [Dub96, ES97, EK96, Elo98, Epp91, EK94, EM94b, FAD99, FM98a, Fie90, Fit91, FLM92, FB96, FW99, FHR98, Fra99, FS95, FT98, GG97a, GPS92, Gel95, GRZA98a, GK96b, GM90b, GS97, GL94, Guo95b, HW99, Han91b, HZL94, Han91d, Han92, HS92b, Hei95a, HD93, Heu92, HF95, HJ95b, HHL97, HLFC99, Hus94, HM97, Ish91, JL92, Jia91, Jód90, Jo’90, JN91, Jód92, JA92, JN92, Joh98, JM96, JKLL95, JKY97, Jun94, JL97, Kan97, KNP97, Kha94, KKL97, KKL98, KT92, Koe93, KPP97, KM99, Krz96, KL90, Kul92, KM94b, KLL97, KY97, Lad97, LD98, Le 98, LC98, Lem94, Li99c, Lim94, Lim96, LR96a, LM92, Liu96, LN96, LC97, Lu91, LR96c, Mag95a, Mar91a, MQ95, Mas99, MLD96, Mau98, McC93, Med95, Med96, Med97, Mel97, Mes96a, Mes96b]. equations [MXZ97, Mic99, MB92, MJ96, MJG96a, MJG96b, MJG98, Moh98, Mon93, Mon98, Mou93, MS92, MV92, NYW98, NS91, Ois95, Osb97, Ozi92, Pao98, PH99, Pop96, RA97, RS95, RZQ99, RP96, RZ99, RVY97, RD96, Rob98, Ron90, RZG95a, Ruo94, Sak99, SS94a, Say97, Sch90, SD98, SY90, SA97, Shu96, SIC92, Som93, SV92b, SDP97, Sto95, Sty96, Sun96, Swa97, Tho90, TK98, Tra96, TLL91, Tsa96, VV94, Ven92, Ven93, WS97, WB94, WL92, Xia96, XC96, XM98, Xu99, YC90, YN95, YMW98, ZD94, ZLQ96, ZZ91, ZW97, ZL90, Bre90a, Ano98p]. Equidistribution [Emm97b]. Equilibrium [HF95, UH98, HH90]. equivalence [Bre90b]. equivalent [HS92a]. equivariant [DM94]. Erdos [LM93b]. Ergodic [GR92d, Drm90]. Erno [Anh99]. Errata [Ano97-31]. Erratum [Ano97-31, Die02]. Error [AC90, AM95, BGF97, Die98, Die02, Dun97, Lub93, SL96, Swa96, Swa97, ZLQ96, AW95a, AP96, AL97a, AE90, AB96, Bal93, Bel90a, Ber99, BM96c, BELN92, CGR99, CH94a, DVVV90, Die95, Die97, Ehr96a, Eir93, Guo95a, Han91b, Han94a, HLFC99, Hug94, Jód90, Jo’90, Jód92, JNC99, Joh98, JL99, Köh93, Köh94, Lim96, MR95b, Mar91b, Mat90, Mat97, McC93, Moh95, Mon98, NYW98, NR90, OF96, PS93a, Rat94, RW98, Seg95, SW95a, Thr99, TB97, Ver94a, Wan97, dS97a, dSS97, SL99]. errors [BJ93, CH98, GM93, GP98, PT96]. especially [SR91]. estimate [AB96, BELN92, VD94, Wan97]. Estimates [För95, AL97b, BP96a, BM96c, Die95, EH90, FG91, Han94a, Isa92, Joh98, Kan94, Köh94, Lub93, Mar91b, Odl92, Ols99, Seg95, SL96, Swa96, Swa97, TB97, Wan94, dS97a]. Estimating [BJ93, DNS98, Miy93, MdCVR99]. Estimation [HHL97, AP96, AL97a, DR90, DVVV90, DF95, Faz92, JL99, MPSM93, NR90, OF96, Osb97, Ver94a, YO98]. Estimations [HC99]. estimator [Hug94]. estimators [CS99, GR95, Par97, Rol99, dSS97]. Euclidean [MMST96]. Euler [Ada97, CQRZ97, Dic91, DR94, DR95, Gil99b, Kju95, OO98, PCV99, RD96, Say98, Sya99, TK98, XM98]. Eulerian [Chr90]. European [Ano91o, Ano91p, Ano97r, Ano97-30]. Euros [Eme94b]. evaluate [MO96b]. Evaluating [LS95, Luc95]. Evaluation [Lew91, Pas92, Ada95, CGR95, DL96, DPS98, Die99, DM97a, EW99, Goo91, Hür97a, Lis95, Mon94a, NPB92, Paq94, San96, Wen97, lX97]. evaluations [Faz90]. event [JDV95]. events [BBH90]. evolution [LP94, MTW96]. evolutionary [DK92, MS92]. Exact [JN91, Val96, Han91b, Hug94, Jo’90, KBÖ99, PM92, ZL90]. exact-explicit [KBÖ99]. example [Gro91b]. examples [GH99b, Le 95, LC99a].

Excel [Ano94l, Eme94f]. excess [DGC95]. exchange [TH94]. excited [ESEHEK93]. exhibiting [FSW94]. Existence [AW98, BPP97, BGM94, CL99, EGBE93, JS97b, Plu95, But92b, CMGP99, GHMZ94, NW94, Ois95, Ozi91b, YC90]. expanded [Ver96]. expanding [HB99]. expansion [AS97, Ber92a, Dun97, GL94, Guo95a, HT94, Hwa96, Jia94, LL99a, ME93, Mat90, MW90a, RZQ99, Rha92, SS94b, Sya99]. Expansions [SRD98, ÁNYD98, DJ98, FS96a, GKH91, GPRV93, Haj95, Lim96, Lóp99, Per92, SL98, SLK91, SLK92, SFB94, Tem95, Tok96b, Wen90, WL92, WZ97]. expectations [Bou90]. expected [HvdV94, Hür97a, Hür97b]. experience [Krí95]. Experiences [Hei95a]. experiment [Atk90]. Experimental [CS97d]. experimentation [Mal90]. Experiments [KNST91]. Explicit [AW95a, CFMR96, Hin92, Jód91b, JA92, Moe95, Odl92, Sim92, Tre91, XZ98, Zho97, CS98, Con99, DLP98, Fra95a, FM98b, Hus94, Kal95a, KBÖ99, LMF99, Mic99, MV92, Sim94, SSV98, vdVNTV97]. Exploiting [JM99]. Exploring [Eme94a]. explosion [Rob98]. exponent [SSS96]. exponential [Dri95, DT98, Dun97, GSSV92, IP99, JS97b, KS96a, LL98a, MB93, OM91, OM99]. exponential-fitting [IP99]. Exponentially [PW92, Sac96, Pfe91, Sim95, Sim99, TSM96, TS97]. Exponentially-improved [PW92]. exponentials [SL96]. expressions [KO99, KO01, Wim90]. Extended [HMT90, Gau92]. extending [Gil99b]. Extension [Ben96, CQRZ97, Cuy92, Krz96, GMS91, Jin96b, Mar95c, Mon94b, Ozi92, Smi91]. Extensions [Pan95a, BEL93b, GH99b, PS99, Sto99]. exterior [Ama94, HSW91]. External [VZ94]. Extracting [HB92a]. extraction [Dub98]. extrapolated [Son99].

Extrapolation

[CRZR96, Guo95b, Mat90, Ano91u, BM96c, JS91, Lim94, Sid91, Sid99, SE99]. Extremal [CG97, DP93a, Njå95, Weg93b, BV95, Leo98a, Peh93]. Extreme [CT99, CI97].

F [Kra96, Lev91, CB92, SS91b]. Faber [DEP97, He94, Kui95b]. factor [Fos97, Fos98]. factorization [Ano96-34, Ban99, BJ99, CS95a, GK97, LOH92, Yun98, vdHS98, Lev90a]. Factorizations [CR97, Hoc97]. factors [FFK97]. fairing [PWT99]. families [ÁN98, RZG95b]. family [AS99, BC99, BCRV91, CB92, CH94b, Gra96, JMMW90, KKZN97, MS98b, SV95a, Sim95, TSM96, TS97, Thu99, Zha98]. farside [WMC97]. Fast [Epp92, Hai92, Jin96b, Lev96, Raf94, Str95, Weg96, lX97, BHS92, CCJ97, Gem96, HB92a, Le 98, MM99b, Wen97, YM99, Kra96]. fault [PdST96]. Favard [BGVHN95b, Kal95a]. FE [CZS96, KKZ95]. fed [KV99]. Feedforward [GR97]. Feintuch [Bul96b]. Fejer [Jia94, JK98b, Sta93, DDR93]. FEM [LLY99, LN95, ND95]. Fer [BJ99]. Fériet [Ext97a]. Fermi [TD90]. Feynman [Kor96]. FFT [HT91]. FFTs [CM96d]. Fibonacci [Ano93-28, NS94]. field [Jer95, VDM96]. fields [AR97, Hei95b, Pro95]. fifth [CMR90, DBFV96, DPTW95]. fifth-order [CMR90, DBFV96, DPTW95]. figs [Die94a, Lev91]. figures [Lan99b, VR91]. filtered [DB99]. filtering [Bes96, Dar98]. filters [JNS90, Pan93a]. financial [Hür97a, Hür97b, Pan96c]. finding [KKZN97, Osa94, PT95, TG91, VV92]. Finite [BAG^{+}96, BT91, Bul90b, lCHcH97b, EBV^{+}99, FFRV92, LT93, LN99, Mat98, MMST96, Ned94, Ned97b, Peh92, SSS96, SH96, Sty95, TF91, TBF99, Wer95, AKV92, AL98, Bän91, BGH98, Ber94, BR92a, BS98b, But93, CDF95, Che97a, CS98, Cri97, CDTV99, DC95, De 91b, DV99, Deh99, DGH95, Ell95, Eme94d, Fas95, FFLM95, FM98a, Fit91, GC97, Heu92, Ish91, Joh98, Jun92, KO99, KO01, KK95, KZ94, KL90, KBÖ99, LM97, Li97, Li99c, LC99a, LW99, LM92, Liu97, LJQ95, LU96, Mai93, MP95, MRT95, MO96b, MVMP96, Mit91, NYW98, OF96, Ors97, PP91, Riv91, RS90, SdO99, Sau91, Sca94, Sca95, SFGF^{+}99, SB99, SW97, Sim98, Smi96, SR91, SLK91, Ste99, Ste95, Sty92, Sty96, Sun96, Sur91, Tre91, TB97, VVD92, VD94, VZ94, Wan94, Wan97]. finite [WF97, Wit96b, Wu90, YN95, Yan90, Yan92, YL97, ZV95, ZZ97, dMM99, vN93]. Finite-difference [BT91, LT93, KL90, SW97, Sty92, VVD92]. Finite-element [FFRV92, Ned94, TF91, Wer95, AKV92, Bän91, Fit91, Heu92, Ish91, LM92, Mit91, Riv91, Sau91, Sca94, SR91, Sur91, VD94, VZ94, Wan94, Wu90, Yan90, Yan92, vN93]. finite-part [BR92a, Cri97, Ell95, MO96b, Ors97]. first [BKM95, BS98a, Cap90, CS96b, Cop94, ES97, FRK99, FN98, Gro91b, KO99, KO01, Köl94a, LW96b, LMF99, LU96, MJG98, RGZA98, Shu96, TD97, Tsa96, VVVD99a, Ano91p, Ano91q]. first-order [BS98a, Cop94, Gro91b, TD97, VVVD99a]. Fisher [KM95]. fitted [Pfe91, Sac96, Sim94, Sim95, Sim99, TSM96, TS97]. Fitting [ANS96, Böc96, HZ96, IP99, Jüt97, LST96, SR92, WD94]. FitzHugh [Ham99]. Fitzroy [Ano94k, Ano94m, Eme94a, Eme94e]. five [AZ97, OS94, VVD92]. five-diagonal [VVD92]. fixed [JS91, Jer91, Sab95]. fixed-point [JS91, Jer91]. flame [EBV^{+}99]. flexing [LM94b]. flexings [Fen94, JM94]. Fliess [Cha98a]. floating [Fen94, Gre93, LM94b]. floating-body [Gre93]. Flow [eM94a, Cha97, eM91, eMB91, ASWO91, ABD^{+}96b, Ari95, BCP99, CP96, De 99, Eme94c, Ezz94, Faz92, FFRV92, FFLM95, HB90, HB92b, Han94b, HAMeM96, HBL96, LLM91, MRT95, OP94, Sac96, Sch99, TF91, TD97, TE97, TOCH99, XM98]. flows [AKM97, CDF95, CFBT99, DLP98, DS99, DGH95, Eme94f, KL99, MR92b, PR93, SDP97, ZZ97, vdVNTV97]. fluid [eMM97a, ADH98b, Ari95, CP96, Eme94b, Eme94c, Ezz94, MRT95, MEES95, Ned95a, SB99, TOCH99]. fluids [AR97]. flux [LR96b, VDM96]. Foata [Han96]. focal [AAP98, HY98]. Fokker [PdA99]. fold [He94]. Forced [ABM96, BBM96, CHM94, JM94, VJ96]. forecasting [And95]. form [BM98, CTV96, ELW95, Jód91a, Jód91b, Kal95a, KO99, KO01, Mar95b, RJ99, Sch94b, Wim90]. formal [HM97, PGGC98]. formalism [Mes96a, Mes96b, Wer95]. format [Paq94]. formation [Cou99, Kle99]. forms [ACD92, BY99, CDW90, KM94b, Lan95, LR92a, RBM91, SV95b, VB93]. formula [Ber93b, Boj96, CZ93, CH96a, CH96b, CH97, EM98b, Gro91c, Koe96, Lew97a, Min98, Müh96, OM91, OM99, Pet99b, RV98, STS91, Say98, SS96b, Voi93]. Formulae [Ano91z, BD90, BH91, BP97a, BC93, CH98, CH94b, Die95, Ext97b, Ext98c, GM97, HX99, HH90, Kza95, MM99b, Nik96, Not90, Not97, Not98, Smi91, Sto97, VDV90]. formulas [AC90, AKV92, BE91, BM96b, BHGVN94, BGVHN97, Die97, För95, GMS94, GS91b, Hin92, JK98a, Köh93, Lau99, Let96, LSSZ90, Ste90b, VPR97]. formulation [ADG99, BCV98, Dic91, Fra95b, HS96, Ros95a, ZZ97]. formulations [GWC96]. Fornberg [DE93]. Fornberg-like [DE93]. Fortran [Eme94e]. Foundation [Ano97-32]. Foundations [Ano94j, RF99b]. Four [IR97, KZ94, ME93, Sab95, Sim90, Sim94, Sim95]. four-node [KZ94]. four-step [Sim90, Sim94, Sim95]. Fourier [AD97, Ano96y, ARS96, AVW99, CZ95, DDR93, FG93, GSSV92, MO96a, MS92, OM99, PB92, Sin94, Slo91, VIA96]. Fourier-type [OM99]. Fourt [ZRG93].

Fourt-order [ZRG93]. Fourth [Ano97-47, FKR98, FRK99, LRV96, MJG98, RBDM90, RZG95a, RGZA98, BS94, Cha98a, CAZ98, IP99, Jon93, Lew97b, Mas93, Not97, PT95, SR92, TM92, Usm92, VVD94, VVVD99b, VVD92]. fourth- [BS94]. Fourth-order [FKR98, LRV96, MJG98, RBDM90, RZG95a, RGZA98, CAZ98, IP99, Jon93, Lew97b, SR92, TM92, Usm92, VVD94, VVVD99b, VVD92]. Fowler [Lim96]. fractal [DE90]. Fractals [Ano91t]. fraction [AP92a, BCH^{+}99, Lor95, MM90]. Fractional [Mac99, ZZ96, AVW99, Bal93, Ell95, Sin94]. fractions [BCCM91, BF94b, CGR95, CJT90, Gil96, Gil99a, Gil99b, GM95, Hov90, Kal95a, Kuc99, Lan99a, Lev90b, Lev94, LB99, Li96, Lor92, Lor93, Mat90, MCM99, Rob99, Sha97a, SG94, Thr90, Thr99, VV95, Waa91, Zen95]. fracture [Ioa90]. fragmentation [Kle99]. Francis [Eme94d]. Fréchet [EH98]. Fréchet-derivative [EH98]. Fredholm [Abd96a, GL94, Guo95a, LN96, Tsa96]. free [eMB91, AKM97, BV99, BCV98, BKF^{+}96, BS98b, CFBT99, Dos95, EH97b, Ezz94, Han94b, HAMeM96, HBL96, LST94, Ozi91b, TBF99]. free-boundary [Ozi91b]. free-convective [eMB91]. free-surface [Han94b, HAMeM96]. French [Bre90a, Han96]. frequency [JTNW93, Pet99a]. Freud [ABM97, Dam98, FHR98, JK98b, LM95, Mat95]. frictionless [BR92b]. Friedr [Eme94c]. frontier [VBK91]. Froude [eM91, eM94a]. Fucik [Esp99, GHMZ94]. Full [LM95, Lio96]. Fully [Deh99, GK96a, Say97, CK92]. Function [Ano97n, Bul91, AS97, Aco97b, Ada95, AE90, Alz97, Alz98, AR99, Ata94, BHS^{+}98, BMS98, BS98a, Boe91, Bus99, CTV96, CQRZ97, CES91, Det95, Doh98, Dri95, DT98, Ext97a, Ext97b, Ext99, FGKP95, FM96, GR95, GLP98, GSSV92, GM90b, Hai92, HvdV94, Hen90b, Jia94, JC98, Köl95a, KS96a, Köl96b, KK97, LW96b, LW91, Lew91, Mav95, MW90a, Mil06, MA98, Mil98, NPB92, NM92, Pal98, PW92, Pre90, Sch94b, Smi96, Tem92, VV94, Van93c, WGK99]. function-valued [GM90b]. function-values [FM96]. functional [BKM96, EK94, JS92, KT92, LSZ96, PP97, PS95b, SC95b]. functional-analytic [JS92]. functionals [CGR99, Die96, DE99, Ehr98, MR95a, MP99a]. Functions [Ano91y, Ano93w, Ano97-53, AS97, Ada95, Ada98, ACM98, All96, ADT95, Arg92a, AAVB98, BZ93, BS93a, Bar99a, Ber95, BP97a, BI96, BGVHN93, BHGVN94, BGVHN95a, BGVO95, BDMGVO97a, BDMGVO97b, BGVHN99, Bus99, BL93, CGR95, CGV97, CZ94, CZ95, CIM94, CADMMB94, Drm90, Dzi99, Ehr95b, ESS99, EM94b, Ext96, FI96, FV96, FKM97, GR92a, GL90a, GP99, GR92b, GVN90b, GMS94, GHH99, GIM96, GH99b, Has97, HRGL98, Hov90, IS90, IKK91, IS92, IKF91, Ill98, JB96, JL97, KKZN97, Kam91, KB91, KW90, Kar91, Köl95b, Köl96a, Kui95c, Kza95, Laf91, LM93a, Lai96, Le 95, Lev96, Lev97, LS99, Li96, Li99a, LR92a, Lóp98, LU96, Loz96, LS95, Luc95, MM99a, Mac94, Mat95, Mel98, MM99b, Mil97c, MM98, Moe95, Moh90, MP93, MS99b, MCM99]. functions [Müh90, Mul93, Mul95, NV94, NS94, NW97, Not90, Odl92, OM91, Osa98, Pan93b, Pan93c, Pan94, Pan95a, Pan95b, Pan96a, Pan96b, Peh95, PH92, Pfe91, PB92, Pon98a, Pon98b, Pro94, Rah96, Rec90, Ren99, Roo90a, Sac96, SL99, SW96, Sch96b, Sid97, Sif98, Sin95, Skr92, Sta93, SLK92, SFB94, Tag98, Thr90, Tok96b, Ury94, Van91a, Van93a, VD94, Wen99, WMC97, YA95, ZZ96, Zei90, Zha92, Zha96b, ZM97, Bul96a]. Fundamental [Bul94c, Chr98, Kit91]. Further [GR92b, Hoc97, Sid99, Tim93, ZL99]. Fuzzy [RTK95].

G [Bul90a, Bul90b, Bul94b, HJ95a]. Galerkin [LLY99, Bar97, BKF^{+}96, Li97, MXZ97, Say97, Shu96, Ven92]. Galloping [CHM94]. gamma [Alz97, Alz98, CZ94, CZ95, CTV96, FV96, GLP98, Lóp98, MM98, Pal98, PW92]. gaps [IL99]. gas [Eme94f, HB90, SR99]. gas-particle [Eme94f]. gaseous [BA98]. gases [AU96, Sch99]. gate [eM91]. gauge [Kal95b]. Gaunt [lX97]. Gauss [Ehr96b, Ext98a, GL90a, GL91, GM97, Hag99, Kza95, Lab98, Let95b, LW95, LMSSD96, Not90, Pet99b, Rad91, Sch96b, Ver97, Zei92]. Gaussian [Die95, Fos97, Fos98, GK96a, GM99, Gra93, GH99b, HN98, Lau99, Pon98a, Sin95]. Gautschi [GLP98]. GB [KS99]. GB-splines [KS99]. GBP [Ano98p, Ano98q, Ano98r]. GCD [NS91]. GCR [dS96a]. Gegenbauer [BHD93, Ext96, FV95, Lew93, MFMBPC97, ME94, VV94, dMPP98]. Gegenbauer-type [ME94]. Geijn [Ano98r]. Gelfand [Kha96]. General [BP96a, DH95, AR92a, ARS96, BGM99, BJM97, Car90, CH96a, Doh98, GR95, RA97, Sak99, Sch94c, Sch94d, SF92b, SBK98, TRBS99, Zha98]. general-order [Doh98]. generalised [ATBB99, CH96b, CH97, DB99, Dun97, Ehr96a, Ext97b, GR92c, GM90b, Lev90b]. Generalization [LMF99, BM96b, EM94b, OSG^{+}93, Pal98, Sal93, SC95b, Val95]. Generalizations [LGR96, GVGPSL93, SF90]. generalize [GHH99]. Generalized [CZ94, DLMT99, FK95, FSV96, Fro91, Ioa90, Lew97a, Pan96c, Sch93, Sif98, Arg92a, Arg92b, Ber93a, BHS92, BH95, BDKM96, CMR95, CZ95, CTV96, Chu97b, Cuy99, Dam99, EFL97, HN90, HT94, IS93, KK97, KL90, LR94, Lev94, Mat97, Mav95, Mil97a, Mil97c, MM98, Mil98, Ned95b, PZ98, RZG95a, RW98, SL97, Say98, SWY98, Sid90, Sid99, SFB94, ST91, WZ97, Yua96, YI96, Zei92]. generate [FG92, Roo94]. generated [EHT90, PP97, QS99]. Generating [Van93c, Ext96, FV91, GR92a, Kem90b, Raf94, Rah96, Sta90, TRBS99]. Generation [ADH98a, Abr91, ATBB99, CW99, DVVV97, Fus90, HT99, Str95]. generator [AS96, EHGNT90, Zie90]. generators [Aff90, Edd90, EHG90, EH92a, Gen90, Nie90, Rip90, Sez92]. Generic [DM94]. Geometric [Ano93-30, MW97, Ver94b, Ber98, GP97]. geometrical [ER96, Sch94c, Sch94d]. geometries [Cha95, Epp92, VIA96]. Geometry [Ano97o, Ver94b, GH99a, Ver94b, Ver96]. Georg [VR91]. Germany [Eme94c, Kra96]. Gibbs [GSSV92, ZM97]. given [FV91, MM99b, Paq94, Pet99b, Ury94]. Glasses [Ano91t]. Global [Ano97-33, LLY99, Mar91b, PP93, Roo90a, Yam96, AL97a, BY96a, Moh90, NR90, Rum95, SFW99, Tre92b, Wol94, Zha92]. GMRES [CF93, EBP96, SG96, Vui95, vG95, vdVV93]. GMRES-like [Vui95]. Gohberg [Bul96a, Bul96b]. good [Joe96, Lan99b]. Gordan [ÁNYD98]. Gordan-type [ÁNYD98]. Gordon [EIM91, KL90]. grade [Ari95, SB99]. gradient [ACGW91, CLW95, FSV96, Fuk90, Heu92, KvdB92, LLP97, NKS98, PP96a, Saa92, SS90, YI96]. gradient-type [Heu92]. gradients [BMP96]. Gradshteyn [BM99]. Graf [Koe96]. Graham [VR91]. Gram [Stu95]. Graphical [FZ96, Sas97]. graphics [Pie96, Sch97]. graphs [FV91, LR92b]. Gravitational [Wam99]. gravity [eM91, eMM97a, Ned95a]. gravity-affected [eM91]. Graz [Ano91s]. greatest [MKK96]. Green [SV92b]. Gregory [BWS99, BDV94, CH97, SS91a]. Gregory- [BDV94]. Gregory-type [BDV94]. GREP [Sid99]. Grid [Seg95, Abr91, ADT95, Cou99, HGL96, QS99]. grids [CP96, Fer97, LL99b, RD96, TRBS99, Yán99, dOP97, vdPBW96]. Groebner [Alw96]. grooved [LLM91]. Grosjean [Ano91h]. Grotefeld [Kra96]. Group [eMB91, Che98a]. groups [GY95, RV98]. growth [CC99, Fos97, Fos98, FSW94, Li96, Rec90]. Gruber [Ver94b]. Grunsky [Gri99]. Gruyter [Bul94a]. guidelines [SK90]. Gunzbuger [Eme94b]. Guy [Eme94e].

H [Eme94e, Eme94c]. Habsieger [Ekh91]. Hadamard [Cri97, Ell95, HT94, LR94, MO96b, Ors97]. Hahn [GIM96, JS93, MP99a, Sab93]. Halanay [Bul96c]. half [BGVO95, BDMGVO97a, BDMGVO97b, OM91, Pon98b]. half-line [BGVO95, BDMGVO97a]. Hall [Uyt96, MR96]. Hamiltonian [AT99, ACD92, BMX97, LM94c, Oku93, Rab94, Sch94b]. Hamiltonians [Die93]. Hammerstein [Buo96, EK96, Guo95b, KNP97]. hand [Jbi99, Lim94, Mus99, SG96]. Handbook [Ano97-34, Ano97-35, Eme94e, Bul90b, Ver94b]. hanging [CHM94]. Hankel [CES91, Fas95, Klu92, KS96a, VB97]. Hard [Bul90b, Mön96a]. Hardback [Ano98q]. Hardbound [Ano98p, Ver94b]. Hardcover [Bul96a, Bul96b, Bul96c, Kra96, Lev90a, Lev91, Pie96, VR91]. Hardy [BZ93]. harm [Oku93]. Harmonic [KN99, Alz97, Alz98, CLL98, DP93a, Dun98, Ked93, LR94, LM90, PS99, Weg93b]. harmonics [Bar90]. having [Ad94, Bav98a, BK99a, KKL98, KLL97, KY97, RA97]. HB [Eme94b]. HC [Bul94b, Die94a]. headlands [EEAES99]. Heat [GL98, Cha99, Epp92, HZ95, IO91, Liu99, LSTZ94, Moo95, MdCVR99, Reg95]. heat-transfer [Moo95]. Heidtmann [BM96b]. Heine [Grü98, GK96c]. Heinrich [Eme94d]. Heisenberg [Grü92]. Hele [MR92b, PR93]. helicopter [SB96]. helioseismology [Kos99]. Helmholtz [MEES95]. hemivariational [Pan95c]. Henrici [Ben96, Sad90, TA95]. hereditary [SWW96]. Hermite [Hag99, AAVB98, BGM93b, BGM94, BGM95, BL93, CLB92, DDL93, DVB96, DJ98, DE97, Dri95, FK95, GM90a, HGL96, Jia94, JK98b, Kha94, KYY97, Man96, MW97, Müh96, Pas90, VMV94, dlCYL98]. Hermite-Fejer [Jia94]. Hermitian [CT99, DL99, Fre92]. Héroult [MR96]. Hessian [FS91]. Heun [AZD91, Val95]. Heverlee [Ano90k]. hexagonal [Cha95]. hidden [BM99]. hierarchic [Sca95]. hierarchical [DOS94, Mit91, Wam99]. High [Alp95, Ano97-36, CPK96, FP90, FP97b, MM97, MJ96, MJG96a, Moh98, Tsi98, eM91, AS96, BP97b, BJM97, Eme94e, FS90, FK95, GM99, JV96, KL99, LJQ95, MJG96b, NRSS99, PP92, RTK95, STS91, Sca95, TF91, VS90, Ver93, XZ98, vdHSvdV95]. high-accuracy [XZ98]. high-dimensional [GM99, PP92]. High-order [Alp95, FP90, MM97, Tsi98, AS96, FS90, JV96, KL99, STS91, Sca95, VS90, Ver93, vdHSvdV95]. Higher [PH92, AW98, AP92b, BK90, BS94, CW99, EM94b, Ext97b, KO99, KO01, TE97, VMV94, WA98, YN95]. higher-index [AP92b]. Higher-order [PH92, AW98, AP92b, BS94, EM94b, TE97, WA98]. Highly [Vec96, DP97, EW99, HS92b]. Hilb [Voi93]. Hilbert [Bea99, BC94, Die99, HT94, RS90, Weg90a, Weg93a, ZV95]. Hill [Wit96a]. hip [BK95a]. historic [DNN92]. hole [Ad94]. Holland [Bul90b, Roo90b, Ver94b]. Höllig [Die94a]. holonomic [Zei90]. homoclinic [ABM96]. homoclinics [DW94]. Homogeneous [Nea96, CDL96a, DK92, WM92]. homotopies [GWC96]. homotopy [HL95, VC94]. Hoog [CRV90]. Hopf [ASW97, LST94, LR96c, Pel94]. Hord [Ano94l]. House [Ano94k, Ano94m, Eme94a, Eme94e]. HPOPT97 [Ano97-36]. Hubbell [GKH91, Let95b, Tim93]. hulls [LR92b]. human [BK95a]. Hurwitz [LR94, MA98]. Hybrid [AH98, JS97a, LS96, BP97b, CR98, DRW95, Fra95a, KKZ95, LB90, MRT95, NM92, SF91, SG96, TS97, XZ98]. hybrid-combined [LB90]. hybrid-displacement [KKZ95]. hydrodynamical [IB99]. hydrodynamics [Dor99a, RL99]. hydromagnetic [Ezz94]. hyper [DM97a]. hyperbolas [Weg98]. Hyperbolic [GH99a, BKM95, BBM96, BKM96, CRZR96, Faz90, HD93, Kan97, Leó99, MB92, MJG96a, MJG96b, MJG98, SdO99]. hypercube [DG92]. Hypergeometric [Orl99, Pon98b, AS97, AS99, ÁNYD98, ADMFSR98, Bar99a, Bus99, DDL93, Ext97b, Ext98a, Ext98b, Ext99, GS91b, GIM96, GK96c, JC98, Kar91, KK97, Mav95, Mil06, Mil97a, Mil97c, NPB92, NV94, Pon98a, SRD98, SL97, SFB94, VPR97, Wen90, ZD94]. hypergeometric-type [ADMFSR98, ZD94]. hypergroup [Flo96]. hypergroups [CS95b]. hyperplane [FMT91]. hypersingular [ES92, Kre95b, Kre11, Mon94a]. hypersphere [Sto97]. hypersurfaces [PP96a]. Hyperterminants [Daa96, Daa98].

Ideal [HBL96, HR94]. Identification [ADT97, Lap93, ADT96, Ano98q, Böc96, FW99]. identify [Rol99]. Identities [Koo98, Ada95, Cal97, DLMT99, Ext98a, Lev94, Rad91, Zei90]. identity [All96, Bre90b, Ekh91]. if [BF94a]. if-problem [BF94a]. IFIP [Ano93s]. ignition [BA98, EBV^{+}99]. Igor [Ver96]. II [And96, Lev91, BR98, BDMGVO97a, CLW95, Daa98, DL94b, DR95, Hür97b, IRZ96, Li99c, LN95, Mas92b, Mes96b, Not90, RF99a, Sch94d, TE97, Lev91]. IIA [dS97b]. IID [Sch98]. III [Ano97-34, BDMGVO97b, GIM96, LC99a].

IIIA [PRT97]. Il’in [Roo94]. Ill [Sha93, EH91, NS91]. Ill-conditioned [Sha93, NS91]. ill-posed [EH91]. illposed [Dar98]. illustrations [Kra96]. illustrative [Gro91b]. ILU [CS97d, Sau91, vdPBW96]. ILU-decomposition [vdPBW96]. IMACS [Ano93a, Ano97a, Bul90a, Ano93t]. Image [Ano91f, LB98, Li99b]. imaging [AEG98]. imbedded [CH94b, Tho90]. imbedding [HS92c]. impact [Pan95c, VV95]. imperfect [Wei96]. Implementation [LYFW95, Tra96, AB97a, AB97b, CV93, Gen90, LD99, RW98, Sez92, Sid91, ST99, Kra96]. implementations [Bah95]. Implicit [Dor99a, AB97a, Arg92a, Bru94, BC93, CMGP99, CV93, CRV90, DVVV97, Deh99, EIM91, GPMR96, JB99, JW91, MS92, RZ99, Ste90b, VVVD99a, VVVD97, VM99, WF97, dSS97]. implicitly [DB99, SS99b, Sya99]. implicitness [MS98b]. Improved [FFH97, Thr99, DR94, Ish99, PW92, Sto95, vN93]. Improvement [Bat98]. Improving [GPMR96, Osa98]. impulse [WRW92]. impulsive [BKM95, BS95, BDS96, BKM96, BM96a, FN98, Kul92, LD98, PP97]. impulsively [GK96b]. inaccurately [MM99b]. including [IS93]. inclusion [AH95, Ois95]. inclusions [LP94]. Incomplete [CGR95, CZ94, CZ95, CTV96, Dup94, MM98, SLK92, Tem92, Yun98]. incompressible [BCP99, CF93, DGH95, Eme94c, LLM91, Ned95a, TLL91, ZZ97, Eme94b]. incremental [BK95b, YR99]. indefinite [CR98, CS97d, LLP97, Mas93]. independence [Nie90]. independent [KV95]. Indeterminate [Ber95, Ber98]. Index [Ano90f, Ano91c, Ano92b, Ano96j, Ano98d, Ano99q, Ano99r, Ano99s, Ano99t, Ano99u, Ano99v, Ano99w, Ano99x, Ano99y, Yak97, Ano90b, Ano90c, Ano90d, Ano90e, Ano90l, Ano91a, Ano91b, Ano91d, Ano91e, Ano92a, Ano92c, Ano93d, Ano93e, Ano93f, Ano93g, Ano93h, Ano94c, Ano94d, Ano94e, Ano94f, Ano94g, Ano94h, Ano94i, Ano94z, Ano95a, Ano95b, Ano95c, Ano95d, Ano95e, Ano95f, Ano95g, Ano95h, Ano95i, Ano95y, Ano96a, Ano96b, Ano96c, Ano96d, Ano96e, Ano96f, Ano96g, Ano96h, Ano96i, Ano96k, Ano96-28, Ano97c, Ano97d, Ano97e, Ano97f, Ano97g, Ano97h, Ano97i, Ano97j, Ano97k, Ano97l, Ano98e, Ano98f, Ano98a, Ano98g, Ano98h, Ano98i, Ano98j, Ano98k, Ano98l, Ano98m, Ano98n, Ano98o, Ano98c, Ano98b, Ano98-34]. index [Ano98-35, Ano99a, Ano99b, Ano99c, Ano99d, Ano99e, AP92b, BK90, CNS97, Ked93, LMM94, Lop90, Ren97a, RS99, YM99]. index- [RS99]. index-2 [Ren97a]. index-reduction [LMM94]. indicators [Mon98, PP96b]. indices [FZ96]. indirect [hC93]. induced [PCC95, Ste90a, dMM99]. Inductance [Kar91]. induction [SK90]. inefficiencies [Ste90a]. Inequalities [För93, JB96, AP98a, BC91, EM98a, Erd93, ESS99, Fuk90, GLP98, IS90, Jhu95, Mas95, NZ90, NNR93, PP97, Pal98, Pan95c, Ury94, Ano93k]. inequality [Alz97, Alz98, BEL93b, Grü92, IS92, JMT94, Tse95]. Inexact [FFM95, KM99, MQ95, PZ98, Arg97, LYFW95]. Inexact-Newton [KM99]. inference [Zel95]. infima [BC98b]. infinite [Ad94, Bou90, CGV97, Cha99, GK96a, JKY97, Kla94, Köl95b, Köl96a, KS96b, LW96b, LYFW95, Liu96, LS95, Luc95, OM91, Ree94, Sid97, Sid99, Tre99]. infinitely [Doh98, FT96, Gro91b, MR96]. inflation [May95]. Influence [Han94b]. information [Art94, Dup94, KM94a, KM95, MM99b, MPSM93]. inhomogeneous [CM92, Ked93, Mac94]. Initial [Kju95, AM97, AFK93, CFMR96, CS93a, DF91, Fra95a, FG97, Fra99, Gro91b, KT97a, Lah96, MdCVR99, SK96, SF92b, SS92, Sim92, Sim94, SDS94, Spi96, Voe95, VM99, hC94a, hC94b, vdHSvdV95]. initial-value [AFK93, CS93a, Fra95a, FG97, Gro91b, KT97a, SF92b, SS92, Sim92, Sim94, SDS94, Spi96, hC94a, hC94b]. initial/boundary [AM97]. injectively [FG99]. injectivity [AZ97]. inner [Bav95a, BP96b, De 93, FKM97, Lov95, Wit95]. input [KM94a]. insights [Vui95]. inspired [Lor95]. Instability [MS98a, AR97, MEES95]. Institut [Ano91s, Ano91r, Ano93u]. Institute [Ano91y]. Institutional [Ano90k, Ano91r, Ano91s, Ano93u, Ano94w, Ano94x, Ano96z, Ano97-37]. integer [HvdV94, TPA95]. integrability [AAVB98, Tre92b]. Integrable [Shi99, NKS98]. Integral [DM97a, FAD99, GMS94, Mat97, Abd90, Abd96a, AE91, ADG99, Alp95, Ano97-40, BGM90, BGM93a, BGM93b, Bar97, Ber92a, Ber92b, BR92a, BM99, BW96, BCRV91, Bru94, BY96a, BY96b, Cah90, CS97b, Cap90, CS96b, Chr98, CH96c, CRV90, CFR96, Del95, DS98, Dun97, Dzi99, ES97, EK96, ES96, ES92, Ext91, Ext95, FB96, Fra95b, GKH91, Goo91, GM90b, Guo95a, HLCW93, HKW99, IRZ96, JM96, Jun94, JL97, KP99, Kha94, KS96a, Köl96a, Kre95b, Kre11, Let95b, Let98, LL98b, Mal90, MR95a, Mar95b, Mas93, MW90a, MVMP96, MXZ97, Pen99, PB98, RZQ99, Ruo94, Say97, Shu96, Tim93, VV94, Ven92, Ven93, Wer95, Yak97, YR99, ZL90]. Integrals [CK99, Ada95, AE90, AGW91, BK99b, Cri97, DS90, DS91, DL96, DPS98, Die94b, Die95, Die97, Die98, Die02, Ell95, EW99, FAD99, GK96a, GM99, HB92a, Ioa90, KL98, Köl94b, Köl95b, KS96b, Kor96, Lew91, LSSZ90, LSZ96, Lóp99, LS95, Luc95, Maa94, MO96b, Mav95, ME94, Mon94a, OM99, Orl99, Ors90, Ors97, Par92, PP92, San96, Sch98, Sid97, Sid99, SS97b, SLK91, SFB94, TD90, Tem95, Ury94]. integrand [Ver97]. integrated [ElM94, SZ97, Sin94]. Integrating [SE99, Die93, Lev97, Li99b]. Integration [Ano97-53, DP97, NS94, AR92a, Bar97, Ber93b, BR92a, Car93, CZ93, CW99, EM98b, HT91, Has97, HM97, JVZ97, Joe90, JM96, KN90, Lau96, Lev96, Lev99, Lóp99, Mas93, MS99b, MV92, NR90, NM92, NRSS99, Owr95, Pag92, Pag98, Rab90, RS92, RA96, Rán92, RZ99, SL98, Sha93, Sim92, Sim94, SDS94, Sim95, SS96b, SLK91, TSM96, TS97, dDGZ99]. integrations [SW95a]. integrator [PCV99]. integrators [Oku93]. integro [ADH98b, CCJ97, Kha94, LN96, MV92]. integro-differential [ADH98b, CCJ97, Kha94, LN96, MV92]. integrodifferential [CRV90, Jód91b, Lad97]. intensity [MB93]. interacting [GY95]. interaction [Kre95a]. interactions [Ewi96, RL99]. Interactive [KKS96b, SU99, GS99]. interconnect [BOS97]. interest [Ask93]. interesting [RV94]. interface [CZS96]. interfaces [KV99, Sas97]. interferometry [Els99]. interior [Ama94, BGH98, Eme94a, Fay97]. interior-point [Eme94a, Fay97]. Interlacing [PS95a]. Intermediate [Köh94]. Internal [Ham99, eMM97a]. International [Ano91q, Ano91u, Ano91-32, Ano93b, Ano93j, Ano93l, Ano93-28, Ano93-30, Ano94l, Ano97-39, Ano97-41, Ano97-47, Ano97-51, Bul90a, Eme94f, Ano91t, Ano93v, Ano93w, Ano98q]. interpolant [CMR90, CMR99, Loe96, McC98]. interpolants [Chu97a, Chu97b, Sab96, Ver93, ZT99]. interpolating [DP94a, DP94b, Gil99a, MR95b, San96]. Interpolation [ES90, Pfe91, Rab96, AC90, Aco97c, AW95a, ADT92, BHS^{+}98, BWS99, BL97, BM92, Ber97, BM97, CH96a, CH96b, CH97, CH98, Col98, CM96b, Cri94, DNZ98, DVVV90, DVV90, DLR92, FI96, FS96b, GM90a, Gau92, HGL96, HMS96, HS94b, Hig92, Jia94, JK98b, KS97, Lah90, Lah92, Lah96, LL91, LL98b, LM93b, Man96, MN91, Mas92c, MO96a, MW97, Mes96a, Mes96b, Müh90, Müh96, Pan94, Pan95b, SU99, VB90a, VB90b, VDV90, VVD92, VMV94, Ven96, Vol90, XGSSKF96, YR99, ZV95, Zha95, ZM97, Zhu97]. Interpolatory [Not97, RS92, BI96, Dem95, FK96, GK96a, Kob95, KvD99, Not98, NRSS99, SV92c]. interpretation [Grü98, Mus99, Sha97a]. interval [Ben94b, ES92, Hai90, Has97, Mat95, Moh90, OM91, Sch94f, SFI95, Wol94]. intervals [Peh93, Van93a, WA98]. Intl [Van91b]. Introduction [Ano97-38, BMV95, Kir92, Ano96y, EHK^{+}95, Mar95a, Van91b, VR91, Lev91]. invariant [BH91, BMX97, HS92c]. Inverse [Ano97-39, CKM97, GM93, LM92, Tag98, ADT96, ADT97, AD92b, BM98, Cha99, Chu97b, CM92, DS99, DDG98, EH97b, Fea91, GM90b, HS92c, Isa92, KNN96, Ked93, KZ92, Lau91, MBP93, Mat90, Müh96, MdCVR99, OO94, SÁFV96, Sha94a, SIC92, ZL96, ZL99]. inverse-free [EH97b]. inverses [Mic92, PRS92]. Inversion [Kos99, VV96, AD97, Ehr95a, GS91a, Iqb95, Iqb97, Tem92, XK92]. inversions [EHT90]. inversive [EH92a, Weg98]. Invertibly [Tre99]. investigation [ASWO91, Chr98]. investigations [AL97a]. invited [Ano91v, Ano95s]. involved [CH98]. involving [ÁNQR99, Bav95a, BM95b, CK99, Ext98c, FFH97, Li92, LS95, Luc95, Mav95, MW90a, PP96b, Sid97]. ion [TH94]. ion-exchange [TH94]. Ionescu [Bul96c]. IRK [Lab97, RH99]. irrational [Hov90]. irrationality [Pré96]. irregular [Hin92]. irregularly [BBH90]. ISBN [Ano93k, Ano94j, Ano94k, Ano94l, Ano94m, Ano98p, Ano98q, Ano98r, Bul90b, Bul94a, Bul94b, Bul96a, Bul96b, Bul96c, Die94a, Eme94a, Eme94d, Eme94e, Eme94b, Eme94c, Eme94f, Kra96, Lev90a, Lev91, Mee96, Pie96, Roo90b, Uyt96, VR91, Ver94b, Ver96]. ISIP’98 [Ano97-39]. island [AHEM97]. isometric [BGF97, Gra93]. isoparametric [ZV95]. Isospectral [DS99, DLP98, LMP97]. isotherm [Woo91]. isotropic [JSWW95]. iterated [BY96a, Cha98b, GlZ94, KNP97, Som93]. iterates [GG97a, GM93]. iteration [AH95, Axe91, AL98, Bat98, CR96, CP93, Dub98, Dub97, Fea91, Hyv97, KKZN97, Osa98, PH99, VS90, Zhe93, hC94b, vdHSvdV95]. iterations [BH98, BEC99, Fre92, JS91, YN95]. Iterative [Bog97, CLR97, EIM91, GK97, LC99b, Lit92, RK90, Weg93a, Arg98, BE98, BKM96, BBD^{+}96, Bec94, Bru96, BC94, CR98, Chr92, CNS97, DDG98, Fie90, Fro91, Gem96, GPMR96, GC97, GH97b, HS92a, JB99, Jbi99, Jin96b, KNO96, KNO97, Mee96, MdSvdV97, Miy92, Nou93, PH92, STS91, Sch90, Sch94c, Sch94d, Son99, Sta93, SBK98, Tse95, WS97, Weg90a, Weg90b, Wei96, YK96, Zha92, Mes96b]. IV [Ano97-35, GR92a]. IVP [Rán92]. IVPs [PRT97, VVVD99a]. ix [Bul91].

J [Alz98, Ano97-31, Ano98p, Ano98q, Bul90b, Bul94b, Die02, KO01, Kre11, KR99, VR91, Ver94b]. J. [DP94b]. Jacobi [AÁNMP98, DR94, DR95, HS94a, Hen90a, Hen90b, Kui95c, Kza95, Let95a, Lew92, Lew93, MP99b, MBMF99, Pan97, Rös95b, WMC97]. Jan [Van91b]. January [Eme94e]. Jarratt [ACQ94, EH97b]. Jarratt-type [EH97b]. JCAM [RW99]. jet [Ezz94]. John [Eme94d, Lev91]. joint [BK95a]. Jones [Pie96]. Jordan [Fay97, SS94c]. Jr [Eme94e]. Juan [Eme94d]. July [Ano91-28]. jump [LL99a]. justification [HZ95].

K-queues [HJ95a]. Kac [CL97, Kor96]. Kadomtsev [FM98a]. Kakutani [Pag92]. Kampé [Ext97a]. Kantorovich [CM95, Lar94, Zhe93]. Kantorovitch [Abd98]. Kárman [Boc95]. Kees [Ano94j]. Kehe [Bul91]. Kelvin [AHEM97, MEES95]. Kent [Ano94l, Eme94f, Van91b]. Kergin [LL98b]. kernel [Abd96a, Alp98, Del95, Epp91, Fre92, Kan98, PS93a, RNM97, RS90, VV94]. kernel-based [Epp91]. Kernels [CS96a, Alp95, ARS96, BGVHN95b, DDR93]. Kerner [Bat98, DF91, Yam96, Yam97]. Kershaw [GLP98]. key [HB92a]. kind [Abd90, Abd96a, AE91, BCRV91, BY96a, Cap90, CS96b, Die95, ES97, GK96c, Köl94a, MO96a, MXZ97, Not97, Peh95, RZQ99, Shu96, Tsa96, Van91a, WMC97]. kinds [Mas93, ME93]. Kinetic [Bab99, Bab98]. kinetics [BV99]. Klein [Bea99]. Kluwer [Ano98p]. knapsack [DRW95, RW97]. knowing [JN92]. known [FG92]. Koekoeks [ELW95]. Koelbel [Eme94e]. Koenig [Sid90]. Kolmogorov [Nee94]. Kontorovich [Ehr95a]. Kopernikusgasse [Ano91s]. Korteweg [BDKM96, DPTW95, FM98a]. Krall [GHH99, Zhe99]. Krawtchouk [CSE95, IS98]. Kron [Lui98]. Kronecker [FFH97]. Kronrod [Ehr96b, Not90]. Krylov [SL96, dS96a]. Kumar [Fra95b]. Kuramoto [SZ98]. Kuramoto-Tsuzuki [SZ98]. Kutta [CB95, DVVV97, OS94, PT96, SDS94, Tsi98, VVVD97, hC93, vdHM97, BJZ93, Bru94, BC93, BJ93, CMR90, CFMR96, CMGP99, Cha98b, CS97c, CV93, CRV90, CRvdHV93, CRV97, DVVV97, GPMR96, HH90, HGC99, JB99, LMP97, MdSvdV97, Owr95, PRT97, SS92, Som93, ST91, Tho90, Tra96, VVVD99a, VS90, Vec96, Ver93, VM99, dSS97, hC94b, vdHSvdV95]. Kutta-type [JB99].

L [Bul90a, Bul90b, Bul94a, Bul96a, Eme94e, VR91]. labelled [FV91]. Lacunary [KS97, Ven96]. Ladyzhenskaya [DGH95]. lag [Liu96, Sim92, XZ98]. lagged [EH97a]. Lagrange [Cri94, Gau92, LM93b, MN91, Mas92c, MO96a, Müh96]. Lagrangean [RW97]. Laguerre [Bel93a, BG98, Che98a, CIM94, ELW95, Ext98c, FHR98, GS91a, GHH99, Haj95, KS96a, Let93, LS98, MPP96, MMPP97, MP99a, Sif98, SS96a, WK98, Zen95]. laminar [eMB91, Eme94c]. lamination [VDM96]. Lanczos [Bah95, Chr94, DK92, DK94, KC92, Krz96, Mus99]. Lanczos-type [Mus99]. Landau [Jhu95, LS99, PdA99]. Landau-type [Jhu95]. Language [Ano93c]. Laplace [Iqb95, Iqb97, LB90, SS91a, Tag98]. Laplacian [Esp99, GMdLSG99, Med98a]. Large [dDGZ99, eM94a, BFG96, BEC99, BK95b, Bou90, Cou99, DNN92, DHP94, FFH97, HS97, HL95, KL99, KKS96b, LDV91, Lit95, MS99a, NPB92, Rum95, SV95a, SSF95, Zha96b, Zin99]. Large-scale [dDGZ99, BFG96, Cou99, KKS96b, Lit95]. larger [Mat95]. Larkfield [Ano94l, Eme94f]. lasers [PCC95]. lateral [Els99]. Lattice [KN90, Lan99b, EHGNT90, Joe90, Kal95b, Lan95, Lew98, Mag95b, NS94, RJ99]. launcher [De 99]. Laurent [BHJ99, CJT90, Hag99, Hen90a, Hen90b, JN99, Zhe98]. Laurent-polynomials [CJT90]. Law [Bul90a, Ezz94]. laws [Leó99, Lin98, SdO99]. layer [eMB91, Ham99]. layered [Pro95]. layers [Bab99, MET94]. Least [BP96b, BM93b, YL97, FFH97, JS97b, Mon96b, PP96a, SF91, SSY98, Sci94, VB94, Van90b, Yua96, YI96]. least-gradient [PP96a]. Least-squares [BP96b, BM93b, YL97, Mon96b, SSY98, Sci94, VB94]. Lebedev [Ehr95a]. Lebesgue [BC98a, Hwa96, Mat95, SC95b]. leg [HLFC99]. Legendre [EK98, Fun92, Let95b, LW95, LMSSD96, MO96a, OSG^{+}93, Pet99b, Sch93, Ver97]. Legendre-based [EK98]. Leja [CR96]. lemma [ZV95]. lemniscates [Yán99]. length [EHT90, EH92a, Faz92, PP96b]. lensing [Wam99]. Less [Bus99]. Letter [Han91c, New91]. Leuven [Ano90k]. level [GP93, KV99, LW96a, Lit95, MJG98, Ols99, ST99, Zha98]. level-set [KV99]. levels [Wit96a, Wit96b, Wit96c]. Levin [Bru96]. Levinson [JNS90, Pan93a]. Levitan [Kha96]. Lévy [Sch98]. Liapunov [NL94]. libraries [Sas97]. Lidstone [AW95a]. lie [FM98b, Cas96, HV91]. like [ANS96, CY94, Chr92, DE93, KV95, LS96, Mey94, Ned97a, Ned97b, OO94, SOMF91, VD91, Vui95]. Limit [AC95, FRR94, Voi95, Bal93, Bla95, CK91, GR92a, Gil90, GRZA98b, LM94c, SG94, Thr99]. limit-periodic [Bal93]. limits [RZAG98]. Lindelöf [Hyv97]. line [Boe91, BGVO95, BDMGVO97a, BDMGVO97b, Die99, Ekh91, JS92, Sin95, SS97b, YR99]. linéaires [Bre90a]. Linear [Ano97-31, Ano97-40, Ano97-49, Bav97, Ben94a, Fay97, LR93, LMJ93, AG98, And92, Ano98r, BE98, Bal93, BBD^{+}96, BK90, Bav95b, Bav98a, Bav98b, Ben97, BP91, Bre97, Bre98, BY96a, Bul94a, Bul96c, BJM97, CSS93, CR98, Cas96, Cel98, Cho92, CNS97, DR90, DHP94, Del95, DOAG91, Die93, Die96, DLR92, Ehr98, EHG90, Eme94a, EH91, FT96, Fis91, FM98b, Fuk90, GG97a, GPS92, GR92c, GL94, Han92, Hug94, IO91, JL92, Jbi99, JN92, KL98, KPP97, Kor94b, KM94b, Li92, LYFW95, Lin99, Liu97, MFI95, MR95a, MP99a, Mas99, MLD96, Mes96a, Mes96b, MdSvdV97, MJG98, Mus99, PACH90, Rha92, RBM91, Rov92, Sch90, Sen99, ST97, Sin92, SS94c, Son99, SV92b, SR91, TPA95, Tre92a, Tse95, TM92, Ush91, Was95, WL92, YR99, ZL96]. linear [ZLQ96, ZL99, vdHM97]. Linearization [ADMFSR98, ÁNQR99, Han94a, Lew96b]. linearized [SIC92, VB94]. linearly [Eln97, LDV91, RZ99]. lines [AR92a, KNST91]. linesearch [SS97c]. Lions [Bul90b]. Liouville [Cha98a, DW94, MP92, Pry94, SV92b, VD91, VD94]. Lipschitz [BDS96, Jia94]. liquid [Kre95a]. liquid-shell [Kre95a]. liquified [AU96]. List [Ano91v, Ano91w, Ano91x, Ano93x, Ano93y, Ano94y, Ano95s, Ano95t, Ano95u, Ano95v, Ano95w, Ano95x, Ano96-27]. little [Sza93]. Lixing [Eme94f]. loaded [AK91, BLG91, CHM94]. loads [GJ95]. Lobatto [GL90a, GL91, Min98, PRT97, Sch96b]. Lobatto-type [Min98]. Local [Ano97-33, OF96, PS93a, Riv91, BJ93, CMR99, DL96, DV99, Deh99, GR92a, Kla94, Kob95, KKZ95, Kos94, SW95a, VIA96, Yan95]. Localization [FS96b]. localizing [Sza95]. Locally [RT91, SS91b]. locating [De 93]. Location [Leo98b]. locations [OO94]. locking [BS98b, Sca95]. locking-free [BS98b]. locus [FW99]. Löf [Van91b]. log [CJT90]. log-normal [CJT90]. logarithmic [Osa90, Osa96, Pré94, Sab95, Sed90]. logarithms [Köl95b, KS96b]. logic [Ano94j]. Logistic [Arn92]. Logistics [Ano94l, Eme94f]. London [Ano94k, Ano94m, Ano98q, Eme94a, Eme94d, Eme94e]. long [MR96, Slo91]. Longman [Mee96]. Look [VB97, Hoc97]. Look-ahead [VB97, Hoc97]. Lotka [HLS94]. Lotka-type [HLS94]. Loveman [Eme94e]. Low [Syd99, Bar99b, SS92, SDP97, Str95]. low-discrepancy [Str95]. Low-noise [Syd99]. low-order [SS92]. low-speed [SDP97]. Lower [AH95, ma92, CW99, LP98]. Löwner [VV96]. LP [CK91]. Ltd [Ano94l, Eme94d, Eme94f]. lumped [AKV92]. Lyapunov [BS95].

M [Eme94b, Eme94c, GR95, Kri92, Lev90a, Roo90b, Van91b, Ver94b, Yun98, HJ95a, Kri92]. M-estimators [GR95]. M-matrix [Yun98]. M.I.T [Ano94k, Ano94m, Eme94a, Eme94e]. M/G/1/K [HJ95a]. M/G/1/K-queues [HJ95a]. M/M/1 [Kri92]. Macdonald [RV98]. machine [GB90a, Kra96]. machining [But93, Chr90, Ozi91a]. Maclaurin [Say98, Sya99]. macro [CM96c]. macro-elements [CM96c]. magnetic [VDM96]. magnetodynamic [Ned95b]. magnetodynamics [Ned95a]. magnetohydrodynamics [Sto99]. magnitude [CDM91, Zha96b]. magnitudes [NPB92]. Magnus [Mee96]. major [FZ96]. majorizer [HN95]. Making [Vol90, Zel95]. Mangler [Ioa90]. Mangler-type [Ioa90]. manifolds [Ver96]. Mannheim [Kra96]. many [FT96, Gro91b]. MAOR [Cve99, Son97]. map [CLL98]. mapping [Ama94, Cha95, DP93a, DE93, DEP97, FPS99, FZ93, HLCW93, How93, PS93a, RNM97, Weg90b, Weg96]. mappings [BGH98, Weg93b]. maps [Jer91]. Marcel [Ano93k, Bul91, Kan98]. March [Ano93-31]. Marcinkiewicz [EM98a]. Markov [KW90, Lin98, Mai93, TG91]. Martin [Van91b]. Martin-Löf [Van91b]. Mary [Eme94e]. Mass [HR94, OP94, AKV92, CG97, GH99b]. Mass- [OP94]. Mass-transshipment [HR94]. masses [OO94]. massive [BW96]. Massively [Ano93s, Ano93z]. Master [Ano90l, Ano94z, Ano95y, Ano96-28]. matching [BOS97]. Matchings [Pal92]. material [KV99, VV95]. material-void [KV99]. materials [Mar96]. Math [Alz98, Ano97-31, Die02, DP94b, KO01, Kre11, KR99]. Mathematica [Pie96]. Mathematical [Ano91q, Ano93-30, Ano96l, Ano98q, CZ90, Die94a, MPS96, NS95, Car96, Ell94, LM90]. Mathematics [Ano90m, Ano91-27, Ano91-34, Ano91-35, Ano93-27, Ano97-30, Bul90a, Bul91, VR91, Ano90o, Ano90n, Ano91-28, Ano91-29, Ano92j, Ano92k, Ano92l, Ano94-27, Ano94-28, Ano95z, Ano95-27, Ano96-30, Ano96-31, Ano96-32, Ano97-32, Ano97-43, Ano97-44, Ano97-45, Ano98-32, Ano98-33, Ano99z, BB95, Lev90a, Mee96, Ano93a, Ano97a, Ano94m]. Mathematik [Ano91r, Ano91s, Ano93u]. Mathieu [SX99]. Matrices [BM97, Ste97, And90, AL97b, AL98, Cal97, CR97, CT99, CS97d, DM94, DS99, Fas95, FP97a, Gra93, HN90, HS92a, Hai90, KNO96, KNO97, LR94, Lio96, MP95, MS99a, Mel98, Mic92, Njå98, Sha95, Sha94b, Sha93, Tre99, Wri93, ma92]. Matrix [BF93, De 90b, Lu95, Mig95, Axe91, BFG96, BL97, Ben96, BMX97, Bul94c, CGR95, CGR99, Chu97a, DJ98, Dei95, DDG98, Fre92, FKM97, FM98b, GMR94, HvdV94, Hoc97, Jód92, JN92, JC98, KM94a, KM95, Lau91, Le 93, MB93, MKK96, Müh96, Mus99, Nie90, Par91, PGGC98, Ros95a, Sin95, SV96, SL96, VV96, Yun98, ZT99, Bul96a]. matrix-based [MKK96]. Matrix-dependent [De 90b]. matrix-valued [Chu97a, FKM97, ZT99]. maxima [BPP97]. maximal [CG97, EH92a, Lan99b]. Maximum [FN98]. Maxwell [Mon93, Mon98, RS95, SIC92]. mbH [Eme94c]. me [Ask93]. ME20 [Ano94l, Eme94f]. Mean [HK94, JK98b, LM93b, MN91, Alz97, Alz98, Gau92, HJ95a, SV92c]. meaningfulness [FR95a]. means [FR95a, Ito94]. measure [DP93a, Mas92a, Mas92b, SC95b, Tsa96]. measure-theoretic [Mas92a, Mas92b]. measured [BM95b]. measurements [Els99]. measures [BV95, Ber98, CM95, DP95, HK94, Kui95a, MMFMB97, Pan93b, Pet99a, VS93, dlCYL98]. measuring [Drm90]. mechanical [Wer95]. Mechanics [Ano97-39, Ano97-41, Ioa90, MOT91, Ros95a, SR91, Eme94c]. mechanisms [Ext95]. media [CC99, HB90, HB92b, KKS96a, Ozi91b, WRW92, Yan90]. medial [RF99b, RF99a]. medium [CM92, CKM97, Ked93, MP97, MET93, MPS96]. Meijer [SFB94]. Meixner [Bav98c, CSE95, CI97, LRV96, Let99, Lew97b]. Mellin [GS91a, Mil97c, OSG^{+}93, Par92]. melting [GB90b]. membrane [BR98]. memory [BT95, CM96d, CFBT99, KNST91, MTW96]. merit [Lan99b]. meromorphic [Li96, Rec90]. Mesh [SR91, Cha95, Kle99, Kos94, Lin99, Sab96, Sty92, Ver94a]. mesh-refinement [Ver94a]. meshes [CLM95, De 91b, ES90, EBV^{+}99, KS97, Riv91]. meshsize [BHS^{+}98]. method [A.94, eMB91, AKM97, AP96, AR92a, Ama94, And96, Anh99, ACQ94, Axe91, Axe96, AL98, BAG^{+}96, Bah95, BOS97, BS95, Ban99, Bar97, BV99, BCV98, Bec92, BKF^{+}96, Ben96, BMX97, Ber92b, BP97b, BS98b, Bre90a, Bru96, BC91, CR98, CMR90, Cap90, Cha96, CAZ98, CQW95, Che97a, Chr90, Chr94, CNS97, Cve99, CPK96, DE93, DC95, DV99, Dic91, Die99, DK92, DK94, Dub98, ES97, ES96, EFL97, Eme94d, Epp91, Ern96, EH98, EHS98, FPS99, FFH97, FM98a, Fra95a, FB90, FG93, Fun92, GW91, Gan93, Gem96, GV96, GL90b, Gre93, GC97, GL94, Guo95a, Guo95b, GB90b, Gut97, HN90, Han91a, Han91b, Hei95a, Hem90a, HMT90, How90, Ish91, Ish99, IO91, IP99, Jer91, JW91, Jo’90, Jon93, KNST91, KC92, KL99, Kit91, Kju95]. method [KvdB92, KV99, KNO96, KNO97, Lar94, LC98, Lem94, Lev97, Li92, LB98, LW99, LOH92, Lit92, Liu96, Lóp99, LMF99, LN99, Lui98, LD99, Lyn92, MFI95, MP95, Mak92, MP92, Mar91b, Men91, Mie98, Mon93, Mor92, MV92, Mus99, NKS98, NL94, Nou93, NRSS99, Ols99, Osa94, Pag98, PRS90, PS93a, PS94, PP96a, Pen99, PK97, PH98, QA97, RE94, RZQ99, RVY97, RD96, RW98, Ruo94, Saa94, SF91, SK96, SS94a, Sau91, Sca94, Sch90, Sch94f, Sha95, Shu96, SS92, SF90, SFW99, SG96, Sim90, SR92, Sim94, SDS94, SW97, Sim98, Sim99, Sin92, SS97a, Smi96, SC95a, Son97, Sta93, Ste95, SOMF91, SFI95, ST91, Sur91, Swa96, Swa97, Sza95, TLL91, VVVD97, VVVD99b, VDM96, VVD92, Ven92, VV98, Vol90, VIA96, WS97, Weg90a]. method [Wei96, Wic90, Woo91, Wu90, XM98, Yam96, Yam97, YO98, Yan92, YZY96, YI96, Zha92, ZZ94, Zhe99, Zit94, hC94a, vN93, BGM96a, BGM96b]. méthode [Bre90a]. methodologies [MKK96]. Methodology [Bre98, TD97]. Methods [Ano93v, Ano93-30, Ano97n, LR95, LJQ95, VR91, AT99, AH98, ACGW91, AGW91, AM96, AB97a, AB97b, AV96, Ano93-29, Ano97-42, ABD^{+}96b, Arg91, Arg92a, Arg92b, Arg97, AH95, AS96, Bai97, Bai98, BE98, BKM96, Bar99b, BJZ93, Ber99, BGM99, BT92a, BT91, BP96b, Bre97, BT96, BCRV91, Bru94, BY96b, Bul90b, Bur93, BJ93, BJM97, BC94, CLR97, CMR93, CFMR96, CMGP99, CP93, Cas96, CRZR96, Cha98b, CY91, Che94, CY94, Che97b, CS98, Che98b, lCHcH97b, CS97c, CL91, CLW95, CW99, CH94a, CS96b, Chr92, CB92, CB95, Col98, CM92, Con99, CV93, Cou99, CRV90, CRvdHV93, CFR96, CRV97, DEP97, DVVV97, Deh99, DL99, Del95, DOAG91, DR94, DR95, DLP98, Die98, Die02, DTB93, DPTW95, Eme94c, EH91, EW99, Faz94, Fen94, FD93, FS90]. methods [FK95, Fit91, FZ93, FFM95, FS91, FG91, FM94, FM96, FM97, FP90, FP97b, FGR97, FG97, FS96b, Fra99, Fro91, FS95, FR99, GY95, Gea91, GBP97, GPMR96, GN96, GH97a, GH97b, HS92a, HW99, Hai92, HZL94, Hei95b, HGC99, HT99, HJ95b, HKW99, HLFC99, HM97, IY95, IY97, IR97, IDV98, JVZ97, JSWW95, JV96, JB99, JS91, Jbi99, Jek91, JC95, JM96, Jun94, JL97, Kal95b, KNP97, KP90, KNN96, KT97a, Kor94b, Kos99, KM99, Krz96, KL90, KBÖ99, Lab97, Lab98, Lad97, LLP97, LDV91, LM97, Lep94, LB99, LB90, Li97, Li99b, Li99c, LC99a, Lim96, LC99b, LLM91, Liu95, LMSSD96, Lop90, LMP97, Lu91, Mal90, MOT91, Mar91a, MQ95, MVMP96, Mel97, MR92a, Mes96a, Mes96b, MdSvdV97, MXZ97, Mic99, Miy92, MJG96b, Moh97]. methods [MS92, MdCVR99, Nie94, OS94, Osa91, OP94, Owr95, OO96, PT96, Pet90, PH92, PT95, PH99, PRT97, PZ98, RA96, RA97, Ree94, RK90, Riv91, Rod99, RH99, SP96, SSY98, Say97, Sch94c, Sch94d, SV92a, SH96, Sid91, Sim92, Sim95, ST99, SS97c, Som93, Son99, SDP97, Ste97, Ste90a, Sty95, ST91, TSM96, TS97, Tho90, TH94, Tra96, Tsa96, Tse95, Tsi98, TM92, UH98, VB97, VVVD99a, VS90, VV92, Vec96, Ver93, Vui95, Weg93a, Weg90b, Wer95, Xia96, XZ98, Xu99, Yan90, YD96, YL97, YK96, Yua96, ZZ97, ZFV98, dO90, dS96a, dSS97, hC93, hC94b, vdHSvdV95, vdHSdS96, vdHM97, vdHM99, Mee96]. metric [CM95, HK94]. metrics [Bea99, HR94]. Micchelli [Boj96]. Michael [Ano94l]. Michele [Ano94m]. microelectronic [SD98]. micropolar [Mar96]. midpoint [EHS98, GVGPSL93, LW91]. midterm [Lis95]. migration [Woo91]. Mild [Tre92b, EHS98]. Miles [Ver96]. Milos [Mar95a]. MIMD [Ano94l, Hei95b]. mind [Ano94m]. Mindlin [lCHcH97b, Sca95]. minima [Sch96a]. Minimal [AGRZ97, AGRZ98, BV90, GRZA97, BP97a, HS97, LLP97, Lan99b, RW98, Sid91, Sim92, XZ98, YA95, Zha98]. minimality [PGGC98]. Minimax [LW95, Aco97b, AP98a, EH90, ME93]. minimization [BT95, Saa92]. minimizers [Kla94]. minimizing [PP96a]. Minimum [FPR92, Rol99, But92b, Det95, Lyn92]. minus [BP96b]. Mirror [Ano94l, Eme94f]. miscible [Yan90]. Mises [HB92b]. Mixed [Col98, LM97, MRT95, Yan90, BGH98, BS93b, BJ99, CH96a, CH96b, CH97, Che94, lCHcH97a, CL91, DVVV90, DVV90, HB92a, JNC99, LOH92, MVMP96, MJ96, Par97, Sca94, TPA95, VDV90, VVD92, Yan92, vN93]. Mixed-hybrid [MRT95]. mixed-type [VDV90, VVD92]. mixture [Par97]. MODCOM [Ano97-41]. MODCOM-14 [Ano97-41]. Model [FR95b, Hol95, BK95a, BM95a, BM95b, BN92, CSS93, CL97, DGC95, DGH95, FSZ94, FSW94, Fun92, GWC96, Gör99, Gro91a, GM95, HB90, Hem96, How90, JS97b, Lis95, LL99b, McC98, Saa92, Saa94, SB99, SGFR94, Ste90b, SBK98, ZZ97]. model-trust [McC98]. modeling [Eme94f, Ewi96, PCC95]. Modelling [Ano97a, KKZ95, Pan95c, Vrb95, AU96, LV99, MPS96, Sch99, TBF99]. models [AD92a, BA98, Cus94, DNS98, KM94a, KM95, Kor94a, Lac98, Mig95, Pan96c, RS95, SV95b, Syd99, Ano97-41]. Modern [AFK93]. modification [Mak92, Riv91, VVVD97]. modifications [ÁNGM95]. Modified [ÁNYD98, CH96b, HS92a, HU99, AS97, BB98b, BDV94, BGVHN95a, CZ93, CH97, Dem95, Die97, HT91, Joe96, KvdB92, Köl95b, Laf91, LD99, Lyn92, RBDM90, SS94a, SS97b, Skr92, Stu95, SA94, VDV90]. modular [Owr95]. modules [FM98b, PS94]. modulus [EHGNT90, EH94]. molecular [Kle99, LMSSD96]. Moment [BOS97, BGVHN93, AD92b, Ber95, Ber98, BHJ99, CM96a, CG97, Fas95, JN99, Kal95a, Njå95, Njå98, RS90, TG93]. Moments [NV94, BB98b, CZ93, Doh98]. momentum [OP94]. momentum-conserving [OP94]. Monge [Abd98]. Mono [VM99, DVVV97, VVVD99a, VVVD97]. Mono-implicit [VM99, DVVV97, VVVD99a, VVVD97]. Monographs [Ano98q, Van91b, Bul91]. Monomial [CR93, Coo99]. monopolar [vdFvGdRdV95]. Monotone [BKM96, BM96b, CG94, CADMMB94, HS94b, RT91, Rov96]. Monotonic [Hig92, LM90]. Monotonicity [KvD99, dOP97, GR92b, LM93a, Moo95]. Monte [Ano96-29, Ano97-42, DT93, DDG98]. Montessus [Cuy90, Cuy92, Sid90]. Morchoisne [Eme94c]. morph [FG99]. morphisms [LR94]. motion [TOCH99]. motions [VJ96]. Moulton [CW99]. movement [DM94]. moving [ADH98a, Chr90, KV99, MR92a, Yan95]. MP [GP93]. MSOR [HS92a, HMT90]. MSPT [ESEHEK93]. MTL [BOS97].

Multi [DR94, DR95, ER96, FM94, Zha98, Ano93-29, BTCC99, DV99, FM96, FM97, LW96a, SJAM98, ST99, WM92, vdVNTV97]. multi- [WM92]. multi-component [DV99]. multi-domain [Ano93-29]. Multi-level [Zha98, LW96a, ST99]. Multi-phase [ER96]. multi-physics [BTCC99]. multi-span [SJAM98]. Multi-staging [DR94, DR95]. Multi-step [FM94, FM96, FM97]. multi-time-stepping [vdVNTV97]. multibody [SGFR94]. multicolour [Gje97, Wu90]. Multiconference [Ano91o]. Multidimensional [MW90b, MJG98, SB94]. Multidisciplinary [Ewi96]. Multidomain [VIA96]. Multigrid [Kal95b, MZ95, SDP97, De 90b, Dic91, DR94, DR95, Hem90a, Hem90b, IG90, Ram99, RD96, Riv91, Sha97b, Zha96a, Zha98]. Multilevel [GS97, LLM91, Axe91, AL98, GN96, Wan94]. multiphase [ABD^{+}96b]. Multiple [Apt98, LM94c, VV92, DRW95, GL90a, GK96a, HC99, IY95, Jbi99, Joe90, KM94a, LSZ96, Miy92, Mus99, Osa94, SG96, Sto95, VZ94, Waz91]. multiple-choice [DRW95]. multiple-input [KM94a]. Multiplicative [Pop96, DEHG92]. Multiplicity [Miy93, Dub97]. multipoint [BGVHN97, GVN90b]. multipoles [Cuy92]. multiresolution [BH95]. multiscale [HS96]. multisplitting [Bai97, BH98, BEC99, Cha96, MFI95]. multistage [JVZ97]. Multistep [FI96, Ver93, FS90, FK95, FP90, MdSvdV97, NR90, dO90]. multitape [Kra96]. Multivariate [Chu97b, LST96, AC90, AB96, Cuy90, Cuy92, CDL96b, Cuy99, FS96b, GR92a, Gui97, LMJ93, Moe95, Müh90, Zho97]. Müntz [MDR98, Zho94]. Murat [Uyt96]. Muskhelishvili [CH96c]. mutual [MPSM93]. mutually [Gro91b].

Naghdi [BS98b]. Nagumo [Ham99]. narrow [ZV95]. NATO [Ano91z, Ano91y]. natural [Gil99a, Sun96]. Navier [AN95, Bän91, CF93, HM97, Ols99, SDP97, TLL91, WS97]. Near [ME93, McC93, ASW97, Has97, Sch94b]. Near-circularity [McC93]. Near-minimax [ME93]. nearby [SLK91]. nearly [AD92a]. nearside [WMC97]. nearside-farside [WMC97]. necessary [CL98, MB92]. Nédélec [Mon93, Mon94b]. needs [Loz96]. negative [Ada98, SS94a]. neighbourhood [GPS92]. Nested [Gui97, dS96a, vdPBW96, Pat99, RW97]. nets [GR97]. network [KK91, PP91]. networks [Ano94k, Ell94, GN96, Ito94, NL94, Ano91t]. Neumann [BGH98, BCP99, CL91]. Neural [Ano91t, Ell94, GR97, Ito94, NL94, Ano94k]. neurocontrol [FR95b]. neutral [BS95, BBM96, BPP97, GS92b, HHL97, HC99, KT92, MB92]. Neutrino [MM99c]. neutron [Cha95, DZY^{+}94]. neutron-transport [DZY^{+}94]. Nevanlinna [BV95, Njå98]. Newsletter [Ano90m, Ano90o, Ano90n, Ano91-27, Ano91-28, Ano91-29, Ano92j, Ano92k, Ano92l, Ano93-27, Ano94-27, Ano94-28, Ano95z, Ano95-27, Ano96-30, Ano96-31, Ano96-32, Ano97-43, Ano97-44, Ano97-45, Ano98-32, Ano98-33, Ano99z]. Newton [Arg92a, Arg97, Bai97, BDV94, CY94, CQW95, Che97b, DP94a, DP94b, Dub98, Ehr96a, EH98, FFM95, FS91, FG91, FM94, FM96, FM97, GW91, Gut97, Ish99, Jer91, Jun94, KM99, Lyn92, Mar91a, MQ95, PZ98, Saa94, VDV90, Weg90b, Xu99, ZDR90, ZZ94, Zhe93]. Newton-like [CY94]. Newton-type [Bai97]. NGILU [vdPBW96]. Ni [PCC95]. Nicolaides [Eme94b]. Nigrin [Ano94k]. Nikishin [BL93]. Nikishin-type [BL93]. nodal [DPS98]. node [KZ94]. nodes [Col98, FP99, SL99]. noise [Bes96, GP97, GP99, SD98, Syd99]. noisy [DDR93, Iqb97]. Non [BSW99, AW98, BM96b, Bre90a, CT99, DV99, Deh99, Fre92, LC97, Zel95]. non-Bayesian [Zel95]. non-Hermitian [CT99, Fre92]. non-local [DV99, Deh99]. non-monotone [BM96b]. non-normality [LC97]. non-positive [AW98]. Non-uniqueness [BSW99]. nonadaptive [HR99]. noncommutative [Flo96, Rad91]. nonconforming [Joh98, Rod99]. nonconvex [Pan95c]. nondefinite [Ehr98]. Nondiagonal [Dri95, MBMF99]. nondifferentiable [CY94]. Nonexistence [Kor94b, YC90]. nongridded [CM96b]. nonhomogeneous [CDL96b, DM97b, GRZA98a, JL92]. Noninverse [SK90]. nonisothermal [eMB91]. Nonlinear [eMM97a, Ano91-34, BR98, CS97c, Chr92, Hei95b, JM94, KK91, LM94b, Ned95b, ZZ97, Zin99, ADH98b, AU96, ACD92, Bab98, Bai97, BH98, Bai98, BEC99, BGM96a, BGM96b, BK95b, BM95b, Bre90a, BY96b, Bru96, BJM97, CNP99, CY91, Che94, CY94, CS98, CHM94, Cho92, CK92, CS99, DKKV98, DTB93, EHGNT90, EGBE93, EK96, EK98, Epp91, Fen94, FSZ94, FB96, Fro91, FS95, FT98, Gan93, GHMZ94, GPCGV91, GS92b, GK96b, Gro91a, Guo95a, HZL94, HS92b, Hem90a, HLFC99, Ish91, JV96, Jek91, Jia91, JM96, Jun94, KNN96, Krz96, Lad97, LP94, Lap93, LM97, LP98, Mar91a, Med96, Mic99, MJ96, MJG96a, MJG96b, MJG98, Nak91, NW94, Ois95, Plu95, Pop96, PdA99, QA97, RIG98, RW97, Rob98, Ruo94, Sab95, SF91, Sci94, SY90]. nonlinear [SA97, SGFR94, SS96b, TB97, Weg90a, Weg93a, Wei97, Xu99, YC90, YMW98, ZZ94, ZCG99, ZYCL97, ZZ91]. Nonlinearity [Gab92]. nonlocal [LT93, Liu99, Pao98, Sun96]. nonnegative [MS98a]. Nonoscillatory [Med98a, BPP97, BMS98, Sak99]. nonoverlapping [RVY97]. nonperfect [CLB92]. nonperiodic [GSSV92, NS94]. nonpolynomial [VVD94, Wit96b]. nonquadratic [Saa94]. nonrectangular [Fer97, Sch94f]. nonregular [Lim94]. Nonselfadjoint [Bul96b]. nonsimple [Mor92]. nonsmooth [Abd90, Che97b, MQ95, Pan95c, Tak92]. nonstandard [Mic99, dMM99]. nonstationary [AD92a]. nonsteady [Ozi91b]. nonstiff [CS93a, hC94b, vdHSvdV95]. nonstrictly [Leó99]. nonsymmetric [ARS96, DKKV98, Jód92, KC92, Li92, Mus99, Sha95, SG96, Wit96c, vdFvGdRdV95]. nonterminating [Koo98]. nonuniform [CLM95, Mag95b, SE99, WRW92, dOP97]. nonunique [Sch94c, Sch94d]. nonvanishing [LS99]. nonzero [Tok96a]. Nordström [Van91b]. norm [Gri99, HS97, VM94]. normal [AR97, ACD92, BY99, CDW90, CJT90, Par97, Rin94, Rol99, Sch94b]. normality [LC97]. normalized [FFK97]. Normals [SDL99]. North [Bul90b, Roo90b, Ver94b]. North-Holland [Bul90b, Roo90b, Ver94b]. nose [ASWO91]. Note [Sha97b, Wen97, hC93, AB97a, Bar97, Cve99, DP94a, DP94b, Ehr96b, Fra95b, Gil96, Gil99b, GVGPSL93, Hai90, Han91a, IG90, KM96, Kan94, LB99, Lor93, Ors97, PK97, PRS92, Rab94, RS99, Woo91, Zhe93, ZT99]. Notes [GG97b, Han91b, Mee96]. notions [Pas90]. novel [Lin99]. November [Ano91-29, Ano92l]. nr [Ano91-28, Ano91-29, Ano92l, Ano93-31]. nth [GS92b]. nth-order [GS92b]. nuclear [HT99]. nucleosynthesis [HT99]. null [YH95]. number [eM91, eM94a, Aff90, BB93, Bur93, CH94a, DEHG92, Edd90, EHT90, EH92a, Fus90, Gen90, GB90a, Joe96, KL99, PP96b, Rip90, Sez92, TF91, Zie90]. number-machine [GB90a]. number-theoretic [Joe96].

Numbers [Ano93-28, Ada97, All97, AL97b, CGV97, EHN91, EH92b, EH94, EH95, EH97a, Emm97b, Köl94a, Nie94, Rad95, Sen99, Weg98, Zen95]. NUMDIFF [Ano97-46]. NUMDIFF-8 [Ano97-46]. Numerical [AL97a, Alv99, AM97, Ano91q, Ano93w, Ano96l, Ano97-30, Ano97-46, ASWO91, BK95a, BR92a, CC99, CH96c, DL96, DEP97, DL99, DKKV98, DTB93, DPTW95, Dub96, Faz90, Faz92, Faz94, GS91a, Has97, HB99, HLCW93, How93, IB99, JSWW95, Jer91, Jer95, JC95, Jun94, KNN96, Krí95, KBÖ99, Le 98, Lim96, Liu99, Lu91, LR96c, LL99b, Mal90, MKK96, Mon94a, MS99b, NPB92, Ois95, PR93, Rab90, RIG98, RNM97, RE94, RS90, SD98, SS99a, SZ97, SB96, SS99b, TOCH99, TM92, WGK99, YN95, YD96, Yua96, ZW97, vdFvGdRdV95, Abd90, AP98b, Ama94, AB96, AKV92, Ano97-34, Ano97-35, Ari95, AS96, BGN^{+}90, BOS97, Bak99, BK90, BV99, Ben97, BJ99, BP97b, BW91, Bru96, BCG98, Bul90b, But92a, BELN92, Cah90, CMR95, Car93, CS93b, CZ93, Cel98]. numerical [CH98, Cha99, CW99, CS96b, DPS98, DP93a, DBFV96, Ehr95a, Eir93, Eme94c, Eme94f, EM98b, FPS99, Fas95, FD93, FW99, GPCGV91, Gov94, Gro91a, GH99b, Han94a, Hem96, Hom96, HJ95b, How90, Iqb95, Jia91, JNC99, Joe90, JL97, Kha96, KSFAH98, KL99, KZ92, Kre95b, Kre11, KBÖ97, Lau96, LW91, LWZ98, LL99a, Mas99, MO96b, Mau98, MM97, Mel97, MJG98, Mön96a, MPS96, NR90, Pag92, Pag98, PS93a, PRT97, RA96, Rán92, Sak99, Sas97, Sca94, Sch99, Sim96, Sim90, Sim92, SR92, Sim94, SDS94, Sim95, SW97, Sim98, Sim99, Sin92, Spi96, TSM96, TS97, Tic90, TLL91, Usm92, VDM96, VV98, Wam99, Zha96b, dDGZ99, NS95, Bul94a, Eme94c, VR91]. Numerically [DD95, HHL97, HC99, dMM99]. Numerics [Ano97-51]. Numerov [Fra95a, Men91, TSM96, VVVD97, VD91]. Numerov-like [VD91]. Numerov-type [TSM96]. Nuttall [CDL96a]. Nyström [PT96, CB95, DVVV97, FP97b, GL94, Guo95a, OS94, SF92a, SF92b, SDS94, Tsi98, VVVD97, hC93, vdHM97].

O [ABM96, SS91b, Ano94l, Eme94f]. Obrechkoff [RA97]. obstacle [Han94a, HAMeM96, HBL96, KZ92, Lov95, MM91, Wes90]. obstacles [Isa92]. obstruction [Han94b]. obtained [LMM94]. occurring [Boe91, BBH90]. ocean [LL99b]. oceanographical [ADH98b]. Octabasic [SS96a]. ODE [BVZ90, Hig92, Hig93, NR90, SW95b, Voe95, dS97b, vdHS98]. ODE-solvers [BVZ90]. ODEs [AB97a, AB97b, Cel98, Cha98b, DS96b, Eir93, FGR97, GHMZ94, GR97, LR93, MS98b, Sha93, SS96b, VM99]. Oeuvres [Bul94b]. offstep [FS90]. old [MAR93a]. omitted [Ren99]. One [DLP98, Lop90, AR92a, BP96b, BCP99, Bre90a, Car93, Ekh91, FHR98, Hem90a, HS94b, HLFC99, Hug94, Jia91, KBÖ97, KBÖ99, Med98a, MJG96b, Nak91, OO98, Pag92, RIG98, RP96, SJAM98, SY90, Sim92, XZ98]. one- [OO98]. one-dimensional [AR92a, Hem90a, Hug94, Jia91, KBÖ99, Med98a, Pag92, RIG98, SJAM98, Sim92, XZ98]. one-leg [HLFC99]. one-line [Ekh91]. one-phase [KBÖ97]. one-space [Nak91]. One-step [Lop90, BCP99]. one-way [RP96]. only [SY90]. Open [Van93b, MP97, Mön96a]. operational [DLMT99]. operative [KKNU97]. Operator [Bul96c, FS96a, Arg92b, Bab99, Bul96a, Bul96b, Buo96, BF94b, CL91, Hyv97, JS93, Kal95a, LV99, Lem94, Ois95, YH95, YA95, Zha96a, ZZ91, Bul91]. Operators [SW96, Alp95, Ano97-31, Bav97, Bav98a, Bav98b, BK99a, BM93a, BM95b, DL96, Dzi99, FM98b, GlZ94, Gör99, Gra93, GH97b, IRZ96, Tri93]. Opportunities [JDV95].

OPTAC [AV96]. optics [ER96]. Optimal [Abd98, BP91, BR92b, Lay93, Lov95, MM99b, Nou93, Pro95, SJAM98, APH94, BHS^{+}98, Ber93b, CSS93, DR90, De 91b, Die94b, DGH95, Eln97, EK98, GWC96, HNT98, HLS94, JS97b, KSFAH98, Lit92, Lit95, Osa94, SS97d, YY98]. Optimal-by-order [MM99b]. optimality [McC99]. Optimization [Ano97-36, Ano97-33, GVO94, KKNU97, AV96, Bul90a, BJM97, DK92, FT96, FM94, FM97, GV96, Kla94, Moh90, Moh95, PZ98, Ree94, Rum95, Sch94a, Wol94, ZZ94, ZP95]. optimized [Hoc97]. Optimum [But93, MS98b]. options [ZFV98]. orbits [CKP94, Mey94, Yam97]. Order [Moh97, Thr90, Ada98, APH94, AW98, Alp95, ÁN98, AZ97, AÁNMP98, AP92b, AH95, AS96, BKM95, BPP97, BMS98, BS98a, BT91, BC93, BJM97, BS94, CL99, CMR90, Cha98a, CAZ98, Che98a, CS98, CW99, CLM95, Cop94, CES91, DNZ98, DBFV96, Dic91, Die94b, DTB93, DPTW95, Doh98, DP93b, DM97b, EK94, EM94b, Ext97b, FS90, FK95, Fit91, FLM92, FKR98, FRK99, FP90, Fra95a, FP97b, FN98, GHMZ94, GWC96, GS92b, Gro91b, GH97b, HU99, Han91b, Hem90b, Hus94, IY95, IP99, JV96, Jód90, Jo’90, JA92, JN92, Jon93, JKY97, KT97a, KKL97, KL99, Kla94, KSK92, KS99, LRV96, Lew96b, Lew97b, LOH92, LN96, LS95, Luc95, MP95, MM97, Med95, MM99b, MJG96a, MJG96b, MJG98, Oku93, Osa94, Osa98, PH92, PT95, Pfe91, PZ98]. order [RA97, RBDM90, RZG95a, RGZA98, Rov96, STS91, SK96, SS94a, Sca94, Sca95, SF92b, ST97, SS92, Sid97, Sim92, SR92, Sim94, SDS94, Sim99, SV92b, Sun96, TD97, TE97, Tre92a, Tsi98, TM92, Usm92, VVD94, VVVD99a, VS90, VVVD99b, VVD92, VMV94, Ver93, WL92, WA98, YN95, YMW98, ZRG93, Zha94, ZLQ96, ZCG99, vN93, vdHSvdV95]. Ordered [Stu95, HN90]. ordering [FMT91]. orderings [Ish99]. orders [IKF91, OS94]. ordinary [AL97a, BJZ93, BT96, But92a, CKP94, Elo98, GPS92, KNST91, LB99, Som93, Tra96, Wri93]. organisational [VBK91]. oriented [GN96, Mah92]. Orthogonal [Abr91, Ano91-36, Ano93w, BM93a, BEL93b, CJT90, DZY^{+}94, DG93, DP95, ELW93, JN99, KKL97, MG91, MS93, MAR93a, Peh93, SV96, Skr92, Tot98, Val93, Van91a, ALR96, ACM98, ÁNGM95, ABM97, Apt98, AGRZ97, AGRZ98, AÁNMP98, Bar90, Bav95a, Bav98a, BK99a, Bel90b, BV95, BC99, BHJ99, BM93b, BI92, BGVHN93, BGVHN95a, BGVHN99, But92a, CSE95, CSvD98, De 90a, De 93, Deh94, Dou96, DM90, DE97, Ehr98, FAD99, FS96a, FG92, FKR98, FHR98, FRK99, GMS95, GRZA97, GRZA98b, Gro91b, GIM96, HKL96, JMMW90, JKY97, Kal95a, KM96, KW90, KvD95, KKL98, Koe93, KS98, KR98, KR99, KLL97, KY97, Le 93, Leo98a, Leo98b, LL98a, Lew95, Lew96a, Lew96b, Lew98, Li99a, Lin98, Mag95a, Mag95b, MDR90, MBP93, MPPP95, MPP96, MMPP97]. orthogonal [MMFMB97, Mar93b, MFMBPC97, MF98, Mei93, MP99b, Mig95, Mil93, MR94, MDR98, Mon96b, MBMF99, Mul93, NW97, Pan93b, Pan93c, Pan95a, Pan95b, Pan96a, Pan96b, Pan97, Peh92, PS95b, PP93, Ran96, Ren96, RM95a, RBDM90, RZG95a, RZG95b, RM95b, RGZA98, SRD98, SS94b, Sla93, Sza95, VV95, Val96, VB93, Voi93, Yán99, dlCYL98, ZDR90, ZRG93, Zhe97]. Orthogonality [CS96a, RBM91, BC98a, GM97, GM95, JKLL95, KR96, Ron90, Sza93, dMPP98]. orthogonalization [Stu95, Wat97]. orthogonalizing [EEKL93]. orthonormal [MP93]. orthotropic [CKM97, Zha91]. oscillating [CFMR96, Dib98, FGR97, SZ97]. Oscillation [BM96a, BBM96, CHM94, CL98, GS92b, KPP97, MB92]. Oscillations [Cus94, APH94, FSZ94, Ham99]. oscillator [Dun98, GK96b, LM90, RE94, Wit96b]. oscillators [JV96, LMF99, Wit95]. oscillatory [CAZ98, DP97, EW99, Lev96, Lev97, MM99b, OM91, Rov92, SV92b, VVVD99b]. osculatory [LL91]. osteotomy [NS95]. other [MKK96]. outer [Ree94]. output [KM94a]. Overconvergence [IL99]. overconvergent [SS91b, Sty92]. overlapping [Mie98]. overrelaxation [Bai98, HS92a, Ish91]. overview [Fra95c]. Owe [Mar95a]. Oxford [Ano98q, Van91b].

P [Ano93k, Ano98p, Bul90b, Roo90b,

Ver94b, BHS^{+}98]. package

[AV96, Mou93, Ano98r]. packages [GH97a]. packet [Mon96b, Sif98]. packing [Hal92]. PaCT [Ano97-47]. PaCT-97 [Ano97-47]. Padé [AC90, AP92a, AW95b, BGM93b, BGM94, BGM95, Bec90, Bec92, Bes96, Bre90c, BV90, BGVO95, BDMGVO97a, BDMGVO97b, BL93, CLB92, CDL96b, CDL96a, Cuy99, DDL93, DVB96, DMGVO99, Dri95, DT98, Fie90, GM93, GP97, GP98, GP99, GVN90a, GVN90b, GVO94, GM90b, GMS91, GMR94, GMBW96, GMR97, Gui97, HRGL98, Hoc97, IL99, KW90, Le 95, Mat97, Pas90, Pre90, Pré96, Rob96, RW92, Sid90, Sta97, Sta98, Thu99, Tok96a, Tok96b, Wim90, dlCYL98, Zho97]. Padé-based [Fie90]. Padé-type [Bre90c, DMGVO99, GP98, GVO94, HRGL98, Le 95, RW92, Thu99]. pages [Ano93k, Ano94j, Ano98p, Ano98q, Ano98r, Bul90b, Bul94a, Bul94b, Bul96a, Bul96b, Bul96c, Die94a, Kra96, Roo90b, Uyt96, VR91, Ver94b, Bul91, Bul94c, Van91b]. Painlevé [Mag95a]. Painlevé-type [Mag95a]. pairs [Bab96, EH97a, MPPP95, MMFMB97, MP99b, Pan97, SF92a, SF92b]. Pan [Bul94c]. Panovsky [CB95]. Paperback [Eme94e, Ano98r]. papers [Ano91w, Ano94y, Ano96-27, Bul94b]. Para [BHJ99]. Para-orthogonal [BHJ99]. parabola [Lor92]. Parabolic [Ano91p, AR92a, BM96a, Bak99, BCG98, Car93, CL98, DH94, DKKV98, Epp91, HLS94, KSFAH98, LC98, LST94, LT93, MJ96, Moh98, MV92, Nak91, NW94, PP97, Sch90, Seg95, ST99, SSV98, Sty96, Sun96, Swa96, XC96]. Parallel [AM96, AB97b, Ano91-30, Ano93s, Ano93z, Ano96-33, Ano97-47, Ano98r, BVZ90, DHP94, GP93, GN96, HS97, HL95, MdSvdV97, Som93, VS90, Wri93, hC94b, vdHSvdV95, vdHSdS96, vdHM97, vdHM99, Anh99, AD92b, Bai97, Bai98, CGR95, Cha96, CRvdHV93, DBFV96, Edd90, Ern96, FFH97, FG97, FS95, FR99, GC97, Hei95b, JW91, KNST91, KC92, Mar91b, Mon96b, Sha94b, ST99, SFI95, Vec96, VM99, YK96, ZP95, dDGZ99, hC94a, Ano94l, Roo90b]. Parallel-iterated [Som93]. parallelepiped [Moh95]. parallelism [Gea91]. parallelizable [JB99]. parameter [Ehr96a, Köh93, Plu95]. parameter-dependent [Köh93, Plu95]. parameterization [Eln97]. parameters [Ano96-34, Ish99, Kob95, Lah92, Nou93, RA96, RA97]. parametric [CM96b, Fis91, HZ96, Man96, SFW99, Zha91, ZC99]. parametrically [ESEHEK93]. parametrized [TB97]. Part [Mas92b, Sch94c, Sch94d, BR92a, Bul90b, Coo99, Cri97, Ell95, JN92, LU96, Mas92a, MO96b, Ors97, BGVO95, BDMGVO97b, Nak91, TD97, TE97]. Partial [Ano91z, KKL98, Van90a, Van90b, AG98, AR92a, BPV93, BJ99, CGR95, Cas96, CS99, CPK96, DK94, GMS94, GSSV92, Han91b, JL92, Jo’90, JNC99, KKL97, MLD96, McC93, MS92, RVY97, Sin92, Tri93, XC96]. partially [Xu99, BGM96a]. participants [Ano93x, Ano95t, Ano95u, Ano95v]. Particle [CL97, Bre90a, Cou99, Eme94f, Hac97, LC98, RL99, Syd99, Vrb95]. particle-grid [Cou99]. particle-in-cell [Syd99]. particles [Jer95]. particulaire [Bre90a]. particular [JL92, MLD96, Sch97]. Partington [Ano98q]. partition [All96]. partitions [SWY98]. past [Cha97, TF91, TOCH99, Tre91]. patched [Roo90a]. patches [LS96]. Patchwork [Kem90a]. Path [Goo91, Ext91, Ext95, Kor96, PP92, SLK91]. Path-integral [Goo91, Ext91, Ext95]. path-integrals [Kor96]. Patricia [Ano98p]. pattern [Ano94k, Gje97]. patterns [PC96]. payoffs [Hür97a, Hür97b]. PCG [CY91]. PDE [ABM96, GN96, Jer91]. PDEs [CM97, Car93, MM97, Sch94f, SSV98, VIA96]. Peano [Ehr96b]. Pearcey [KP99]. PECE [Zha94]. Penalty [Li97, ZFV98, Li99c, LC99a]. pencil [MKK96]. pencils [Par91, ma92]. Pepper [Eme94d]. Performance [Ano97-36, FMT91, CPK96, DT93, Eme94e, MKK96, hC93]. period [EHT90, EH92a, JV96]. Periodic [Alw96, FSZ94, HLS94, LP94, APH94, Bal93, BM93a, BF94b, CDF95, DH94, ES90, FZ93, Fra95a, FN98, GL99, Gil90, Han91d, KPP97, Kor94a, LM94b, LP98, Mey94, Rab94, SSS96, SG94, SS92, SDS94, Thr99]. periodic-parabolic [DH94]. periodically [CHM94, JM94]. Periodizing [Lau96]. permittivity [HS92c]. permutation [SS96a]. permutations [Fau90]. Perturbation [IS95b, Arg91, Dei95, GMdLSG99, Wen90]. Perturbations [MP99a, Ano97-31, BDS96, Bav97, Bav98a, Peh92, Rha92, SS97c, Wei97, Wit95]. perturbed [Bog98, BEM97a, Che97a, CLM95, Fit91, GPCGV91, Hem96, JP91, Kul92, Leo98a, LMF99, MDR90, Med98b, QS99, Roo90a, SV95a, Sin92, ST99, Wan97]. Petersson [Van91b]. Petviashvili [FM98a]. Pfaff [And96]. Phase [Mai93, CDM91, ER96, KBÖ97, MW90b, ND95, PRS90, PT96, Sim92, Sim94, XZ98]. phase-errors [PT96]. phase-fitted [Sim94]. phase-lag [Sim92, XZ98]. Phase-type [Mai93]. phases [VV98]. phenomena [BTCC99, HS96, Yan95, Ano91q]. phenomenon [GSSV92, Par92, ZM97]. Phillips [Rha92]. physical [Böc96]. physics [BTCC99, LSZ96]. Picard [Hyv97, Ozi92]. Piecewise [Ill98, AW95a, DLR92, HMS96, Hig92, KL98, SS94c, YD96]. piecewise-Lidstone [AW95a]. pinched [Zha91]. pivoting [Fos97, Fos98]. placement [CSS93]. planar [eM91, MW95, MW99b, RF99b, RF99a, SS94c, WM96a, WM96b]. Planck [PdA99]. plane [ADG99, Alv99, Ata94, CH96c, Ehr94, FT98, GM93, HB92a, Han91d, JK98a, Mah92, McC93, Pon98b, TRBS99]. plane-oriented [Mah92]. planning [TPA95]. PLAPACK [Ano98r]. plasma [Syd99]. plastic [CZS96, GK95, Ned97a]. plasticity [Ned97b]. plate [Ad94, ADH98a, AK91, BG95, Ber94, lCHcH97b, Dor99b, Rab96]. plates [Boc95, KZ94, Sca95, Wes90]. plots [EHK^{+}95, FW99, Weg98]. plotting [Hig92]. plus [Sca95]. Pochhammer [ÁNQR99]. POCS [SÁFV96]. Poincaré [Bea99]. point [Ari95, ASW97, BJZ93, BGVO95, BDMGVO97a, BDMGVO97b, CS93b, Che98b, DF95, DMGVO99, EGBE93, Eme94a, Fay97, GG97a, GVN90a, GVO94, GN96, HB92a, HS92c, JS91, Jer91, Jon93, Liu97, MM99a, MW90b, OO94, Paq94, Pry94, QS99, QA97, Sab95, SH96, SA97, Sin92, Tok96a, Tok96b, VVD92, Wic90]. point- [GN96]. point-like [OO94]. points [CR96, FS90, FW99, GL90a, GL91, How90, JP91, Lev99, LR96c, Mag95b, Mor92, RS99, SX99, Sta93, SLK91, Sty92, Tic90, Waz91]. pointwise [BZ93, Hug94]. poised [MN96]. Poisson [BELN92, GR92d, IG90, Kem90b, LOH92, OO94, PCV99, SS97a, YO98]. polar [MJG98]. poles [Alv99, Ber97, Has97, Kza95, Li99a, Müh90, Pan95b, Sta98]. Pollaczek [CI97, Dam99, Wim90]. pollution [LV99, OF96]. poly [BBD^{+}96]. poly-iterative [BBD^{+}96]. Pólya [MH94]. polygamma [Köl96b, Lóp98, Ada98]. polygon [CL91]. polygonal [DM97a, Han94b]. polygons [How93]. polyhedra [DL94a, DL94b]. polyhedral [LN95]. Polynomial [Ano97-48, Ano97-49, BM92, BW96, KT97b, PP91, Van93d, WM92, Aco97a, AH95, BM98, Bla95, CP93, CS95a, CS95b, CM96c, DC95, Ext98c, FS96a, FB90, Gem96, GRZA98a, GPRV93, Haj95, IY97, JS92, Jia94, JMMW90, KKZN97, KKL98, KLL97, KY97, LR95, Loe96, Lyn92, ME93, Mas95, MLD96, Miy92, Miy93, Mus99, Paq94, Pas91, Per92, Pet90, PK97, STS91, Sch90, Sid91, Sto95, vG95, Bul94c]. Polynomials [Ano91-36, De 90a, AS99, AP98a, ALR96, ACM98, ÁNGM95, ÁNYD98, ÁN98, AGR93, Ano97-31, ABM97, Apt98, AGRZ97, AGRZ98, ADMFSR98, AÁNMP98, Ban99, BHD93, Bav95a, Bav95b, Bav97, Bav98a, Bav98c, BK99a, Bel90b, Bel93a, BV95, BC99, BM93a, BHJ99, BG98, BM93b, BEL93b, BI92, But92a, CS97a, CSE95, Che98a, CIM94, CI97, CSvD98, CJT90, CDM91, DLMT99, De 93, DJ98, DZY^{+}94, Deh94, DG93, DP94a, DP94b, DP95, Dou96, DM90, DE97, DT98, Ehr95b, Ehr98, ES92, EEKL93, ELW93, Ext96, FAD99, FG92, FV95, Flo96, För93, FKR98, FHR98, FRK99, Fre92, FG93, GS91a, GMS95, GL99, GRZA97, GRZA98b, GR92c, GS99, Gro91b, Grü98, GHH99, GIM96, HKL96, HS94a, He94, Hen90a, Hen90b, HS97, HV91, HN98, IS93, IS95b]. polynomials [IY95, IS98, JR93, JNS90, JTNW93, JN99, JKY97, Kal95a, KR96, KM96, KW90, KvD95, KKL97, Koe93, KS98, Köl94a, KS96a, KS96b, Kui95b, KR98, KR99, KYY97, Lan98, Lap93, Le 93, Leo98a, Leo98b, Let93, Let95a, LRV96, Let99, LL98a, LS98, Lew93, Lew95, Lew96a, Lew96b, Lew97b, Lew98, Lin98, LMJ93, LM95, Mag95a, Mag95b, MDR90, MG91, MBP93, MS93, MAR93a, MPPP95, MPP96, MMPP97, MMFMB97, Mar93b, MFMBPC97, MF98, Mas93, MO96a, McN93, McN97, McN99, Mei93, MP99b, Mig95, Mil93, MR94, MKK96, MBMF99, Mul93, OSG^{+}93, Pan93a, Pan93b, Pan97, PSG91, Peh92, Peh93, PS95b, PS95a, Peh95, PP93, Rad91, Raf94, Rah96, Ran96, Ren96, RBDM90, RZG95a, RZG95b, RM95b, RGZA98, Rös95b]. polynomials [RV98, Sab93, SRD98, SV95a, SS94b, Sch93, Sch98, SS96a, SV96, Skr92, Sla93, Sza93, Sza95, Tot98, Tri93, Val93, Val95, VV95, Val96, Van91a, VS93, Voi93, Wen97, Wim90, WK98, WZ97, dlCYL98, ZDR90, ZRG93, Zen95, Zhe97, Zhe98, Zhe99, Zhu97, dR99, dMPP98, vD95, Ano93w]. polytopes [BC91]. Pommerenke [CDL96a]. Pontryagin [Alp98]. population [Cus94, FSW94]. porous [CC99, HB90, HB92b, MP97, MPS96, Ozi91b, Yan90]. portable [AV96]. posed [EH91]. Positive [AO98, BS98a, Dib98, Elo98, FT98, Gra93, HS94b, AW98, AAP98, CLL98, CMR99, DH94, För95, IS90, IKK91, IS92, KvD95, Lah96, LW96b, Let96, LU96, MTW96, Odl92, Pop96, Sha94b, Wen99, ZCG99]. positive-type [MTW96]. Positivity [LD98, CGEP96, DE99]. possible [Hug94, LW96b, LU96]. posteriori [AP96, CS99, Hug94, Joh98, JL99, Mon98, NYW98, Seg95, Ver94a, Wan97]. Postfach [Kra96]. potential [Abd96a, CL93, FAD99, Goo91, HB92a, MRT95, MR92a, Wit96a, Wit96c]. potential-theoretic [FAD99]. potentials [BEM97b, BEM98, MZ95, RS94, Wit96c, WGK99]. Pounds [Ano94k, Ano94l, Ano94m]. pour [Bre90a]. Powell [Lai96, WD94]. Power [Ezz94, EH94, Gil99b, Hai90, NKS98, PGGC98, PRS92, RJ99, SF90, Van93d]. power-boundedness [Hai90]. Power-law [Ezz94]. powers [Köl95b, KS96b, Lan98, LM95]. pp [Lev90a, Lev91, Pie96]. Practical [Ano96-33, Die99, Ish99]. Prandtl [CMR95, CCJ97, Dem95]. precipitation [MP97]. Precise [Paq94, Sch94a]. precision [Pet99b]. Preconditioned [YI96, AL97b, BMP96, Che98b, EBP96, KNO97, Mee96]. preconditioner [JB99, Ram99, SS90, vG95]. Preconditioners [Rod99, CS97d, Jin96a, Kin90, Mar95c, Vui95, Yun98]. preconditioning [AM96, BT92b, GS97, JW91, MS92, SSF95]. predator [MT98]. predicting [Tre91]. prediction [Vek97]. predictor [Bur93, Kha94, Sim95, TSM96, TS97, UH98, vdHSdS96]. Preface [Aga98, Ano91-31, Ano97-50, Bre90d, BGPW91, BGPW94b, BGPW96, BGPW97, BBS98, BV97, BCvdH93, Deh93, Dur98, GS92a, GWP95, IM96, LM94a, LN90, MW99a, Mon92, Ned95c, OC96, PS93b, PC99, RR94b, RW99, Ver91, Yam91, Yam95, dB95]. premiums [BLG91]. Prentice [Uyt96]. Prentice-Hall [Uyt96]. prescribed [Ber97, Müh90, Rab94, Van93d]. presence [EEAES99, RS92, SS97c]. presentation [LGT90]. presented [Ano91w, Ano93y, Ano94y, Ano95x, Ano96-27]. preserving [CMR99, KvD99, Lah90, Lah92, SU99]. Press [Ano94j, Ano94k, Ano94m, Ano98q, Ano98r, Eme94a, Eme94e, Eme94b]. pressure [BCP99, ZZ97]. prey [MT98]. Price [Ano94j, Ano94l, Bul94a, Bul94b, Die94a, Eme94e, Eme94b, Lev90a, Lev91, Roo90b, Ver94b]. primal [CK91, Fay97]. primal-dual [Fay97]. primality [Lev90a]. prime [RJ99]. prime-power [RJ99]. Prince [CMR90]. principal [BK99b, Cri97, CES91, DS91, DL96, DPS98, Die94b, Die95, Die96, Die97, Ioa90, Ors90, San96]. principles [BKM95, FN98]. print [Lev91]. priori [JNC99, MM99b, NYW98, Ols99, TB97]. priority [Cam91, Rad97]. Probabilistic [LB98, TG91]. probabilities [DH95]. probability [CM95, LJQ95, Tag98, Ury94]. Problem [BP96a, A.94, Abd98, ADH98b, AH98, ÁNQR99, AAP98, ADT96, ADT97, Anh99, AD92b, BPV93, BK95a, BV99, BCV98, BF94a, Bec90, BM98, Boe91, Bog97, Bog98, BHJ99, BGVHN99, But92b, BR92b, BELN92, Che94, Che97a, lCHcH97a, CL91, Chr98, CLM95, CM92, CM96a, DS99, DDG98, DKKV98, DRW95, EFL97, Ern96, FKM97, Fun92, GWC96, GGF99, Gre93, GB90b, Hal92, HvdV94, HB99, HS92c, Hem96, HSW91, Ish91, Kal95a, Ked93, KSFAH98, Kre95a, KZ92, Krí95, KBÖ97, LR94, Lau91, LST94, Leó99, LS99, LC99b, Lin99, MRT95, Mey94, Moh95, Mön96a, ND95, Ned95b, Ned97a, Ned97b, Njå95, Njå98, OO94, Ozi91a, Ozi91b, PP91, PKB98, Pas91, Pen99, QS99, RIG98, Roo90a, RS90, SF91, Sca94, Sch94b, Sci94, Sha94a, SF92b, Sim96, SIC92, SR99, SS97d]. problem [TD97, TE97, TPA95, TG93, Tse95, Ush91, Usm92, VB90a, VB90b, VVD94, Ven96, Wan97, Was95, WD99, Wes90, YL97, YY98]. Problems [Ano97-39, Ask93, GMR97, APH94, AO98, AP98b, ACGW91, AB96, AKV92, Ano96-34, ABD^{+}96b, AFK93, BAG^{+}96, BFG96, Bai98, BE98, BG95, BGH98, BR98, BL97, Ben97, Ber95, Ber98, Ber94, BS93b, BF93, BP97b, BGVHN93, BS94, CNP99, CL99, CFMR96, CG94, Car96, CZS96, CS93a, CS93b, CZ90, Cha98a, Cha99, CAZ98, Che98b, lCHcH97b, CLW95, CF93, CK92, Chr90, CG97, DH94, Dar98, DVB96, DB99, DV99, DOAG91, DM97a, DTB93, Dor99a, Dos95, DGH95, EGBE93, Eln97, EK98, Elo98, EH91, EK94, FT96, Fas95, FFH97, Fea91, FFRV92, Fra95a, FG97, FN98, FP99, GM90a, GK95, GBP97, GK97, GRZA98a, Gro91b, HB92a, Han94a, HZ95, HR94, HY98, Hug94, Ira94, JSWW95, Jód90, Jon93, KNST91, Kar94, KNN96, KT97a, KvdB92, KKS96b, LW96a]. problems [LV99, Lar94, Lay93, LDV91, LM97, Li97, LLY99, Lim94, LG95, LT93, LYFW95, Lit92, Lit95, Liu97, LN96, LP98, Lov95, LN95, LN99, Lui98, MFI95, MP95, MBP93, MP92, Mar93b, Mas99, Mat98, MS98a, MR92a, MM91, Moh97, Moo95, Nak91, NW94, Ned94, OF96, PP97, Pan95c, PACH90, Plu95, Pry94, QA97, Ram99, Ree94, Ren97a, RS99, RW97, Ros95a, Rum95, Sac96, SK96, SÁFV96, SSY98, Sch94c, Sch94d, SH96, SWW96, SA97, ST97, SS92, Sim92, Sim94, SDS94, Sin92, ST99, Spi96, SSF95, Ste99, Sty95, SFI95, Swa96, Tak92, Tem95, TB97, TM92, Van93b, VD91, VVD92, Voe95, Weg90a, Weg93a, Weg93b, Wei97, Wei96, Wim91, WA98, Wri93, Yan92, Yua96, YI96, ZV95, ZL96, ZL99, ZP95, hC94a, hC94b, vdHSvdV95, Ano91p, VVVD99b]. procedure [FT96, Leo98b, Moo95, TK98, Zha94]. procedures [Arg98, BTCC99, BS93b, Hem90b, KKS96b]. process [Bel90a, BGF97, Dub98, Gra93, Hyv97, Sid99, Val93]. processed [Sha94b]. processes [Arn92, Bec94, BV95, Gra96, SV92c, Val96, Wat97]. Processing [Ano93z, Uyt96, Ano91f, Uyt96]. processor [FG97]. processors [Edd90, SOMF91]. produced [Nie90]. Product [BR92a, JM96, LL99a, SL99, Spa99, All96, Bar97, Bav95a, BM98, BP97a, BP96b, CH94b, DS90, DS91, De 93, HT91, Jüt97, RS92, SL98, Wit95]. production [Lis95, TPA95]. products [ÁNYD98, FFH97, GMS94, KM96, Köl95b, Luc95, Mav95, Mil97c, Tre99]. Professor [Ano91h, Mar95a]. Programming [Ano93s, Ano96-33, CZ90, Dup94, DRW95, Eme94a, Kan94, Lep94, LYFW95, SFI95, TPA95, Vog94, Was95, WM92, YZY96, ZL96, ZL99, Ano94j, Van91b]. programming/branch [DRW95]. programming/branch-and-bound [DRW95]. programs [Art94, CQW95, Fra94, Sch94e]. Progress [Bul94c, GMR97]. progression [Ber98]. progressive [eMM97a, AE90]. Projected [AP92b, Ish91]. Projection [Bre97, Lep94, Arg91, Arg92b, BC94, CZ90, HM97, Jun94, Lar94, Liu95, Zha96a]. projective [Lub93, Ver96, ZZ94]. prolongations [De 90b, Hem90b]. proof [Ber99, Bre90b, Cap90, Ekh91, Han96, IS95a, JMT94, Pré96, Sya99]. proofs [NW94]. Propagation [Ano91q, KKS96a, MET93]. propagators [Kal95b]. Properties [Mul93, ALR96, ÁNGM95, And92, AZD91, CG97, DR94, Die95, Emm97b, Fas95, FD93, GR92b, GR92d, JVZ97, JMMW90, Kuc99, MF98, Med96, Moo95, MBMF99, Nie90, NW97, Pan95b, PS95a, PP93, Pon98a, Rad95, RV94, WK98, ZD94, dO90]. property [AR99]. proportion [Par97]. proportionality [Hig93]. proportioning [Dos95]. proposed [Gro91c]. protein [BB95]. protein-dependent [BB95]. proved [Nee94]. Pruess [MP92]. pseudo [Con99, Fra99, GHMZ94]. pseudo-spectral [Fra99]. pseudodifferential [SS94a]. pseudorandom [DEHG92, EHT90, EHN91, EH92a, EH92b, EH94, EH95, EH97a, Emm97b, Nie90, Nie94, Rip90, Weg98, Zie90]. Pseudospectra [DT96]. Pseudospectral [EK98, MS92, RE94, Ren97b, Slo91]. psi [AR99]. PTR [Uyt96]. Publications [Van91b]. Publishers [Ano94l, Ano98p, Bul90b, Eme94f, Ver94b]. Publishing [Bul90b]. PULSE [Lis95]. pulsed [PCC95]. Pure [Bul91, Mas99]. purpose [Zha98].

QD [dR99]. QD-algorithms [dR99]. QMR [Cel98]. Quadratic [BF94b, EH95, EH97a, CQW95, CMR99, Dri95, EHN91, EH92b, EH94, ES90, Eln97, Emm97b, FK96, HMS96, KK95, Lah90, Lah92, Lah96, McC98, Rol99, SFI95, Val96, WM94, YZY96, ZC99]. quadratically [Sto95]. Quadrature [Ano91z, Ano93w, BHGVN94, BGVO95, BDMGVO97a, BDMGVO97b, Hag99, SS91a, AKV92, BES91, BK99b, BDV94, Boj96, BGVHN97, CH96b, CH98, CG97, DS90, Die94b, Die95, Die96, Die97, Die98, Die02, Ehr95a, Ehr96b, För93, För95, GL90a, GM97, Gra93, GH99b, HN98, KL98, Köh93, Kui95a, Kui95c, Kui95b, Kza95, Lan95, Lau99, LW91, LW95, Let98, LM95, LD99, MM99b, Min98, Nik96, Not90, Not97, Not98, Pat99, Pet99b, RV94, SS94a, Sin95, Smi91, Sya99, Tsa96, VDV90, Ver97]. quadratures [Alp95, DS91, GL91, Sch96b, Spa99]. quadrilateral [KZ94, ZV95]. quadruple [MW90a]. Qualitative [FB96]. Quantitative [HZ95, Sid90]. quantization [Pag98]. Quantum [FV95, Wit96c, FV96, Jer95, LSZ96, MOT91]. Quartic [XGSSKF96, Alw96, HS94b, KS97, SK96, Usm92, Val93, Val96, Wit95]. Quasi [Fre92, Mar91a, Ron90, Sab96, And92, Ano97-42, CDF95, Che97b, CMR99, Dem95, FS91, FG91, FM94, FM96, FM97, Li96, Saa94, San96, Sen99, SWY98, Xu99, ZZ94]. quasi- [FG91]. quasi-cross-cut [SWY98]. quasi-interpolant [CMR99]. Quasi-interpolants [Sab96]. quasi-interpolating [San96]. quasi-interpolatory [Dem95]. Quasi-kernel [Fre92]. quasi-linear [And92, Sen99]. quasi-Monte [Ano97-42]. Quasi-Newton [Mar91a, Che97b, FS91, FM94, FM96, FM97, Saa94, Xu99, ZZ94]. Quasi-orthogonality [Ron90]. quasi-periodic [CDF95]. quasi-regular [Li96]. Quasiconformality [PS99]. quasilinear [Anh99, Bog97, BCG98, Ern96, Faz90, KSFAH98, Mat98, Sty96]. quasistability [BDS96]. question [Lor95]. questions [Ven96]. queueing [BHD93, Kri92]. queues [HJ95a]. Quick [Lew92]. Quintan [Roo90b]. quintic [HS94b, Lai96]. quintuple [All96]. Quotient [SF90, NKS98]. Quotient-difference [SF90].

R [Ano94l, Ano98q, Ano98r, Bul94a, Eme94b, Lev91, VR91, Ver96, BD90]. radar [AEG98]. Radau [GL90a, GL91, HW99, dS97b]. radial [FI96, SW96, TSM96, TS97, Wen99]. radially [DKKV98, Rab96]. radiating [GL98]. radiation [Dor99a]. radiative [DZY^{+}94]. radiative-transfer [DZY^{+}94]. radii [Tok96a]. radius [HHL97]. Radon [AZ97, Ito94, Lub93, Tri93]. Ramanujan [AP92a, BE91, BCH^{+}99, BB93, FGKP95, Gro91c]. Random [Edd90, Fus90, LSTZ94, Tic90, VS93, Aff90, And90, BM93a, CSvD98, Gen90, Ham94, HvdV94, Mah92, Pal92, PKB98, Rin94, Sch97, Sez92, SC95a, Sta90, Voi95]. Randomization [Joe90]. Range [BWS99, HMS96, LR95, Sch94a]. rank [Lan99b, SSY98, Sid91]. rank-deficient [SSY98]. Raphson [Lyn92]. rapidly [BEM97b, Lev96, Lev97]. rarefied [Sch99]. Rate [BL93, HRGL98, IY97, KV95, Lin98, Zho94]. Rates [BGVHN97, HS94a, Fie90, Ham94, HK94, Rin94]. ratio [Sta90]. Rational [CGV97, Let96, MR92b, PACH90, RP96, VM94, Zhe97, Zho94, AE91, AC90, Ano91u, BSW99, BWS99, BL97, Ber97, BM97, BHGVN94, BGVHN95a, BGVHN97, BGVHN99, CGR95, CGEP96, CS95a, Chu97a, Chu97b, CDM91, GP99, Hai92, Ill98, KT93, Li99a, LR92a, MA98, Min98, Müh90, NW97, NM92, Pan93b, Pan93c, Pan94, Pan95a, Pan95b, Pan96a, Pan96b, Saa92, SU99, Sta93, VB90a, VB90b, VB94, ZT99]. Rationality [PGGC98]. ratios [EHG90]. Ravi [Ano93k]. Ravl [Ano98p]. Rayleigh [NKS98, Zit94]. Raynal [Roo90b]. reaction [Bog98, BN92, HMV98, JMT94, LD98, LV99, Pao98, PC96]. reactions [PCC95]. reactive [ABD^{+}96b, MP97, MPS96]. reactors [Faz92]. Real [EEKL93, BMX97, BI96, CNS97, CES91, Die99, JS92, KSK92, Mei93, Mel98, Sha94a, Sin95, Tok96a, Tok96b]. real-valued [BI96, Tok96a]. recirculating [KL99]. Recognition [Bab96, Ano94k]. reconstruction [HS92c]. records [DNS98]. recourse [CQW95, HvdV94, Sch94e]. Recovering [Els99, GSSV92]. recovery [BZ93, Lau99]. rectangular [BHS^{+}98, GKH91, Kar94, Let95b, Let98, Tim93]. rectangular-source [GKH91, Let95b]. Recurrence [Lew96a, Lew98, RZG95b, ABM97, AGRZ97, AGRZ98, Bre90b, CZ93, FHR98, GL99, GRZA97, IS95b, Lew92, Lew96b, Mag95a, MDR90, Sla93, Van93c, dR99, vVBZ93]. recurrences [Hoc97]. recurring [LMJ93]. recursion [BW91, Fus90, Lau99]. Recursive [JL92, Mes96a, Mes96b, CZ90, FRR94, HGL96, Kos94, Let93, Let95a, Let99, Mah92, MLD96, RZG95a, lX97]. Recursiveness [BL97]. rediscovered [PH99]. reduce [Fau90, Mou93]. reduced [Axe91, BMP96, PT96, Sid91]. reducibility [Ext97a]. reducing [GV96]. Reduction [ADH98b, BC93, GWC96, GS91b, JDV95, LMM94, Osb97]. redundant [FR95b]. refinement [Kle99, Kos94, Mit91, Ver94a]. reflection [RV98]. refraction [Ked93]. regarding [Iqb95]. region [Cha99, HZ96, McC98, Waa91, Weg96]. regions [CH94b, DE93, FZ93, GK96a, GM99, KW99, Lan99a, TRBS99, Weg93b]. registered [Ano95v]. regression [Atk90, DGC95, JDV95]. regridding [FD93]. regular [GPS92, Kha94, Li96, RBM91]. regularised [Slo91]. Regularity [JVZ97, Li99a, Gra96, Kla94]. regularization [EH91, FG93, Ren97a, SÁFV96]. regularized [AD97, Iqb97, PB98, Voe95]. Reichel [Bul94a]. Reid [Ver96]. Reissner [lCHcH97b, Sca95]. rejection [Kem90a]. Related [Ano91z, Ver94b, AH95, BV95, Bul96b, Faz90, GKH91, Gro91b, IS95a, JS91, Kan97, Kan98, Maa94, Mul93, OO98, Pat99, Roo94, SBK98, Val96, Ven96, Wen90]. relating [AS99]. relation [Dou96, Ext99, IS95b, LB99, Lew96b, Mil06, TB99, vVBZ93]. relations [AGRZ97, AGRZ98, Bus99, GL99, GRZA97, GRZA98b, GM95, GIM96, Han96, Lew92, Lew96a, Lew98, MDR90, RZG95b, Sla93, Van93c, dR99]. relationship [Bre90b, Lev90b]. Relationships [IY95]. relative [LM94c]. relativistic [BGM99, IM99, Syd99]. relativity [SS99a]. relaxation [BJZ93, CRZR96, Cho92, CRV97, DR94, DR95, GG97a, Han91a, Mie98, SFI95]. relaxed [Arg98]. Relaxing [EH97b]. release [AU96]. relevant [Sch97]. reliability [DNN92]. reliable [Bec92, YO98]. remainder [GL90a, Sch96b]. Remark [EHS98, All97, EH92b]. Remarkable [Cal97]. Remarks [Ben97, Gri99, Mie98, Mil98]. remediation [ABD^{+}96b]. Remez [Erd93]. Remez-type [Erd93]. repeated [AW95b]. reports [Ano90k, Ano91r, Ano91s, Ano93u, Ano94w, Ano94x, Ano96z, Ano97-37]. representation [CDTV99, FV96, FM98b, Mar95b, Mat97, PGGC98, Rob99, SFW99]. Representations [KS98, KM96, MR95a]. reproducing [Alp98, BGVHN95b, CMR99, RS90]. Research [Ano91z, Mee96, Pie96]. residual [AP96, LLP97, YN95, Zha98]. residue [PP96b, TB99]. resolution [Kha96]. resonance [Alv99]. Resources [Ano93v]. respect [Art94, Bav95a, Che98a, Pan93b, PS95b, RBM91, Tot98]. responses [WRW92]. Restarted [EBP96]. Restricted [Per92, BWS99, Sch94b]. restrictions [DH95, De 90b, Hem90b, HMS96]. result [DJR91, GHMZ94, Han96, Joe96, MMPP97, Nak91]. resultant [LMJ93]. résultat [Han96]. Results [BGM93b, Lew93, Lew95, AKV92, Bav98c, BM92, CG97, CFR96, Hem90a, IS95a, JS91, LM93a, Let99, NS95, Plu95, Sid99, TD97, Tim93, vD95]. returns [Pan96c]. reversal [AT99]. Review [Ano93k, Ano94j, Ano94k, Ano94l, Ano94m, Ano98p, Ano98q, Ano98r, Bul90a, Bul90b, Bul91, Bul94a, Bul94b, Bul94c, Bul96a, Bul96b, Bul96c, Die94a, Eme94a, Eme94d, Eme94e, Eme94b, Eme94c, Eme94f, Kra96, Lev90a, Lev91, Mee96, Pie96, Roo90b, Uyt96, Van91b, VR91, Ver94b, Ver96]. revised [Stu95, Ver96]. revisited [IDV98, McC99, Men91, Ven92]. Reynolds [KL99, TF91]. rheology [Ned97a]. Riccati [Ber93a, Die93, JN91, Jód92]. Richardson [CR96, CB95, Lim94, Sid99]. Riemann [CES91, IM99, Weg90a, Weg93a]. Riemenschneider [Die94a]. Riesz [Kan98]. right [HY98, Jbi99, Lim94, Mus99, SG96]. right-hand [Jbi99, Lim94, Mus99, SG96]. Rigorous [Ber99, Dei95]. RILU [BT92b]. ring [Bab96, SR99]. Ritz [Li97, LLY99, Zit94]. Rivlin [Boj96]. RK [Bar99b, Ber99, Con99, Lab98]. RKC [SSV98]. RKS [DB99]. Robert [Eme94e, Roo90b]. Robin [Wei97]. Robust [SSF95, JDV95, OM99, RT91]. robustness [Sur91]. Rogers [BCH^{+}99]. role [But92a, MS92]. rook [Fos97, Fos98]. root [Gri99, KKZN97, Osa94, VV92, Dub98]. root-finding [Osa94, VV92]. roots [AH95, Lyn92, McN93, McN97, McN99, PK97, SV95b, Yey92]. rotating [Dun98, vdFvGdRdV95]. rotation [GGF99]. rotationally [CLL98]. rotations [Sch97]. row [GMS91, Liu95, GH97a]. rows [Hoc97, IL99, Rob96]. rule [Ehr96a, GVGPSL93, LW91, Let98, RV94, SS97b]. ruled [CP99]. rules [AE90, CH96b, CH97, CR93, Coo99, DS90, DPS98, Die94b, GL90a, GK96a, GH99b, Joe90, Joe96, Lan95, Lan99b, Lev99, NS94, Pat99, RJ99, San96, SL98, SL99, SS91a, ZDR90]. Runge [CMR90, CRvdHV93, CRV97, BJZ93, Bru94, BC93, BJ93, CFMR96, CMGP99, Cha98b, CS97c, CB95, CV93, CRV90, DVVV97, GPMR96, HH90, HGC99, JB99, LMP97, MdSvdV97, OS94, Owr95, PT96, PRT97, SS92, SDS94, Som93, ST91, Tho90, Tra96, Tsi98, VVVD99a, VS90, VVVD97, Vec96, Ver93, VM99, dSS97, hC93, hC94b, vdHSvdV95, vdHM97]. Ruttan [Bul94a]. Ryzhik [BM99].

S [Bul94a, Die94a, Eme94e]. S. [Gro91c]. Sabin [Lai96, WD94]. saddle [Che98b, SLK91]. saddle-point [Che98b]. safe [DF91]. Sag [LD99]. Sakhnovich [Bul96a]. same [SLK91]. Sampling [BS94, Hin92, Klu92, MPSM93, Ano98q]. Sarafyan [Tho90]. satellite [De 99]. satisfied [RZG95a, ZRG93]. satisfying [EM94b]. Scalable [AN95]. scale [BFG96, Cou99, GR95, KKS96b, Lit95, dDGZ99]. scaling [AC95, CK91]. SCAN [Ano91-32, Ano97-51]. SCAN-1991 [Ano91-32]. SCAN-97 [Ano97-51]. scatter [Weg98]. Scattered [HMS96, ANS96, ADT95, FI96, LST96, Lev99, PdST96]. scattering [Ano97m, AD92b, CM92, CKM97, DOAG91, Grü92, HKW99, Isa92, KB91, Ked93, KZ92, Kre95b, Kre11, Mön96a, PRS90, WMC97]. schedules [BBH90]. Scheiber [Anh99]. scheme [BH95, Car93, GPCGV91, Jia91, Kha94, Lin99, Liu95, Sak99, SdO99, Sha94b, SDL99, SS91b, Sty92, Sty96, Sun96, SZ98, Zha95]. schemes [AT99, BW91, CV93, DP97, Ehr96b, EIM91, GR97, Gje97, GPMR96, KvD99, Lay93, LM92, MM97, MR92b, MS98b, MJ96, MJG96a, Moh98, PdA99, Roo90a, Roo94, SV92a, SWW96, VM99, ZCG99, dMM99, Ver96]. Schlömilch [Mil97b]. Schlömilch-type [Mil97b]. Schmidt [Stu95]. Schoenberg [dBS99]. Schoenberg-type [dBS99]. Schönhage [Kra96]. Schreiber [Eme94e]. Schröder [Ise93]. Schrödinger [AS96, BM93a, IDV98, Sim90, Sim92, SR92, Sim95, SW97, Sim98, Sim99, TSM96, TS97, XZ98, ZT98]. Schur [Ano96-34, BS93a]. Schur-concave [BS93a]. Schwarz [FZ93, Han91a, ST99]. SciCADE97 [Ano97-52]. Science [Ano90k, Ano91i, Ano93m, Bul90b, Bul94c, Van91b, Ver94b]. Sciences [Die94a, vVBZ93]. Scientific [Ano91-32, Ano97a, Ano97-51, Ano97-52, Car96, Eme94e, Ano93i, Ano97-42]. SCPACK [Li92]. SCV057SCV [Ano96l]. SCV057SCV19 [Ano96l]. SCV164S [Ano97p]. SCV164SCV [Ano97p]. SCV164SCV12 [Ano97p]. SDIRK [FGR97, FG97]. search [Koe93, LR96a, SB94, Was95, YZY96]. searches [Joe96]. searching [Lyn92]. secant [Mar95c]. Second [ÁN98, Ano91-33, Ano97-30, Ano97-36, AÁNMP98, Dic91, Lew96b, Ver96, Abd90, Abd96a, AE91, APH94, Ano97-40, Ari95, BPP97, BMS98, BCRV91, BY96a, DP93b, EK94, Ext98a, EH98, Fit91, FLM92, Fra95a, GHMZ94, GH97b, Han91b, IKK91, Jód90, Jo’90, JA92, JN92, KT97a, KKL97, Kla94, LOH92, LN96, MP95, Mas93, MO96a, Med95, MXZ97, MJG96a, MJG96b, MJG98, Oku93, Peh95, Pfe91, PZ98, RA97, RZQ99, SK96, SB99, SF92b, Sim92, Sim94, SDS94, SV92b, Sun96, Tre92a, Van91a, VV98, WMC97, WL92, YMW98, ZLQ96, ZCG99, vN93]. second-grade [SB99]. second-kind [AE91, BY96a, MXZ97]. Second-order [AÁNMP98, Dic91, Lew96b, APH94, BPP97, EK94, Fit91, FLM92, Fra95a, GHMZ94, Han91b, Jód90, Jo’90, JA92, JN92, KT97a, KKL97, Kla94, LOH92, Med95, MJG98, Oku93, Pfe91, RA97, SK96, SF92b, Sim92, Sim94, SDS94, Sun96, Tre92a, WL92, ZLQ96, ZCG99, vN93]. Section [Han91c, New91]. secular [Mel97, VBK91]. selection [SW95b, Tem95]. Self [HMV98, BCG98, BS94, DM94, Kam91]. self-adjoint [BS94, DM94]. Self-similar [HMV98, BCG98, Kam91]. Semi [Mon96b, Arn92, BM98, BY99, Cha99, CNS97, DLP98, FHR98, Kla94, LYFW95, Mag95a, Ned94, Ree94, RBDM90, SV95a, Ste90b, Wic90]. semi-classical [BM98, FHR98, Mag95a, RBDM90, SV95a]. semi-coercive [Ned94]. semi-conductor [Wic90]. semi-explicit [DLP98]. semi-implicit [Ste90b]. semi-infinite [Cha99, Kla94, LYFW95, Ree94]. semi-iterative [CNS97]. semi-logistic [Arn92]. Semi-orthogonal [Mon96b]. semi-simple [BY99]. semicircle [Mil93]. Semiclassical [Ext91, CS97a, MR95a, RBM91, ZDR90]. semiconductor [FLM92, Hem90a, LM92]. Semiconvergence [Son99]. semidiscrete [SS91b]. semigroups [ElM94, Rha92]. semilinear [MS98a, Smi96]. semilocal [Gut97]. Semiorthogonal [ACM98]. semismooth [KM99]. Sensitivity [Art94]. separable [BF93, Fra94]. separate [Lor93]. separated [ASWO91]. separately [Thr90]. separation [Han91b, Jo’90]. sequence [BCCM91, BP91, Bre98, Fan95, Osa96, PM92, vD95]. Sequences [BW91, BGN^{+}90, BGM93b, BGM94, Ben94a, Ben96, Ben94b, CGV97, Cop94, DEHG92, Drm90, EHT90, Fdi99, GB90a, Gro91b, GR92d, JMMW90, Kam91, KT97b, LMJ93, Osa90, Osa91, Pag92, Pré94, RT91, Sab95, Sad90, Sed90, Str95, Sza95, Thu99, Van93d, Yey92]. SERC [Ano91-34]. Series [Ano94j, Ano98q, Van91b, VR91, Aco97b, AP98a, AD97, AD92a, And95, BG95, Boe91, BB93, Bul91, Cha98a, CDTV99, DDL93, De 91a, DEP97, DDR93, DJ98, Eme94e, GS91a, GP97, Gil99b, GVN90a, GS91b, GK96c, KO99, KO01, KM94a, KM95, Klu92, Koo98, Kri92, Le 93, Lóp98, LM90, Mal90, MO96a, Mil97b, OO98, PGGC98, PRS92, RZQ99, SRD98, SL97, Tok96a, Tre91, Uyt96, VPR97, Wen90, TE97]. SERM [AP96]. Set [Ver94b, BC91, KV99]. Sets [Ver94b, BW96, DE90, Koe93, Lor95, Mag95b, MH94, Pel94, Ury94]. setting [ACQ94]. seventh [Sto97]. Several [Lóp98, BPV93, CL98, GS91b, HJ95b, LR92a, Lu91, Moh90, Nie95, Peh93, Van93a, Zhu97]. sextic [Wit95]. Shadowing [CKP94]. Shafarevich [Ver96]. shallow [ADH98a, BR98]. Shape [Lah90, Ad94, BR92b, CMR99, Lah92, Sac96, SU99]. shape-preserving [CMR99, Lah92]. shared [KNST91]. shared-memory [KNST91]. Sharp [MR95b, AB96, BELN92]. Shaw [MR92b, PR93]. shearing [Els99]. sheet [CES91]. Sheffer [Sch98]. Sheldon [Ano94l, Eme94f]. shell [Ber94, Kre95a, SR91]. shells [BS98b, Zha91]. shift [NKS98]. Shifted [Let98, All96]. shifts [PRS90]. Shishkin [Lin99]. Shishkin-type [Lin99]. shock [Faz90]. shocks [IB99]. shooting [Jon93, LB98, Li99b, Mar91b, Pry94]. shooting-type [Pry94]. Short [Ano91-35]. shortcut [Deh94]. Si [PCC95]. side [Lim94, Ozi92]. sided [Tag98, ZT98]. sides [Jbi99, Mus99, SG96]. Sideways [Reg95]. sign [Gra96]. signal [Uyt96]. Signorini [AP98b]. Siljak [Sto95]. similar [BCG98, HMV98, Kam91]. similarity [Faz90]. Simple [SL97, Ata94, BK95a, BY99, Bus99, Epp92, Gro91a, HvdV94, KZ94, MS98b, RZQ99, SF91, SFGF^{+}99, SW95b, WGK99, dS97b]. simplex [Nea96, Spa99]. simplices [Wen97]. simplifications [HZ95]. simplified [LB90]. simply [Weg96]. simulated [TPA95]. Simulating [Cou99]. Simulation [Ano91o, Bul90a, CFBT99, Ama94, AD92a, Atk90, BGN^{+}90, CC99, DP97, EBV^{+}99, Eme94c, GH97a, Hei95b, Hem90a, JDV95, Kor96, NS95, Nie94, PC96, RK90, SD98, Sch97, Syd99, TOCH99]. simulations [Alv99, CL97, Jer95, MPS96, Spu99, SB96, Wam99, XM98]. Simultaneous [CS95a, Gel95, VMV94, Ata94, AH95, Bel90a, BC94, CP93, CDM91, DDL93, KR96, Mar91a, Rán92, STS91, SP96]. sinc [BK99b, Ste97]. since [CR93, Coo99]. Sine [EIM91, KL90]. Sine-Gordon [EIM91, KL90]. single [KM94a]. single-output [KM94a]. Singular [HY98, Wei97, AO98, AE90, Alp95, And90, ASW97, BAG^{+}96, Bar97, Bec90, Ber92b, BR92a, CT99, CS93b, CNS97, DS90, DW94, Die98, Die02, DM97a, EGBE93, EK94, GBP97, Gov94, Jun94, JL97, KL98, Lad97, LC99b, Moh97, Moh98, PB92, Pry94, QA97, Rab94, SH96, Son99, Van90a, Ven92, Ven93, Waz91, ZL90, ma92]. singular-point [Pry94]. singularities [CL91, HB92a, Hug94, Ill98, Lew91, LB90, LG95, LN99, MS99b, RS92, SE99]. singularity [GPS92, Kui95a, Li97, LLY99, Ver97]. singularly [Bog98, Che97a, CLM95, Fit91, GPCGV91, Hem96, QS99, Roo90a, Sin92, ST99, Wan97]. sinusoidal [Loe96]. six [OS94]. Sixth [Ano93-28]. size [HJ95a, HHL97, HC99, LMP97, Mau98, Oku93, SW95b]. Skorohod [Tak92]. slab [Gro91a, HS92c]. slackened [GJ95, GK95]. slender [DE93]. Sloan [Fra95b]. slow [Le 98]. slowly [BEM98]. sluice [eM91]. Smale [DF95]. small [Bur93, JMT94]. Smith [Van91b]. smooth [ADT95, Ata94, MW95, SS94a, Shu96, TRBS99, VVD94, Zha92]. Smoothed [RL99]. Smoothing [BKF^{+}96, Gje97, Jbi99, Par92, BC93, Che97b, DR94, FS96b, Zha98]. Smoothness [Dam99, Dam98, DNZ98, GW91, Wen99]. snul [Mag95b]. Sobolev [ALR96, AÁNMP98, BZ93, BM92, BC98a, BC99, De 93, DE97, DE99, ELW93, HKL96, JKLL95, JKY97, MAR93a, MPP96, MMPP97, MMFMB97, MFMBPC97, MF98, Mei93, MP99b, MBMF99, Pan97, PP93, WK98, dMPP98]. Sobolev-type [ALR96, AÁNMP98, JKY97, PP93]. Society [Ano98q]. Softcover [Ver96].

Software [Loz96, AV96, Eme94a]. Sohn [Eme94c]. solar [Kos99]. solid [JSWW95, PCC95]. solid-state [PCC95]. solute [TH94]. Solution [AU96, AEG98, Ano97-46, BK95b, Bul90b, Cas96, CP96, DOAG91, Gov94, GM90b, GB90b, eM91, Abd90, Abd98, AP98b, AM96, AM97, AB96, AS96, AN95, BOS97, Bak99, BBD^{+}96, BK95a, Bec90, Ben97, Bog97, Bru96, Buo96, BC91, BELN92, CL99, CMR95, CMGP99, Car93, CS93b, Cel98, Cha99, Chr90, CPK96, DBFV96, Dem95, DM97a, DPTW95, Dos95, DK94, Elb92, EK96, FFRV92, FFLM95, GK97, GP93, GPCGV91, Han94a, HS92b, HB99, HJ95b, Hug94, Hus94, JB99, JC95, Jód91b, JN91, JA92, JS97b, JL97, KNST91, Kit91, KZ92, Kre95b, Kre11, Kri92, KBÖ97, KBÖ99, Le 98, Liu99, MP92, Mas99, Mau98, MM97, MJG98, Mön96a, Moo95, MS92, ND95, Ozi91a, Ozi91b, PRT97, RIG98, RE94, Ren97b, RS90, Rov96, Rum95, Sak99, Sca94]. solution [SD98, SB99, Sim96, Sim90, SR92, SW97, Sim98, Sim99, SS97a, Slo91, Ste99, Usm92, VVD94, VM99, Wei96, Wri93]. Solutions [Ano91z, SA97, AS99, AG98, AW98, AO98, Alw96, AAP98, And92, BAG^{+}96, BBM96, BM96a, BPP97, BMS98, BS98a, BK90, BJ99, BDKM96, BY96a, CNP99, CFMR96, CM97, CLL98, CK92, DH94, Dib98, DKKV98, DP93b, DM97b, EGBE93, Elo98, Fas95, Faz90, Fie90, FB96, FGR97, FS95, FT98, GPS92, GBP97, GRZA98a, GK96b, Gro91b, Han91b, Han91d, Han92, Hei95a, HLCW93, Ish91, JL92, Jo’90, Jód91a, Jód92, JNC99, KSFAH98, KKL98, KT92, KLL97, KY97, LP94, LD98, Leó99, LP98, Lub93, LM94c, Mar97, MLD96, MS98a, McC93, Med95, Med96, Med98a, MS98b, MM91, NW94, NYW98, Njå95, Ois95, Ozi92, Pao98, PR93, PB98, Plu95, Pop96, Rab94, Rov92, ST97, TE97, Tre92a, Tre92b, TB97, Val96, Voe95, Waz91, Wer95, Wim91, YC90, YN95, YMW98]. solutions [ZD94, ZLQ96, ZCG99, ZZ91, ZW97, ZL90]. solvability [BGH98, GM90a]. solvable [Sch94c, Sch94d]. solve [CZ90, DR90, DK92]. solved [HW99]. solver [De 90b, GC97, HB92a, KNST91, SSV98, dS97b]. solvers [Epp92, Hig93, IM99, Jin96b, LR93, MdSvdV97, vdHM97, vdHS98, BVZ90]. solves [Fis91]. Solving [JN92, Lau91, MV92, Nak91, SY90, A.94, AH98, AD92b, Arg92b, BAG^{+}96, BHS92, CS93a, CY94, CM92, DHP94, DLP98, FT96, Han91b, HZL94, IY97, Jo’90, JM96, LLP97, Lem94, Li92, LYFW95, Mar91a, MQ95, Mel97, Mes96a, Mes96b, Mou93, Mus99, PH99, QA97, SK96, Sto95, TPA95, Weg90a]. Some [ALR96, AKV92, Arg92b, BFG96, BGM95, Bav98c, BE91, CV93, CADMMB94, Dar98, DJ98, Ext97b, GKH91, GS91b, GHH99, JS91, KS96b, KK97, Kuc99, Let99, Lev94, Mar96, Mas95, MO96b, MDR98, NNR93, Rad91, Rad95, Rah96, Ree94, Sez92, SF92b, Tim93, WK98, Wit95, vD95, APH94, Arg91, AL97b, Ben94a, BV95, Ber98, Car96, lCHcH97a, CS96b, De 91a, DT93, DE99, DK94, ES97, EW99, FP97a, GPRV93, HZ95, IB99, JB96, Kha94, KPP97, LC98, Mat98, MdCVR99, Ols99, Pal98, Par97, Pet90, Pré94, RV94, Sab95, SP96, Sed90, Sen99, ST99, Ste95, Swa96, Val96, Ven96, Wen90, Wim90, Wim91, WGK99, ZL96]. Sonderforschungsbereich [Ano91r, Ano93u]. Sonine [ME94]. SOR [Che98b, HNT96, Ish99, Nou93, PK97, SSY98, SOMF91, Wu90, Yam97]. SOR-like [SOMF91]. sound [MET93, Mön96a]. sound-hard [Mön96a]. source [GKH91, HS92c, Let95b, MLD96, OO94, SdO99, Tim93]. space [ACQ94, BF93, BW96, Car93, FR99, Gea91, HS96, MMST96, Mei93, MJ96, MJG96a, MJG96b, Moh97, Moh98, Nak91, Pag98, RS90, Yan92, Ver96]. space-time [HS96]. spaces [Aco97a, Aco97c, Alp98, BM92, BC94, De 93, DE99, ELW93, HKL96, JL97, Lad97, MAR93a, Min98, Mit91, Voi95, Yan90, Bul91]. span [ATBB99, SJAM98]. Spanish [Ano91-36]. spanning [PP96a]. sparse [BEC99, Fuk90, HL95, LOH92, SSF95]. spatial [Kre95a, Mon93]. Special [Ano93w, Mag95b, DTB93, Ext91, Ext95, JB96, KK97, Lem94, Loz96, MP93, RW99, Sci94, Sim92, SDS94, Zei90, vVBZ93]. spectra [Böc96, BEM97a, Sza95]. Spectral [Ano93-29, AZD91, BGM99, BEM97a, BEM97b, Fas95, Jek91, MOT91, Mel98, ZD94, BEM98, CDF95, CPK96, DW94, DP95, EK96, Eln97, Fra99, HHL97, KP90, Kar94, OP94, XM98, Zhe97]. spectral-finite [CDF95]. spectrum [CNS97, Esp99, GHMZ94, JN92]. speed [SDP97]. SPH [SR99]. SPH-code [SR99]. sphere [ANS96, DS98, Hal92, KN90, LS96, Spa99, Tic90, TOCH99]. sphere-like [ANS96, LS96]. spheres [HX99]. spherical [ANS96, Bar90]. spherically [ZT98]. Spin [Ano91t]. Spiral [MW99b, ATBB99, WM96a, WM96b, Yam97]. Spline [Ano91y, BY96b, DS90, FP99, Ors90, ST97, ADT92, BS93b, DS91, DL96, DNZ98, Die98, Die02, ES90, HS94b, Iqb97, Lah92, Lah96, Lai96, Loe96, MW99b, Ors97, Pfe91, PWT99, Roo90a, SK96, Voe95, XGSSKF96, ZM97]. spline-based [Die98, Die02]. spline-functions [Roo90a]. spline-on-spline [XGSSKF96]. Splines [SWY98, ANS96, BWS99, CMR95, CMR99, DPS98, Dem95, Jun92, Jüt97, KS97, Kob95, Lah90, MR95b, MW95, Nea96, NV94, Rab90, Rab96, Sab96, San96, SU99, Str94, Usm92, VVD94, Vol90, WM94, WD94, dS97a, KS99, Die94a]. split [Fra95c]. splitting [DK92, LV99, LB98, Li99b, LR96a]. splitting-shooting [LB98]. splitting-shooting-integrating [Li99b]. splitting-up [DK92]. splittings [Mie98]. Springer [Bul94b, Die94a, Lev90a, Lev91, Pie96, VR91, Ver96]. Springer-Verlag [Die94a, Lev90a, VR91, Ver96]. Spurious [Sta98]. square [AAVB98, Yey92]. squared [Chr94, FSV96]. squares [BP96b, BM93b, FFH97, JS97b, Mon96b, SF91, SSY98, Sci94, VB94, Van90b, YL97, Yua96, YI96]. SSOPAS [Ano91-36]. SSOR [HN90, HN95, HNT98]. stabilised [Ram99]. Stability [CS97b, CRV90, CRvdHV93, Han91d, HLFC99, Isa92, LC97, MT98, Med97, Med98b, MM91, Owr95, Ren97b, SV92a, XC96, BS95, Ber92b, BT96, Cah90, CMR93, CS97c, Chr90, DSH95, DJR91, FB96, Gem96, HJ95b, IO91, Kit91, Kul92, LMM94, LC99a, Lu91, LR96c, MR96, MS99a, Mas99, MJG98, SZ97, Sid99, Ste90a, Ste90b, Vog94, Zin99]. stability-induced [Ste90a]. Stabilized [AP96, ZZ97, Axe96]. stabilizes [SW95a]. Stable [Kla94, Lev99, Yam97, AS96, Bar99b, BMX97, CAZ98, FLM92, FP90, HH90, HHL97, HC99, IP99, Pan96c, RW98, VVVD99b, Vec96, Vol90]. stage [BH98, BEC99, CMGP99, CAZ98, FSW94, VVVD99b, dS97b]. staggered [Paq94]. staging [DR94, DR95]. stagnation [Ari95]. staircases [Bec92]. standard [Mar91b]. Star [Kle99]. Starting [Lab97, Lab98, RH99, Zha94]. State [Eln97, CK91, PCC95, RS94, Swa97, WB94]. State-control [Eln97]. state-dependent [WB94]. states [HH90, HF95, UH98, ZT98]. stationary [DGH95, GR92d, LM92, MW90b, TG91]. statistic [Drm90]. Statistical [Nie90, Zel95]. statistics [SS96a]. steady [eMB91, Bab98, Dic91, DR94, DR95, RD96, SB99, Swa97, WS97]. steady-state [Swa97]. Steele [Eme94e]. steepest [SLK91]. Stefan [ADT96, ADT97, KBÖ97, ND95, Ned97a, Ned97b]. Stefan-like [Ned97a, Ned97b]. Steffensen [Nie95]. stellar [HB99, RL99, WD99]. Step [Mau98, BCP99, Con99, DLP98, FM94, FM96, FM97, HS92a, HHL97, HC99, IR97, KT97a, Lop90, LMP97, Oku93, SW95b, Sim90, Sim92, Sim94, Sim95, Sza93, Tra96, XZ98]. step-size [HHL97, HC99, LMP97]. steplength [EM98b]. stepping [WF97, vdVNTV97]. steps [MTW96, vdHSvdV95]. Stepsize [Tho90, WB94, BVZ90, CMR93, Con99, FS90, FK95, HU99, VS90, Zha94]. Sterling [Ano94k, Ano94l, Ano94m]. Stieltjes [Bul94b, BHJ99, BGVO95, BDMGVO97a, BDMGVO97b, Ehr95b, EM98a, GM93, GVN90a, Grü98, HRGL98, IS95a, KW90, Lor95, Njå95, Njå98, Peh95, RM95b, Tok96a, Tok96b, Val95, VV95]. Stiff [HW99, AFK93, FGR97, FG97, PRT97, Sha93, Som93, hC94a, vdHS98]. Stiffness [Spi96]. still [Zit94]. Stirling [Ada97, Köl94a, Zen95]. Stochastic [BT95, Art94, CQW95, Dup94, Fra94, GY95, Gel95, GM99, Kan94, Lep94, LL99a, Mau98, NR90, SD98, Sch98, Sch94e, VBK91, Vog94, ZFV98]. stochastics [CADMMB94]. stock [KKS96b]. Stoer [VR91]. Stokes [Bän91, BP96b, lCHcH97a, CL91, CF93, Heu92, HM97, Joh98, NYW98, Ols99, OP94, Par92, SDP97, TD97, TE97, TLL91, WS97]. Stone [Pro94]. stopped [Ham94]. Störmer [Fra95c, LMF99]. straight [Fra95c]. strategies [HU99]. strategy [Bän91, BES91]. Stratified [Pat99, ADH98a, MEES95, MPSM93, TOCH99]. streamline [HBL96]. Street [Ano94k, Ano94m, Eme94a, Eme94d, Eme94e]. strength [DNN92]. Stress [DNN92, ZZ97]. Stress-strength [DNN92]. stretched [Nou93]. Strong [BG98, BC94, IS98, MFMBPC97, Pan93c, Sch94e, dlCYL98, BHJ99, BMR99, CM96a, CG97, DMGVO99, GH99b, JN99, Kui95a, Njå95, Njå98, dO90]. strongly [Die98, Die02, GHMZ94, JN91, JNC99, TB97]. Stroud [CR93, Coo99]. structural [SR91]. structure [Alp98, BGM93a, Bec90, Cou99, EHGNT90, FSW94, Gil99a, JM99, KM99, Mai93]. structured [Cus94, HS92b]. structures [Pro95, PCV99, Rös95b, vdFvGdRdV95]. studies [Hom96]. Study [Ano91y, AD92a, ADT92, BT92b, CS97d, Dub96, Fen94, GMS95, JV96, JMT94, RZG95a, Smi91, ZRG93, Zha91, ZL99]. Sturm [Cha98a, DW94, MP92, Pry94, VD91, VD94]. subdivision [BES91, FPS99, KvD99, SDL99, SE99]. Subject [Ano98-34, Ano98-35, HMS96]. submerged [DS98]. submultiplicative [AR99]. subroutine [BK99b]. subsequences [IL99]. Subsets [CS95b, Osa90]. subspace [BMX97, Kin90, SL96]. subsurface [ABD^{+}96b]. successive [Anh99, Ern96, HS92a, Ish91, Ozi92, YZY96]. suction [Ari95]. sufficient [CL98, MB92]. suited [SOMF91]. sum [Bav98b, CDTV99, GSSV92, LGR96, RBM91, ZDR90]. summability [BT92a, Kor94b, RW92]. summary [SA97]. Summation [Ext98c, Ext97b, Ext98a, Lew97a, VPR97, Wen90]. summations [Mil97a]. summed [Fra95c]. Summing [OO98]. sums [Ada97, BPV93, GlZ94, Ham94, KvD95, Lan98, LM95, Rin94, RM95b, Van93d]. Super [HAMeM96]. Super-critical [HAMeM96]. superalgebras [HV91]. supercomputer [FFM95]. supercomputing [Ano94l]. superconductivity [Ext91, Ext95]. Superconvergence [KNP97, BY96a, LLY99, Shu96, LC99a]. Superior [HN95]. Superlinear [Che97b, AO98, BS98a, vdVV93]. supernovae [MM99c]. superposed [AR97]. superposition [Nee94]. superpositions [Pal92]. supersonic [De 99]. supplementary [McN97, McN99]. support [CS95b, CMR99, MMFMB97, YA95]. supported [AK91, FI96]. Suppression [dMM99]. Surface [Jüt97, WD94, BN92, CFBT99, Han94b, HAMeM96, HZ96, Maa94, Spa99, TBF99]. Surfaces [Ano91f, ANS96, CP99, Li99c, LC99a, LS96, SDL99]. survey [AL97b, BB98a, Jin96b, JN99, Mee96, SV92c]. survival [BS93a]. suspended [JM94]. suspension [FSZ94, Sim96]. Switching [BGM96a, BGM96b]. Switching-Method [BGM96a, BGM96b]. Sylvester [DSH95]. Symbolic [Ano97-53, ACD92, BY99, Zha96b]. symbols [ÁNQR99]. Symm [ES96, HLCW93]. Symmetric [AT99, BS93b, Par91, VC94, Bar99b, BK99a, CR98, CLL98, CM96a, Dam99, Die93, ELW95, GK96a, He94, HL95, Jin96b, LLP97, Lui98, MFI95, Mel98, Rab96, Sha94a, SSF95, VB93, Wit96c, Yun98, ZT98, hC94b]. symmetrically [MMFMB97]. Symmetrizability [KY97]. symmetrization [GS97]. symmetry [AT99, ABM96, BMR99, DH94]. Symposium [Ano91-32, Ano91-36, Ano93w, Ano97-39, Ano97-41, Ano97-51, Ano91x, Ano93y]. synchrotron [PB98]. system [ADH98b, BMP96, BL93, BC91, CY91, Fis91, Gil90, HJ95a, HMV98, How90, Hwa96, Ish91, Jer91, JM96, KR96, Kri92, Leó99, Leo98a, LWZ98, LR93, MT98, MLD96, MJ96, MJG96a, MJG96b, MP93, NKS98, SGFR94, Sty96, Wit96a, Wit96c, Zho94, vdHM97]. Systems [Ano91t, AG98, AT99, AR92a, ATV98, And92, ACD92, BH98, BEC99, BBD^{+}96, BK90, BGM96a, BGM96b, BK95b, BJ99, Bre97, Bre98, Bru96, CR98, Car93, CRZR96, Cel98, CNS97, DW94, DR90, DHP94, DL99, Fay97, Faz90, FR95b, Gab92, GG97a, GPS92, GJ95, GK95, GP93, Gov94, Han91b, HLS94, Hei95a, HC99, Jbi99, JC95, Jin96a, Jin96b, Jód90, Jo’90, Jód91a, Jód91b, JA92, JN92, JNC99, KT92, KKNU97, KM99, Krz96, Kul92, LLP97, LMM94, Lap93, Li92, LR96a, Lis95, LP92, Lop90, LMP97, Lu91, LM94c, Med98b, Mes96a, Mes96b, MDR98, Müh96, Mus99, Oku93, Pan93b, Pan93c, PC96, PP96b, PCC95, Rab94, RM95a, Rum95, SJAM98, Say97, SS90, SFGF^{+}99, Seg95, SY90, SG96, Sin92, Son99, VB97, VJ96, VB93, Xu99, YC90, Zei90]. Systems [Ano91-34, Ano93s, Bul96c]. systolic [Mie91]. Szego [AGR93, GVN90b, JR93, JNS90, JTNW93, Not90, Pan93a, Pan95a]. Szekeres [LD99].

T [Bul94b, Eme94c, Pie96, SS91b]. T.-H [Eme94c]. table [BGM93a, Bec90, Hoc97, Pas92]. tableau [RZAG98]. tables [Kra96]. tail [DH95]. Takens [JP91]. talks [Ano91v, Ano91x, Ano93y, Ano95s, Ano95w, Ano95x]. tallest [McC99]. tanh [AE90]. Tau [SS97a]. Tauberian [Moe95, Odl92]. Taylor [Eme94d, KO01, KO99, Kri92, RZQ99]. Taylor-series [RZQ99]. Tchebychev [AP98a]. technique [Ano97-40, Iqb95, Kar94, KV95, Rad91, Wit96b]. Techniques [Ano91g, Ell94, CZ90, Cho92, Dar98, Drm90, KP90, LMP97, Pie96, SFGF^{+}99, Ste95, Ver94a, Wit95, Zha98, Ano97-36]. Technische [Ano91s]. Technologies [Ano97-47]. technology [HvdV94, Ano93j]. Tekalp [Uyt96]. temperature [CSS93]. Ten [Roo94]. tension [Han94b, MR95b]. tensor [Jüt97, KM96]. tensor-product [Jüt97]. term [Ekh91, Ext99, GL90a, Lóp99, MLD96, MTW96, Mil06, Sch96b]. term-by-term [Lóp99]. terms [AS97, CY94, Epp91, HJ95b, Köl94a, Lu91, SdO99]. ternary [VV98]. test [SR99]. testing [AH98, Lev90a]. tests [Atk90, Gab92]. tetrahedral [LM92]. textbooks [Bul91]. texts [Lev90a]. th [DM97b, Dub98, Gri99, LM95, TB99]. Their [Ano93-28, ÁNQR99, Ano91-36, CH98, Erd93, Fre92, JL97, NV94, RZG95a, SDL99, SF90, Sim92, Sta93, XZ98, ZRG93]. theme [Grü98]. theorem [Alp98, Bla95, BGVHN95b, Cuy90, Cuy92, Ext98a, GMS91, GMBW96, GH96, Gut97, IS95a, Kal95a, Klu92, LGR96, Lor92, Nee94, Pro94, RW92, Waa91, YC90, Zhe93]. theorems [BGM95, BT92a, BS94, CDL96a, Dam99, FRR94, GR92a, GlZ94, Mar96, Pas90, Sid90, Voi95]. theoretic [FAD99, Joe96, Mas92a, Mas92b]. Theoretical [Bul94c, LWZ98, RF99b]. theories [DZY^{+}94, Kal95b]. Theory [Ano91y, Ano97n, Ano97-40, Bul91, Bul96b, Bul96c, Van91b, ADG99, Aco97c, AFK93, BHD93, Ber95, Bul90a, Bul96a, But92b, CLB92, DW94, DF95, FV96, GGF99, GMBW96, Jer95, JN99, Kre95b, Kre11, Lor95, Mar95c, Ned97b, Pan95a, PS94, RS92, SV92a, WMC97, Wit96c, Eme94f]. there [Weg98]. thermal [Moo95]. thermistor [ZW97]. thermo [LWZ98, Ned95a, ND95, Ned95b, Ned97a, Ned97b]. thermo-Bingham [Ned95a]. thermo-elastic [LWZ98]. thermo-elasticity [ND95]. thermo-magnetodynamic [Ned95b]. thermo-plasticity [Ned97b]. thermo-visco-plastic [Ned97a]. thermoelasticity [Jia91, Ned94, RIG98]. thermostatic [Jek91]. theses [Ano90k, Ano91r, Ano91s, Ano93u, Ano94w, Ano94x, Ano96z, Ano97-37]. Thiele [Chu97a]. Thiele-type [Chu97a]. thin [AK91, Rab96, SR91]. Third [Ano93-30, GH97b, CL99, Die95, DPTW95, Hus94, Mas93, Not97, Rov96, Sca94]. third- [DPTW95]. Third-order [GH97b, Hus94, Rov96, Sca94]. Thoughts [Rip90]. Three [AAP98, De 99, Ell95, Grü92, XK92, Bän91, BH91, BB93, Chr98, EIM91, FG97, Goo91, IG90, Köl94b, Kos94, KZ92, Krí95, LDV91, Moh98, Osa94, Sab96, Sch94b, TLL91, WS97, Wit95, Wit96a, Wit96c, YO98]. three-body [Krí95]. Three-dimensional [Grü92, XK92, Bän91, BH91, EIM91, Goo91, IG90, KZ92, LDV91, TLL91, WS97, Wit95, Wit96a, Wit96c, YO98]. three-direction [Sab96]. three-parallel [FG97]. three-processor [FG97]. tilings [FG99]. Time [BJZ93, KB91, AT99, AR92a, AD92a, And95, Bän91, CDF95, CFR96, CRV97, FLM92, GG97a, Gea91, Gel95, HS92c, HS96, Ked93, KM94a, KM95, Lap93, Mal90, MTW96, MJG98, Ned97b, Tre91, XM98, vdVNTV97, Bul96c]. Time-dependent [KB91, AR92a, Bän91, XM98]. time-harmonic [Ked93]. Time-point [BJZ93, GG97a]. time-series [And95, Mal90]. time-varying [Lap93, Bul96c]. Times [Ano94l, Eme94f, Yan95]. Timoshenko [lCHcH97b]. Titchmarsh [WGK99, Yak97]. Toda [ABM97]. Toda-type [ABM97]. Toeplitz [BW96, CT99, DS99, Gra93, Hoc97, Jin96a, Jin96b, Mel98, SS90]. tolerance [Hig93]. tomography [KvdB92, Kos99, Rat94]. tool [De 93, GS99, Rol99, SS99a]. Topics [Ver94b, Ano98p, Ano91z, Bul96b]. TOR [Cha96]. torsional [JM94]. Total [Rat94, CGEP96, Van90b, VDM96]. trailing [FFRV92]. Transcendental [Ano97-53]. transfer [DZY^{+}94, FP97a, Moo95]. transform [AZ97, Bar90, DMGVO99, DLP98, Die99, Ehr95a, GS91a, Iqb95, KS96a, Mil97c, Sin94, Tag98, Tri93]. Transformation [Hac97, VPR97, AE90, ABM97, ARS96, Atk90, CRD94, Faz94, Gri99, GS91b, Hom96, Ito94, Sad90, SG94, Sid97, Sid99, TA95]. transformations [Bal93, BP91, BW91, Bre98, Ext98b, Fan95, GK96c, JS97a, Lau96, LB98, Li99b, MS93, MN96, Osa96, Sen99, Yak97, Zhe97]. transformed [BI92, SF92a]. transforms [Ano96y, AVW99, CZ95, DS98, HT94, Iqb97, Klu92, Mac99, Mas93, OSG^{+}93, Pag92, PB92, Pon98b, Sab95]. transient [Ozi91b, SD98]. transition [DV99]. transitional [Sch99]. translated [Ver96]. transmission [BGH98, Boe91, KB91]. transmission-line [Boe91]. transport [ABD^{+}96b, DZY^{+}94, KK91, Lay93, MM99c, TH94, WF97]. transportation [Abd98]. transporters [BB95]. transshipment [HR94]. Transverse [RZAG98, MM97]. trapezoidal [eMM97a, GL94, HAMeM96, SL98, SS97b]. trapping [Alv99]. traveling [CM97]. treating [JSWW95, SWW96]. treatment [Ext91, GLP98, GL90b, SGFR94, Wri93]. Treatments [Ano96l]. tree [PP91, Wam99]. trees [Mah92, Pal92]. trend [BM93a]. Trends [Mie91, Eme94b, Wim91]. trial [AKM97]. trial-free-boundary [AKM97]. triangles [CRD94, Say98]. triangular [BW91, CM96c, HGL96]. triangulations [DLR92, Lai96]. tridiagonal [EFL97, Yun98]. Trigonometric [Rös95b, Cal97, Cap90, Köl94b, VD94]. triple [AE90]. triples [EH95]. truck [SGFR94]. truncated [CG97]. Truncation [Bal93, BJ93, Thr99]. trust [HZ96, McC98]. Tschirnhausen [MW97]. Tsertsvadze [SZ98]. Tsuzuki [SZ98]. tubular [Faz92, MH94]. tuple [DDL93]. tuples [Ren99]. Turán [GM97, Spa99, Van93a]. Turán-type [GM97]. Turbulence [Ano93t]. turbulent [Eme94f, Kha94, SC95a]. Turing [Kra96]. turning [How90, Mor92]. twelfth [BT91, ST97]. twelfth-order [BT91, ST97]. twin [Waa91, WA98]. Two [DMGVO99, FS90, FG97, Köl95b, KL90, Le 95, MM99a, Mav95, Ols99, Tok96b, Ver94b, ZT98, ZL90, AR97, Abr91, Alv99, BH98, BEC99, Bav98a, BF93, BGVO95, BDMGVO97a, BDMGVO97b, CS93b, CAZ98, CM92, Con99, Deh99, Ehr95a, Ehr96a, EH94, EGBE93, EIM91, EEAES99, Ext99, Fdi99, GBP97, GVN90a, GVO94, GS99, GB90b, HS94b, HR99, Ira94, Jon93, KKL97, KvdB92, KKS96b, KS96a, Kor94a, Kos94, Kuc99, LW96b, LB99, Lit95, Liu97, Mil06, MJ96, MJG96a, Moh97, ND95, NS94, OO98, Ozi91a, Ozi92, QS99, QA97, RBM91, RZG95b, Sac96, SH96, Sim92, Sin92, SFB94, Sty96, Tag98, Tra96, VVVD99b, VVD92, XZ98, XM98, Yey92]. two- [EIM91, KvdB92]. Two-dimensional [ZL90, Abr91, Deh99, GS99, GB90b, HR99, Ira94, KKS96b, Kuc99, OO98, Ozi91a, XM98]. Two-level [Ols99, Lit95]. two-parameter [Ehr96a]. Two-point [DMGVO99, MM99a, Tok96b, BGVO95, BDMGVO97a, BDMGVO97b, CS93b, EGBE93, GVN90a, GVO94, Jon93, Liu97, QS99, QA97, SH96, Sin92, VVD92]. two-side [Ozi92]. Two-sided [ZT98, Tag98]. two-stage [BH98, BEC99, CAZ98, VVVD99b]. two-step [Con99, Sim92, Tra96, XZ98]. two-term [Ext99, Mil06]. Two-Volume [Ver94b]. Type [Van91b, ALR96, ÁNYD98, ABM97, ADMFSR98, AÁNMP98, Bai97, BS95, BBM96, Ber92b, BDV94, BB93, Bre90c, Buo96, BL93, CS97b, Chu97a, Cuy90, Cuy92, DVVV90, DVV90, DMGVO99, Erd93, ELW95, EH97b, FAD99, FPR92, FB96, Fra95a, FP97b, FP99, GM97, GP98, GLP98, GVO94, Gör99, Guo95b, HT91, HLS94, HRGL98, Heu92, Ioa90, JL92, JB99, JS93, Jhu95, JKY97, Kar94, KW90, Kui95c, Kui95b, Lai96, Le 95, Lin99, LM92, LN96, Mag95a, Mai93, Mat97, MTW96, MP99b, ME94, Mil97b, Min98, MB92, Mus99, Not90, OM99, PP93, Pry94, RV94, RW92, SL98, SF90, TSM96, Thr90, Thu99, VDV90, VVD92, Voi93, Weg90b, XC96, Yak97, YK96, ZZ96, ZD94, dBS99].

U [Ano90k, SS91b]. U.S [Roo90b]. UK [Ano98q]. ultra [Kan97]. ultra-hyperbolic [Kan97]. ultraspherical [Doh98, Ehr95b, För93, SS94b]. ultratriangular [RJ99]. unbiased [Rol99]. unbounded [BDMGVO97b, GM99, Rha92]. unconstrained [FM97, GV96]. unconventional [Ven93]. uncoupled [Jia91]. underdetermined [CY94, Mar91a]. Undergraduate [Lev90a]. underlying [Art94]. Unified [GLP98, FR99]. Uniform [Ano91-35, BS95, BG95, Boe91, CLM95, Die94b, EH91, Kul92, Mas92c, SLK91, SLK92, TD90, Tem95, Dun97, ES90, HS97, KW99, KS97, VM94]. uniformly [Drm90, Ehr94, Hac97, Roo90a, SL99]. uniforms [Sta90]. Unifying [Cha95]. unilateral [GK95, Lov95]. Unimodal [HV91]. Uniqueness [CLL98, BGM90, BSW99, EGBE93, Ozi91b, PGGC98]. unit [ACM98, BG95, BC98a, BC99, BHGVN94, BGVHN97, CS96a, CS97a, Ext99, Gra93, KN99, Li99a, MG91, Mig95, Mil06, Pan93b, Pan93c, Pan95a, Pan96a, Pan96b, SL99, SV95b, Tri93, VB94, Weg93b]. Unitary [Wat97, DL99]. univalent [MM99a, Ren99]. universally [Kor94b]. Universität [Ano91s, Ano91r, Ano93u]. University [Ano98q, Eme94b]. unknown [BMS98, BS98a, BY96b, FG92]. unstable [MR92b, PR93]. unsteady [Eme94c, Ols99, Sak99, TLL91]. unstructured [EBV^{+}99, RD96, WF97]. unsymmetric [Chr94]. Updated [BGM96a, McN99]. upon [VDV90]. Upper [EHG90, Jód90, LW96b, LS98, LP98, LU96]. upwind [AM95, FP99, Lin99]. US$308.50 [Ver94b]. USA [Pie96]. Use [Jon93, BCCM91, Dem95, FG91, FW99, Fre92, MJG96b, Sta93, Usm92, WMC97, Woo91, vdHS98]. used [LL99b, Owr95]. useful [Dar98, Joe96, LMSSD96]. user [Sas97]. uses [YZY96]. Using [DB99, Fau90, FM96, AR92a, ANS96, AKV92, ADT95, BWS99, BJ99, CMR95, CRD94, Cha98a, CS98, CW99, CK91, CMR99, CSvD98, CPK96, DOAG91, Fea91, FI96, FP99, GP93, Gra96, GM90b, HS92c, JC95, Jüt97, Lap93, LG95, Lyn92, MP92, Mic99, Par92, Pré96, Roo90a, Saa94, Sch94a, Sin92, SE99, Swa96, Swa97, Tho90, WM92, WD94, Wit96a, Wit96b, YN95, Ano98r]. utilizing [McC98]. Uzawa [Che98b].

valid [Ehr94]. Validated [Ano97-51]. value [AO98, AM96, AB97b, AAP98, ADT96, ADT97, Anh99, AFK93, BV99, BR98, BS93b, BF93, BK99b, BT96, CL99, CFMR96, CG94, CS93a, CS93b, Che97a, CLM95, Cri97, DS91, DL96, DPS98, Die94b, Die95, Die96, Die97, DTB93, EGBE93, Elo98, EK94, Ern96, Fea91, Fra95a, FG97, Gre93, Gro91b, HvdV94, HY98, Ioa90, Ira94, Jód90, Jód91a, Jon93, KT97a, Lah96, LW96a, Lim94, LG95, LC99b, LT93, LR93, Liu97, LN96, LP98, Lor95, Lov95, MS98a, Moh97, OF96, Ors90, PM92, PACH90, Plu95, QS99, QA97, Roo90a, SK96, San96, SH96, SF92b, SA97, ST97, SS92, Sim92, Sim94, SDS94, Sin92, SIC92, Spi96, Ste99, SS97d, TB97, TM92, Usm92, VVD94, Van90a, VVD92, Voe95, VM99, WA98, Wri93, hC94a, hC94b, vdHSvdV95]. valued [BI96, Bul96a, Chu97a, Chu97b, FKM97, GM90b, Pro94, Sin95, Tok96a, ZT99]. values [Aco97b, And90, BZ93, CT99, DNN92, De 91a, FM96, Hai92]. Vandermonde [Müh96]. Varga [Bul94a, VR91]. Variable [KT97a, LMP97, Oku93, BDS96, BK90, CSS93, CMR93, Con99, Dub97, FS90, FK95, GJ95, Gre93, KM94b, Lew96a, Lew98, MTW96, MJ96, MJG96a, MJG96b, Moh98, Moo95, SFGF^{+}99, SY90, Vui95, Zha94]. Variable-order [KT97a, Zha94]. variable-step [KT97a]. variable-stepsize [Zha94]. variables [ÁNGM95, BPV93, GS91b, Han91b, Jo’90, KKL97, LR92a, Moh90, NS94, Nie95, Rin94, Zhu97]. variance [För95, GR95, JDV95, Rol99, SV95b]. variates [Kem90b, Sta90]. variation [CMR95]. variation-diminishing [CMR95]. Variational [Ros95a, CP96, EK98, Fen94, NZ90, NNR93, OO96, Rán92, Tse95].

Variations [Grü98, Mar93b, BC98b]. Varieties [Ver96]. varying [BG98, Lap93, Pan93b, Tot98, dlCYL98, Bul96c]. Veber [PP91]. Vector [Bre98, DM90, Le 93, Osa96, Ben94a, BV90, Chu97b, DVB96, Fan95, Fis91, FFM95, FMT91, GMS91, GMR94, GMBW96, GMR97, JS91, JS97a, Kal95a, KC92, MZ95, Osa91, Pry94, Rob96, Rob99, Sad90, SOMF91, VB90b]. vector-valued [Chu97b]. vectors [Nie90]. vehicle [De 99]. velocity [Faz90, MEES95, ZZ97]. velocity-pressure-stress [ZZ97]. Verification [Sch96a, May95, Nak91, Ois95]. verifications [YN95]. vérificative [Han96]. Verified [Rum95]. verifying [Han96]. Verlag [Bul94b, Bul96a, Bul96b, Bul96c, Die94a, Lev90a, Lev91, VR91, Ver96]. Verlagsgesellschaft [Eme94c]. Versal [Sch94b]. version [Anh99, Che97a, Ern96, GH96, GC97, LM97, MP95, Waa91, vN93]. versions [Nee94, Sur91, Weg90b]. versus [IY97]. vertex [Joe96]. vertex-modified [Joe96]. Vertical [PdST96, eMB91, TOCH99]. very [MN96]. very-well-poised [MN96]. Vetter [Kra96]. via [AP98a, Ada95, BGH98, Cha95, HLCW93, HHL97, Jer91, Lau91, LM94c, Man96, Pel94, RNM97, SFB94]. vibrating [SJAM98]. Victor [Bul94c]. video [Uyt96]. Vieweg [Eme94c]. vii [Bul94a]. viii [Bul90b]. visco [Ned97a]. viscoelastic [Boc95, JSWW95, KKS96a]. viscoelasticity [SWW96]. viscoplastic [CZS96]. viscosity [MET93]. viscous [DGH95, Ehr94, Eme94c, SR99]. vision [BM95b]. visual [Ano94m]. visualization [AV96]. visualize [GS99]. vitae [Ano91h]. Vlad [Bul96c]. void [KV99]. voids [Mar96, Mar97]. Voigt [LW91]. Vol [Ano91-28, Ano92l, Ano93-31, Ano97-34, Ano97-35, Bul91, Lev91, Van91b, Ano91-29, Die94a]. volatility [ZFV98]. Volterra [BD90, Bru94, BY96a, BY96b, Cah90, CS97b, CRV90, CRvdHV93, CFR96, CRV97, DBFV96, EK96, FB96, GL94, Guo95a, HLS94, Jód91b, JM96, Kha94, Rob98, Vec96]. Volume [AGW91, Ano90f, Ano91c, Ano92b, Ano96j, Ano98d, Lev91, VR91, Ver94b, Ver96, Ano90b, Ano90c, Ano90d, Ano90e, Ano91a, Ano91b, Ano91d, Ano91e, Ano92c, Ano93d, Ano93e, Ano93f, Ano93g, Ano93h, Ano94c, Ano94d, Ano94e, Ano94f, Ano94g, Ano94h, Ano94i, Ano95a, Ano95b, Ano95c, Ano95d, Ano95e, Ano95f, Ano95g, Ano95h, Ano95i, Ano96a, Ano96b, Ano96c, Ano96d, Ano96e, Ano96f, Ano96g, Ano96h, Ano96i, Ano96k, Ano97c, Ano97d, Ano97e, Ano97f, Ano97g, Ano97h, Ano97i, Ano97j, Ano97k, Ano97l, Ano98e, Ano98f, Ano98g, Ano98h, Ano98i, Ano98j, Ano98k, Ano98l, Ano98m, Ano98n, Ano98o, Ano98c, Ano99a, Ano99b, Ano99c, Ano99d, Ano99e, Bul90b, Bul94c, EBV^{+}99, Eme94c, FFLM95, LW99, SdO99, Sty95]. volume [TBF99]. volumes [Ano90l, Ano90o, Ano94z, Ano95y, Ano96-28, Ano98a, Ano98b, Ano98-34, Ano98-35, Bul94b, WF97]. Voronoi [RF99b, RF99a]. vortex [vdFvGdRdV95]. vortices [TBF99]. Vries [BDKM96, DPTW95, FM98a]. vs [AVW99]. Vykhandu [Bru96].

W [Bul94a, Eme94d, Kra96]. walk [DG92, VS93]. walks [CSvD98, LSTZ94, Voi95]. Walsh [Hwa96, Sta93]. Wang [Bul90a, PH98]. warming [Gro91a]. Wasserstein [HK94]. Water [Ano93v, DS98, Dor99b, ADH98a, EEAES99, KKNU97, Ush91]. waterfall [eM94a]. Watson [Lew97a]. Wave [Ano91q, eMM97a, AHEM97, CM97, DT96, Han91d, Ira94, KKS96a, LL99b, RP96, Ren97b, Slo91, Ush91]. Waveform [Cho92, Gea91, CFR96, CRV97, Mie98]. wavefronts [Els99]. Wavelet [MXZ97, AY94, Mon96b, Sif98]. wavelets [BH95, EHK^{+}95, Reg95]. waves [ADH98a, ADH98b, CM92, DS98, Dor99b, Ehr94, EEAES99, Ked93]. Way [Ano94l, Eme94f, RP96]. ways [Roo94]. WC1E [Ano94k, Ano94m, Eme94a, Eme94e]. WC1N [Eme94d]. Weak [BGH98, GW91, GPRV93, Mar97, Ozi91b, Pan93c]. weakened [Ad94]. weakly [BH98, BEC99, Bar97, GB90b]. Wegelerstr [Ano93u]. Wegelerstrasse [Ano91r]. Weierstrass [PK97, Pro94, SP96, ZZ96]. Weierstrass-type [ZZ96]. weight [BP97a, Ehr95b, GK96a, GM99, Hen90b, Kui95c, LSSZ90, Not90, Pre90, Skr92]. Weighted [Jhu95, KvD95, HLCW93, JS92, Mas95, Mat95, PCV99, VM94]. weighting [Tre91]. weights [Ber97, BM97, BG98, CG97, Dam98, Dam99, EEKL93, JK98b, LL98a, LM93b, LM95, Mat95, Tot98]. Weisbaden [Eme94c]. Weiss [CRV90]. well [BGN^{+}90, Cuy99, MN96, Wit96a, Wit96c]. well-distributed [BGN^{+}90]. Weyl [WGK99]. WG10.3 [Ano93s]. wheel [Sim96]. which [Ask93, CS95b]. Whipple [LGR96]. whole [GM93]. Wickham [Pie96]. Wickham-Jones [Pie96]. Wiener [JNS90, LSSZ90, Pan93a]. Wills [Ver94b]. Wilson [Ban99, Rah96]. Wimp [Fan95]. Wissenschaftsverlag [Kra96]. with/without [Ari95]. within [DEHG92, TOCH99]. without [Ari95, CP93, LMF99, RS94, Zha92]. Wolfram [Pie96]. Wong [Ano98p]. work [VV95]. Working [Ano93s]. Workshop [Ano91z, Ano93t, Ano97-36, Ano93-29]. workstation [GN96]. World [Ano93a, Ano97a].

X [Ano94m, Bul96b, CB92, SS91b, Eme94b]. xii [Ano94j, Ano98q]. xvi [Bul94c]. xvii [Die94a]. xx [Ano98r]. xxii [Uyt96]. xxvii [Bul96c]. xxviii [Bul96a].

Y-MP [GP93]. yields [LM90]. York

[Ano93k, Bul90b, Bul91, Lev91].

Zeilberger [Ekh91, Han96, Koo93, Koo98]. Zero [KR98, KR99, DGC95, DT98, Miy92, Pel94, PT95, Tsi98]. zero-dissipative [Tsi98]. zero-excess [DGC95]. zerofinding [PT95]. zeroes [KP99]. Zeros [CES91, GL99, HKL96, ALR96, AÁNMP98, Ata94, Bav95b, CP93, CI97, De 93, EM98a, FB90, GR92b, GR92c, GS99, Grü98, IS90, IKK91, IS92, IS93, IY95, IKF91, KvD95, KSK92, Kui95b, LM93a, LW96b, Lan98, Leo98a, Leo98b, LW95, Let96, LU96, MG91, Mei93, Miy93, Mul93, Mul95, NW97, PSG91, PS95a, PP93, Pet90, PH92, RZG95a, RM95b, STS91, Sto95, Van93d, ZDR90, ZRG93]. Zeta [MA98]. ZEUS [Sto99]. Zheng [PH98]. Zhou [Eme94f]. Zhu [Bul91]. Zlámal [Mar95a]. Zosel [Eme94e].

References

A:1994:AMS

[A.94] Dang Quang A. Approximate method for solving an elliptic problem with discontinuous coefficients. Journal of Computational and Applied Mathematics, 51(2): 193–203, June 20, 1994. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279200006U.

Arvesu:1998:JST

[AÁNMP98] J. Arvesú, R. Álvarez-Nodarse, F. Marcellán, and K. Pan. Jacobi–Sobolevtype orthogonal polynomials: Second-order differential equation and zeros. Journal of Computational and Applied Mathematics, 90(2):135–156, April 17, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042798000053.

Anderson:1998:TPS

[AAP98] D. Anderson, R. Avery, and A. Peterson. Three positive solutions to a discrete focal boundary value problem. Journal of Computational and Applied Mathematics, 88(1):103–118, February 23, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704279700201X.

Atakishiyeva:1998:SIQ

[AAVB98] M. K. Atakishiyeva, N. M. Atakishiyev, and C. Villegas-Blas. On the square integrability of the q-Hermite functions. Journal of Computational and Applied Mathematics, 99(1–2):27– 35, November 16, 1998. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042798001423.

Anastassiou:1996:SEE

[AB96] George A. Anastassiou and Alexander Bendikov. A sharp error estimate for the numerical solution of multivariate Dirichlet problems. Journal of Computational and Applied Mathematics, 75(2):215–229, November 28, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042796000520.

Amodio:1997:NEI

[AB97a] Pierluigi Amodio and Luigi Brugnano. A note on the efficient implementation of implicit methods for ODEs. Journal of Computational and Applied Mathematics, 87(1):1–9, December 18, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042797001672.

Amodio:1997:PIB

[AB97b] Pierluigi Amodio and Luigi Brugnano. Parallel implementation of block boundary value methods for ODEs. Journal of Computational and Applied Mathematics, 78(2):197–211, February 17, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042796001124.

Abdalkhani:1990:NAS

[Abd90] J. Abdalkhani. A numerical approach to the solution of Abel integral equations of the second kind with nonsmooth solution. Journal of Computational and Applied Mathematics, 29(3):249– 255, March 15, 1990. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/037704279090011N.

Abdou:1996:FIE

[Abd96a]  M. A. Abdou.  Fred-

holm integral equation of the second kind with potential kernel. Journal of Computational and Applied Mathematics, 72(1): 161–167, July 30, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0377042795002693.

Arbogast:1996:CMM

[ABD^{+}96b] Todd Arbogast, Steve Bryant, Clint Dawson, Fredrik Saaf, Chong Wang, and Mary Wheeler. Computational methods for multiphase flow and reactive transport problems arising in subsurface contaminant remediation. Journal of Computational and Applied Mathematics, 74(1–2):19–32, November 5, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042796000155.

Abdellaoui:1998:OSM

[Abd98] Taoufiq Abdellaoui. Optimal solution of a Monge– Kantorovitch transportation problem. Journal of Computational and Applied Mathematics, 96(2):149– 161, September 15, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042798000946.

Ashwin:1996:FSB

[ABM96] Peter Ashwin, Klaus Böhmer, and Z. Mei. Forced symmetry breaking of homoclinic cycles in a PDE with o(2) symmetry. Journal of Computational and Applied Mathematics, 70(2): 297–310, June 28, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279500212X.

Aptekarev:1997:TTD

[ABM97] A. I. Aptekarev, A. Branquinho, and F. Marcellán. Toda-type differential equations for the recurrence coefficients of orthogonal polynomials and Freud transformation. Journal of Computational and Applied Mathematics, 78(1):139– 160, February 3, 1997. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042796001380.

Abrahamsson:1991:OGG

[Abr91] Leif Abrahamsson. Orthogonal grid generation for twodimensional ducts. Journal of Computational and Applied Mathematics, 34(3): 305–314, April 26, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
037704279190091W.

Abouir:1990:EFM

[AC90] J. Abouir and A. Cuyt. Error formulas for multivariate rational interpolation and Padé approximation. Journal of Computational and Applied Mathematics, 31(2):233–241, August 31, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279090166W.

Andersen:1995:LAD

[AC95] Knud D. Andersen and Edmund Christiansen. Limit analysis with the dual affine scaling algorithm. Journal of Computational and Applied Mathematics, 59(2): 233–243, May 19, 1995. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279400031U.

Ashkenazi:1992:SCN

[ACD92] Max Ashkenazi, Shui-Nee Chow, and Byron Drachman. Symbolic computation of normal forms for nonlinear Hamiltonian systems. Journal of Computational and Applied Mathematics, 39(3):363– 378, May 8, 1992. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279290210O.

Allgower:1991:CGM

[ACGW91] E. L. Allgower, C.-S. Chien, K. Georg, and C.-F. Wang. Conjugate gradient methods for continuation problems. Journal of Computational and Applied Mathematics, 38(1–3):1–16, December 23, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279190157F.

Alfaro:1998:SFO

[ACM98] Manuel Alfaro, María José Cantero, and Leandro Moral. Semiorthogonal functions and orthogonal polynomials on the unit circle. Journal of Computational and Applied Mathematics, 99(1–2):3–14, November 16, 1998. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S037704279800140X.

Acosta:1997:BPA

[Aco97a] Francisco Pérez Acosta. Best polynomial approximation in Besov spaces. Journal of Computational and Applied Mathematics, 85(2):315–323, November 12, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042797001362.

Acosta:1997:BMS

[Aco97b] Francisco Pérez Acosta. Boundedness of minimax series from values of the function and its derivative. Journal of Computational and Applied Mathematics, 81(1): 101–106, June 1, 1997. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042797000204.

Acosta:1997:CBS

[Aco97c] Francisco Pérez Acosta. On certain Banach spaces in connection with interpolation theory. Journal of Computational and Applied Mathematics, 83(1):55–69, September 30, 1997. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042797000605.

Argyros:1994:JMB

[ACQ94] I. K. Argyros, Dong Chen, and Qingshan Qian. The Jarratt method in Banach space setting. Journal of Computational and Applied Mathematics, 51(1): 103–106, May 30, 1994. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042794900930.

Anderson:1992:DBN

[AD92a] Oliver D. Anderson and Jan G. De Gooijer. Discriminating between nonstationary and nearly nonstationary time series models: a simulation study. Journal of Computational and Applied Mathematics, 41(3):265– 280, August 28, 1992. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279290135K.

Apostolopoulos:1992:PAS

[AD92b] Theodore Apostolopoulos and George Dassios. A parallel algorithm for solving the inverse scattering moment problem. Journal of Computational and Applied Mathematics, 42(1):63–77, September 8, 1992. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279290163R.

Abdou:1994:IPW

[Ad94] M. A. Abdou and E. A. Khar-El din. An infinite plate weakened by a hole having arbitrary shape. Journal of Computational and Applied Mathematics, 56(3):341–351, December 30, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042794900884.

Amato:1997:CRI

[AD97] Umberto Amato and Italia De Feis. Convergence for the regularized inversion of Fourier series. Journal of Computational and Applied Mathematics, 87(2):261–284, December 23, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042797001933.

Adamchik:1995:EIB

[Ada95]  Victor Adamchik.  The

evaluation of integrals of Bessel functions via Gfunction identities. Journal of Computational and Applied Mathematics, 64 (3):283–290, December 20, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042795001530.

Adamchik:1997:SNE

[Ada97]  Victor Adamchik.  On

Stirling numbers and Euler sums. Journal of Computational and Applied Mathematics, 79(1):119– 130, March 3, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042796001677.

Adamchik:1998:PFN

[Ada98] Victor S. Adamchik. Polygamma functions of negative order. Journal of Computational and Applied Mathematics, 100(2):191–199, December 21, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042798001927.

Abou-Dina:1999:BIF

[ADG99] M. S. Abou-Dina and A. F.

Ghaleb. On the boundary integral formulation of the plane theory of elasticity with applications (analytical aspects). Journal of Computational and Applied Mathematics, 106(1): 55–70, June 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042799000527.

Abou-Dina:1998:GWM

[ADH98a] M. S. Abou-Dina and F. M. Hassan. Generation of waves by a moving plate in stratified shallow water. Journal of Computational and Applied Mathematics, 88(2):239– 253, March 2, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042797001866.

Abou-Dina:1998:RNP

[ADH98b] M. S. Abou-Dina and M. A. Helal. Reduction for the nonlinear problem of fluid waves to a system of integrodifferential equations with an oceanographical application. Journal of Computational and Applied Mathematics, 95(1–2):65–81, August 28, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042798000727.

Artes:1998:LCC

[ADMFSR98] P. L. Artés, J. S. Dehesa, A. Martínez-Finkelshtein, and J. Sánchez-Ruiz. Linearization and connection coefficients for hypergeometrictype polynomials. Journal of Computational and Applied Mathematics, 99(1–2): 15–26, November 16, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042798001411.

Arge:1992:BSI

[ADT92] Erlend Arge, Morten Dæhlen, and Aslak Tveito. Box spline interpolation; a computational study. Journal of Computational and Applied Mathematics, 44 (3):303–329, December 24, 1992. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
037704279290003G.

Arge:1995:ASD

[ADT95] Erlend Arge, Morten Dæhlen, and Aslak Tveito. Approximation of scattered data using smooth grid functions. Journal of Computational and Applied Mathematics, 59(2):191– 205, May 19, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279400033W.

Ang:1996:ISP

[ADT96] D. D. Ang, A. Pham Ngoc Dinh, and D. N. Thanh. An inverse Stefan problem: identification of boundary value. Journal of Computational and Applied Mathematics, 66(1–2):75–84, January 31, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279500159X.

Ang:1997:BIS

[ADT97] D. D. Ang, A. Pham Ngoc Dinh, and D. N. Thanh. A bidimensional inverse Stefan problem: Identification of boundary value. Journal of Computational and Applied Mathematics, 80(2): 227–240, May 5, 1997. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042797000216.

Aihie:1990:CEF

[AE90] V. U. Aihie and G. A. Evans. A comparison of the error function and the tanh transformation as progressive rules for double and triple singular integrals. Journal of Computational and Applied Mathematics, 30(2): 145–154, May 28, 1990. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279090023S.

Abelman:1991:RBS

[AE91] S. Abelman and D. Eyre.

A rational basis for secondkind Abel integral equations. Journal of Computational and Applied Mathematics, 34(3):281–290, April 26, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042791900893.

Andersson:1998:SBE

[AEG98] Lars-Erik Andersson, Tommy Elfving, and Gene H. Golub. Solution of biharmonic equations with application to radar imaging. Journal of Computational and Applied Mathematics, 94(2):153–180, August 3, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704279800079X.

Afflerbach:1990:CAR

[Aff90] Lothar Afflerbach. Criteria for the assessment of random number generators.

Journal of Computational and Applied Mathematics, 31(1):3–10, July 24, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042790903303.

Auzinger:1993:MCT

[AFK93] W. Auzinger, R. Frank, and G. Kirlinger. Modern convergence theory for stiff initial-value problems. Journal of Computational and Applied Mathematics, 45(1– 2):5–16, April 8, 1993. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/037704279390260I.

Abdel-Gawad:1998:ASS

[AG98] H. I. Abdel-Gawad. An approach to solutions of systems of linear partial differential equations with applications. Journal of Computational and Applied Mathematics, 95(1–2):13–27, August 28, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042798000612.

Agarwal:1998:P

[Aga98] Ravi P. Agarwal. Preface. Journal of Computational and Applied Mathematics, 88(1):1–2, February 23, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042797002331.

Ammar:1993:ASP

[AGR93] Gregory S. Ammar, William B.

Gragg, and Lothar Reichel. An analogue for Szegő polynomials of the Clenshaw algorithm. Journal of Computational and Applied Mathematics, 46(1–2):211–216, June 14, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279390296N.

Area:1997:MRR

[AGRZ97] I. Area, E. Godoy, A. Ronveaux, and A. Zarzo. Minimal recurrence relations for connection coefficients between classical orthogonal polynomials: Discrete case. Journal of Computational and Applied Mathematics, 87(2):321–337, December 23, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042797001969.

Area:1998:MRR

[AGRZ98] I. Area, E. Godoy, A. Ronveaux, and A. Zarzo. Minimal recurrence relations for connection coefficients between classical orthogonal polynomials: Discrete case. Journal of Computational and Applied Mathematics, 89(2):309– 325, March 16, 1998. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042798000028.

Allgower:1991:VIB

[AGW91] Eugene L. Allgower, Kurt Georg, and Ralf Widmann. Volume integrals for boundary element methods. Journal of Computational and Applied Mathematics, 38(1–3):17–29, December 23, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279190158G.

Atanassova:1995:LBO

[AH95] Lidiya Y. Atanassova and Jürgen P. Herzberger. Lower bounds for the R-order of convergence of simultaneous inclusion methods for polynomial roots and related iteration methods. Journal of Computational and Applied Mathematics, 60(1–2):3–12, June 20, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279400080K.

Al-Homidan:1998:HMS

[AH98] Suliman Al-Homidan. Hybrid methods for solving the educational testing problem. Journal of Computational and Applied Mathematics, 91(1):31–45, April 24, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042798000119.

Abdel-Hafz:1997:DKW

[AHEM97] A. M. Abdel-Hafz, A. H.

Essawy, and M. A. M. Moubark. The diffraction of Kelvin wave due to an island. Journal of Computational and Applied Mathematics, 84(2):147–160, October 28, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042797000976.

Anwar:1991:DCS

[AK91] M. N. Y. Anwar and M. Y. Z. Koujouk. Deflection of concentrically supported and eccentrically loaded thin circular plate. Journal of Computational and Applied Mathematics, 36(2):189–207, August 27, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279190026G.

Ahn:1997:ACA

[AK97] Young Joon Ahn and Hong Oh Kim. Approximation of circular arcs by Bézier curves. Journal of Computational and Applied Mathematics, 81(1): 145–163, June 1, 1997. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S037704279700037X.

Acker:1997:TFB

[AKM97] Andrew Acker, Ercan Kadakal, and Kenneth G. Miller. A trial-free-boundary method for computing Batchelor flows. Journal of Computational and Applied Mathematics, 80(1):31–48, April 14, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042796001525.

Andreev:1992:SRL

[AKV92] A. B. Andreev, V. A. Kascieva, and M. Vanmaele. Some results in lumped mass finite-element approximation of eigenvalue problems using numerical quadrature formulas. Journal of Computational and Applied Mathematics, 43(3):291– 311, December 2, 1992. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279290016Q.

Aid:1997:NIG

[AL97a] René Aïd and Laurent Levacher. Numerical investigations on global error estimation for ordinary differential equations. Journal of Computational and Applied Mathematics, 82(1–2):21– 39, September 15, 1997. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042797000794.

Axelsson:1997:SSE

[AL97b] Owe Axelsson and Hao Lu.

A survey of some estimates of eigenvalues and condition numbers for certain preconditioned matrices. Journal of Computational and Applied Mathematics, 80(2): 241–264, May 5, 1997. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S037704279700023X.

Axelsson:1998:AMI

[AL98] O. Axelsson and M. Larin.

An algebraic multilevel iteration method for finite element matrices. Journal of Computational and Applied Mathematics, 89(1):135– 153, March 9, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042797002410.

Alladi:1996:QPI

[All96] Krishnaswami Alladi. The quintuple product identity and shifted partition functions. Journal of Computational and Applied Mathematics, 68(1–2):3–13, April 22, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042795002510.

Allouche:1997:RAN

[All97] Jean-Paul Allouche. A remark on Apéry’s numbers. Journal of Computational and Applied Mathematics, 83(1):123–125, September 30, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042797000745.

Alpert:1995:HOQ

[Alp95] Bradley K. Alpert. Highorder quadratures for integral operators with singular kernels. Journal of Computational and Applied Mathematics, 60(3):

367–378, June 30, 1995. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042794000408.

Alpay:1998:STR

[Alp98]  D. Alpay.  A structure

theorem for reproducing kernel Pontryagin spaces. Journal of Computational and Applied Mathematics, 99(1–2):413–422, November 16, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042798001745.

Alfaro:1996:SPZ

[ALR96] M. Alfaro, G. López, and M. L. Rezola. Some properties of zeros of Sobolevtype orthogonal polynomials. Journal of Computational and Applied Mathematics, 69(1):171–179, April 30, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042795000348.

Alves:1999:NSR

[Alv99] Carlos J. S. Alves. Numerical simulations on resonance poles — the trapping case of two plane cracks. Journal of Computational and Applied Mathematics, 111(1–2):267–279, November 15, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704279900148X.

Alwash:1996:PSQ

[Alw96] M. A. M. Alwash. Periodic solutions of a quartic differential equation and Groebner bases. Journal of Computational and Applied Mathematics, 75(1):67–76, November 12, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042796000593.

Alzer:1997:HMI

[Alz97]  Horst Alzer.  A har-

monic mean inequality for the gamma function. Journal of Computational and Applied Mathematics, 87(2):195–198, December 23, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042796001811.

See corrigendum [Alz98].

Alzer:1998:CHM

[Alz98] Horst Alzer. Corrigendum: A harmonic mean inequality for the gamma function [J. Comput. Appl. Math. 87 (1997) 195–198]. Journal of Computational and Applied Mathematics, 90(2): 265, April 17, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042798000405.

See [Alz97].

Ali:1995:EUD

[AM95] Farhad Ali and D. F. Mayers. Error in the upwind differencing of the convection– diffusion equation. Journal of Computational and Applied Mathematics, 58(2): 201–224, March 30, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042793E0271M.

Amodio:1996:PBP

[AM96] Pierluigi Amodio and Francesca Mazzia. Parallel block preconditioning for the solution of boundary value methods. Journal of Computational and Applied Mathematics, 69(2):191– 206, May 23, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042795000399.

Amodio:1997:NSD

[AM97] Pierluigi Amodio and Francesca Mazzia. Numerical solution of differential algebraic equations and computation of consistent initial/boundary conditions. Journal of Computational and Applied Mathematics, 87(1):135–146, December 18, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042797001787.

Amano:1994:CSM

[Ama94]  K. Amano.  A charge

simulation method for the numerical conformal mapping of interior, exterior and doubly-connected domains. Journal of Computational and Applied Mathematics, 53(3):353–370, August 30, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042794900639.

Axelsson:1995:SAS

[AN95] Owe Axelsson and Maya Neytcheva. Scalable algorithms for the solution of Navier’s equations of elasticity. Journal of Computational and Applied Mathematics, 63(1–3):149–178, November 20, 1995. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042795000542.

Alvarez-Nodarse:1998:SOD

[ÁN98] R. Álvarez-Nodarse. Second order difference equations for certain families of ‘discrete’ polynomials. Journal of Computational and Applied Mathematics, 99(1–2):37–45, November 16, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042798001435.

Andrew:1990:ESV

[And90] Alan L. Andrew. Eigenvalues and singular values of certain random matrices. Journal of Computational and Applied Mathematics, 30(2):165– 171, May 28, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279090025U.

Andres:1992:APS

[And92] Jan Andres. Asymptotic properties of solutions to quasi-linear differential systems. Journal of Computational and Applied Mathematics, 41(1–2):57–64, August 20, 1992. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279290238S.

Anderson:1995:MET

[And95] Oliver D. Anderson. More effective time-series analysis and forecasting. Journal of Computational and Applied Mathematics, 64(1–2):117– 147, November 30, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042795000119.

Andrews:1996:PMI

[And96] George E. Andrews. Pfaff’s method (II): diverse applications. Journal of Computational and Applied Mathematics, 68(1–2):15– 23, April 22, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042795002588.

Alvarez-Nodarse:1995:PMC

[ÁNGM95] R. Álvarez-Nodarse, A. G. García, and F. Marcellán. On the properties for modifications of classical orthogonal polynomials of discrete variables. Journal of Computational and Applied Mathematics, 65(1–3):3–18, December 29, 1995. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042795000976.

Anh:1999:CPV

[Anh99] Pham Ky Anh. Comments on “The parallel version of the successive approximation method for quasilinear boundary-value problem”, by Scheiber Ernő. Journal of Computational and Applied Mathematics, 102(2):333– 334, February 26, 1999. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042798002271.

See [Ern96].

Anonymous:1990:AC

[Ano90a]  Anonymous.  1. An-

nouncements of conferences. Journal of Computational and Applied Mathematics, 32(1–2):N2–N5, November 26, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042790904423.

Anonymous:1990:AIVa

[Ano90b] Anonymous. Author index to volume 29 (1990). Journal of Computational and Applied Mathematics, 29(3): 373–374, March 15, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        037704279090019V.

Anonymous:1990:AIVb

[Ano90c] Anonymous. Author index to volume 30 (1990). Journal of Computational and Applied Mathematics, 30(3): 359–360, July 10, 1990. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0377042790902869.

Anonymous:1990:AIVc

[Ano90d] Anonymous. Author index to volume 31 (1990). Journal of Computational and Applied Mathematics, 31(3):397–398, October 1, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042790900407.

Anonymous:1990:AIVd

[Ano90e] Anonymous. Author index to volume 32 (1990). Journal of Computational and Applied Mathematics, 32(3):431–432, December 10, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042790900485.

Anonymous:1990:AIVe

[Ano90f] Anonymous. Author index to volume 33 (1990).

Journal of Computational and Applied Mathematics, 33(3):347, December 31, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042705800093.

Anonymous:1990:EBa

[Ano90g]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 29(1):ii–iii, January 10, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042790901886.

Anonymous:1990:EBb

[Ano90h]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 31(1):ii–iii, July 24, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
037704279090328W.

Anonymous:1990:EBc

[Ano90i]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 32(1–2):vi–vii, November 26, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279090409S.

Anonymous:1990:EBd

[Ano90j]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 33(1):ii–iii, December 17, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279090251T.

Anonymous:1990:IRD

[Ano90k] Anonymous. Institutional reports and doctoral theses: Department of Computer Science, K. U. Leuven, Celestijnenland 200A, B-3030 Heverlee (Leuven), Belgium. Journal of Computational and Applied Mathematics, 32(1–2):N8, November 26, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042790904478.

Anonymous:1990:MIV

[Ano90l] Anonymous. Master index to volumes 21–30. Journal of Computational and Applied Mathematics, 30(3): 361–379, July 10, 1990. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/037704279090287A.

Anonymous:1990:NCAa

[Ano90m] Anonymous. Newsletter on computational and applied mathematics. Journal of Computational and Applied Mathematics, 29(3):n1–n15, March 15, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279090020Z.

Anonymous:1990:NCAc

[Ano90n] Anonymous. Newsletter on computational and applied mathematics. Journal of Computational and Applied Mathematics, 32(1–2):N1– N9, November 26, 1990. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042790904412.

Anonymous:1990:NCAb

[Ano90o]  Anonymous.  Newslet-

ter on computational and applied mathematics volumes 21–30. Journal of Computational and Applied Mathematics, 30(3): N1–N10, July 10, 1990. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279090288B.

Anonymous:1991:AIVa

[Ano91a] Anonymous. Author index to volume 34 (1991). Journal of Computational and Applied Mathematics, 34(3): 397–398, April 26, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042791900974.

Anonymous:1991:AIVb

[Ano91b] Anonymous. Author index to volume 35 (1991). Journal of Computational and Applied Mathematics, 35(1– 3):335–336, June 26, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042791902215.

Anonymous:1991:AIVc

[Ano91c] Anonymous. Author index to volume 36 (1991). Journal of Computational and Applied Mathematics, 36(3):409–410, September 24, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279190020K.

Anonymous:1991:AIVd

[Ano91d] Anonymous. Author index to volume 37 (1991). Journal of Computational and Applied Mathematics, 37(1–3):315–316, November 18, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042791901276.

Anonymous:1991:AIVe

[Ano91e] Anonymous. Author index to volume 38 (1991). Journal of Computational and Applied Mathematics, 38(1–3):477–478, December 23, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279190192M.

Anonymous:1991:CCS

[Ano91f]  Anonymous.  Conference

on Curves, Surfaces, CAGD and Image Processing. Journal of Computational and Applied Mathematics, 34(2): N7–N8, April 4, 1991. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279190057Q.

Anonymous:1991:CCT

[Ano91g]  Anonymous.  CTAC 91

Computational Techniques and Applications Conference. Journal of Computational and Applied Mathematics, 34(2):N10, April 4, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        037704279190061N.

Anonymous:1991:CVP

[Ano91h] Anonymous. Curriculum vitae of Professor C. C. Grosjean. Journal of Computational and Applied Mathematics, 37(1–3):3–4, November 18, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279190100X.

Anonymous:1991:DCS

[Ano91i] Anonymous. Department of computer science. Journal of Computational and Applied Mathematics, 34(2): N14, April 4, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704279190073S.

Anonymous:1991:EBa

[Ano91j]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 34(1):ii–iii, February 10, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042791901427.

Anonymous:1991:EBb

[Ano91k]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 35(1–3):ii–iii, June 26, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
037704279190193N.

Anonymous:1991:EBc

[Ano91l]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 36(1):ii–iii, August 20, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042791902226.

Anonymous:1991:EBd

[Ano91m]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 37(1–3):ii–iii, November 18, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042791900985.

Anonymous:1991:EBe

[Ano91n]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 38(1–3):ii–iii, December 23, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279190155D.

Anonymous:1991:EES

[Ano91o] Anonymous. EMS’91 European Simulation Multiconference. Journal of Computational and Applied Mathematics, 34(2):N5–N6, April 4, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279190055O.

Anonymous:1991:FEC

[Ano91p] Anonymous. First European Conference on Elliptic and Parabolic Problems. Journal of Computational and Applied Mathematics, 34(2): N6–N7, April 4, 1991. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/037704279190056P.

Anonymous:1991:FIC

[Ano91q] Anonymous. First International Conference on Mathematical and Numerical Aspects of Wave Propagation Phenomena. Journal of Computational and Applied Mathematics, 34(2):N4–N5, April 4, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279190053M.

Anonymous:1991:IRDa

[Ano91r] Anonymous. Institutional reports and doctoral theses: Institut für Angewandte Mathematik der Universität Bonn Sonderforschungsbereich 256 Wegelerstrasse 6 D-3500 Bonn 1. Journal of Computational and Applied Mathematics, 34(2):N15– N16, April 4, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704279190074T.

Anonymous:1991:IRDb

[Ano91s] Anonymous. Institutional reports and doctoral theses: Technische Universität Graz, Institut für Mathematik B Kopernikusgasse 24 8010 Graz, Austria. Journal of Computational and Applied Mathematics, 34(2): N16, April 4, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704279190075U.

Anonymous:1991:ICC

[Ano91t] Anonymous. International Conference on Complex Systems: Fractals, Spin Glasses and Neural Networks. Journal of Computational and Applied Mathematics, 34(2): N8–N9, April 4, 1991. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/037704279190059S.

Anonymous:1991:ICE

[Ano91u]  Anonymous.  Interna-

tional Congress on extrapolation and rational approximation. Journal of Computational and Applied Mathematics, 34(2):N14, April 4, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704279190072R.

Anonymous:1991:LIT

[Ano91v] Anonymous. List of invited talks. Journal of Computational and Applied Mathematics, 38(1–3):473–474, December 23, 1991. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/037704279190190U.

Anonymous:1991:LPP

[Ano91w]  Anonymous.  List of

papers presented at the congress. Journal of Computational and Applied Mathematics, 35(1–3):331–334, June 26, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279190220E.

Anonymous:1991:LTS

[Ano91x] Anonymous. List of talks at the symposium. Journal of Computational and Applied Mathematics, 38(1–3):475, December 23, 1991. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279190191L.

Anonymous:1991:NAS

[Ano91y] Anonymous. NATO — Advanced Study Institute Approximation Theory, Spline Functions and Applications. Journal of Computational and Applied Mathematics, 34(2):N5, April 4, 1991. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279190054N.

Anonymous:1991:NAR

[Ano91z] Anonymous. NATO Advanced Research Workshop on Approximation by Solutions of Partial Differential Equations, Quadrature Formulae, and Related Topics. Journal of Computational and Applied Mathematics, 34(2): N9–N10, April 4, 1991. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279190060W.

Anonymous:1991:NCAa

[Ano91-27]  Anonymous.  Newsletter

on computational and applied mathematics. Journal of Computational and Applied Mathematics, 34(2): N1, April 4, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279190048O.

Anonymous:1991:NCAb

[Ano91-28] Anonymous. Newsletter on computational and applied mathematics: Vol. 7, nr. 2, July 1991. Journal of Computational and Applied Mathematics, 36(1):N1– N12, August 20, 1991. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/037704279190230H.

Anonymous:1991:NCAc

[Ano91-29]  Anonymous.  Newsletter

on computational and applied mathematics: vol. 7, nr. 3, November 1991. Journal of Computational and Applied Mathematics, 37(1–3):N1–N10, November 18, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042791901287.

Anonymous:1991:PC

[Ano91-30] Anonymous. Parallel computing 91. Journal of Computational and Applied Mathematics, 34(2):N12, April 4, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279190064Q.

Anonymous:1991:P

[Ano91-31]  Anonymous.  Preface.

Journal of Computational and Applied Mathematics, 37(1–3):1–2, November 18, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042791900996.

Anonymous:1991:SIS

[Ano91-32] Anonymous. SCAN-1991:

International Symposium on Computer Arithmetic and Scientific Computation. Journal of Computational and Applied Mathematics, 34(2):N13–N14, April 4, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704279190071Q.

Anonymous:1991:SCD

[Ano91-33] Anonymous. Second colloquium on differential equations. Journal of Computational and Applied Mathematics, 34(2):N11–N12, April 4, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279190063P.

Anonymous:1991:SCM

[Ano91-34] Anonymous. SERC Conference: The Mathematics of Nonlinear Systems. Journal of Computational and Applied Mathematics, 34(2): N8, April 4, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279190058R.

Anonymous:1991:SCU

[Ano91-35] Anonymous. Short Conference on Uniform Mathematics and Applications. Journal of Computational and Applied Mathematics, 34(2): N11, April 4, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279190062O.

Anonymous:1991:SSS

[Ano91-36] Anonymous. SSOPAS: 7th Spanish Symposium on Orthogonal Polynomials and their Applications. Journal of Computational and Applied Mathematics, 34(2):

N13, April 4, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279190065R.

Anonymous:1992:AI

[Ano92a] Anonymous. Author index.

Journal of Computational and Applied Mathematics, 39(3):393, May 8, 1992. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279290214I.

Anonymous:1992:AIVa

[Ano92b] Anonymous. Author index to volume 42 (1992). Journal of Computational and Applied Mathematics, 42(3): 395, October 30, 1992. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279290087E.

Anonymous:1992:AIVb

[Ano92c] Anonymous. Author index to volume 44 (1992). Journal of Computational and Applied Mathematics, 44(3):391, December 24, 1992. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704279290011L.

Anonymous:1992:EBa

[Ano92d]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 39(1):ii–iii, February 28, 1992. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279290215J.

Anonymous:1992:EBb

[Ano92e]  Anonymous.  Editorial

board. Journal of Computational and Applied Mathematics, 40(1):ii–iii, June 12, 1992. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        037704279290037X.

Anonymous:1992:EBc

[Ano92f]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 41(1–2):ii–iii, August 20, 1992. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279290231L.

Anonymous:1992:EBd

[Ano92g]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 42(1):ii–iii, September 8, 1992. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279290157S.

Anonymous:1992:EBe

[Ano92h]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 43(1–2):ii–iii, November 25, 1992. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279290254U.

Anonymous:1992:EBf

[Ano92i]  Anonymous.  Editorial

board. Journal of Computational and Applied Mathematics, 44(1):ii–iii, December 9, 1992. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042792900494.

Anonymous:1992:NCAa

[Ano92j] Anonymous. Newsletter on computational and applied mathematics. Journal of Computational and Applied Mathematics, 39(1):n1–n12, February 28, 1992. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279290230U.

Anonymous:1992:NCAb

[Ano92k] Anonymous. Newsletter on computational and applied mathematics. Journal of Computational and Applied Mathematics, 41(1–2):N1– N16, August 20, 1992. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279290253T.

Anonymous:1992:NCAc

[Ano92l]  Anonymous.  Newsletter

on computational and applied mathematics: Vol. 8, nr. 3, November 1992. Journal of Computational and Applied Mathematics, 43(1–2):N1–N14, November 25, 1992. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279290271X.

Anonymous:1993:IWC

[Ano93a] Anonymous. 14th IMACS World Congress on Computation and Applied Mathematics. Journal of Computational and Applied Mathematics, 47(3):N8, September 30, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279390078P.

Anonymous:1993:BEI

[Ano93b] Anonymous. 16th Boundary Element International Conference. Journal of Computational and Applied Mathematics, 47(3):N8–N9, September 30, 1993. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279390079Q.

Anonymous:1993:ALA

[Ano93c] Anonymous. APL 94 — the language and its applications. Journal of Computational and Applied Mathematics, 47(3):N11, September 30, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279390082M.

Anonymous:1993:AIVa

[Ano93d] Anonymous. Author index to volume 45 (1993). Journal of Computational and Applied Mathematics, 45(3): 371–372, May 3, 1993. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279390056H.

Anonymous:1993:AIVb

[Ano93e] Anonymous. Author index to volume 46 (1993). Journal of Computational and Applied Mathematics, 46(3): 465–466, June 22, 1993. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279390042A.

Anonymous:1993:AIVc

[Ano93f] Anonymous. Author index to volume 47 (1993). Journal of Computational and Applied Mathematics, 47(3):401–402, September 30, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279390066K.

Anonymous:1993:AIVd

[Ano93g] Anonymous. Author index to volume 48 (1993). Journal of Computational and Applied Mathematics, 48(3):343, November 30, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279390029B.

Anonymous:1993:AIVe

[Ano93h] Anonymous. Author index to volume 49 (1993). Journal of Computational and Applied Mathematics, 49(1–3):361–362, December 31, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279390169C.

Anonymous:1993:BCS

[Ano93i] Anonymous. Benin conference on scientific computing. Journal of Computational and Applied Mathematics, 47(3):N3–N4, September 30, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279390070R.

Anonymous:1993:BIC

[Ano93j] Anonymous. BETECH 94:

9th International conference on Boundary Element Technology. Journal of Computational and Applied Mathematics, 47(3):N4, September 30, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279390072J.

Anonymous:1993:BRD

[Ano93k]  Anonymous.  Book re-

view: Difference equations and inequalities: Ravi P. Agarwal: Marcel Dekker, Inc., New York, 1992, 777 pages, ISBN: 0-8247-8676-

9. Journal of Computational and Applied Mathematics, 47(3):N2, September 30, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279390068M.

Anonymous:1993:CIC

[Ano93l]  Anonymous.  CEEDE

’94: The International Conference on Concurrent Engineering and Electronic Design Automation. Journal of Computational and Applied Mathematics, 47(3):N5, September 30, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279390073K.

Anonymous:1993:DCS

[Ano93m] Anonymous. Department of computer science. Journal of Computational and Applied Mathematics, 47(3):N12, September 30, 1993. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/037704279390083N.

Anonymous:1993:EBa

[Ano93n]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 45(1–2):ii–iii, April 8, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.

com/science/article/pii/
037704279390258D.

Anonymous:1993:EBb

[Ano93o]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 46(1–2):iv–v, June 14, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
037704279390282G.

Anonymous:1993:EBc

[Ano93p]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 47(1):iv–v, June 28, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
037704279390085P.

Anonymous:1993:EBd

[Ano93q]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 48(1–2):ii–iii, October 29, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042793903108.

Anonymous:1993:EBe

[Ano93r]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 49(1–3):ii–iii, December 31, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279390127W.

Anonymous:1993:IWW

[Ano93s] Anonymous. IFIP WG10.3 Working Conference on Programming Environments for Massively Parallel Distributed Systems. Journal of Computational and Applied Mathematics, 47(3): N5, September 30, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
037704279390074L.

Anonymous:1993:IWT

[Ano93t] Anonymous. IMACS Workshop on Turbulence. Journal of Computational and Applied Mathematics, 47(3): N4, September 30, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
037704279390071I.

Anonymous:1993:IRD

[Ano93u] Anonymous. Institutional reports and doctoral theses: Institut für Angewandte Mathematik: Der Universität Bonn Sonderforschungsbereich 256, Wegelerstr. 6, D-5300 Bonn 1.

Journal of Computational and Applied Mathematics, 47(3):N12–N13, September 30, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279390084O.

Anonymous:1993:ICC

[Ano93v] Anonymous. International Conference on Computational Methods in Water Resources. Journal of Computational and Applied Mathematics, 47(3):N10, September 30, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279390081L.

Anonymous:1993:ISS

[Ano93w] Anonymous. International Symposium on Special Functions, Approximation, Numerical Quadrature and Orthogonal Polynomials. Journal of Computational and Applied Mathematics, 47(3):N2–N3, September 30, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279390069N.

Anonymous:1993:LP

[Ano93x] Anonymous. List of participants. Journal of Computational and Applied Mathematics, 48(1–2):248–252, October 29, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042793903267.

Anonymous:1993:LTP

[Ano93y] Anonymous. List of talks presented at the symposium. Journal of Computational and Applied Mathematics, 48(1–2):245–247, October 29, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042793903256.

Anonymous:1993:MPP

[Ano93z] Anonymous. Massively parallel processing applications and development. Journal of Computational and Applied Mathematics, 47(3):N7–N8, September 30, 1993. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/037704279390077O.

Anonymous:1993:NCA

[Ano93-27]  Anonymous.  Newsletter

on computational and applied mathematics. Journal of Computational and Applied Mathematics, 47(3): N1, September 30, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
037704279390067L.

Anonymous:1993:SIC

[Ano93-28]  Anonymous.  Sixth In-

ternational Conference on Fibonacci Numbers and Their Applications. Journal of Computational and Applied Mathematics, 47(3): N9, September 30, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
037704279390080U.

Anonymous:1993:SMD

[Ano93-29] Anonymous. Spectral multidomain methods workshop. Journal of Computational and Applied Mathematics, 47(3):N6, September 30, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279390075M.

Anonymous:1993:TIC

[Ano93-30] Anonymous. The Third International Conference on Mathematical Methods in Computer Aided Geometric Design. Journal of Computational and Applied Mathematics, 47(3):N6–N7, September 30, 1993. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/037704279390076N.

Anonymous:1993:VNM

[Ano93-31] Anonymous. Vol. 9, nr. 1, March 1993. Journal of Computational and Applied Mathematics, 45(1–2):N1– N18, April 8, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279390278J.

Anonymous:1994:AC

[Ano94a]  Anonymous.  Announce-

ments of conferences. Journal of Computational and Applied Mathematics, 50(1– 3):N5–N19, May 20, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042794903417.

Anonymous:1994:A

[Ano94b]  Anonymous.  Appendix.

Journal of Computational and Applied Mathematics, 51(2):261–262, June 20, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042794900574.

Anonymous:1994:AIVa

[Ano94c] Anonymous. Author index to volume 50 (1994). Journal of Computational and Applied Mathematics, 50(1– 3):631–632, May 20, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042794903344.

Anonymous:1994:AIVb

[Ano94d] Anonymous. Author index to volume 51 (1994). Journal of Computational and Applied Mathematics, 51(3): 413–414, June 30, 1994. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042794900140.

Anonymous:1994:AIVc

[Ano94e] Anonymous. Author index to volume 52 (1994). Journal of Computational and Applied Mathematics, 52(1– 3):353–354, July 20, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042794903662.

Anonymous:1994:AIVd

[Ano94f] Anonymous. Author index to volume 53 (1994). Journal of Computational and Applied Mathematics, 53(3):397–398, August 30, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042794900671.

Anonymous:1994:AIVe

[Ano94g] Anonymous. Author index to volume 54 (1994). Journal of Computational and Applied Mathematics, 54(3):389–390, October 20, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042794902593.

Anonymous:1994:AIVf

[Ano94h] Anonymous. Author index to volume 55 (1994). Journal of Computational and Applied Mathematics, 55(3):385, November 30, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042794900418.

Anonymous:1994:AIVg

[Ano94i] Anonymous. Author index to volume 56 (1994). Journal of Computational and Applied Mathematics, 56(3):353, December 30, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042794900892.

Anonymous:1994:BRL

[Ano94j] Anonymous. Book review:

From logic to logic programming, Kees Doets: Foundations of Computing Series, The MIT Press, ISBN 0-262-04142-1, xii + 214 pages, Price: $43.95 / £29.95. Journal of Computational and Applied Mathematics, 52(1–3):N2–N3, July 20, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042794903689.

Anonymous:1994:BRN

[Ano94k] Anonymous. Book review:

Neural networks for pattern recognition: Albert Nigrin: M.I.T. Press, Fitzroy House, 11, Chenies Street, London WC1E 7ET, ISBN 0-262-14054-3, Cost 40.50 Pounds Sterling. Journal of Computational and Applied Mathematics, 52(1–3):N3– N4, July 20, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042794903700.

Anonymous:1994:BRP

[Ano94l]  Anonymous.  Book re-

view: Parallel supercomputing in MIMD architectures R. Michael Hord (ed.): Times Mirror International Publishers Ltd., c/o Excel Logistics, 3 Sheldon Way, Larkfield, Aylesford, Kent, ME20 6SF, England, ISBN 0-8493-4417-4, Price 61,00 Pounds Sterling. Journal of Computational and Applied Mathematics, 52(1–3): N3, July 20, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042794903697.

Anonymous:1994:BRV

[Ano94m] Anonymous. Book review: The visual mind: Art and mathematics: Michele Emmer (ed.): M.I.T. Press, Fitzroy House, 11, Chenies Street, London WC1E 7ET, ISBN 0-262-05048-X, Cost 35.95 Pounds Sterling. Journal of Computational and Applied Mathematics, 52(1– 3):N2, July 20, 1994. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042794903670.

Anonymous:1994:Ca

[Ano94n] Anonymous. Conferences. Journal of Computational and Applied Mathematics, 52(1–3):N4–N16, July 20, 1994. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042794903719.

Anonymous:1994:Cb

[Ano94o] Anonymous. Conferences.

Journal of Computational and Applied Mathematics, 54(3):N6–N21, October 20, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042794902674.

Anonymous:1994:EBa

[Ano94p]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 50(1–3):ii–iii, May 20, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042794902844.

Anonymous:1994:EBb

[Ano94q]  Anonymous.  Editorial

board. Journal of Computational and Applied Mathematics, 51(1):ii–iii, May 30, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042794900906.

Anonymous:1994:EBc

[Ano94r]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 52(1–3):iv–v, July 20, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042794903433.

Anonymous:1994:EBd

[Ano94s]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 53(1):ii–iii, July 29, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042794901953.

Anonymous:1994:EBe

[Ano94t]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 54(1):iv–v, September 20, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042794903891.

Anonymous:1994:EBf

[Ano94u]  Anonymous.  Editorial

board. Journal of Computational and Applied Mathematics, 55(1):v–vi, October 31, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042794901813.

Anonymous:1994:EBg

[Ano94v]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 56(1–2):ii–iii, December 20, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042794903735.

Anonymous:1994:IRDa

[Ano94w] Anonymous. Institutional reports and doctoral theses. Journal of Computational and Applied Mathematics, 50(1–3):N19–N21, May 20, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042794903425.

Anonymous:1994:IRDb

[Ano94x] Anonymous. Institutional reports and doctoral theses. Journal of Computational and Applied Mathematics, 52(1–3):N16–N17, July 20, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042794903727.

Anonymous:1994:LPP

[Ano94y]  Anonymous.  List of

papers presented at the congress. Journal of Computational and Applied Mathematics, 50(1–3):625–630, May 20, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042794903336.

Anonymous:1994:MIV

[Ano94z] Anonymous. Master index to volumes 41–50. Journal of Computational and Applied Mathematics, 50(1–3):633– 652, May 20, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042794903352.

Anonymous:1994:NCAa

[Ano94-27]  Anonymous.  Newsletter

on computational and applied mathematics. Journal of Computational and Applied Mathematics, 51(1): N1–N18, May 30, 1994. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0377042794900957.

Anonymous:1994:NCAb

[Ano94-28] Anonymous. Newsletter on computational and applied mathematics. Journal of Computational and Applied Mathematics, 54(3):N1, October 20, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042794902607.

Anonymous:1995:AIVa

[Ano95a] Anonymous. Author index to volume 57 (1995). Journal of Computational and Applied Mathematics, 57(3):449–450, February 28, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042795900810.

Anonymous:1995:AIVb

[Ano95b] Anonymous. Author index to volume 58 (1995). Journal of Computational and Applied Mathematics, 58(3): 363–364, April 20, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042795900632.

Anonymous:1995:AIVc

[Ano95c] Anonymous. Author index to volume 59 (1995). Journal of Computational and Applied Mathematics, 59(3): 385–386, May 30, 1995. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/037704279590073X.

Anonymous:1995:AIVd

[Ano95d] Anonymous. Author index to volume 60 (1995). Journal of Computational and Applied Mathematics, 60(3): 389–390, June 30, 1995. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0377042795901094.

Anonymous:1995:AIVe

[Ano95e] Anonymous. Author index to volume 61 (1995). Journal of Computational and Applied Mathematics, 61(3):361–362, August 21, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042795900756.

Anonymous:1995:AIVf

[Ano95f] Anonymous. Author index to volume 62 (1995). Journal of Computational and Applied Mathematics, 62(3):371–372, September 29, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042795901736.

Anonymous:1995:AIVg

[Ano95g] Anonymous. Author index to volume 63 (1995). Journal of Computational and Applied Mathematics, 63(1–3):485–486, November 20, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042795900519.

Anonymous:1995:AIVh

[Ano95h] Anonymous. Author index to volume 64 (1995). Journal of Computational and Applied Mathematics, 64(3):299–300, December 20, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042795900411.

Anonymous:1995:AIVi

[Ano95i] Anonymous. Author index to volume 65 (1995). Journal of Computational and Applied Mathematics, 65(1–3):469–470, December 29, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042795900616.

Anonymous:1995:EBa

[Ano95j]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 57(1–2):ii–iii, February 20, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042795900314.

Anonymous:1995:EBb

[Ano95k]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 58(1):ifc, March 20, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042795900837.

Anonymous:1995:EBc

[Ano95l]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 59(1):ii–iii, April 28, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042795900004.

Anonymous:1995:EBi

[Ano95m]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 60(1–2):ifc, June 20, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042795900950.

Anonymous:1995:EBd

[Ano95n]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 61(1):ifc, July 20, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042795901698.

Anonymous:1995:EBe

[Ano95o]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 62(1):ii–iii, August 31, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042795900934.

Anonymous:1995:EBf

[Ano95p]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 63(1–3):ii–iii, November 20, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042795900438.

Anonymous:1995:EBg

[Ano95q]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 64(1–2):ifc, November 30, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042795900373.

Anonymous:1995:EBh

[Ano95r]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 65(1–3):ii–iii, December 29, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042795900535.

Anonymous:1995:LIT

[Ano95s] Anonymous. List of invited talks. Journal of Computational and Applied Mathematics, 60(1–2):281–282, June 20, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042795900993.

Anonymous:1995:LPa

[Ano95t] Anonymous. List of participants. Journal of Computational and Applied Mathematics, 57(1–2):319–321, February 20, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042795900357.

Anonymous:1995:LPb

[Ano95u] Anonymous. List of participants. Journal of Computational and Applied Mathematics, 63(1–3):481–484, November 20, 1995. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042795900497.

Anonymous:1995:LRP

[Ano95v] Anonymous. List of registered participants. Journal of Computational and Applied Mathematics, 65(1– 3):464–468, December 29, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042795900594.

Anonymous:1995:LT

[Ano95w] Anonymous. List of talks.

Journal of Computational and Applied Mathematics, 60(1–2):283, June 20, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042795901019.

Anonymous:1995:LTP

[Ano95x] Anonymous. List of talks presented at the conference. Journal of Computational and Applied Mathematics, 65(1–3):461–463, December 29, 1995. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042795900578.

Anonymous:1995:MIV

[Ano95y] Anonymous. Master index to volumes 51–60. Journal of Computational and Applied Mathematics, 60(3): 391–407, June 30, 1995. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042795901116.

Anonymous:1995:NCAa

[Ano95z] Anonymous. Newsletter on computational and applied mathematics. Journal of Computational and Applied Mathematics, 59(1):N1– N17, April 28, 1995. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042795900020.

Anonymous:1995:NCAb

[Ano95-27] Anonymous. Newsletter on computational and applied mathematics. Journal of Computational and Applied Mathematics, 61(3):N1– N14, August 21, 1995. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042795900772.

Anonymous:1996:AIVa

[Ano96a] Anonymous. Author index to volume 66 (1996). Journal of Computational and Applied Mathematics, 66(1–2):589–590, January 31, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042796900230.

Anonymous:1996:AIVb

[Ano96b] Anonymous. Author index to volume 67 (1996). Journal of Computational and Applied Mathematics, 67(2): 381–382, March 29, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042796900497.

Anonymous:1996:AIVc

[Ano96c] Anonymous. Author index to volume 68 (1996). Journal of Computational and Applied Mathematics, 68(1– 2):331, April 22, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042796900357.

Anonymous:1996:AIVd

[Ano96d] Anonymous. Author index to volume 69 (1996). Journal of Computational and Applied Mathematics, 69(2): 379–380, May 23, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279690059X.

Anonymous:1996:AIVe

[Ano96e] Anonymous. Author index to volume 70 (1996). Journal of Computational and Applied Mathematics, 70(2): 357–358, June 28, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042796900412.

Anonymous:1996:AIVf

[Ano96f] Anonymous. Author index to volume 71 (1996). Journal of Computational and Applied Mathematics, 71(2): 395–396, July 27, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042796900771.

Anonymous:1996:AIVg

[Ano96g] Anonymous. Author index to volume 72 (1996). Journal of Computational and Applied Mathematics, 72(2):433–434, August 13, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042796900576.

Anonymous:1996:AIVh

[Ano96h] Anonymous. Author index to volume 73 (1996). Journal of Computational and Applied Mathematics, 73(1–2):271, October 5, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704279690014X.

Anonymous:1996:AIVi

[Ano96i] Anonymous. Author index to volume 74 (1996). Journal of Computational and Applied Mathematics, 74(1–2):345–346, November 5, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042796900758.

Anonymous:1996:AIVj

[Ano96j] Anonymous. Author index to volume 75 (1996). Journal of Computational and Applied Mathematics, 75(2):365–366, November 28, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042796900795.

Anonymous:1996:AIVk

[Ano96k] Anonymous. Author index volume 76 (1996). Journal of Computational and Applied Mathematics, 76(1–2):335–336, December 17, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042796900655.

Anonymous:1996:CNM

[Ano96l] Anonymous. Conference On Numerical and Mathematical Treatments for Elasticity SCV057SCV, SCV057SCV19. Journal of Computational and Applied Mathematics, 75(2):N5–N14, November 28, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042796900084.

Anonymous:1996:C

[Ano96m] Anonymous. Conferences. Journal of Computational and Applied Mathematics, 66(1–2):N6–N19, January 31, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042796900278.

Anonymous:1996:EBa

[Ano96n]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 66(1–2):iv–v, January 31, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042796900175.

Anonymous:1996:EBb

[Ano96o]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 67(1):ifc, February 20, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279690045X.

Anonymous:1996:EBc

[Ano96p]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 68(1–2):ii–iii, April 22, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
037704279690031X.

Anonymous:1996:EBd

[Ano96q]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 69(1):ii–iii, April 30, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042796900394.

Anonymous:1996:EBe

[Ano96r]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 70(1):ii–iii, June 14, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042796900618.

Anonymous:1996:EBf

[Ano96s]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 71(1):ii–iii, July 10, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042796900370.

Anonymous:1996:EBg

[Ano96t]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 72(1):ii–iii, July 30, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S037704279690000X.

Anonymous:1996:EBh

[Ano96u]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 73(1–2):ii–iii, October 5, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042796900138.

Anonymous:1996:EBi

[Ano96v]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 74(1–2):ii–iii, November 5, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042796900710.

Anonymous:1996:EBj

[Ano96w]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 75(1):ii–iii, November 12, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042796901077.

Anonymous:1996:EBk

[Ano96x]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 76(1–2):iii–iv, December 17, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042796900631.

Anonymous:1996:FTI

[Ano96y] Anonymous. Fourier transforms: an introduction for engineers. Journal of Computational and Applied Mathematics, 75(2):N2–N3, November 28, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042797808205.

Anonymous:1996:IRD

[Ano96z] Anonymous. Institutional reports and doctoral theses. Journal of Computational and Applied Mathematics, 66(1–2):N19–N20, January 31, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042796900291.

Anonymous:1996:LPP

[Ano96-27] Anonymous. List of papers presented at the congress. Journal of Computational and Applied Mathematics, 66(1–2):585–588, January 31, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042796900217.

Anonymous:1996:MIV

[Ano96-28] Anonymous. Master index to volumes 61–70. Journal of Computational and Applied Mathematics, 70(2): 359–378, June 28, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042796900436.

Anonymous:1996:MCC

[Ano96-29] Anonymous. Monte Carlo:

Concepts, algorithms, and applications. Journal of Computational and Applied Mathematics, 75(2):N3–N4, November 28, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042797808229.

Anonymous:1996:NCAa

[Ano96-30] Anonymous. Newsletter on computational and applied mathematics. Journal of Computational and Applied Mathematics, 66(1–2):N1, January 31, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042796900254.

Anonymous:1996:NCAb

[Ano96-31]  Anonymous.  Newsletter

on computational and applied mathematics. Journal of Computational and Applied Mathematics, 72(1): N1–N15, July 30, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042796900023.

Anonymous:1996:NCAc

[Ano96-32]  Anonymous.  Newsletter

on computational and applied mathematics. Journal of Computational and Applied Mathematics, 75(2): N1, November 28, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0377042796900813.

Anonymous:1996:PPP

[Ano96-33] Anonymous. Practical parallel programming. Journal of Computational and Applied Mathematics, 75(2): N3, November 28, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042797808217.

Anonymous:1996:SPF

[Ano96-34] Anonymous. Schur parameters, factorization and dilation problems. Journal of Computational and Applied Mathematics, 75(2): N2, November 28, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042797808199.

Anonymous:1997:IWC

[Ano97a] Anonymous. 15th IMACS World Congress on Scientific Computation, Modelling and Applied Mathematics. Journal of Computational and Applied Mathematics, 81(2):N11–N12, July 8, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042797900686.

Anonymous:1997:AC

[Ano97b]  Anonymous.  Animat-

ing calculus. Journal of Computational and Applied Mathematics, 81(2): N2, July 8, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042797894654.

Anonymous:1997:AIVa

[Ano97c] Anonymous. Author index volume 77 (1997). Journal of Computational and Applied Mathematics, 77(1–2):289, January 13, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042797827616.

Anonymous:1997:AIVb

[Ano97d] Anonymous. Author index volume 79 (1997). Journal of Computational and Applied Mathematics, 79(2): 323, March 17, 1997. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042797801227.

Anonymous:1997:AIVc

[Ano97e] Anonymous. Author index volume 80 (1997). Journal of Computational and Applied Mathematics, 80(2): 359, May 5, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042797900108.

Anonymous:1997:AIVd

[Ano97f] Anonymous. Author index volume 81 (1997). Journal of Computational and Applied Mathematics, 81(2): 375–376, July 8, 1997. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042797894642.

Anonymous:1997:AIVe

[Ano97g] Anonymous. Author index volume 82 (1997). Journal of Computational and Applied Mathematics, 82(1–2):477–478, September 15, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042797807819.

Anonymous:1997:AIVf

[Ano97h] Anonymous. Author index volume 83 (1997). Journal of Computational and Applied Mathematics, 83(2):289, October 7, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042797807881.

Anonymous:1997:AIVg

[Ano97i] Anonymous. Author index volume 84 (1997). Journal of Computational and Applied Mathematics, 84(2):281, October 28, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042797865156.

Anonymous:1997:AIVh

[Ano97j] Anonymous. Author index volume 85 (1997). Journal of Computational and Applied Mathematics, 85(2):355, November 12, 1997. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042797889558.

Anonymous:1997:AIVi

[Ano97k] Anonymous. Author index volume 86 (1997). Journal of Computational and Applied Mathematics, 86(2):427, December 10, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042797822571.

Anonymous:1997:AIVj

[Ano97l] Anonymous. Author index volume 87 (1997). Journal of Computational and Applied Mathematics, 87(2):385–386, December 23, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042797888255.

Anonymous:1997:CSA

[Ano97m] Anonymous. Chain-scattering approach to H^{∞}-control. Journal of Computational and Applied Mathematics, 81(2):N4–N5, July 8, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704279789468X.

Anonymous:1997:CCM

[Ano97n] Anonymous. CMFT97 computational methods and function theory. Journal of Computational and Applied Mathematics, 81(2): N15–N16, July 8, 1997. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042797900789.

Anonymous:1997:CCA

[Ano97o]  Anonymous.  Combina-

tional convexity and algebraic geometry. Journal of Computational and Applied Mathematics, 79(1):N2–N3, March 3, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042797816135.

Anonymous:1997:CSS

[Ano97p] Anonymous. Conferences SCV164S, SCV164SCV, SCV164SCV12. Journal of Computational and Applied Mathematics, 79(1):N5– N16, March 3, 1997. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042797900200.

Anonymous:1997:CA

[Ano97q] Anonymous. Congress announcement. Journal of Computational and Applied Mathematics, 86(2):429, December 10, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042797900431.

Anonymous:1997:EEC

[Ano97r] Anonymous. ECC97 4th European Control Conference. Journal of Computational and Applied Mathematics, 81(2):N8–N9, July 8, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042797900601.

Anonymous:1997:EBk

[Ano97s]  Anonymous.  Editional

board. Journal of Computational and Applied Mathematics, 87(1):ii–iii, December 18, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042797900315.

Anonymous:1997:EBa

[Ano97t]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 77(1–2):v–vi, January 13, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042797900479.

Anonymous:1997:EBb

[Ano97u]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 78(1):ii–iii, February 3, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042797900832.

Anonymous:1997:EBc

[Ano97v]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 79(1):v–vi, March 3, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042797900169.

Anonymous:1997:EBd

[Ano97w]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 80(1):v–vi, April 14, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042797900261.

Anonymous:1997:EBe

[Ano97x]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 81(1):ii–iii, June 1, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042797900376.

Anonymous:1997:EBf

[Ano97y]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 82(1–2):v–vi, September 15, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042797900170.

Anonymous:1997:EBg

[Ano97z]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 83(1):ii–iii, September 30, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042797900418.

Anonymous:1997:EBh

[Ano97-27]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 84(1):ii–iii, October 21, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042797900352.

Anonymous:1997:EBi

[Ano97-28]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 85(1):iii–iv, November 6, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042797900157.

Anonymous:1997:EBj

[Ano97-29]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 86(1):v–vi, November 28, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704279790039X.

Anonymous:1997:ESE

[Ano97-30] Anonymous. ENUMATH 97: Second European Conference on Numerical Mathematics and Advanced Applications. Journal of Computational and Applied Mathematics, 81(2):N16, July 8, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042797900807.

Anonymous:1997:EAE

[Ano97-31] Anonymous. Erratum: Addenda and Errata to “Linear perturbations of differential or difference operators with polynomials as eigenfunctions” [J. Comput. Appl. Math. 78 (1997) 179–195]. Journal of Computational and Applied Mathematics, 83(1):145–146, September 30, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042797001167.

See [Bav97].

Anonymous:1997:FCM

[Ano97-32] Anonymous. Foundation of computational mathematics. Journal of Computational and Applied Mathematics, 81(2):N6–N7, July 8, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042797894708.

Anonymous:1997:LGO

[Ano97-33] Anonymous. From local to global optimization. Journal of Computational and Applied Mathematics, 81(2): N10, July 8, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042797900649.

Anonymous:1997:HNAa

[Ano97-34] Anonymous. Handbook of numerical analysis, vol. III. Journal of Computational and Applied Mathematics, 81(2):N3–N4, July 8, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042797894666.

Anonymous:1997:HNAb

[Ano97-35] Anonymous. Handbook of numerical analysis, vol. IV. Journal of Computational and Applied Mathematics, 81(2):N4, July 8, 1997. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042797894678.

Anonymous:1997:HSW

[Ano97-36] Anonymous. HPOPT97:

Second Workshop on High Performance Optimization Techniques. Journal of Computational and Applied Mathematics, 81(2):N10, July 8, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042797900662.

Anonymous:1997:IRD

[Ano97-37] Anonymous. Institutional reports and doctoral theses. Journal of Computational and Applied Mathematics, 81(2):N17, July 8, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042797900844.

Anonymous:1997:IDE

[Ano97-38] Anonymous. An introduction to different equations. Journal of Computational and Applied Mathematics, 79(1):N4, March 3, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042797816159.

Anonymous:1997:IIS

[Ano97-39] Anonymous. ISIP’98 international symposium on inverse problems in engineering mechanics. Journal of Computational and Applied Mathematics, 81(2): N16–N17, July 8, 1997. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042797900820.

Anonymous:1997:LIE

[Ano97-40] Anonymous. Linear integral equations (theory and technique), second edition. Journal of Computational and Applied Mathematics, 81(2): N7–N8, July 8, 1997. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S037704279789471X.

Anonymous:1997:MIS

[Ano97-41] Anonymous. MODCOM-14: The 14th International Symposium of Continuum Mechanics Models. Journal of Computational and Applied Mathematics, 81(2): N9, July 8, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042797900625.

Anonymous:1997:MCQ

[Ano97-42] Anonymous. Monte Carlo and quasi-Monte Carlo methods in scientific computing. Journal of Computational and Applied Mathematics, 81(2):N5–N6, July 8, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042797894691.

Anonymous:1997:NCAa

[Ano97-43]  Anonymous.  Newsletter

on computational and applied mathematics. Journal of Computational and Applied Mathematics, 79(1): N1, March 3, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042797900182.

Anonymous:1997:NCAb

[Ano97-44]  Anonymous.  Newsletter

on computational and applied mathematics. Journal of Computational and Applied Mathematics, 81(2): N1, July 8, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279790011X.

Anonymous:1997:NCAc

[Ano97-45]  Anonymous.  Newsletter

on computational and applied mathematics. Journal of Computational and Applied Mathematics, 87 (1):N1–N16, December 18, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042797900339.

Anonymous:1997:NNS

[Ano97-46] Anonymous. NUMDIFF-8: Numerical solution of differential and differentialalgebraic equations. Journal of Computational and Applied Mathematics, 81(2): N12–N13, July 8, 1997. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042797900704.

Anonymous:1997:PFI

[Ano97-47]  Anonymous.  PaCT-97

Fourth International Conference on Parallel Computing Technologies. Journal of Computational and Applied Mathematics, 81(2):

N13, July 8, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042797900728.

Anonymous:1997:PAC

[Ano97-48]  Anonymous.  Polyno-

mial algorithms in computer algebra. Journal of Computational and Applied Mathematics, 79(1): N2, March 3, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042797816123.

Anonymous:1997:PAL

[Ano97-49]  Anonymous.  A poly-

nomial approach to linear algebra. Journal of Computational and Applied Mathematics, 79(1): N4, March 3, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042797816160.

Anonymous:1997:P

[Ano97-50] Anonymous. Preface. Journal of Computational and Applied Mathematics, 86(1): 1, November 28, 1997. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042797001441.

Anonymous:1997:SIS

[Ano97-51] Anonymous. SCAN-97 international symposium on scientific computing computer arithmetic and validated numerics. Journal of Computational and Applied Mathematics, 81(2): N13–N14, July 8, 1997. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042797900741.

Anonymous:1997:SSC

[Ano97-52] Anonymous. SciCADE97 scientific computing and differential equations. Journal of Computational and Applied Mathematics, 81(2): N14–N15, July 8, 1997. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042797900765.

Anonymous:1997:SIT

[Ano97-53] Anonymous. Symbolic integration I: Transcendental functions. Journal of Computational and Applied Mathematics, 79(1):N3–N4, March 3, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042797816147.

Anonymous:1998:AIVd

[Ano98a] Anonymous. Author index to volumes 81–90. Journal of Computational and Applied Mathematics, 90(2):269– 285, April 17, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042798000569.

Anonymous:1998:AIVo

[Ano98b] Anonymous. Author index to volumes 91–100 (1998). Journal of Computational and Applied Mathematics, 100(2):231–242, December 21, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042798002556.

Anonymous:1998:AIVn

[Ano98c] Anonymous. Author index volume 100 (1998). Journal of Computational and Applied Mathematics, 100(2):229, December 21, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042798002544.

Anonymous:1998:AIVa

[Ano98d] Anonymous. Author index volume 88 (1997). Journal of Computational and Applied Mathematics, 88(2):401– 402, March 2, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042797832463.

Anonymous:1998:AIVb

[Ano98e] Anonymous. Author index volume 89 (1998). Journal of Computational and Applied Mathematics, 89(2): 327, March 16, 1998. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042798000399.

Anonymous:1998:AIVc

[Ano98f] Anonymous. Author index volume 90 (1998). Journal of Computational and Applied Mathematics, 90(2): 267, April 17, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042798000557.

Anonymous:1998:AIVe

[Ano98g] Anonymous. Author index volume 91 (1998). Journal of Computational and Applied Mathematics, 91(2): 275, May 4, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042798000715.

Anonymous:1998:AIVf

[Ano98h] Anonymous. Author index volume 92 (1998). Journal of Computational and Applied Mathematics, 92(2): 139, June 26, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042798800198.

Anonymous:1998:AIVg

[Ano98i] Anonymous. Author index volume 93 (1998). Journal of Computational and Applied Mathematics, 93(2): 145, July 13, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042798001204.

Anonymous:1998:AIVh

[Ano98j] Anonymous. Author index volume 94 (1998). Journal of Computational and Applied Mathematics, 94(2): 215, August 3, 1998. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042798800149.

Anonymous:1998:AIVi

[Ano98k] Anonymous. Author index volume 95 (1998). Journal of Computational and Applied Mathematics, 95(1–2):163, August 28, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704279800123X.

Anonymous:1998:AIVj

[Ano98l] Anonymous. Author index volume 96 (1998). Journal of Computational and Applied Mathematics, 96(2):163, September 15, 1998. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042798800010.

Anonymous:1998:AIVk

[Ano98m] Anonymous. Author index volume 97 (1998). Journal of Computational and Applied Mathematics, 97(1–2):167, September 24, 1998. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042798001824.

Anonymous:1998:AIVl

[Ano98n] Anonymous. Author index volume 98 (1998). Journal of Computational and Applied Mathematics, 98(2):317, October 19, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042798001848.

Anonymous:1998:AIVm

[Ano98o] Anonymous. Author index volume 99 (1998). Journal of Computational and Applied Mathematics, 99(1–2):535–536, November 16, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042798002349.

Anonymous:1998:BRA

[Ano98p] Anonymous. Book review:

Advance topics in difference equations: Ravl P. Agarwal and Patricia J. Y. Wong Kluwer Academic Publishers, Dordrecht 1997, ISBN: 0-7923-4521-5, 520 pages, Hardbound GBP 146. Journal of Computational and Applied Mathematics, 90(1): N2, April 6, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042798800034.

Anonymous:1998:BRI

[Ano98q] Anonymous. Book review:

International, identification and sampling: J. R. Partington, London Mathematical Society Monographs, New Series, 17, Oxford University Press, Oxford, UK, 1997, ISBN 0-19-850024-6, xii + 267 pages, Hardback GBP 60. Journal of Computational and Applied Mathematics, 90(1): N2, April 6, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042798800046.

Anonymous:1998:BRU

[Ano98r]  Anonymous.  Book re-

view: Using PLAPACK: Parallel linear algebra package: R. A. van de Geijn, The MIT Press, Cambridge, 1997, ISBN 0-262-72026-4, xx + 194 pages, GBP 23.50 paperback. Journal of Computational and Applied Mathematics, 90(1): N3, April 6, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042798800058.

Anonymous:1998:EBa

[Ano98s]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 88(1):v–vi, February 23, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042798900139.

Anonymous:1998:EBb

[Ano98t]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 89(1):v–vi, March 9, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042798900450.

Anonymous:1998:EBc

[Ano98u]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 90(1):v–vi, April 6, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042798900541.

Anonymous:1998:EBd

[Ano98v]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 91(1):v–vi, April 24, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042798900310.

Anonymous:1998:EBe

[Ano98w]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 92(1):v–vi, May 20, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042798900255.

Anonymous:1998:EBf

[Ano98x]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 93(1):v–vi, July 2, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042798900358.

Anonymous:1998:EBg

[Ano98y]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 94(1):ifc, July 31, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042798900000.

Anonymous:1998:EBh

[Ano98z]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 95(1–2):v–vi, August 28, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042798900218.

Anonymous:1998:EBi

[Ano98-27]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 96(1):ii, September 3, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704279890053X.

Anonymous:1998:EBj

[Ano98-28]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 97(1–2):ii, September 24, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042798900231.

Anonymous:1998:EBk

[Ano98-29]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 98(1):ii, October 5, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042798900176.

Anonymous:1998:EBl

[Ano98-30]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 99(1–2):ifc, November 16, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704279890019X.

Anonymous:1998:EBm

[Ano98-31]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 100(1):ii, November 30, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042798900279.

Anonymous:1998:NCAa

[Ano98-32]  Anonymous.  Newsletter

on computational and applied mathematics. Journal of Computational and Applied Mathematics, 92(2): N1–N20, June 26, 1998. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042798000983.

Anonymous:1998:NCAb

[Ano98-33] Anonymous. Newsletter on computational and applied mathematics. Journal of Computational and Applied Mathematics, 100(1):N1– N14, November 30, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042798900292.

Anonymous:1998:SIVa

[Ano98-34] Anonymous. Subject index to volumes 81–90 (1998). Journal of Computational and Applied Mathematics, 90(2):287–308, April 17, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042798000570.

Anonymous:1998:SIVb

[Ano98-35] Anonymous. Subject index to volumes 91–100 (1998). Journal of Computational and Applied Mathematics, 100(2):243–256, December 21, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042798002568.

Anonymous:1999:AIVa

[Ano99a] Anonymous. Author index volume 101 (1999). Journal of Computational and Applied Mathematics, 101(1–2):265, January 15, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042799800016.

Anonymous:1999:AIVb

[Ano99b] Anonymous. Author index volume 102 (1999). Journal of Computational and Applied Mathematics, 102(2):335–336, February 26, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042799000023.

Anonymous:1999:AIVc

[Ano99c] Anonymous. Author index volume 103 (1999). Journal of Computational and Applied Mathematics, 103(2): 337–338, March 31, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042799900894.

Anonymous:1999:AIVd

[Ano99d] Anonymous. Author index volume 104 (1999). Journal of Computational and Applied Mathematics, 104(2): 215, May 17, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042799900857.

Anonymous:1999:AIVe

[Ano99e] Anonymous. Author index volume 106 (1999). Journal of Computational and Applied Mathematics, 106(2): 399–400, June 30, 1999. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042799001132.

Anonymous:1999:Ca

[Ano99f] Anonymous. Conference.

Journal of Computational and Applied Mathematics, 105(1–2):x, May 1999. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042799000060.

Anonymous:1999:Cb

[Ano99g] Anonymous. Conference.

Journal of Computational and Applied Mathematics, 105(1–2):477–478, May 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042799000084.

Anonymous:1999:Cc

[Ano99h] Anonymous. Conference.

Journal of Computational and Applied Mathematics, 105(1–2):479–482, May 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042799000072.

Anonymous:1999:Cd

[Ano99i]  Anonymous.  Correspon-

dence. Journal of Computational and Applied Mathematics, 106(1):197–202, June 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042799000783.

Anonymous:1999:Ea

[Ano99j]  Anonymous.  Editorial.

Journal of Computational and Applied Mathematics, 105(1–2):xi, May 1999. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042799000035.

Anonymous:1999:Ec

[Ano99k]  Anonymous.  Editorial.

Journal of Computational and Applied Mathematics, 111(1–2):ix, November 15, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042799003283.

Anonymous:1999:Ed

[Ano99l]  Anonymous.  Editorial.

Journal of Computational and Applied Mathematics, 112(1–2):1–2, November 30, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042799002095.

Anonymous:1999:EBa

[Ano99m]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 101(1–2):ii, January 15, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042799900833.

Anonymous:1999:EBb

[Ano99n]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 102(1):ii, February 15, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042799900870.

Anonymous:1999:EBc

[Ano99o]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 103(1):ii, March 15, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042799900912.

Anonymous:1999:EBd

[Ano99p]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 104(1):ii, April 15, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042799900006.

Anonymous:1999:Ia

[Ano99q] Anonymous. Index. Journal of Computational and Applied Mathematics, 105(1– 2):483–484, May 1999. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042799000977.

Anonymous:1999:Ib

[Ano99r] Anonymous. Index. Journal of Computational and Applied Mathematics, 107(2): 285, July 31, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042799001892.

Anonymous:1999:Ic

[Ano99s] Anonymous. Index. Journal of Computational and Applied Mathematics, 108 (1–2):233, August 15, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042799002113.

Anonymous:1999:Id

[Ano99t] Anonymous. Index. Journal of Computational and Applied Mathematics, 109(1– 2):527, September 30, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704279900309X.

Anonymous:1999:Ie

[Ano99u] Anonymous. Index. Journal of Computational and Applied Mathematics, 110(2): 307, October 30, 1999. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042799003106.

Anonymous:1999:If

[Ano99v] Anonymous. Index. Journal of Computational and Applied Mathematics, 111(1– 2):319, November 15, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042799003295.

Anonymous:1999:Ig

[Ano99w] Anonymous. Index. Journal of Computational and Applied Mathematics, 111(1– 2):321–335, November 15, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042799003301.

Anonymous:1999:Ih

[Ano99x] Anonymous. Index. Journal of Computational and Applied Mathematics, 111(1– 2):337–356, November 15, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042799003313.

Anonymous:1999:Ii

[Ano99y] Anonymous. Index. Journal of Computational and Applied Mathematics, 112(1– 2):309, November 30, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042799003325.

Anonymous:1999:NCA

[Ano99z] Anonymous. Newsletter on computational and applied mathematics. Journal of Computational and Applied Mathematics, 104(1):N1– N18, April 15, 1999. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042799000680.

Alvarez-Nodarse:1999:LPI

[ÁNQR99] R. Álvarez-Nodarse, N. R. Quintero, and A. Ronveaux. On the linearization problem involving Pochhammer symbols and their qanalogues. Journal of Computational and Applied Mathematics, 107(1):133– 146, July 15, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042799000874.

Alfeld:1996:FSD

[ANS96] Peter Alfeld, Marian Neamtu, and Larry L. Schumaker. Fitting scattered data on sphere-like surfaces using spherical splines. Journal of Computational and Applied Mathematics, 73(1–2):5–43, October 5, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042796000349.

Alvarez-Nodarse:1998:MCG

[ÁNYD98] Renato Álvarez-Nodarse, R. J. Yáñez, and Jesús S. Dehesa. Modified Clebsch– Gordan-type expansions for products of discrete hypergeometric polynomials. Journal of Computational and Applied Mathematics, 89(1):171–197, March 9, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042797002446.

Agarwal:1998:PSS

[AO98] Ravi P. Agarwal and Donal O’Regan. Positive solutions to superlinear singular boundary value problems. Journal of Computational and Applied Mathematics, 88(1):129–147, February 23, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042797002057.

Achuthan:1992:PAE

[AP92a] P. Achuthan and S. Ponnuswamy. Padé approximants and Eisenstein– Ramanujan continued fraction. Journal of Computational and Applied Mathematics, 41(3):247–264, August 28, 1992. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279290134J.

Ascher:1992:PCH

[AP92b] Uri M. Ascher and Linda R. Petzold. Projected collocation for higher-order higherindex differential-algebraic equations. Journal of Computational and Applied Mathematics, 43(1–2):243– 259, November 25, 1992. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042792902694.

Agarwal:1996:SER

[AP96] Amit N. Agarwal and Peter M. Pinsky. Stabilized element residual method (SERM): a posteriori error estimation for the advection–diffusion equation. Journal of Computational and Applied Mathematics, 74(1–2):3–17, November 5, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042796000143.

Acosta:1998:CIM

[AP98a] F. Pérez Acosta and J. M. Almira Picazo. Comparison of inequalities for the minimax series via Tchebychev polynomials. Journal of Computational and Applied Mathematics, 99(1–2):319–325, November 16, 1998. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042798001666.

Aitchison:1998:NAS

[AP98b] J. M. Aitchison and M. W.

Poole. A numerical algorithm for the solution of Signorini problems. Journal of Computational and Applied Mathematics, 94(1):55–67, July 31, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042798000302.

Aftabizadeh:1994:APO

[APH94] A. R. Aftabizadeh, N. H. Pavel, and Y. K. Huang. Anti-periodic oscillations of some second-order differential equations and optimal control problems. Journal of Computational and Applied Mathematics, 52(1–3):3–21, July 20, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279490345X.

Aptekarev:1998:MOP

[Apt98] A. I. Aptekarev. Multiple orthogonal polynomials. Journal of Computational and Applied Mathematics, 99(1–2):423–447, November 16, 1998. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042798001757.

Al-Rabeh:1992:TGI

[AR92a] Ala Al-Rabeh. Towards a general integration algorithm for time-dependent one-dimensional systems of parabolic partial differential equations using the method of lines. Journal of Computational and Applied Mathematics, 42(2):187–198, October 12, 1992. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042792900737.

Anastassiou:1992:E

[AR92b] George A. Anastassiou and Svetlozar T. Rachev. Editorial. Journal of Computational and Applied Mathematics, 40(2): 137, June 24, 1992. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279290100C.

Abdella:1997:EIT

[AR97] K. Abdella and H. Rasmussen. Electrohydrodynamic instability of two superposed fluids in normal electric fields. Journal of Computational and Applied Mathematics, 78(1):33–61, February 3, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042796001069.

Alzer:1999:SPP

[AR99] Horst Alzer and O. G.

Ruehr. A submultiplicative property of the psi function. Journal of Computational and Applied Mathematics, 101(1–2):53–60, January 15, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042798001903.

Argyros:1991:CSP

[Arg91] Ioannis K. Argyros. On the convergence of some projection methods with perturbation. Journal of Computational and Applied Mathematics, 36(2):255–258, August 27, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279190032F.

Argyros:1992:CGN

[Arg92a] Ioannis K. Argyros. On the convergence of generalized Newton methods and implicit functions. Journal of Computational and Applied Mathematics, 43(3):335– 342, December 2, 1992. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/037704279290019T.

Argyros:1992:SGP

[Arg92b] Ioannis K. Argyros. Some generalized projection methods for solving operator equations. Journal of Computational and Applied Mathematics, 39(1):1– 6, February 28, 1992. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279290216K.

Argyros:1997:CCI

[Arg97] Ioannis K. Argyros. Concerning the convergence of inexact Newton methods. Journal of Computational and Applied Mathematics, 79(2):235–247, March 17, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042796001628.

Argyros:1998:CCC

[Arg98] Ioannis K. Argyros. On the convergence of a certain class of iterative procedures under relaxed conditions with applications. Journal of Computational and Applied Mathematics, 94(1): 13–21, July 31, 1998. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042798000600.

Ariel:1995:NAC

[Ari95] P. D. Ariel. A numerical algorithm for computing the stagnation point flow of a second grade fluid with/ without suction. Journal of Computational and Applied Mathematics, 59(1):9– 24, April 28, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279400012P.

Arnold:1992:LSL

[Arn92]  Barry C. Arnold.  Lo-

gistic and semi-logistic processes. Journal of Computational and Applied Mathematics, 40(2):139– 149, June 24, 1992. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042792901013.

Askey:1996:GFT

[ARS96] Richard A. Askey, Mizan Rahman, and Sergeĭ K. Suslov. On a general q-Fourier transformation with nonsymmetric kernels. Journal of Computational and Applied Mathematics, 68(1– 2):25–55, April 22, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042795002596.

Artstein:1994:SRU

[Art94]  Zvi Artstein.  Sensi-

tivity with respect to the underlying information in stochastic programs. Journal of Computational and Applied Mathematics, 56(1–2):127–136, December 20, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042794903832.

Avdelas:1996:GHO

[AS96] G. Avdelas and T. E. Simos.

A generator of high-order embedded P -stable methods for the numerical solution of the Schrödinger equation. Journal of Computational and Applied Mathematics, 72(2):345–358, August 13, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042796000052.

Abad:1997:NEC

[AS97] Julio Abad and Javier Sesma. A new expansion of the confluent hypergeometric function in terms of modified Bessel functions. Journal of Computational and Applied Mathematics, 78(1):97–101, February 3, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042796001331.

Abad:1999:BPF

[AS99] Julio Abad and Javier Sesma. Buchholz polynomials: a family of polynomials relating solutions of confluent hypergeometric and Bessel equations. Journal of Computational and Applied Mathematics, 101(1–2):237– 241, January 15, 1999. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042799002265.

Askey:1993:PWI

[Ask93] Richard Askey. Problems which interest and/or annoy me. Journal of Computational and Applied Mathematics, 48(1–2):3–15, October 29, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279390312Y.

Aston:1997:HBN

[ASW97] Philip J. Aston, Alastair Spence, and Wu Wei. Hopf bifurcation near a double singular point. Journal of Computational and Applied Mathematics, 80(2): 277–297, May 5, 1997. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042797000319.

Arai:1991:NIS

[ASWO91] Norio Arai, Syuji Sakaguchi, Yasuhiro Wada, and Satoru Ogawa. Numerical investigation of separated flow around a bentnose biconic. Journal of Computational and Applied Mathematics, 35(1–3):99– 108, June 26, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279190199T.

Aceto:1999:SST

[AT99] Lidia Aceto and Donato Trigiante. Symmetric schemes, time reversal symmetry and conservative methods for Hamiltonian systems. Journal of Computational and Applied Mathematics, 107(2):257–274, July 31, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704279900093X.

Atanassova:1994:SDZ

[Ata94] L. Atanassova. On the simultaneous determination of the zeros of an analytic function inside a simple smooth closed contour in the complex plane. Journal of Computational and Applied Mathematics, 50(1–3): 99–107, May 20, 1994. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042794902925.

Ali:1999:GCS

[ATBB99] J. M. Ali, R. M. Tookey, J. V. Ball, and A. A. Ball. The generalised Cornu spiral and its application to span generation. Journal of Computational and Applied Mathematics, 102(1):37–47, February 15, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042798002076.

Atkinson:1990:SEC

[Atk90] A. C. Atkinson. A simulation experiment for comparing tests of regression transformation. Journal of Computational and Applied Mathematics, 31(1):11–14, July 24, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279090331S.

Alonso:1998:CDD

[ATV98] Ana Alonso, R. Loredana Trotta, and Alberto Valli. Coercive domain decomposition algorithms for advection–diffusion equations and systems. Journal of Computational and Applied Mathematics, 96(1): 51–76, September 3, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042798000910.

Aitchison:1996:SCN

[AU96] J. M. Aitchison and N. K.

Upton. Solution of constrained nonlinear equations in modelling the release of liquified gases. Journal of Computational and Applied Mathematics, 67(1):1– 14, February 20, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042794001200.

Androulakis:1996:OPS

[AV96] G. S. Androulakis and M. N. Vrahatis. OPTAC: a portable software package for analyzing and comparing optimization methods by visualization. Journal of Computational and Applied Mathematics, 72(1): 41–62, July 30, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042795002448.

Atakishiyev:1999:CVD

[AVW99] Natig M. Atakishiyev, Luis Edgar Vicent, and Kurt Bernardo Wolf. Continuous vs. discrete fractional Fourier transforms. Journal of Computational and Applied Mathematics, 107(1):73–95, July 15, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042799000825.

Agarwal:1995:EEB

[AW95a] Ravi P. Agarwal and Patricia J. Y. Wong. Explicit error bounds for the derivatives of piecewise-Lidstone interpolation. Journal of Computational and Applied Mathematics, 58(1):67–81, March 20, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042793E0262K.

Ambroladze:1995:ARP

[AW95b] Amiran Ambroladze and Hans Wallin. Approximation by repeated Padé approximants. Journal of Computational and Applied Mathematics, 62(3):353– 358, September 29, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042794001123.

Agarwal:1998:EPS

[AW98] Ravi P. Agarwal and Fu-Hsiang Wong. Existence of positive solutions for non-positive higher-order BVPs. Journal of Computational and Applied Mathematics, 88(1):3–14, February 23, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042797002112.

Axelsson:1991:MDC

[Axe91] O. Axelsson. The method of diagonal compensation of reduced matrix entries and multilevel iteration. Journal of Computational and Applied Mathematics, 38(1–3):31–43, December 23, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279190159H.

Axelsson:1996:SCM

[Axe96] O. Axelsson. The stabilized V -cycle method. Journal of Computational and Applied Mathematics, 74(1–2):33– 50, November 5, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042796000167.

Anastassiou:1994:BCW

[AY94] George Anastassiou and X. M. Yu. Bivariate constrained wavelet approximation. Journal of Computational and Applied Mathematics, 53(1):1– 9, July 29, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279200131R.

Amdeberhan:1997:ICR

[AZ97] Tewodros Amdeberhan and Melkamu Zeleke. On injectivity of the combinatorial Radon transform of order five. Journal of Computational and Applied Mathematics, 83(2):251–255, October 7, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704279700109X.

Arriola:1991:SPB

[AZD91] E. R. Arriola, A. Zarzo, and J. S. Dehesa. Spectral properties of the biconfluent Heun differential equation. Journal of Computational and Applied Mathematics, 37(1–3):161–169, November 18, 1991. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/037704279190114Y.

Bricher:1998:ABG

[BA98] Stephen Bricher and Evrim Akdoğu. The asymptotic behavior for gaseous ignition models. Journal of Computational and Applied Mathematics, 97(1–2):23– 37, September 24, 1998. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042798001010.

Babovsky:1996:RRP

[Bab96] Hans Babovsky. Recognition of ring pairs in the data analysis of Cherenkov detectors. Journal of Computational and Applied Mathematics, 72(2):227–234, August 13, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279500274X.

Babovsky:1998:CAS

[Bab98] Hans Babovsky. A constructive approach to steady nonlinear kinetic equations. Journal of Computational and Applied Mathematics, 89(2):199–211, March 16, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042797002264.

Babovsky:1999:KBL

[Bab99] Hans Babovsky. Kinetic boundary layers: on the adequate discretization of the Boltzmann collision operator. Journal of Computational and Applied Mathematics, 110(2):225–239, October 30, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704279900196X.

Babuska:1996:FEM

[BAG^{+}96] I. Babuska, B. Andersson, B. Guo, J. M. Melenk, and H. S. Oh. Finite element method for solving problems with singular solutions. Journal of Computational and Applied Mathematics, 74(1–2):51– 70, November 5, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0377042796000179.

Baheux:1995:NIL

[Bah95]  C. Baheux.  New im-

plementations of Lanczos method. Journal of Computational and Applied Mathematics, 57(1–2):3–15, February 20, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042793E0230J.

Bai:1997:CPC

[Bai97] Zhong-Zhi Bai. On the convergence of parallel chaotic nonlinear multisplitting Newton-type methods. Journal of Computational and Applied Mathematics, 80(2):317–334, May 5, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042797000368.

Bai:1998:CAP

[Bai98] Zhong-Zhi Bai. A class of asynchronous parallel nonlinear accelerated overrelaxation methods for the nonlinear complementarity problems. Journal of Computational and Applied Mathematics, 93(1):35–44, July 2, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042798002805.

Baker:1999:AAN

[Bak99] George A. Baker, Jr. Algebraic approximants and the numerical solution of parabolic equations. Journal of Computational and Applied Mathematics, 105(1– 2):129–139, May 1999. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042799000424.

Baltus:1993:TEB

[Bal93] C. Baltus. Truncation error bounds for the composition of limit-periodic linear fractional transformations. Journal of Computational and Applied Mathematics, 46(3):395– 404, June 22, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279390035A.

Bansch:1991:AFE

[Bän91] Eberhard Bänsch. An adaptive finite-element strategy for the three-dimensional time-dependent Navier– Stokes equations. Journal of Computational and Applied Mathematics, 36(1):3– 28, August 20, 1991. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042791902248.

Bangerezako:1999:FMA

[Ban99] Gaspard Bangerezako. The factorization method for the Askey–Wilson polynomials. Journal of Computational and Applied Mathematics, 107(2):219–232, July 31, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042799000904.

Barone:1990:DOT

[Bar90]  Piero Barone.  A dis-

crete orthogonal transform based on spherical harmonics. Journal of Computational and Applied Mathematics, 33(1):29–34, December 17, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279090253V.

Baratella:1997:NCP

[Bar97] Paola Baratella. A note on the convergence of product integration and Galerkin method for weakly singular integral equations. Journal of Computational and Applied Mathematics, 85(1):11–18, November 6, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042797001088.

Barnard:1999:AHF

[Bar99a] Roger W. Barnard. On applications of hypergeometric functions. Journal of Computational and Applied Mathematics, 105(1– 2):1–8, May 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042799000102.

Barrio:1999:CLD

[Bar99b] Roberto Barrio. Characterization of low degree Astable symmetric RK collocation methods. Journal of Computational and Applied Mathematics, 111(1–2):1– 11, November 15, 1999. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042799001272.

Batra:1998:ICC

[Bat98]  Prashant Batra.  Im-

provement of a convergence condition for the Durand–Kerner iteration. Journal of Computational and Applied Mathematics, 96(2):117–125, September 15, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042798001095.

Bavinck:1995:POR

[Bav95a] Herman Bavinck. On polynomials orthogonal with respect to an inner product involving differences. Journal of Computational and Applied Mathematics, 57(1–2):17–27, February 20, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042793E0231A.

Bavinck:1995:ZCL

[Bav95b] Herman Bavinck. On the zeros of certain linear combinations of Chebyshev polynomials. Journal of Computational and Applied Mathematics, 65(1–3):19–26, December 29, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042795000984.

Bavinck:1997:LPD

[Bav97] H. Bavinck. Linear perturbations of differential of difference operators with polynomials as eigenfunctions. Journal of Computational and Applied Mathematics, 78(1):179–195, February 3, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042796001409.

See erratum [Ano97-31].

Bavinck:1998:DDO

[Bav98a] H. Bavinck. Differential and difference operators having orthogonal polynomials with two linear perturbations as eigenfunctions. Journal of Computational and Applied Mathematics, 92(2):85–95, June 26, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042798000107.

Bavinck:1998:SCC

[Bav98b]  H. Bavinck.  On the

sum of the coefficients of certain linear differential operators. Journal of Computational and Applied Mathematics, 89(2):213– 217, March 16, 1998. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042797002318.

Bavinck:1998:SNR

[Bav98c] H. Bavinck. Some new results for Charlier and Meixner polynomials. Journal of Computational and Applied Mathematics, 90(1): 51–56, April 6, 1998. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042798000168.

Borwein:1993:CNT

[BB93] J. M. Borwein and P. B. Borwein. Class number three Ramanujan type series for 1/π. Journal of Computational and Applied Mathematics, 46(1–2):281– 290, June 14, 1993. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://docserver.carma.
                                        newcastle.edu.au/1560/;
                                         http://www.sciencedirect.
                                        com/science/article/pii/

037704279390302R. Com-

putational complex analysis.

Bohl:1995:BPD

[BB95] Erich Bohl and Winfried Boos. Binding proteindependent transporters: an answer of mathematics to biology. Journal of Computational and Applied Mathematics, 63(1–3):11– 25, November 20, 1995. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042795000550.

Bandle:1998:BDE

[BB98a] Catherine Bandle and Hermann Brunner. Blowup in diffusion equations: a survey. Journal of Computational and Applied Mathematics, 97(1–2):3–22, September 24, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042798001009.

Beckermann:1998:HCM

[BB98b] Bernhard Beckermann and Emmanuel Bourreau. How to choose modified moments? Journal of Computational and Applied Mathematics, 98(1):81–98, October 5, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042798001162.

Barrett:1996:ABI

[BBD^{+}96] Richard Barrett, Michael Berry, Jack Dongarra, Victor Eijkhout, and Charles Romine. Algorithmic bombardment for the iterative solution of linear systems: a poly-iterative approach. Journal of Computational and Applied Mathematics, 74(1–2):91–109, November 5, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042796000192.

Brucker:1990:CSI

[BBH90] Peter Brucker, Rainer E. Burkard, and Johann Hurink. Cyclic schedules for r irregularly occurring events. Journal of Computational and Applied Mathematics, 30(2): 173–189, May 28, 1990. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/037704279090026V.

Bainov:1996:FOS

[BBM96] Drumi Bainov, Cui Baotong, and Emil Minchev. Forced oscillation of solutions of certain hyperbolic equations of neutral type. Journal of Computational and Applied Mathematics, 72(2):309–318, August 13, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042796000027.

Brunner:1998:P

[BBS98] Hermann Brunner, Chris J.

Budd, and David M. Sloan. Preface. Journal of Computational and Applied Mathematics, 97(1–2):1–2, September 24, 1998. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042798000995.

Butnariu:1991:MAS

[BC91] Dan Butnariu and Yair Censor. A method for approximating the solution set of a system of convex inequalities by polytopes. Journal of Computational and Applied Mathematics, 36(3):289–304, September 24, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042791900129.

Burrage:1993:SOR

[BC93] Kevin Burrage and R. P. K.

Chan. On smoothing and order reduction effects for implicit Runge– Kutta formulae. Journal of Computational and Applied Mathematics, 45(1–2): 17–27, April 8, 1993. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0377042793902619.

Butnariu:1994:SCA

[BC94] Dan Butnariu and Yair Censor. Strong convergence of almost simultaneous blockiterative projection methods in Hilbert spaces. Journal of Computational and Applied Mathematics, 53(1):33–42, July 29, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279200123Q.

Berriochoa:1998:LSO

[BC98a] E. Berriochoa and A. Cachafeiro. Lebesgue Sobolev orthogonality on the unit circle. Journal of Computational and Applied Mathematics, 96(1):27–34, September 3, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042798000892.

Brighi:1998:AIC

[BC98b] Bernard Brighi and Michel Chipot. Approximation of infima in the calculus of variations. Journal of Computational and Applied Mathematics, 98(2):273–287, October 19, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042798001125.

Berriochoa:1999:FSO

[BC99] E. Berriochoa and A. Cachafeiro.

A family of Sobolev orthogonal polynomials on the unit circle. Journal of Computational and Applied Mathematics, 105(1–2):163– 173, May 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042799000400.

Baltus:1991:UCS

[BCCM91] C. Baltus, S. Clement Cooper, C. Craviotto, and J. H. McCabe. On the use of a corresponding sequence algorithm for δ-fractions. Journal of Computational and Applied Mathematics, 37(1–3):57–69, November 18, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279190106T.

Budd:1998:AND

[BCG98] C. J. Budd, G. J. Collins, and V. A. Galaktionov. An asymptotic and numerical description of selfsimilar blow-up in quasilinear parabolic equations. Journal of Computational and Applied Mathematics, 97(1–2):51–80, September 24, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042798001022.

Berndt:1999:RRC

[BCH^{+}99] Bruce C. Berndt, Heng Huat Chan, Sen-Shan Huang, Soon-Yi Kang, Jaebum Sohn, and Seung Hwan Son. The Rogers–Ramanujan continued fraction. Journal of Computational and Applied Mathematics, 105(1– 2):9–24, May 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042799000333.

Bravo:1999:DOS

[BCP99] Elba Bravo, Julio R. Claeyssen, and Rodrigo B. Platte. A direct onestep pressure actualization for incompressible flow with pressure Neumann condition. Journal of Computational and Applied Mathematics, 103(1):43–53, March 15, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042798002398.

Brunner:1991:FMA

[BCRV91] H. Brunner, M. R. Crisci, E. Russo, and A. Vecchio. A family of methods for Abel integral equations of the second kind. Journal of Computational and Applied Mathematics, 34(2): 211–219, April 4, 1991. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279190043J.

Bayada:1998:CMF

[BCV98] G. Bayada, M. Chambat, and C. Vázquez. Characteristics method for the formulation and computation of a free boundary cavitation problem. Journal of Computational and Applied Mathematics, 98(2):191–212, October 19, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042798001265.

Butcher:1993:P

[BCvdH93] John C. Butcher, Jeff R.

Cash, and Pieter J. van der Houwen. Preface. Journal of Computational and Applied Mathematics, 45(1–2): 1–2, April 8, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279390259E.

Baker:1990:RKF

[BD90] Christopher T. H. Baker and Mir S. Derakhshan. R–K formulae applied to Volterra equations with delay. Journal of Computational and Applied Mathematics, 29(3): 293–310, March 15, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        037704279090013P.

Bona:1996:EDS

[BDKM96] J. L. Bona, V. A. Dougalis, O. A. Karakashian, and W. R. McKinney. The effect of dissipation on solutions of the generalized Korteweg–de Vries equation. Journal of Computational and Applied Mathematics, 74(1–2):127–154, November 5, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042796000210.

Bultheel:1997:QHLa

[BDMGVO97a] A. Bultheel, C. Díaz-Mendoza, P. González-Vera, and R. Orive. Quadrature on the half-line and twopoint Padé approximants to Stieltjes functions — II:

Convergence. Journal of Computational and Applied Mathematics, 77(1–2):53– 76, January 13, 1997. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042796001227.

Bultheel:1997:QHLb

[BDMGVO97b] A. Bultheel, C. Díaz-Mendoza, P. González-Vera, and R. Orive. Quadrature on the half line and two-point Padé approximants to Stieltjes functions. Part III. The unbounded case. Journal of Computational and Applied Mathematics, 87(1):95–117, December 18, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042797001805.

Bainov:1996:LQI

[BDS96] D. D. Bainov, A. B.

Dishliev, and I. M. Stamova. Lipschitz quasistability of impulsive differentialdifference equations with variable impulsive perturbations. Journal of Computational and Applied Mathematics, 70(2):267– 277, June 28, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042795002073.

Bocher:1994:GMG

[BDV94] P. Bocher, H. De Meyer, and G. Vanden Berghe. On Gregory- and modified Gregory-type corrections to Newton–Cotes quadrature. Journal of Computational and Applied Mathematics, 50(1–3):145–158, May 20, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042794902968.

Berndt:1991:SEA

[BE91] Bruce C. Berndt and Ronald J. Evans. Some elegant approximations and asymptotic formulas of Ramanujan. Journal of Computational and Applied Mathematics, 37(1–3):35– 41, November 18, 1991. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/037704279190104R.

Bai:1998:CIM

[BE98] Zhong-Zhi Bai and D. J. Evans. Chaotic iterative methods for the linear complementarity problems. Journal of Computational and Applied Mathematics, 96(2):127–138, September 15, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042798001113.

Beardon:1999:KHP

[Bea99] A. F. Beardon. The Klein, Hilbert and Poincaré metrics of a domain. Journal of Computational and Applied Mathematics, 105(1–2):155– 162, May 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042799000126.

Beckermann:1990:SSS

[Bec90] Bernhard Beckermann. The structure of the singular solution table of the M-Padé approximation problem. Journal of Computational and Applied Mathematics, 32(1–2):3–15, November 26, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279090411R.

Beckermann:1992:RMC

[Bec92] Bernhard Beckermann. A reliable method for computing M-Padé approximants on arbitrary staircases. Journal of Computational and Applied Mathematics, 40(1):19–42, June 12, 1992. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
037704279290039Z.

Beck:1994:ADI

[Bec94]  Thomas Beck.  Auto-

matic differentiation of iterative processes. Journal of Computational and Applied Mathematics, 50(1–3):109– 118, May 20, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042794902933.

Bai:1999:CAM

[BEC99] Zhong-Zhi Bai, D. J. Evans, and R. C. Calinescu. A class of asynchronous multisplitting two-stage iterations for large sparse block systems of weakly nonlinear equations. Journal of Computational and Applied Mathematics, 110(2):271–286, October 30, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042799002368.

Bellalij:1990:SPC

[Bel90a] M. Bellalij. A simultaneous process for convergence acceleration and error control. Journal of Computational and Applied Mathematics, 33(2):217–231, December 21, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279090370F.

Belmehdi:1990:AOP

[Bel90b] S. Belmehdi. On the associated orthogonal polynomials. Journal of Computational and Applied Mathematics, 32(3):311–319, December 10, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279090041W.

Belmehdi:1993:ALP

[Bel93a] S. Belmehdi. On Appell– Laguerre polynomials. Journal of Computational and Applied Mathematics, 49 (1–3):11–17, December 31, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
037704279390129Y.

Brown:1993:OPE

[BEL93b] B. M. Brown, W. D. Evans, and L. L. Littlejohn. Orthogonal polynomials and extensions of Copson’s inequality. Journal of Computational and Applied Mathematics, 48(1–2):33–48, October 29, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042793903142.

Buttgenbach:1992:SEE

[BELN92] B. Büttgenbach, H. Esser, G. Lüttgens, and R. J. Nessel. A sharp error estimate for the numerical solution of a Dirichlet problem for the Poisson equation. Journal of Computational and Applied Mathematics, 44(3):331–337, December 24, 1992. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279290004H.

Brown:1997:SCP

[BEM97a] B. M. Brown, M. S. P. Eastham, and D. K. R. McCormack. Spectral concentration and perturbed discrete spectra. Journal of Computational and Applied Mathematics, 86(2):415–425, December 10, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042797001738.

Brown:1997:SCR

[BEM97b] B. M. Brown, M. S. P. Eastham, and D. K. R. Mc-Cormack. Spectral concentration and rapidly decaying potentials. Journal of Computational and Applied Mathematics, 81(2): 333–348, July 8, 1997. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042797000721.

Brown:1998:ACS

[BEM98] B. M. Brown, M. S. P. Eastham, and D. K. R. McCormack. Absolute continuity and spectral concentration for slowly decaying potentials. Journal of Computational and Applied Mathematics, 94(2):181–197, August 3, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042798000879.

Benchiboun:1994:LCV

[Ben94a] M. D. Benchiboun. Linear convergence for vector sequences and some applications. Journal of Computational and Applied Mathematics, 55(1):81–97, October 31, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042794901864.

Bentbib:1994:ACI

[Ben94b] A. H. Bentbib. Acceleration of convergence of interval sequences. Journal of Computational and Applied Mathematics, 51(3): 395–409, June 30, 1994. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/037704279200120X.

Benchiboun:1996:EHM

[Ben96] M. D. Benchiboun. Extension of Henrici’s method to matrix sequences. Journal of Computational and Applied Mathematics, 75(1):1– 21, November 12, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0377042796000313.

Benzi:1997:RNS

[Ben97] M. Benzi. Remarks on the numerical solution of certain linear complementarity problems. Journal of Computational and Applied Mathematics, 83 (1):137–143, September 30, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042797001039.

Berg:1992:AEC

[Ber92a] Lothar Berg. The asymptotic expansion of a convolution integral. Journal of Computational and Applied Mathematics, 41(1–2):159– 161, August 20, 1992. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279290245S.

Berthold:1992:SCC

[Ber92b] Dietmar Berthold. On the stability of the classical collocation method for Cauchytype singular integral equations. Journal of Computational and Applied Mathematics, 43(1–2):171–184, November 25, 1992. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279290265Y.

Berg:1993:DGR

[Ber93a] Lothar Berg. On the discretization of generalized Riccati differential equations with constant coefficients. Journal of Computational and Applied Mathematics, 47(3):277– 283, September 30, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
037704279390057I.

Berrut:1993:FOI

[Ber93b] Jean-Paul Berrut. A formula for optimal integration in H ^{2}. Journal of Computational and Applied Mathematics, 46(1–2):199– 210, June 14, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279390295M.

Bernadou:1994:CFE

[Ber94] M. Bernadou. C ^{1}-curved finite elements and applications to plate and shell problems. Journal of Computational and Applied Mathematics, 50(1–3):133–144, May 20, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279490295X.

Berg:1995:IMP

[Ber95] Christian Berg. Indeterminate moment problems and the theory of entire functions. Journal of Computational and Applied Mathematics, 65(1–3):27–55, December 29, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042795000992.

Berrut:1997:BWR

[Ber97]  Jean-Paul Berrut.  The

barycentric weights of rational interpolation with prescribed poles. Journal of Computational and Applied Mathematics, 86(1):45–52, November 28, 1997. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042797001477.

Berg:1998:SIM

[Ber98] Christian Berg. On some indeterminate moment problems for measures on a geometric progression. Journal of Computational and Applied Mathematics, 99(1–2): 67–75, November 16, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042798001460.

Berger:1999:REB

[Ber99] Arno Berger. Rigorous error bounds for RK methods in the proof of chaotic behaviour. Journal of Computational and Applied Mathematics, 111(1–2):13– 24, November 15, 1999. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042799001284.

Berntsen:1991:SSA

[BES91] Jarle Berntsen, Terje O. Espelid, and Tor Sørevik. On the subdivision strategy in adaptive quadrature algorithms. Journal of Computational and Applied Mathematics, 35(1–3):119–132, June 26, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279190201T.

Bessis:1996:PAN

[Bes96] Daniel Bessis. Padé approximations in noise filtering. Journal of Computational and Applied Mathematics, 66(1–2):85–88, January 31, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042795001778.

Bialecki:1993:MDA

[BF93] Bernard Bialecki and Graeme Fairweather. Matrix decomposition algorithms for separable elliptic boundary value problems in two space dimensions. Journal of Computational and Applied Mathematics, 46(3): 369–386, June 22, 1993. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0377042793900338.

Beck:1994:IPA

[BF94a] Thomas Beck and Herbert Fischer. The if-problem in automatic differentiation. Journal of Computational and Applied Mathematics, 50(1–3):119–131, May 20, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042794902941.

Busby:1994:QOE

[BF94b] Robert C. Busby and Wyman Fair. Quadratic operator equations and periodic operator continued fractions. Journal of Computational and Applied Mathematics, 54(3):377– 387, October 20, 1994. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042794902585.

Bai:1996:SLS

[BFG96] Zhaojun Bai, Mark Fahey, and Gene Golub. Some large-scale matrix computation problems. Journal of Computational and Applied Mathematics, 74(1–2):71– 89, November 5, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042796000180.

Baratella:1995:UAU

[BG95] Paola Baratella and Bruno Gabutti. Uniform approximation on unit disk of series arising in plate contact problems. Journal of Computational and Applied Mathematics, 62(2):181– 196, September 20, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042794001019.

Bosbach:1998:SAL

[BG98] Christof Bosbach and Wolfgang Gawronski. Strong asymptotics for Laguerre polynomials with varying weights. Journal of Computational and Applied Mathematics, 99(1–2):77– 89, November 16, 1998. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042798001472.

Bunse-Gerstner:1997:EBI

[BGF97] A. Bunse-Gerstner and H. Faßbender. Error bounds in the isometric Arnoldi process. Journal of Computational and Applied Mathematics, 86(1):53–72, November 28, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042797001489.

Barrenechea:1998:WSI

[BGH98] Gabriel R. Barrenechea, Gabriel N. Gatica, and George C. Hsiao. Weak solvability of interior transmission problems via mixed finite elements and Dirichletto-Neumann mappings. Journal of Computational and Applied Mathematics, 100 (2):145–160, December 21, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704279800185X.

Baker:1990:DUI

[BGM90] George A. Baker, Jr.

and P. R. Graves-Morris. Definition and uniqueness of integral approximants. Journal of Computational and Applied Mathematics, 31(3):357–372, October 1, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279090036Y.

Baker:1993:DSI

[BGM93a] George A. Baker, Jr. and P. R. Graves-Morris. Determinants and the structure of the integral approximant table. Journal of Computational and Applied Mathematics, 47(1):67–81, June 28, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279390091O.

Baker:1993:RSH

[BGM93b] George A. Baker, Jr. and P. R. Graves-Morris. Results on sequences of Hermite–Padé integral approximants. Journal of Computational and Applied Mathematics, 47(2):241– 252, ???? 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279390006W.

Baker:1994:ECS

[BGM94] George A. Baker, Jr. and P. R. Graves-Morris. Existence of certain sequences of Hermite–Padé approximants. Journal of Computational and Applied Mathematics, 54(1):45–49, September 20, 1994. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279490393X.

Baker:1995:SCD

[BGM95] George A. Baker, Jr. and P. R. Graves-Morris. Some convergence and divergence theorems for Hermite–Padé approximants. Journal of Computational and Applied Mathematics, 59(3): 285–293, May 30, 1995. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279400035Y.

Bellavia:1996:PUS

[BGM96a] S. Bellavia, M. G. Gasparo, and M. Macconi. Partially Updated Switching-Method for systems of nonlinear equations. Journal of Computational and Applied Mathematics, 76(1–2):77– 88, December 17, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S037704279600091X.

Bellavia:1996:SMN

[BGM96b] S. Bellavia, M. G. Gasparo, and M. Macconi. A Switching-Method for nonlinear systems. Journal of Computational and Applied Mathematics, 71(1):83–93, July 10, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279500226X.

Bonazzola:1999:SMG

[BGM99] S. Bonazzola, E. Gourgoulhon, and J.-A. Marck. Spectral methods in general relativistic astrophysics. Journal of Computational and Applied Mathematics, 109(1–2):433–473, September 30, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042799001673.

Babovsky:1990:AWD

[BGN^{+}90] Hans Babovsky, Frank Gropengießer, Helmut Neunzert, Jens Struckmeier, and Bernd Wiesen. Application of well-distributed sequences to the numerical simulation of the Boltzmann equation. Journal of Computational and Applied Mathematics, 31(1):15–22, July 24, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279090332T.

Broeckx:1991:P

[BGPW91] F. Broeckx, M. J. Goovaerts, R. Piessens, and L. Wuytack. Preface. Journal of Computational and Applied Mathematics, 35(1–3): 1, June 26, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279190194O.

Broeckx:1994:E

[BGPW94a] F. Broeckx, M. J. Goovaerts, R. Piessens, and L. Wuytack. Editorial. Journal of Computational and Applied Mathematics, 50(1–3): 1–3, May 20, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042794902852.

Broeckx:1994:P

[BGPW94b] F. Broeckx, M. J. Goovaerts, R. Piessens, and L. Wuytack. Preface. Journal of Computational and Applied Mathematics, 50(1–3):5–6, May 20, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042794902860.

Broeckx:1996:P

[BGPW96] F. Broeckx, M. J. Goovaerts, R. Piessens, and L. Wuytack. Preface. Journal of Computational and Applied Mathematics, 66(1–2):1–2, January 31, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042796900199.

Broeckx:1997:P

[BGPW97] F. Broeckx, M. J. Goovaerts, R. Piessens, and L. Wuytack. Preface. Journal of Computational and Applied Mathematics, 82(1–2):1–2, September 15, 1997. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042797000897.

Bultheel:1993:MPO

[BGVHN93] A. Bultheel, P. González-Vera, E. Hendriksen, and O. Njåstad. Moment problems and orthogonal functions. Journal of Computational and Applied Mathematics, 48(1–2):49–68, October 29, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042793903153.

Bultheel:1995:CMA

[BGVHN95a] A. Bultheel, P. González-Vera, E. Hendriksen, and O. Njåstad. Convergence of modified approximants associated with orthogonal rational functions. Journal of Computational and Applied Mathematics, 57(1–2):77– 86, February 20, 1995. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042793E0235E.

Bultheel:1995:FTR

[BGVHN95b] Adhemar Bultheel, Pablo González-Vera, Erik Hendriksen, and Olav Njåstad. Favard theorem for reproducing kernels. Journal of Computational and Applied Mathematics, 57(1–2):57– 76, February 20, 1995. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042793E0234D.

Bultheel:1997:RCM

[BGVHN97] A. Bultheel, P. González-Vera, E. Hendriksen, and O. Njåstad. Rates of convergence of multipoint rational approximants and quadrature formulas on the unit circle. Journal of Computational and Applied Mathematics, 77(1–2):77– 101, January 13, 1997. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042796001239.

Bultheel:1999:DPO

[BGVHN99] A. Bultheel, P. González-Vera, E. Hendriksen, and O. Njåstad. A density problem for orthogonal rational functions. Journal of Computational and Applied Mathematics, 105(1–2):199– 212, May 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042799000382.

Bultheel:1995:QHL

[BGVO95] A. Bultheel, P. González-Vera, and R. Orive. Quadrature on the half-line and two-point Padé approximants to Stieltjes functions. Part I. Algebraic aspects. Journal of Computational and Applied Mathematics, 65(1–3):57–72, December 29, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279500100X.

Beckers:1991:CTD

[BH91] Marc Beckers and Ann Haegemans. The construction of three-dimensional invariant cubature formulae. Journal of Computational and Applied Mathematics, 35(1–3):109–118, June 26, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042791902004.

Bihari:1995:AGW

[BH95] Barna L. Bihari and Ami Harten. Application of generalized wavelets: an adaptive multiresolution scheme. Journal of Computational and Applied Mathematics, 61(3):275–321, August 21, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042794000701.

Bai:1998:AMT

[BH98] Zhong-Zhi Bai and Yu-Guang Huang. Asynchronous multisplitting twostage iterations for systems of weakly nonlinear equations. Journal of Computational and Applied Mathematics, 93(1):13–33, July 2, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042798000636.

Bavinck:1993:AGP

[BHD93] Herman Bavinck, Gerard Hooghiemstra, and Ellen De Waard. An application of Gegenbauer polynomials in queueing theory. Journal of Computational and Applied Mathematics, 49(1–3):1–10, December 31, 1993. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279390128X.

Bultheel:1994:QFU

[BHGVN94] Adhemar Bultheel, Erik Hendriksen, Pablo González-Vera, and Olav Njåstad. Quadrature formulas on the unit circle based on rational functions. Journal of Computational and Applied Mathematics, 50(1–3):159– 170, May 20, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042794902976.

Bonan-Hamada:1999:POL

[BHJ99] Catherine M. Bonan-Hamada and William B. Jones. Paraorthogonal Laurent polynomials and the strong Stieltjes moment problem. Journal of Computational and Applied Mathematics, 105(1–2):175–185, May 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042799000242.

Berthold:1992:FAS

[BHS92] D. Berthold, W. Hoppe, and B. Silbermann. A fast algorithm for solving the generalized airfoil equation. Journal of Computational and Applied Mathematics, 43(1–2):185–219, November 25, 1992. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279290266Z.

Babuska:1998:AOM

[BHS^{+}98] I. Babuska, Jens Hugger, T. Strouboulis, K. Copps, and S. K. Gangaraj. The asymptotically optimal meshsize function for bi-p degree interpolation over rectangular elements. Journal of Computational and Applied Mathematics, 90(2):185– 221, April 17, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042798000181.

Buhmann:1992:OPT

[BI92] M. D. Buhmann and A. Iserles. On orthogonal polynomials transformed by the QR algorithm. Journal of Computational and Applied Mathematics, 43(1–2):117– 134, November 25, 1992. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
037704279290262V.

Bokhari:1996:ARV

[BI96] M. A. Bokhari and M. Iqbal.

L_{2}-approximation of realvalued functions with interpolatory constraints. Journal of Computational and Applied Mathematics, 70(2): 201–205, June 28, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042795002103.

Butcher:1993:ELT

[BJ93] J. C. Butcher and P. B.

Johnston. Estimating local truncation errors for Runge– Kutta methods. Journal of Computational and Applied Mathematics, 45(1–2):203– 212, April 8, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279390275G.

Blanes:1999:CNS

[BJ99] Sergio Blanes and Lucas Jódar. Continuous numerical solutions of coupled mixed partial differential systems using Fer’s factorization. Journal of Computational and Applied Mathematics, 101(1–2):189–202, January 15, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042798002192.

Butcher:1997:NOA

[BJM97] J. C. Butcher, Z. Jackiewicz, and H. D. Mittelmann. A nonlinear optimization approach to the construction of general linear methods of high order. Journal of Computational and Applied Mathematics, 81(2): 181–196, July 8, 1997. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042797000393.

Bellen:1993:TPR

[BJZ93] A. Bellen, Z. Jackiewicz, and M. Zennaro. Timepoint relaxation Runge– Kutta methods for ordinary differential equations. Journal of Computational and Applied Mathematics, 45(1–2):121–137, April 8, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279390269H.

Barrlund:1990:ANS

[BK90] Anders Barrlund and Bo Kågström.

Analytical and numerical solutions to higher index linear variable coefficient DAE systems. Journal of Computational and Applied Mathematics, 31(3):305–330, October 1, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279090033V.

Bartos:1995:NSC

[BK95a] Miroslav Bartos and Zdenek Kestránek. Numerical solution of the contact problem. Application to a simple model of the human hip joint. Journal of Computational and Applied Mathematics, 63(1–3):439–447, November 20, 1995. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042795000690.

Blaheta:1995:SLN

[BK95b] R. Blaheta and R. Kohut. Solution of large nonlinear systems in elastoplastic incremental analysis. Journal of Computational and Applied Mathematics, 63(1–3):255–264, November 20, 1995. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0377042795000798.

Bavinck:1999:DOH

[BK99a] H. Bavinck and J. Koekoek. Differential operators having symmetric orthogonal polynomials as eigenfunctions. Journal of Computational and Applied Mathematics, 106(2):369–393, June 30, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042799000941.

Bialecki:1999:SQS

[BK99b] B. Bialecki and P. Keast. A sinc quadrature subroutine for Cauchy principal value integrals. Journal of Computational and Applied Mathematics, 112(1–2):3– 20, November 30, 1999. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042799002198.

Belytschko:1996:SAC

[BKF^{+}96] Ted Belytschko, Yury Krongauz, Mark Fleming, Daniel Organ, and Wing Kam Snm Liu. Smoothing and accelerated computations in the element free Galerkin method. Journal of Computational and Applied Mathematics, 74(1–2):111–126, November 5, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042796000209.

Bainov:1995:CPI

[BKM95] Drumi D. Bainov, Zdzisław Kamont, and Emil Minchev. Comparison principles for impulsive hyperbolic equations of first order. Journal of Computational and Applied Mathematics, 60(3): 379–388, June 30, 1995. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042794000464.

Bainov:1996:MIM

[BKM96] Drumi D. Bainov, Zdzisław Kamont, and Emil Minchev. Monotone iterative methods for impulsive hyperbolic differential-functional equations. Journal of Computational and Applied Mathematics, 70(2):329– 347, June 28, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279500209X.

Bustamante:1993:RCH

[BL93] J. Bustamante and G. López Lagomasino. Rate of convergence of Hermite–Padé approximants to a Nikishintype system of analytic functions. Journal of Computational and Applied Mathematics, 49(1–3):19–25, December 31, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042793901304.

Beckermann:1997:RMR

[BL97] Bernhard Beckermann and George Labahn. Recursiveness in matrix rational interpolation problems. Journal of Computational and Applied Mathematics, 77(1–2): 5–34, January 13, 1997. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042796001203.

Blacher:1995:CLT

[Bla95]  René Blacher.  Central

limit theorem by polynomial dependence coefficients. Journal of Computational and Applied Mathematics, 57(1–2):45– 56, February 20, 1995. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042793E0233C.

Bauwelinckx:1991:NAL

[BLG91] T. Bauwelinckx, E. Labie, and M. J. Goovaerts. A new approach for loaded credibility premiums. Journal of Computational and Applied Mathematics, 37(1–3):301– 314, November 18, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042791901265.

Bernardi:1992:PIR

[BM92] Christine Bernardi and Yvon Maday. Polynomial interpolation results in Sobolev spaces. Journal of Computational and Applied Mathematics, 43(1–2):53– 80, November 25, 1992. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279290259Z.

Bessis:1993:OPA

[BM93a] Daniel Bessis and Giorgio Mantica. Orthogonal polynomials associated to almost periodic Schrödinger operators. A trend towards random orthogonal polynomials. Journal of Computational and Applied Mathematics, 48(1–2):17–32, October 29, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279390313Z.

Brezinski:1993:LSO

[BM93b] Claude Brezinski and Ana C. Matos. Least-squares orthogonal polynomials. Journal of Computational and Applied Mathematics, 46(1– 2):229–239, June 14, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        037704279390298P.

Bohl:1995:MA

[BM95a] Erich Bohl and Ivo Marek. A model of amplification. Journal of Computational and Applied Mathematics, 63(1–3):27–47, November 20, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042795000526.

Bohl:1995:NMI

[BM95b] Erich Bohl and Ivo Marek. A nonlinear model involving M-operators. An amplification effect measured in the cascade of vision. Journal of Computational and Applied Mathematics, 60(1–2): 13–28, June 20, 1995. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279400081B.

Bainov:1996:OSI

[BM96a] Drumi Bainov and Emil Minchev. Oscillation of the solutions of impulsive parabolic equations. Journal of Computational and Applied Mathematics, 69(2): 207–214, May 23, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042795000402.

Bertschy:1996:GAH

[BM96b] R. Bertschy and P. A. Monney. A generalization of the algorithm of heidtmann to non-monotone formulas. Journal of Computational and Applied Mathematics, 76(1–2):55–76, December 17, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042796000891.

Brezinski:1996:DEA

[BM96c] Claude Brezinski and Ana C.

Matos. A derivation of extrapolation algorithms based on error estimates. Journal of Computational and Applied Mathematics, 66(1–2):5–26, January 31, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042795001573.

Berrut:1997:MDD

[BM97] Jean-Paul Berrut and Hans D. Mittelmann. Matrices for the direct determination of the barycentric weights of rational interpolation. Journal of Computational and Applied Mathematics, 78(2):355–370, February 17, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704279600163X.

Beghdadi:1998:IPP

[BM98] Driss Beghdadi and Pascal Maroni. On the inverse problem of the product of a semi-classical form by a polynomial. Journal of Computational and Applied Mathematics, 88(2):377– 399, March 2, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042797002276.

Boros:1999:IHG

[BM99] George Boros and Victor H. Moll. An integral hidden in gradshteyn and ryzhik. Journal of Computational and Applied Mathematics, 106(2):361– 368, June 30, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042799000813.

Bielawski:1996:CAR

[BMP96] Stanislav S. Bielawski, Stepan G. Mulyarchik, and Andrew V. Popov. The construction of an algebraically reduced system for the acceleration of preconditioned conjugate gradients. Journal of Computational and Applied Mathematics, 70(2): 189–200, June 28, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042795002049.

Bracciali:1999:SSD

[BMR99] C. F. Bracciali, J. H. Mc-Cabe, and A. Sri Ranga. On a symmetry in strong distributions. Journal of Computational and Applied Mathematics, 105(1–2):187– 198, May 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042799000461.

Bainov:1998:ABN

[BMS98] D. D. Bainov, N. T.

Markova, and P. S. Simeonov. Asymptotic behaviour of the nonoscillatory solutions of differential equations of second order with delay depending on the unknown function. Journal of Computational and Applied Mathematics, 91(1):87–96, April 24, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042798000247.

Brezinski:1995:I

[BMV95] C. Brezinski, P. Maroni, and J. Van Iseghem. Introduction. Journal of Computational and Applied Mathematics, 57(1–2):1, February 20, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042795900330.

Benner:1997:NMC

[BMX97] Peter Benner, Volker Mehrmann, and Hongguo Xu. A new method for computing the stable invariant subspace of a real Hamiltonian matrix. Journal of Computational and Applied Mathematics, 86(1):17–43, November 28, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042797001465.

Bramson:1992:CSR

[BN92] Maury Bramson and Claudia Neuhauser. A catalytic surface reaction model. Journal of Computational and Applied Mathematics, 40(2):157–161, June 24, 1992. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042792901035.

Bock:1995:KEV

[Boc95] Igor Bock. On von Kárman equations for viscoelastic plates. Journal of Computational and Applied Mathematics, 63(1–3):277–282, November 20, 1995. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042795000828.

Bockmann:1996:CFI

[Böc96] Christine Böckmann. Curve fitting and identification of physical spectra. Journal of Computational and Applied Mathematics, 70(2): 207–224, June 28, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042795002022.

Boersma:1991:UAB

[Boe91]  J. Boersma.  Uniform

asymptotics of a Besselfunction series occurring in a transmission-line problem. Journal of Computational and Applied Mathematics, 37(1–3):143–159, November 18, 1991. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279190113X.

Boglaev:1997:IAD

[Bog97] Igor P. Boglaev. Iterative algorithms of domain decomposition for the solution of a quasilinear elliptic problem. Journal of Computational and Applied Mathematics, 80(2): 299–316, May 5, 1997. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042797000320.

Boglaev:1998:DDA

[Bog98] Igor Boglaev. On a domain decomposition algorithm for a singularly perturbed reaction–diffusion problem. Journal of Computational and Applied Mathematics, 98(2):213–232, October 19, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042798001277.

Bojanov:1996:QFM

[Boj96] Borislav Bojanov. On a quadrature formula of micchelli and Rivlin. Journal of Computational and Applied Mathematics, 70(2): 349–356, June 28, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042795002065.

Bai:1997:MMM

[BOS97] Zhaojun Bai, Antonio Orlandi, and William T. Smith. Moment matching method for numerical solution of MTL equations in interconnect analysis. Journal of Computational and Applied Mathematics, 86(1):3– 16, November 28, 1997. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042797821218.

Bouleau:1990:ECE

[Bou90] Nicolas Bouleau. On effective computation of expectations in large or infinite dimension. Journal of Computational and Applied Mathematics, 31(1):23–34, July 24, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279090333U.

Brezinski:1991:OLC

[BP91] Claude Brezinski and Stefan Paszkowski. Optimal linear contractive sequence transformations. Journal of Computational and Applied Mathematics, 38(1–3):45– 59, December 23, 1991. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/037704279190160L.

Boneh:1996:GAE

[BP96a] Shahar Boneh and Vassilis G. Papanicolaou. General asymptotic estimates for the coupon collector problem. Journal of Computational and Applied Mathematics, 67(2):277– 289, March 29, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0377042795000208.

Bramble:1996:LSM

[BP96b] James H. Bramble and Joseph E. Pasciak. Leastsquares methods for Stokes equations based on a discrete minus one inner product. Journal of Computational and Applied Mathematics, 74(1–2):155–173, November 5, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042796000222.

Bojanov:1997:MCF

[BP97a] Borislav Bojanov and Guergana Petrova. On minimal cubature formulae for product weight functions. Journal of Computational and Applied Mathematics, 85(1):113–121, November 6, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042797001337.

Borcea:1997:HNM

[BP97b] Liliana Borcea and George C.

Papanicolaou. A hybrid numerical method for high contrast conductivity problems. Journal of Computational and Applied Mathematics, 87(1):61–77, December 18, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042797001751.

Bainov:1997:EAB

[BPP97] D. D. Bainov, V. A. Petrov, and V. S. Proytcheva. Existence and asymptotic behaviour of nonoscillatory solutions of second-order neutral differential equations with “maxima”. Journal of Computational and Applied Mathematics, 83(2):237– 249, October 7, 1997. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042797001052.

Barnard:1993:APS

[BPV93] Roger W. Barnard, Kent Pearce, and Richard S. Varga. An application from partial sums of e^{z} to a problem in several complex variables. Journal of Computational and Applied Mathematics, 46(1–2):271–279, June 14, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279390301Q.

Bertram:1992:PIF

[BR92a] B. Bertram and O. Ruehr. Product integration for finite-part singular integral equations: Numerical asymptotics and convergence acceleration. Journal of Computational and Applied Mathematics, 41(1–2): 163–173, August 20, 1992. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        037704279290246T.

Butt:1992:OSD

[BR92b] R. Butt and J. E. Rubio. Optimal shape design for a frictionless contact problem. Journal of Computational and Applied Mathematics, 40(1):1–17, June 12, 1992. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279290038Y.

Baxley:1998:NBV

[BR98] John V. Baxley and Stephen B.

Robinson. Nonlinear boundary value problems for shallow membrane caps, II. Journal of Computational and Applied Mathematics, 88(1):203–224, February 23, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042797002161.

Brenier:1990:MPP

[Bre90a] Yann Brenier. Une méthode particulaire pour les équations non linéaires de diffusion convection en dimension un. (French) [A particle method for nonlinear convection diffusion equations in dimension one]. Journal of Computational and Applied Mathematics, 31(1): 35–56, July 24, 1990. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279090334V.

Brezinski:1990:DPC

[Bre90b] Claude Brezinski. A direct proof of the Christoffel– Darboux identity and its equivalence to the recurrence relationship. Journal of Computational and Applied Mathematics, 32(1–2): 17–25, November 26, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        037704279090412S.

Brezinski:1990:DPT

[Bre90c]  Claude Brezinski.  Du-

ality in Padé-type approximation. Journal of Computational and Applied Mathematics, 30(3):351– 357, July 10, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042790902858.

Brezinski:1990:P

[Bre90d] Claude Brezinski. Preface. Journal of Computational and Applied Mathematics, 32(1–2):1, November 26, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042790904102.

Brezinski:1997:PML

[Bre97] C. Brezinski. Projection methods for linear systems. Journal of Computational and Applied Mathematics, 77(1–2):35–51, January 13, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042796001215.

Brezinski:1998:VST

[Bre98]  C. Brezinski.  Vector

sequence transformations: Methodology and applications to linear systems. Journal of Computational and Applied Mathematics, 98(1):149–175, October 5, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042798001198.

Brunner:1994:CCI

[Bru94] Hermann Brunner. Collocation and continuous implicit Runge–Kutta methods for a class of delay Volterra integral equations. Journal of Computational and Applied Mathematics, 53(1):61–72, July 29, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279200125S.

Brutman:1996:VLI

[Bru96]  L. Brutman.  On the

Vykhandu–Levin iterative method for numerical solution of nonlinear systems. Journal of Computational and Applied Mathematics, 70(1):51–56, June 14, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042795001360.

Barlow:1993:SCS

[BS93a] Richard E. Barlow and Fabio Spizzichino. Schurconcave survival functions and survival analysis. Journal of Computational and Applied Mathematics, 46(3): 437–447, June 22, 1993. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279390039E.

Bhatta:1993:SSP

[BS93b] Sukanta Kumar Bhatta and K. Suryanarayana Sastri. Symmetric spline procedures for boundary value problems with mixed boundary conditions. Journal of Computational and Applied Mathematics, 45(3): 237–250, May 3, 1993. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279390043B.

Butzer:1994:STA

[BS94] P. L. Butzer and G. Schöttler.

Sampling theorems associated with fourth- and higher-order self-adjoint eigenvalue problems. Journal of Computational and Applied Mathematics, 51(2): 159–177, June 20, 1994. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042792000107.

Bainov:1995:UAS

[BS95] D. D. Bainov and I. M.

Stamova. Uniform asymptotic stability of impulsive differential-difference equations of neutral type by Lyapunov’s direct method. Journal of Computational and Applied Mathematics, 62(3):359–369, September 29, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042794001093.

Bainov:1998:PSS

[BS98a] Drumi Bainov and Pavel Simeonov. Positive solutions of a superlinear first-order differential equation with delay depending on the unknown function. Journal of Computational and Applied Mathematics, 88(1):95–101, February 23, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042797002136.

Bramble:1998:LFF

[BS98b] James H. Bramble and Tong Sun. A locking-free finite element method for Naghdi shells. Journal of Computational and Applied Mathematics, 89(1):119– 133, March 9, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042797002343.

Baratchart:1999:NUR

[BSW99] L. Baratchart, Herbert Stahl, and F. Wielonsky. Non-uniqueness of rational best approximants. Journal of Computational and Applied Mathematics, 105(1– 2):141–154, May 1999. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042799000321.

Boutayeb:1991:FDM

[BT91] A. Boutayeb and E. H. Twizell. Finite-difference methods for twelfth-order. Journal of Computational and Applied Mathematics, 35(1–3):133–138, June 26, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
037704279190202U.

Boos:1992:CTC

[BT92a] Johann Boos and Hubert Tietz. Convexity theorems for the circle methods of summability. Journal of Computational and Applied Mathematics, 40(2): 151–155, June 24, 1992. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042792901024.

Bruaset:1992:RPC

[BT92b] A. M. Bruaset and A. Tveito.

RILU preconditioning; a computational study. Journal of Computational and Applied Mathematics, 39(3): 259–275, May 8, 1992. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279290203A.

Brunelli:1995:SMA

[BT95] R. Brunelli and G. P. Tecchiolli. Stochastic minimization with adaptive memory. Journal of Computational and Applied Mathematics, 57(3):329–343, February 28, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042793E0203X.

Brugnano:1996:CSB

[BT96] L. Brugnano and D. Trigiante. Convergence and stability of boundary value methods for ordinary differential equations. Journal of Computational and Applied Mathematics, 66(1–2):97– 109, January 31, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0377042795001662.

Bailey:1999:DPM

[BTCC99] C. Bailey, G. A. Taylor, M. Cross, and P. Chow. Discretisation procedures for multi-physics phenomena. Journal of Computational and Applied Mathematics, 103(1):3–17, March 15, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042798002362.

Bultheel:1990:BRA

[Bul90a] A. Bultheel. Book review: Approximation, optimization and computing (theory and applications), A. G. Law and C. L. Wang (eds.): International Association for Mathematics and Computers in Simulation (IMACS). Journal of Computational and Applied Mathematics, 32(1–2): N7, November 26, 1990. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042790904456.

Bultheel:1990:BRH

[Bul90b]  A. Bultheel.  Book re-

view: Handbook of numerical analysis, volume I. (Finite difference methods — part 1; Solution of equations in R^{n} — part 1): P. G. Ciarlet and J. L. Lions: North-Holland, Elsevier Science Publishers, Amsterdam and Elsevier Science Publishing Co. Inc., New York, 1990. viii + 652 pages. Hard cover, Dfl.220.00/US $107.25. ISBN 0-444-70366-

7. Journal of Computational and Applied Mathematics, 32(1–2):N7–N8, November 26, 1990. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042790904467.

Bultheel:1991:BRO

[Bul91]  A. Bultheel.  Book re-

view: Operator Theory in Function Spaces: Kehe Zhu (Vol. 139 in Pure and Applied Mathematics, A series of monographs and textbooks) Marcel Dekker, Inc., New York — Basel, 1990, ix + 258 pages. Journal of Computational and Applied Mathematics, 34(2):

N2, April 4, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279190049P.

Bultheel:1994:BRN

[Bul94a]  A. Bultheel.  Book re-

view: Numerical linear algebra: L. Reichel, A. Ruttan, R. S. Varga (Eds.): W. de Gruyter and Co., Berlin, 1993, vii + 199 pages, Price: DM 168, ISBN 3-11-013784-4. Journal of Computational and Applied Mathematics, 50(1–3):N2– N3, May 20, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042794903379.

Bultheel:1994:BRO

[Bul94b] A. Bultheel. Book review: Oeuvres complètes; collected papers, volumes 1 & 2, T. J. Stieltjes: G. van Dijk (ed.), Springer Verlag, 1993, 566 + 766 pages, Price: DM 348 (HC), ISBN 3-540-55560-9. Journal of Computational and Applied Mathematics, 50(1–3):N2, May 20, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042794903360.

Bultheel:1994:BRP

[Bul94c]  A. Bultheel.  Book re-

view: Polynomial and matrix computations. volume 1: Fundamental algorithm: Dario Bini and Victor Pan Progress in Theoretical Computer Science, Birkhäuser, 1994, xvi + 415 pages. Journal of Computational and Applied Mathematics, 54(3):N3–N4, October 20, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042794902623.

Bultheel:1996:BRM

[Bul96a]  A. Bultheel.  Book re-

view: Matrix and operator valued functions: I. Gohberg and L. A. Sakhnovich (eds.) Operator theory advances and applications, 72, Birkhäuser Verlag, Basel, Boston, Berlin, 1994, xxviii + 211 pages, ISBN 3-7643-5091-1, Hardcover, DM 118, £48. Journal of Computational and Applied Mathematics, 66(1–2):N2, January 31, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042796804699.

Bultheel:1996:BRN

[Bul96b]  A. Bultheel.  Book re-

view: Nonselfadjoint operator and related topics: A. Feintuch and I. Gohberg (eds.) Operator Theory Advances and Applications, 73, Birkhäuser Verlag, Basel, Boston, Berlin, 1994, x + 416 pages, ISBN 3-7643-5097-0, Hardcover, DM 178, £78. Journal of Computational and Applied Mathematics, 66(1–2):N2– N3, January 31, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042796804705.

Bultheel:1996:BRT

[Bul96c] A. Bultheel. Book review: Time-varying discrete linear systems: Aristide Halanay and Vlad Ionescu Operator Theory Advances and Applications, 72, Birkhäuser Verlag, Basel, Boston, Berlin, 1994, xxvii + 211 pages, ISBN 3-7643-5091-1, Hardcover, DM 118, £48. Journal of Computational and Applied Mathematics, 66(1–2): N2, January 31, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042796804687.

Buong:1996:ASO

[Buo96] Nguyen Buong. On approximate solution for operator equations of Hammerstein type. Journal of Computational and Applied Mathematics, 75(1):77–86, November 12, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S037704279600060X.

Burrage:1993:EBP

[Bur93] Kevin Burrage. Efficient block predictor–corrector methods with a small number of corrections. Journal of Computational and Applied Mathematics, 45(1–2): 139–150, April 8, 1993. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279390270L.

Buschman:1999:LSC

[Bus99] R. G. Buschman. Less simple contiguous function relations for hypergeometric functions. Journal of Computational and Applied Mathematics, 107(1):127– 131, July 15, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042799000850.

Butcher:1992:ROP

[But92a]  J. C. Butcher.  The

role of orthogonal polynomials in numerical ordinary differential equations. Journal of Computational and Applied Mathematics, 43(1–2):231–242, November 25, 1992. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042792902683.

Butt:1992:ETM

[But92b]  R. Butt.  An existence

theory for a minimum energy problem. Journal of Computational and Applied Mathematics, 40(3): 253–257, July 24, 1992. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/037704279290181V.

Butt:1993:ODF

[But93] Rizwan Butt. Optimum design with finite elements: design of electrochemical machining. Journal of Computational and Applied Mathematics, 47(2):151– 165, ???? 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279390002S.

Bultheel:1990:MVP

[BV90] Adhemar Bultheel and Marc Van Barel. Minimal vector Padé approximation. Journal of Computational and Applied Mathematics, 32(1–2):27–37, November 26, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279090413T.

Berg:1995:NEM

[BV95] Christian Berg and Galliano Valent. Nevanlinna extremal measures for some orthogonal polynomials related to birth and death processes. Journal of Computational and Applied Mathematics, 57(1–2):29– 43, February 20, 1995. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042793E0232B.

Bultheel:1997:P

[BV97] A. Bultheel and M. Van Barel. Preface. Journal of Computational and Applied Mathematics, 77(1–2):1–3, January 13, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042796001197.

Batens:1999:NMF

[BV99] Nicholas Batens and Roger Van Keer. A numerical method for a free boundary value problem arising from chemical kinetics. Journal of Computational and Applied Mathematics, 111(1–2):187–199, November 15, 1999. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042799001429.

Bellen:1990:POS

[BVZ90] A. Bellen, R. Vermiglio, and M. Zennaro. Parallel ODEsolvers with stepsize control. Journal of Computational and Applied Mathematics, 31(2):277–293, August 31, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042790901705.

Brezinski:1991:STT

[BW91] Claude Brezinski and Guido Walz. Sequences of transformations and triangular recursion schemes, with applications in numerical analysis. Journal of Computational and Applied Mathematics, 34(3):361– 383, April 26, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042791900952.

Bottcher:1996:PCM

[BW96] Albrecht Böttcher and Hartmut Wolf. Polynomial collocation over massive sets for Toeplitz integral equations on the Bergman space. Journal of Computational and Applied Mathematics, 66(1–2):89–96, January 31, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042795001727.

Bastian-Walther:1999:RRI

[BWS99] Marion Bastian-Walther and Jochen W. Schmidt. Range restricted interpolation using Gregory’s rational cubic splines. Journal of Computational and Applied Mathematics, 103(2):221– 237, March 31, 1999. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S037704279800257X.

Brunner:1996:GSI

[BY96a] Hermann Brunner and Ningning Yan. On global superconvergence of iterated collocation solutions to linear second-kind Volterra integral equations. Journal of Computational and Applied Mathematics, 67(1):185– 189, February 20, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279600012X.

Brunner:1996:SCM

[BY96b] Hermann Brunner and Yuri Yatsenko. Spline collocation methods for nonlinear Volterra integral equations with unknown delay. Journal of Computational and Applied Mathematics, 71(1): 67–81, July 10, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0377042795002286.

Bi:1999:SCN

[BY99] Qinsheng Bi and Pei Yu. Symbolic computation of normal forms for semisimple cases. Journal of Computational and Applied Mathematics, 102(2):195– 220, February 26, 1999. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042798002222.

Baratchart:1993:RFP

[BZ93] Laurent Baratchart and Martin Zerner. On the recovery of functions from pointwise boundary values in a Hardy–Sobolev class of the disk. Journal of Computational and Applied Mathematics, 46(1–2):255– 269, June 14, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279390300Z.

Cuesta-Albertos:1994:SSM

[CADMMB94] J. A. Cuesta-Albertos, J. S.

Domínguez-Menchero, and C. Matrán-Bea. Some stochastics on monotone functions. Journal of Computational and Applied Mathematics, 55(2):165– 182, November 21, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042794900183.

Cahlon:1990:NSV

[Cah90] Baruch Cahlon. On the numerical stability of Volterra integral equations with delay argument. Journal of Computational and Applied Mathematics, 33(1):97–104, December 17, 1990. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042790902593.

Calogero:1997:RMT

[Cal97] F. Calogero. Remarkable matrices and trigonometric identities. Journal of Computational and Applied Mathematics, 83(1):127– 130, September 30, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042797001015.

Campbell:1991:AP

[Cam91] L. Lorne Campbell. Acknowledgement of priority. Journal of Computational and Applied Mathematics, 34(1):131, February 10, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279190153B.

Capobianco:1990:NPC

[Cap90] Maria Rosaria Capobianco. A new proof for the convergence of the trigonometric method for the Abel integral equations of the first kind. Journal of Computational and Applied Mathematics, 30(2):247– 250, May 28, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279090032U.

Carstensen:1990:GA

[Car90] C. Carstensen. On a general ρ-algorithm. Journal of Computational and Applied Mathematics, 33(1):61–71, December 17, 1990. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279090256Y.

Carroll:1993:CIS

[Car93] John Carroll. A composite integration scheme for the numerical solution of systems of parabolic PDEs in one space dimension. Journal of Computational and Applied Mathematics, 46(3): 327–343, June 22, 1993. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279390030F.

Carpenter:1996:SCS

[Car96] Amos J. Carpenter. Scientific computation on some mathematical problems. Journal of Computational and Applied Mathematics, 66(1–2):111–122, January 31, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042795001875.

Casas:1996:SLP

[Cas96]  Fernando Casas.  So-

lution of linear partial differential equations by Lie algebraic methods. Journal of Computational and Applied Mathematics, 76(1–2):159–170, December 17, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042796000994.

Chawla:1998:TSF

[CAZ98] M. M. Chawla and M. A. Al-Zanaidi. A two-stage fourth-order “almost” P - stable method for oscillatory problems. Journal of Computational and Applied Mathematics, 89(1):115– 118, March 9, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704279700232X.

See comment [VVVD99b].

Coleman:1992:AFC

[CB92] John P. Coleman and Andrew S. Booth. Analysis of a family of Chebyshev methods for y” = f(x, y). Journal of Computational and Applied Mathematics, 44(1):95–114, December 9, 1992. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042792900542.

Coleman:1995:CMP

[CB95] John P. Coleman and Andrew S. Booth. The Chebyshev methods of panovsky and Richardson as Runge– Kutta–Nyström methods. Journal of Computational and Applied Mathematics, 61(3):245–261, August 21, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042794000749.

Chen-Charpentier:1999:NSB

[CC99] Benito Chen-Charpentier.

Numerical simulation of biofilm growth in porous media. Journal of Computational and Applied Mathematics, 103(1):55–66, March 15, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042798002404.

Capobianco:1997:FAP

[CCJ97] Maria Rosaria Capobianco, Giuliana Criscuolo, and Peter Junghanns. A fast algorithm for Prandtl’s integro-differential equation. Journal of Computational and Applied Mathematics, 77(1–2):103–128, January 13, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042796001240.

Carte:1995:CSF

[CDF95] Gilles Carte, Jan Dusek, and Philippe Fraunié. Combined spectral-finite difference time discretization for periodic and quasi-periodic flows. Journal of Computational and Applied Mathematics, 63(1–3):245–254, November 20, 1995. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/037704279500078X.

Cuyt:1996:NPT

[CDL96a] A. Cuyt, K. Driver, and D. S. Lubinsky. Nuttall– Pommerenke theorems for homogeneous Padé approximants. Journal of Computational and Applied Mathematics, 67(1):141– 146, February 20, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0377042794001154.

Cuyt:1996:DAC

[CDL96b] A. Cuyt, K. A. Driver, and D. S. Lubinsky. A direct approach to convergence of multivariate, nonhomogeneous, Padé approximants. Journal of Computational and Applied Mathematics, 69(2):353– 366, May 23, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042795000445.

Cortelazzo:1991:ASM

[CDM91] G. Cortelazzo, C. De Simone, and G. A. Mian. An algorithm for simultaneous magnitude and phase approximation of bivariate rational polynomials. Journal of Computational and Applied Mathematics, 34(2): 145–160, April 4, 1991. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279190038L.

Cuyt:1999:FSR

[CDTV99] A. Cuyt, K. Driver, J. Tan, and B. Verdonk. A finite sum representation of the Appell series F_{1}(a, b, b′; c; x, y). Journal of Computational and Applied Mathematics, 105(1–2):213–219, May 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042799000357.

Chow:1990:CNF

[CDW90] Shui-Nee Chow, Byron Drachman, and Duo Wang. Computation of normal forms. Journal of Computational and Applied Mathematics, 29(2):129–143, February 15, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042790903532.

Celledoni:1998:DQB

[Cel98] Elena Celledoni. Discrete QMR and BCG in the numerical solution of linear systems of ODEs. Journal of Computational and Applied Mathematics, 91(1):159– 177, April 24, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042798000363.

Cruz:1991:ZHF

[CES91] Andrés Cruz, Javier Esparza, and Javier Sesma. Zeros of the Hankel function of real order out of the principal Riemann sheet. Journal of Computational and Applied Mathematics, 37(1–3):89–99, November 18, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279190109W.

Chin:1993:CGC

[CF93] P. Chin and P. A. Forsyth. A comparison of GMRES and CGSTAB accelerations for incompressible Navier– Stokes problems. Journal of Computational and Applied Mathematics, 46(3): 415–426, June 22, 1993. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/037704279390037C.

Cuminato:1999:SFS

[CFBT99] JoséA. Cuminato, Antonio Castelo Filho, Maurilio Boaventura, and Murilo F. Tomé. Simulation of free surface flows in a distributed memory environment. Journal of Computational and Applied Mathematics, 103(1):77–92, March 15, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042798002428.

Calvo:1996:ERK

[CFMR96] M. Calvo, J. M. Franco, J. I.

Montijano, and L. Rández. Explicit Runge–Kutta methods for initial value problems with oscillating solutions. Journal of Computational and Applied Mathematics, 76(1–2):195–212, December 17, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042796001033.

Crisci:1996:CRC

[CFR96] M. R. Crisci, N. Ferraro, and E. Russo. Convergence results for continuoustime waveform methods for Volterra integral equations. Journal of Computational and Applied Mathematics, 71(1):33–45, July 10, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042795002251.

Carl:1994:BVP

[CG94] S. Carl and C. Grossmann. Boundary value problems with discontinuities and monotone discretization. Journal of Computational and Applied Mathematics, 51(3):293– 303, June 30, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279200113N.

Cooper:1997:EPS

[CG97] S. Clement Cooper and Philip E. Gustafson. Extremal properties of strong quadrature weights and maximal mass results for truncated strong moment problems. Journal of Computational and Applied Mathematics, 80(2):197– 208, May 5, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042797000125.

Carnicer:1996:CRC

[CGEP96] J. M. Carnicer, M. García-Esnaola, and J. M. Peña. Convexity of rational curves and total positivity. Journal of Computational and Applied Mathematics, 71(2):

365–382, July 27, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042795002405.

Calvetti:1995:IPF

[CGR95] D. Calvetti, E. Gallopoulos, and L. Reichel. Incomplete partial fractions for parallel evaluation of rational matrix functions. Journal of Computational and Applied Mathematics, 59(3): 349–380, May 30, 1995. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042794000372.

Calvetti:1999:CEB

[CGR99] D. Calvetti, G. H. Golub, and L. Reichel. A computable error bound for matrix functionals. Journal of Computational and Applied Mathematics, 103(2):301– 306, March 31, 1999. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042798002672.

Camacho:1997:RFA

[CGV97] M. Camacho and P. González-Vera. Rational functions associated with double infinite sequences of complex numbers. Journal of Computational and Applied Mathematics, 85(1):37–51, November 6, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042797001258.

Christiansen:1994:ECN

[CH94a] Søren Christiansen and Per Christian Hansen. The effective condition number applied to error analysis of certain boundary collocation methods. Journal of Computational and Applied Mathematics, 54(1):15–36, September 20, 1994. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0377042794903913.

Cools:1994:IFC

[CH94b] Ronald Cools and Ann Haegemans. An imbedded family of cubature formulae for n-dimensional product regions. Journal of Computational and Applied Mathematics, 51(2): 251–260, June 20, 1994. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/037704279200007V.

Chakrabarti:1996:DGM

[CH96a] A. Chakrabarti and Hamsapriye. Derivation of a general mixed interpolation formula. Journal of Computational and Applied Mathematics, 70(1): 161–172, June 14, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042796001458.

Chakrabarti:1996:MQR

[CH96b] A. Chakrabarti and Hamsapriye. Modified quadrature rules based on a generalised mixed interpolation formula. Journal of Computational and Applied Mathematics, 76(1–2):239–254, December 17, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042796001070.

Chuang:1996:NCM

[CH96c] J. M. Chuang and S. Z.

Hu. Numerical computation of Muskhelishvili’s integral equation in plane elasticity. Journal of Computational and Applied Mathematics, 66(1–2):123–138, January 31, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042795001905.

Chakrabarti:1997:MGR

[CH97] A. Chakrabarti and Hamsapriye. On modified Gregory rules based on a generalised mixed interpolation formula. Journal of Computational and Applied Mathematics, 78(1):103– 124, February 3, 1997. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042796001355.

Chakrabarti:1998:DEI

[CH98] A. Chakrabarti and Hamsapriye. Derivation of the errors involved in interpolation and their application to numerical quadrature formulae. Journal of Computational and Applied Mathematics, 92(1):59–68, May 20, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042798000454.

Chao:1995:UCM

[Cha95]  Y. A. Chao.  Unify-

ing via conformal mapping coarse-mesh neutron diffusion calculations in Cartesian and hexagonal geometries. Journal of Computational and Applied Mathematics, 63(1–3):449–456, November 20, 1995. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042795000909.

Chang:1996:CAP

[Cha96] Da-Wei Chang. Convergence analysis of the parallel multisplitting TOR method. Journal of Computational and Applied Mathematics, 72(1):169–177, July 30, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042795002707.

Chandna:1997:FPE

[Cha97] A. Chandna. Flow past an elliptic cylinder. Journal of Computational and Applied Mathematics, 85 (2):203–214, November 12, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042797000356.

Chanane:1998:EFO

[Cha98a]  B. Chanane.  Eigenval-

ues of fourth order Sturm– Liouville problems using Fliess series. Journal of Computational and Applied Mathematics, 96(2):91–97, September 15, 1998. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042798000867.

Chartier:1998:DIR

[Cha98b]  P. Chartier.  On di-

agonally iterated Runge– Kutta methods for dissipative ODEs. Journal of Computational and Applied Mathematics, 89(1):73–85, March 9, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704279700229X.

Chapko:1999:NSD

[Cha99] Roman Chapko. On the numerical solution of direct and inverse problems for the heat equation in a semiinfinite region. Journal of Computational and Applied Mathematics, 108(1–2):41– 55, August 15, 1999. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042799000990.

Chen:1994:BMM

[Che94]  Zhangxin Chen.  BDM

mixed methods for a nonlinear elliptic problem. Journal of Computational and Applied Mathematics, 53(2):207–223, August 19, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042794900469.

Chen:1997:VFE

[Che97a]  Jiuhua Chen.  The p-

version of the finite element method for a singularly perturbed boundary value problem. Journal of Computational and Applied Mathematics, 84(1):119–135, October 21, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042797001192.

Chen:1997:SCS

[Che97b] Xiaojun Chen. Superlinear convergence of smoothing quasi-Newton methods for nonsmooth equations. Journal of Computational and Applied Mathematics, 80(1): 105–126, April 14, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042797801331.

Cheikh:1998:DLP

[Che98a] Youssèf Ben Cheikh. Decomposition of Laguerre polynomials with respect to the cyclic group of order n. Journal of Computational and Applied Mathematics, 99(1–2):55–66, November 16, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042798001459.

Chen:1998:PUM

[Che98b] Xiaojun Chen. On preconditioned Uzawa methods and SOR methods for saddle-point problems. Journal of Computational and Applied Mathematics, 100(2):207–224, December 21, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042798001976.

Choi:1994:GNO

[CHM94] Y. S. Choi, Hyeyoung Oh Her, and P. J. McKenna. Galloping: nonlinear oscillation in a periodically forced loaded hanging cable. Journal of Computational and Applied Mathematics, 52(1– 3):23–34, July 20, 1994. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0377042794903468.

Choudhury:1992:WRT

[Cho92] S. Roy Choudhury. Waveform relaxation techniques for linear and nonlinear diffusion equations. Journal of Computational and Applied Mathematics, 42(2):253– 267, October 12, 1992. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/037704279290079D.

Christiansen:1990:SAE

[Chr90] Søren Christiansen. A stability analysis of a Eulerian solution method for moving boundary problems in electrochemical machining. Journal of Computational and Applied Mathematics, 33(3):269–296, December 31, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042705800044.

Chronopoulos:1992:NCL

[Chr92] A. T. Chronopoulos. Nonlinear CG-like iterative methods. Journal of Computational and Applied Mathematics, 40(1):73–89, June 12, 1992. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279290043W.

Chronopoulos:1994:SUL

[Chr94] A. T. Chronopoulos. On the squared unsymmetric Lanczos method. Journal of Computational and Applied Mathematics, 54(1):65–78, September 20, 1994. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042794903956.

Christiansen:1998:DAI

[Chr98] Søren Christiansen. Derivation and analytical investigation of three direct boundary integral equations for the fundamental biharmonic problem. Journal of Computational and Applied Mathematics, 91(2): 231–247, May 4, 1998. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042798000417.

Chuanqing:1997:BTT

[Chu97a] Gu Chuanqing. Bivariate Thiele-type matrix-valued rational interpolants. Journal of Computational and Applied Mathematics, 80(1): 71–82, April 14, 1997. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S037704279700006X.

Chuanqing:1997:MGI

[Chu97b]  Gu Chuanqing.  Multi-

variate generalized inverse vector-valued rational interpolants. Journal of Computational and Applied Mathematics, 84(2):137–146, October 28, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042797865144.

Chen:1997:AEZ

[CI97] Yang Chen and Mourad E. H. Ismail. Asymptotics of extreme zeros of the Meixner–Pollaczek polynomials. Journal of Computational and Applied Mathematics, 82(1–2):59– 78, September 15, 1997. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042797000939.

Chen:1994:ABB

[CIM94] Yang Chen, Mourad E. H.

Ismail, and K. A. Muttalib. Asymptotics of basic Bessel functions and q-Laguerre polynomials. Journal of Computational and Applied Mathematics, 54(3):263– 272, October 20, 1994. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279200128V.

Cooper:1990:OLP

[CJT90] S. Clement Cooper, William B.

Jones, and W. J. Thron. Orthogonal Laurent-polynomials and continued fractions associated with log-normal distributions. Journal of Computational and Applied Mathematics, 32(1–2):39– 46, November 26, 1990. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/037704279090414U.

Christiansen:1991:CCS

[CK91] E. Christiansen and K. O. Kortanek. Computation of the collapse state in limit analysis using the LP primal affine scaling algorithm. Journal of Computational and Applied Mathematics, 34(1):47–63, February 10, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279190147C.

Choudury:1992:CSF

[CK92] Gilbert Choudury and Philip Korman. On computation of solutions of fully nonlinear elliptic problems. Journal of Computational and Applied Mathematics, 41(3):301– 311, August 28, 1992. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/037704279290137M.

Cvijovic:1999:IIC

[CK99] Djurdje Cvijovic and Jacek Klinowski. Integrals involving complete elliptic integrals. Journal of Computational and Applied Mathematics, 106(1):169– 175, June 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042799000667.

Colton:1997:ISO

[CKM97] David Colton, Rainer Kress, and Peter Monk. Inverse scattering from an orthotropic medium. Journal of Computational and Applied Mathematics, 81(2): 269–298, July 8, 1997. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042797000654.

Coomes:1994:SOO

[CKP94] Brian A. Coomes, Hüseyin Koçak, and Kenneth J. Palmer. Shadowing orbits of ordinary differential equations. Journal of Computational and Applied Mathematics, 52(1–3):35– 43, July 20, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042794903476.

Chettab:1991:CSM

[CL91] Mahieddine Chettab and Mbaro-Saman Lubuma. Coefficients of singularities and mixed methods for the mixed Dirichlet–Neumann problem for the Stokes operator on a polygon. Journal of Computational and Applied Mathematics, 35(1–3): 139–157, June 26, 1991. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/037704279190203V.

Caudill:1993:CPC

[CL93] Lester F. Caudill, Jr. and Bruce D. Lowe. On the construction of a potential from Cauchy data. Journal of Computational and Applied Mathematics, 47(3):323–333, September 30, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279390060O.

Coulibaly:1997:PSK

[CL97] Ibrahim Coulibaly and Christian Lécot. Particle simulations of the Kac model of the Boltzmann equation. Journal of Computational and Applied Mathematics, 87 (1):169–193, December 18, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042797001842.

Cui:1998:NSC

[CL98] Bao Tong Cui and Wei Nian Li. A necessary and sufficient condition for oscillation of parabolic equations with several delays. Journal of Computational and Applied Mathematics, 95(1–2):153–156, August 28, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042798000697.

Cabada:1999:ESD

[CL99] Alberto Cabada and Susana Lois. Existence of solution for discontinuous third order boundary value problems. Journal of Computational and Applied Mathematics, 110(1):105–114, October 15, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042799001995.

Cabay:1992:TCN

[CLB92] Stan Cabay, George Labahn, and Bernhard Beckermann. On the theory and computation of nonperfect Padé– Hermite approximants. Journal of Computational and Applied Mathematics, 39(3): 295–313, May 8, 1992. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279290206D.

Cheung:1998:UPS

[CLL98] Leung-Fu Cheung, Chun-Kong Law, and Man-Chun Leung. Uniqueness of positive solutions to the rotationally symmetric p-harmonic map equations. Journal of Computational and Applied Mathematics, 88(1):45–56, February 23, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042797002045.

Clavero:1995:UCA

[CLM95] Carmelo Clavero, Francisco Lisbona, and John J. H. Miller. Uniform convergence of arbitrary order on nonuniform meshes for a singularly perturbed boundary value problem. Journal of Computational and Applied Mathematics, 59(2): 155–171, May 19, 1995. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/037704279400023T.

Calvetti:1997:IMA

[CLR97] D. Calvetti, N. Levenberg, and L. Reichel. Iterative methods for X-AXB = C. Journal of Computational and Applied Mathematics, 86(1):73–101, November 28, 1997. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042797001490.

Chien:1995:CGM

[CLW95] C.-S. Chien, N.-H. Lu, and Z.-L. Weng. Conjugate gradient methods for continuation problems II. Journal of Computational and Applied Mathematics, 62(2):197–216, September 20, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042794001044.

Colton:1992:CTM

[CM92] David Colton and Peter Monk. A comparison of two methods for solving the inverse scattering problem for acoustic waves in an inhomogeneous medium. Journal of Computational and Applied Mathematics, 42(1):5– 16, September 8, 1992. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279290159U.

Cabrelli:1995:KMP

[CM95] Carlos A. Cabrelli and Ursula M. Molter. The Kantorovich metric for probability measures on the circle. Journal of Computational and Applied Mathematics, 57(3):345–361, February 28, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042793E02136.

Common:1996:SSM

[CM96a] A. K. Common and J. H. McCabe. The symmetric strong moment problem. Journal of Computational and Applied Mathematics, 67(2):327–341, March 29, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042794001251.

Costantini:1996:CPI

[CM96b] P. Costantini and C. Manni. Comonotone parametric C 1 interpolation of nongridded data. Journal of Computational and Applied Mathematics, 75(1):147– 169, November 12, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042796000842.

Costantini:1996:CPT

[CM96c] Paolo Costantini and Carla Manni. On a class of polynomial triangular macroelements. Journal of Computational and Applied Mathematics, 73(1–2):45– 64, October 5, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042796000337.

Costian:1996:DMA

[CM96d] Calin R. Costian and Dan C.

Marinescu. A distributed memory algorithm for 3-D FFTs. Journal of Computational and Applied Mathematics, 66(1–2):139–151, January 31, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042795001638.

Campbell:1997:DAT

[CM97] Stephen L. Campbell and Wiesław Marszalek. DAEs arising from traveling wave solutions of PDEs. Journal of Computational and Applied Mathematics, 82(1–2): 41–58, September 15, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042797000848.

Calvo:1999:ESS

[CMGP99] M. Calvo, J. I. Montijano, and S. Gonzalez-Pinto. On the existence of solution of stage equations in implicit Runge–Kutta methods. Journal of Computational and Applied Mathematics, 111(1–2):25–36, November 15, 1999. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042799001296.

Calvo:1990:FOI

[CMR90] M. Calvo, J. I. Montijano, and L. Randez. A fifth-order interpolant for the Dormand and Prince Runge–Kutta method. Journal of Computational and Applied Mathematics, 29(1):91–100, January 10, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042790901989.

Calvo:1993:SVS

[CMR93] M. Calvo, J. I. Montijano, and L. Rández. A_{0}- stability of variable stepsize BDF methods. Journal of Computational and Applied Mathematics, 45(1–2): 29–39, April 8, 1993. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/037704279390262A.

Calio:1995:NSG

[CMR95] F. Caliò, E. Marchetti, and P. Rabinowitz. On the numerical solution of the generalized Prandtl equation using variationdiminishing splines. Journal of Computational and Applied Mathematics, 60(3): 297–307, June 30, 1995. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279400024U.

Conti:1999:RSP

[CMR99] C. Conti, R. Morandi, and C. Rabut. P _{1}-reproducing shape-preserving quasi-interpolant using positive quadratic splines with local support. Journal of Computational and Applied Mathematics, 104(2):111–122, May 17, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042798002593.

Cabada:1999:ASN

[CNP99] Alberto Cabada, Juan J. Nieto, and Rodrigo L. Pouso. Approximate solutions to a new class of nonlinear diffusion problems. Journal of Computational and Applied Mathematics, 108(1–2):219– 231, August 15, 1999. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042799001120.

Climent:1997:SIM

[CNS97] Joan-Josep Climent, Michael Neumann, and Avram Sidi. A semi-iterative method for real spectrum singular linear systems with an arbitrary index. Journal of Computational and Applied Mathematics, 87(1):21–38, December 18, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042797001696.

Coleman:1998:MIM

[Col98] John P. Coleman. Mixed interpolation methods with arbitrary nodes. Journal of Computational and Applied Mathematics, 92(1):69–83, May 20, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042798000478.

Cong:1999:CVS

[Con99] Nguyen Huu Cong. Continuous variable stepsize explicit pseudo two-step RK methods. Journal of Computational and Applied Mathematics, 101(1–2):105–116, January 15, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704279800199X.

Cools:1999:MCR

[Coo99]  Ronald Cools.  Mono-

mial cubature rules since “Stroud”: a compilation

— part 2. Journal of Computational and Applied Mathematics, 112(1–2):21– 27, November 30, 1999. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042799002290.

Cope:1994:CFO

[Cop94] Davis K. Cope. Characterization of first-order convergent sequences. Journal of Computational and Applied Mathematics, 54(1):51–63, September 20, 1994. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042794903948.

Couchman:1999:SFL

[Cou99] H. M. P. Couchman. Simulating the formation of large-scale cosmic structure with particle-grid methods. Journal of Computational and Applied Mathematics, 109(1–2):373–406, September 30, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704279900165X.

Carstensen:1993:IMD

[CP93] Carsten Carstensen and Miodrag S. Petković. On iteration methods without derivatives for the simultaneous determination of polynomial zeros. Journal of Computational and Applied Mathematics, 45(3): 251–266, May 3, 1993. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279390044C.

Castillo:1996:SAD

[CP96] José Castillo and Erik M. Pedersen. Solution adaptive direct variational grids for fluid flow calculations. Journal of Computational and Applied Mathematics, 67(2): 343–370, March 29, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042795000178.

Chen:1999:ARS

[CP99] Horng-Yang Chen and Helmut Pottmann. Approximation by ruled surfaces. Journal of Computational and Applied Mathematics, 102(1):143–156, February 15, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704279800212X.

Cyphers:1996:HPS

[CPK96] Cliff Cyphers, Marcin Paprzycki, and Andreas Karageorghis. High performance solution of partial differential equations discretized using a Chebyshev spectral collocation method. Journal of Computational and Applied Mathematics, 69(1): 71–80, April 30, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042795000291.

Chaudhry:1997:EEB

[CQRZ97] M. Aslam Chaudhry, Asghar Qadir, M. Rafique, and S. M. Zubair. Extension of Euler’s beta function. Journal of Computational and Applied Mathematics, 78(1):19–32, February 3, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042796001021.

Chen:1995:NMQ

[CQW95] Xiaojun Chen, Liqun Qi, and Robert S. Womersley. Newton’s method for quadratic stochastic programs with recourse. Journal of Computational and Applied Mathematics, 60(1– 2):29–46, June 20, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
037704279400082C.

Cools:1993:MCR

[CR93] Ronald Cools and Philip Rabinowitz. Monomial cubature rules since “Stroud”: a compilation. Journal of Computational and Applied Mathematics, 48(3):309– 326, November 30, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042793900279.

Calvetti:1996:ARI

[CR96] D. Calvetti and L. Reichel. Adaptive Richardson iteration based on Leja points. Journal of Computational and Applied Mathematics, 71(2):267– 286, July 27, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042796871627.

Calvetti:1997:FCM

[CR97] D. Calvetti and L. Reichel. Factorizations of Cauchy matrices. Journal of Computational and Applied Mathematics, 86 (1):103–123, November 28, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042797001507.

Calvetti:1998:HIM

[CR98] D. Calvetti and L. Reichel. A hybrid iterative method for symmetric indefinite linear systems. Journal of Computational and Applied Mathematics, 92(2): 109–133, June 26, 1998. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S037704279800048X.

Carino:1994:ACC

[CRD94] Ricolindo Cariño, Ian Robinson, and Elise De Doncker. Adaptive cubature over a collection of triangles using the dtransformation. Journal of Computational and Applied Mathematics, 50(1–3):171– 183, May 20, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042794902984.

Criscuolo:1994:LIC

[Cri94] Giuliana Criscuolo. Lagrange interpolation with constraints. Journal of Computational and Applied Mathematics, 50(1–3):185– 196, May 20, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042794902992.

Criscuolo:1997:NAC

[Cri97] Giuliana Criscuolo. A new algorithm for Cauchy principal value and Hadamard finite-part integrals. Journal of Computational and Applied Mathematics, 78(2): 255–275, February 17, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042796001422.

Crisci:1990:SAH

[CRV90] M. R. Crisci, E. Russo, and A. Vecchio. Stability analysis of the de hoog and Weiss implicit Runge–Kutta methods for the Volterra integral and integrodifferential equations. Journal of Computational and Applied Mathematics, 29(3):329– 341, March 15, 1990. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/037704279090015R.

Crisci:1997:DTW

[CRV97] M. R. Crisci, E. Russo, and A. Vecchio. Discretetime waveform relaxation Volterra–Runge–Kutta methods: Convergence analysis. Journal of Computational and Applied Mathematics, 86(2):359–374, December 10, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042797001684.

Crisci:1993:SPV

[CRvdHV93] M. R. Crisci, E. Russo, P. J. van der Houwen, and A. Vecchio. Stability of parallel Volterra–Runge– Kutta methods. Journal of Computational and Applied Mathematics, 45(1–2):169– 180, April 8, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279390272D.

Cecchi:1996:EMH

[CRZR96] M. Morandi Cecchi, M. Redivo-Zaglia, and G. Russo. Extrapolation methods for hyperbolic systems with relaxation. Journal of Computational and Applied Mathematics, 66(1–2):359–375, January 31, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042795001948.

Cash:1993:NAS

[CS93a] J. R. Cash and S. Semnani. A new approach to solving nonstiff initialvalue problems. Journal of Computational and Applied Mathematics, 45(1–2): 41–46, April 8, 1993. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279390263B.

Cash:1993:NSC

[CS93b] J. R. Cash and H. H. M. Silva. On the numerical solution of a class of singular two-point boundary value problems. Journal of Computational and Applied Mathematics, 45(1–2): 91–102, April 8, 1993. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279390267F.

Carstensen:1995:SFP

[CS95a] Carsten Carstensen and Tetsuya Sakurai. Simultaneous factorization of a polynomial by rational approximation. Journal of Computational and Applied Mathematics, 61(2):165– 178, July 31, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042794000637.

Connett:1995:SWS

[CS95b] William C. Connett and Alan L. Schwartz. Subsets of R which support hypergroups with polynomial characters. Journal of Computational and Applied Mathematics, 65(1–3):73– 84, December 29, 1995. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042795001018.

Cachafeiro:1996:KUC

[CS96a] A. Cachafeiro and C. Suárez.

Kernels on the unit circle. orthogonality. Journal of Computational and Applied Mathematics, 72(2):407– 420, August 13, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042796000106.

Christiansen:1996:CSN

[CS96b] Søren Christiansen and Jukka Saranen. The conditioning of some numerical methods for first kind boundary integral equations. Journal of Computational and Applied Mathematics, 67(1):43–58, February 20, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042794001219.

Cachafeiro:1997:ASP

[CS97a] A. Cachafeiro and C. Suárez. About semiclassical polynomials on the unit circle corresponding to the class (2,2). Journal of Computational and Applied Mathematics, 85(1):123– 144, November 6, 1997. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042797001349.

Cahlon:1997:SCC

[CS97b] Baruch Cahlon and Darrell Schmidt. Stability criteria for certain delay integral equations of Volterra type. Journal of Computational and Applied Mathematics, 84(2):161–188, October 28, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042797001155.

Chengjian:1997:NSC

[CS97c] Zhang Chengjian and Zhou Shuzi. Nonlinear stability and D-convergence of Runge–Kutta methods for delay differential equations. Journal of Computational and Applied Mathematics, 85(2):225–237, November 12, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042797001180.

Chow:1997:ESI

[CS97d] Edmond Chow and Yousef Saad. Experimental study of ILU preconditioners for indefinite matrices. Journal of Computational and Applied Mathematics, 86(2):387–414, December 10, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042797001714.

Chen:1998:DND

[CS98] Benito Chen and Francisco Solis. Discretizations of nonlinear differential equations using explicit finite order methods. Journal of Computational and Applied Mathematics, 90(2):171– 183, April 17, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704279800017X.

Christian:1999:PEN

[CS99] Felix Christian and Guimarães Santos. A posteriori estimators for nonlinear elliptic partial differential equations. Journal of Computational and Applied Mathematics, 103(1):99– 114, March 15, 1999. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042798002441.

Campigotto:1995:AKM

[CSE95] C. Campigotto, Yu. F. Smirnov, and S. G. Enikeev. q-analogue of the Krawtchouk and Meixner orthogonal polynomials. Journal of Computational and Applied Mathematics, 57(1–2):87–97, February 20, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042793E0236F.

Cahlon:1993:CCO

[CSS93] B. Cahlon, I. E. Schochetman, and M. Shillor. Convective cooling and optimal placement of electronic components with variable ambient temperature. I. The linear model. Journal of Computational and Applied Mathematics, 47(3):351–367, September 30, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279390062G.

Coolen-Schrijner:1998:ARW

[CSvD98] Pauline Coolen-Schrijner and Erik A. van Doorn. Analysis of random walks using orthogonal polynomials. Journal of Computational and Applied Mathematics, 99(1–2):387–399, November 16, 1998. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042798001721.

Capizzano:1999:ESV

[CT99] Stefano Serra Capizzano and Paolo Tilli. Extreme singular values and eigenvalues of non-Hermitian block Toeplitz matrices. Journal of Computational and Applied Mathematics, 108(1–2):113–130, August 15, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042799001041.

Chaudhry:1996:ACF

[CTV96] M. A. Chaudhry, N. M.

Temme, and E. J. M. Veling. Asymptotics and closed form of a generalized incomplete gamma function. Journal of Computational and Applied Mathematics, 67(2):371– 379, March 29, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042795000186.

Cushing:1994:OAS

[Cus94] J. M. Cushing. Oscillations in age-structured population models with an Allee effect. Journal of Computational and Applied Mathematics, 52(1–3):71– 80, July 20, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042794903492.

Cuyt:1990:MCT

[Cuy90]  A. Cuyt.  A multivari-

ate convergence theorem of the “de Montessus de Ballore” type. Journal of Computational and Applied Mathematics, 32(1–2):47– 57, November 26, 1990. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/037704279090415V.

See extension [Cuy92].

Cuyt:1992:EMC

[Cuy92] A. Cuyt. Extension of “A multivariate convergence theorem of the ‘de Montessus de Ballore’ type” to multipoles. Journal of Computational and Applied Mathematics, 41(3):323–330, August 28, 1992. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/

science/article/pii/037704279290139O.

See [Cuy90].

Cuyt:1999:HWC

[Cuy99] Annie Cuyt. How well can the concept of Padé approximant be generalized to the multivariate case? Journal of Computational and Applied Mathematics, 105(1– 2):25–50, May 1999. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S037704279900028X.

Cooper:1993:SSI

[CV93] G. J. Cooper and R. Vignesvaran. Some schemes for the implementation of implicit Runge–Kutta methods. Journal of Computational and Applied Mathematics, 45(1–2):213–225, April 8, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279390276H.

Cvetkovic:1999:NCM

[Cve99] Ljiljana Cvetković. A note on the convergence of the MAOR method. Journal of Computational and Applied Mathematics, 103(2):281– 285, March 31, 1999. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042798002635.

Chiou:1999:GHO

[CW99] J. C. Chiou and S. D. Wu. On the generation of higher order numerical integration methods using lower order Adams–Bashforth and Adams–Moulton methods. Journal of Computational and Applied Mathematics, 108(1–2):19–29, August 15, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042799000965.

Chen:1991:PMA

[CY91] Xiaojun Chen and Tetsuro Yamamoto. PCG methods applied to a system of nonlinear equations. Journal of Computational and Applied Mathematics, 38(1–3):61– 75, December 23, 1991. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279190161C.

Chen:1994:NLM

[CY94] Xiaojun Chen and Tetsuro Yamamoto. Newtonlike methods for solving underdetermined nonlinear equations with nondifferentiable terms. Journal of Computational and Applied Mathematics, 55(3):311– 324, November 30, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042794900361.

Cecchi:1990:MPT

[CZ90] M. Morandi Cecchi and M. Redivo Zaglia. Mathematical programming techniques to solve biharmonic problems by a recursive projection algorithm. Journal of Computational and Applied Mathematics, 32(1–2):191–201, November 26, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042790904308.

Cecchi:1993:CCR

[CZ93] M. Morandi Cecchi and M. Redivo Zaglia. Computing the coefficients of a recurrence formula for numerical integration by moments and modified moments. Journal of Computational and Applied Mathematics, 49(1–3):207–216, December 31, 1993. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042793901522.

Chaudhry:1994:GIG

[CZ94] M. Aslam Chaudhry and S. M. Zubair. Generalized incomplete gamma functions with applications. Journal of Computational and Applied Mathematics, 55(1):99–123, October 31, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042794901872.

Chaudhry:1995:DGI

[CZ95] M. Aslam Chaudhry and S. M. Zubair. On the decomposition of generalized incomplete gamma functions with applications to Fourier transforms. Journal of Computational and Applied Mathematics, 59(3): 253–284, May 30, 1995. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279400026W.

Carstensen:1996:ACF

[CZS96] Carsten Carstensen, Darius Zarrabi, and Ernst P. Stephan. On the h-adaptive coupling of FE and BE for viscoplastic and elastoplastic interface problems. Journal of Computational and Applied Mathematics, 75(2):345–363, November 28, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042796000908.

Daalhuis:1996:H

[Daa96] A. B. Olde Daalhuis. Hyperterminants I. Journal of Computational and Applied Mathematics, 76(1–2):255– 264, December 17, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042796001082.

Daalhuis:1998:HI

[Daa98] A. B. Olde Daalhuis. Hyperterminants II. Journal of Computational and Applied Mathematics, 89(1):87–95, March 9, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042797002203.

Damelin:1998:CSF

[Dam98] S. B. Damelin. A characterization of smoothness for Freud weights. Journal of Computational and Applied Mathematics, 99(1–2):463–473, November 16, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042798001770.

Damelin:1999:STG

[Dam99] S. B. Damelin. Smoothness theorems for generalized symmetric Pollaczek weights on (-1,1). Journal of Computational and Applied Mathematics, 101(1–2):87– 103, January 15, 1999. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042798001964.

Daripa:1998:SUF

[Dar98] Prabir Daripa. Some useful filtering techniques for illposed problems. Journal of Computational and Applied Mathematics, 100(2):161– 171, December 21, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042798001861.

deBruin:1995:P

[dB95] Marcel G. de Bruin. Preface. Journal of Computational and Applied Mathematics, 65(1–3):1, December 29, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042795900551.

DeSamblanx:1999:UIF

[DB99] Gorik De Samblanx and Adhemar Bultheel. Using implicitly filtered RKS for generalised eigenvalue problems. Journal of Computational and Applied Mathematics, 107(2):195–218, July 31, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042799000898.

DeMeyer:1996:PFO

[DBFV96] H. De Meyer, P. Bocher, V. Fack, and G. Vanden Berghe. A parallel fifth-order algorithm for the numerical solution of Volterra equations. Journal of Computational and Applied Mathematics, 71(1): 115–124, July 10, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042795002294.

deBruin:1999:STC

[dBS99] M. G. de Bruin and A. Sharma. On a Schoenbergtype conjecture. Journal of Computational and Applied Mathematics, 105(1–2):221– 228, May 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042799000138.

DeMarchi:1995:PAF

[DC95] S. De Marchi and M. Morandi Cecchi. The polynomial approximation in the finite element method. Journal of Computational and Applied Mathematics, 57(1–2):99– 114, February 20, 1995. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0377042793E0237G.

Deakin:1995:NDB

[DD95] A. S. Deakin and J. R. Dryden. Numerically derived boundary conditions on artificial boundaries. Journal of Computational and Applied Mathematics, 58(1):1– 16, March 20, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042793E0261J.

Dimov:1998:NIM

[DDG98] I. T. Dimov, T. T. Dimov, and T. V. Gurov. A new iterative Monte Carlo approach for inverse matrix problem. Journal of Computational and Applied Mathematics, 92(1):15–35, May 20, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042798000430.

deDoncker:1999:LSP

[dDGZ99] Elise de Doncker, Ajay Gupta, and Rodger R. Zanny. Large-scale parallel numerical integration. Journal of Computational and Applied Mathematics, 112(1–2):29–44, November 30, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042799002101.

DeBruin:1993:CSH

[DDL93] M. G. De Bruin, K. A.

Driver, and D. S. Lubinsky. Convergence of simultaneous Hermite–Padé approximants to the n-tuple of q-hypergeometric series {_{2}Φ_{0}((A, α_{j}), (1, 1); z)}_{j}^{n} =

1. Journal of Computational and Applied Mathematics, 49(1–3):37–43, December 31, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279390132U.

DeMichele:1993:FKN

[DDR93] L. De Michele, M. Di Natale, and D. Roux. Féjer kernels and noisy Fourier series. Journal of Computational and Applied Mathematics, 49(1–3):45–50, December 31, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279390133V.

DeBruin:1990:POC

[De 90a] Marcel G. De Bruin. Polynomials orthogonal on a circular arc. Journal of Computational and Applied Mathematics, 31(2):253–266, August 31, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279090168Y.

DeZeeuw:1990:MDP

[De 90b] P. M. De Zeeuw. Matrixdependent prolongations and restrictions in a blackbox multigrid solver. Journal of Computational and Applied Mathematics, 33(1): 1–27, December 17, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        037704279090252U.

DeDoelder:1991:SSC

[De 91a] P. J. De Doelder. On some series containing ψ(x) − ψ(y >) and (ψ(x) − ψ(y > ))^{2} for certain values of x and y. Journal of Computational and Applied Mathematics, 37(1–3):125–141, November 18, 1991. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279190112W.

DeOliveira:1991:CFD

[De 91b] Paula De Oliveira. On the characterization of finite differences “optimal” meshes. Journal of Computational and Applied Mathematics, 36(2):137–148, August 27, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279190022C.

DeBruin:1993:TLZ

[De 93] Marcel G. De Bruin. A tool for locating zeros of orthogonal polynomials in Sobolev inner product spaces. Journal of Computational and Applied Mathematics, 49(1–3):27–35, December 31, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279390131T.

DeBortoli:1999:TDS

[De 99] A. L. De Bortoli. Three dimensional supersonic flow analysis over a satellite vehicle launcher. Journal of Computational and Applied Mathematics, 103(1):33–41, March 15, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042798002386.

Dubuc:1990:AFS

[DE90] Serge Dubuc and Abdelkader Elqortobi. Approximations of fractal sets. Journal of Computational and Applied Mathematics, 29(1):79–89, January 10, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042790901978.

DeLillo:1993:FLC

[DE93] Thomas K. DeLillo and Alan R. Elcrat. A Fornberglike conformal mapping method for slender regions. Journal of Computational and Applied Mathematics, 46(1–2):49–64, June 14, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        037704279390286K.

Draux:1997:HSC

[DE97] André Draux and Charaf El Hami. Hermite–Sobolev and closely connected orthogonal polynomials. Journal of Computational and Applied Mathematics, 81(1): 165–179, June 1, 1997. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042797000381.

Draux:1999:PSB

[DE99] André Draux and Charaf Elhami. On the positivity of some bilinear functionals in Sobolev spaces. Journal of Computational and Applied Mathematics, 106(2): 203–243, June 30, 1999. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042799000631.

Dehesa:1993:P

[Deh93] Jesús S. Dehesa. Preface. Journal of Computational and Applied Mathematics, 48(1–2):1, October 29, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279390311X.

Dehn:1994:SAO

[Deh94] Thomas Dehn. A shortcut to asymptotics for orthogonal polynomials. Journal of Computational and Applied Mathematics, 50(1–3):207– 219, May 20, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042794903018.

Dehghan:1999:FIF

[Deh99] Mehdi Dehghan. Fully implicit finite differences methods for two-dimensional diffusion with a non-local boundary condition. Journal of Computational and Applied Mathematics, 106 (2):255–269, June 30, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042799000655.

DeMatteis:1992:CCD

[DEHG92] A. De Matteis, J. Eichenauer-Herrmann, and H. Grothe. Computation of critical distances within multiplicative congruential pseudorandom number sequences. Journal of Computational and Applied Mathematics, 39(1):49–55, February 28, 1992. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279290221I.

Deif:1995:RPB

[Dei95] A. S. Deif. Rigorous perturbation bounds for eigenvalues and eigenvectors of a matrix. Journal of Computational and Applied Mathematics, 57(3):403– 412, February 28, 1995. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042793E02084.

Dellwo:1995:ADK

[Del95]  David R. Dellwo.  Ac-

celerated degenerate-kernel methods for linear integral equations. Journal of Computational and Applied Mathematics, 58(2):135– 149, March 30, 1995. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042793E0273O.

Demichelis:1995:UMQ

[Dem95] V. Demichelis. The use of modified quasi-interpolatory splines for the solution of the Prandtl equation. Journal of Computational and Applied Mathematics, 59(3): 329–338, May 30, 1995. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279400036Z.

DeLillo:1997:NCM

[DEP97] Thomas K. DeLillo, Alan R.

Elcrat, and John A. Pfaltzgraff. Numerical conformal mapping methods based on Faber series. Journal of Computational and Applied Mathematics, 83(2):205– 236, October 7, 1997. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S037704279700099X.

Dette:1995:MCF

[Det95] Holger Dette. On the minimum of the Christoffel function. Journal of Computational and Applied Mathematics, 65(1–3):85–96, December 29, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042795001026.

Deren:1991:DSI

[DF91] Wang Deren and Zhao Fengguang. On the determination of the safe initial approximation for the Durand–Kerner algorithm. Journal of Computational and Applied Mathematics, 38(1–3):447–456, December 23, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279190188P.

Deren:1995:TSP

[DF95] Wang Deren and Zhao Fengguang. The theory of Smale’s point estimation and its applications. Journal of Computational and Applied Mathematics, 60(1–2): 253–269, June 20, 1995. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279400095I.

Diaconis:1992:AWH

[DG92] Persi Diaconis and Ron Graham. An affine walk on the hypercube. Journal of Computational and Applied Mathematics, 41(1–2):215– 235, August 20, 1992. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279290251R.

Delgado:1993:OPC

[DG93] Francisco-Javier Muñoz Delgado and Victoriano Ramírez González. Orthogonal polynomials and conservative approximation. Journal of Computational and Applied Mathematics, 49(1–3):217– 223, December 31, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042793901533.

DeVylder:1995:CRM

[DGC95] F. De Vylder, M. Goovaerts, and H. Cossette. Classical regression model under zero-excess assumptions. Journal of Computational and Applied Mathematics, 64(1–2):189–196, November 30, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042795000151.

Du:1995:AFE

[DGH95] Qiang Du, Max D. Gunzburger, and L. Steven Hou. Analysis and finite element approximation of optimal control problems for a Ladyzhenskaya model for stationary, incompressible, viscous flows. Journal of Computational and Applied Mathematics, 61(3):323– 343, August 21, 1995. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0377042795000720.

Dancer:1994:SPS

[DH94] E. N. Dancer and P. Hess. The symmetry of positive solutions of periodicparabolic problems. Journal of Computational and Applied Mathematics, 52(1–3): 81–89, July 20, 1994. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0377042794903506.

DeSchepper:1995:GRT

[DH95] A. De Schepper and B. Heijnen. General restrictions on tail probabilities. Journal of Computational and Applied Mathematics, 64(1–2):177–188, November 30, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042795000143.

Dekker:1994:PAS

[DHP94] T. J. Dekker, W. Hoffmann, and K. Potma. Parallel algorithms for solving large linear systems. Journal of Computational and Applied Mathematics, 50(1–3):221– 232, May 20, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042794903026.

Diblik:1998:POS

[Dib98] Josef Diblík. Positive and oscillating solutions of differential equations with delay in critical case. Journal of Computational and Applied Mathematics, 88(1):185– 202, February 23, 1998. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042797002173.

Dick:1991:SOF

[Dic91] Erik Dick. Second-order formulation of a multigrid method for steady Euler equations through defectcorrection. Journal of Computational and Applied Mathematics, 35(1–3):159– 168, June 26, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279190204W.

Dieci:1993:DDL

[Die93] Luca Dieci. On the decoupling of dichotomic linear Hamiltonians. Considerations on integrating symmetric differential Riccati equations. Journal of Computational and Applied Mathematics, 45(1–2):47– 63, April 8, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279390264C.

Dierckx:1994:BRB

[Die94a] P. Dierckx. Book review: Box splines C. de Boor, K. Höllig, S. Riemenschneider: Applied Mathematical Sciences, vol. 98 Springer-Verlag, Berlin, 1993, xvii + 200 pages, 57 figs. Price: DM 60 (HC) ISBN 3-540-94101-0. Journal of Computational and Applied Mathematics, 50(1–3):N4– N5, May 20, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042794903409.

Diethelm:1994:UCO

[Die94b]  Kai Diethelm.  Uniform

convergence of optimal order quadrature rules for Cauchy principal value integrals. Journal of Computational and Applied Mathematics, 56(3):321–329, December 30, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042794900868.

Diethelm:1995:GQF

[Die95] Kai Diethelm. Gaussian quadrature formulae of the third kind for Cauchy principal value integrals: Basic properties and error estimates. Journal of Computational and Applied Mathematics, 65(1–3):97–114, December 29, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042795001034.

Diethelm:1996:DCL

[Die96] Kai Diethelm. A definiteness criterion for linear functionals and its application to Cauchy principal value quadrature. Journal of Computational and Applied Mathematics, 66(1–2):167– 176, January 31, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0377042795001751.

Diethelm:1997:NEB

[Die97] Kai Diethelm. New error bounds for modified quadrature formulas for Cauchy principal value integrals. Journal of Computational and Applied Mathematics, 82(1–2):93–104, September 15, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042797000459.

Diethelm:1998:EBS

[Die98] Kai Diethelm. Error bounds for spline-based quadrature methods for strongly singular integrals. Journal of Computational and Applied Mathematics, 89(2):257– 261, March 16, 1998. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042797002458.

See erratum [Die02].

Diethelm:1999:MPE

[Die99] Kai Diethelm. A method for the practical evaluation of the Hilbert transform on the real line. Journal of Computational and Applied Mathematics, 112(1–2):45– 53, November 30, 1999. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042799002125.

Diethelm:2002:EEB

[Die02] Kai Diethelm. Erratum to “Error bounds for splinebased quadrature methods for strongly singular integrals” [J. Comput. Appl. Math. 89 (1998) 257]. Journal of Computational and Applied Mathematics, 142(2):449–450, May 15, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042702003564.

See [Die98].

Defez:1998:SAH

[DJ98] Emilio Defez and Lucas Jódar. Some applications of the Hermite matrix polynomials series expansions. Journal of Computational and Applied Mathematics, 99(1–2):105–117, November 16, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042798001496.

Descloux:1991:BSR

[DJR91] J. Descloux, Y. Jaccard, and M. V. Romerio. A bidimensional stability result for aluminium electrolytic cells. Journal of Computational and Applied Mathematics, 38(1–3):77–85, December 23, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279190162D.

Druskin:1992:LOS

[DK92] Vladimir Druskin and Leonid Knizhnerman. The Lanczos optimization of a splittingup method to solve homogeneous evolutionary equations. Journal of Computational and Applied Mathematics, 42(2):221–231, October 12, 1992. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279290076A.

Druskin:1994:ALM

[DK94] Vladimir Druskin and Leonid Knizhnerman. On application of the Lanczos method to solution of some partial differential equations. Journal of Computational and Applied Mathematics, 50(1–3):255–262, May 20, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042794903050.

Dimova:1998:NAR

[DKKV98] Stefka Dimova, Michael Kaschiev, Milena Koleva, and Daniela Vasileva. Numerical analysis of radially nonsymmetric blowup solutions of a nonlinear parabolic problem. Journal of Computational and Applied Mathematics, 97(1–2):81–97, September 24, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042798001034.

Deza:1994:ACPa

[DL94a] Michel Deza and Monique Laurent. Applications of cut polyhedra — I. Journal of Computational and Applied Mathematics, 55(2):191–216, November 21, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042794900205.

Deza:1994:ACPb

[DL94b] Michel Deza and Monique Laurent. Applications of cut polyhedra — II. Journal of Computational and Applied Mathematics, 55(2):217–247, November 21, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042794900213.

Dagnino:1996:NEC

[DL96] C. Dagnino and P. Lamberti. Numerical evaluation of Cauchy principal value integrals based on local spline approximation operators. Journal of Computational and Applied Mathematics, 76(1–2):231–238, December 17, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042796001057.

DelBuono:1999:NMH

[DL99] N. Del Buono and L. Lopez. Numerical methods for Hermitian unitary differential systems. Journal of Computational and Applied Mathematics, 111(1–2):133–145, November 15, 1999. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042799001375.

Ysern:1998:SAO

[dlCYL98] B. de la Calle Ysern and G. López Lagomasino. Strong asymptotics of orthogonal polynomials with varying measures and Hermite– Padé approximants. Journal of Computational and Applied Mathematics, 99(1–2): 91–103, November 16, 1998.

CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042798001484.

Dattoli:1999:GPA

[DLMT99] G. Dattoli, S. Lorenzutta, A. M. Mancho, and A. Torre. Generalized polynomials and associated operational identities. Journal of Computational and Applied Mathematics, 108(1–2):209– 218, August 15, 1999. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042799001119.

Diele:1998:OSS

[DLP98] F. Diele, L. Lopez, and T. Politi. One step semiexplicit methods based on the Cayley transform for solving isospectral flows. Journal of Computational and Applied Mathematics, 89(2):219–223, March 16, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042797002367.

Dyn:1992:BCP

[DLR92] Nira Dyn, David Levin, and Samuel Rippa. Boundary correction for piecewise linear interpolation defined over data-dependent triangulations. Journal of Computational and Applied Mathematics, 39(2):179– 192, March 30, 1992. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/037704279290128K.

Draux:1990:VOP

[DM90] André Draux and Abdallah Maanaoui. Vector orthogonal polynomials. Journal of Computational and Applied Mathematics, 32(1–2):59– 68, November 26, 1990. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/037704279090416W.

Dellnitz:1994:GME

[DM94] Michael Dellnitz and Ian Melbourne. Generic movement of eigenvalues for equivariant self-adjoint matrices. Journal of Computational and Applied Mathematics, 55(3):249–259, November 30, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042794900329.

Diligenti:1997:IEB

[DM97a] M. Diligenti and G. Monegato. Integral evaluation in the BEM solution of (hyper)singular integral equations. 2D problems on polygonal domains. Journal of Computational and Applied Mathematics, 81(1):29–57, June 1, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042797000071.

Drozdowicz:1997:ABS

[DM97b] Andrzej Drozdowicz and Małgorzata Migda. Asymptotic behavior of solutions of the m-th-order nonhomogeneous difference equations. Journal of Computational and Applied Mathematics, 81(1):75–81, June 1, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042797000113.

Diaz-Mendoza:1999:TPP

[DMGVO99] C. Díaz-Mendoza, P. González-Vera, and R. Orive. Twopoint Padé-type approximation to the Cauchy transform of certain strong distributions. Journal of Computational and Applied Mathematics, 105(1–2):229– 243, May 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042799000436.

deMarkus:1999:SNI

[dMM99] Alicia Serfaty de Markus and Ronald E. Mickens. Suppression of numerically induced chaos with nonstandard finite difference schemes. Journal of Computational and Applied Mathematics, 106(2):317–324, June 30, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704279900076X.

deMorales:1998:SOG

[dMPP98] María Álvarez de Morales, Teresa E. Pérez, and Miguel A. Piñar. Sobolev orthogonality for the Gegenbauer polynomials {C_{n}^{(−N+12)}}_{n} ≥

0. Journal of Computational and Applied Mathematics, 100(1):111–120, November 30, 1998. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042798001241.

Dargahi-Noubary:1992:SSR

[DNN92] G. R. Dargahi-Noubary and A. Nanthakumar. Stressstrength reliability for designs based on large historic values of stress. Journal of Computational and Applied Mathematics, 40(1):43–53, June 12, 1992. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042792900405.

Dargahi-Noubary:1998:DEM

[DNS98] G. R. Dargahi-Noubary and Yixun Shi. Difference equation models for estimating athletic records. Journal of Computational and Applied Mathematics, 91(1):107– 122, April 24, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042798000326.

Davydov:1998:AOB

[DNZ98] O. V. Davydov, G. Nürnberger, and F. Zeilfelder. Approximation order of bivariate spline interpolation for arbitrary smoothness. Journal of Computational and Applied Mathematics, 90(2): 117–134, April 17, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042798000041.

deOliveira:1990:CMM

[dO90] Paula de Oliveira. On a class of multistep methods with strong contractivity properties. Journal of Computational and Applied Mathematics, 30(1):11–19, April 10, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279090002H.

Demkowicz:1991:SES

[DOAG91] L. Demkowicz, J. T. Oden, M. Ainsworth, and P. Geng. Solution of elastic scattering problems in linear acoustics using h–p boundary element methods. Journal of Computational and Applied Mathematics, 36(1):29–63, August 20, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042791902259.

Doha:1998:UCM

[Doh98]  E. H. Doha.  The ul-

traspherical coefficients of the moments of a generalorder derivative of an infinitely differentiable function. Journal of Computational and Applied Mathematics, 89(1):53–72, March 9, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042797002288.

deOliveira:1997:MNG

[dOP97] P. de Oliveira and F. Patricio. Monotonicity on nonuniform grids. Journal of Computational and Applied Mathematics, 80(2):

171–181, May 5, 1997. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042797000034.

Dorfi:1999:IRH

[Dor99a]  Ernst A. Dorfi.  Im-

plicit radiation hydrodynamics for 1D-problems. Journal of Computational and Applied Mathematics, 109(1–2):153–171, September 30, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042799001570.

Dorfmann:1999:WWD

[Dor99b] A. A. Dorfmann. Water waves diffraction by a circular plate. Journal of Computational and Applied Mathematics, 110(2):287–304, October 30, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042799002721.

Dahmen:1994:HB

[DOS94] Wolfgang Dahmen, Peter Oswald, and Xi-Quan Shi. C 1-hierarchical bases. Journal of Computational and Applied Mathematics, 51(1): 37–56, May 30, 1994. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/037704279200002Q.

Dostal:1995:DBD

[Dos95]  Zdenek Dostál.  Dual-

ity based domain decomposition with proportioning for the solution of free boundary problems. Journal of Computational and Applied Mathematics, 63(1–3):203–208, November 20, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042796809125.

Douak:1996:ROP

[Dou96] Khalfa Douak. The relation of the d-orthogonal polynomials to the Appell polynomials. Journal of Computational and Applied Mathematics, 70(2): 279–295, June 28, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0377042795002111.

DeLillo:1993:EDH

[DP93a] Thomas K. DeLillo and John A. Pfaltzgraff. Extremal distance, harmonic measure and numerical conformal mapping. Journal of Computational and Applied Mathematics, 46(1–2):103– 113, June 14, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279390289N.

Drozdowicz:1993:ABS

[DP93b] Andrzej Drozdowicz and Jerzy Popenda. Asymptotic behavior of solutions of difference equations of second order. Journal of Computational and Applied Mathematics, 47(2): 141–149, ???? 1993. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279390001R.

Dimitrov:1994:NCNa

[DP94a] Dimitar K. Dimitrov and George M. Philipps. A note on convergence of Newton interpolating polynomials. Journal of Computational and Applied Mathematics, 51(1):127–130, May 30, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042794000573.  See

note [DP94b].

Dimitrov:1994:NCNb

[DP94b] Dimitar K. Dimitrov and George M. Phillips. A note on convergence of Newton interpolating polynomials J. Comput. Appl. Math. 51 (1) (1994) 127–130. Journal of Computational and Applied Mathematics, 51(3): 411, June 30, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042794000611.

See [DP94a].

Dombrowski:1995:OPS

[DP95] J. Dombrowski and S. Pedersen. Orthogonal polynomials, spectral measures, and absolute continuity. Journal of Computational and Applied Mathematics, 65(1–3):115–124, December 29, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042795001042.

Denk:1997:ISH

[DP97] G. Denk and Chr. Penski. Integration schemes for highly oscillatory DAEs with applications to circuit simulation. Journal of Computational and Applied Mathematics, 82(1–2):79– 91, September 15, 1997. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042797000423.

Dagnino:1998:CRB

[DPS98] C. Dagnino, S. Perotto, and E. Santi. Convergence of rules based on nodal splines for the numerical evaluation of certain 2D Cauchy principal value integrals. Journal of Computational and Applied Mathematics, 89(2):225– 235, March 16, 1998. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042797002379.

Djidjeli:1995:NMS

[DPTW95] K. Djidjeli, W. G. Price, E. H. Twizell, and Y. Wang. Numerical methods for the solution of the thirdand fifth-order dispersive Korteweg–de Vries equations. Journal of Computational and Applied Mathematics, 58(3):307–336, April 20, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
037704279400005L.

DeFreitasPinto:1990:OLE

[DR90] Ricardo L. U. De Freitas Pinto and Atair Rios Neto. An optimal linear estimation approach to solve systems of linear algebraic equations. Journal of Computational and Applied Mathematics, 33(3):261–268, December 31, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042705800032.

Dick:1994:MSJ

[DR94] Erick Dick and Kris Riemslagh. Multi-staging of Jacobi relaxation to improved smoothing properties of multigrid methods for steady Euler equations. Journal of Computational and Applied Mathematics, 50(1–3):241–254, May 20, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042794903042.

Dick:1995:MSJ

[DR95] Erik Dick and Kris Riemslagh. Multi-staging of Jacobi relaxation in multigrid methods for steady Euler equations, II. Journal of Computational and Applied Mathematics, 59(3): 339–348, May 30, 1995. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/037704279400022S.

daRocha:1999:QAR

[dR99]  Zélia da Rocha.  QD-

algorithms and recurrence relations for biorthogonal polynomials. Journal of Computational and Applied Mathematics, 107(1):53–72, July 15, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042799000758.

Driver:1995:NQH

[Dri95] Kathy Driver. Nondiagonal quadratic Hermite– Padé approximation to the exponential function. Journal of Computational and Applied Mathematics, 65(1–3):125–134, December 29, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042795001069.

Drmota:1990:AUD

[Drm90] Michael Drmota. Applications of uniformly distributed functions and sequences in statistic ergodic measuring techniques. Journal of Computational and Applied Mathematics, 31(1): 57–62, July 24, 1990. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279090335W.

Dyer:1995:HDP

[DRW95] M. E. Dyer, W. O. Riha, and J. Walker. A hybrid dynamic programming/branch-and-bound algorithm for the multiplechoice knapsack problem. Journal of Computational and Applied Mathematics, 58(1):43–54, March 20, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042793E0264M.

Dagnino:1990:SPQ

[DS90] C. Dagnino and E. Santi. Spline product quadrature rules for Cauchy singular integrals. Journal of Computational and Applied Mathematics, 33(2):133–140, December 21, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042790903635.

Dagnino:1991:CSP

[DS91] C. Dagnino and E. Santi. On the convergence of spline product quadratures for Cauchy principal value integrals. Journal of Computational and Applied Mathematics, 36(2):181–187, August 27, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279190025F.

deSturler:1996:NKM

[dS96a]  E. de Sturler.  Nested

Krylov methods based on GCR. Journal of Computational and Applied Mathematics, 67(1):15–41, February 20, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042794001235.

Dzurina:1996:AAO

[DS96b] Jozef Dzurina and Vincent Soltés. Asymptotic analysis of ODEs. Journal of Computational and Applied Mathematics, 67(2):301– 307, March 29, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042794001308.

deSilanes:1997:CEE

[dS97a] M. C. López de Silanes.

Convergence and error estimates for ( m, l, s )- splines. Journal of Computational and Applied Mathematics, 87(2):373–384, December 23, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042797002008.

deSwart:1997:SOS

[dS97b] Jacques J. B. de Swart. A simple ODE solver based on 2-stage Radau IIA. Journal of Computational and Applied Mathematics, 84(2):277–280, October 28, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042797001416.

Dorfmann:1998:WWD

[DS98] A. A. Dorfmann and A. A. Savvin. Water waves diffraction by a submerged sphere and dual integral transforms. Journal of Computational and Applied Mathematics, 91(1):1–30, April 24, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042798000090.

Diele:1999:IFI

[DS99] Fasma Diele and Ivonne Sgura. Isospectral flows and the inverse eigenvalue problem for Toeplitz matrices. Journal of Computational and Applied Mathematics, 110(1):25–43, October 15, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042799001934.

Deif:1995:SEA

[DSH95] A. S. Deif, N. P. Seif, and S. A. Hussein. Sylvester’s equation: accuracy and computational stability. Journal of Computational and Applied Mathematics, 61(1): 1–11, July 20, 1995. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042794000534.

deSwart:1997:CEE

[dSS97] Jacques J. B. de Swart and Gustaf Söderlind. On the construction of error estimators for implicit Runge–Kutta methods. Journal of Computational and Applied Mathematics, 86(2):347–358, December 10, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042797001660.

Dimov:1993:MCA

[DT93] Ivan T. Dimov and Ognyan I.

Tonev. Monte Carlo algorithms: performance analysis for some computer architectures. Journal of Computational and Applied Mathematics, 48(3):253– 277, November 30, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042793900246.

Driscoll:1996:PWE

[DT96] Tobin A. Driscoll and Lloyd N. Trefethen. Pseudospectra for the wave equation with an absorbing boundary. Journal of Computational and Applied Mathematics, 69(1):125– 142, April 30, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042795000216.

Driver:1998:AZD

[DT98] Kathy A. Driver and Nico M. Temme. Asymptotics and zero distribution of Padé polynomials associated with the exponential function. Journal of Computational and Applied Mathematics, 89(1):97–114, March 9, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042797002306.

Djidjeli:1993:NMS

[DTB93] K. Djidjeli, E. H. Twizell, and A. Boutayeb. Numerical methods for special nonlinear boundary-value problems of order 2 m. Journal of Computational and Applied Mathematics, 47(1):35–45, June 28, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279390088S.

Dubach:1996:ABC

[Dub96]  Eric Dubach.  Artificial

boundary conditions for diffusion equations: Numerical study. Journal of Computational and Applied Mathematics, 70(1): 127–144, June 14, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042795001409.

Dubrulle:1997:AVI

[Dub97] Augustin A. Dubrulle. A QR algorithm with variable iteration multiplicity. Journal of Computational and Applied Mathematics, 86(1):125–139, November 28, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042797001519.

Dubeau:1998:RED

[Dub98]  François Dubeau.  n-th

Root extraction: Double iteration process and Newton’s method. Journal of Computational and Applied Mathematics, 91(2): 191–198, May 4, 1998. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042798000338.

Dunster:1997:EAU

[Dun97]  T. M. Dunster.  Er-

ror analysis in a uniform asymptotic expansion for the generalised exponential integral. Journal of Computational and Applied Mathematics, 80(1):127– 161, April 14, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042797000265.

Dunster:1998:AER

[Dun98] T. M. Dunster. Asymptotics of the eigenvalues of the rotating harmonic oscillator. Journal of Computational and Applied Mathematics, 93(1):45–73, July 2, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042798000703.

Dupacova:1994:ASP

[Dup94] Jitka Dupacová. Applications of stochastic programming under incomplete information. Journal of Computational and Applied Mathematics, 56(1–2):113– 125, December 20, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042794903824.

Duran:1998:P

[Dur98] Antonio J. Duran. Preface. Journal of Computational and Applied Mathematics, 99(1–2):1, November 16, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042798002337.

DeSchepper:1999:FEM

[DV99] Hennie De Schepper and Roger Van Keer. A finite element method for elliptic eigenvalue problems in a multi-component domain in 2D with non-local Dirichlet transition conditions. Journal of Computational and Applied Mathematics, 111(1–2):253–265, November 15, 1999. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042799001478.

DeSamblanx:1996:DVP

[DVB96] Gorik De Samblanx, Marc Van Barel, and Adhemar Bultheel. Duality in vector Padé–Hermite approximation problems. Journal of Computational and Applied Mathematics, 66(1–2):153– 166, January 31, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0377042795001867.

DeMeyer:1990:NTM

[DVV90] H. De Meyer, J. Vanthournout, and G. Vanden Berghe. On a new type of mixed interpolation. Journal of Computational and Applied Mathematics, 30(1): 55–69, April 10, 1990. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/037704279090005K.

DeMeyer:1990:EEM

[DVVV90] H. De Meyer, J. Vanthournout, G. Vanden Berghe, and A. Vanderbauwhede. On the error estimation for a mixed type of interpolation. Journal of Computational and Applied Mathematics, 32(3):407–415, December 10, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042790900452.

DeMeyer:1997:GMI

[DVVV97] H. De Meyer, G. Vanden Berghe, T. Van Hecke, and M. Van Daele. On the generation of mono-implicit Runge–Kutta–Nyström methods by mono-implicit Runge– Kutta methods. Journal of Computational and Applied Mathematics, 87 (1):147–167, December 18, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042797001830.

DeCoster:1994:DST

[DW94] C. De Coster and M. Willem.

Density, spectral theory and homoclinics for singular Sturm–Liouville systems. Journal of Computational and Applied Mathematics, 52(1–3):45–70, July 20, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042794903484.

Dziok:1999:CFD

[Dzi99] J. Dziok. Classes of functions defined by certain differential-integral operators. Journal of Computational and Applied Mathematics, 105(1–2):245–255, May 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704279900014X.

Dehesa:1994:OPD

[DZY^{+}94] J. S. Dehesa, A. Zarzo, R. J.

Yañez, B. Germano, and P. E. Ricci. Orthogonal polynomials and differential equations in neutrontransport and radiativetransfer theories. Journal of Computational and Applied Mathematics, 50(1–3):197– 206, May 20, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279490300X.

Erhel:1996:RGP

[EBP96] Jocelyne Erhel, Kevin Burrage, and Bert Pohl. Restarted GMRES preconditioned by deflation. Journal of Computational and Applied Mathematics, 69(2):303– 318, May 23, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279500047X.

Elmahi:1999:FVS

[EBV^{+}99] I. Elmahi, F. Benkhaldoun, R. Vilsmeier, O. Gloth, A. Patschull, and D. Hänel. Finite volume simulation of a droplet flame ignition on unstructured meshes. Journal of Computational and Applied Mathematics, 103(1):187–205, March 15, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042798002520.

Eddy:1990:RNG

[Edd90] William F. Eddy. Random number generators for parallel processors. Journal of Computational and Applied Mathematics, 31(1):63–71, July 24, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279090336X.

Essawy:1999:DWW

[EEAES99] A. H. Essawy, H. El-Arabawy, and Sh. M. El-Sawaby. Diffraction of water waves due to a presence of two headlands. Journal of Computational and Applied Mathematics, 104(1):27–39, April 15, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042798002003.

Evans:1993:ROW

[EEKL93] W. D. Evans, W. N. Everitt, K. H. Kwon, and L. L. Littlejohn. Real orthogonalizing weights for Bessel polynomials. Journal of Computational and Applied Mathematics, 49(1–3):51–57, December 31, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279390134W.

Elsner:1997:DCM

[EFL97] L. Elsner, A. Fasse, and E. Langmann. A divideand-conquer method for the tridiagonal generalized eigenvalue problem. Journal of Computational and Applied Mathematics, 86(1):141–148, November 28, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042797001520.

El-Gebeily:1993:EUS

[EGBE93] M. A. El-Gebeily, A. Boumenir, and M. B. M. Elgindi. Existence and uniqueness of solutions of a class of two-point singular nonlinear boundary value problems. Journal of Computational and Applied Mathematics, 46(3):345–355, June 22, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042793900316.

Eichenauer-Herrmann:1990:CME

[EH90] J. Eichenauer-Herrmann. Conditional minimax estimates. Journal of Computational and Applied Mathematics, 29(3):369–371, March 15, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.

com/science/article/pii/
037704279090018U.

Engl:1991:UCR

[EH91] Heinz W. Engl and Günther Hodina. Uniform convergence of regularization methods for linear ill-posed problems. Journal of Computational and Applied Mathematics, 38(1–3):87– 103, December 23, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
037704279190163E.

Eichenauer-Herrmann:1992:CIC

[EH92a] Jürgen Eichenauer-Herrmann.

Construction of inversive congruential pseudorandom number generators with maximal period length. Journal of Computational and Applied Mathematics, 40(3):345–349, July 24, 1992. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042792901909.

Eichenauer-Herrmann:1992:RDQ

[EH92b] Jürgen Eichenauer-Herrmann.

A remark on the discrepancy of quadratic congruential pseudorandom numbers. Journal of Computational and Applied Mathematics, 43(3):383–387, December 2, 1992. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279290023Q.

Eichenauer-Herrmann:1994:DQC

[EH94] Jürgen Eichenauer-Herrmann. On the discrepancy of quadratic congruential pseudorandom numbers with power of two modulus. Journal of Computational and Applied Mathematics, 53(3):371–376, August 30, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042794900647.

Eichenauer-Herrmann:1995:QCP

[EH95] Jürgen Eichenauer-Herrmann. Quadratic congruential pseudorandom numbers: distribution of triples. Journal of Computational and Applied Mathematics, 62 (2):239–253, September 20, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042794001056.

Eichenauer-Herrmann:1997:QCP

[EH97a] Jürgen Eichenauer-Herrmann. Quadratic congruential pseudorandom numbers: Distribution of lagged pairs. Journal of Computational and Applied Mathematics, 79(1): 75–85, March 3, 1997. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042796001586.

Ezquerro:1997:RCC

[EH97b] J. A. Ezquerro and M. A.

Hernández. Relaxing convergence conditions for an inverse-free Jarratt-type approximation. Journal of Computational and Applied Mathematics, 83(1):131– 135, September 30, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042797001027.

Ezquerro:1998:ACS

[EH98] J. A. Ezquerro and M. A.

Hernández. Avoiding the computation of the second Fréchet-derivative in the convex acceleration of Newton’s method. Journal of Computational and Applied Mathematics, 96(1):1– 12, September 3, 1998. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042798000831.

Eichenauer-Herrmann:1990:UBB

[EHG90] J. Eichenauer-Herrmann and H. Grothe. Upper bounds for the Beyer ratios of linear congruential generators. Journal of Computational and Applied Mathematics, 31(1):73–80, July 24, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279090337Y.

Eichenauer-Herrmann:1990:LSN

[EHGNT90] J. Eichenauer-Herrmann, H. Grothe, H. Niederreiter, and A. Topuzoğlu. On the lattice structure of a nonlinear generator with modulus 2^{α}. Journal of Computational and Applied Mathematics, 31(1):81–85, July 24, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279090338Z.

Embrechts:1995:IWA

[EHK^{+}95] Paul Embrechts, Agnes M. Herzberg, Heidi K. Kalbfleisch, William N. Traves, and J. Robertson Whitla. An introduction to wavelets with applications to Andrews; plots. Journal of Computational and Applied Mathematics, 64(1–2):41– 56, November 30, 1995. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042795000054.

Eichenauer-Herrmann:1991:DQC

[EHN91] Jürgen Eichenauer-Herrmann and Harald Niederreiter. On the discrepancy of quadratic congruential pseudorandom numbers. Journal of Computational and Applied Mathematics, 34(2): 243–249, April 4, 1991. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279190046M.

Ehrenmark:1994:CUV

[Ehr94] Ulf Ehrenmark. On computing uniformly valid approximations for viscous waves on a plane beach. Journal of Computational and Applied Mathematics, 50(1–3):263– 281, May 20, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042794903069.

Ehrenmark:1995:NIT

[Ehr95a] Ulf Torsten Ehrenmark.

The numerical inversion of two classes of Kontorovich– Lebedev transform by direct quadrature. Journal of Computational and Applied Mathematics, 61(1):43–72, July 20, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042794000523.

Ehrich:1995:ABS

[Ehr95b]  Sven Ehrich.  Asymp-

totic behaviour of Stieltjes polynomials for ultraspherical weight functions. Journal of Computational and Applied Mathematics, 65(1–3):135–144, December 29, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042795001077.

Ehrenmark:1996:ECT

[Ehr96a] Ulf Torsten Ehrenmark. On the error and its control in a two-parameter generalised Newton–Cotes rule. Journal of Computational and Applied Mathematics, 75(1):171–195, November 12, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042796000854.

Ehrich:1996:NPC

[Ehr96b]  Sven Ehrich.  A note

on Peano constants of Gauss–Kronrod quadrature schemes. Journal of Computational and Applied Mathematics, 66(1–2):177–183, January 31, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042795001735.

Ehrich:1998:OPC

[Ehr98] Sven Ehrich. On orthogonal polynomials for certain nondefinite linear functionals. Journal of Computational and Applied Mathematics, 99(1–2):119–128, November 16, 1998. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042798001502.

Ezquerro:1998:RCM

[EHS98] J. A. Ezquerro, M. A.

Hernández, and M. A. Salanova. Remark on the convergence of the midpoint method under mild differentiability conditions. Journal of Computational and Applied Mathematics, 98(2):305–309, October 19, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042798001307.

Eichenauer-Herrmann:1990:PLC

[EHT90] J. Eichenauer-Herrmann and A. Topuzoğlu. On the period length of congruential pseudorandom number sequences generated by inversions. Journal of Computational and Applied Mathematics, 31(1):87–96, July 24, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042790903392.

Elzoheiry:1991:IIS

[EIM91] H. Elzoheiry, L. Iskandar, and M. Sh. El-Deen Mohamedein. Iterative implicit schemes for the twoand three-dimensional sine-Gordon equation. Journal of Computational and Applied Mathematics, 34(2): 161–170, April 4, 1991. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/037704279190039M.

Eirola:1993:ABE

[Eir93] Timo Eirola. Aspects of backward error analysis of numerical ODEs. Journal of Computational and Applied Mathematics, 45(1–2): 65–73, April 8, 1993. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/037704279390265D.

Erbe:1994:BVP

[EK94] L. H. Erbe and Qingkai Kong. Boundary value problems for singular secondorder functional differential equations. Journal of Computational and Applied Mathematics, 53(3):377– 388, August 30, 1994. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042794900655.

Elnagar:1996:CSS

[EK96] Gamal N. Elnagar and M. Kazemi. Chebyshev spectral solution of nonlinear Volterra–Hammerstein integral equations. Journal of Computational and Applied Mathematics, 76(1– 2):147–158, December 17, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042796000982.

Elnagar:1998:PLB

[EK98] Gamal N. Elnagar and Mohammad A. Kazemi. Pseudospectral Legendre-based optimal computation of nonlinear constrained variational problems. Journal of Computational and Applied Mathematics, 88(2):363– 375, March 2, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042797002252.

Ekhad:1991:OLP

[Ekh91] Shalosh B. Ekhad. A oneline proof of the Habsieger– Zeilberger G_{2} constant term identity. Journal of Computational and Applied Mathematics, 34(1):133– 134, February 10, 1991. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/037704279190154C.

Elbert:1992:ABA

[Elb92] Á. Elbert. Asymptotic behaviour of the analytic solution of the differential equation y′(t) + y(qt) = 0 as q → 1^{−}. Journal of Computational and Applied Mathematics, 41(1–2):5–22, August 20, 1992. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279290233N.

Ellacott:1994:TMA

[Ell94]  S. W. Ellacott.  Tech-

niques for the mathematical analysis of neural networks. Journal of Computational and Applied Mathematics, 50(1–3):283–297, May 20, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042794903077.

Elliott:1995:TAH

[Ell95] David Elliott. Three algorithms for Hadamard finitepart integrals and fractional derivatives. Journal of Computational and Applied Mathematics, 62 (3):267–283, September 29, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042794001135.

ElMennaoui:1994:ABI

[ElM94] Omar ElMennaoui. Asymptotic behaviour of integrated semigroups. Journal of Computational and Applied Mathematics, 54(3):351– 369, October 20, 1994. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042794902569.

Elnagar:1997:SCS

[Eln97] Gamal N. Elnagar. Statecontrol spectral Chebyshev parameterization for linearly constrained quadratic optimal control problems. Journal of Computational and Applied Mathematics, 79(1): 19–40, March 3, 1997. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042796001343.

Eloe:1998:PSB

[Elo98] Paul W. Eloe. Positive solutions of boundary-value problems for disfocal ordinary differential equations. Journal of Computational and Applied Mathematics, 88(1):71–78, February 23, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042797002069.

Elster:1999:RWD

[Els99] Clemens Elster. Recovering wavefronts from difference measurements in lateral shearing interferometry. Journal of Computational and Applied Mathematics, 110(1):177–180, October 15, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042799002320.

Everitt:1993:OPA

[ELW93] W. N. Everitt, L. L. Littlejohn, and S. C. Williams. Orthogonal polynomials and approximation in Sobolev spaces. Journal of Computational and Applied Mathematics, 48(1–2):69–90, October 29, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042793903164.

Everitt:1995:SFK

[ELW95] W. N. Everitt, L. L. Littlejohn, and R. Wellman. The symmetric form of the Koekoeks’ Laguerre type differential equation. Journal of Computational and Applied Mathematics, 57(1–2):115–121, February 20, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042793E0238H.

Abd-el-Malek:1991:ASG

[eM91] Mina B. Abd el Malek.

Approximate solution of gravity-affected flow from planar sluice gate at high Froude number. Journal of Computational and Applied Mathematics, 35(1–3): 83–97, June 26, 1991. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279190198S.

Abd-el-Malek:1994:FWL

[eM94a] Mina B. Abd el Malek. Flow in a waterfall with large Froude number. Journal of Computational and Applied Mathematics, 50(1–3): 87–98, May 20, 1994. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042794902917.

Everitt:1994:GBF

[EM94b] W. N. Everitt and C. Markett. On a generalization of Bessel functions satisfying higher-order differential equations. Journal of Computational and Applied Mathematics, 54(3):325– 349, October 20, 1994. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0377042794902550.

Ehrich:1998:MIB

[EM98a] Sven Ehrich and Giuseppe Mastroianni. Marcinkiewicz inequalities based on Stieltjes zeros. Journal of Computational and Applied Mathematics, 99(1–2):129–141, November 16, 1998. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042798001514.

Endresen:1998:FSC

[EM98b] Lars Petter Endresen and Jan Myrheim. A formula for steplength control in numerical integration. Journal of Computational and Applied Mathematics, 90(2):263– 264, April 17, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042797002380.

Abd-el-Malek:1991:GMA

[eMB91] Mina B. Abd el Malek and Nagwa A. Badran. Group method analysis of steady free-convective laminar boundary-layer flow on a nonisothermal vertical circular cylinder. Journal of Computational and Applied Mathematics, 36(2):227– 238, August 27, 1991. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279190028I.

Emerson:1994:BRE

[Eme94a] D. R. Emerson. Book review: Exploring interiorpoint linear programming: Algorithms and software: Ami Arbel, M.I.T. Press, Fitzroy House, 11, Chenies Street, London WC1E 7ET, ISBN 0-262-51073-

1. Journal of Computational and Applied Mathematics, 54(3):N5–N6, October 20, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042794902666.

Emerson:1994:BRI

[Eme94b] D. R. Emerson. Book review: Incompressible computational fluid dynamics: trends and advances: M. D. Gunzbuger and R. A.

Nicolaides: Cambridge University Press, The Edinburgh Building, Cambridge CB2 2RU, England, Price: 35 Euros (HB), ISBN 0-521-40407-X. Journal of Computational and Applied Mathematics, 50(1–3):N3, May 20, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042794903387.

Emerson:1994:BRN

[Eme94c]  D. R. Emerson.  Book

review: Numerical simulation of 3-D incompressible unsteady viscous laminar flow numerical methods in fluid mechanics, volume 36 M. Deville, T.- H. Le and Y. Morchoisne: Vieweg, Friedr. Vieweg & Sohn, Verlagsgesellschaft mbH, Braunschweig, Weisbaden, Germany, ISBN 3-528-07636-4. Journal of Computational and Applied Mathematics, 50(1–3):N3– N4, May 20, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042794903395.

Emerson:1994:BRF

[Eme94d] D. R. Emerson. Book review: The finite element method: Basic concepts and applications: Darrell W.

Pepper and Juan C. Heinrich Taylor and Francis Ltd., 4 John Street, London WC1N 2ET, ISBN 1-56032-104-0. Journal of Computational and Applied Mathematics, 54(3):N4–N5, October 20, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279490264X.

Emerson:1994:BRH

[Eme94e] D. R. Emerson. Book review: The high performance Fortran handbook, Scientific and engineering computation series: Charles H. Koelbel, David B. Loveman, Robert S. Schreiber, Guy L. Steele Jr. and Mary E. Zosel, M.I.T. Press, Fitzroy House, 11, Chenies Street, London WC1E 7ET, England, ISBN 0-262-61094-9, January 1994, Price: $24.95 (Paperback). Journal of Computational and Applied Mathematics, 54(3):N4, October 20, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042794902631.

Emerson:1994:BRT

[Eme94f] D. R. Emerson. Book review: Theory and numerical modeling of turbulent gasparticle flows and combustion: Zhou Lixing, Times Mirror International Publishers Ltd., c/o Excel Logistics,, 3 Sheldon Way, Larkfield, Aylesford, Kent ME20 6SF, England, ISBN 0-8493-7721-8. Journal of Computational and Applied Mathematics, 54(3):N5, October 20, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042794902658.

Abd-el-Malek:1997:NPI

[eMM97a] Mina B. Abd el Malek and Malak N. Makar. Nonlinear progressive internal gravity wave on fluid of trapezoidal bottom. Journal of Computational and Applied Mathematics, 82(1–2):3–19, September 15, 1997. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042797000782.

Emmerich:1997:EPQ

[Emm97b] Frank Emmerich. Equidistribution properties of quadratic congruential pseudorandom numbers. Journal of Computational and Applied Mathematics, 79(2):207– 214, March 17, 1997. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042796001574.

Epperson:1991:KBM

[Epp91] James F. Epperson. A kernel-based method for parabolic equations with nonlinear convection terms. Journal of Computational and Applied Mathematics, 36(3):275–288, September 24, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042791900118.

Epperson:1992:FDS

[Epp92] James F. Epperson. Fast direct solvers for the heat equation in simple geometries. Journal of Computational and Applied Mathematics, 39(2):213– 225, March 30, 1992. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279290131G.

Engquist:1996:MPC

[ER96] Björn Engquist and Olof Runborg. Multi-phase computations in geometrical optics. Journal of Computational and Applied Mathematics, 74(1–2):175–192, November 5, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042796000234.

Erdelyi:1993:RTI

[Erd93] Tamás Erdélyi. Remeztype inequalities and their applications. Journal of Computational and Applied Mathematics, 47(2):167– 209, ???? 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279390003T.

Erno:1996:PVS

[Ern96] Scheiber Ernő. On the parallel version of the successive approximation method for quasilinear boundary value problem. Journal of Computational and Applied Mathematics, 72(2):335–343, August 13, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042796000040.

See comments [Anh99].

ElTarazi:1990:IQS

[ES90] M. N. El Tarazi and S. Sallam. Interpolation by quadratic spline with periodic derivative on uniform meshes. Journal of Computational and Applied Mathematics, 33(3):307–314, December 31, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042705800068.

Ervin:1992:CCP

[ES92] V. J. Ervin and E. P.

Stephan. Collocation with Chebyshev polynomials for a hypersingular integral equation on an interval. Journal of Computational and Applied Mathematics, 43(1–2):221–229, November 25, 1992. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042792902672.

Elschner:1996:DCM

[ES96] Johannes Elschner and Ernst P. Stephan. A discrete collocation method for Symm’s integral equation on curves with corners. Journal of Computational and Applied Mathematics, 75(1):131–146, November 12, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042796000702.

El-Seoud:1997:ACC

[ES97] M. S. Abou El-Seoud. On the asymptotic convergence of a collocation method for some boundary integral equations of the first kind. Journal of Computational and Applied Mathematics, 78(2):329–353, February 17, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042797001611.

El-Serafi:1993:MPE

[ESEHEK93] S. A. El-Serafi, F. Z. El-Halafawy, M. H. Eissa, and M. M. Kamel. An MSPT of a parametrically excited cantilever beam. Journal of Computational and Applied Mathematics, 47(2): 219–239, ???? 1993. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/037704279390005V.

Espinoza:1999:DAF

[Esp99] Pedro C. Espinoza. Discrete analogue of fucik spectrum of the Laplacian. Journal of Computational and Applied Mathematics, 103(1):93–97, March 15, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704279800243X.

Essen:1999:BCI

[ESS99] Matts Essén, Daniel Shea, and Charles Stanton. Best constant inequalities for conjugate functions. Journal of Computational and Applied Mathematics, 105(1– 2):257–264, May 1999. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042799000151.

Evans:1999:CSM

[EW99] G. A. Evans and J. R.

Webster. A comparison of some methods for the evaluation of highly oscillatory integrals. Journal of Computational and Applied Mathematics, 112(1–2):55– 69, November 30, 1999. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042799002137.

Ewing:1996:MIE

[Ewi96] Richard E. Ewing. Multidisciplinary interactions in energy and environmental modeling. Journal of Computational and Applied Mathematics, 74(1–2):193– 215, November 5, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042796000246.

Extremera:1991:STP

[Ext91] A. Extremera. Semiclassical treatment of the pathintegral approach to special effects in superconductivity. Journal of Computational and Applied Mathematics, 37(1–3):71–79, November 18, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279190107U.

Extremera:1995:DMP

[Ext95]  A. Extremera.  Damp-

ing mechanisms in the path-integral approach to special effects in superconductivity. Journal of Computational and Applied Mathematics, 58(1):17–26, March 20, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042793E0265N.

Exton:1996:NGF

[Ext96]  Harold Exton.  New

generating functions for Gegenbauer polynomials. Journal of Computational and Applied Mathematics, 67(1):191–193, February 20, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042795002375.

Exton:1997:RKD

[Ext97a] Harold Exton. On the reducibility of the Kampé de Fériet function. Journal of Computational and Applied Mathematics, 83 (1):119–121, September 30, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042797865971.

Exton:1997:SNS

[Ext97b]  Harold Exton.  Some

new summation formulae for the generalised hypergeometric function of higher order. Journal of Computational and Applied Mathematics, 79(2):183– 187, March 17, 1997. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042796001549.

Exton:1998:NHI

[Ext98a] H. Exton. New hypergeometric identities arising from Gauss’s second summation theorem. Journal of Computational and Applied Mathematics, 88(2):269– 274, March 2, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042797001891.

Exton:1998:NHT

[Ext98b] H. Exton. New hypergeometric transformations. Journal of Computational and Applied Mathematics, 92(2):135–137, June 26, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042797002355.

Exton:1998:SFI

[Ext98c] H. Exton. Summation formulae involving the Laguerre polynomial. Journal of Computational and Applied Mathematics, 100 (2):225–227, December 21, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042798001320.

Exton:1999:NTT

[Ext99] H. Exton. A new twoterm relation for the _{3}F_{2} hypergeometric function of unit argument. Journal of Computational and Applied Mathematics, 106(2):395– 397, June 30, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042799000771.

See comment [Mil06].

Ezzat:1994:PLF

[Ezz94] Magdy A. Ezzat. Power-law fluid flow of a hydromagnetic free jet. Journal of Computational and Applied Mathematics, 54(1):37–43, September 20, 1994. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042794903921.

Fahmy:1999:IEP

[FAD99] M. H. Fahmy, M. A. Abdou, and M. A. Darwish. Integral equations and potential-theoretic type integrals of orthogonal polynomials. Journal of Computational and Applied Mathematics, 106(2):245–254, June 30, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042799000643.

Fang:1995:WVS

[Fan95] Chen Fang. On Wimp’s vector sequence transformations. Journal of Computational and Applied Mathematics, 57(3):425– 447, February 28, 1995. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042793E02125.

Fasino:1995:SPH

[Fas95]  Dario Fasino.  Spectral

properties of Hankel matrices and numerical solutions of finite moment problems. Journal of Computational and Applied Mathematics, 65(1–3):145–155, December 29, 1995. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0377042795001085.

Faure:1990:UPR

[Fau90]  H. Faure.  Using per-

mutations to reduce discrepancy. Journal of Computational and Applied Mathematics, 31(1):97–103, July 24, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042790903406.

Faybusovich:1997:LSJ

[Fay97] Leonid Faybusovich. Linear systems in Jordan algebras and primal-dual interior-point algorithms. Journal of Computational and Applied Mathematics, 86(1):149–175, November 28, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042797001532.

Fazio:1990:NES

[Faz90] Riccardo Fazio. Numerical evaluations for shock and discontinuity velocity related to similarity solutions of quasilinear hyperbolic systems. Journal of Computational and Applied Mathematics, 30(3): 341–349, July 10, 1990. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042790902847.

Fazio:1992:NLE

[Faz92] Riccardo Fazio. Numerical length estimation for tubular flow reactors. Journal of Computational and Applied Mathematics, 41(3):313– 321, August 28, 1992. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279290138N.

Fazio:1994:NTM

[Faz94] Riccardo Fazio. Numerical transformation methods: a constructive approach. Journal of Computational and Applied Mathematics, 50(1–3):299–303, May 20, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042794903085.

Freeman:1990:DCM

[FB90] T. L. Freeman and R. W.

Brankin. A divide and conquer method for polynomial zeros. Journal of Computational and Applied Mathematics, 30(1):71–79, April 10, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279090006L.

Ford:1996:QBS

[FB96] Neville J. Ford and Christopher T. H. Baker. Qualitative behaviour and stability of solutions of discretised nonlinear Volterra integral equations of convolution type. Journal of Computational and Applied Mathematics, 66(1–2):213– 225, January 31, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0377042795001808.

Ferreira:1993:CPN

[FD93] J. A. Ferreira and Paula De Oliveira. Convergence properties of numerical discretizations and regridding methods. Journal of Computational and Applied Mathematics, 45(3): 321–330, May 3, 1993. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/037704279390049H.

Fdil:1999:ATC

[Fdi99] A. Fdil. On the acceleration of two classes of sequences by the θ̂-algorithm. Journal of Computational and Applied Mathematics, 106(1): 71–85, June 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042799000539.

Fearn:1991:EBV

[Fea91] David R. Fearn. Eigensolutions of boundary value problems using inverse iteration. Journal of Computational and Applied Mathematics, 34(2):201– 209, April 4, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279190042I.

Feng:1994:SNF

[Fen94] Y. Feng. The study of nonlinear flexings in a floating beam by variational methods. Journal of Computational and Applied Mathematics, 52(1–3):91–112, July 20, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042794903514.

Ferreira:1997:CNG

[Fer97] J. A. Ferreira. On the convergence on nonrectangular grids. Journal of Computational and Applied Mathematics, 85(2):333– 344, November 12, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042797001398.

Fausett:1997:IPM

[FFH97] Donald W. Fausett, Charles T. Fulton, and Hany Hashish. Improved parallel QR method for large least squares problems involving Kronecker products. Journal of Computational and Applied Mathematics, 78(1):63–78, February 3, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042796001094.

Fritzsche:1997:ANE

[FFK97] Bernd Fritzsche, Stefan Fuchs, and Bernd Kirstein. On Arov-normalized j_{pq}- elementary factors. Journal of Computational and Applied Mathematics, 77(1–2): 129–156, January 13, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042796001252.

Feistauer:1995:CFE

[FFLM95] Miloslav Feistauer, Jirí Felcman, and Mária Lukácová-Medvid’ová. Combined finite element-finite volume solution of compressible flow. Journal of Computational and Applied Mathematics, 63(1–3):179–199, November 20, 1995. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042795000518.

Foerster:1995:INM

[FFM95] Hildegard Foerster, Andreas Frommer, and Günter Mayer. Inexact Newton methods on a vector supercomputer. Journal of Computational and Applied Mathematics, 58(3):237– 253, April 20, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279400002I.

Feistauer:1992:FES

[FFRV92] Miloslav Feistauer, Jirí Felcman, Mirko Rokyta, and Zdenek Vlásek. Finiteelement solution of flow problems with trailing conditions. Journal of Computational and Applied Mathematics, 44(2):131–166, December 17, 1992. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279290008L.

Ford:1991:UCE

[FG91] J. A. Ford and R.-A. Ghandhari. On the use of curvature estimates in quasi-Newton methods. Journal of Computational and Applied Mathematics, 35(1–3):185– 196, June 26, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279190206Y.

Fischer:1992:HGU

[FG92] Bernd Fischer and Gene H. Golub. How to generate unknown orthogonal polynomials out of known orthogonal polynomials. Journal of Computational and Applied Mathematics, 43(1–2):99–115, November 25, 1992. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/037704279290261U.

Frontini:1993:RMD

[FG93] M. Frontini and L. Gotusso. A regularization method for discrete Fourier polynomials. Journal of Computational and Applied Mathematics, 49(1–3):71–78, December 31, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279390136Y.

Franco:1997:TTP

[FG97] J. M. Franco and I. Gomez.

Two three-parallel and three-processor SDIRK methods for stiff initial-value problems. Journal of Computational and Applied Mathematics, 87(1):119– 134, December 18, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042797001829.

Floater:1999:HMT

[FG99] Michael S. Floater and Craig Gotsman. How to morph tilings injectively. Journal of Computational and Applied Mathematics, 101(1–2):117– 129, January 15, 1999. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042798002027.

Flajolet:1995:RF

[FGKP95] Philippe Flajolet, Peter J.

Grabner, Peter Kirschenhofer, and Helmut Prodinger. On Ramanujan’s Q-function. Journal of Computational and Applied Mathematics, 58(1):103–116, March 20, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042793E0258N.

Franco:1997:SMS

[FGR97] J. M. Franco, I. Gómez, and L. Rández. SDIRK methods for stiff ODEs with oscillating solutions. Journal of Computational and Applied Mathematics, 81(2): 197–209, July 8, 1997. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042797000563.

Foupouagnigni:1998:LFE

[FHR98] M. Foupouagnigni, M. N. Hounkonnou, and A. Ronveaux. Laguerre–Freud equations for the recurrence coefficients of D_{ω} semi-classical orthogonal polynomials of class one. Journal of Computational and Applied Mathematics, 99(1–2):143–154, November 16, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042798001526.

Floater:1996:MSD

[FI96] Michael S. Floater and Armin Iske. Multistep scattered data interpolation using compactly supported radial basis functions. Journal of Computational and Applied Mathematics, 73(1–2): 65–78, October 5, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042796000350.

Field:1990:CRP

[Fie90]  David A. Field.  Con-

vergence rates for Padébased iterative solutions of equations. Journal of Computational and Applied Mathematics, 32(1–2):69– 75, November 26, 1990. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279090417X.

Fischer:1991:ADV

[Fis91] Herbert Fischer. Automatic differentiation of the vector that solves a parametric linear system. Journal of Computational and Applied Mathematics, 35(1–3):169– 184, June 26, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279190205X.

Fitzsimons:1991:CFE

[Fit91] Conor J. Fitzsimons. A class of finite-element methods for singularly perturbed second-order differential equations. Journal of Computational and Applied Mathematics, 34(3):269– 280, April 26, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042791900882.

Filippi:1995:GHM

[FK95] S. Filippi and T. Knecht. Generalized Hermite multistep methods of high order with a variable stepsize. Journal of Computational and Applied Mathematics, 59(2):173–180, May 19, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
037704279400032V.

Feng:1996:CIC

[FK96] Yu Yu Feng and Jernej Kozak. On G^{2} continuous interpolatory composite quadratic Bézier curves. Journal of Computational and Applied Mathematics, 72(1):141–159, July 30, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042795002685.

Fritzsche:1997:BCP

[FKM97] B. Fritzsche, B. Kirstein, and K. Müller. A block completion problem for matrix-valued inner functions. Journal of Computational and Applied Mathematics, 77(1–2):157–172, January 13, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042796001264.

Foupouagnigni:1998:FOD

[FKR98] M. Foupouagnigni, W. Koepf, and A. Ronveaux. Fourthorder difference equation for the associated classical discrete orthogonal polynomials. Journal of Computational and Applied Mathematics, 92(2):103– 108, June 26, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042798000508.

Fitzsimons:1992:SOS

[FLM92] Conor J. Fitzsimons, Fawang Liu, and John J. H. Miller. A second-order L-stable time discretisation of the semiconductor device equations. Journal of Computational and Applied Mathematics, 42(2):175–186, October 12, 1992. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042792900726.

Floris:1996:NDH

[Flo96] Paul G. A. Floris. A noncommutative discrete hypergroup associated with qdisk polynomials. Journal of Computational and Applied Mathematics, 68(1–2): 69–78, April 22, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0377042795002561.

Ford:1994:MSQ

[FM94] J. A. Ford and I. A. Moghrabi. Multi-step quasi-Newton methods for optimization. Journal of Computational and Applied Mathematics, 50(1–3):305– 323, May 20, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042794903093.

Ford:1996:UFV

[FM96] J. A. Ford and I. A. Moghrabi. Using functionvalues in multi-step quasi-Newton methods. Journal of Computational and Applied Mathematics, 66(1–2):201– 211, January 31, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0377042795001786.

Ford:1997:AMS

[FM97] J. A. Ford and I. A. Moghrabi. Alternating multi-step quasi-Newton methods for unconstrained optimization. Journal of Computational and Applied Mathematics, 82(1–2):105– 116, September 15, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042797000757.

Feng:1998:FDM

[FM98a] Bao-Feng Feng and Taketomo Mitsui. A finite difference method for the Korteweg–de Vries and the Kadomtsev–Petviashvili equations. Journal of Computational and Applied Mathematics, 90(1):95–116, April 6, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042798000065.

Fu:1998:CMR

[FM98b] Jyun-Horng Fu and David F.

Miller. Computational matrix representation modules for linear operators with explicit constructions for a class of lie operators. Journal of Computational and Applied Mathematics, 98(1): 1–26, October 5, 1998. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042798000880.

Fujino:1991:PHO

[FMT91] Seiji Fujino, Masatake Mori, and Toshiki Takeuchi. Performance of hyperplane ordering on vector computers. Journal of Computational and Applied Mathematics, 38(1–3):125–136, December 23, 1991. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/037704279190165G.

Franco:1998:MPP

[FN98] Daniel Franco and Juan J. Nieto. Maximum principles for periodic impulsive first order problems. Journal of Computational and Applied Mathematics, 88(1):149– 159, February 23, 1998. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042797002124.

Forster:1993:IUP

[För93] Klaus-Jürgen Förster. Inequalities for ultraspherical polynomials and application to quadrature. Journal of Computational and Applied Mathematics, 49(1–3):59– 70, December 31, 1993. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/037704279390135X.

Forster:1995:EVP

[För95] Klaus-Jürgen Förster. Estimates for the variance of positive quadrature formulas. Journal of Computational and Applied Mathematics, 57(1–2):135–146, February 20, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042793E0240M.

Foster:1997:GFE

[Fos97]  Leslie V. Foster.  The

growth factor and efficiency of Gaussian elimination with rook pivoting. Journal of Computational and Applied Mathematics, 86(1):177–194, November 28, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042797001544.

Foster:1998:GFE

[Fos98]  Leslie V. Foster.  The

growth factor and efficiency of Gaussian elimination with rook pivoting. Journal of Computational and Applied Mathematics, 98(1):177, October 5, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042798000934.

Franco:1990:HOS

[FP90] J. M. Franco and M. Palacios. High-order P -stable multistep methods. Journal of Computational and Applied Mathematics, 30(1):1– 10, April 10, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279090001G.

Fernandez:1997:EST

[FP97a] Bastien Fernandez and Maciej Pindor. On the eigenvalues of some transfer matrices. Journal of Computational and Applied Mathematics, 81(2):249– 255, July 8, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042797000617.

Franco:1997:HOA

[FP97b] J. M. Franco and J. F. Palacian. High order adaptive methods of Nyström– Cowell type. Journal of Computational and Applied Mathematics, 81(1):115– 134, June 1, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042797000307.

Funaro:1999:SAA

[FP99] Daniele Funaro and Giuseppe Pontrelli. Spline approximation of advection– diffusion problems using upwind type collocation nodes. Journal of Computational and Applied Mathematics, 110(1):141–153, October 15, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042799002071.

Fang:1992:MCE

[FPR92] S. C. Fang, E. L. Peterson, and J. R. Rajasekera. Minimum cross-entropy analysis with entropy-type constraints. Journal of Computational and Applied Mathematics, 39(2):165– 178, March 30, 1992. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279290127J.

Falcao:1999:CCS

[FPS99] M. I. Falcão, N. Papamichael, and N. S. Stylianopoulos. Curvilinear crosscuts of subdivision for a domain decomposition method in numerical conformal mapping. Journal of Computational and Applied Mathematics, 106(1): 177–196, June 1999. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042799000679.

Fodor:1995:MM

[FR95a] János Fodor and Marc Roubens. On meaningfulness of means. Journal of Computational and Applied Mathematics, 64(1–2):103– 115, November 30, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042795000100.

Frolov:1995:MNR

[FR95b] Alexander Frolov and Stanislav Rízek. Model of neurocontrol of redundant systems. Journal of Computational and Applied Mathematics, 63(1–3):465–473, November 20, 1995. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0377042795000712.

Frommer:1999:UAP

[FR99] Andreas Frommer and Rosemary A. Renaut. A unified approach to parallel space decomposition methods. Journal of Computational and Applied Mathematics, 110(1):205–223, October 15, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042799002356.

Frauendorfer:1994:ASS

[Fra94] Karl Frauendorfer. The approximation of separable stochastic programs. Journal of Computational and Applied Mathematics, 56(1–2):23–44, December 20, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042794903778.

Franco:1995:EHM

[Fra95a] J. M. Franco. An explicit hybrid method of Numerov type for secondorder periodic initial-value problems. Journal of Computational and Applied Mathematics, 59(1):79–90, April 28, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279400011O.

Frankel:1995:NIF

[Fra95b] J. I. Frankel. A note on the integral formulation of Kumar and Sloan. Journal of Computational and Applied Mathematics, 61(3):263– 274, August 21, 1995. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042794000713.

Frankena:1995:SCS

[Fra95c] J. F. Frankena. Störmer– Cowell: straight, summed and split. An overview. Journal of Computational and Applied Mathematics, 62(2):129–154, September 20, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042794001020.

Frauendiener:1999:CID

[Fra99] Jörg Frauendiener. Calculating initial data for the conformal Einstein equations by pseudo-spectral methods. Journal of Computational and Applied Mathematics, 109(1–2):475– 491, September 30, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042799001685.

Freund:1992:QKP

[Fre92] Roland W. Freund. Quasikernel polynomials and their use in non-Hermitian matrix iterations. Journal of Computational and Applied Mathematics, 43(1–2):135– 158, November 25, 1992. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.

com/science/article/pii/
037704279290263W.

Foupouagnigni:1999:FOD

[FRK99] M. Foupouagnigni, A. Ronveaux, and W. Koepf. Fourth order q-difference equation for the first associated of the q-classical orthogonal polynomials. Journal of Computational and Applied Mathematics, 101(1–2):231–236, January 15, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042798002258.

Frommer:1991:GND

[Fro91] Andreas Frommer. Generalized nonlinear diagonal dominance and applications to asynchronous iterative methods. Journal of Computational and Applied Mathematics, 38(1–3):105– 124, December 23, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
037704279190164F.

Feldman:1994:LTR

[FRR94] P. Feldman, S. T. Rachev, and L. Rüschendorf. Limit theorems for recursive algorithms. Journal of Computational and Applied Mathematics, 56(1–2):169–182, December 20, 1994. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0377042794903867.

Filippi:1990:THO

[FS90] S. Filippi and F. Schöne. Two high-order multistep methods with offstep points applied with a variable stepsize. Journal of Computational and Applied Mathematics, 30(2):155–164, May 28, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        037704279090024T.

Ford:1991:CHA

[FS91] John A. Ford and Adel F. Saadallah. On the conditioning of the Hessian approximation in quasi-Newton methods. Journal of Computational and Applied Mathematics, 35(1–3):197– 206, June 26, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279190207Z.

Frommer:1995:APM

[FS95] Andreas Frommer and Hartmut Schwandt. Asynchronous parallel methods for enclosing solutions of nonlinear equations. Journal of Computational and Applied Mathematics, 60(1–2): 47–62, June 20, 1995. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279400083D.

Feinsilver:1996:OCA

[FS96a] P. Feinsilver and R. Schott.

Operator calculus approach to orthogonal polynomial expansions. Journal of Computational and Applied Mathematics, 66(1–2):185– 199, January 31, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042795001611.

Franke:1996:LMI

[FS96b] R. Franke and K. Salkauskas. Localization of multivariate interpolation and smoothing methods. Journal of Computational and Applied Mathematics, 73(1–2):79–94, October 5, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042796000362.

Fokkema:1996:GCG

[FSV96] Diederik R. Fokkema, Gerard L. G. Sleijpen, and Henk A. Van der Vorst. Generalized conjugate gradient squared. Journal of Computational and Applied Mathematics, 71(1): 125–146, July 10, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0377042795002278.

Freedman:1994:MGP

[FSW94] H. I. Freedman, Joseph W.- H So, and Jianhong Wu. A model for the growth of a population exhibiting stage structure: cannibalism and cooperation. Journal of Computational and Applied Mathematics, 52(1–3):177– 198, July 20, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042794903565.

Fonda:1994:PON

[FSZ94] Alessandro Fonda, Zdenek Schneider, and Fabio Zanolin. Periodic oscillations for a nonlinear suspension bridge model. Journal of Computational and Applied Mathematics, 52(1–3):113– 140, July 20, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042794903522.

Fang:1996:ECP

[FT96] Shu-Cherng Fang and H.- S. Jacob Tsao. An efficient computational procedure for solving entropy optimization problems with infinitely many linear constraints. Journal of Computational and Applied Mathematics, 72(1): 127–139, July 30, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042795002669.

Furusho:1998:PES

[FT98] Yasuhiro Furusho and Kusano Takaši. Positive entire solutions to nonlinear biharmonic equations in the plane. Journal of Computational and Applied Mathematics, 88(1):161– 173, February 23, 1998. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042797002148.

Fukushima:1990:CGA

[Fuk90]  Masao Fukushima.  A

conjugate gradient algorithm for sparse linear inequalities. Journal of Computational and Applied Mathematics, 30(3):329– 339, July 10, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042790902836.

Funaro:1992:ALC

[Fun92] Daniele Funaro. Approximation by the Legendre collocation method of a model problem in electrophysiology. Journal of Computational and Applied Mathematics, 43(1–2):261–271, November 25, 1992. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0377042792902708.

Fushimi:1990:RNG

[Fus90] Masanori Fushimi. Random number generation with the recursion X_{t} = X_{t−3p} ⊕ X_{t−3q}. Journal of Computational and Applied Mathematics, 31(1): 105–118, July 24, 1990. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/037704279090341V.

Fack:1991:AGL

[FV91] V. Fack and J. Van der Jeugt. Algorithms for generating labelled graphs with given degree. Journal of Computational and Applied Mathematics, 37(1–3):187– 194, November 18, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042791901173.

Floreanini:1995:QAA

[FV95] Roberto Floreanini and Luc Vinet. Quantum algebra approach to q Gegenbauer polynomials. Journal of Computational and Applied Mathematics, 57(1–2):123– 133, February 20, 1995. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042793E0239I.

Floreanini:1996:GBF

[FV96] Roberto Floreanini and Luc Vinet. q-gamma and qbeta functions in quantum algebra representation theory. Journal of Computational and Applied Mathematics, 68(1–2):57– 68, April 22, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042795002537.

Ford:1999:UBL

[FW99] Neville J. Ford and Volker Wulf. The use of boundary locus plots in the identification of bifurcation points in numerical approximation of delay differential equations. Journal of Computational and Applied Mathematics, 111(1–2):153–162, November 15, 1999. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042799001399.

Floryan:1993:SCM

[FZ93] J. M. Floryan and Charles Zemach. Schwarz–Christoffel methods for conformal mapping of regions with a periodic boundary. Journal of Computational and Applied Mathematics, 46(1–2): 77–102, June 14, 1993. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/037704279390288M.

Foata:1996:GMI

[FZ96] Dominique Foata and Doron Zeilberger. Graphical major indices. Journal of Computational and Applied Mathematics, 68(1–2):79– 101, April 22, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042795002545.

Gabr:1992:NTB

[Gab92]  M. M. Gabr.  Non-

linearity tests for bilinear systems. Journal of Computational and Applied Mathematics, 40(3):313– 322, July 24, 1992. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042792901873.

Ganesh:1993:BMN

[Gan93] M. Ganesh. A BIE method for a nonlinear BVP. Journal of Computational and Applied Mathematics, 45(3): 299–308, May 3, 1993. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279390047F.

Gautschi:1992:MCE

[Gau92]  Walter Gautschi.  On

mean convergence of extended Lagrange interpolation. Journal of Computational and Applied Mathematics, 43(1–2):19– 35, November 25, 1992. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279290257X.

Germain-Bonne:1990:CAN

[GB90a] B. Germain-Bonne. Convergence acceleration of number-machine sequences. Journal of Computational and Applied Mathematics, 32(1–2):83–88, November 26, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279090419Z.

Gupta:1990:SWT

[GB90b] R. S. Gupta and N. C. Banik. Solution of a weakly two-dimensional melting problem by an approximate method. Journal of Computational and Applied Mathematics, 31(3):351–356, October 1, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279090035X.

Georgiou:1997:CTM

[GBP97] Georgios Georgiou, Andreas Boudouvis, and Andreas Poullikkas. Comparison of two methods for the computation of singular solutions in elliptic problems. Journal of Computational and Applied Mathematics, 79(2):277– 287, March 17, 1997. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042796001732.

Guo:1997:IPS

[GC97] Benqi Guo and Weiming Cao. An iterative and parallel solver based on domain decomposition for the h- p version of the finite element method. Journal of Computational and Applied Mathematics, 83(1):71–85, September 30, 1997. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042797000630.

Gear:1991:WMS

[Gea91] C. W. Gear. Waveform methods for space and time parallelism. Journal of Computational and Applied Mathematics, 38(1–3):137– 147, December 23, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
037704279190166H.

Gelbrich:1995:STC

[Gel95] Matthias Gelbrich. Simultaneous time and chance discretization for stochastic differential equations. Journal of Computational and Applied Mathematics, 58(3): 255–289, April 20, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
037704279400003J.

Gemignani:1996:FIM

[Gem96] Luca Gemignani. A fast iterative method for determining the stability of a polynomial. Journal of Computational and Applied Mathematics, 76(1–2):1–11, December 17, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042796000489.

Gentle:1990:CIR

[Gen90] James E. Gentle. Computer implementation of random number generators. Journal of Computational and Applied Mathematics, 31(1): 119–125, July 24, 1990. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/037704279090342W.

Galeone:1997:CAT

[GG97a] Luciano Galeone and Roberto Garrappa. Convergence analysis of time-point relaxation iterates for linear systems of differential equations. Journal of Computational and Applied Mathematics, 80(2): 183–195, May 5, 1997. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042797000046.

Gates:1997:NA

[GG97b] K. Gates and W. B. Gragg. Notes on T QR algorithms. Journal of Computational and Applied Mathematics, 86(1):195–203, November 28, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042797001556.

Gonzalez:1999:AEP

[GGF99] Ana B. González, Juan Getino, and JoséM. Farto. An algorithm for an eigenvalues problem in the Earth rotation theory. Journal of Computational and Applied Mathematics, 101(1–2):243– 254, January 15, 1999. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042798002301.

Grunbaum:1996:VTB

[GH96] F. Alberto Grünbaum and Luc Haine. The qversion of a theorem of Bochner. Journal of Computational and Applied Mathematics, 68(1–2):103–114, April 22, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042795002626.

Gunther:1997:RMA

[GH97a] M. Günther and M. Hoschek.

ROW methods adapted to electric circuit simulation packages. Journal of Computational and Applied Mathematics, 82(1–2):159– 170, September 15, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042797000435.

Gutierrez:1997:TOI

[GH97b] JoséM. Gutiérrez and Miguel A. Hernández. Third-order iterative methods for operators with bounded second derivative. Journal of Computational and Applied Mathematics, 82(1–2):171– 183, September 15, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042797000769.

Gehring:1999:HGD

[GH99a] F. W. Gehring and K. Hag. Hyperbolic geometry and disks. Journal of Computational and Applied Mathematics, 105(1–2):275–284, May 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042799000163.

Gustafson:1999:GQR

[GH99b] Philip E. Gustafson and Brian A. Hagler. Gaussian quadrature rules and numerical examples for strong extensions of mass distribution functions. Journal of Computational and Applied Mathematics, 105(1–2):317– 326, May 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042799000230.

Grunbaum:1999:SFG

[GHH99] F. Alberto Grünbaum, Luc Haine, and Emil Horozov. Some functions that generalize the Krall–Laguerre polynomials. Journal of Computational and Applied Mathematics, 106(2):271– 297, June 30, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042799000692.

Garcia-Huidobro:1994:PFS

[GHMZ94]  Marta  García-Huidobro,

Raúl Manásevich, and Fabio Zanolin. On a pseudo fucik spectrum for strongly nonlinear secondorder ODEs and an existence result. Journal of Computational and Applied Mathematics, 52(1–3):219– 239, July 20, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042794903581.

Gill:1990:CDL

[Gil90] John Gill. Complex dynamics of the limit periodic system F_{n}(z) = F_{n−1}(f_{n}(z)), f_{n} → f.

Journal of Computational and Applied Mathematics, 32(1–2):89–96, November 26, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042790904205.

Gill:1996:NBD

[Gil96] John Gill. A note on bounds for the derivatives of continued fractions. Journal of Computational and Applied Mathematics, 72(2):301– 307, August 13, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0377042795002804.

Gill:1999:NCI

[Gil99a]  John Gill.  A natu-

ral continuous interpolating structure for continued fractions. Journal of Computational and Applied Mathematics, 105(1–2):299– 309, May 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042799000370.

Gill:1999:NEE

[Gil99b] John Gill. A note on extending Euler’s connection between continued fractions and power series. Journal of Computational and Applied Mathematics, 106(2): 299–305, June 30, 1999. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042799000710.

Gupta:1996:CRB

[GIM96] D. P. Gupta, M. E. H. Ismail, and D. R. Masson. Contiguous relations, basic hypergeometric functions, and orthogonal polynomials. III. Associated continuous dual q-Hahn polynomials. Journal of Computational and Applied Mathematics, 68(1–2):115–149, April 22, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042795002642.

Gawecki:1995:CAS

[GJ95] A. Gawecki and B. Janińska.

Computer analysis of slackened systems under variable loads. Journal of Computational and Applied Mathematics, 63(1–3):325–332, November 20, 1995. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042795000925.

Gjesdal:1997:SAM

[Gje97]  Thor Gjesdal.  Smooth-

ing analysis of multicolour pattern schemes. Journal of Computational and Applied Mathematics, 85(2):345–350, November 12, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042797001404.

Gawecki:1995:EPU

[GK95] A. Gawecki and M. S. Kuczma. Elastic-plastic unilateral contact problems for slackened systems. Journal of Computational and Applied Mathematics, 63(1–3):313–323, November 20, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042795000879.

Genz:1996:FSI

[GK96a] Alan Genz and B. D. Keister. Fully symmetric interpolatory rules for multiple integrals over infinite regions with Gaussian weight. Journal of Computational and Applied Mathematics, 71(2): 299–309, July 27, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0377042795002324.

Graef:1996:ABS

[GK96b] John R. Graef and János Karsai. On the asymptotic behavior of solutions of impulsively damped nonlinear oscillator equations. Journal of Computational and Applied Mathematics, 71(1): 147–162, July 10, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042795002154.

Gustafson:1996:HTN

[GK96c] Robert A. Gustafson and C. Krattenthaler. Heine transformations for a new kind of basic hypergeometric series in U ( n ). Journal of Computational and Applied Mathematics, 68(1–2):151– 158, April 22, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279500260X.

Giladi:1997:ISE

[GK97] Eldar Giladi and Joseph B.

Keller. Iterative solution of elliptic problems by approximate factorization. Journal of Computational and Applied Mathematics, 85(2):287–313, November 12, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042797001325.

Gabutti:1991:SER

[GKH91] B. Gabutti, S. L. Kalla, and J. H. Hubbell. Some expansions related to the Hubbell rectangular-source integral. Journal of Computational and Applied Mathematics, 37(1–3):273–285, November 18, 1991. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0377042791901243.

Gautschi:1990:RTA

[GL90a] Walter Gautschi and Shikang Li. The remainder term for analytic functions of Gauss–Radau and Gauss– Lobatto quadrature rules with multiple end points. Journal of Computational and Applied Mathematics, 33(3):315–329, December 31, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704270580007X.

Gray:1990:TCB

[GL90b] L. J. Gray and E. Lutz. On the treatment of corners in the boundary element method. Journal of Computational and Applied Mathematics, 32(3):369–386, December 10, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279090043Y.

Gautschi:1991:GRG

[GL91] Walter Gautschi and Shikang Li. Gauss–Radau and Gauss–Lobatto quadratures with double end points. Journal of Computational and Applied Mathematics, 34(3):343–360, April 26, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
037704279190094Z.

Guoqiang:1994:AET

[GL94] Han Guoqiang and Zhang Liqing. Asymptotic expansion for the trapezoidal Nyström method of linear Volterra–Fredholm equations. Journal of Computational and Applied Mathematics, 51(3):339– 348, June 30, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279200013Y.

Guenther:1998:HCR

[GL98] R. B. Guenther and J. W.

Lee. Heat conduction with radiating boundary conditions. Journal of Computational and Applied Mathematics, 88(1):119– 124, February 23, 1998. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042797002082.

Gilewicz:1999:ZPR

[GL99] Jacek Gilewicz and Elie Leopold. Zeros of polynomials and recurrence relations with periodic coefficients. Journal of Computational and Applied Mathematics, 107(2):241–255, July 31, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042799000928.

Giordano:1998:UTG

[GLP98] C. Giordano, A. Laforgia, and J. Pecarić. Unified treatment of Gautschi– Kershaw type inequalities for the gamma function. Journal of Computational and Applied Mathematics, 99(1–2):167–175, November 16, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704279800154X.

Gonska:1994:ATI

[GlZ94] Heinz H. Gonska and Xin long Zhou. Approximation theorems for the iterated Boolean sums of Bernstein operators. Journal of Computational and Applied Mathematics, 53(1):21–31, July 29, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279200133T.

Gasca:1990:SBH

[GM90a] M. Gasca and J. J. Martínez.

On the solvability of bivariate Hermite–Birkhoff interpolation problems. Journal of Computational and Applied Mathematics, 32(1–2): 77–82, November 26, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
037704279090418Y.

Graves-Morris:1990:SIE

[GM90b] P. R. Graves-Morris. Solution of integral equations using generalised inverse, function-valued Padé approximants, I. Journal of Computational and Applied Mathematics, 32(1–2):117–124, November 26, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279090423W.

Gilewicz:1993:ISI

[GM93] J. Gilewicz and A. P. Magnus. Inverse Stieltjes iterates and errors of Padé approximants in the whole complex plane. Journal of Computational and Applied Mathematics, 49(1–3):79–84, December 31, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279390137Z.

Gupta:1995:CRC

[GM95] Dharma P. Gupta and David R. Masson. Contiguous relations, continued fractions and orthogonality: an _{8}Ø_{7} model. Journal of Computational and Applied Mathematics, 65(1–3):157–164, December 29, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042795001158.

Gautschi:1997:OCG

[GM97] Walter Gautschi and Gradimir V. Milovanovic. S-orthogonality and construction of Gauss– Turán-type quadrature formulae. Journal of Computational and Applied Mathematics, 86(1):205–218, November 28, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042797001568.

Genz:1999:SAH

[GM99] Alan Genz and John Monahan. A stochastic algorithm for high-dimensional integrals over unbounded regions with Gaussian weight. Journal of Computational and Applied Mathematics, 112(1–2):71–81, November 30, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042799002149.

Graves-Morris:1996:CTA

[GMBW96] P. R. Graves-Morris, George A.

Baker, Jr., and C. F. Woodcock. Cayley’s theorem and its application in the theory of vector Padé approximants. Journal of Computational and Applied Mathematics, 66(1–2):255–265, January 31, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279500176X.

Garcia-Melian:1999:EAL

[GMdLSG99] J. García-Melián, J. Sabina de Lis, and M. Sanabria-García. Eigenvalue analysis for the p-Laplacian under convective perturbation. Journal of Computational and Applied Mathematics, 110(1):73–91, October 15, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042799001971.

Graves-Morris:1994:MVP

[GMR94] P. R. Graves-Morris and D. E. Roberts. From matrix to vector Padé approximants. Journal of Computational and Applied Mathematics, 51(2):205– 236, June 20, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279200008W.

Graves-Morris:1997:PPV

[GMR97] P. R. Graves-Morris and D. E. Roberts. Problems and progress in vector Padé approximation. Journal of Computational and Applied Mathematics, 77(1–2):173– 200, January 13, 1997. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042796001276.

Graves-Morris:1991:ERC

[GMS91] P. R. Graves-Morris and E. B. Saff. An extension of a row convergence theorem for vector Padé approximants. Journal of Computational and Applied Mathematics, 34(3):315– 324, April 26, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279190092X.

Gorlich:1994:IFA

[GMS94] E. Görlich, C. Markett, and O. Stüpp. Integral formulas associated with products of Bessel functions: a new partial differential equation approach. Journal of Computational and Applied Mathematics, 51(2): 135–157, June 20, 1994. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279200011W.

Garcia:1995:DSD

[GMS95] A. G. García, F. Marcellán, and L. Salto. A distributional study of discrete classical orthogonal polynomials. Journal of Computational and Applied Mathematics, 57(1–2):147–162, February 20, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042793E0241D.

Griebel:1996:PPD

[GN96] Michael Griebel and Tilman Neunhoeffer. Parallel pointand domain-oriented multilevel methods for elliptic PDE’s on workstation networks. Journal of Computational and Applied Mathematics, 66(1–2):267–278, January 31, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042795001956.

Goovaerts:1991:PIE

[Goo91] M. J. Goovaerts. Pathintegral evaluation for the three-dimensional potential γδ(r−a). Journal of Computational and Applied Mathematics, 37(1–3):113–124, November 18, 1991. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/037704279190111V.

Gorsch:1999:DMC

[Gör99] Daniel Görsch. Discrete model collision operators of Boltzmann type. Journal of Computational and Applied Mathematics, 104(2): 145–162, May 17, 1999. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042799000503.

Govaerts:1994:SBS

[Gov94] W. Govaerts. Solution of bordered singular systems in numerical continuation and bifurcation. Journal of Computational and Applied Mathematics, 50(1–3):339– 347, May 20, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042794903115.

Gladwell:1993:PSA

[GP93] Ian Gladwell and Marcin Paprzycki. Parallel solution of almost block diagonal systems on the CRAY Y-MP using level 3 BLAS. Journal of Computational and Applied Mathematics, 45(1–2): 181–189, April 8, 1993. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279390273E.

Gilewicz:1997:PAN

[GP97] Jacek Gilewicz and Maciej Pindor. Padé approximants and noise: a case of geometric series. Journal of Computational and Applied Mathematics, 87(2):199–214, December 23, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042797001854.

Gilewicz:1998:PTA

[GP98] Jacek Gilewicz and Maciej Pindor. Padé-type approximants and errors of Padé approximants. Journal of Computational and Applied Mathematics, 99(1–2):155–165, November 16, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042798001538.

Gilewicz:1999:PAN

[GP99] Jacek Gilewicz and Maciej Pindor. Padé approximants and noise: rational functions. Journal of Computational and Applied Mathematics, 105(1–2):285– 297, May 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042799000412.

Gonzalez-Pinto:1991:NSA

[GPCGV91] S. González-Pinto, L. Casasús, and P. González-Vera. A numerical scheme to approximate the solution of a singularly perturbed nonlinear differential equation. Journal of Computational and Applied Mathematics, 35(1–3):217–225, June 26, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042791902093.

Gonzalez-Pinto:1996:IEI

[GPMR96] S. González-Pinto, J. I. Montijano, and L. Rández. Improving the efficiency of the iterative schemes for implicit Runge–Kutta methods. Journal of Computational and Applied Mathematics, 66(1–2):227–238, January 31, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042795001840.

Guadalupe:1993:EWB

[GPRV93] JoséJ. Guadalupe, Mario Pérez, Francisco J. Ruiz, and Juan L. Varona. Endpoint weak boundedness of some polynomial expansions. Journal of Computational and Applied Mathematics, 49(1–3):93–102, December 31, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042793901393.

Galiev:1992:CSS

[GPS92] V. I. Galiev, A. F. Polupanov, and I. E. Shparlinski. On the construction of solutions of systems of linear ordinary differential equations in the neighbourhood of a regular singularity. Journal of Computational and Applied Mathematics, 39(2):151– 163, March 30, 1992. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/037704279290126I.

Gao:1992:CLL

[GR92a] Zhicheng Gao and L. Bruce Richmond. Central and local limit theorems applied to asymptotic enumeration IV: multivariate generating functions. Journal of Computational and Applied Mathematics, 41(1–2):177– 186, August 20, 1992. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/037704279290247U.

Giordano:1992:FMC

[GR92b] Carla Giordano and Lucia G. Rodonò. Further monotonicity and convexity properties of the zeros of cylinder functions. Journal of Computational and Applied Mathematics, 42(2):245–251, October 12, 1992. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279290078C.

Grant:1992:DZL

[GR92c] J. A. Grant and A. A. Rahman. Determination of the zeros of a linear combination of generalised polynomials. Journal of Computational and Applied Mathematics, 42(2):269–278, October 12, 1992. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279290080H.

Gross:1992:EPS

[GR92d] Aaron Gross and James B.

Robertson. Ergodic properties of stationary Poisson sequences. Journal of Computational and Applied Mathematics, 40(2): 163–175, June 24, 1992. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042792901046.

Genton:1995:CVF

[GR95] Marc G. Genton and Peter J. Rousseeuw. The change-of-variance function of m-estimators of scale under general contamination. Journal of Computational and Applied Mathematics, 64(1–2):69–80, November 30, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042795000070.

Gerstberger:1997:FNN

[GR97] R. Gerstberger and P. Rentrop. Feedforward neural nets as discretization schemes for ODEs and DAEs. Journal of Computational and Applied Mathematics, 82(1–2):117–128, September 15, 1997. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S037704279700085X.

Gragg:1993:PDT

[Gra93] William B. Gragg. Positive definite Toeplitz matrices, the Arnoldi process for isometric operators, and Gaussian quadrature on the unit circle. Journal of Computational and Applied Mathematics, 46(1–2):183–198, June 14, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279390294L.

Graca:1996:SRA

[Gra96] Mário M. Graça. The sign regularity of the auxiliary family g_{i}(x) = x^{α_{i}}(− ln x)^{β_{i}} in convergence acceleration processes using the E-algorithm. Journal of Computational and Applied Mathematics, 67(2):237– 253, March 29, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/037704279500016X.

Greenhow:1993:CVM

[Gre93] Martin Greenhow. A complex variable method for the floating-body boundaryvalue problem. Journal of Computational and Applied Mathematics, 46(1–2):115– 128, June 14, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279390290R.

Grinshpan:1999:RGN

[Gri99] Arcadii Z. Grinshpan. Remarks on the Grunsky norm and p-th root transformation. Journal of Computational and Applied Mathematics, 105(1–2):311– 315, May 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042799000175.

Groetsch:1991:SNM

[Gro91a] C. W. Groetsch. A simple numerical model for nonlinear warming of a slab. Journal of Computational and Applied Mathematics, 38(1–3):149–156, December 23, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279190167I.

Grosjean:1991:DIM

[Gro91b] C. C. Grosjean. Deducing infinitely many mutually related sequences of orthogonal polynomials from solutions of first-order initialvalue problems: an illustrative example. Journal of Computational and Applied Mathematics, 35(1–3):227– 250, June 26, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279190210B.

Grosjean:1991:FPR

[Gro91c] C. C. Grosjean. On a formula proposed by S. Ramanujan. Journal of Computational and Applied Mathematics, 37(1–3):43– 56, November 18, 1991. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/037704279190105S.

Grunbaum:1992:TDS

[Grü92] F. Alberto Grünbaum. Three-dimensional scattering and the Heisenberg inequality. Journal of Computational and Applied Mathematics, 42(1):37–47, September 8, 1992. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279290161P.

Grunbaum:1998:VTH

[Grü98] F. Alberto Grünbaum. Variations on a theme of Heine and Stieltjes: an electrostatic interpretation of the zeros of certain polynomials. Journal of Computational and Applied Mathematics, 99(1–2):189–194, November 16, 1998. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042798001563.

Godoy:1997:MRR

[GRZA97] E. Godoy, A. Ronveaux, A. Zarzo, and I. Area. Minimal recurrence relations for connection coefficients between classical orthogonal polynomials: Continuous case. Journal of Computational and Applied Mathematics, 84(2):257–275, October 28, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042797001374.

Godoy:1998:CPP

[GRZA98a] E. Godoy, A. Ronveaux, A. Zarzo, and I. Area. Connection problems for polynomial solutions of nonhomogeneous differential and difference equations. Journal of Computational and Applied Mathematics, 99(1–2):177–187, November 16, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042798001551.

Godoy:1998:LRB

[GRZA98b] E. Godoy, A. Ronveaux, A. Zarzo, and I. Area. On the limit relations between classical continuous and discrete orthogonal polynomials. Journal of Computational and Applied Mathematics, 91(1):97–105, April 24, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042798000260.

Gabutti:1991:NIM

[GS91a] B. Gabutti and L. Sacripante. Numerical inversion of the Mellin transform by accelerated series of Laguerre polynomials. Journal of Computational and Applied Mathematics, 34(2): 191–200, April 4, 1991. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/037704279190041H.

Grosjean:1991:STR

[GS91b] C. C. Grosjean and H. M. Srivastava. Some transformation and reduction formulas for hypergeometric series in several variables. Journal of Computational and Applied Mathematics, 37(1–3):287–299, November 18, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042791901254.

Goh:1992:P

[GS92a] William M. Y. Goh and Eric Schmutz. Preface. Journal of Computational and Applied Mathematics, 41(1–2): 1–2, August 20, 1992. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279290232M.

Graef:1992:ONO

[GS92b] John R. Graef and Paul W.

Spikes. On the oscillation of an nth-order nonlinear neutral delay differential equation. Journal of Computational and Applied Mathematics, 41(1–2):35–40, August 20, 1992. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279290235P.

Griebel:1997:MPB

[GS97] Michael Griebel and Gerhard Starke. Multilevel preconditioning based on discrete symmetrization for convection–diffusion equations. Journal of Computational and Applied Mathematics, 83(2):165–183, October 7, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042797000952.

Griener:1999:ITV

[GS99] Bernhard Griener and Hans Joachim Schmid. An interactive tool to visualize common zeros of two-dimensional polynomials. Journal of Computational and Applied Mathematics, 112(1–2):83– 94, November 30, 1999. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042799002150.

Gottlieb:1992:GPR

[GSSV92] David Gottlieb, Chi-Wang Shu, Alex Solomonoff, and Hervé Vandeven. On the Gibbs phenomenon I: recovering exponential accuracy from the Fourier partial sum of a nonperiodic analytic function. Journal of Computational and Applied Mathematics, 43(1–2):81– 98, November 25, 1992. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042792902605.

Guillaume:1997:NMP

[Gui97] Philippe Guillaume. Nested multivariate Padé approximants. Journal of Computational and Applied Mathematics, 82(1–2):149–158, September 15, 1997. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042797000812.

Guoqiang:1995:AEE

[Guo95a] Han Guoqiang. Asymptotic error expansion for the Nyström method for a nonlinear Volterra–Fredholm integral equation. Journal of Computational and Applied Mathematics, 59(1):49–59, April 28, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279400021R.

Guoqiang:1995:EDC

[Guo95b] Han Guoqiang. Extrapolation of a discrete collocationtype method of Hammerstein equations. Journal of Computational and Applied Mathematics, 61(1):73–86, July 20, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042794000497.

Gutierrez:1997:NSC

[Gut97] JoséM Gutiérrez. A new semilocal convergence theorem for Newton’s method. Journal of Computational and Applied Mathematics, 79(1):131–145, March 3, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042797816111.

Grapsa:1996:DRM

[GV96] T. N. Grapsa and M. N. Vrahatis. A dimension-reducing method for unconstrained optimization. Journal of Computational and Applied Mathematics, 66(1–2):239– 253, January 31, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042795001743.

Gonzalez-Vera:1993:NCG

[GVGPSL93] P. González-Vera, S. González-Pinto, and J. C. Santos-León. A note on certain generalizations of the midpoint rule. Journal of Computational and Applied Mathematics, 49(1–3):85–91, December 31, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042793901382.

Gonzalez-Vera:1990:CTP

[GVN90a] Pablo González-Vera and Olav Njåstad. Convergence of two-point Padé approximants to series of Stieltjes. Journal of Computational and Applied Mathematics, 32(1–2):97–105, November 26, 1990. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279090421U.

Gonzalez-Vera:1990:SFM

[GVN90b] Pablo González-Vera and Olav Njåstad. Szegő functions and multipoint Padé approximation. Journal of Computational and Applied Mathematics, 32(1–2):107– 116, November 26, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
037704279090422V.

Gonzalez-Vera:1994:OTP

[GVO94] Pablo González-Vera and Ramón Orive. Optimization of two-point Padé-type approximants. Journal of Computational and Applied Mathematics, 50(1–3):325– 337, May 20, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042794903107.

Galperin:1991:NMU

[GW91] A. Galperin and Z. Waksman. Newton’s method under a weak smoothness assumption. Journal of Computational and Applied Mathematics, 35(1–3):207– 215, June 26, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042791902082.

Ge:1996:CHA

[GWC96] Y. Ge, L. T. Watson, and E. G. Collins, Jr. A comparison of homotopies for alternative formulations of the L^{2} optimal model order reduction problem. Journal of Computational and Applied Mathematics, 69(2): 215–241, May 23, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0377042795000453.

Goovaerts:1995:P

[GWP95] M. Goovaerts, L. Wuytack, and R. Piessens. Preface. Journal of Computational and Applied Mathematics, 64(1–2):1, November 30, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279590039X.

Gani:1995:CSM

[GY95] J. Gani and S. Yakowitz.

Computational and stochastic methods for interacting groups in the AIDS epidemic. Journal of Computational and Applied Mathematics, 59(2):207– 220, May 19, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279400034X.

Hack:1997:TUD

[Hac97] Michael Hack. Transformation of uniformly distributed particle ensembles. Journal of Computational and Applied Mathematics, 83(1):87–113, September 30, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704279700068X.

Hagler:1999:LHG

[Hag99] Brian A. Hagler. Laurent– Hermite–Gauss quadrature. Journal of Computational and Applied Mathematics, 104(2):163–171, May 17, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042799000540.

Haiyan:1990:NAP

[Hai90] Zhu Haiyan. A note about power-boundedness of interval matrices. Journal of Computational and Applied Mathematics, 31(2):295– 298, August 31, 1990. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/037704279090171U.

Haiyan:1992:FMC

[Hai92] Zhu Haiyan. Fast methods for computing the values of the derivatives of a rational function. Journal of Computational and Applied Mathematics, 39(3): 387–391, May 8, 1992. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/037704279290213H.

Hajmirzaahmad:1995:LPE

[Haj95] Mojdeh Hajmirzaahmad. Laguerre polynomial expansions. Journal of Computational and Applied Mathematics, 59(1):25–37, April 28, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042794000202.

Hales:1992:SPP

[Hal92] Thomas C. Hales. The sphere packing problem.

Journal of Computational and Applied Mathematics, 44(1):41–76, December 9, 1992. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279290052Y.

Hamboker:1994:CRS

[Ham94] Ralf Hamböker. Convergence rates for stopped random sums. Journal of Computational and Applied Mathematics, 54(2):159– 184, September 30, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042794901767.

Ham:1999:ILO

[Ham99]  YoonMee Ham.  Inter-

nal layer oscillations in FitzHugh–Nagumo equation. Journal of Computational and Applied Mathematics, 103(2):287–295, March 31, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042798002647.

Hanna:1996:SCF

[HAMeM96] S. N. Hanna, M. N. Abdel-Malek, and M. B. Abd el Malek. Super-critical freesurface flow over a trapezoidal obstacle. Journal of Computational and Applied Mathematics, 66(1–2):279– 291, January 31, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0377042795001603.

Han:1991:NRS

[Han91a] Weimin Han. A note on a relaxation Schwarz alternating method. Journal of Computational and Applied Mathematics, 34(1):125– 130, February 10, 1991. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/037704279190152A.

Han:1991:NSV

[Han91b] Weimin Han. Notes on “A separation of the variables method for solving coupled systems of secondorder partial differential equations: exact, approximate solutions and error bounds”. Journal of Computational and Applied Mathematics, 36(2):259–263, August 27, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279190033G.

See [Jo’90].

Hanna:1991:LS

[Han91c] S. N. Hanna. Letter section. Journal of Computational and Applied Mathematics, 36(2):274, August 27, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279190036J.

Hanna:1991:SPP

[Han91d] Sarwat N. Hanna. Stability of periodic plane wave solutions of wave equations. Journal of Computational and Applied Mathematics, 34(1):41–46, February 10, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279190146B.

Hanschke:1992:CAD

[Han92] Thomas Hanschke. Characterization of anti dominant solutions of linear differential equations. Journal of Computational and Applied Mathematics, 40(3): 351–361, July 24, 1992. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042792901929.

Han:1994:CEE

[Han94a] Weimin Han. Computable error estimates for linearization and numerical solution of obstacle problems. Journal of Computational and Applied Mathematics, 55(1):69–79, October 31, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042794901856.

Hanna:1994:IST

[Han94b]  Sarwat N. Hanna.  In-

fluence of surface tension on free-surface flow over a polygonal and curved obstruction. Journal of Computational and Applied Mathematics, 51(3):357– 374, June 30, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279200111L.

Han:1996:DVR

[Han96] Gou-Niu Han. Une démonstration “vérificative” d’un résultat de Foata–Zeilberger sur les relations bipartitionnaires. (French) [A “verifying” proof of a Foata– Zeilberger result on bipartitioning relations]. Journal of Computational and Applied Mathematics, 68(1–2):159– 162, April 22, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042795002553.

Hasegawa:1997:NIF

[Has97] Takemitsu Hasegawa. Numerical integration of functions with poles near the interval of integration. Journal of Computational and Applied Mathematics, 87 (2):339–357, December 23, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042797001970.

Hamdan:1990:DGF

[HB90] M. H. Hamdan and R. M.

Barron. A dusty gas flow model in porous media. Journal of Computational and Applied Mathematics, 30(1):21–37, April 10, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
037704279090003I.

Haas:1992:ESC

[HB92a] Roland Haas and Hans Brauchli. Extracting singularities of Cauchy integrals — a key point of a fast solver for plane potential problems with mixed boundary conditions. Journal of Computational and Applied Mathematics, 44(2):167–186, December 17, 1992. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279290009M.

Hamdan:1992:AMC

[HB92b] M. H. Hamdan and R. M. Barron. Applications of von Mises coordinates in porous media flow. Journal of Computational and Applied Mathematics, 39(3): 353–361, May 8, 1992. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/037704279290209G.

Hauschildt:1999:NSE

[HB99] Peter H. Hauschildt and E. Baron. Numerical solution of the expanding stellar atmosphere problem. Journal of Computational and Applied Mathematics, 109(1–2):41–63, September 30, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042799001533.

Hureau:1996:IFS

[HBL96] J. Hureau, E. Brunon, and Ph. Legallais. Ideal free streamline flow over a curved obstacle. Journal of Computational and Applied Mathematics, 72(1): 193–214, July 30, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042795002723.

huuCong:1993:NPD

[hC93] Nguyen huu Cong. Note on the performance of direct and indirect Runge–Kutta– Nyström methods. Journal of Computational and Applied Mathematics, 45(3): 347–355, May 3, 1993. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279390053E.

huuCong:1994:PDM

[hC94a]  Nguyen huu Cong.  A

parallel DIRK method for stiff initial-value problems. Journal of Computational and Applied Mathematics, 54(1):121–127, September 20, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042794903999.

huuCong:1994:PIS

[hC94b] Nguyen huu Cong. Parallel iteration of symmetric Runge–Kutta methods for nonstiff initialvalue problems. Journal of Computational and Applied Mathematics, 51(1): 117–125, May 30, 1994. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0377042794900949.

Hu:1999:ENS

[HC99] Guang-Da Hu and Baruch Cahlon. Estimations on numerically stable step-size for neutral delay differential systems with multiple delays. Journal of Computational and Applied Mathematics, 102(2):221– 234, February 26, 1999. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042798002155.

Hemker:1993:CBD

[HD93] P. W. Hemker and J.- A. Désidéri. Convergence behaviour of defect correction for hyperbolic equations. Journal of Computational and Applied Mathematics, 45(3): 357–365, May 3, 1993. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/037704279390054F.

He:1994:FPF

[He94] Matthew He. The Faber polynomials for m-fold symmetric domains. Journal of Computational and Applied Mathematics, 54(3):313– 324, October 20, 1994. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042794902542.

Heindl:1995:EME

[Hei95a] Gerhard Heindl. Experiences with a method for enclosing solutions of systems of equations. Journal of Computational and Applied Mathematics, 60(1–2): 63–76, June 20, 1995. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279400084E.

Heise:1995:NSE

[Hei95b]  Bodo Heise.  Nonlinear

simulation of electromagnetic fields with domain decomposition methods on MIMD parallel computers. Journal of Computational and Applied Mathematics, 63(1–3):373–381, November 20, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042795000658.

Hemker:1990:NMM

[Hem90a] P. W. Hemker. A nonlinear multigrid method for onedimensional semiconductor device simulation: results for the diode. Journal of Computational and Applied Mathematics, 30(1):117– 126, April 10, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279090010W.

Hemker:1990:OPR

[Hem90b] P. W. Hemker. On the order of prolongations and restrictions in multigrid procedures. Journal of Computational and Applied Mathematics, 32(3):423–429, December 10, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042790900474.

Hemker:1996:SPM

[Hem96] P. W. Hemker. A singularly perturbed model problem for numerical computation. Journal of Computational and Applied Mathematics, 76(1–2):277–285, December 17, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042796001136.

Hendriksen:1990:AJL

[Hen90a] Erik Hendriksen. Associated Jacobi–Laurent polynomials. Journal of Computational and Applied Mathematics, 32(1–2):125–141, November 26, 1990. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279090424X.

Hendriksen:1990:WFA

[Hen90b] Erik Hendriksen. A weight function for the associated Jacobi–Laurent polynomials. Journal of Computational and Applied Mathematics, 33(2):171–180, December 21, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042790903679.

Heusser:1992:CGT

[Heu92] Carmen Heusser. Conjugate gradient-type algorithms for a finite-element discretization of the Stokes equations. Journal of Computational and Applied Mathematics, 39(1):23–37, February 28, 1992. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279290219N.

Higham:1995:ESA

[HF95] Desmond J. Higham and Ioannis Th. Famelis. Equilibrium states of adaptive algorithms for delay differential equations. Journal of Computational and Applied Mathematics, 58(2):151– 169, March 30, 1995. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0377042793E0268Q.

Higueras:1999:RKM

[HGC99] Inmaculada Higueras and Berta García-Celayeta. Runge– Kutta methods for DAEs. a new approach. Journal of Computational and Applied Mathematics, 111(1–2):49– 61, November 15, 1999. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042799001314.

Habib:1996:RAH

[HGL96] A. W. Habib, R. N. Goldman, and T. Lyche. A recursive algorithm for Hermite interpolation over a triangular grid. Journal of Computational and Applied Mathematics, 73(1–2):95–118, October 5, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042796000386.

Higham:1990:ERK

[HH90] D. J. Higham and G. Hall. Embedded Runge–Kutta formulae with stable equilibrium states. Journal of Computational and Applied Mathematics, 29(1):25–33, January 10, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042790901923.

Hu:1997:ENS

[HHL97] Guang-Di Hu, Guang-Da Hu, and Ming-Zhu Liu. Estimation of numerically stable step-size for neutral delaydifferential equations via spectral radius. Journal of Computational and Applied Mathematics, 78(2):311– 316, February 17, 1997. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042796001537.

Higham:1992:MPC

[Hig92] D. J. Higham. Monotonic piecewise cubic interpolation, with applications to ODE plotting. Journal of Computational and Applied Mathematics, 39(3): 287–294, May 8, 1992. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279290205C.

Higham:1993:TPA

[Hig93] Desmond J. Higham. The tolerance proportionality of adaptive ODE solvers. Journal of Computational and Applied Mathematics, 45(1– 2):227–236, April 8, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
037704279390277I.

Hinsen:1992:EIS

[Hin92] G. Hinsen. Explicit irregular sampling formulas. Journal of Computational and Applied Mathematics, 40(2): 177–198, June 24, 1992. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042792901057.

Heijnen:1995:BMS

[HJ95a] B. Heijnen and G. K. Janssens. Bounds for the mean system size in M/G/1/K-queues. Journal of Computational and Applied Mathematics, 64(1– 2):149–161, November 30, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042795000127.

Hongjiong:1995:SMN

[HJ95b] Tian Hongjiong and Kuang Jiaoxun. The stability of the θ-methods in the numerical solution of delay differential equations with several delay terms. Journal of Computational and Applied Mathematics, 58(2):171– 181, March 30, 1995. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042793E0269R.

Horowitz:1994:MRC

[HK94] Joseph Horowitz and Rajeeva L. Karandikar. Mean rates of convergence of empirical measures in the Wasserstein metric. Journal of Computational and Applied Mathematics, 55 (3):261–273, November 30, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042794900337.

Han:1996:ZOP

[HKL96] S. S. Han, K. H. Kwon, and L. L. Littlejohn. Zeros of orthogonal polynomials in certain discrete Sobolev spaces. Journal of Computational and Applied Mathematics, 67(2):309– 325, March 29, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042794001316.

Hsiao:1999:ABI

[HKW99] G. C. Hsiao, R. E. Kleinman, and D.-Q. Wang. Applications of boundary integral equation methods in 3D electromagnetic scattering. Journal of Computational and Applied Mathematics, 104(2): 89–110, May 17, 1999. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042798002246.

Huang:1995:PHA

[HL95] Liang Jiao Huang and Tien-Yien Li. Parallel homotopy algorithm for symmetric large sparse eigenproblems. Journal of Computational and Applied Mathematics, 60(1–2):77–100, June 20, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279400085F.

Hough:1993:NCM

[HLCW93] D. M. Hough, J. Levesley, and S. N. Chandler-Wilde. Numerical conformal mapping via Chebyshev weighted solutions of Symm’s integral equation. Journal of Computational and Applied Mathematics, 46(1–2):29–48, June 14, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
037704279390285J.

Huang:1999:SEA

[HLFC99] Chengming Huang, Shoufu Li, Hongyuan Fu, and Guangnan Chen. Stability and error analysis of one-leg methods for nonlinear delay differential equations. Journal of Computational and Applied Mathematics, 103(2):263–279, March 31, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042798002623.

He:1994:POC

[HLS94] Feiyue He, Anthony Leung, and Srdjan Stojanovic. Periodic optimal control for competing parabolic Volterra–Lotka-type systems. Journal of Computational and Applied Mathematics, 52(1–3):199–217, July 20, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042794903573.

Hylin:1997:DPM

[HM97] E. C. Hylin and J. M. Mc-Donough. Derivation of projection methods from formal integration of the Navier– Stokes equations. Journal of Computational and Applied Mathematics, 81(2): 349–374, July 8, 1997. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042797000733.

Herrmann:1996:SDI

[HMS96] Matthias Herrmann, Bernd Mulansky, and Jochen W. Schmidt. Scattered data interpolation subject to piecewise quadratic range restrictions. Journal of Computational and Applied Mathematics, 73(1–2):209–223, October 5, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042796000441.

Herceg:1990:ECA

[HMT90] Dragoslav Herceg, M. Madalena Martins, and M. Estela Trigo. Extended convergence area for the (MSOR) method. Journal of Computational and Applied Mathematics, 33(2):123–132, December 21, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042790903624.

Herrero:1998:SSB

[HMV98] M. A. Herrero, E. Medina, and J. J. L. Velázquez. Self-similar blow-up for a reaction–diffusion system. Journal of Computational and Applied Mathematics, 97(1–2):99–119, September 24, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042798001046.

Hadjidimos:1990:CDS

[HN90] A. Hadjidimos and M. Neumann. Convergence domains of the SSOR method for a class of generalized consistently ordered matrices. Journal of Computational and Applied Mathematics, 33(1):35–52, December 17, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279090254W.

Hadjidimos:1995:SCD

[HN95] Apostolos Hadjidimos and Michael Neumann. Superior convergence domains for the p-cyclic SSOR majorizer. Journal of Computational and Applied Mathematics, 62(1):27–40, August 31, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042794000981.

Hunter:1998:GQC

[HN98] D. B. Hunter and Geno Nikolov. Gaussian quadrature of Chebyshev polynomials. Journal of Computational and Applied Mathematics, 94(2):123–131, August 3, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042798000740.

Hadjidimos:1996:CDC

[HNT96] A. Hadjidimos, D. Noutsos, and M. Tzoumas. On the convergence domains of the p-cyclic SOR. Journal of Computational and Applied Mathematics, 72(1):63–83, July 30, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042795002456.

Hadjidimos:1998:TDO

[HNT98] A. Hadjidimos, D. Noutsos, and M. Tzoumas. Towards the determination of the optimal p-cyclic SSOR. Journal of Computational and Applied Mathematics, 90(1): 1–14, April 6, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042797002471.

Hochbruck:1997:FOL

[Hoc97] Marlis Hochbruck. Further optimized look-ahead recurrences for adjacent rows in the Padé table and Toeplitz matrix factorizations. Journal of Computational and Applied Mathematics, 86 (1):219–236, November 28, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704279700157X.

Holecek:1995:MCB

[Hol95] M. Holeček. Model of critical behaviour in elasticity. Journal of Computational and Applied Mathematics, 63(1–3):271–275, November 20, 1995. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279500081X.

Homeier:1996:ANS

[Hom96] Herbert H. H. Homeier.

Analytical and numerical studies of the convergence behavior of the j transformation. Journal of Computational and Applied Mathematics, 69(1):81–112, April 30, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042795000224.

Hovstad:1990:CFI

[Hov90] R. M. Hovstad. Continued fractions and irrational functions. Journal of Computational and Applied Mathematics, 33(2):157–162, December 21, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042790903657.

Howes:1990:ANM

[How90] F. A. Howes. An asymptoticnumerical method for a model system with turning points. Journal of Computational and Applied Mathematics, 29(3):343– 356, March 15, 1990. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/037704279090016S.

Howell:1993:NCM

[How93] Louis H. Howell. Numerical conformal mapping of circular arc polygons. Journal of Computational and Applied Mathematics, 46(1–2):7–28, June 14, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279390284I.

Hanin:1994:MTP

[HR94] L. G. Hanin and S. T. Rachev. Mass-transshipment problems and ideal metrics. Journal of Computational and Applied Mathematics, 56(1–2):183–196, December 20, 1994. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042794903875.

Hill:1999:DNA

[HR99] Michael Hill and Ian Robinson. d2lri: a nonadaptive algorithm for twodimensional cubature. Journal of Computational and Applied Mathematics, 112 (1–2):121–145, November 30, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042799002174.

Hernandez:1998:CRP

[HRGL98] M. Bello Hernández, F. Cala Rodríguez, J. J. Guadalupe, and G. López Lagomasino. Convergence rate of Padétype approximants for Stieltjes functions. Journal of Computational and Applied Mathematics, 99(1–2):47– 53, November 16, 1998. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042798001447.

Hadjidimos:1992:MSO

[HS92a] Apostolos Hadjidimos and Yiannis G. Saridakis. Modified successive overrelaxation (MSOR) and equivalent 2-step iterative methods for collocation matrices. Journal of Computational and Applied Mathematics, 42(3):375–393, October 30, 1992. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279290086D.

Hart:1992:SHS

[HS92b] W. E. Hart and F. Soesianto. On the solution of highly structured nonlinear equations. Journal of Computational and Applied Mathematics, 40(3): 285–296, July 24, 1992. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/037704279290184Y.

He:1992:EIP

[HS92c] Sailing He and Staffan Ström. The electromagnetic inverse problem in the time domain for a dissipative slab and a point source using invariant imbedding: reconstruction of the permittivity and conductivity. Journal of Computational and Applied Mathematics, 42(1):137– 155, September 8, 1992. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/037704279290168W.

Hartmann:1994:RCC

[HS94a] T. Hartmann and E. P.

Stephan. Rates of convergence for collocation with Jacobi polynomials for the airfoil equation. Journal of Computational and Applied Mathematics, 51(2): 179–191, June 20, 1994. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279200005T.

Hess:1994:PQM

[HS94b] Walter Heß and Jochen W.

Schmidt. Positive quartic, monotone quintic C ^{2}-spline interpolation in one and two dimensions. Journal of Computational and Applied Mathematics, 55(1):51–67, October 31, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042794901848.

Hughes:1996:STF

[HS96] Thomas J. R. Hughes and James R. Stewart. A spacetime formulation for multiscale phenomena. Journal of Computational and Applied Mathematics, 74(1–2):217– 229, November 5, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042796000258.

Heuveline:1997:PCP

[HS97] Vincent Heuveline and Miloud Sadkane. Parallel computation of polynomials with minimal uniform norm and its application to large eigenproblems. Journal of Computational and Applied Mathematics, 82(1–2):185–198, September 15, 1997. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042797000460.

Hsiao:1991:DPE

[HSW91] George C. Hsiao, Ernst P. Stephan, and Wolfgang L. Wendland. On the Dirichlet problem in elasticity for a domain exterior to an arc. Journal of Computational and Applied Mathematics, 34(1):1–19, February 10, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042791901438.

Hasegawa:1991:AMF

[HT91] Takemitsu Hasegawa and Tatsuo Torii. Application of a modified FFT to product type integration. Journal of Computational and Applied Mathematics, 38(1–3):157–168, December 23, 1991. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279190168J.

Hasegawa:1994:HHT

[HT94] Takemitsu Hasegawa and Tatsuo Torii. Hilbert and Hadamard transforms by generalized Chebyshev expansion. Journal of Computational and Applied Mathematics, 51(1):71–83, May 30, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279200003R.

Hix:1999:CMN

[HT99] W. Raphael Hix and Friedrich-Karl Thielemann. Computational methods for nucleosynthesis and nuclear energy generation. Journal of Computational and Applied Mathematics, 109(1–2):321–351, September 30, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042799001636.

Hall:1999:MOS

[HU99] George Hall and Anila Usman. Modified order and stepsize strategies in Adams codes. Journal of Computational and Applied Mathematics, 111(1–2):113–122, November 15, 1999. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042799001351.

Hugger:1994:AEP

[Hug94] Jens Hugger. An asymptotically exact, pointwise, a posteriori error estimator for linear, elliptic, one-dimensional problems with possible singularities in the solution. Journal of Computational and Applied Mathematics, 54(2):185–220, September 30, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042794901775.

Hurlimann:1997:BBEa

[Hür97a] Werner Hürlimann. Best bounds for expected financial payoffs I algorithmic evaluation. Journal of Computational and Applied Mathematics, 82(1–2):199– 212, September 15, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042797000538.

Hurlimann:1997:BBEb

[Hür97b] Werner Hürlimann. Best bounds for expected financial payoffs II: Applications.

Journal of Computational and Applied Mathematics, 82(1–2):213–227, September 15, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042797000551.

Hussein:1994:EST

[Hus94]  H. Attia Hussein.  An

explicit solution of thirdorder difference equations. Journal of Computational and Applied Mathematics, 54(3):307–311, October 20, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042794902534.

Hughes:1991:UPA

[HV91] J. W. B. Hughes and J. Van der Jeugt. Unimodal polynomials associated with Lie algebras and superalgebras. Journal of Computational and Applied Mathematics, 37(1–3):81– 88, November 18, 1991. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279190108V.

Haneveld:1994:EVF

[HvdV94] Willem K. Klein Haneveld and Maarten H. van der Vlerk. On the expected value function of a simple integer recourse problem with random technology matrix. Journal of Computational and Applied Mathematics, 56(1–2):45–53, December 20, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042794903786.

Hairer:1999:SDE

[HW99] Ernst Hairer and Gerhard Wanner. Stiff differential equations solved by Radau methods. Journal of Computational and Applied Mathematics, 111(1–2):93– 111, November 15, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704279900134X.

Hwang:1996:AEL

[Hwa96] Hsien-Kuei Hwang. Asymptotic expansion for the Lebesgue constants of the Walsh system. Journal of Computational and Applied Mathematics, 71(2): 237–243, July 27, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042795002367.

Heo:1999:CCF

[HX99] Sangwoo Heo and Yuan Xu. Constructing cubature formulae for spheres and balls. Journal of Computational and Applied Mathematics, 112(1–2):95–119, November 30, 1999. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042799002162.

Henderson:1998:SBV

[HY98] Johnny Henderson and William Yin. Singular ( k, n- k ) boundary value problems between conjugate and right focal. Journal of Computational and Applied Mathematics, 88(1):57–69, February 23, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042797002070.

Hyvonen:1997:CAP

[Hyv97] Saara Hyvönen. Convergence of the Arnoldi process when applied to the Picard– Lindelöf iteration operator. Journal of Computational and Applied Mathematics, 87(2):303–320, December 23, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042797001957.

Han:1995:QJS

[HZ95] Weimin Han and Jinshi Zhou. Quantitative justification of simplifications in some heat conduction problems. Journal of Computational and Applied Mathematics, 62(1):75–88, August 31, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042794000762.

Helfrich:1996:TRA

[HZ96] H.-P. Helfrich and D. Zwick. A trust region algorithm for parametric curve and surface fitting. Journal of Computational and Applied Mathematics, 73(1–2):119– 134, October 5, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0377042796000398.

Han:1994:NCM

[HZL94] Bo Han, Dali Zhang, and Jiaqi Liu. A new class of methods for solving nonlinear equations. Journal of Computational and Applied Mathematics, 51(1): 107–111, May 30, 1994. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0377042794000093.

Iordan:1999:NCS

[IB99] Victoria Iordan and Stefan Balint. Numerical capturing of some hydrodynamical shocks. Journal of Computational and Applied Mathematics, 111(1–2):297–303, November 15, 1999. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042799001508.

Ixaru:1998:CMS

[IDV98] L. Gr. Ixaru, H. De Meyer, and G. Vanden Berghe. CP methods for the Schrödinger equation revisited. Journal of Computational and Applied Mathematics, 88(2):289– 314, March 2, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042797002185.

Iyengar:1990:NMT

[IG90] Satteluri R. K. Iyengar and Anu Goyal. A note on multigrid for the threedimensional Poisson equation in cylindrical coordinates. Journal of Computational and Applied Mathematics, 33(2):163–169, December 21, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042790903668.

Ikebe:1991:CZO

[IKF91] Yasuhiko Ikebe, Yasushi Kikuchi, and Issei Fujishiro. Computing zeros and orders of Bessel functions. Journal of Computational and Applied Mathematics, 38(1–3):169–184, December 23, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279190169K.

Ifantis:1991:PZS

[IKK91] E. K. Ifantis, C. G. Kokologiannaki, and C. B. Kouris. On the positive zeros of the second derivative of Bessel functions. Journal of Computational and Applied Mathematics, 34(1):21–31, February 10, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042791901449.

Infante:1999:OSR

[IL99] A. Fernández Infante and G. López Lagomasino. Overconvergence of subsequences of rows of Padé approximants with gaps. Journal of Computational and Applied Mathematics, 105(1–2):265– 273, May 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042799000448.

Illan:1998:PRA

[Ill98]  J. Illán.  Piecewise ra-

tional approximation to continuous functions with characteristic singularities. Journal of Computational and Applied Mathematics, 99(1–2):195–203, November 16, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042798001575.

Ismail:1996:P

[IM96] Mourad E. H. Ismail and David R. Masson. Preface. Journal of Computational and Applied Mathematics, 68(1–2):1–2, April 22, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042796900333.

Ibanez:1999:RSR

[IM99] J. M^{a}. Ibáñez and J. M^{a}.

Martí. Riemann solvers in relativistic astrophysics. Journal of Computational and Applied Mathematics, 109(1–2):173–211, September 30, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042799001582.

Iso:1991:SBE

[IO91] Yuusuke Iso and Kazuei Onishi. On the stability of the boundary element collocation method applied to the linear heat equation. Journal of Computational and Applied Mathematics, 38(1–3):201–209, December 23, 1991. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/037704279190171F.

Ioakimidis:1990:GMT

[Ioa90] N. I. Ioakimidis. Generalized mangler-type principal value integrals with an application to fracture mechanics. Journal of Computational and Applied Mathematics, 30(2): 227–234, May 28, 1990. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0377042790900304.

Ixaru:1999:CSF

[IP99] L. Gr. Ixaru and B. Paternoster. A conditionally P -stable fourth-order exponential-fitting method for y′′ = f(x, y). Journal of Computational and Applied Mathematics, 106(1): 87–98, June 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042799000552.

Iqbal:1995:NTR

[Iqb95]  M. Iqbal.  On a nu-

merical technique regarding inversion of the Laplace transform. Journal of Computational and Applied Mathematics, 59(2):145– 154, May 19, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279400028Y.

Iqbal:1997:SRI

[Iqb97]  M. Iqbal.  On spline

regularized inversion of noisy Laplace transforms. Journal of Computational and Applied Mathematics, 83(1):39–54, September 30, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042797000289.

Ixaru:1997:FSM

[IR97] L. Gr Ixaru and M. Rizea.

Four step methods for y′′ = f(x, y). Journal of Computational and Applied Mathematics, 79(1):87–99, March 3, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042796001653.

Iraniparast:1994:BVP

[Ira94] Nezam Iraniparast. Boundary value problems for a two-dimensional wave equation. Journal of Computational and Applied Mathematics, 55(3):349– 356, November 30, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042794900388.

Ismail:1996:DCI

[IRZ96] Mourad E. H. Ismail, Mizan Rahman, and Ruiming Zhang. Diagonalization of certain integral operators II. Journal of Computational and Applied Mathematics, 68(1–2):163–196, April 22, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042795002634.

Ifantis:1990:DIP

[IS90] E. K. Ifantis and P. D. Siafarikas. Differential inequalities for the positive zeros of Bessel functions. Journal of Computational and Applied Mathematics, 30(2): 139–143, May 28, 1990. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279090022R.

Ifantis:1992:DIP

[IS92] E. K. Ifantis and P. D.

Siafarikas. A differential inequality for the positive zeros of Bessel functions. Journal of Computational and Applied Mathematics, 44(1):115–120, December 9, 1992. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042792900553.

Ifantis:1993:ZCP

[IS93] E. K. Ifantis and P. D.

Siafarikas. On the zeros of a class of polynomials including the generalized Bessel polynomials. Journal of Computational and Applied Mathematics, 49(1–3):103–109, December 31, 1993. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042793901407.

Ifantis:1995:APT

[IS95a] E. K. Ifantis and P. D.

Siafarikas. An alternative proof of a theorem of Stieltjes and related results. Journal of Computational and Applied Mathematics, 65(1–3):165–172, December 29, 1995. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0377042795001239.

Ifantis:1995:PCR

[IS95b] E. K. Ifantis and P. D. Siafarikas. Perturbation of the coefficients in the recurrence relation of a class of polynomials. Journal of Computational and Applied Mathematics, 57(1–2):163–170, February 20, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042793E0242E.

Ismail:1998:SAK

[IS98] Mourad E. H. Ismail and Plamen Simeonov. Strong asymptotics for Krawtchouk polynomials. Journal of Computational and Applied Mathematics, 100(2):121– 144, December 21, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042798001836.

Isakov:1992:SEO

[Isa92] Victor Isakov. Stability estimates for obstacles in inverse scattering. Journal of Computational and Applied Mathematics, 42(1):79–88, September 8, 1992. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279290164S.

Iserles:1993:CAS

[Ise93] A. Iserles. A constructive approach to the Schröder equation. Journal of Computational and Applied Mathematics, 46(1–2):301– 314, June 14, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279390304T.

Ishihara:1991:PSO

[Ish91] Kazuo Ishihara. Projected successive overrelaxation method for finite-element solutions to the Dirichlet problem for a system of nonlinear elliptic equations. Journal of Computational and Applied Mathematics, 38(1–3):185–200, December 23, 1991. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279190170O.

Ishiwata:1999:PCP

[Ish99] Emiko Ishiwata. A practical choice of parameters in improved SOR– Newton method with orderings. Journal of Computational and Applied Mathematics, 102(2):315–331, February 26, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042798002325.

Ito:1994:DAM

[Ito94] Yoshifusa Ito. Differentiable approximation by means of the Radon transformation and its applications to neural networks. Journal of Computational and Applied Mathematics, 55(1):31–50, October 31, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279490183X.

Igarashi:1995:RBO

[IY95] Masao Igarashi and Tjalling Ypma. Relationships between order and efficiency of a class of methods for multiple zeros of polynomials. Journal of Computational and Applied Mathematics, 60(1–2):101–113, June 20, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279400086G.

Igarashi:1997:EVA

[IY97] Masao Igarashi and Tjalling Ypma. Empirical versus asymptotic rate of convergence of a class of methods for solving a polynomial equation. Journal of Computational and Applied Mathematics, 82(1–2):229–237, September 15, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042797000770.

Jodar:1992:ESS

[JA92] L. Jódar and J. A. Martin Alustiza. Explicit solution of systems of second-order differential-difference equations. Journal of Computational and Applied Mathematics, 43(3):323–334, December 2, 1992. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279290018S.

Joshi:1996:ISS

[JB96] C. M. Joshi and S. K.

Bissu. Inequalities for some special functions. Journal of Computational and Applied Mathematics, 69(2): 251–259, May 23, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042795000429.

Jay:1999:PPI

[JB99] Laurent O. Jay and Thierry Braconnier. A parallelizable preconditioner for the iterative solution of implicit Runge–Kutta-type methods. Journal of Computational and Applied Mathematics, 111(1–2):63–76, November 15, 1999. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042799001326.

Jbilou:1999:SIB

[Jbi99] K. Jbilou. Smoothing iterative block methods for linear systems with multiple right-hand sides. Journal of Computational and Applied Mathematics, 107(1): 97–109, July 15, 1999. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042799000837.

Jiaxiang:1995:NSD

[JC95] Xiang Jiaxiang and Ian T. Cameron. Numerical solution of DAE systems using block BDF methods. Journal of Computational and Applied Mathematics, 62(3):255–266, September 29, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042794001111.

Jodar:1998:HMF

[JC98] L. Jódar and J. C. Cortés.

On the hypergeometric matrix function. Journal of Computational and Applied Mathematics, 99(1–2):205– 217, November 16, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042798001587.

Janssens:1995:ORR

[JDV95] Gerrit K. Janssens, Wim Deceuninck, and Alex Van Breedam. Opportunities of robust regression for variance reduction in discrete event simulation. Journal of Computational and Applied Mathematics, 64(1–2):163–176, November 30, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042795000135.

Jekot:1991:SMN

[Jek91] T. Jekot. Spectral methods for nonlinear thermostatic bimetal. Journal of Computational and Applied Mathematics, 35(1–3):251– 258, June 26, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042791902112.

Jerome:1991:NAP

[Jer91] Joseph W. Jerome. Numerical approximation of PDE system fixed-point maps via Newton’s method. Journal of Computational and Applied Mathematics, 38(1–3):211–230, December 23, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279190172G.

Jersak:1995:NSQ

[Jer95] J. Jersák. Numerical simulations in quantum field theory of elementary particles. Journal of Computational and Applied Mathematics, 63(1–3):49–56, November 20, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042795000607.

Jhunjhunwala:1995:WLT

[Jhu95] N. Jhunjhunwala. Weighted Landau-type inequalities. Journal of Computational and Applied Mathematics, 57(3):323–328, February 28, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042793E0184N.

Jiang:1991:UNS

[Jia91] Song Jiang. An uncoupled numerical scheme for the equations of nonlinear one-dimensional thermoelasticity. Journal of Computational and Applied Mathematics, 34(2):135– 144, April 4, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279190037K.

Jiang:1994:AEA

[Jia94] Tianzi Jiang. Asymptotic expansion for the approximation of the Lipschitz function by the Hermitefejer interpolation polynomial. Journal of Computational and Applied Mathematics, 54(3):285–296, October 20, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042794902518.

Jin:1996:BTP

[Jin96a]  Xiao-Qing Jin.  Band

Toeplitz preconditioners for block Toeplitz systems. Journal of Computational and Applied Mathematics, 70(2):225–230, June 28, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042795002057.

Jin:1996:FIS

[Jin96b] Xiao-Qing Jin. Fast iterative solvers for symmetric Toeplitz systems — a survey and an extension. Journal of Computational and Applied Mathematics, 66(1–2):315– 321, January 31, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042795001565.

Jeon:1998:NBE

[JK98a] Youngmok Jeon and Ha-Jine Kimn. New boundary element formulas for plane elasticity. Journal of Computational and Applied Mathematics, 89(2):263– 274, March 16, 1998. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S037704279700246X.

Jung:1998:MCH

[JK98b] H. S. Jung and K. H. Kwon. Mean convergence of Hermite–Fejér and Hermite interpolation for Freud weights. Journal of Computational and Applied Mathematics, 99(1–2):219–238, November 16, 1998. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042798001599.

Jung:1995:DES

[JKLL95] I. H. Jung, K. H. Kwon, D. W. Lee, and L. L. Littlejohn. Differential equations and Sobolev orthogonality. Journal of Computational and Applied Mathematics, 65(1–3):173–180, December 29, 1995. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042795001115.

Jung:1997:DEI

[JKY97] I. H. Jung, K. H. Kwon, and G. J. Yoon. Differential equations of infinite order for Sobolev-type orthogonal polynomials. Journal of Computational and Applied Mathematics, 78(2):277– 293, February 17, 1997. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042796001434.

Janssen:1992:RCP

[JL92] Hugo L. Janssen and Hendrik L. Lambert. Recursive construction of particular solutions to nonhomogeneous linear partial differential equations of elliptic type. Journal of Computational and Applied Mathematics, 39(2):227– 242, March 30, 1992. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/037704279290132H.

Junghanns:1997:CSI

[JL97] Peter Junghanns and Uwe Luther. Cauchy singular integral equations in spaces of continuous functions and methods for their numerical solution. Journal of Computational and Applied Mathematics, 77(1–2):201– 237, January 13, 1997. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042796001288.

Jou:1999:PBE

[JL99] Jang Jou and Jinn-Liang Liu. A posteriori boundary element error estimation. Journal of Computational and Applied Mathematics, 106(1):1–19, June 1999. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042799000497.

Jacover:1994:NTF

[JM94] Daniela Jacover and P. J. McKenna. Nonlinear torsional flexings in a periodically forced suspended beam. Journal of Computational and Applied Mathematics, 52(1–3):241–265, July 20, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279490359X.

Jumarhon:1996:PIM

[JM96] Bartur Jumarhon and Sean McKee. Product integration methods for solving a system of nonlinear Volterra integral equations. Journal of Computational and Applied Mathematics, 69(2): 285–301, May 23, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042795000410.

Jackiewicz:1999:ESC

[JM99] Z. Jackiewicz and H. D.

Mittelmann. Exploiting structure in the construction of DIMSIMs. Journal of Computational and Applied Mathematics, 107(2): 233–239, July 31, 1999. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042799000916.

Jones:1990:APF

[JMMW90] William B. Jones, Arne Magnus, John H. McCabe, and Nancy J. Wyshinski. Asymptotic properties of a family of orthogonal polynomial sequences. Journal of Computational and Applied Mathematics, 32(1–2):143–151, November 26, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279090425Y.

Jumarhon:1994:PIA

[JMT94] B. Jumarhon, S. McKee, and T. Tang. The proof of an inequality arising in a reaction–diffusion study in a small cell. Journal of Computational and Applied Mathematics, 51(1):99–101, May 30, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042794900922.

Jodar:1991:ECS

[JN91] Lucas Jódar and Enrique Navarro. Exact computable solution of a class of strongly coupled Riccati equations. Journal of Computational and Applied Mathematics, 36(2):265– 271, August 27, 1991. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279190034H.

Jodar:1992:SCS

[JN92] Lucas Jódar and Enrique Navarro. Solving coupled systems of linear secondorder differential equations knowing a part of the spectrum of the companion matrix. Journal of Computational and Applied Mathematics, 39(1):115– 119, February 28, 1992. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279290228P.

Jones:1999:OLP

[JN99] William B. Jones and Olav Njåstad. Orthogonal Laurent polynomials and strong moment theory: a survey. Journal of Computational and Applied Mathematics, 105(1–2):51– 91, May 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042799000278.

Jodar:1999:ANS

[JNC99] L. Jódar, E. Navarro, and J. Camacho. Analyticnumerical solutions with a priori error bounds for a class of strongly coupled mixed partial differential systems. Journal of Computational and Applied Mathematics, 104(2):123– 143, May 17, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042799000485.

Jones:1990:SPA

[JNS90] William B. Jones, Olav Njåstad, and E. B. Saff. Szegő polynomials associated with Wiener–Levinson filters. Journal of Computational and Applied Mathematics, 32(3):387–406, December 10, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279090044Z.

Jodar:1990:SVM

[Jo’90] Lucas Jo’dar. A separation of the variables method for solving coupled systems of second-order partial differential equations: exact, approximate solutions and error bounds. Journal of Computational and Applied Mathematics, 31(2):243– 251, August 31, 1990. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/037704279090167X.

See notes [Han91b].

Jodar:1990:BVP

[Jód90] Lucas Jódar. Boundary value problems for systems of second-order differential equations: approximations, convergence and upper error bounds. Journal of Computational and Applied Mathematics, 30(1):39–54, April 10, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279090004J.

Jodar:1991:CFS

[Jód91a] Lucas Jódar. Closed-form solutions for boundary value coupled differential and difference systems. Journal of Computational and Applied Mathematics, 35(1–3):259– 268, June 26, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042791902123.

Jodar:1991:ECF

[Jód91b]  Lucas Jódar.  Explicit

closed-form solution of coupled systems of Volterra integrodifferential systems. Journal of Computational and Applied Mathematics, 36(3):339–348, September 24, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279190014B.

Jodar:1992:AAS

[Jód92] Lucas Jódar. Analytical approximate solutions and error bounds for nonsymmetric Riccati matrix differential equations. Journal of Computational and Applied Mathematics, 41(3):377– 383, August 28, 1992. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/037704279290144M.

Joe:1990:RLR

[Joe90]  S. Joe.  Randomization

of lattice rules for numerical multiple integration. Journal of Computational and Applied Mathematics, 31(2):299–304, August 31, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279090172V.

Joe:1996:URS

[Joe96] Stephen Joe. A useful result in searches for good vertexmodified number-theoretic rules. Journal of Computational and Applied Mathematics, 72(2):371–377, August 13, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042796000076.

John:1998:PEE

[Joh98] Volker John. A posteriori L^{2}-error estimates for the nonconforming P_{1}P_{0}-finite element discretization of the Stokes equations. Journal of Computational and Applied Mathematics, 96(2):99–116, September 15, 1998. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042798000958.

Jones:1993:USM

[Jon93] D. J. Jones. Use of a shooting method to compute eigenvalues of fourthorder two-point boundary value problems. Journal of Computational and Applied Mathematics, 47 (3):395–400, September 30, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
037704279390065J.

Janovsky:1991:AAP

[JP91] Vladimír Janovský and Petr Plechác. Asymptotic analysis of perturbed Takens–Bogdanov points. Journal of Computational and Applied Mathematics, 36(3):349–359, September 24, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279190015C.

Jagels:1993:CSP

[JR93] Carl Jagels and Lothar Reichel. On the construction of Szegő polynomials. Journal of Computational and Applied Mathematics, 46(1–2): 241–254, June 14, 1993. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/037704279390299Q.

Jbilou:1991:SRA

[JS91] Khalide Jbilou and Hassane Sadok. Some results about vector extrapolation methods and related fixed-point iterations. Journal of Computational and Applied Mathematics, 36(3):385–398, September 24, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279190018F.

Jansche:1992:BWP

[JS92] S. Jansche and R. L. Stens. Best weighted polynomial approximation on the real line; a functionalanalytic approach. Journal of Computational and Applied Mathematics, 40(2): 199–213, June 24, 1992. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0377042792901068.

Jayasri:1993:BTO

[JS93] C. Jayasri and Y. Sitaraman. On a Bernsteintype operator of bleimann, butzer and Hahn. Journal of Computational and Applied Mathematics, 47(2): 267–272, ???? 1993. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279390008Y.

Jbilou:1997:HVT

[JS97a] K. Jbilou and H. Sadok.

Hybrid vector transformations. Journal of Computational and Applied Mathematics, 81(2):257– 267, July 8, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042797000642.

Jukic:1997:EOS

[JS97b] D. Jukić and R. Scitovski.

Existence of optimal solution for exponential model by least squares. Journal of Computational and Applied Mathematics, 78(2):317– 328, February 17, 1997. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042796001604.

Janovsky:1995:NMT

[JSWW95] V. Janovský, S. Shaw, M. K. Warby, and J. R. Whiteman. Numerical methods for treating problems of viscoelastic isotropic solid deformation. Journal of Computational and Applied Mathematics, 63(1–3):91– 107, November 20, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042795000593.

Jones:1993:SPA

[JTNW93] William B. Jones, W. J. Thron, Olav Njåstad, and Haakon Waadeland. Szegő polynomials applied to frequency analysis. Journal of Computational and Applied Mathematics, 46(1–2):217– 228, June 14, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279390297O.

Junbin:1992:NFE

[Jun92] Gao Junbin. A new finite element of C ^{1} cubic splines. Journal of Computational and Applied Mathematics, 40(3):305– 312, July 24, 1992. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042792901862.

Junghanns:1994:NAN

[Jun94] Peter Junghanns. Numerical analysis of Newton projection methods for nonlinear singular integral equations. Journal of Computational and Applied Mathematics, 55(2):145–163, November 21, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042794900175.

Juttler:1997:SFU

[Jüt97] Bert Jüttler. Surface fitting using convex tensor-product splines. Journal of Computational and Applied Mathematics, 84(1):23–44, October 21, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042797001003.

Janssen:1996:AHO

[JV96] Hugo Janssen and René Van Dooren. Application of highorder difference methods for the study of period doubling bifurcations in nonlinear oscillators. Journal of Computational and Applied Mathematics, 66(1–2):293– 313, January 31, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0377042795001883.

Jackiewicz:1997:RPM

[JVZ97] Z. Jackiewicz, R. Vermiglio, and M. Zennaro. Regularity properties of multistage integration methods. Journal of Computational and Applied Mathematics, 87(2):285–302, December 23, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042797001945.

Jiang:1991:PAD

[JW91] Hong Jiang and Yau Shu Wong. A parallel alternating direction implicit preconditioning method. Journal of Computational and Applied Mathematics, 36(2):209– 226, August 27, 1991. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/037704279190027H.

Kaliaguine:1995:OMP

[Kal95a] V. Kaliaguine. The operator moment problem, vector continued fractions and an explicit form of the Favard theorem for vector orthogonal polynomials. Journal of Computational and Applied Mathematics, 65(1–3):181–193, December 29, 1995. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042795001093.

Kalkreuter:1995:MMP

[Kal95b] Thomas Kalkreuter. Multigrid methods for propagators in lattice gauge theories. Journal of Computational and Applied Mathematics, 63(1–3):57–68, November 20, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042795000496.

Kametaka:1991:CFS

[Kam91] Yoshinori Kametaka. On counting functions of selfsimilar sequences. Journal of Computational and Applied Mathematics, 38(1– 3):255–262, December 23, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
037704279190174I.

Kankova:1994:NES

[Kan94] Vlasta Kanková. A note on estimates in stochastic programming. Journal of Computational and Applied Mathematics, 56(1–2):97– 112, December 20, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042794903816.

Kananthai:1997:DRU

[Kan97] Amnuay Kananthai. On the distribution related to the ultra-hyperbolic equations. Journal of Computational and Applied Mathematics, 84(1):101–106, October 21, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042797001143.

Kananthai:1998:CER

[Kan98] Amnuay Kananthai. On the convolution equation related to the diamond kernel of marcel Riesz. Journal of Computational and Applied Mathematics, 100(1):33–39, November 30, 1998. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042798001289.

Karlsson:1991:IHF

[Kar91]  Per W. Karlsson.  In-

ductance and hypergeometric functions. Journal of Computational and Applied Mathematics, 37(1–3):171– 177, November 18, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
037704279190115Z.

Karageorghis:1994:CST

[Kar94] Andreas Karageorghis. A conforming spectral technique for biharmonic-type problems in rectangular domains. Journal of Computational and Applied Mathematics, 51(2):275– 278, June 20, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042794000433.

Karanjai:1991:TDS

[KB91] S. Karanjai and G. Biswas.

Time-dependent scattering and transmission functions. Journal of Computational and Applied Mathematics, 37(1–3):221–235, November 18, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042791901209.

Kutluay:1997:NSO

[KBÖ97] S. Kutluay, A. R. Bahadir, and A. Özdes. The numerical solution of one-phase classical Stefan problem. Journal of Computational and Applied Mathematics, 81(1):135– 144, June 1, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042797000344.

Kutluay:1999:NSO

[KBÖ99] S. Kutluay, A. R. Bahadir, and A. Özdes. Numerical solution of one-dimensional Burgers equation: explicit and exact-explicit finite difference methods. Journal of Computational and Applied Mathematics, 103(2):251– 261, March 31, 1999. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042798002611.

Kim:1992:ENL

[KC92] S. K. Kim and A. T. Chronopoulos. An efficient nonsymmetric Lanczos method on parallel vector computers. Journal of Computational and Applied Mathematics, 42(3):357– 374, October 30, 1992. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/037704279290085C.

Kedzierawski:1993:ISP

[Ked93] Andrzej W. Kedzierawski. The inverse scattering problem for time-harmonic acoustic waves in an inhomogeneous medium with complex refraction index. Journal of Computational and Applied Mathematics, 47(1):83–100, June 28, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279390092P.

Kemp:1990:PRA

[Kem90a] A. W. Kemp. Patchwork rejection algorithms. Journal of Computational and Applied Mathematics, 31(1): 127–131, July 24, 1990. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279090343X.

Kemp:1990:NAG

[Kem90b] C. D. Kemp. New algorithms for generating Poisson variates. Journal of Computational and Applied Mathematics, 31(1):133– 137, July 24, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279090344Y.

Khanh:1994:HPC

[Kha94] Bui Doan Khanh. Hermite predictor–corrector scheme for regular Volterra integral equations and for some integro-differential equations of turbulent diffusion. Journal of Computational and Applied Mathematics, 51(3):305–316, June 30, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        037704279200012X.

Khanh:1996:NRG

[Kha96] Bui Doan Khanh. A numerical resolution of the Gelfand–Levitan equation. Journal of Computational and Applied Mathematics, 72(2):235–244, August 13, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042795002758.

King:1990:CPS

[Kin90] J. Thomas King. On the construction of preconditioners by subspace decomposition. Journal of Computational and Applied Mathematics, 29(2):195– 205, February 15, 1990. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0377042790903576.

Kirsch:1992:I

[Kir92] Andreas Kirsch. Introduction. Journal of Computational and Applied Mathematics, 42(1):1–3, September 8, 1992. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279290158T.

Kitagawa:1991:ASF

[Kit91] Takashi Kitagawa. Asymptotic stability of the fundamental solution method. Journal of Computational and Applied Mathematics, 38(1–3):263–269, December 23, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279190175J.

Kjurkchiev:1995:IAE

[Kju95]  N. Kjurkchiev.  Initial

approximations in Euler– Chebyshev’s method. Journal of Computational and Applied Mathematics, 58(2): 233–236, March 30, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042795000319.

Klempous:1991:NTN

[KK91] Ryszard Klempous and Jerzy Kotowski. Nonlinear transport network design. Journal of Computational and Applied Mathematics, 35(1–3):269–275, June 26, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042791902134.

Koukal:1995:CAQ

[KK95] Stanislav Koukal and Michal Krízek. Curved affine quadratic finite elements. Journal of Computational and Applied Mathematics, 63(1–3):333–339, November 20, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042795000615.

Krupnikov:1997:SSC

[KK97] E. D. Krupnikov and K. S. Kölbig. Some special cases of the generalized hypergeometric function _{q+1}F_{q}. Journal of Computational and Applied Mathematics, 78(1):79–95, February 3, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042796001112.

Kim:1997:OPT

[KKL97] Y. J. Kim, K. H. Kwon, and J. K. Lee. Orthogonal polynomials in two variables and second-order partial differential equations. Journal of Computational and Applied Mathematics, 82(1–2):239–260, September 15, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042797000824.

Kim:1998:PDE

[KKL98] Y. J. Kim, K. H. Kwon, and J. K. Lee. Partial differential equations having orthogonal polynomial solutions. Journal of Computational and Applied Mathematics, 99(1–2):239–253, November 16, 1998. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042798001605.

Klempous:1997:OAO

[KKNU97] Ryszard Klempous, Jerzy Kotowski, Jan Nikodem, and Jedrzej Ulasiewicz. Optimization algorithms of operative control in water distribution systems. Journal of Computational and Applied Mathematics, 84(1):81–99, October 21, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042797001131.

Kim:1996:ABC

[KKS96a] Dongryeol Kim, Jinsoo Kim, and Dongwoo Sheen. Absorbing boundary conditions for wave propagation in viscoelastic media. Journal of Computational and Applied Mathematics, 76(1–2):301–314, December 17, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704279600115X.

Klempous:1996:IPL

[KKS96b] Ryszard Klempous, Jerzy Kotowski, and Ewa Szlachcic. Interactive procedures in large-scale twodimensional cutting stock problems. Journal of Computational and Applied Mathematics, 66(1–2):323– 331, January 31, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042795002006.

Kompis:1995:MLE

[KKZ95] V. Kompis, M. Kaukic, and M. Zmindák. Modelling of local effects by hybrid-displacement FE. Journal of Computational and Applied Mathematics, 63(1–3):265–269, November 20, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042796809137.

Kalantari:1997:BFI

[KKZN97] Bahman Kalantari, Iraj Kalantari, and Rahim Zaare-Nahandi. A basic family of iteration functions for polynomial root finding and its characterizations. Journal of Computational and Applied Mathematics, 80(2): 209–226, May 5, 1997. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042797000149.

Kuang:1990:TCF

[KL90] Jiaoxun Kuang and Lianhua Lu. Two classes of finitedifference methods for generalized sine-Gordon equations. Journal of Computational and Applied Mathematics, 31(3):389–396, October 1, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042790900393.

Kim:1998:PLQ

[KL98] Philsu Kim and Sungyun Lee. A piecewise linear quadrature of Cauchy singular integrals. Journal of Computational and Applied Mathematics, 95(1–2):101– 115, August 28, 1998. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042798000752.

Kim:1999:HOA

[KL99] Sun-Chul Kim and June-Yub Lee. A high-order adaptive numerical method for recirculating flows at large Reynolds number. Journal of Computational and Applied Mathematics, 108(1–2):75–86, August 15, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042799001016.

Klatte:1994:SLM

[Kla94] Diethard Klatte. Stable local minimizers in semiinfinite optimization: regularity and second-order conditions. Journal of Computational and Applied Mathematics, 56(1–2):137–157, December 20, 1994. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0377042794903840.

Klein:1999:SFD

[Kle99]  Richard I. Klein.  Star

formation with 3-D adaptive mesh refinement: the collapse and fragmentation of molecular clouds. Journal of Computational and Applied Mathematics, 109(1–2):123–152, September 30, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042799001569.

Kwon:1997:DEH

[KLL97] K. H. Kwon, D. W. Lee, and L. L. Littlejohn. Differential equations having orthogonal polynomial solutions. Journal of Computational and Applied Mathematics, 80(1): 1–16, April 14, 1997. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042796000969.

Klusch:1992:STD

[Klu92] Dieter Klusch. The sampling theorem, Dirichlet series and Hankel transforms. Journal of Computational and Applied Mathematics, 44(3):261–273, December 24, 1992. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279290001E.

Klein:1994:IMM

[KM94a] A. A. B. Klein and G. Mélard.

The information matrix of multiple-input single-output time series models. Journal of Computational and Applied Mathematics, 51(3): 349–356, June 30, 1994. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/037704279200116Q.

Kunkel:1994:CFL

[KM94b] Peter Kunkel and Volker Mehrmann. Canonical forms for linear differentialalgebraic equations with variable coefficients. Journal of Computational and Applied Mathematics, 56 (3):225–251, December 30, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042794900809.

Klein:1995:CFI

[KM95] André Klein and Guy Mélard. Computation of the Fisher information matrix for time series models. Journal of Computational and Applied Mathematics, 64(1–2):57–68, November 30, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042795000062.

Kalnins:1996:NTP

[KM96] E. G. Kalnins and Willard Miller, Jr. A note on tensor products of q-algebra representations and orthogonal polynomials. Journal of Computational and Applied Mathematics, 68(1–2):197– 207, April 22, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042795002529.

Krejic:1999:INM

[KM99] Nataša Krejić and José Mario Martínez. Inexact-Newton methods for semismooth systems of equations with block-angular structure. Journal of Computational and Applied Mathematics, 103(2):239–249, March 31, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042798002581.

Kallrath:1990:LIS

[KN90] Josef Kallrath and Wolfram Neutsch. Lattice integration on the 7-sphere. Journal of Computational and Applied Mathematics, 29(1):9– 14, January 10, 1990. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279090190B.

Krzyz:1999:HAU

[KN99] Jan G. Krzyz and Maria Nowak. Harmonic automorphisms of the unit disk. Journal of Computational and Applied Mathematics, 105(1–2):337–346, May 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042799000199.

Karkkainen:1996:NMN

[KNN96] T. Kärkkäinen, P. Neittaanmäki, and A. Niemistö. Numerical methods for nonlinear inverse problems. Journal of Computational and Applied Mathematics, 74(1–2):231–244, November 5, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279600026X.

Kotakemori:1996:AIM

[KNO96] Hisashi Kotakemori, Hiroshi Niki, and Naotaka Okamoto. Accelerated iterative method for Z-matrices. Journal of Computational and Applied Mathematics, 75(1):87–97, November 12, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042796000611.

Kotakemori:1997:CPI

[KNO97] Hisashi Kotakemori, Hiroshi Niki, and Naotaka Okamoto. Convergence of a preconditioned iterative method for H-matrices. Journal of Computational and Applied Mathematics, 83(1):115–118, September 30, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042797001453.

Kaneko:1997:SIC

[KNP97] Hideaki Kaneko, Richard D.

Noren, and Peter A. Padilla. Superconvergence of the iterated collocation methods for Hammerstein equations. Journal of Computational and Applied Mathematics, 80(2):335–349, May 5, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704279700040X.

Kahaner:1991:EOD

[KNST91] D. K. Kahaner, E. Ng, W. E.

Schiesser, and S. Thompson. Experiments with an ordinary differential equation solver in the parallel solution of method of lines problems on a sharedmemory parallel computer. Journal of Computational and Applied Mathematics, 38(1–3):231–253, December 23, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279190173H.

Khan:1999:CFE

[KO99] Ishtiaq Rasool Khan and Ryoji Ohba. Closed-form expressions for the finite difference approximations of first and higher derivatives based on Taylor series. Journal of Computational and Applied Mathematics, 107(2): 179–193, July 31, 1999. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042799000886.

See corrigendum [KO01].

Khan:2001:CCF

[KO01] Ishtiaq Rasool Khan and Ryoji Ohba. Corrigendum to: “Closed-form expressions for the finite difference approximations of first and higher derivatives based on Taylor series” [J. Comput. Appl. Math. 107 (1999) 179–193]. Journal of Computational and Applied Mathematics, 130(1– 2):385, May 1, 2001. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042700002636.

See [KO99].

Kobza:1995:CLP

[Kob95] Jirí Kobza. Computing local parameters of biquartic interpolatory splines. Journal of Computational and Applied Mathematics, 63(1–3):229–236, November 20, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042795000763.

Koekoek:1993:SDE

[Koe93]  Roelof Koekoek.  The

search for differential equations for certain sets of orthogonal polynomials. Journal of Computational and Applied Mathematics, 49(1–3):111–119, December 31, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279390141W.

Koelink:1996:YAB

[Koe96] H. T. Koelink. Yet another basic analogue of Graf’s addition formula. Journal of Computational and Applied Mathematics, 68(1–2):209– 220, April 22, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279500257X.

Kohler:1993:EPD

[Köh93] Peter Köhler. On the error of parameter-dependent compound quadrature formulas. Journal of Computational and Applied Mathematics, 47(1):47–60, June 28, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        037704279390089T.

Kohler:1994:IEE

[Köh94] Peter Köhler. Intermediate error estimates. Journal of Computational and Applied Mathematics, 50(1–3):349– 360, May 20, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042794903123.

Kolbig:1994:CBP

[Köl94a] K. S. Kölbig. The complete Bell polynomials for certain arguments in terms of Stirling numbers of the first kind. Journal of Computational and Applied Mathematics, 51(1):113– 116, May 30, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042794000107.

Kolbig:1994:TTI

[Köl94b] K. S. Kölbig. On three trigonometric integrals of ln Γ(x) or its derivative. Journal of Computational and Applied Mathematics, 54(1):129–131, September 20, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042794904006.

Kolbig:1995:CCC

[Köl95a]  K. S. Kölbig.  Cheby-

shev coefficients for the Clausen function Cl _{2} ( x ). Journal of Computational and Applied Mathematics, 64(3):295–297, December 20, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042795001506.

Kolbig:1995:TII

[Köl95b] K. S. Kölbig. Two infinite integrals of products of modified Bessel functions and powers of logarithms. Journal of Computational and Applied Mathematics, 62(1): 41–65, August 31, 1995. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042794000750.

Kolbig:1996:IIB

[Köl96a] K. S. Kölbig. An infinite integral of Bessel functions. Journal of Computational and Applied Mathematics, 67(1):181–183, February 20, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042795002383.

Kolbig:1996:PF

[Köl96b]  K. S. Kölbig.  The

polygamma function ψ^{(k)}(x) for x = 1/4 and x = 3/4. Journal of Computational and Applied Mathematics, 75(1):43–46, November 12, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042796000556.

Koornwinder:1993:ZAA

[Koo93] Tom H. Koornwinder. On Zeilberger’s algorithm and its q-analogue. Journal of Computational and Applied Mathematics, 48(1–2):91– 111, October 29, 1993. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0377042793903175.

Koornwinder:1998:INS

[Koo98] Tom H. Koornwinder. Identities of nonterminating series by Zeilberger’s algorithm. Journal of Computational and Applied Mathematics, 99(1–2):449–461, November 16, 1998. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042798001769.

Korman:1994:DTC

[Kor94a] Philip Korman. On the dynamics of two classes of periodic ecological models. Journal of Computational and Applied Mathematics, 52(1–3):267–275, July 20, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042794903603.

Kornfeld:1994:NUA

[Kor94b]  Isaac Kornfeld.  Nonex-

istence of universally accelerating linear summability methods. Journal of Computational and Applied Mathematics, 53(3):309– 321, August 30, 1994. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042794900590.

Korzeniowski:1996:CSF

[Kor96] Andrzej Korzeniowski. On computer simulation of Feynman–Kac path-integrals. Journal of Computational and Applied Mathematics, 66(1–2):333–336, January 31, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042795001700.

Kossaczky:1994:RAL

[Kos94]  Igor Kossaczký.  A re-

cursive approach to local mesh refinement in two and three dimensions. Journal of Computational and Applied Mathematics, 55(3):275–288, November 30, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042794900345.

Kosovichev:1999:IMH

[Kos99]  A. G. Kosovichev.  In-

version methods in helioseismology and solar tomography. Journal of Computational and Applied Mathematics, 109(1–2):1– 39, September 30, 1999. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042799001521.

Karageorghis:1990:EDM

[KP90] Andreas Karageorghis and Timothy N. Phillips. On efficient direct methods for conforming spectral domain decomposition techniques. Journal of Computational and Applied Mathematics, 33(2):141–155, December 21, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042790903646.

Kaminski:1999:ZPI

[KP99] D. Kaminski and R. B.

Paris. On the zeroes of the Pearcey integral. Journal of Computational and Applied Mathematics, 107(1):31–52, July 15, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042799000746.

Kordonis:1997:OSL

[KPP97] I.-G. E. Kordonis, Ch. G.

Philos, and I. K. Purnaras. On the oscillation of some linear difference equations with periodic coefficients. Journal of Computational and Applied Mathematics, 84(2):219–241, October 28, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704279700126X.

Kaliaguine:1996:SCP

[KR96] V. Kaliaguine and A. Ronveaux. On an system of “classical” polynomials of simultaneous orthogonality. Journal of Computational and Applied Mathematics, 67(2):207– 217, March 29, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0377042794001294.

Kuijlaars:1998:ZDD

[KR98] A. B. J. Kuijlaars and E. A. Rakhmanov. Zero distributions for discrete orthogonal polynomials. Journal of Computational and Applied Mathematics, 99(1–2):255–274, November 16, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042798001617.

See corrigendum [KR99].

Kuijlaars:1999:CZD

[KR99] A. B. J. Kuijlaars and E. A. Rakhmanov. Corrigendum to “Zero distributions for discrete orthogonal polynomials” [J. Comput. Appl. Math. 99 (1998) 255–274]. Journal of Computational and Applied Mathematics, 104(2): 213, May 17, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704279900059X.

See [KR98].

Kravanja:1996:BRF

[Kra96]  P. Kravanja.  Book re-

view: Fast algorithms: a multitape Turing machine implementation: Arnold Schönhage, Andreas F. W. Grotefeld and Ekkehart Vetter, B.I. Wissenschaftsverlag, Dudenstraße 6, Postfach 100311, 6800 Mannheim 1, Germany, 1994, 297 pages with 24 illustrations and 5 tables, ISBN 3-411-16891-9, Hardcover, DM 68. Journal of Computational and Applied Mathematics, 66(1–2): N5, January 31, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042796804730.

Kren:1995:SPL

[Kre95a] Jirí Kren. A spatial problem of liquid-shell interaction. Journal of Computational and Applied Mathematics, 63(1–3):301–306, November 20, 1995. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042795000852.

Kress:1995:NSH

[Kre95b] Rainer Kress. On the numerical solution of a hypersingular integral equation in scattering theory. Journal of Computational and Applied Mathematics, 61(3):345– 360, August 21, 1995. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0377042794000737.

See corrigendum [Kre11].

Kress:2011:CNS

[Kre11] Rainer Kress. Corrigendum to “On the numerical solution of a hypersingular integral equation in scattering theory” [J. Comput. Appl. Math. 61 (1995) 345–360]. Journal of Computational and Applied Mathematics, 235(9):3109, March 1, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042710005364.

See [Kre95b].

Krinik:1992:TSS

[Kri92] Alan Krinik. Taylor series solution of the M/M/1 queueing system. Journal of Computational and Applied Mathematics, 44 (3):371–380, December 24, 1992. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        037704279290007K.

Krizek:1995:NET

[Krí95] Michal Krízek. Numerical experience with the threebody problem. Journal of Computational and Applied Mathematics, 63(1–3):403– 409, November 20, 1995.

CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042795000674.

Krzyworzcka:1996:ELC

[Krz96] Sonia Krzyworzcka. Extension of the Lanczos and CGS methods to systems of nonlinear equations. Journal of Computational and Applied Mathematics, 69(1):181– 190, April 30, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042795000321.

Kolbig:1996:HTI

[KS96a] K. S. Kölbig and H. Scherb.

On a Hankel transform integral containing an exponential function and two Laguerre polynomials. Journal of Computational and Applied Mathematics, 71(2): 357–363, July 27, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042795002243.

Kolbig:1996:SII

[KS96b] K. S. Kölbig and W. Strampp.

Some infinite integrals with powers of logarithms and the complete Bell polynomials. Journal of Computational and Applied Mathematics, 69(1):39–47, April 30, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042795000283.

Karaballi:1997:LIQ

[KS97] A. A. Karaballi and S. Sallam. Lacunary interpolation by quartic splines on uniform meshes. Journal of Computational and Applied Mathematics, 80(1):97–104, April 14, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042797000150.

Koepf:1998:ROP

[KS98] Wolfram Koepf and Dieter Schmersau. Representations of orthogonal polynomials. Journal of Computational and Applied Mathematics, 90(1):57–94, April 6, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042798000235.

Ksasov:1999:GSA

[KS99] Boris I. Ksasov and Pairote Sattayatham. GB-splines of arbitrary order. Journal of Computational and Applied Mathematics, 104(1):63–88, April 15, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042798002659.

Khater:1998:ANS

[KSFAH98] A. H. Khater, A. B. Shamardan, M. H. Farag, and A. H. Abel-Hamid. Analytical and numerical solutions of a quasilinear parabolic optimal control problem. Journal of Computational and Applied Mathematics, 95(1–2):29–43, August 28, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042798000661.

Kokologiannaki:1992:CZR

[KSK92] C. G. Kokologiannaki, P. D.

Siafarikas, and C. B. Kouris. On the complex zeros of

H_{μ}(z), J^{′_{μ}(z)}, Jμ^{′′}(z) for real

or complex order. Journal of Computational and Applied Mathematics, 40(3): 337–344, July 24, 1992. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042792901895.

Kitamura:1992:AAS

[KT92] Yuichi Kitamura and Kusano Takaši. Asymptotic analysis of solutions of systems of neutral functional differential equations. Journal of Computational and Applied Mathematics, 41(1–2):23–33, August 20, 1992. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279290234O.

Kaufman:1993:CCR

[KT93] E. H. Kaufman, Jr. and G. D. Taylor. Coupled constrained rational approximation. Journal of Computational and Applied Mathematics, 45(3):343– 346, May 3, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279390052D.

Khiyal:1997:VOV

[KT97a] M. S. H. Khiyal and R. M. Thomas. Variable-order, variable-step methods for second-order initial-value problems. Journal of Computational and Applied Mathematics, 79(2):263– 276, March 17, 1997. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042796001707.

Klinger:1997:PDS

[KT97b] Bernhard Klinger and Robert F. Tichy. Polynomial discrepancy of sequences. Journal of Computational and Applied Mathematics, 84(1):107–117, October 21, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042797001179.

Kuchminska:1999:SPT

[Kuc99] Khrystyna Kuchmins’ka.

Some properties of twodimensional continued fractions. Journal of Computational and Applied Mathematics, 105(1–2):347–353, May 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042799000308.

Kuijlaars:1995:CQM

[Kui95a] A. B. J. Kuijlaars. Chebyshev quadrature for measures with a strong singularity. Journal of Computational and Applied Mathematics, 65(1–3):207–214, December 29, 1995. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042795001107.

Kuijlaars:1995:CTQb

[Kui95b] A. B. J. Kuijlaars. Chebyshevtype quadrature and zeros of Faber polynomials.

Journal of Computational and Applied Mathematics, 62(2):155–179, September 20, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042795001312.

Kuijlaars:1995:CTQa

[Kui95c] Arno Kuijlaars. Chebyshevtype quadrature for Jacobi weight functions. Journal of Computational and Applied Mathematics, 57(1–2):171– 180, February 20, 1995. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0377042793E0243F.

Kulev:1992:UAS

[Kul92] G. K. Kulev. Uniform asymptotic stability in impulsive perturbed systems of differential equations. Journal of Computational and Applied Mathematics, 41(1–2):49–55, August 20, 1992. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279290237R.

Krizkova:1995:DDL

[KV95] Jitka Krízková and Petr Vanek. Domain decomposition like technique with H/h-independent rate of convergence. Journal of Computational and Applied Mathematics, 63(1–3):215– 220, November 20, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042795000747.

Koren:1999:FBL

[KV99] B. Koren and A. C. J. Venis.

A fed back level-set method for moving material-void interfaces. Journal of Computational and Applied Mathematics, 101(1–2):131–152, January 15, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042798002039.

Kijima:1995:WSO

[KvD95] Masaaki Kijima and Erik A.

van Doorn. Weighted sums of orthogonal polynomials with positive zeros. Journal of Computational and Applied Mathematics, 65(1–3):195–206, December 29, 1995. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042795001050.

Kuijt:1999:MPI

[KvD99] Frans Kuijt and Ruud van Damme. Monotonicity preserving interpolatory subdivision schemes. Journal of Computational and Applied Mathematics, 101(1–2):203– 229, January 15, 1999. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042798002209.

Kleinman:1992:MGM

[KvdB92] R. E. Kleinman and P. M. van den Berg. A modified gradient method for two- dimensional problems in tomography. Journal of Computational and Applied Mathematics, 42(1):17–35, September 8, 1992. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279290160Y.

Karlberg:1990:PTA

[KW90] Lennart Karlberg and Hans Wallin. Padé type approximants and orthogonal polynomials for Markov– Stieltjes functions. Journal of Computational and Applied Mathematics, 32(1– 2):153–157, November 26, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
037704279090426Z.

Kanas:1999:CRU

[KW99] Stanisława Kanas and Agnieszka Wiśniowska. Conic regions and k-uniform convexity. Journal of Computational and Applied Mathematics, 105(1–2):327–336, May 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042799000187.

Kwon:1997:SDE

[KY97] K. H. Kwon and G. J. Yoon.

Symmetrizability of differential equations having orthogonal polynomial solutions. Journal of Computational and Applied Mathematics, 83(2):257–268, October 7, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042797001106.

Kwon:1997:CHP

[KYY97] K. H. Kwon, B. H. Yoo, and G. J. Yoon. A characterization of Hermite polynomials. Journal of Computational and Applied Mathematics, 78(2):295–299, February 17, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042796001446.

Kress:1992:NST

[KZ92] Rainer Kress and Axel Zinn.

On the numerical solution of the three-dimensional inverse obstacle scattering problem. Journal of Computational and Applied Mathematics, 42(1):49–61, September 8, 1992. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279290162Q.

Kratzig:1994:SFN

[KZ94] Wilfried B. Krätzig and Jian-Wu Zhang. A simple four-node quadrilateral finite element for plates. Journal of Computational and Applied Mathematics, 50(1–3):361–373, May 20, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042794903131.

Kzaz:1995:CAJ

[Kza95] M. Kzaz. Convergence acceleration of Jacobi–Gauss quadrature formulae for analytic functions with poles. Journal of Computational and Applied Mathematics, 57(1–2):181–192, February 20, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042793E0244G.

Laburta:1997:SAI

[Lab97] M. P. Laburta. Starting algorithms for IRK methods. Journal of Computational and Applied Mathematics, 83(2):269–288, October 7, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704279700112X.

Laburta:1998:CSA

[Lab98]  M. P. Laburta.  Con-

struction of starting algorithms for the RK–Gauss methods. Journal of Computational and Applied Mathematics, 90(2):239– 261, April 17, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704279800003X.

Lacey:1998:DMB

[Lac98] A. A. Lacey. Diffusion models with blow-up. Journal of Computational and Applied Mathematics, 97(1–2):39– 49, September 24, 1998. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042798001058.

Ladopoulos:1997:CAM

[Lad97] E. G. Ladopoulos. Collocationapproximation methods for nonlinear singular integrodifferential equations in Banach spaces. Journal of Computational and Applied Mathematics, 79(2):289– 297, March 17, 1997. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042796001744.

Laforgia:1991:BMB

[Laf91] Andrea Laforgia. Bounds for modified Bessel functions. Journal of Computational and Applied Mathematics, 34(3):263–267, April 26, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
037704279190087Z.

Lahtinen:1990:SPI

[Lah90]  Aatos Lahtinen.  Shape

preserving interpolation by quadratic splines. Journal of Computational and Applied Mathematics, 29(1):15–24, January 10, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042790901912.

Lahtinen:1992:CPS

[Lah92] Aatos Lahtinen. On the choice of parameters in shape-preserving quadratic spline interpolation. Journal of Computational and Applied Mathematics, 39(1): 109–113, February 28, 1992. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
037704279290227O.

Lahtinen:1996:CIV

[Lah96] Aatos Lahtinen. On the choice of the initial value of a quadratic spline in positive interpolation. Journal of Computational and Applied Mathematics, 69(1):13–25, April 30, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042795000240.

Lai:1996:QSF

[Lai96] Ming-Jun Lai. On C^{2} quintic spline functions over triangulations of Powell– Sabin’s type. Journal of Computational and Applied Mathematics, 73(1–2):135– 155, October 5, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042796000404.

Langtry:1995:DCF

[Lan95] T. N. Langtry. The determination of canonical forms for lattice quadrature rules. Journal of Computational and Applied Mathematics, 59(2):129– 143, May 19, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279400025V.

Lang:1998:SPZ

[Lan98]  Wolfdieter Lang.  On

sums of powers of zeros of polynomials. Journal of Computational and Applied Mathematics, 89(2):237– 256, March 16, 1998. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042797002409.

Lange:1999:CRB

[Lan99a] L. J. Lange. Convergence regions with bounded convex complements for continued fractions K(1/b_{n}). Journal of Computational and Applied Mathematics, 105(1– 2):355–366, May 1999. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S037704279900031X.

Langtry:1999:LRM

[Lan99b] T. N. Langtry. Lattice rules of minimal and maximal rank with good figures of merit. Journal of Computational and Applied Mathematics, 112(1–2):147–164, November 30, 1999. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042799002204.

Lapin:1993:ITV

[Lap93] Sergei V. Lapin. Identification of time-varying nonlinear systems using Chebyshev polynomials. Journal of Computational and Applied Mathematics, 49(1–3):121–126, December 31, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279390142X.

Lardy:1994:PKM

[Lar94] L. J. Lardy. The projection Kantorovich method for eigenvalue problems. Journal of Computational and Applied Mathematics, 50(1– 3):375–384, May 20, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
037704279490314X.

Laurie:1991:SIE

[Lau91] Dirk P. Laurie. Solving the inverse eigenvalue problem via the eigenvector matrix. Journal of Computational and Applied Mathematics, 35(1–3):277–289, June 26, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042791902145.

Laurie:1996:PTN

[Lau96] Dirk P. Laurie. Periodizing transformations for numerical integration. Journal of Computational and Applied Mathematics, 66(1–2):337– 344, January 31, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0377042795001964.

Laurie:1999:ARR

[Lau99]  Dirk P. Laurie.  Accu-

rate recovery of recursion coefficients from Gaussian quadrature formulas. Journal of Computational and Applied Mathematics, 112(1–2):165–180, November 30, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042799002289.

Layton:1993:ODS

[Lay93] William Layton. Optimal difference schemes for 2-D transport problems. Journal of Computational and Applied Mathematics, 45(3): 337–341, May 3, 1993. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279390051C.

Li:1990:SHC

[LB90] Z. C. Li and T. D. Bui. The simplified hybrid-combined methods for Laplace’s equation with singularities. Journal of Computational and Applied Mathematics, 29(2):171–193, February 15, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042790903565.

Li:1998:PAS

[LB98] Zi-Cai Li and Z. D. Bai.

Probabilistic analysis on the splitting-shooting method for image transformations. Journal of Computational and Applied Mathematics, 94(2):69–121, August 3, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042798000624.

Levrie:1999:NRB

[LB99] Paul Levrie and Adhemar Bultheel. A note on the relation between two convergence acceleration methods for ordinary continued fractions. Journal of Computational and Applied Mathematics, 101(1–2):167–175, January 15, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042798002179.

Lotstedt:1997:SNN

[LC97] Per Lötstedt and Per Carlbom. Stability and non-normality of the kϵ equations. Journal of Computational and Applied Mathematics, 83(1):11–37, September 30, 1997. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042797000241.

Lecot:1998:PMS

[LC98] Christian Lécot and Ibrahim Coulibaly. A particle method for some parabolic equations. Journal of Computational and Applied Mathematics, 90(1):25–44, April 6, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042798000089.

Li:1999:BPFb

[LC99a] Zi-Cai Li and Chia-Shen Chang. Boundary penalty finite element methods for blending surfaces, III. Superconvergence and stability and examples. Journal of Computational and Applied Mathematics, 110(2):241– 270, October 30, 1999. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042799002319.

Lima:1999:IMS

[LC99b] P. M. Lima and M. P. Carpentier. Iterative methods for a singular boundaryvalue problem. Journal of Computational and Applied Mathematics, 111(1–2):173– 186, November 15, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042799001417.

Cheng:1997:ASM

[lCHcH97a] Xiao liang Cheng, Weimin Han, and Hong ci Huang. Analysis of some mixed elements for the Stokes problem. Journal of Computational and Applied Mathematics, 85(1):19–35, November 6, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042797001209.

Cheng:1997:FEM

[lCHcH97b] Xiao liang Cheng, Weimin Han, and Hong ci Huang. Finite element methods for Timoshenko beam, circular arch and Reissner–Mindlin plate problems. Journal of Computational and Applied Mathematics, 79(2):215– 234, March 17, 1997. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042797001593.

Lakshmikantham:1998:PBS

[LD98] V. Lakshmikantham and Z. Drici. Positivity and boundedness of solutions of impulsive reaction–diffusion equations. Journal of Computational and Applied Mathematics, 88(1):175– 184, February 23, 1998. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042797002100.

Lyness:1999:IMS

[LD99] J. N. Lyness and L. M. Delves. On the implementation of a modified Sag–Szekeres quadrature method. Journal of Computational and Applied Mathematics, 112(1–2):189–200, November 30, 1999. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042799002216.

LeTallec:1991:DDM

[LDV91] P. Le Tallec, Y. H. De Roeck, and M. Vidrascu. Domain decomposition methods for large linearly elliptic three-dimensional problems. Journal of Computational and Applied Mathematics, 34(1):93–117, February 10, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279190150I.

LeFerrand:1993:VOP

[Le 93] Hervé Le Ferrand. Vector orthogonal polynomials and matrix series. Journal of Computational and Applied Mathematics, 45(3): 267–282, May 3, 1993. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279390045D.

LeFerrand:1995:PTA

[Le 95] Hervé Le Ferrand. Padétype approximants of entire functions and acceleration of convergence: Two examples. Journal of Computational and Applied Mathematics, 57(1–2):193–202, February 20, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042793E0245H.

LeRoux:1998:NSF

[Le 98] Marie-Noëlle Le Roux. Numerical solution of fast diffusion or slow diffusion equations. Journal of Computational and Applied Mathematics, 97(1–2):121–136, September 24, 1998. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S037704279800106X.

Lemke:1994:SOM

[Lem94] Mathias Lemke. A special operator method for solving differential equations. Journal of Computational and Applied Mathematics, 53(2):139–161, August 19, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042794900426.

Leopold:1998:EZP

[Leo98a] Elie Leopold. The extremal zeros of a perturbed orthogonal polynomials system. Journal of Computational and Applied Mathematics, 98(1):99–120, October 5, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042798001174.

Leopold:1998:LZO

[Leo98b]  Elie Leopold.  Location

of the zeros of orthogonal polynomials with an automatic procedure I. Journal of Computational and Applied Mathematics, 92(1):1– 14, May 20, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042798000429.

Leon:1999:CAS

[Leó99] N. Luis E. León. Convergence of approximate solutions of the Cauchy problem for a 2 × 2 nonstrictly hyperbolic system of conservation laws. Journal of Computational and Applied Mathematics, 103(1):139– 144, March 15, 1999. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042798002477.

Lepp:1994:PDM

[Lep94]  Riho Lepp.  Projection

and discretization methods in stochastic programming. Journal of Computational and Applied Mathematics, 56(1–2):55–64, December 20, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042794903794.

Letessier:1993:CRA

[Let93] Jean Letessier. On corecursive associated Laguerre polynomials. Journal of Computational and Applied Mathematics, 49(1– 3):127–136, December 31, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        037704279390143Y.

Letessier:1995:CRA

[Let95a] Jean Letessier. Co-recursive associated Jacobi polynomials. Journal of Computational and Applied Mathematics, 57(1–2):203–213, February 20, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042793E0246I.

Lether:1995:GLA

[Let95b]  F. G. Lether.  Gauss–

Legendre approximations for the Hubbell rectangularsource integral. Journal of Computational and Applied Mathematics, 57(3):393– 402, February 28, 1995. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0377042793E0210D.

Lether:1996:RAF

[Let96] Frank G. Lether. Rational approximation formulas for computing the positive zeros of J_{0}(x). Journal of Computational and Applied Mathematics, 67(1):167– 172, February 20, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042795002197.

Lether:1998:SRQ

[Let98] Frank G. Lether. Shifted rectangular quadrature rule approximations to Dawson’s integral F ( x ). Journal of Computational and Applied Mathematics, 92(2):97–102, June 26, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042798000193.

Letessier:1999:SRC

[Let99] Jean Letessier. Some results on co-recursive associated Meixner and Charlier polynomials. Journal of Computational and Applied Mathematics, 103(2):323– 335, March 31, 1999. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042798002696.

Levrie:1990:BRF

[Lev90a] P. Levrie. Book review:

Factorization and primality testing: David M. Bressoud: Springer-Verlag, Undergraduate texts in mathematics, 1989, 237 pp., Price: DM 98 (Hardcover), ISBN 3-540-97040-1. Journal of Computational and Applied Mathematics, 32(1–2):N6– N7, November 26, 1990. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042790904445.

Levrie:1990:RBG

[Lev90b] Paul Levrie. On the relationship between generalised continued fractions and G-continued fractions. Journal of Computational and Applied Mathematics, 32(1–2):159–167, November 26, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042790904272.

Levrie:1991:BRI

[Lev91] P. Levrie. Book review: Introduction to Calculus and Analysis, Volume I and II: R. Courant and F. John. Springer Verlag, New York. Vol. I: 1989, 2nd print (1st ed. 1965), 660 pp., 240 figs. Hardcover. Price: DM 96,- ISBN 3-540-97151-3. Vol. II: 1989, 2nd print (1st ed. 1974), 955 pp., 120 figs. Hardcover. Price: DM 126,- ISBN 3-540-97152-1. Journal of Computational and Applied Mathematics, 34(2): N3–N4, April 4, 1991. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279190051K.

Levrie:1994:SIC

[Lev94] Paul Levrie. Some identities of G-continued fractions and generalized continued fractions. Journal of Computational and Applied Mathematics, 51(1):85–97, May 30, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279200080S.

Levin:1996:FIR

[Lev96] David Levin. Fast integration of rapidly oscillatory functions. Journal of Computational and Applied Mathematics, 67(1):95–101, February 20, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042794001189.

Levin:1997:ACM

[Lev97] David Levin. Analysis of a collocation method for integrating rapidly oscillatory functions. Journal of Computational and Applied Mathematics, 78(1):131– 138, February 3, 1997. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042796001379.

Levin:1999:SIR

[Lev99] David Levin. Stable integration rules with scattered integration points. Journal of Computational and Applied Mathematics, 112(1–2):181–187, November 30, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042799002186.

Lewanowicz:1991:EBF

[Lew91] Stanisław Lewanowicz. Evaluation of Bessel function integrals with algebraic singularities. Journal of Computational and Applied Mathematics, 37(1–3):101–112, November 18, 1991. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0377042791901106.

Lewanowicz:1992:QCR

[Lew92] Stanisław Lewanowicz. Quick construction of recurrence relations for the Jacobi coefficients. Journal of Computational and Applied Mathematics, 43(3):355–372, December 2, 1992. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279290021O.

Lewanowicz:1993:RAJ

[Lew93] Stanisław Lewanowicz. Results on the associated Jacobi and Gegenbauer polynomials. Journal of Computational and Applied Mathematics, 49(1–3):137–143, December 31, 1993. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279390144Z.

Lewanowicz:1995:RAC

[Lew95] Stanisław Lewanowicz. Results on the associated classical orthogonal polynomials. Journal of Computational and Applied Mathematics, 65(1–3):215–231, December 29, 1995. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042795001123.

Lewanowicz:1996:RRC

[Lew96a] Stanisław Lewanowicz. Recurrence relations for the connection coefficients of orthogonal polynomials of a discrete variable. Journal of Computational and Applied Mathematics, 76(1–2):213– 229, December 17, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042796001045.

Lewanowicz:1996:SOR

[Lew96b] Stanisław Lewanowicz. Secondorder recurrence relation for the linearization coefficients of the classical orthogonal polynomials. Journal of Computational and Applied Mathematics, 69(1):159– 170, April 30, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279500033X.

Lewanowicz:1997:GWS

[Lew97a] Stanisław Lewanowicz. Generalized Watson’s summation formula for _{3} F _{2} (1). Journal of Computational and Applied Mathematics, 86(2):375–386, December 10, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042797001702.

Lewanowicz:1997:FOD

[Lew97b] Stanisław Lewanowicz. On the fourth-order difference equation for the associated Meixner polynomials. Journal of Computational and Applied Mathematics, 80(2): 351–358, May 5, 1997. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S037704279782986X.

Lewanowicz:1998:RRC

[Lew98] Stanisław Lewanowicz. Recurrence relations for the connection coefficients of orthogonal polynomials of a discrete variable on the lattice x(s) = q^{2s}. Journal of Computational and Applied Mathematics, 99(1–2):275–286, November 16, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042798001629.

Lima:1995:CAB

[LG95] Pedro M. Lima and Mário M.

Graça. Convergence acceleration for boundary value problems with singularities using the e-algorithm. Journal of Computational and Applied Mathematics, 61(2): 139–164, July 31, 1995. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042794000659.

Lavoie:1996:GWT

[LGR96] J. L. Lavoie, F. Grondin, and A. K. Rathie. Generalizations of Whipple’s theorem on the sum of a _{3} F _{2}. Journal of Computational and Applied Mathematics, 72(2):293–300, August 13, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042795002790.

Laurent-Gengoux:1990:NPC

[LGT90] Pascal Laurent-Gengoux and Denis Trystram. A new presentation of the conjugate direction algorithm. Journal of Computational and Applied Mathematics, 32(3):417–422, December 10, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042790900463.

Li:1992:AMS

[Li92] Xiezhang Li. An adaptive method for solving nonsymmetric linear systems involving applications of SC-PACK. Journal of Computational and Applied Mathematics, 44(3):351–370, December 24, 1992. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279290006J.

Li:1996:GMF

[Li96]  Jia-Liang Li.  On the

growth of meromorphic functions defined by quasiregular c-fractions. Journal of Computational and Applied Mathematics, 69(2): 367–377, May 23, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042795000437.

Li:1997:PCR

[Li97] Zi-Cai Li. Penalty combinations of the Ritz– Galerkin and finite difference methods for singularity problems. Journal of Computational and Applied Mathematics, 81(1):1– 17, June 1, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042796001483.

Li:1999:ROR

[Li99a]  Xin Li.  Regularity of

orthogonal rational functions with poles on the unit circle. Journal of Computational and Applied Mathematics, 105(1–2):371– 383, May 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704279900045X.

Li:1999:ACS

[Li99b]  Zi-Cai Li.  Advanced

combinations of splittingshooting-integrating methods for digital image transformations. Journal of Computational and Applied Mathematics, 107(2):147– 177, July 31, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042799000862.

Li:1999:BPFa

[Li99c]  Zi-Cai Li.  Boundary

penalty finite element methods for blending surfaces

— II: Biharmonic equations. Journal of Computational and Applied Mathematics, 110(1):155–176, October 15, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042799002083.

Lima:1994:REB

[Lim94] Pedro M. Lima. Richardson extrapolation in boundary value problems for differential equations with nonregular right-hand side. Journal of Computational and Applied Mathematics, 50(1–3): 385–400, May 20, 1994. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042794903158.

Lima:1996:NMA

[Lim96] Pedro M. Lima. Numerical methods and asymptotic error expansions for the Emden–Fowler equations. Journal of Computational and Applied Mathematics, 70(2):245–266, June 28, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042795002030.

Lindlbauer:1998:RCL

[Lin98] Marc Lindlbauer. On the rate of convergence of the laws of Markov chains associated with orthogonal polynomials. Journal of Computational and Applied Mathematics, 99(1–2):287–297, November 16, 1998. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042798001630.

Linss:1999:UDS

[Lin99] Torsten Linß. An upwind difference scheme on a novel Shishkin-type mesh for a linear convection–diffusion problem. Journal of Computational and Applied Mathematics, 110(1):93–104, October 15, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042799001983.

Lioen:1996:DAF

[Lio96] Walter M. Lioen. On the diagonal approximation of full matrices. Journal of Computational and Applied Mathematics, 75(1):35–42, November 12, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042796000544.

Lissens:1995:ESA

[Lis95] Gilbert Lissens. Electricity systems analyses with EPLAN–PULSE, a midterm model for the evaluation of electricity production. Journal of Computational and Applied Mathematics, 64(1–2):21–40, November 30, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042795000046.

Litvinchev:1992:IAM

[Lit92] I. S. Litvinchev. Iterative aggregation method for constrained optimal control problems. Journal of Computational and Applied Mathematics, 39(3): 315–328, May 8, 1992. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279290207E.

Litvinchev:1995:TLD

[Lit95]  I. S. Litvinchev.  A

two-level decompositionaggregation approach for large-scale optimal control problems. Journal of Computational and Applied Mathematics, 61(2): 117–137, July 31, 1995. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042794000626.

Liu:1995:ASR

[Liu95] Changwen Liu. An acceleration scheme for row projection methods. Journal of Computational and Applied Mathematics, 57(3):363– 391, February 28, 1995. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042793E02095.

Liu:1996:MDD

[Liu96] Yunkang Liu. On the θmethod for delay differential equations with infinite lag. Journal of Computational and Applied Mathematics, 71(2):177– 190, July 27, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042795002227.

Liu:1997:LFE

[Liu97]  Bo Liu.  On the lin-

ear finite element approximation for two-point boundary value problems. Journal of Computational and Applied Mathematics, 85(2):215–223, November 12, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042797000629.

Liu:1999:NSH

[Liu99] Yunkang Liu. Numerical solution of the heat equation with nonlocal boundary conditions. Journal of Computational and Applied Mathematics, 110(1):115–127, October 15, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042799002009.

Long:1995:MHA

[LJQ95] Peng Long, Wang Jinliang, and Zhu Qiding. Methods with high accuracy for finite element probability computing. Journal of Computational and Applied Mathematics, 59(2): 181–189, May 19, 1995. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279400027X.

Liang:1991:BOI

[LL91] Xue-Zhang Liang and Luo-Qing Li. On bivariate osculatory interpolation. Journal of Computational and Applied Mathematics, 38(1–3):271–282, December 23, 1991. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279190176K.

Levin:1998:BOP

[LL98a] A. L. Levin and D. S. Lubinsky. Bounds for orthogonal polynomials for exponential weights. Journal of Computational and Applied Mathematics, 99(1–2):475–490, November 16, 1998. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042798001782.

Liang:1998:ICK

[LL98b] Xue-Zhang Liang and Chun-Mei Lü. On the integral convergence of Kergin interpolation on the disk. Journal of Computational and Applied Mathematics, 95(1–2): 45–63, August 28, 1998. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042798000673.

Liu:1999:PES

[LL99a] X. Q. Liu and C. W. Li. Product expansion for stochastic jump diffusions and its application to numerical approximation. Journal of Computational and Applied Mathematics, 108(1–2):1–17, August 15, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042799000953.

Ly:1999:NGU

[LL99b] Le Ngoc Ly and Phu Luong. Numerical grids used in a coastal ocean model with breaking wave effects. Journal of Computational and Applied Mathematics, 103(1):125–137, March 15, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042798002465.

Liu:1991:MAM

[LLM91] C. Liu, Z. Liu, and S. Mc-Cormick. Multilevel adaptive methods for incompressible flow in grooved channels. Journal of Computational and Applied Mathematics, 38(1–3):283– 295, December 23, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
037704279190177L.

Lai:1997:CGM

[LLP97] Yu-Ling Lai, Wen-Wei Lin, and Dan’l Pierce. Conjugate gradient and minimal residual methods for solving symmetric indefinite systems. Journal of Computational and Applied Mathematics, 84(2):243–256, October 28, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042797001271.

Li:1999:GSC

[LLY99] Zi-Cai Li, Qun Lin, and Ning-Ning Yan. Global superconvergence in combinations of Ritz–Galerkin– FEM for singularity problems. Journal of Computational and Applied Mathematics, 106(2):325–344, June 30, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042799000795.

Lynch:1990:DHO

[LM90] R. Lynch and H. A. Mavromatis. N-dimensional harmonic oscillator yields monotonic series for the mathematical constant π. Journal of Computational and Applied Mathematics, 30(2):127–137, May 28, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
037704279090021Q.

Liu:1992:IAT

[LM92] Fawang Liu and John J. H. Miller. Inverse average type tetrahedral finite-element schemes for the stationary semiconductor device equations. Journal of Computational and Applied Mathematics, 44(1):77–94, December 9, 1992. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279290053Z.

Laforgia:1993:AMR

[LM93a] Andrea Laforgia and Maria Luisa Mathis. Additional monotonicity results for the zeros of Bessel functions. Journal of Computational and Applied Mathematics, 47(1): 135–139, June 28, 1993. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279390095S.

Lubinsky:1993:MCL

[LM93b] D. S. Lubinsky and Thandwa Mthembu. Mean convergence of Lagrange interpolation for Erdős weights. Journal of Computational and Applied Mathematics, 47(3):369–390, September 30, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279390063H.

Lazer:1994:P

[LM94a] Alan Lazer and Joe McKenna.

Preface. Journal of Computational and Applied Mathematics, 52(1–3):1–2, July 20, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042794903441.

Lazer:1994:NPF

[LM94b] Alan C. Lazer and P. J.

McKenna. Nonlinear periodic flexing in a floating beam. Journal of Computational and Applied Mathematics, 52(1–3):287–303, July 20, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279490362X.

Lupo:1994:MSH

[LM94c] D. Lupo and A. M. Micheletti.

Multiple solutions of Hamiltonian systems via limit relative category. Journal of Computational and Applied Mathematics, 52(1–3):325– 335, July 20, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042794903646.

Lubinsky:1995:FQS

[LM95] D. S. Lubinsky and D. M. Matjila. Full quadrature sums for p-th powers of polynomials with Freud weights. Journal of Computational and Applied Mathematics, 60(3): 285–296, June 30, 1995. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042794000453.

Lee:1997:MFE

[LM97] Miyoung Lee and Fabio A. Milner. Mixed finite element methods for nonlinear elliptic problems: the h- p version. Journal of Computational and Applied Mathematics, 85(2):239– 261, November 12, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042797001210.

Lopez:1999:GSM

[LMF99] David J. López, Pablo Martín, and José M. Farto. Generalization of the Störmer method for perturbed oscillators without explicit first derivatives. Journal of Computational and Applied Mathematics, 111(1–2):123– 132, November 15, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042799001363.

Llovet:1993:LRS

[LMJ93] J. Llovet, R. Martínez, and J. A. Jaén. Linear recurring sequences for computing the resultant of multivariate polynomials. Journal of Computational and Applied Mathematics, 49(1–3):145–152, December 31, 1993. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042793901452.

Lamour:1994:SBS

[LMM94] R. Lamour, R. März, and R. M. M. Mattheij. On the stability behaviour of systems obtained by indexreduction. Journal of Computational and Applied Mathematics, 56(3):305– 319, December 30, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        037704279490085X.

Lopez:1997:VSS

[LMP97] L. Lopez, C. Mastroserio, and T. Politi. Variable stepsize techniques in continuous Runge–Kutta methods for isospectral dynamical systems. Journal of Computational and Applied Mathematics, 82(1–2):261–278, September 15, 1997. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042797000484.

Lopez-Marcos:1996:GLM

[LMSSD96] M. A. López-Marcos, J. M. Sanz-Serna, and J. C. Díaz. Are Gauss–Legendre methods useful in molecular dynamics? Journal of Computational and Applied Mathematics, 67(1):173– 179, February 20, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0377042795001514.

Lehn:1990:P

[LN90] J. Lehn and H. Neunzert. Preface. Journal of Computational and Applied Mathematics, 31(1):1, July 24, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279090329X.

Lubuma:1995:DPP

[LN95] Jean M.-S. Lubuma and Serge Nicaise. Dirichlet problems in polyhedral domains II: Approximation by FEM and BEM. Journal of Computational and Applied Mathematics, 61(1):13–27, July 20, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279400050B.

Liz:1996:BVP

[LN96] E. Liz and Juan J. Nieto.

Boundary value problems for second order integrodifferential equations of Fredholm type. Journal of Computational and Applied Mathematics, 72(2):215– 225, August 13, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042795002731.

Lubuma:1999:FEM

[LN99] Jean M.-S. Lubuma and Serge Nicaise. Finite element method for elliptic problems with edge singularities. Journal of Computational and Applied Mathematics, 106(1):145– 168, June 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704279900062X.

Loe:1996:SPS

[Loe96] K. F. Loe. A sinusoidal polynomial spline and its Bézier blended interpolant. Journal of Computational and Applied Mathematics, 71(2): 383–393, July 27, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0377042795002413.

Liniger:1992:SOS

[LOH92] Werner Liniger, Farouk Odeh, and Veikko Hara. A second-order sparse factorization method for Poisson’s equation with mixed boundary conditions. Journal of Computational and Applied Mathematics, 44 (2):201–218, December 17, 1992. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        037704279290011L.

Lopez:1990:OSC

[Lop90]  Luciano Lopez.  One-

step collocation methods for differential-algebraic systems of index 1. Journal of Computational and Applied Mathematics, 29(2):145– 159, February 15, 1990. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042790903543.

Lopez:1998:SSC

[Lóp98] JoséL. López. Several series containing gamma and polygamma functions. Journal of Computational and Applied Mathematics, 90(1): 15–23, April 6, 1998. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042798000077.

Lopez:1999:AEI

[Lóp99]  JoséL. López.  Asymp-

totic expansions of integrals: the term-by-term integration method. Journal of Computational and Applied Mathematics, 102(2):181– 194, February 26, 1999. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042798002167.

Lorentzen:1990:CC

[Lor90]  Lisa Lorentzen.  Com-

positions of contractions. Journal of Computational and Applied Mathematics, 32(1–2):169–178, November 26, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042790904283.

Lorentzen:1992:BPT

[Lor92] Lisa Lorentzen. Bestness of the parabola theorem for continued fractions. Journal of Computational and Applied Mathematics, 40(3): 297–304, July 24, 1992. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/037704279290185Z.

Lorentzen:1993:NSC

[Lor93] Lisa Lorentzen. A note on separate convergence for continued fractions. Journal of Computational and Applied Mathematics, 47(1): 123–133, June 28, 1993. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/037704279390094R.

Lorentzen:1995:CQI

[Lor95] Lisa Lorentzen. A convergence question inspired by Stieltjes and by value sets in continued fraction theory. Journal of Computational and Applied Mathematics, 65(1–3):233–251, December 29, 1995. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042795001131.

Lovisek:1995:OCC

[Lov95] Ján Lovísek. Optimal control in coefficients of boundary value problems with a unilateral inner obstacle. Journal of Computational and Applied Mathematics, 63(1–3):283–290, November 20, 1995. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042795000836.

Lozier:1996:SNS

[Loz96] Daniel W. Lozier. Software needs in special functions. Journal of Computational and Applied Mathematics, 66(1–2):345–358, January 31, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042795001816.

Lloyd:1992:CCC

[LP92] N. G. Lloyd and J. M.

Pearson. Computing centre conditions for certain cubic systems. Journal of Computational and Applied Mathematics, 40(3): 323–336, July 24, 1992. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042792901884.

Lakshmikantham:1994:PSN

[LP94] V. Lakshmikantham and Nikolaos S. Papageorgiou. Periodic solutions of nonlinear evolution inclusions. Journal of Computational and Applied Mathematics, 52(1–3):277–286, July 20, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042794903611.

Liz:1998:ASN

[LP98] Eduardo Liz and Rodrigo L. Pouso. Approximation of solutions for nonlinear periodic boundary value problems with discontinuous upper and lower solutions. Journal of Computational and Applied Mathematics, 95(1–2):127–138, August 28, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042798000806.

Lin:1992:CCF

[LR92a] Qun Lin and J. G. Rokne. Circular centered forms for rational functions in several complex variables. Journal of Computational and Applied Mathematics, 41(3):347–357, August 28, 1992. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279290141J.

Luczak:1992:CHD

[LR92b] Tomasz Luczak and Andrzej Ruciński. Convex hulls of dense balanced graphs. Journal of Computational and Applied Mathematics, 41(1–2):205–213, August 20, 1992. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042792902502.

Liu:1993:LSS

[LR93] Lixin Liu and Robert D.

Russell. Linear system solvers for boundary value ODEs. Journal of Computational and Applied Mathematics, 45(1–2):103–117, April 8, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279390268G.

Lambert:1994:HPH

[LR94] Dominique Lambert and André Ronveaux. Hurwitz problem, harmonic morphisms and generalized Hadamard matrices. Journal of Computational and Applied Mathematics, 54(3):273–283, October 20, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279490250X.

Lin:1995:MBR

[LR95] Qun Lin and J. G. Rokne. Methods for bounding the range of a polynomial. Journal of Computational and Applied Mathematics, 58(2): 193–199, March 30, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042793E0270V.

Lin:1996:CSS

[LR96a] Qun Lin and J. G. Rokne. A circular splitting search algorithm for systems of complex equations. Journal of Computational and Applied Mathematics, 75 (1):119–129, November 12, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042796000696.

Lowe:1996:DCE

[LR96b] Bruce D. Lowe and William Rundell. The determination of a coefficient in an elliptic equation from average flux data. Journal of Computational and Applied Mathematics, 70(1):

173–187, June 14, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042795001557.

Luzyanina:1996:NSA

[LR96c] Tatyana Luzyanina and Dirk Roose. Numerical stability analysis and computation of Hopf bifurcation points for delay differential equations. Journal of Computational and Applied Mathematics, 72(2):379– 392, August 13, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042796000088.

Letessier:1996:FOD

[LRV96] Jean Letessier, André Ronveaux, and Galliano Valent. Fourth-order difference equation for the associated Meixner and Charlier polynomials. Journal of Computational and Applied Mathematics, 71(2): 331–341, July 27, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042796871639.

Lucas:1995:EIIa

[LS95] S. K. Lucas and H. A.

Stone. Evaluating infinite integrals involving Bessel functions of arbitrary order. Journal of Computational and Applied Mathematics, 64(3):217–231, December 20, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042795001425.

Liu:1996:HBP

[LS96] Xiaoyu Liu and Larry L. Schumaker. Hybrid Bézier patches on sphere-like surfaces. Journal of Computational and Applied Mathematics, 73(1–2):157–172, October 5, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042796000416.

Lewandowski:1998:UBL

[LS98] Zdzisław Lewandowski and Jan Szynal. An upper bound for the Laguerre polynomials. Journal of Computational and Applied Mathematics, 99(1–2):529–533, November 16, 1998. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042798001812.

Lewandowski:1999:LPB

[LS99] Zdzisław Lewandowski and Jan Szynal. The Landau problem for bounded nonvanishing functions. Journal of Computational and Applied Mathematics, 105(1–2):367– 369, May 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042799000205.

Lobanov:1990:CCW

[LSSZ90] Yu. Yu. Lobanov, R. R.

Shahbagian, O. V. Sidorova, and E. P. Zhidkov. Computation of conditional Wiener integrals by the composite approximation formulas with weight. Journal of Computational and Applied Mathematics, 29(1):51–60, January 10, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042790901945.

Lee:1994:HBP

[LST94] Yoon-Mee Lee, Renate Schaaf, and Russell C. Thompson. A Hopf bifurcation in a parabolic free boundary problem. Journal of Computational and Applied Mathematics, 52(1–3): 305–324, July 20, 1994. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042794903638.

LeMehaute:1996:MSD

[LST96] Alain Le Méhauté, Larry L. Schumaker, and Leonardo Traversoni. Multivariate scattered data fitting. Journal of Computational and Applied Mathematics, 73(1–2):1–4, October 5, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042796000325.

Louchard:1994:RWH

[LSTZ94] G. Louchard, R. Schott, M. Tolley, and P. Zimmermann. Random walks, heat equation and distributed algorithms. Journal of Computational and Applied Mathematics, 53(2):243– 274, August 19, 1994. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0377042794900485.

Lobanov:1996:CMF

[LSZ96] Yu. Yu. Lobanov, R. R. Shahbagian, and E. P. Zhidkov. Computation of multiple functional integrals in quantum physics. Journal of Computational and Applied Mathematics, 70(1): 145–160, June 14, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042795001387.

Lin:1993:FDA

[LT93] Yanping Lin and R. J.

Tait. Finite-difference approximations for a class of nonlocal parabolic boundary value problems. Journal of Computational and Applied Mathematics, 47 (3):335–350, September 30, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
037704279390061F.

Lu:1991:NSM

[Lu91] Lianhua Lu. Numerical stability of the θ-methods for systems of differential equations with several delay terms. Journal of Computational and Applied Mathematics, 34(3):291– 304, April 26, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042791900907.

Lu:1995:MCD

[Lu95] Hao Lu. Matrix compensation and diagonal compensation. Journal of Computational and Applied Mathematics, 63(1–3):237–244, November 20, 1995. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0377042795000771.

Lorch:1996:BPU

[LU96] Lee Lorch and Riccardo Uberti. “Best possible” upper bounds for the first positive zeros of Bessel functions — the finite part. Journal of Computational and Applied Mathematics, 75(2):249–258, November 28, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042796000659.

Lubuma:1993:EEP

[Lub93] Jean Mbaro-Saman Lubuma. Error estimates in projective solutions of the Radon equation. Journal of Computational and Applied Mathematics, 45(3): 309–319, May 3, 1993. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/037704279390048G.

Lucas:1995:EIIb

[Luc95]  S. K. Lucas.  Evaluat-

ing infinite integrals involving products of Bessel functions of arbitrary order. Journal of Computational and Applied Mathematics, 64(3):269–282, December 20, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042795001433.

Lui:1998:KMS

[Lui98] S. H. Lui. Kron’s method for symmetric eigenvalue problems. Journal of Computational and Applied Mathematics, 98(1):35–48, October 5, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042798001101.

Lanser:1999:AOS

[LV99] D. Lanser and J. G. Verwer. Analysis of operator splitting for advection– diffusion-reaction problems from air pollution modelling. Journal of Computational and Applied Mathematics, 111(1–2):201–216, November 15, 1999. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042799001430.

Lether:1991:NCV

[LW91] F. G. Lether and P. R.

Wenston. The numerical computation of the Voigt function by a corrected midpoint quadrature rule for (−∞, ∞). Journal of Computational and Applied Mathematics, 34(1):75–92, February 10, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279190149E.

Lether:1995:MAZ

[LW95] F. G. Lether and P. R. Wenston. Minimax approximations to the zeros of P_{n}(x) and Gauss– Legendre quadrature. Journal of Computational and Applied Mathematics, 59(2): 245–252, May 19, 1995. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0377042794000305.

Lai:1996:MLB

[LW96a] Ming-Jun Lai and Paul Wenston. On multi-level bases for elliptic boundary value problems. Journal of Computational and Applied Mathematics, 71(1):95–113, July 10, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042795002146.

Lang:1996:BPU

[LW96b] T. Lang and R. Wong. “Best possible” upper bounds for the first two positive zeros of the Bessel function J_{v}(x): the infinite case. Journal of Computational and Applied Mathematics, 71(2): 311–329, July 27, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042795002200.

Li:1999:FVM

[LW99] Zi-Cai Li and Song Wang.

The finite volume method and application in combinations. Journal of Computational and Applied Mathematics, 106(1):21– 53, June 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042799000515.

Li:1998:TNA

[LWZ98] Zhilin Li, Deshen Wang, and Jun Zou. Theoretical and numerical analysis on a thermo-elastic system with discontinuities. Journal of Computational and Applied Mathematics, 92(1):37–58, May 20, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042798000442.

Xu:1997:FEG

[lX97] Yu lin Xu. Fast evaluation of Gaunt coefficients: recursive approach. Journal of Computational and Applied Mathematics, 85(1):53–65, November 6, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042797001283.

Lin:1995:IIA

[LYFW95] Chih-Jen Lin, Eugene K. Yang, Shu-Cherng Fang, and Soon-Yi Wu. Implementation of an inexact approach to solving linear semi-infinite programming problems. Journal of Computational and Applied Mathematics, 61(1):87–103, July 20, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042794000512.

Lynge:1992:CMP

[Lyn92] Walter C. Lynge. A combined method for polynomial roots using modified Newton–Raphson with minimum searching. Journal of Computational and Applied Mathematics, 42(3):339– 355, October 30, 1992. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/037704279290084B.

monga-ade:1992:LEB

[ma92] Magolu monga ade. Lower eigenvalue bounds for singular pencils of matrices. Journal of Computational and Applied Mathematics, 39(3): 329–351, May 8, 1992. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279290208F.

Miller:1998:DHZ

[MA98] Jeff Miller and Victor S.

Adamchik. Derivatives of the Hurwitz zeta function for rational arguments. Journal of Computational and Applied Mathematics, 100(2):201–206, December 21, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042798001939.

Maas:1994:SAE

[Maa94] Leo R. M. Maas. On the surface area of an ellipsoid and related integrals of elliptic integrals. Journal of Computational and Applied Mathematics, 51(2): 237–249, June 20, 1994. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279200009X.

MacLeod:1994:CIA

[Mac94] Allan J. MacLeod. Computation of inhomogeneous Airy functions. Journal of Computational and Applied Mathematics, 53(1): 109–116, July 29, 1994. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0377042794901961.

MacGregor:1999:FCT

[Mac99] Thomas H. MacGregor. Fractional Cauchy transforms. Journal of Computational and Applied Mathematics, 105(1–2):93–108, May 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042799000229.

Magnus:1995:PTD

[Mag95a] Alphonse P. Magnus. Painlevétype differential equations for the recurrence coefficients of semi-classical orthogonal polynomials. Journal of Computational and Applied Mathematics, 57(1–2):215–237, February 20, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042793E0247J.

Magnus:1995:SNL

[Mag95b] Alphonse P. Magnus. Special nonuniform lattice ( snul ) orthogonal polynomials on discrete dense sets of points. Journal of Computational and Applied Mathematics, 65(1–3):253–265, December 29, 1995. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279500114X.

Mahmoud:1992:DRP

[Mah92] Hosam M. Mahmoud. Distances in random planeoriented recursive trees. Journal of Computational and Applied Mathematics, 41(1–2):237–245, August 20, 1992. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279290252S.

Maier:1993:PTD

[Mai93] Robert S. Maier. Phasetype distributions and the structure of finite Markov chains. Journal of Computational and Applied Mathematics, 46(3):449–453, June 22, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
037704279390040I.

Makrelov:1992:MCM

[Mak92] I. Makrelov. On a modification of Chebyshev’s method. Journal of Computational and Applied Mathematics, 41(3):373–375, August 28, 1992. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279290143L.

Maleknejad:1990:NET

[Mal90] K. Maleknejad. Numerical experimentation with time-series methods for convolution integral equation. Journal of Computational and Applied Mathematics, 31(2):211–225, August 31, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279090164U.

Manni:1996:CHI

[Man96] Carla Manni. C ^{1} comonotone Hermite interpolation via parametric cubics. Journal of Computational and Applied Mathematics, 69(1): 143–157, April 30, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042794000263.

Martinez:1991:QNM

[Mar91a] JoséMario Martínez. Quasi-Newton methods for solving underdetermined nonlinear simultaneous equations. Journal of Computational and Applied Mathematics, 34(2):171–190, April 4, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
037704279190040Q.

Marzulli:1991:GEE

[Mar91b] Pietro Marzulli. Global error estimates for the standard parallel shooting method. Journal of Computational and Applied Mathematics, 34(2):233– 241, April 4, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279190045L.

Marcellan:1993:OPS

[MAR93a] F. Marcellán, M. Alfaro, and M. L. Rezola. Orthogonal polynomials on Sobolev spaces: old and new directions. Journal of Computational and Applied Mathematics, 48(1–2):113–131, October 29, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042793903186.

Maroni:1993:VAC

[Mar93b]  P. Maroni.  Variations

around classical orthogonal polynomials. Connected problems. Journal of Computational and Applied Mathematics, 48(1–2):133– 155, October 29, 1993. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0377042793903197.

Marek:1995:IPO

[Mar95a] I. Marek. An introduction of Professor Owe Axelsson and Professor Milos Zlámal. Journal of Computational and Applied Mathematics, 63(1–3):5–8, November 20, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042795900470.

Maroni:1995:IRB

[Mar95b] P. Maroni. An integral representation for the Bessel form. Journal of Computational and Applied Mathematics, 57(1–2):251–260, February 20, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042793E0249L.

Martinez:1995:ETS

[Mar95c] JoséMario Martínez. An extension of the theory of secant preconditioners. Journal of Computational and Applied Mathematics, 60(1–2):115–125, June 20, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279400087H.

Marin:1996:SBT

[Mar96]  M. Marin.  Some ba-

sic theorems in elastostatics of micropolar materials with voids. Journal of Computational and Applied Mathematics, 70(1): 115–126, June 14, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042795001379.

Marin:1997:WSE

[Mar97] Marin Marin. On weak solutions in elasticity of dipolar bodies with voids. Journal of Computational and Applied Mathematics, 82(1–2):291–297, September 15, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042797000472.

Masani:1992:MTAa

[Mas92a] P. R. Masani. The measuretheoretic aspects of entropy, part I. Journal of Computational and Applied Mathematics, 40(2): 215–232, June 24, 1992. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0377042792901079.

Masani:1992:MTAb

[Mas92b] P. R. Masani. The measuretheoretic aspects of entropy, part II. Journal of Computational and Applied Mathematics, 44(2):245–260, December 17, 1992. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279290014O.

Mastroianni:1992:UCD

[Mas92c] Giuseppe Mastroianni. Uniform convergence of derivatives of Lagrange interpolation. Journal of Computational and Applied Mathematics, 43(1–2):37– 51, November 25, 1992. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/037704279290258Y.

Mason:1993:CPS

[Mas93] J. C. Mason. Chebyshev polynomials of the second, third and fourth kinds in approximation, indefinite integration, and integral transforms. Journal of Computational and Applied Mathematics, 49(1–3):169–178, December 31, 1993. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042793901485.

Mastroianni:1995:SWP

[Mas95]  G. Mastroianni.  Some

weighted polynomial inequalities. Journal of Computational and Applied Mathematics, 65(1–3):279– 292, December 29, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042795001166.

Maset:1999:ASN

[Mas99] S. Maset. Asymptotic stability in the numerical solution of linear pure delay differential equations as abstract Cauchy problems. Journal of Computational and Applied Mathematics, 111(1–2):163–172, November 15, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042799001405.

Matos:1990:EAB

[Mat90] Ana C. Matos. Extrapolation algorithms based on the asymptotic expansion of the inverse of the error; application to continued fractions. Journal of Computational and Applied Mathematics, 32(1–2):179–190, November 26, 1990. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0377042790904294.

Matjila:1995:BWL

[Mat95] D. M. Matjila. Bounds for the weighted Lebesgue functions for Freud weights on a larger interval. Journal of Computational and Applied Mathematics, 65(1–3):293– 298, December 29, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042795001174.

Matos:1997:IRE

[Mat97] Ana C. Matos. Integral representation of the error and asymptotic error bounds for generalized Padé type approximants. Journal of Computational and Applied Mathematics, 77(1–2):239– 254, January 13, 1997. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S037704279600129X.

Matsuzawa:1998:FEA

[Mat98]  Yuki Matsuzawa.  Fi-

nite element approximation for some quasilinear elliptic problems. Journal of Computational and Applied Mathematics, 96(1):13–25, September 3, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042798000855.

Mauthner:1998:SSC

[Mau98] Susanne Mauthner. Step size control in the numerical solution of stochastic differential equations. Journal of Computational and Applied Mathematics, 100(1):93–109, November 30, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042798001393.

Mavromatis:1995:TII

[Mav95] Harry A. Mavromatis. Two integrals involving products of two Bessel and a generalized hypergeometric function. Journal of Computational and Applied Mathematics, 59(3):381– 384, May 30, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279500035D.

Mayer:1995:EIV

[May95] Günter Mayer. Epsiloninflation in verification algorithms. Journal of Computational and Applied Mathematics, 60(1–2):147–169, June 20, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279400089J.

Mishev:1992:NSC

[MB92] D. P. Mishev and D. D. Bainov. A necessary and sufficient condition for oscillation of neutral type hyperbolic equations. Journal of Computational and Applied Mathematics, 42(2):215– 220, October 12, 1992. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0377042792900759.

Melloy:1993:CEI

[MB93] Brian J. Melloy and G. Kemble Bennett. Computing the exponential of an intensity matrix. Journal of Computational and Applied Mathematics, 46(3): 405–413, June 22, 1993. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/037704279390036B.

Moreno-Balcazar:1999:APN

[MBMF99] Juan J. Moreno-Balcázar and Andrei Martínez-Finkelshtein. Analytic properties of nondiagonal Jacobi–Sobolev orthogonal polynomials. Journal of Computational and Applied Mathematics, 105 (1–2):393–401, May 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042799000473.

Marcellan:1993:IPO

[MBP93] F. Marcellán, A. Branquinho, and J. Petronilho. On inverse problems for orthogonal polynomials, I. Journal of Computational and Applied Mathematics, 49(1–3):153–160, December 31, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042793901463.

McCoy:1993:NCE

[McC93] Peter A. McCoy. Nearcircularity of the error curve in Chebyshev approximation of solutions to a class of elliptic partial differential equations in the complex plane. Journal of Computational and Applied Mathematics, 46(1–2):315–326, June 14, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279390305U.

McCartin:1998:MTR

[McC98]  Brian J. McCartin.  A

model-trust region algorithm utilizing a quadratic interpolant. Journal of Computational and Applied Mathematics, 91(2):249– 259, May 4, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042798000521.

McCarthy:1999:TCO

[McC99] C. Maeve McCarthy. The tallest column — optimality revisited. Journal of Computational and Applied Mathematics, 101(1–2):27– 37, January 15, 1999. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042798001885.

Moussa:1999:CFB

[MCM99] Pierre Moussa, Andrea Cassa, and Stefano Marmi. Continued fractions and brjuno functions. Journal of Computational and Applied Mathematics, 105(1–2):403– 415, May 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042799000291.

McNamee:1993:BRP

[McN93] John Michael McNamee. A bibliography on roots of polynomials. Journal of Computational and Applied Mathematics, 47(3):391– 394, September 30, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
037704279390064I.

McNamee:1997:SBR

[McN97] John Michael McNamee.

A supplementary bibliography on roots of polynomials. Journal of Computational and Applied Mathematics, 78(1):1, February 3, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042796001471.

McNamee:1999:USB

[McN99] John Michael McNamee. An updated supplementary bibliography on roots of polynomials. Journal of Computational and Applied Mathematics, 110(2):305–306, October 30, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042799002733.

Muniz:1999:CSI

[MdCVR99] Wagner Barbosa Muniz, Haroldo F. de Campos Velho, and Fernando Manuel Ramos. A comparison of some inverse methods for estimating the initial condition of the heat equation. Journal of Computational and Applied Mathematics, 103(1):145–163, March 15, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042798002490.

Marcellan:1990:OPP

[MDR90] F. Marcellan, J. S. Dehesa, and A. Ronveaux. On orthogonal polynomials with perturbed recurrence relations. Journal of Computational and Applied Mathematics, 30(2): 203–212, May 28, 1990. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279090028X.

Milovanovic:1998:SMO

[MDR98] Gradimir V. Milovanović, Bratislav Danković, and Snezana Lj. Rancić. Some Müntz orthogonal systems. Journal of Computational and Applied Mathematics, 99(1–2):299–310, November 16, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042798001642.

Messina:1997:PIL

[MdSvdV97] Eleonora Messina, Jacques J. B. de Swart, and Wolter A. van der Veen. Parallel iterative linear solvers for multistep Runge– Kutta methods. Journal of Computational and Applied Mathematics, 85(1):145– 167, November 6, 1997. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042797001350.

Mason:1993:NMC

[ME93] J. C. Mason and G. H. Elliott. Near-minimax complex approximation by four kinds of Chebyshev polynomial expansion. Journal of Computational and Applied Mathematics, 46(1–2):291– 300, June 14, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279390303S.

Miller:1994:SGT

[ME94] Allen R. Miller and Harold Exton. Sonine–Gegenbauertype integrals. Journal of Computational and Applied Mathematics, 55(3):289– 310, November 30, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.

com/science/article/pii/
0377042794900353.

Medina:1995:ABS

[Med95] R. Medina. Asymptotic behavior of the solutions of second-order difference equations. Journal of Computational and Applied Mathematics, 64(3):233– 246, December 20, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042795001522.

Medina:1996:APS

[Med96] Rigoberto Medina. Asymptotic properties of solutions of nonlinear difference equations. Journal of Computational and Applied Mathematics, 70(1):57–66, June 14, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042795001395.

Medina:1997:SAB

[Med97] Rigoberto Medina. Stability and asymptotic behavior of difference equations. Journal of Computational and Applied Mathematics, 80(1): 17–30, April 14, 1997. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042796001513.

Medina:1998:NSO

[Med98a] Rigoberto Medina. Nonoscillatory solutions for the onedimensional p-Laplacian. Journal of Computational and Applied Mathematics, 98(1):27–33, October 5, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704279800096X.

Medina:1998:SAB

[Med98b] Rigoberto Medina. Stability and asymptotic behavior of perturbed difference systems. Journal of Computational and Applied Mathematics, 98(2):179–189, October 19, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042798001228.

Meerbergen:1996:BRS

[Mee96] K. Meerbergen. Book review: A survey of preconditioned iterative methods: Are Magnus Bruaset, Research notes in mathematics, 328, Longman, 1995, ISBN 0-582-27654-

3. Journal of Computational and Applied Mathematics, 66(1–2):N4–N5, January 31, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042796804729.

Mohamed:1995:EKH

[MEES95] A. A. Mohamed, E. F. Elshehawey, and M. F. El-Sayed. Electrohydrodynamic Kelvin–Helmholtz instability for a velocity stratified fluid. Journal of Computational and Applied Mathematics, 60(3): 331–346, June 30, 1995. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0377042794000486.

Meijer:1993:RCZ

[Mei93] H. G. Meijer. On real and complex zeros of orthogonal polynomials in a discrete Sobolev space. Journal of Computational and Applied Mathematics, 49(1–3):179–191, December 31, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042793901496.

Melman:1997:NCM

[Mel97] A. Melman. A numerical comparison of methods for solving secular equations. Journal of Computational and Applied Mathematics, 86(1):237–249, November 28, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704279782122X.

Melman:1998:SFR

[Mel98]  A. Melman.  Spectral

functions for real symmetric Toeplitz matrices. Journal of Computational and Applied Mathematics, 98(2):233–243, October 19, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042798001290.

Meneguette:1991:CNM

[Men91] M. Meneguette. Chawla– Numerov method revisited. Journal of Computational and Applied Mathematics, 36(2):247–250, August 27, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279190030N.

Messaoudi:1996:RIAa

[Mes96a] A. Messaoudi. Recursive interpolation algorithm: a formalism for solving systems of linear equations — I. Direct methods. Journal of Computational and Applied Mathematics, 76(1–2):13– 30, December 17, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042796000660.

Messaoudi:1996:RIAb

[Mes96b] A. Messaoudi. Recursive interpolation algorithm: a formalism for solving systems of linear equations — II. Iterative methods. Journal of Computational and Applied Mathematics, 76(1–2):31– 53, December 17, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042796000714.

Mikhail:1993:EMV

[MET93] Mikhail Naguib Mikhail and Mostafa R. El-Tantawy. Effect of the medium viscosity on sound propagation and attenuation in ducts. Journal of Computational and Applied Mathematics, 45(3): 283–298, May 3, 1993. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/037704279390046E.

Mikhail:1994:ABL

[MET94] Mikhail Naguib Mikhail and Mostafa R. El-Tantawy. The acoustic boundary layers: a detailed analysis. Journal of Computational and Applied Mathematics, 51(1):15–36, May 30, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042794900914.

Meyer:1994:CLP

[Mey94] Kenneth R. Meyer. Comet like periodic orbits in the Nbody problem. Journal of Computational and Applied Mathematics, 52(1–3):337– 351, July 20, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042794903654.

Martinez-Finkelshtein:1998:APS

[MF98] Andrei Martínez-Finkelshtein.

Asymptotic properties of Sobolev orthogonal polynomials. Journal of Computational and Applied Mathematics, 99(1–2):491–510, November 16, 1998. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042798001794.

Machida:1995:MMS

[MFI95] Naoki Machida, Masao Fukushima, and Toshihide Ibaraki. A multisplitting method for symmetric linear complementarity problems. Journal of Computational and Applied Mathematics, 62(2):217–227, September 20, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042794001032.

Martinez-Finkelshtein:1997:SAG

[MFMBPC97] Andrei Martínez-Finkelshtein, Juan J. Moreno-Balcázar, and Héctor Pijeira-Cabrera. Strong asymptotics for Gegenbauer–Sobolev orthogonal polynomials. Journal of Computational and Applied Mathematics, 81(2): 211–216, July 8, 1997. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042797000599.

Marcellan:1991:OPU

[MG91] F. Marcellán and E. Godoy. Orthogonal polynomials on the unit circle: distribution of zeros. Journal of Computational and Applied Mathematics, 37(1–3):195–208, November 18, 1991. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0377042791901184.

Marano:1994:PAT

[MH94] Miguel Marano and Robert Huotari. The Pólya algorithm on tubular sets. Journal of Computational and Applied Mathematics, 54(2):151–157, September 30, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042794901759.

Micchelli:1992:BMB

[Mic92] Charles A. Micchelli. Banded matrices with banded inverses. Journal of Computational and Applied Mathematics, 41(3):281–300, August 28, 1992. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279290136L.

Mickens:1999:DND

[Mic99] Ronald E. Mickens. Discretizations of nonlinear differential equations using explicit nonstandard methods. Journal of Computational and Applied Mathematics, 110(1):181–185, October 15, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042799002332.

Miel:1991:TSC

[Mie91] George Miel. Trends in systolic and cellular computation. Journal of Computational and Applied Mathematics, 38(1–3):297–321, December 23, 1991. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/037704279190178M.

Miekkala:1998:RWR

[Mie98] Ulla Miekkala. Remarks on waveform relaxation method with overlapping splittings. Journal of Computational and Applied Mathematics, 88(2):349– 361, March 2, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042797002240.

Mignolet:1995:MPO

[Mig95] Marc P. Mignolet. Matrix polynomials orthogonal on the unit circle and accuracy of autoregressive models. Journal of Computational and Applied Mathematics, 62(2):229–238, September 20, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042795001300.

Milovanovic:1993:POS

[Mil93] Gradimir V. Milovanović. On polynomials orthogonal on the semicircle and applications. Journal of Computational and Applied Mathematics, 49(1–3):193–199, December 31, 1993. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279390150A.

Miller:1997:CGH

[Mil97a] Allen R. Miller. A class of generalized hypergeometric summations. Journal of Computational and Applied Mathematics, 87(1):79–85, December 18, 1997. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042797001763.

Miller:1997:CST

[Mil97b] Allen R. Miller. On certain Schlömilch-type series. Journal of Computational and Applied Mathematics, 80(1):83–95, April 14, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042797000137.

Miller:1997:MTP

[Mil97c] Allen R. Miller. On the Mellin transform of products of Bessel and generalized hypergeometric functions. Journal of Computational and Applied Mathematics, 85(2):271–286, November 12, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042797001295.

Miller:1998:RGB

[Mil98]  Allen R. Miller.  Re-

marks on a generalized beta function. Journal of Computational and Applied Mathematics, 100(1):23–32, November 30, 1998. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042798001216.

Milgram:2006:CNT

[Mil06] M. Milgram. Comment on “A new two-term relation for the _{3}F_{2} hypergeometric function of unit argument”. Journal of Computational and Applied Mathematics, 196(2):437–438, November 15, 2006. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042705005868.

See [Ext99].

Min:1998:LTQ

[Min98]  G. Min.  Lobatto-type

quadrature formula in rational spaces. Journal of Computational and Applied Mathematics, 94(1):1–12, July 31, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042798000685.

Mitchell:1991:ARA

[Mit91] William F. Mitchell. Adaptive refinement for arbitrary finite-element spaces with hierarchical bases. Journal of Computational and Applied Mathematics, 36(1): 65–78, August 20, 1991. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279190226A.

Miyakoda:1992:BCI

[Miy92] Tsuyako Miyakoda. Balanced convergence of iterative methods to a multiple zero of a complex polynomial. Journal of Computational and Applied Mathematics, 39(2):201– 212, March 30, 1992. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279290130P.

Miyakoda:1993:MEA

[Miy93] Tsuyako Miyakoda. Multiplicity estimating algorithm for zeros of a complex polynomial and its applications. Journal of Computational and Applied Mathematics, 46(3):357– 368, June 22, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042793900327.

Mohanty:1996:HAD

[MJ96] R. K. Mohanty and M. K. Jain. High accuracy difference schemes for the system of two space nonlinear parabolic differential equations with mixed derivatives and variable coefficients. Journal of Computational and Applied Mathematics, 70(1):15–32, June 14, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042795001352.

Mohanty:1996:HOD

[MJG96a] R. K. Mohanty, M. K. Jain, and Kochurani George. High order difference schemes for the system of two space second order nonlinear hyperbolic equations with variable coefficients. Journal of Computational and Applied Mathematics, 70(2): 231–243, June 28, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0377042795002014.

Mohanty:1996:UHO

[MJG96b] R. K. Mohanty, M. K. Jain, and Kochurani George. On the use of high order difference methods for the system of one space second order nonlinear hyperbolic equations with variable coefficients. Journal of Computational and Applied Mathematics, 72(2):421–431, August 13, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042796000118.

Mohanty:1998:FOA

[MJG98] R. K. Mohanty, M. K. Jain, and Kochurani George. Fourth-order approximations at first time level, linear stability analysis and the numerical solution of multidimensional second-order nonlinear hyperbolic equations in polar coordinates. Journal of Computational and Applied Mathematics, 93(1):1– 12, July 2, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042798000545.

Mitrouli:1996:NPM

[MKK96] M. Mitrouli, N. Karcanias, and C. Koukouvinos. Numerical performance of the matrix pencil algorithm computing the greatest common divisor of polynomials and comparison with other matrix-based methodologies. Journal of Computational and Applied Mathematics, 76(1–2):89–112, December 17, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042796000921.

Matthys:1996:RCP

[MLD96] Lieven Matthys, Hendrik Lambert, and Gilbert De Mey. A recursive construction of particular solutions to a system of coupled linear partial differential equations with polynomial source term. Journal of Computational and Applied Mathematics, 69(2): 319–329, May 23, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0377042795000380.

Magnus:1990:CFL

[MM90] Arne Magnus and John Mc-Cabe. On a continued fraction for log_{e}^{2}(1 + x). Journal of Computational and Applied Mathematics, 30(1): 81–86, April 10, 1990. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/037704279090007M.

Miersemann:1991:SCS

[MM91] E. Miersemann and H. D. Mittelmann. Stability and continuation of solutions to obstacle problems. Journal of Computational and Applied Mathematics, 35(1–3): 5–31, June 26, 1991. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279190195P.

Mazzia:1997:HOT

[MM97] Annamaria Mazzia and Francesca Mazzia. Highorder transverse schemes for the numerical solution of PDEs. Journal of Computational and Applied Mathematics, 82(1–2):299–311, September 15, 1997. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042797000903.

Miller:1998:CGI

[MM98] Allen R. Miller and Ira S.

Moskowitz. On certain generalized incomplete gamma functions. Journal of Computational and Applied Mathematics, 91(2):179– 190, May 4, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042798000314.

Ma:1999:TPD

[MM99a] William Ma and David Minda. Two-point distortion for univalent functions. Journal of Computational and Applied Mathematics, 105(1–2):385–392, May 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042799000217.

Melnik:1999:OOQ

[MM99b] K. N. Melnik and R. V. N. Melnik. Optimal-by-order quadrature formulae for fast oscillatory functions with inaccurately given a priori information. Journal of Computational and Applied Mathematics, 110(1):45–72, October 15, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042799001958.

Mezzacappa:1999:NTC

[MM99c] Anthony Mezzacappa and O. E. B. Messer. Neutrino transport in core collapse supernovae. Journal of Computational and Applied Mathematics, 109(1–2):281– 319, September 30, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042799001624.

Marcellan:1997:ASO

[MMFMB97] Francisco Marcellán, Andrei Martínez-Finkelshtein, and Juan J. Moreno-Balcázar. Asymptotics of Sobolev orthogonal polynomials for symmetrically coherent pairs of measures with compact support. Journal of Computational and Applied Mathematics, 81(2):217–227, July 8, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042797000575.

Marcellan:1997:ARL

[MMPP97] Francisco Marcellán, Henk G.

Meijer, Teresa E. Pérez, and Miguel A. Piñar. An asymptotic result for Laguerre– Sobolev orthogonal polynomials. Journal of Computational and Applied Mathematics, 87(1):87–94, December 18, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042797001799.

Medrano:1996:FAE

[MMST96] A. Medrano, P. Myers, H. M. Stark, and A. Terras. Finite analogues of Euclidean space. Journal of Computational and Applied Mathematics, 68(1–2):221– 238, April 22, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042795002618.

Mastroianni:1991:MCD

[MN91] Giuseppe Mastroianni and Paul Nevai. Mean convergence of derivatives of Lagrange interpolation. Journal of Computational and Applied Mathematics, 34(3): 385–396, April 26, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042791900963.

Milne:1996:VWP

[MN96] Stephen C. Milne and John W. Newcomb. U(n) very-well-poised_{10} ø_{9} transformations. Journal of Computational and Applied Mathematics, 68(1–2):239– 285, April 22, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042795002480.

Mastroianni:1996:LPS

[MO96a] G. Mastroianni and D. Occorsio. Legendre polynomials of the second kind, Fourier series and Lagrange interpolation. Journal of Computational and Applied Mathematics, 75(2):305– 327, November 28, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.

com/science/article/pii/
S0377042796000830.

Mastronardi:1996:SNA

[MO96b] N. Mastronardi and D. Occorsio. Some numerical algorithms to evaluate Hadamard finite-part integrals. Journal of Computational and Applied Mathematics, 70(1):75–93, June 14, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042795001344.

Moews:1995:ETB

[Moe95]  David Moews.  Explicit

Tauberian bounds for multivariate functions. Journal of Computational and Applied Mathematics, 58(3):355– 361, April 20, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279400008O.

Mohd:1990:IGO

[Moh90] Ismail Bin Mohd. An interval global optimization algorithm for a class of functions with several variables. Journal of Computational and Applied Mathematics, 31(3):373–382, October 1, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279090037Z.

Mohd:1995:CEB

[Moh95] Ismail Bin Mohd. Computable error bounds for an optimization problem with parallelepiped constraint. Journal of Computational and Applied Mathematics, 58(2):183–192, March 30, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042793E0201V.

Mohanty:1997:ODM

[Moh97]  R. K. Mohanty.  Or-

der h^{4} difference methods for a class of singular two space elliptic boundary value problems. Journal of Computational and Applied Mathematics, 81(2): 229–247, July 8, 1997. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042797000587.

Mohanty:1998:HAD

[Moh98] R. K. Mohanty. High accuracy difference schemes for a class of three space dimensional singular parabolic equations with variable coefficients. Journal of Computational and Applied Mathematics, 89(1):39–51, March 9, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042797002239.

Monegato:1992:P

[Mon92] Giovanni Monegato. Preface. Journal of Computational and Applied Mathematics, 43(1–2):1–2, November 25, 1992. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279290255V.

Monk:1993:ANM

[Mon93] Peter Monk. An analysis of Nédélec’s method for the spatial discretization of Maxwell’s equations. Journal of Computational and Applied Mathematics, 47(1): 101–121, June 28, 1993. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279390093Q.

Monegato:1994:NEH

[Mon94a] Giovanni Monegato. Numerical evaluation of hypersingular integrals. Journal of Computational and Applied Mathematics, 50(1–3):9–31, May 20, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042794902879.

Monk:1994:ENC

[Mon94b]  Peter Monk.  On the

p- and hp-extension of Nédélec’s curl-conforming elements. Journal of Computational and Applied Mathematics, 53(1):117– 137, July 29, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279200127U.

Monch:1996:NSD

[Mön96a] Lars Mönch. On the numerical solution of the direct scattering problem for an open sound-hard arc. Journal of Computational and Applied Mathematics, 71(2): 343–356, July 27, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042795002308.

Montefusco:1996:SOW

[Mon96b] Laura B. Montefusco. Semiorthogonal wavelet packet bases for parallel leastsquares approximation. Journal of Computational and Applied Mathematics, 73(1– 2):191–208, October 5, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.

com/science/article/pii/
037704279600043X.

Monk:1998:PEI

[Mon98]  Peter Monk.  A poste-

riori error indicators for Maxwell’s equations. Journal of Computational and Applied Mathematics, 100 (2):173–190, December 21, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042798001873.

Mooney:1995:MPA

[Moo95] J. W. Mooney. A monotonicity procedure for the accurate, efficient solution of certain heat-transfer problems with variable thermal properties. Journal of Computational and Applied Mathematics, 58(1):89–101, March 20, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042793E0255K.

Moret:1992:MCN

[Mor92] Igor Moret. A method for the computation of nonsimple turning points corresponding to cusps. Journal of Computational and Applied Mathematics, 40(3): 269–284, July 24, 1992. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/037704279290183X.

Maraner:1991:SMC

[MOT91] P. Maraner, E. Onofri, and G. P. Tecchioli. Spectral methods in computational quantum mechanics. Journal of Computational and Applied Mathematics, 37(1–3):209–219, November 18, 1991. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0377042791901195.

Moussiaux:1993:CRP

[Mou93]  A. Moussiaux.  CON-

VODE: a reduce package for solving differential equations. Journal of Computational and Applied Mathematics, 48(1–2):157–165, October 29, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279390320B.

Marletta:1992:ASS

[MP92] Marco Marletta and John D. Pryce. Automatic solution of Sturm–Liouville problems using the Pruess method. Journal of Computational and Applied Mathematics, 39(1):57–78, February 28, 1992. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279290222J.

Moiseev:1993:COS

[MP93] E. I. Moiseev and A. P.

Prudnikov. On a complete orthonormal system of special functions. Journal of Computational and Applied Mathematics, 49(1–3):201–206, December 31, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279390151Z.

Maitre:1995:ACM

[MP95] Jean-François Maitre and Olivier Pourquier. About the conditioning of matrices in the p-version of the finite element method for second order elliptic problems. Journal of Computational and Applied Mathematics, 63(1–3):341–348, November 20, 1995. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042795000623.

Maisse:1997:DDP

[MP97] E. Maisse and J. Pousin.

Diffusion and dissolution/ precipitation in an open porous reactive medium. Journal of Computational and Applied Mathematics, 82(1–2):279–290, September 15, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042797000496.

Marcellan:1999:PLH

[MP99a] F. Marcellán and E. Prianes. Perturbations of Laguerre–Hahn linear functionals. Journal of Computational and Applied Mathematics, 105(1–2):109–128, May 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042799000254.

Meijer:1999:ASO

[MP99b] Henk G. Meijer and Miguel A. Piñar. Asymptotics of Sobolev orthogonal polynomials for coherent pairs of Jacobi type. Journal of Computational and Applied Mathematics, 108(1–2):87– 97, August 15, 1999. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042799001028.

Marcellan:1996:LSO

[MPP96] Francisco Marcellán, Teresa E. Pérez, and Miguel A. Piñar. Laguerre–Sobolev orthogonal polynomials. Journal of Computational and Applied Mathematics, 71(2): 245–265, July 27, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042795002340.

Marcellan:1995:WBC

[MPPP95] Francisco Marcellán, JoséC.

Petronilho, Teresa E. Pérez, and Miguel A. Piñar. What is beyond coherent pairs of orthogonal polynomials? Journal of Computational and Applied Mathematics, 65(1–3):267–277, December 29, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042795001212.

Moszkowicz:1996:DDR

[MPS96] P. Moszkowicz, J. Pousin, and F. Sanchez. Diffusion and dissolution in a reactive porous medium: Mathematical modelling and numerical simulations. Journal of Computational and Applied Mathematics, 66(1–2):377– 389, January 31, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042795001921.

Morales:1993:DMI

[MPSM93] D. Morales, L. Pardo, M. Salicrú, and M. L. Menéndez. The λ-divergence and the λ-mutual information: estimation in the stratified sampling. Journal of Computational and Applied Mathematics, 47(1):1– 10, June 28, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279390086Q.

Martinez:1995:INM

[MQ95] JoséMario Martínez and Liqun Qi. Inexact Newton methods for solving nonsmooth equations. Journal of Computational and Applied Mathematics, 60(1–2):127– 145, June 20, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279400088I.

Meredith:1992:BAM

[MR92a] D. Meredith and H. Rasmussen. Boundary approximation methods for potential problems associated with moving boundary problems. Journal of Computational and Applied Mathematics, 39(2):133– 149, March 30, 1992. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279290125H.

Meredith:1992:RAS

[MR92b] D. Meredith and H. Rasmussen. Rational approximation schemes for unstable Hele–Shaw flows. Journal of Computational and Applied Mathematics, 42(2):199– 214, October 12, 1992. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042792900748.

Milovanovic:1994:POC

[MR94] Gradimir V. Milovanović and Predrag M. Rajković. On polynomials orthogonal on a circular arc. Journal of Computational and Applied Mathematics, 51(1):1– 13, May 30, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279200059I.

Marcellan:1995:SLF

[MR95a] F. Marcellán and I. A.

Rocha. On semiclassical linear functionals: integral representations. Journal of Computational and Applied Mathematics, 57(1–2):239– 249, February 20, 1995. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042793E0248K.

Marusic:1995:SEB

[MR95b] Miljenko Marusić and Mladen Rogina. Sharp error bounds for interpolating splines in tension. Journal of Computational and Applied Mathematics, 61(2):205– 223, July 31, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042794000589.

Maillard:1996:SCI

[MR96] P. Maillard and M. V. Romerio. A stability criterion for an infinitely long Hall–Héroult cell. Journal of Computational and Applied Mathematics, 71(1):47–65, July 10, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042795002170.

Maryska:1995:MHF

[MRT95] J. Maryska, M. Rozlozník, and M. Tuma. Mixedhybrid finite element approximation of the potential fluid flow problem. Journal of Computational and Applied Mathematics, 63(1–3):383–392, November 20, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042795000666.

Mulholland:1992:RPS

[MS92] L. S. Mulholland and D. M.

Sloan. The role of preconditioning in the solution of evolutionary partial differential equations by implicit Fourier pseudospectral methods. Journal of Computational and Applied Mathematics, 42(2):157– 174, October 12, 1992. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042792900715.

Marcellan:1993:OPC

[MS93] Francisco Marcellán and Gabriela Sansigre. Orthogonal polynomials and cubic transformations. Journal of Computational and Applied Mathematics, 49(1–3):161– 168, December 31, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042793901474.

Maya:1998:INS

[MS98a] C. Maya and R. Shivaji.

Instability of nonnegative solutions for a class of semilinear elliptic boundary value problems. Journal of Computational and Applied Mathematics, 88(1):125– 128, February 23, 1998. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042797002094.

Meyer-Spasche:1998:DSO

[MS98b] Rita Meyer-Spasche. Difference schemes of optimum degree of implicitness for a family of simple ODEs with blow-up solutions. Journal of Computational and Applied Mathematics, 97(1–2):137–152, September 24, 1998. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042798001071.

MalysheV:1999:SLM

[MS99a] A. N. MalysheV and M. Sadkane. On the stability of large matrices. Journal of Computational and Applied Mathematics, 102(2):303– 313, February 26, 1999. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042798002313.

Monegato:1999:NIF

[MS99b] G. Monegato and L. Scuderi. Numerical integration of functions with boundary singularities. Journal of Computational and Applied Mathematics, 112(1–2):201– 214, November 30, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042799002307.

Ma:1998:SAP

[MT98] Wanbiao Ma and Yasuhiro Takeuchi. Stability analysis on a predator–prey system with distributed delays. Journal of Computational and Applied Mathematics, 88(1):79–94, February 23, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042797002033.

McLean:1996:DVT

[MTW96] W. McLean, V. Thomée, and L. B. Wahlbin. Discretization with variable time steps of an evolution equation with a positivetype memory term. Journal of Computational and Applied Mathematics, 69(1): 49–69, April 30, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042795000259.

Muhlbach:1990:IRF

[Müh90]  G. Mühlbach.  On in-

terpolation by rational functions with prescribed poles with applications to multivariate interpolation. Journal of Computational and Applied Mathematics, 32(1–2):203–216, November 26, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279090431X.

Muhlbach:1996:HIC

[Müh96] G. Mühlbach. On Hermite interpolation by Cauchy– Vandermonde systems: the Lagrange formula, the adjoint and the inverse of a Cauchy–Vandermonde matrix. Journal of Computational and Applied Mathematics, 67(1):147–159, February 20, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042794001162.

Muldoon:1993:PZO

[Mul93] Martin E. Muldoon. Properties of zeros of orthogonal polynomials and related functions. Journal of Computational and Applied Mathematics, 48(1–2):167– 186, October 29, 1993. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0377042793903212.

Muldoon:1995:EZB

[Mul95] Martin E. Muldoon. Electrostatics and zeros of Bessel functions. Journal of Computational and Applied Mathematics, 65(1–3):299– 308, December 29, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042795001182.

Musschoot:1999:LTM

[Mus99] C. Musschoot. A Lanczostype method for solving nonsymmetric linear systems with multiple righthand sides — matrix and polynomial interpretation. Journal of Computational and Applied Mathematics, 101(1–2):61–85, January 15, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042798001952.

Murthy:1992:SPI

[MV92] A. S. Vasudeva Murthy and J. G. Verwer. Solving parabolic integro-differential equations by an explicit integration method. Journal of Computational and Applied Mathematics, 39(1): 121–132, February 28, 1992. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
037704279290229Q.

Meddahi:1996:CBI

[MVMP96] Salim Meddahi, Javier Valdés, Omar Menéndez, and Pablo Pérez. On the coupling of boundary integral and mixed finite element methods. Journal of Computational and Applied Mathematics, 69(1):113– 124, April 30, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042795000232.

McClure:1990:AEQ

[MW90a] J. P. McClure and R. Wong. Asymptotic expansion of a quadruple integral involving a Bessel function. Journal of Computational and Applied Mathematics, 33(2):199–215, December 21, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279090369B.

McClure:1990:MSP

[MW90b] J. P. McClure and R. Wong. Multidimensional stationary phase approximation: boundary stationary point. Journal of Computational and Applied Mathematics, 30(2):213–225, May 28, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
037704279090029Y.

Meek:1995:ASP

[MW95] D. S. Meek and D. J.

Walton. Approximating smooth planar curves by arc splines. Journal of Computational and Applied Mathematics, 59(2):221– 231, May 19, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279400029Z.

Meek:1997:GHI

[MW97] D. S. Meek and D. J. Walton. Geometric Hermite interpolation with tschirnhausen cubics. Journal of Computational and Applied Mathematics, 81(2): 299–309, July 8, 1997. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042797000666.

Meek:1999:P

[MW99a] D. S. Meek and D. J. Walton. Preface. Journal of Computational and Applied Mathematics, 102(1): 1–2, February 15, 1999. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042798002040.

Meek:1999:SAS

[MW99b] D. S. Meek and D. J. Walton. Spiral arc spline approximation to a planar spiral. Journal of Computational and Applied Mathematics, 107(1):21–30, July 15, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042799000722.

Micchelli:1997:WGM

[MXZ97] Charles A. Micchelli, Yuesheng Xu, and Yunhe Zhao. Wavelet Galerkin methods for second-kind integral equations. Journal of Computational and Applied Mathematics, 86(1):251– 270, November 28, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704279700160X.

Monk:1995:MCV

[MZ95] Peter Monk and Shangyou Zhang. Multigrid computation of vector potentials. Journal of Computational and Applied Mathematics, 62(3):301–320, September 29, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042794001068.

Nakao:1991:SNP

[Nak91] Mitsuhiro T. Nakao. Solving nonlinear parabolic problems with result verification. Part I: onespace dimensional case. Journal of Computational and Applied Mathematics, 38(1–3):323–334, December 23, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279190179N.

Nedoma:1995:FSC

[ND95] Jirí Nedoma and Jan Dvorák. On the FEM solution of a coupled contacttwo-phase Stefan problem in thermo-elasticity. Coercive case. Journal of Computational and Applied Mathematics, 63(1–3):411–420, November 20, 1995. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042795000917.

Neamtu:1996:HSS

[Nea96] Marian Neamtu. Homogeneous simplex splines. Journal of Computational and Applied Mathematics, 73(1–2):173–189, October 5, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042796000428.

Nedoma:1994:FEA

[Ned94] Jirí Nedoma. Finite-element analysis of contact problems in thermoelasticity. The semi-coercive case. Journal of Computational and Applied Mathematics, 50(1–3): 411–423, May 20, 1994. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0377042794903174.

Nedoma:1995:EMI

[Ned95a]  Jirí Nedoma.  Equa-

tions of magnetodynamics of incompressible thermo-Bingham’s fluid under the gravity effect. Journal of Computational and Applied Mathematics, 59(1):109– 128, April 28, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279400019W.

Nedoma:1995:NAG

[Ned95b] Jirí Nedoma. Nonlinear analysis of the generalized thermo-magnetodynamic problem. Journal of Computational and Applied Mathematics, 63(1–3):393–402, November 20, 1995. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042795000887.

Nedoma:1995:P

[Ned95c]  Jirí Nedoma.  Preface.

Journal of Computational and Applied Mathematics, 63(1–3):1–3, November 20, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042795900454.

Nedoma:1997:CSL

[Ned97a] J. Nedoma. On a coupled Stefan-like problem in thermo-visco-plastic rheology. Journal of Computational and Applied Mathematics, 84(1):45–80, October 21, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042797001118.

Nedoma:1997:FEA

[Ned97b] Jirí Nedoma. Finite element approximation of a coupled contact Stefan-like problem arising from the time discretization in deformation theory of thermo-plasticity. Journal of Computational and Applied Mathematics, 82(1–2):313–331, September 15, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042797000447.

Nees:1994:AVK

[Nee94] Manuela Nees. Approximative versions of Kolmogorov’s superposition theorem, proved constructively. Journal of Computational and Applied Mathematics, 54(2):239–250, September 30, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042794901791.

Newton:1991:LS

[New91]  Paul K. Newton.  Let-

ter section. Journal of Computational and Applied Mathematics, 36(2):273, August 27, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279190035I.

Niederreiter:1990:SIP

[Nie90] H. Niederreiter. Statistical independence properties of pseudorandom vectors produced by matrix generators. Journal of Computational and Applied Mathematics, 31(1):139– 151, July 24, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279090345Z.

Niederreiter:1994:NCP

[Nie94] Harald Niederreiter. On a new class of pseudorandom numbers for simulation methods. Journal of Computational and Applied Mathematics, 56(1–2):159– 167, December 20, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042794903859.

Nievergelt:1995:CSA

[Nie95] Yves Nievergelt. The condition of Steffensen’s acceleration in several variables. Journal of Computational and Applied Mathematics, 58(3):291–305, April 20, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
037704279400004K.

Nikolov:1996:CDQ

[Nik96] Geno Nikolov. On certain definite quadrature formulae. Journal of Computational and Applied Mathematics, 75(2):329– 343, November 28, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0377042796000866.

Njaastad:1995:ESS

[Njå95] Olav Njåstad. Extremal solutions of the strong Stieltjes moment problem. Journal of Computational and Applied Mathematics, 65(1–3):309–318, December 29, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042795001190.

Njaastad:1998:NMS

[Njå98]  Olav Njåstad.  Nevan-

linna matrices for the strong Stieltjes moment problem. Journal of Computational and Applied Mathematics, 99(1–2):311–318, November 16, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042798001654.

Nakamura:1998:IDR

[NKS98] Y. Nakamura, K. Kajiwara, and H. Shiotani. On an integrable discretization of the Rayleigh quotient gradient system and the power method with a shift. Journal of Computational and Applied Mathematics, 96(2):77–90, September 15, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042798000843.

Noldus:1994:ALM

[NL94] E. Noldus and M. Loccufier. An application of Liapunov’s method for the analysis of neural networks. Journal of Computational and Applied Mathematics, 50(1–3):425–432, May 20, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042794903182.

Noda:1992:HAI

[NM92] Matu-Tarow Noda and Eiichi Miyahiro. A hybrid approach for the integration of a rational function. Journal of Computational and Applied Mathematics, 40(3): 259–268, July 24, 1992. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279290182W.

Noor:1993:SAV

[NNR93] Muhammad Aslam Noor, Khalida Inayat Noor, and Themistocles M. Rassias. Some aspects of variational inequalities. Journal of Computational and Applied Mathematics, 47(3):285– 312, September 30, 1993.

CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        037704279390058J.

Notaris:1990:GKQ

[Not90] Sotirios E. Notaris. Gauss– Kronrod quadrature formulae for weight functions of Bernstein–Szegő type, II. Journal of Computational and Applied Mathematics, 29(2):161–169, February 15, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042790903554.

Notaris:1997:IQF

[Not97] Sotirios E. Notaris. Interpolatory quadrature formulae with Chebyshev abscissae of the third or fourth kind. Journal of Computational and Applied Mathematics, 81(1):83–99, June 1, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042797000186.

Notaris:1998:NIQ

[Not98] Sotirios E. Notaris. New interpolatory quadrature formulae with Chebyshev abscissae. Journal of Computational and Applied Mathematics, 94(2):199–214, August 3, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042798000466.

Noutsos:1993:OSP

[Nou93] Dimitrios Noutsos. Optimal stretched parameters for the SOR iterative method. Journal of Computational and Applied Mathematics, 48(3):293–308, November 30, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042793900268.

Nardin:1992:NEC

[NPB92] Mark Nardin, W. F. Perger, and Atul Bhalla. Numerical evaluation of the confluent hypergeometric function for complex arguments of large magnitudes. Journal of Computational and Applied Mathematics, 39(2):193– 200, March 30, 1992. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279290129L.

Neto:1990:SAG

[NR90] Atair Rios Neto and Kondapalli Rama Rao. A stochastic approach to global error estimation in ODE multistep numerical integration. Journal of Computational and Applied Mathematics, 30(3):257– 281, July 10, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042790902799.

Novak:1999:IMH

[NRSS99] Erich Novak, Klaus Ritter, Richard Schmitt, and Achim Steinbauer. On an interpolatory method for high dimensional integration. Journal of Computational and Applied Mathematics, 112(1–2):215–228, November 30, 1999. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042799002228.

Noda:1991:AGA

[NS91] Matu-Tarow Noda and Tateaki Sasaki. Approximate GCD and its application to ill-conditioned equations. Journal of Computational and Applied Mathematics, 38(1–3):335–351, December 23, 1991. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279190180R.

Niederreiter:1994:INF

[NS94] Harald Niederreiter and Ian H. Sloan. Integration of nonperiodic functions of two variables by Fibonacci lattice rules. Journal of Computational and Applied Mathematics, 51(1):57–70, May 30, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279200004S.

Nedoma:1995:MSO

[NS95] Jirí Nedoma and Jirí Stehlík.

Mathematical simulation of osteotomy. Numerical analysis and results. Journal of Computational and Applied Mathematics, 63(1–3):421– 438, November 20, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042795000682.

Neuman:1994:MDS

[NV94] Edward Neuman and Patrick J.

Van Fleet. Moments of Dirichlet splines and their applications to hypergeometric functions. Journal of Computational and Applied Mathematics, 53(2):225– 241, August 19, 1994. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0377042794900477.

Nakao:1994:CPE

[NW94] Mitsuhiro T. Nakao and Yoshitaka Watanabe. On computational proofs of the existence of solutions to nonlinear parabolic problems. Journal of Computational and Applied Mathematics, 50(1–3):401–410, May 20, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042794903166.

Njaastad:1997:APZ

[NW97] Olav Njåstad and Haakon Waadeland. Asymptotic properties of zeros of orthogonal rational functions. Journal of Computational and Applied Mathematics, 77(1–2):255–275, January 13, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042796001306.

Nakao:1998:PCP

[NYW98] Mitsuhiro T. Nakao, Nobito Yamamoto, and Yoshitaka Watanabe. A posteriori and constructive a priori error bounds for finite element solutions of the Stokes equations. Journal of Computational and Applied Mathematics, 91(1):137– 158, April 24, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042798000351.

Nagurney:1990:DVI

[NZ90] Anna Nagurney and Lan Zhao. Disequilibrium and variational inequalities. Journal of Computational and Applied Mathematics, 33(2):181–198, December 21, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279090368A.

Oden:1996:P

[OC96] J. Tinsley Oden and Graham F. Carey. Preface. Journal of Computational and Applied Mathematics, 74(1–2):1, November 5, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042796900734.

Odlyzko:1992:ETE

[Odl92] A. M. Odlyzko. Explicit Tauberian estimates for functions with positive coefficients. Journal of Computational and Applied Mathematics, 41(1–2):187–197, August 20, 1992. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279290248V.

Oden:1996:LPE

[OF96] J. Tinsley Oden and Yusheng Feng. Local and pollution error estimation for finite element approximations of elliptic boundary value problems. Journal of Computational and Applied Mathematics, 74(1–2):245–293, November 5, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042796000271.

Oishi:1995:NVE

[Ois95] Shin’ichi Oishi. Numerical verification of existence and inclusion of solutions for nonlinear operator equations. Journal of Computational and Applied Mathematics, 60(1–2):171–185, June 20, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279400090N.

Okunbor:1993:VSS

[Oku93] Daniel I. Okunbor. Variable step size does not harm second-order integrators for Hamiltonian systems. Journal of Computational and Applied Mathematics, 47(2): 273–275, ???? 1993. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279390009Z.

Olshanskii:1999:TLM

[Ols99] Maxim A. Olshanskii. Twolevel method and some a priori estimates in unsteady Navier–Stokes calculations. Journal of Computational and Applied Mathematics, 104(2):173–191, May 17, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042799000564.

Ooura:1991:DEF

[OM91] Takuya Ooura and Masatake Mori. The double exponential formula for oscillatory functions over the half infinite interval. Journal of Computational and Applied Mathematics, 38(1–3):353–360, December 23, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279190181I.

Ooura:1999:RDE

[OM99] Takuya Ooura and Masatake Mori. A robust double exponential formula for Fourier-type integrals. Journal of Computational and Applied Mathematics, 112(1–2):229–241, November 30, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704279900223X.

Ohe:1994:DLP

[OO94] Takashi Ohe and Kohzaburo Ohnaka. Determination of locations of point-like masses in an inverse source problem of the Poisson equation. Journal of Computational and Applied Mathematics, 54(2):251–261, September 30, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042794901805.

Ozis:1996:DVM

[OO96] Turgut Ozis and Ali Ozdes. A direct variational methods applied to Burgers’ equation. Journal of Computational and Applied Mathematics, 71(2):163– 175, July 27, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042795002219.

Ogreid:1998:SOT

[OO98] Odd Magne Ogreid and Per Osland. Summing oneand two-dimensional series related to the Euler series. Journal of Computational and Applied Mathematics, 98(2):245–271, October 19, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704279800137X.

Owens:1994:MMC

[OP94] Robert G. Owens and Timothy N. Phillips. Massand momentum-conserving spectral methods for Stokes flow. Journal of Computational and Applied Mathematics, 53(2):185–206, August 19, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042794900450.

Orlik:1999:HIA

[Orl99]  Peter Orlik.  Hypergeo-

metric integrals and arrangements. Journal of Computational and Applied Mathematics, 105(1–2):417– 424, May 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042799000369.

Orsi:1990:SAC

[Ors90] Annamaria Palamara Orsi. Spline approximation for Cauchy principal value integrals. Journal of Computational and Applied Mathematics, 30(2):191– 201, May 28, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279090027W.

Orsi:1997:NSA

[Ors97] Annamaria Palamara Orsi. A note on spline approximation for Hadamard finitepart integrals. Journal of Computational and Applied Mathematics, 79(1):67–73, March 3, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042796001550.

Okunbor:1994:CRK

[OS94] Daniel I. Okunbor and Robert D. Skeel. Canonical Runge–Kutta–Nyström methods of orders five and six. Journal of Computational and Applied Mathematics, 51(3):375– 382, June 30, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279200119T.

Osada:1990:ASL

[Osa90] Naoki Osada. Accelerable subsets of logarithmic sequences. Journal of Computational and Applied Mathematics, 32(1–2):217–227, November 26, 1990. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279090432Y.

Osada:1991:AMV

[Osa91] Naoki Osada. Acceleration methods for vector sequences. Journal of Computational and Applied Mathematics, 38(1–3):361–371, December 23, 1991. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279190182J.

Osada:1994:OMR

[Osa94] Naoki Osada. An optimal multiple root-finding method of order three. Journal of Computational and Applied Mathematics, 51(1): 131–133, May 30, 1994. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042794000441.

Osada:1996:VST

[Osa96] Naoki Osada. Vector sequence transformations for the acceleration of logarithmic convergence. Journal of Computational and Applied Mathematics, 66(1–2):391– 400, January 31, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042795001832.

Osada:1998:IOC

[Osa98]  Naoki Osada.  Improv-

ing the order of convergence of iteration functions. Journal of Computational and Applied Mathematics, 98(2):311–315, October 19, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042798001319.

Osborne:1997:CRD

[Osb97] M. R. Osborne. Cyclic reduction, dichotomy, and the estimation of differential equations. Journal of Computational and Applied Mathematics, 86 (1):271–286, November 28, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042797821231.

Oberle:1993:MTG

[OSG^{+}93] Michael K. Oberle, Steven L. Scott, George T. Gilbert, Rhonda L. Hatcher, and David F. Addis. Mellin transforms of a generalization of Legendre polynomials. Journal of Computational and Applied Mathematics, 45(3):367–369, May 3, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
037704279390055G.

Owren:1995:SRK

[Owr95]  B. Owren.  Stability

of Runge–Kutta methods used in modular integration. Journal of Computational and Applied Mathematics, 62(1):89–101, August 31, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042794000786.

Ozis:1991:EAS

[Ozi91a] Turgut Ozis. An efficient approach to the solution of the two-dimensional electrochemical machining problem. Journal of Computational and Applied Mathematics, 36(2):239–246, August 27, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279190029J.

Ozis:1991:EUW

[Ozi91b]  Turgut Ozis.  On the

existence and uniqueness of a weak solution of a nonsteady transient freeboundary problem in porous media. Journal of Computational and Applied Mathematics, 34(1):33–39, February 10, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279190145A.

Ozis:1992:EPS

[Ozi92]  Turgut Ozis.  The ex-

tension of Picard’s successive approximation for constructing two-side bounds for the solutions of differential equations. Journal of Computational and Applied Mathematics, 39(1):7– 14, February 28, 1992. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/037704279290217L.

Perez-Acosta:1990:RCL

[PACH90] Francisco Pérez-Acosta, Luis Casasús, and Nácere Hayek. Rational collocation for linear boundary value problems. Journal of Computational and Applied Mathematics, 33(3):297–305, December 31, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042705800056.

Pages:1992:VCS

[Pag92]  Gilles Pagès.  Van der

corput sequences, Kakutani transforms and onedimensional numerical integration. Journal of Computational and Applied Mathematics, 44(1):21–39, December 9, 1992. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279290051X.

Pages:1998:SQM

[Pag98] Gilles Pagès. A space quantization method for numerical integration. Journal of Computational and Applied Mathematics, 89(1):1– 38, March 9, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042797001908.

Palmer:1992:MRS

[Pal92] E. M. Palmer. Matchings in random superpositions of bipartite trees. Journal of Computational and Applied Mathematics, 41(1–2):199– 204, August 20, 1992. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/037704279290249W.

Palumbo:1998:GSI

[Pal98] Biagio Palumbo. A generalization of some inequalities for the gamma function. Journal of Computational and Applied Mathematics, 88(2):255– 268, March 2, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042797001878.

Pan:1993:ASP

[Pan93a] K. Pan. Asymptotics for Szegő polynomials associated with Wiener–Levinson filters. Journal of Computational and Applied Mathematics, 46(3):387–394, June 22, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042793900349.

Pan:1993:OSR

[Pan93b] K. Pan. On orthogonal systems of rational functions on the unit circle and polynomials orthogonal with respect to varying measures. Journal of Computational and Applied Mathematics, 47(3):313–322, September 30, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279390059K.

Pan:1993:SWC

[Pan93c] K. Pan. Strong and weak convergence of orthogonal systems of rational functions on the unit circle. Journal of Computational and Applied Mathematics, 46(3): 427–436, June 22, 1993. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279390038D.

Pan:1994:CRI

[Pan94]  K. Pan.  On the con-

vergence of the rational interpolation approximant of Carathéodory functions. Journal of Computational and Applied Mathematics, 54(3):371–376, October 20, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042794902577.

Pan:1995:EST

[Pan95a]  K. Pan.  Extensions

of Szegő’s theory of rational functions orthogonal on the unit circle. Journal of Computational and Applied Mathematics, 62(3):321–331, September 29, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279400107X.

Pan:1995:ORF

[Pan95b] K. Pan. On the orthogonal rational functions with arbitrary poles and interpolation properties. Journal of Computational and Applied Mathematics, 60(3): 347–355, June 30, 1995. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0377042794000394.

Panagiotopoulos:1995:MNN

[Pan95c] P. D. Panagiotopoulos. Modelling of nonconvex nonsmooth energy problems. Dynamic hemivariational inequalities with impact effects. Journal of Computational and Applied Mathematics, 63(1–3):123– 138, November 20, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        037704279500050X.

Pan:1996:COR

[Pan96a] K. Pan. Convergence of orthogonal rational functions outside of the unit circle. Journal of Computational and Applied Mathematics, 69(2):345–351, May 23, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042795000461.

Pan:1996:CRF

[Pan96b] K. Pan. On the convergence of rational functions orthogonal on the unit circle. Journal of Computational and Applied Mathematics, 76(1–2):315–324, December 17, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042796001161.

Panorska:1996:GSM

[Pan96c] Anna K. Panorska. Generalized stable models for financial asset returns. Journal of Computational and Applied Mathematics, 70(1): 111–114, June 14, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042795001441.

Pan:1997:SOP

[Pan97] K. Pan. On Sobolev orthogonal polynomials with coherent pairs the Jacobi case. Journal of Computational and Applied Mathematics, 79(2):249– 262, March 17, 1997. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042796001689.

Pao:1998:ABS

[Pao98] C. V. Pao. Asymptotic behavior of solutions of reaction–diffusion equations with nonlocal boundary conditions. Journal of Computational and Applied Mathematics, 88(1):225– 238, February 23, 1998. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S037704279700215X.

Paquet:1994:PEP

[Paq94] Luc Paquet. Precise evaluation of a polynomial at a point given in staggered correction format. Journal of Computational and Applied Mathematics, 50(1–3):433– 454, May 20, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042794903190.

Parlett:1991:SMP

[Par91] Beresford N. Parlett. Symmetric matrix pencils. Journal of Computational and Applied Mathematics, 38(1– 3):373–385, December 23, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.

com/science/article/pii/
037704279190183K. Pro-

ceedings of the International Symposium on Computational Mathematics (Matsuyama, 1990).

Paris:1992:SSP

[Par92] R. B. Paris. Smoothing of the Stokes phenomenon using Mellin–Barnes integrals. Journal of Computational and Applied Mathematics, 41(1–2):117–133, August 20, 1992. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279290242P.

Pardo:1997:CSE

[Par97] M. C. Pardo. A comparison of some estimators of the mixture proportion of mixed normal distributions. Journal of Computational and Applied Mathematics, 84(2):207–217, October 28, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042797001246.

Paszkowski:1990:HPA

[Pas90] Stefan Paszkowski. Hermite– Padé approximation (basic notions and theorems). Journal of Computational and Applied Mathematics, 32(1–2):229–236, November 26, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279090433Z.

Paszkowski:1991:ACP

[Pas91]  Stefan Paszkowski.  An

application of Chebyshev polynomial to a problem of electrical engineering. Journal of Computational and Applied Mathematics, 37(1–3):5–17, November 18, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279190101O.

Paszkowski:1992:ET

[Pas92] Stefan Paszkowski. Evaluation of a C-table. Journal of Computational and Applied Mathematics, 44 (2):219–234, December 17, 1992. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        037704279290012M.

Patterson:1999:SNR

[Pat99] T. N. L. Patterson. Stratified nested and related quadrature rules. Journal of Computational and Applied Mathematics, 112(1–2):243– 251, November 30, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042799002241.

Piessens:1992:CFT

[PB92] Robert Piessens and Maria Branders. On the computation of Fourier transforms of singular functions. Journal of Computational and Applied Mathematics, 43(1–2):159–169, November 25, 1992. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279290264X.

Piana:1998:ARS

[PB98] M. Piana and J. C. Brown.

Analytical and regularized solutions of the synchrotron integral equation. Journal of Computational and Applied Mathematics, 93(2):75–88, July 13, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704279800020X.

Pardhanani:1996:ESC

[PC96] Anand L. Pardhanani and Graham F. Carey. Efficient simulation of complex patterns in reaction–diffusion systems. Journal of Computational and Applied Mathematics, 74(1–2):295–311, November 5, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042796000283.

Pereyra:1999:P

[PC99] V. Pereyra and J. Casillo. Preface. Journal of Computational and Applied Mathematics, 103(1):1–2, March 15, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042798002350.

Prikryl:1995:CMS

[PCC95] P. Prikryl, V. Cháb, and R. Cerný. Computational modeling of solid-state reactions in the Ni–Si systems induced by pulsed lasers. Journal of Computational and Applied Mathematics, 63(1–3):357–363, November 20, 1995. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279500064X.

Puta:1999:CPS

[PCV99] M. Puta, I. Casu, and A. Voitecovici. Constant Poisson structures and the weighted Euler integrator. Journal of Computational and Applied Mathematics, 111(1–2):147–152, November 15, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042799001387.

Potapenko:1999:CCD

[PdA99] I. F. Potapenko and C. A.

de Azevedo. The completely conservative difference schemes for the nonlinear Landau–Fokker– Planck equation. Journal of Computational and Applied Mathematics, 103(1):115– 123, March 15, 1999. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042798002453.

Parra:1996:VFD

[PdST96] M. C. Parra, M. C. López de Silanes, and J. J. Torrens. Vertical fault detection from scattered data. Journal of Computational and Applied Mathematics, 73(1–2):225– 239, October 5, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042796000453.

Peherstorfer:1992:FPO

[Peh92]  Franz Peherstorfer.  Fi-

nite perturbations of orthogonal polynomials. Journal of Computational and Applied Mathematics, 44 (3):275–302, December 24, 1992. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        037704279290002F.

Peherstorfer:1993:OEP

[Peh93] Franz Peherstorfer. Orthogonal and extremal polynomials on several intervals. Journal of Computational and Applied Mathematics, 48(1–2):187–205, October 29, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042793903223.

Peherstorfer:1995:SPF

[Peh95] Franz Peherstorfer. Stieltjes polynomials and functions of the second kind. Journal of Computational and Applied Mathematics, 65(1–3):319–338, December 29, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042795001204.

Pelikan:1994:HBZ

[Pel94] Steve Pelikan. Hopf bifurcations via zero sets. Journal of Computational and Applied Mathematics, 52(1–3): 141–145, July 20, 1994. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042794903530.

Penzel:1999:BIM

[Pen99] Frank Penzel. A boundary integral method applied to a convection–diffusion problem. Journal of Computational and Applied Mathematics, 111(1–2):217–226, November 15, 1999. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042799001442.

Perlstadt:1992:RPE

[Per92] Marci A. Perlstadt. Restricted polynomial expansions and approximation. Journal of Computational and Applied Mathematics, 40(1):127–135, June 12, 1992. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042792900472.

Petkovic:1990:ESC

[Pet90] M. S. Petković. On the efficiency of some combined methods for polynomial complex zeros. Journal of Computational and Applied Mathematics, 30(1): 99–115, April 10, 1990. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/037704279090009O.

Petersen:1999:MFA

[Pet99a] Vigdis Petersen. On measures in frequency analysis. Journal of Computational and Applied Mathematics, 105(1–2):437–443, May 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042799000266.

Petras:1999:CGL

[Pet99b]  Knut Petras.  On the

computation of the Gauss– Legendre quadrature formula with a given precision. Journal of Computational and Applied Mathematics, 112(1–2):253–267, November 30, 1999. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042799002253.

Pfeifer:1991:IEF

[Pfe91]  E. Pfeifer.  Interpola-

tion with exponentially fitted second-order C^{1}-spline functions. Journal of Computational and Applied Mathematics, 34(1):119– 124, February 10, 1991. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042791901519.

Pestano-Gabino:1998:RMU

[PGGC98] Celina Pestano-Gabino and Concepción González-Concepción. Rationality, minimality and uniqueness of representation of matrix formal power series. Journal of Computational and Applied Mathematics, 94(1):23–38, July 31, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042798000594.

Petkovic:1992:HOI

[PH92] Miodrag Petković and Dragoslav Herceg. Higher-order iterative methods for approximating zeros of analytic functions. Journal of Computational and Applied Mathematics, 39(2):243– 258, March 30, 1992. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279290133I.

Petkovic:1998:CWZ

[PH98] Miodrag S. Petković and – Dor– de D. Herceg. On the convergence of Wang– Zheng’s method. Journal of Computational and Applied Mathematics, 91(1):123– 135, April 24, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704279800034X.

Petkovic:1999:RIM

[PH99] Miodrag Petković and Dragoslav Herceg. On rediscovered iteration methods for solving equations. Journal of Computational and Applied Mathematics, 107(2):275– 284, July 31, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042799001053.

Piessens:1996:BRM

[Pie96] R. Piessens. Book review: Mathematica graphics techniques and applications: T. Wickham-Jones, Wolfram Research Inc., Champaign, IL, USA Springer, Berlin 1994, 721 pp., ISBN 3-540-94047-2, Hardcover, £35, DM 80. Journal of Computational and Applied Mathematics, 66(1–2):N6, January 31, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042796804754.

Petkovic:1997:NCW

[PK97] M. S. Petković and N. Kjurkchiev. A note on the convergence of the Weierstrass SOR method for polynomial roots. Journal of Computational and Applied Mathematics, 80(1):163– 168, April 14, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042797000162.

Papanicolaou:1998:ARC

[PKB98] Vassilis G. Papanicolaou, George E. Kokolakis, and Shahar Boneh. Asymptotics for the random coupon collector problem. Journal of Computational and Applied Mathematics, 93(2):95–105, July 13, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042798000582.

Plum:1995:EER

[Plu95] Michael Plum. Existence and enclosure results for continua of solutions of parameter-dependent nonlinear boundary value problems. Journal of Computational and Applied Mathematics, 60(1–2):187–200, June 20, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279400091E.

Peart:1992:CEV

[PM92] Paul B. Peart and Ricardo A. Mitchell. On computing the exact value of dispersion of a sequence. Journal of Computational and Applied Mathematics, 42(3):309–337, October 30, 1992. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279290083A.

Ponnusamy:1998:CCP

[Pon98a]  S. Ponnusamy.  Close-

to-convexity properties of Gaussian hypergeometric functions. Journal of Computational and Applied Mathematics, 88(2):327– 337, March 2, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042797002215.

Ponnusamy:1998:HTF

[Pon98b] S. Ponnusamy. Hypergeometric transforms of functions with derivative in a half plane. Journal of Computational and Applied Mathematics, 96(1):35–49, September 3, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042798000909.

Popkov:1996:MAD

[Pop96] Yu. S. Popkov. Multiplicative algorithms for determining positive solutions of nonlinear equations. Journal of Computational and Applied Mathematics, 69(1): 27–37, April 30, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042795000305.

Panyukov:1991:PAF

[PP91] Anatoly V. Panyukov and Boris V. Pelzwerger. Polynomial algorithms to finite veber problem for a tree network. Journal of Computational and Applied Mathematics, 35(1–3):291–296, June 26, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042791902156.

Pemantle:1992:PIH

[PP92] Robin Pemantle and Mathew D.

Penrose. On path integrals for the high-dimensional Brownian bridge. Journal of Computational and Applied Mathematics, 44 (3):381–390, December 24, 1992. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704279290010U.

Perez:1993:GPZ

[PP93] Teresa E. Pérez and Miguel A.

Piñar. Global properties of zeros for Sobolevtype orthogonal polynomials. Journal of Computational and Applied Mathematics, 49(1–3):225–232, December 31, 1993. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042793901544.

Parks:1996:LGM

[PP96a] H. R. Parks and J. T. Pitts. The least-gradient method for computing area minimizing hypersurfaces spanning arbitrary boundaries. Journal of Computational and Applied Mathematics, 66(1–2):401–409, January 31, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042795001972.

Posch:1996:DRN

[PP96b] K. C. Posch and R. Posch. Division in residue number systems involving length indicators. Journal of Computational and Applied Mathematics, 66(1–2):411–419, January 31, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042795001646.

Paczkowska-Przadka:1997:DIG

[PP97] Katarzyna Paczkowska-Przadka.

Difference inequalities generated by impulsive parabolic differential-functional problems. Journal of Computational and Applied Mathematics, 84(2):189–206, October 28, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042797001234.

Pettigrew:1993:NSU

[PR93] M. F. Pettigrew and H. Rasmussen. Numerical solutions for unstable Hele– Shaw flows. Journal of Computational and Applied Mathematics, 47(2):253– 266, ???? 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279390007X.

Prevost:1990:AWF

[Pre90] Marc Prevost. Approximation of weight function and approached Padé approximants. Journal of Computational and Applied Mathematics, 32(1–2):237–252, November 26, 1990. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042790904342.

Prevost:1994:ASL

[Pré94]  Marc Prévost.  Acceler-

ation of some logarithmic sequences. Journal of Computational and Applied Mathematics, 55(3):357– 367, November 30, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042794900396.

Prevost:1996:NPI

[Pré96] M. Prévost. A new proof of the irrationality of ζ(2) and ζ(3) using Padé approximants. Journal of Computational and Applied Mathematics, 67(2):219– 235, March 29, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0377042795000194.

Prolla:1994:WST

[Pro94]  João B. Prolla.  The

Weierstrass–Stone theorem for convex-cone-valued functions. Journal of Computational and Applied Mathematics, 53(2):171–183, August 19, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042794900442.

Prochazka:1995:OEF

[Pro95] Petr Procházka. Optimal eigenstress fields in layered structures. Journal of Computational and Applied Mathematics, 63(1–3):475– 480, November 20, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042795000933.

Papageorgiou:1990:MCP

[PRS90] C. D. Papageorgiou, A. D.

Raptis, and T. E. Simos. A method for computing phase shifts for scattering. Journal of Computational and Applied Mathematics, 29(1):61–67, January 10, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042790901956.

Ptak:1992:NIP

[PRS92] M. Ptak, A. Rutkowska, and J. Szczurek. A note on inverses of power series. Journal of Computational and Applied Mathematics, 39(1):95–101, February 28, 1992. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279290225M.

Pinto:1997:NSS

[PRT97] S. González Pinto, S. Pérez Rodríguez, and J. I. Montijano Torcal. On the numerical solution of stiff IVPs by Lobatto IIIA Runge– Kutta methods. Journal of Computational and Applied Mathematics, 82(1–2):129– 148, September 15, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042797000861.

Pryce:1994:CVS

[Pry94]  John D. Pryce.  Clas-

sical and vector Sturm– Liouville problems: recent advances in singular-point analysis and shooting-type algorithms. Journal of Computational and Applied Mathematics, 50(1–3):455– 470, May 20, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042794903204.

Papamichael:1993:LBE

[PS93a] N. Papamichael and E. B. Saff. Local behaviour of the error in the Bergman kernel method for numerical conformal mapping. Journal of Computational and Applied Mathematics, 46(1–2): 65–75, June 14, 1993. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279390287L.

Papamichael:1993:P

[PS93b] N. Papamichael and E. B.

Saff. Preface. Journal of Computational and Applied Mathematics, 46(1–2):1–5, June 14, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279390283H.

Papamichael:1994:TAD

[PS94] N. Papamichael and N. S.

Stylianopoulos. On the theory and application of a domain decomposition method for computing conformal modules. Journal of Computational and Applied Mathematics, 50(1–3): 33–50, May 20, 1994. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042794902887.

Peherstorfer:1995:IPZ

[PS95a] Franz Peherstorfer and Michael Schmuckenschläger. Interlacing properties of zeros of associated polynomials. Journal of Computational and Applied Mathematics, 59(1):61–78, April 28, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        037704279400014R.

Peherstorfer:1995:COP

[PS95b] Franz Peherstorfer and Robert Steinbauer. Characterization of orthogonal polynomials with respect to a functional. Journal of Computational and Applied Mathematics, 65(1–3):339–355, December 29, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042795001255.

Partyka:1999:QHE

[PS99] Dariusz Partyka and Ken-Ichi Sakan. Quasiconformality of harmonic extensions. Journal of Computational and Applied Mathematics, 105(1–2):425–436, May 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042799000096.

Papamichael:1991:ABZ

[PSG91] N. Papamichael, E. B. Saff, and J. Gong. Asymptotic behaviour of zeros of Bieberbach polynomials. Journal of Computational and Applied Mathematics, 34(3): 325–342, April 26, 1991.

CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
037704279190093Y.

Petkovic:1995:ZFM

[PT95] M. S. Petković and S. Tricković. On zero-finding methods of the fourth order. Journal of Computational and Applied Mathematics, 64(3):291–294, December 20, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042795001484.

Papageorgiou:1996:CRK

[PT96] G. Papageorgiou and Ch.

Tsitouras. Continuous Runge–Kutta(–Nyström) methods with reduced phaseerrors. Journal of Computational and Applied Mathematics, 69(1):1–11, April 30, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042795000275.

Paluszny:1999:CCB

[PTP99] Marco Paluszny, Francisco Tovar, and Richard R. Patterson. G^{2} composite cubic Bézier curves. Journal of Computational and Applied Mathematics, 102(1):49–71, February 15, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042798002088.

Paris:1992:EIA

[PW92] R. B. Paris and A. D. Wood. Exponentiallyimproved asymptotics for the gamma function. Journal of Computational and Applied Mathematics, 41 (1–2):135–143, August 20, 1992. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        037704279290243Q.

Poliakoff:1999:ACF

[PWT99] Janet F. Poliakoff, Yew Kee Wong, and Peter D. Thomas. An automated curve fairing algorithm for cubic Bspline curves. Journal of Computational and Applied Mathematics, 102(1):73–85, February 15, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704279800209X.

Pu:1998:IGN

[PZ98] Dingguo Pu and Jianzhong Zhang. Inexact generalized Newton methods for second order c-differentiable optimization. Journal of Computational and Applied Mathematics, 93(2):107– 122, July 13, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042798000648.

Qu:1997:CMS

[QA97] Ruibin Qu and Ravi P.

Agarwal. A collocation method for solving a class of singular nonlinear twopoint boundary value problems. Journal of Computational and Applied Mathematics, 83(2):147–163, October 7, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042797000708.

Qiu:1999:ADA

[QS99] Y. Qiu and D. M. Sloan.

Analysis of difference approximations to a singularly perturbed two-point boundary value problem on an adaptively generated grid. Journal of Computational and Applied Mathematics, 101(1–2):1–25, January 15, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042798001368.

Rai:1996:APM

[RA96] A. Sesappa Rai and U. Ananthakrishnaiah. Additive parameters methods for the numerical integration of y′′ = f(t, y, y′). Journal of Computational and Applied Mathematics, 67(2):271– 276, March 29, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042794001278.

Rai:1997:OMH

[RA97] A. Sesappa Rai and U. Ananthakrishnaiah. Obrechkoff methods having additional parameters for general second-order differential equations. Journal of Computational and Applied Mathematics, 79(2):167– 182, March 17, 1997. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S037704279600132X.

Rabinowitz:1990:NIB

[Rab90] Philip Rabinowitz. Numerical integration based on approximating splines. Journal of Computational and Applied Mathematics, 33(1):73–83, December 17, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279090257Z.

Rabinowitz:1994:NPS

[Rab94] Paul H. Rabinowitz. A note on periodic solutions of prescribed energy for singular Hamiltonian systems. Journal of Computational and Applied Mathematics, 52(1–3):147–154, July 20, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042794903549.

Rabut:1996:IRS

[Rab96] Christophe Rabut. Interpolation with radially symmetric thin plate splines. Journal of Computational and Applied Mathematics, 73(1–2):241–256, October 5, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042796000465.

Radoux:1991:SCI

[Rad91] Christian Radoux. Some congruences and identities for Gauss polynomials: a noncommutative technique. Journal of Computational and Applied Mathematics, 37(1–3):19–25, November 18, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279190102P.

Radoux:1995:SAP

[Rad95] Christian Radoux. Some arithmetical properties of Apéry numbers. Journal of Computational and Applied Mathematics, 64(1–2):11– 19, November 30, 1995. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0377042796809113.

Radoux:1997:AP

[Rad97] Christian Radoux. Acknowledgment of priority. Journal of Computational and Applied Mathematics, 80(1): 169, April 14, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042797000277.

Rafajlowicz:1994:FAG

[Raf94] Ewaryst Rafajlowicz. Fast algorithm for generating Bernstein–Bézier polynomials. Journal of Computational and Applied Mathematics, 51(3):279–292, June 30, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        037704279200110U.

Rahman:1996:SGF

[Rah96] Mizan Rahman. Some generating functions for the associated Askey–Wilson polynomials. Journal of Computational and Applied Mathematics, 68(1–2):287– 296, April 22, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042795002499.

Ramage:1999:MPS

[Ram99]  A. Ramage.  A multi-

grid preconditioner for stabilised discretisations of advection–diffusion problems. Journal of Computational and Applied Mathematics, 110(1):187–203, October 15, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042799002344.

Randez:1992:SNI

[Rán92] L. Rández. On the simultaneous numerical integration of an IVP and its associated variational equation. Journal of Computational and Applied Mathematics, 39(1):103–108, February 28, 1992. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279290226N.

Ranga:1996:COP

[Ran96] A. Sri Ranga. Companion orthogonal polynomials. Journal of Computational and Applied Mathematics, 75(1):23–33, November 12, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704279600074X.

Rathore:1994:TED

[Rat94] R. K. S. Rathore. Total error in the discrete convolution backprojection algorithm in computerized tomography. Journal of Computational and Applied Mathematics, 54(1):79–97, September 20, 1994. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0377042794903964.

Ronveaux:1990:FOD

[RBDM90] A. Ronveaux, S. Belmehdi, J. Dini, and P. Maroni. Fourth-order differential equation for the co-modified of semi-classical orthogonal polynomials. Journal of Computational and Applied Mathematics, 29(2):225– 231, February 15, 1990. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0377042790903598.

Ronveaux:1991:ORS

[RBM91] A. Ronveaux, S. Belmehdi, and F. Marcellán. Orthogonality with respect to the sum of two semiclassical regular linear forms. Journal of Computational and Applied Mathematics, 37(1–3):265–272, November 18, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042791901232.

Riemslagh:1996:MMU

[RD96] Kris Riemslagh and Erik Dick. A multigrid method with unstructured adaptive grids for steady Euler equations. Journal of Computational and Applied Mathematics, 67(1):73–93, February 20, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042794001170.

Razzaghi:1994:NSC

[RE94] Mohsen Razzaghi and Gamal Elnagar. Numerical solution of the controlled Duffing oscillator by the pseudospectral method. Journal of Computational and Applied Mathematics, 56 (3):253–261, December 30, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        0377042794900817.

Reczek:1990:GMF

[Rec90]  Krzysztof Reczek.  On

the growth of meromorphic functions. Journal of Computational and Applied Mathematics, 29(1):1– 7, January 10, 1990. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0377042790901897.

Reemtsen:1994:SOA

[Ree94] Rembert Reemtsen. Some outer approximation methods for semi-infinite optimization problems. Journal of Computational and Applied Mathematics, 53(1): 87–108, July 29, 1994. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/037704279200122P.

Reginska:1995:SHE

[Reg95] Teresa Regińska. Sideways heat equation and wavelets. Journal of Computational and Applied Mathematics, 63(1–3):209–214, November 20, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042795000739.

Renault:1996:NAC

[Ren96]  O. Renault.  A new

algorithm for computing orthogonal polynomials. Journal of Computational and Applied Mathematics, 75(2):231–248, November 28, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042796000647.

Ren:1997:ARI

[Ren97a] Yuhe Ren. The approximate regularization of index-2 differential-algebraic problems. Journal of Computational and Applied Mathematics, 78(2):301–310, February 17, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042796001501.

Renaut:1997:SCP

[Ren97b] Rosemary Renaut. Stability of a Chebychev pseudospectral solution of the wave equation with absorbing boundaries. Journal of Computational and Applied Mathematics, 87 (2):243–259, December 23, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042797001921.

Rensaa:1999:OTU

[Ren99] Ragnhild Johanne Rensaa. On omitted tuples for univalent functions. Journal of Computational and Applied Mathematics, 105(1–2):445– 452, May 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042799000114.

Ramamurthy:1999:VDMb

[RF99a] Rajesh Ramamurthy and Rida T. Farouki. Voronoi diagram and medial axis algorithm for planar domains with curved boundaries

— II: Detailed algorithm description. Journal of Computational and Applied Mathematics, 102(2):253– 277, February 26, 1999. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042798002234.

Ramamurthy:1999:VDMa

[RF99b] Rajesh Ramamurthy and Rida T. Farouki. Voronoi diagram and medial axis algorithm for planar domains with curved boundaries. I. Theoretical foundations. Journal of Computational and Applied Mathematics, 102(1):119–141, February 15, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042798002118.

Ronveaux:1998:FOD

[RGZA98] A. Ronveaux, E. Godoy, A. Zarzo, and I. Area. Fourth-order difference equation for the first associated of classical discrete orthogonal polynomials. Journal of Computational and Applied Mathematics, 90(1):45–50, April 6, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042798000156.

Roldan:1999:IMD

[RH99] Teo Roldán and Inmaculada Higueras. IRK methods for DAE: starting algorithms. Journal of Computational and Applied Mathematics, 111(1–2):77–92, November 15, 1999. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042799001338.

Rhandi:1992:DPE

[Rha92] Abdelaziz Rhandi. Dyson– Phillips expansion and unbounded perturbations of linear c_{0}-semigroups. Journal of Computational and Applied Mathematics, 44(3):339–349, December 24, 1992. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279290005I.

Rawy:1998:NSN

[RIG98] E. K. Rawy, Labib Iskandar, and A. F. Ghaleb. Numerical solution for a nonlinear, one-dimensional problem of thermoelasticity. Journal of Computational and Applied Mathematics, 100(1):53–76, November 30, 1998. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042798001344.

Rinott:1994:NAR

[Rin94]  Yosef Rinott.  On nor-

mal approximation rates for certain sums of dependent random variables. Journal of Computational and Applied Mathematics, 55(2):135–143, November 21, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042794900167.

Ripley:1990:TPN

[Rip90] B. D. Ripley. Thoughts on pseudorandom number generators. Journal of Computational and Applied Mathematics, 31(1):

153–163, July 24, 1990. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042790903462.

Rivara:1991:LMM

[Riv91] María-Cecilia Rivara. Local modification of meshes for adaptive and/or multigrid finite-element methods. Journal of Computational and Applied Mathematics, 36(1):79–89, August 20, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279190227B.

Reddy:1999:UFP

[RJ99] Muni V. Reddy and Stephen Joe. The ultratriangular form for prime-power lattice rules. Journal of Computational and Applied Mathematics, 104(1):49–61, April 15, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704279800260X.

Ripley:1990:ISM

[RK90] B. D. Ripley and M. D.

Kirkland. Iterative simulation methods. Journal of Computational and Applied Mathematics, 31(1): 165–172, July 24, 1990. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0377042790903473.

Rasio:1999:SPH

[RL99] Frederic A. Rasio and James C. Lombardi, Jr. Smoothed particle hydrodynamics calculations of stellar interactions. Journal of Computational and Applied Mathematics, 109(1–2):213– 230, September 30, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042799001594.

Riera:1995:COS

[RM95a] Mario Pérez Riera and Juan L. Varona Malumbres. On completeness of orthogonal systems and Dirac deltas. Journal of Computational and Applied Mathematics, 58(2):225– 231, March 30, 1995. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0377042793E0272N.

Ronveaux:1995:SSZ

[RM95b] André Ronveaux and Martin E. Muldoon. Stieltjes sums for zeros of orthogonal polynomials. Journal of Computational and Applied Mathematics, 57(1–2):261– 269, February 20, 1995. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042793E0250P.

Razali:1997:NCM

[RNM97] M. R. M. Razali, M. Z.

Nashed, and A. H. M. Murid. Numerical conformal mapping via the Bergman kernel. Journal of Computational and Applied Mathematics, 82(1–2):333–350, September 15, 1997. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042797000915.

Roberts:1996:CRV

[Rob96] D. E. Roberts. On the convergence of rows of vector Padé approximants. Journal of Computational and Applied Mathematics, 70(1): 95–109, June 14, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042795001336.

Roberts:1998:AEN

[Rob98] Catherine A. Roberts. Analysis of explosion for nonlinear Volterra equations. Journal of Computational and Applied Mathematics, 97(1–2):153–166, September 24, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042798001083.

Roberts:1999:RVC

[Rob99] D. E. Roberts. On a representation of vector continued fractions. Journal of Computational and Applied Mathematics, 105(1–2):453– 466, May 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042799000345.

Rodrigues:1999:PND

[Rod99] José Alberto Rodrigues. Preconditioners for nonconforming domain decomposition methods. Journal of Computational and Applied Mathematics, 111(1–2):227– 237, November 15, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042799001454.

Rolle:1999:MVQ

[Rol99] Jean-Daniel Rolle. Minimum variance quadratic unbiased estimators as a tool to identify compound normal distributions. Journal of Computational and Applied Mathematics, 102(2):279– 285, February 26, 1999. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042798002283.

Ronveaux:1990:QOD

[Ron90]  A. Ronveaux.  Quasi-

orthogonality and differential equations. Journal of Computational and Applied Mathematics, 29(2):243– 247, February 15, 1990. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042790903613.

Roos:1990:GUC

[Roo90a] H.-G. Roos. Global uniformly convergent schemes for a singularly perturbed boundary-value problem using patched base splinefunctions. Journal of Computational and Applied Mathematics, 29(1):69–77, January 10, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042790901967.

Roose:1990:BRP

[Roo90b] D. Roose. Book review:

Parallel and distributed algorithms: M. Cosnard, Y. Robert, P. Quintan and M. Raynal (Eds.): North-Holland, 1989, 344 pages, Price: U.S. $92.00, ISBN 0-444-87367-8. Journal of Computational and Applied Mathematics, 32(1–2):N6, November 26, 1990. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0377042790904434.

Roos:1994:TWG

[Roo94] Hans-Görg Roos. Ten ways to generate the Il’in and related schemes. Journal of Computational and Applied Mathematics, 53(1):43–59, July 29, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279200124R.

Rosenberg:1995:VFP

[Ros95a] J. Rosenberg. Variational formulation of the problems of mechanics and its matrix analogy. Journal of Computational and Applied Mathematics, 63(1–3):307–311, November 20, 1995. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0377042795000860.

Rosler:1995:TCS

[Rös95b] Margit Rösler. Trigonometric convolution structures on Z derived from Jacobi polynomials. Journal of Computational and Applied Mathematics, 65(1–3):357–368, December 29, 1995. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042795001220.

Rovder:1992:AOB

[Rov92] Jozef Rovder. Asymptotic and oscillatory behaviour of solutions of a linear differential equation. Journal of Computational and Applied Mathematics, 41(1–2): 41–47, August 20, 1992. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279290236Q.

Rovder:1996:MST

[Rov96] Jozef Rovder. On monotone solution of the thirdorder differential equation. Journal of Computational and Applied Mathematics, 66(1–2):421–432, January 31, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042795001654.

Renaut:1996:RAO

[RP96] R. A. Renaut and J. S.

Parent. Rational approximations to 1/\sqrt{1 − s^{2}} oneway wave equations and absorbing boundary conditions. Journal of Computational and Applied Mathematics, 72(2):245–259, August 13, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042795002766.

Rachev:1994:E

[RR94a] Svetlozar T. Rachev and Werner Römisch. Editorial. Journal of Computational and Applied Mathematics, 56(1–2):2, December 20, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042794903751.

Rachev:1994:P

[RR94b] Svetlozar T. Rachev and Werner Römisch. Preface. Journal of Computational and Applied Mathematics, 56(1–2):1, December 20, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042794903743.

Rordriguez:1990:NSF

[RS90] G. Rordriguez and S. Seatzu. Numerical solution of the finite moment problem in a reproducing kernel Hilbert space. Journal of Computational and Applied Mathematics, 33(3):233–244, December 31, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042705800019.

Rabinowitz:1992:IPI

[RS92] Philip Rabinowitz and William E. Smith. Interpolatory product integration in the presence of singularities: L_{2} theory. Journal of Computational and Applied Mathematics, 39(1):79–87, February 28, 1992. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279290223K.

Rundell:1994:DPB

[RS94] William Rundell and Paul Sacks. On the determination of potentials without bound state data. Journal of Computational and Applied Mathematics, 55(3):325–347, November 30, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279490037X.

Raviart:1995:AMM

[RS95] Pierre-Arnaud Raviart and Eric Sonnendrücker. Approximate models for the Maxwell equations. Journal of Computational and Applied Mathematics, 63(1–3): 69–81, November 20, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042795000585.

Ren:1999:NBP

[RS99] Y. Ren and A. Spence. A note on bifurcation points for index-1 differentialalgebraic problems. Journal of Computational and Applied Mathematics, 110(1): 1–14, October 15, 1999. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042799001910.

Rohwer:1991:LMR

[RT91] C. H. Rohwer and L. M. Toerien. Locally monotone robust approximation of sequences. Journal of Computational and Applied Mathematics, 36(3):399– 408, September 24, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
037704279190019G.

Rousseeuw:1995:FCH

[RTK95] P. J. Rousseeuw, E. Trauwaert, and L. Kaufman. Fuzzy clustering with high contrast. Journal of Computational and Applied Mathematics, 64(1–2):81–90, November 30, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042795000089.

Rump:1995:VSL

[Rum95] Siegfried M. Rump. Verified solution of large systems and global optimization problems. Journal of Computational and Applied Mathematics, 60(1–2):201– 218, June 20, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279400092F.

Ruotsalainen:1994:CCM

[Ruo94] K. Ruotsalainen. On the convergence of the collocation method for nonlinear boundary integral equations. Journal of Computational and Applied Mathematics, 50(1–3):471–483, May 20, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042794903212.

Ranga:1994:CTQ

[RV94] A. Sri Ranga and Marisa Veiga. A Chebyshev-type quadrature rule with some interesting properties. Journal of Computational and Applied Mathematics, 51(2): 263–265, June 20, 1994. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0377042794000417.

Rosler:1998:BPA

[RV98] Margit Rösler and Michael Voit. Biorthogonal polynomials associated with reflection groups and a formula of Macdonald. Journal of Computational and Applied Mathematics, 99(1–2):337–351, November 16, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704279800168X.

Rice:1997:AND

[RVY97] J. R. Rice, E. A. Vavalis, and Daoqi Yang. Analysis of a nonoverlapping domain decomposition method for elliptic partial differential equations. Journal of Computational and Applied Mathematics, 87(1):11–19, December 18, 1997. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042797001726.

Rodriguez:1992:PTA

[RW92] Francisco Cala Rodriguez and Hans Wallin. Padé-type approximants and a summability theorem by eiermann. Journal of Computational and Applied Mathematics, 39(1):15–21, February 28, 1992. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279290218M.

Riha:1997:EAL

[RW97] W. O. Riha and J. Walker.

An efficient algorithm for the Lagrangean dual of nonlinear knapsack problems with additional nested constraints. Journal of Computational and Applied Mathematics, 78(1):9–18, February 3, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704279600101X.

Rozloznik:1998:SIG

[RW98] Miroslav Rozlozník and Rüdiger Weiss. On the stable implementation of the generalized minimal error method. Journal of Computational and Applied Mathematics, 98(1):49–62, October 5, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042798001137.

Riffert:1999:JSE

[RW99] Harald Riffert and Klaus Werner. JCAM special edition: Computational astrophysics: Preface. Journal of Computational and Applied Mathematics, 109(1–2):ix–x, September 30, 1999. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042799002058.

Riaza:1999:ACI

[RZ99] Ricardo Riaza and Pedro J. Zufiria. Adaptive cellular integration of linearly implicit differential equations. Journal of Computational and Applied Mathematics, 111(1–2):305–317, November 15, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704279900151X.

Ronveaux:1998:TLA

[RZAG98] A. Ronveaux, A. Zarzo, I. Area, and E. Godoy. Transverse limits in the Askey tableau. Journal of Computational and Applied Mathematics, 99(1–2):327– 335, November 16, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042798001678.

Ronveaux:1995:FOD

[RZG95a] A. Ronveaux, A. Zarzo, and E. Godoy. Fourthorder differential equations satisfied by the generalized co-recursive of all classical orthogonal polynomials. A study of their distribution of zeros. Journal of Computational and Applied Mathematics, 59(3): 295–328, May 30, 1995. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279400006M.

Ronveaux:1995:RRC

[RZG95b] A. Ronveaux, A. Zarzo, and E. Godoy. Recurrence relations for connection coefficients between two families of orthogonal polynomials. Journal of Computational and Applied Mathematics, 62(1):67–73, August 31, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042794000798.

Ren:1999:STS

[RZQ99] Yuhe Ren, Bo Zhang, and Hong Qiao. A simple Taylorseries expansion method for a class of second kind integral equations. Journal of Computational and Applied Mathematics, 110(1):15–24, October 15, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042799001922.

Szidarovszky:1994:CMC

[SA94] Ferenc Szidarovszky and Ioannis K. Argyros. On the convergence of modified contractions. Journal of Computational and Applied Mathematics, 55 (2):183–189, November 21, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042794900191.

Sheng:1997:SPB

[SA97] Qin Sheng and Ravi P. Agarwal. Solutions of npoint boundary value problems associated with nonlinear summary difference equations. Journal of Computational and Applied Mathematics, 80(1):49–70, April 14, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042796001719.

Saadallah:1992:RGM

[Saa92] Adel F. Saadallah. A rational gradient model for minimization. Journal of Computational and Applied Mathematics, 39(3):277– 286, May 8, 1992. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279290204B.

Saadallah:1994:QNM

[Saa94] Adel F. Saadallah. A quasi-Newton method using a nonquadratic model. Journal of Computational and Applied Mathematics, 51(3): 327–338, June 30, 1994. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279200115P.

Sablonniere:1993:DBB

[Sab93] Paul Sablonnière. Discrete Bernstein bases and Hahn polynomials. Journal of Computational and Applied Mathematics, 49(1–3):233– 241, December 31, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042793901555.

Sablonniere:1995:CFN

[Sab95] Paul Sablonnière. Comparison of four nonlinear transforms on some classes of logarithmic fixed point sequences. Journal of Computational and Applied Mathematics, 62(1):103–127, August 31, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042796800136.

Sablonniere:1996:QIA

[Sab96] Paul Sablonnière. Quasiinterpolants associated with H-splines on a threedirection mesh. Journal of Computational and Applied Mathematics, 66(1–2):433– 442, January 31, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0377042795001689.

Sacco:1996:EFS

[Sac96] Riccardo Sacco. Exponentially fitted shape functions for advection-dominated flow problems in two dimensions. Journal of Computational and Applied Mathematics, 67(1):161– 165, February 20, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0377042795001492.

Sadok:1990:AHT

[Sad90]  Hassane Sadok.  About

Henrici’s transformation for accelerating vector sequences. Journal of Computational and Applied Mathematics, 29(1):101–110, January 10, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279090199A.

Sanchez-Avila:1996:NPA

[SÁFV96] C. Sánchez-Ávila and A. R.

Figueiras-Vidal. New POCS algorithms for regularization of inverse problems. Journal of Computational and Applied Mathematics, 72(1): 21–39, July 30, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279500243X.

Sakai:1999:NNS

[Sak99] Katsuhiro Sakai. A nonoscillatory numerical scheme based on a general solution of 2-D unsteady advection–diffusion equations. Journal of Computational and Applied Mathematics, 108(1–2):145–156, August 15, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042799001077.

Salam:1993:GA

[Sal93] A. Salam. On a generalization of the ϵ-algorithm. Journal of Computational and Applied Mathematics, 46(3):455–464, June 22, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042793900419.

Santi:1996:ECP

[San96] E. Santi. On the evaluation of Cauchy principal value integrals by rules based on quasi-interpolating splines. Journal of Computational and Applied Mathematics, 71(1):1–14, July 10, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042795002138.

Sastry:1997:GUI

[Sas97] Venkat V. S. S. Sastry. Graphical user interfaces for numerical libraries. Journal of Computational and Applied Mathematics, 82(1–2):351–366, September 15, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042797000927.

Sauter:1991:IMF

[Sau91] Stefan Sauter. The ILU method for finite-element discretizations. Journal of Computational and Applied Mathematics, 36(1):91–106, August 20, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279190228C.

Sayas:1997:FDG

[Say97] Francisco-Javier Sayas. Fully discrete Galerkin methods for systems of boundary integral equations. Journal of Computational and Applied Mathematics, 81(2): 311–331, July 8, 1997. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042797000691.

Sayas:1998:GEM

[Say98] Francisco-Javier Sayas. A generalized Euler–Maclaurin formula on triangles. Journal of Computational and Applied Mathematics, 93(2): 89–93, July 13, 1998. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042798000491.

Sobol:1994:CMS

[SB94] Ilya M. Sobol’ and Sergey G.

Bakin. On the crude multidimensional search. Journal of Computational and Applied Mathematics, 56(3):283–293, December 30, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042794900833.

Strawn:1996:NSH

[SB96] Roger C. Strawn and Rupak Biswas. Numerical simulations of helicopter aerodynamics and acoustics. Journal of Computational and Applied Mathematics, 66(1–2):471–483, January 31, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042795001719.

Sequeira:1999:FEA

[SB99] Adélia Sequeira and Margarida Baía. A finite element approximation for the steady solution of a second-grade fluid model. Journal of Computational and Applied Mathematics, 111(1–2):281–295, November 15, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042799001491.

Su:1998:GAB

[SBK98] Yangfeng Su, Amit Bhaya, and Eugenius Kaszkurewicz. A general asynchronous block iterative model with related convergence conditions. Journal of Computational and Applied Mathematics, 91(2):261–273, May 4, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042798000533.

Song:1995:DRC

[SC95a] Yu Song and Maria Calzada.

A dictionary random choice method for turbulent combustion. Journal of Computational and Applied Mathematics, 58(1):27–42, March 20, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042793E0256L.

Suarez:1995:GAF

[SC95b] M. C. Suárez and A. Cachafeiro.

A generalization of the associated functional to the Lebesgue measure. Journal of Computational and Applied Mathematics, 57(1–2):283–291, February 20, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042793E0252H.

Scapolla:1994:TOM

[Sca94] Terenzio Scapolla. A thirdorder mixed finite-element method for the numerical solution of the biharmonic problem. Journal of Computational and Applied Mathematics, 50(1–3):497–507, May 20, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042794903239.

Scapolla:1995:ALR

[Sca95] Terenzio Scapolla. Avoiding locking in Reissner–Mindlin plates with high-order hierarchic bubble plus finite elements. Journal of Computational and Applied Mathematics, 63(1–3):291–299, November 20, 1995. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0377042795000844.

Schaefer:1990:PBI

[Sch90] Mark J. Schaefer. A polynomial based iterative method for linear parabolic equations. Journal of Computational and Applied Mathematics, 29(1):35–50, January 10, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042790901934.

Schmidt:1993:GLP

[Sch93] Asmus L. Schmidt. Generalized q-Legendre polynomials. Journal of Computational and Applied Mathematics, 49(1–3):243–249, December 31, 1993. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042793901566.

Schaefer:1994:POU

[Sch94a] Mark J. Schaefer. Precise optimization using range arithmetic. Journal of Computational and Applied Mathematics, 53(3):341– 351, August 30, 1994. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042794900620.

Schmidt:1994:VNF

[Sch94b] Dieter Schmidt. Versal normal form of the Hamiltonian function of the restricted problem of three bodies near L _{4}. Journal of Computational and Applied Mathematics, 52(1–3):155–176, July 20, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042794903557.

Schott:1994:AGCa

[Sch94c] Dieter Schott. About geometrical convergence of general iterative methods applied to nonunique solvable convex problems, part I. Journal of Computational and Applied Mathematics, 54(1):1–14, September 20, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042794903905.

Schott:1994:AGCb

[Sch94d] Dieter Schott. About geometrical convergence of general iterative methods applied to nonunique solvable convex problems, part II. Journal of Computational and Applied Mathematics, 54(2):133–150, September 30, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042794901740.

Schultz:1994:SCS

[Sch94e] Rüdiger Schultz. Strong convexity in stochastic programs with complete recourse. Journal of Computational and Applied Mathematics, 56(1–2):3–22, December 20, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279490376X.

Schwandt:1994:IAD

[Sch94f] Hartmut Schwandt. An interval arithmetic domain decomposition method for a class of elliptic PDEs on nonrectangular domains. Journal of Computational and Applied Mathematics, 50(1–3):509–521, May 20, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042794903247.

Schaefer:1996:VCM

[Sch96a] Mark J. Schaefer. Verification of constrained minima. Journal of Computational and Applied Mathematics, 67(2):195–205, March 29, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042794001324.

Schira:1996:RTA

[Sch96b] Thomas Schira. The remainder term for analytic functions of Gauss–Lobatto quadratures. Journal of Computational and Applied Mathematics, 76(1–2):171– 193, December 17, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042796001008.

Schindler:1997:ESP

[Sch97] W. Schindler. On the efficient simulation of a particular class of random rotations relevant to computer graphics. Journal of Computational and Applied Mathematics, 81(1):

107–114, June 1, 1997. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042797000290.

Schoutens:1998:LSI

[Sch98]  Wim Schoutens.  Lévy–

Sheffer and IID–Sheffer polynomials with applications to stochastic integrals. Journal of Computational and Applied Mathematics, 99(1–2):365–372, November 16, 1998. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042798001708.

Scherer:1999:NMT

[Sch99] G. Scherer. On the numerical modelling of the transitional flow in rarefied gases. Journal of Computational and Applied Mathematics, 103(1):165– 173, March 15, 1999. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042798002507.

Scitovski:1994:SNL

[Sci94] Rudolf Scitovski. A special nonlinear least-squares problem. Journal of Computational and Applied Mathematics, 53(3):323–331, August 30, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042794900604.

Schein:1998:NSS

[SD98] O. Schein and G. Denk. Numerical solution of stochastic differential-algebraic equations with applications to transient noise simulation of microelectronic circuits. Journal of Computational and Applied Mathematics, 100(1):77–92, November 30, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042798001381.

Shenkman:1999:NBS

[SDL99] P. Shenkman, N. Dyn, and D. Levin. Normals of the butterfly subdivision scheme surfaces and their applications. Journal of Computational and Applied Mathematics, 102(1):157– 180, February 15, 1999. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042798002131.

Santos:1999:CFV

[SdO99] J. Santos and P. de Oliveira.

A converging finite volume scheme for hyperbolic conservation laws with source terms. Journal of Computational and Applied Mathematics, 111(1–2):239–251, November 15, 1999. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042799001466.

Steelant:1997:MMC

[SDP97] J. Steelant, E. Dick, and S. Pattijn. Multigrid methods for compressible Navier–Stokes equations in low-speed flows. Journal of Computational and Applied Mathematics, 82(1–2):379–388, September 15, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042797000800.

Simos:1994:RKN

[SDS94] T. E. Simos, E. Dimas, and A. B. Sideridis. A Runge–Kutta–Nyström method for the numerical integration of special second-order periodic initial-value problems. Journal of Computational and Applied Mathematics, 51(3): 317–326, June 30, 1994. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279200114O.

Singstad:1999:ICS

[SE99] Kristoffer Singstad and Terje O. Espelid. Integrating composed singularities using nonuniform subdivision and extrapolation. Journal of Computational and Applied Mathematics, 112(1–2):291–308, November 30, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042799002277.

Sedogbo:1990:CAS

[Sed90] Guy Antoine Sedogbo. Convergence acceleration of some logarithmic sequences. Journal of Computational and Applied Mathematics, 32(1–2):253–260, November 26, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042790904353.

Segeth:1995:GAP

[Seg95] Karel Segeth. Grid adjustment for parabolic systems based on a posteriori error estimates. Journal of Computational and Applied Mathematics, 63(1–3):349– 355, November 20, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        0377042795000631.

Senhadji:1999:CNS

[Sen99] M. N. Senhadji. On condition numbers of some quasilinear transformations. Journal of Computational and Applied Mathematics, 104 (1):1–17, April 15, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042798001940.

Sezgin:1992:SCC

[Sez92] Fatin Sezgin. Some comments on computer implementation of random number generators. Journal of Computational and Applied Mathematics, 39(3): 383–386, May 8, 1992. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/037704279290212G.

Sidi:1990:QDT

[SF90] Avram Sidi and William F. Ford. Quotient-difference type generalizations of the power method and their analysis. Journal of Computational and Applied Mathematics, 32(1–2):261–272, November 26, 1990. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042790904364.

Sagara:1991:HMN

[SF91] Nobuko Sagara and Masao Fukushima. A hybrid method for the nonlinear least squares problem with simple bounds. Journal of Computational and Applied Mathematics, 36(2):149– 157, August 27, 1991. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279190023D.

Sharp:1992:CDT

[SF92a] P. W. Sharp and J. M.

Fine. A contrast of direct and transformed Nyström pairs. Journal of Computational and Applied Mathematics, 42(3):293–308, October 30, 1992. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042792900829.

Sharp:1992:SNP

[SF92b] P. W. Sharp and J. M. Fine.

Some Nyström pairs for the general second-order initialvalue problem. Journal of Computational and Applied Mathematics, 42(3):279– 291, October 30, 1992. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0377042792900818.

Stoyanov:1994:AEI

[SFB94] B. J. Stoyanov, R. A. Farrell, and J. F. Bird. Asymptotic expansions of integrals of two Bessel functions via the generalized hypergeometric and Meijer functions. Journal of Computational and Applied Mathematics, 50(1–3):533–543, May 20, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042794903263.

Schweizer:1999:DVF

[SFGF^{+}99] W. Schweizer, P. Faßbinder, R. González-Ferez, M. Braun, S. Kulla, and M. Stehle. Discrete variable and finite element techniques applied to simple atomic systems. Journal of Computational and Applied Mathematics, 109(1–2):95–122, September 30, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042799001557.

Sugimoto:1995:PRM

[SFI95] Takanobu Sugimoto, Masao Fukushima, and Toshihide Ibaraki. A parallel relaxation method for quadratic programming problems with interval constraints. Journal of Computational and Applied Mathematics, 60(1–2): 219–236, June 20, 1995. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279400093G.

Simon:1999:CGB

[SFW99] Peter L. Simon, Henrik Farkas, and Mária Wittmann. Constructing global bifurcation diagrams by the parametric representation method. Journal of Computational and Applied Mathematics, 108(1–2):157– 176, August 15, 1999. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042799001089.

Shuo:1994:ACL

[SG94] Tang Shuo and Zhu Gongqin.

On the acceleration convergence of limit periodic continued fractions by the T_{+m} transformation. Journal of Computational and Applied Mathematics, 51(2): 267–274, June 20, 1994. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042794000425.

Simoncini:1996:HBG

[SG96] V. Simoncini and E. Gallopoulos. A hybrid block GMRES method for nonsymmetric systems with multiple right-hand sides. Journal of Computational and Applied Mathematics, 66(1–2):457–469, January 31, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042795001980.

Simeon:1994:NTM

[SGFR94] B. Simeon, F. Grupp, C. Führer, and P. Rentrop. A nonlinear truck model and its treatment as a multibody system. Journal of Computational and Applied Mathematics, 50(1–3):523– 532, May 20, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042794903255.

Sen:1996:FDM

[SH96] Rabindra Nath Sen and Md. Bellal Hossain. Finite difference methods for certain singular two-point boundary value problems. Journal of Computational and Applied Mathematics, 70(1):33–50, June 14, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.

com/science/article/pii/
0377042795001468.

Shampine:1993:ICM

[Sha93] Lawrence F. Shampine. Illconditioned matrices and the integration of stiff ODEs. Journal of Computational and Applied Mathematics, 48(3):279–292, November 30, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042793900257.

Shalaby:1994:RSI

[Sha94a]  M. A. Shalaby.  On

the real symmetric inverse eigenvalue problem. Journal of Computational and Applied Mathematics, 56(3):331–340, December 30, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042794900876.

Shalaby:1994:PPS

[Sha94b] M. A. Shalaby. A parallel processed scheme for the eigenproblem of positive definite matrices. Journal of Computational and Applied Mathematics, 54(1):99–106, September 20, 1994. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042794903972.

Shah:1995:NME

[Sha95] A. A. Shah. A new method for eigenvalues of nonsymmetric matrices. Journal of Computational and Applied Mathematics, 58(1):83–88, March 20, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042793E0263L.

Shapira:1997:AIC

[Sha97a] Yair Shapira. Algebraic interpretation of continued fractions. Journal of Computational and Applied Mathematics, 78(1):3– 8, February 3, 1997. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042796000945.

Shapira:1997:NMC

[Sha97b] Yair Shapira. Note on the multigrid W -cycle. Journal of Computational and Applied Mathematics, 85(2):351–353, November 12, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042797001428.

Shibutani:1999:IDA

[Shi99] Norimasa Shibutani. Integrable discretization and its applications. Journal of Computational and Applied Mathematics, 108(1–2):57– 73, August 15, 1999. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042799001004.

Shu:1996:SGC

[Shu96] H. Z. Shu. The superconvergence of a Galerkin collocation method for first kind boundary integral equations on smooth curves. Journal of Computational and Applied Mathematics, 72(1):1– 20, July 30, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042795002421.

Somersalo:1992:LIB

[SIC92] Erkki Somersalo, David Isaacson, and Margaret Cheney. A linearized inverse boundary value problem for Maxwell’s equations. Journal of Computational and Applied Mathematics, 42(1):123–136, September 8, 1992. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279290167V.

Sidi:1990:QCA

[Sid90]  Avram Sidi.  Quantita-

tive and constructive aspects of the generalized Koenig’s and de Montessus’s theorems for Padé approximants. Journal of Computational and Applied Mathematics, 29(3):257– 291, March 15, 1990. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279090012O.

Sidi:1991:EIM

[Sid91] Avram Sidi. Efficient implementation of minimal polynomial and reduced rank extrapolation methods. Journal of Computational and Applied Mathematics, 36(3):305–337, September 24, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279190013A.

Sidi:1997:CII

[Sid97] Avram Sidi. Computation of infinite integrals involving Bessel functions of arbitrary order by the Dtransformation. Journal of Computational and Applied Mathematics, 78(1):125– 130, February 3, 1997. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042796001367.

Sidi:1999:FCS

[Sid99] Avram Sidi. Further convergence and stability results for the generalized Richardson extrapolation process GREP ^{(1)} with an application to the D (1)-transformation for infinite integrals. Journal of Computational and Applied Mathematics, 112(1–2):269– 290, November 30, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042799902261.

Sifi:1998:GWP

[Sif98] Mohamed Sifi. Generalized wavelet packet associated with Laguerre functions. Journal of Computational and Applied Mathematics, 99(1–2):353–364, November 16, 1998. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042798001691.

Simos:1990:FSM

[Sim90]  T. E. Simos.  A four-

step method for the numerical solution of the Schrödinger equation. Journal of Computational and Applied Mathematics, 30(3): 251–255, July 10, 1990. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0377042790902788.

Simos:1992:ETS

[Sim92] T. E. Simos. Explicit twostep methods with minimal phase-lag for the numerical integration of special second-order initialvalue problems and their application to the onedimensional Schrödinger equation. Journal of Computational and Applied Mathematics, 39(1):89–94, February 28, 1992. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279290224L.

Simos:1994:EFS

[Sim94]  T. E. Simos.  An ex-

plicit four-step phase-fitted method for the numerical integration of secondorder initial-value problems. Journal of Computational and Applied Mathematics, 55(2):125–133, November 21, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042794900159.

Simos:1995:FFS

[Sim95] T. E. Simos. A family of four-step exponentially fitted predictor–corrector methods for the numerical integration of the Schrödinger equation. Journal of Computational and Applied Mathematics, 58(3): 337–344, April 20, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042793E0274P.

Simeon:1996:NSW

[Sim96] B. Simeon. On the numerical solution of a wheel suspension benchmark problem. Journal of Computational and Applied Mathematics, 66(1–2):443–456, January 31, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042796001690.

Simos:1998:AFD

[Sim98] T. E. Simos. An accurate finite difference method for the numerical solution of the Schrödinger equation. Journal of Computational and Applied Mathematics, 91(1): 47–61, April 24, 1998. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042798000144.

Simos:1999:EFE

[Sim99] T. E. Simos. An exponentially fitted eighth-order method for the numerical solution of the Schrödinger equation. Journal of Computational and Applied Mathematics, 108(1–2):177– 194, August 15, 1999. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042799001090.

Singh:1992:NMS

[Sin92] Arindama Singh. A numerical method for singularly perturbed systems of linear two-point boundaryvalue problems using partial decoupling. Journal of Computational and Applied Mathematics, 40(1):55–62, June 12, 1992. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279290041U.

Singer:1994:IFF

[Sin94] Peter Singer. An integrated fractional Fourier transform. Journal of Computational and Applied Mathematics, 54(2):221–237, September 30, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042794901783.

Sinap:1995:GQM

[Sin95]  Ann Sinap.  Gaussian

quadrature for matrix valued functions on the real line. Journal of Computational and Applied Mathematics, 65(1–3):369–385, December 29, 1995. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042795001247.

Sadek:1998:OBC

[SJAM98] I. S. Sadek, L. Jamiiru, and H. A. Al-Mohamad. Optimal boundary control of one-dimensional multi-span vibrating systems. Journal of Computational and Applied Mathematics, 94(1): 39–54, July 31, 1998. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042798800204.

Schmeiser:1990:NCI

[SK90] Bruce Schmeiser and Voratas Kachitvichyanukul. Noninverse correlation induction: guidelines for algorithm development. Journal of Computational and Applied Mathematics, 31(1): 173–180, July 24, 1990. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042790903484.

Sallam:1996:QSC

[SK96] S. Sallam and A. A. Karaballi. A quartic C ^{3}- spline collocation method for solving second-order initial value problems. Journal of Computational and Applied Mathematics, 75 (2):295–304, November 28, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042796000829.

Skrzipek:1992:OPM

[Skr92] M.-R. Skrzipek. Orthogonal polynomials for modified weight functions. Journal of Computational and Applied Mathematics, 41(3):331– 346, August 28, 1992. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/037704279290140S.

Stewart:1996:EEK

[SL96] D. E. Stewart and T. S. Leyk. Error estimates for Krylov subspace approximations of matrix exponentials. Journal of Computational and Applied Mathematics, 72(2):359–369, August 13, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042796000064.

Sanchis-Lozano:1997:SCB

[SL97] Miguel Angel Sanchis-Lozano.

Simple connections between generalized hypergeometric series and dilogarithms. Journal of Computational and Applied Mathematics, 85(2):325–331, November 12, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042797001386.

Santos-Leon:1998:AET

[SL98] J. C. Santos-León. Asymptotic expansions for trapezoidal type product integration rules. Journal of Computational and Applied Mathematics, 91(2): 219–230, May 4, 1998. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042798000211.

Santos-Leon:1999:PRU

[SL99] J. C. Santos-León. Product rules on the unit circle with uniformly distributed nodes. Error bounds for analytic functions. Journal of Computational and Applied Mathematics, 108(1–2):195– 208, August 15, 1999. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042799001107.

Slavyanov:1993:DDR

[Sla93] S. Yu. Slavyanov. A “differential” derivation of the recurrence relations for the classical orthogonal polynomials. Journal of Computational and Applied Mathematics, 49(1–3):251–254, December 31, 1993. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0377042793901577.

Steinacker:1991:UAE

[SLK91] Ulrike Steinacker, C. Leubner, and S. L. Kalla. Uniform asymptotic expansions of a class of integrals with finite endpoints of integration on the same path of steepest descent and with nearby saddle points. Journal of Computational and Applied Mathematics, 35(1–3):297– 301, June 26, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042791902167.

Steinacker:1992:UAE

[SLK92] Ulrike Steinacker, C. Leubner, and S. L. Kalla. Uniform asymptotic expansions of a class of incomplete cylindrical functions. Journal of Computational and Applied Mathematics, 44(1):121–130, December 9, 1992. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042792900564.

Sloan:1991:FPS

[Slo91] D. M. Sloan. Fourier pseudospectral solution of the regularised long wave equation. Journal of Computational and Applied Mathematics, 36(2):159–179, August 27, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279190024E.

Smith:1991:ASE

[Smi91] H. V. Smith. An algebraic study of an extension of a class of quadrature formulae. Journal of Computational and Applied Mathematics, 37(1–3):27– 33, November 18, 1991. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279190103Q.

Smiley:1996:FEM

[Smi96]  M. W. Smiley.  A fi-

nite element method for computing the bifurcation function for semilinear elliptic BVPs. Journal of Computational and Applied Mathematics, 70(2):311– 327, June 28, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042795002081.

Sommeijer:1993:PIR

[Som93] B. P. Sommeijer. Paralleliterated Runge–Kutta methods for stiff ordinary differential equations. Journal of Computational and Applied Mathematics, 45(1–2):151– 168, April 8, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279390271C.

Sugihara:1991:ESL

[SOMF91] Masaaki Sugihara, Yoshio Oyanagi, Masatake Mori, and Seiji Fujino. On the efficiency of a SORlike method suited to vector processors. Journal of Computational and Applied Mathematics, 35(1–3):33– 51, June 26, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279190196Q.

Song:1997:CMM

[Son97]  Yongzhong Song.  On

the convergence of the MAOR method. Journal of Computational and Applied Mathematics, 79(2):299– 317, March 17, 1997. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042797000083.

Song:1999:SEI

[Son99] Yongzhong Song. Semiconvergence of extrapolated iterative methods for singular linear systems. Journal of Computational and Applied Mathematics, 106(1): 117–129, June 1999. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042799000606.

Sakurai:1996:SSM

[SP96] T. Sakurai and M. S.

Petković. On some simultaneous methods based on Weierstrass’ correction. Journal of Computational and Applied Mathematics, 72(2):275–291, August 13, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042795002782.

Spalevic:1999:PTQ

[Spa99] Miodrag M. Spalević. Product of Turán quadratures for cube, simplex, surface of the sphere, E_{n}^{r}, E_{n}^{r^{2}}. Journal of Computational and Applied Mathematics, 106(1): 99–115, June 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042799000588.

Spijker:1996:SNI

[Spi96] M. N. Spijker. Stiffness in numerical initial-value problems. Journal of Computational and Applied Mathematics, 72(2):393–406, August 13, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279600009X.

Spurzem:1999:DBS

[Spu99]  Rainer Spurzem.  Di-

rect N-body simulations. Journal of Computational and Applied Mathematics, 109(1–2):407–432, September 30, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042799001661.

Stein:1991:MAL

[SR91] Erwin Stein and Wilhelm Rust. Mesh adaptations for linear 2d finite-element discretizations in structural mechanics, especially in thin shell analysis. Journal of Computational and Applied Mathematics, 36(1):107– 129, August 20, 1991. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279190229D.

Simos:1992:FOB

[SR92] T. E. Simos and A. D.

Raptis. A fourth-order Bessel fitting method for the numerical solution of the Schrödinger equation. Journal of Computational and Applied Mathematics, 43(3):313–322, December 2, 1992. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279290017R.

Speith:1999:VGR

[SR99] R. Speith and H. Riffert.

The viscous gas ring as an astrophysical test problem for a viscous SPH-code. Journal of Computational and Applied Mathematics, 109(1–2):231–242, September 30, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042799001600.

Sanchez-Ruiz:1998:ESO

[SRD98] Jorge Sánchez-Ruiz and Jesús S. Dehesa. Expansions in series of orthogonal hypergeometric polynomials. Journal of Computational and Applied Mathematics, 89(1):155– 170, March 9, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042797002434.

Schafer:1990:ECG

[SS90] Eugen Schäfer and Ingrid Sporer. An efficient conjugate gradient preconditioner for Toeplitz systems. Journal of Computational and Applied Mathematics, 29(2):233–241, February 15, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279090360C.

Solak:1991:QRG

[SS91a] W. Solak and Z. Szydelko. Quadrature rules with Gregory–Laplace end corrections. Journal of Computational and Applied Mathematics, 36(2):251– 253, August 27, 1991. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/037704279190031E.

Stys:1991:LOS

[SS91b] Tadeusz Stys and Krystyna Stys. O( h ^{4} ) locally overconvergent semidiscrete scheme for the equation u _{t} = u _{xx} + f(t, x, u). Journal of Computational and Applied Mathematics, 34(2): 221–231, April 4, 1991. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279190044K.

Sideridis:1992:LOE

[SS92] A. B. Sideridis and T. E.

Simos. A low-order embedded Runge–Kutta method for periodic initial-value problems. Journal of Computational and Applied Mathematics, 44(2):235– 244, December 17, 1992. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
037704279290013N.

Saranen:1994:MQM

[SS94a] J. Saranen and L. Schroderus.

The modified quadrature method for classical pseudodifferential equations of negative order on smooth closed curves. Journal of Computational and Applied Mathematics, 50(1–3):485– 495, May 20, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042794903220.

Sartoretto:1994:CEC

[SS94b] F. Sartoretto and R. Spigler.

Computing expansion coefficients in orthogonal bases of ultraspherical polynomials. Journal of Computational and Applied Mathematics, 56(3):295–303, December 30, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042794900841.

Sloboda:1994:PLA

[SS94c] Fridrich Sloboda and Josef Stoer. On piecewise linear approximation of planar Jordan curves. Journal of Computational and Applied Mathematics, 55 (3):369–383, November 30, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        037704279490040X.

Simion:1996:OLP

[SS96a] R. Simion and D. Stanton. Octabasic Laguerre polynomials and permutation statistics. Journal of Computational and Applied Mathematics, 68(1–2):297– 329, April 22, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042795002502.

Sivakumar:1996:NNI

[SS96b] T. R. Sivakumar and S. Savithri. A new nonlinear integration formula for ODEs. Journal of Computational and Applied Mathematics, 67(2):291– 299, March 29, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042794001286.

Siyyam:1997:ASP

[SS97a] H. I. Siyyam and M. I.

Syam. An accurate solution of the Poisson equation by the Chebyshev–Tau method. Journal of Computational and Applied Mathematics, 85(1):1–10, November 6, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042797001040.

Siyyam:1997:MTR

[SS97b] H. I. Siyyam and M. I.

Syam. The modified trapezoidal rule for line integrals. Journal of Computational and Applied Mathematics, 84(1):1–14, October 21, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704279700071X.

Solodov:1997:DML

[SS97c] M. V. Solodov and B. F.

Svaiter. Descent methods with linesearch in the presence of perturbations. Journal of Computational and Applied Mathematics, 80(2): 265–275, May 5, 1997. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042797000253.

Stys:1997:OAC

[SS97d] Krystyna Styś and Tadeusz Styś. An optimal algorithm for certain boundary value problem. Journal of Computational and Applied Mathematics, 83(2):195–203, October 7, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042797000988.

Seidel:1999:NRT

[SS99a] Edward Seidel and Wai-Mo Suen. Numerical relativity as a tool for computational astrophysics. Journal of Computational and Applied Mathematics, 109(1–2):493–525, September 30, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042799001697.

Syam:1999:NDI

[SS99b] Muhammed I. Syam and Hani I. Siyyam. Numerical differentiation of implicitly defined curves. Journal of Computational and Applied Mathematics, 108(1–2):131– 144, August 15, 1999. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042799001065.

Stathopoulos:1995:RPL

[SSF95] Andreas Stathopoulos, Yousef Saad, and Charlotte Froese Fischer. Robust preconditioning of large, sparse, symmetric eigenvalue problems. Journal of Computational and Applied Mathematics, 64(3):197–215, December 20, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042795001417.

Saridakis:1996:FEC

[SSS96] Y. G. Saridakis, C. G.

Sifniotopoulos, and D. A. Sotiropoulos. Finite element collocation approximation of the characteristic exponent in BVPs with periodic coefficients. Journal of Computational and Applied Mathematics, 70(1):1– 14, June 14, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042795001549.

Sommeijer:1998:RES

[SSV98] B. P. Sommeijer, L. F. Shampine, and J. G. Verwer. RKC: an explicit solver for parabolic PDEs. Journal of Computational and Applied Mathematics, 88(2):315– 326, March 2, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042797002197.

Santos:1998:BSM

[SSY98] C. H. Santos, B. P. B. Silva, and J. Y. Yuan. Block SOR methods for rankdeficient least-squares problems. Journal of Computational and Applied Mathematics, 100(1):1–9, November 30, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042798001149.

Sugiura:1991:MCG

[ST91] Hiroshi Sugiura and Tatsuo Torii. A method for constructing generalized Runge–Kutta methods. Journal of Computational and Applied Mathematics, 38(1–3):399–410, December 23, 1991. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/037704279190185M.

Siddiqi:1997:SSL

[ST97] Shahid S. Siddiqi and E. H.

Twizell. Spline solutions of linear twelfth-order boundary-value problems. Journal of Computational and Applied Mathematics, 78(2):371–390, February 17, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042796001641.

Sirotkin:1999:CPI

[ST99] V. Sirotkin and P. Tarvainen. Comparison of parallel implementation of some multi-level Schwarz methods for singularly perturbed parabolic problems. Journal of Computational and Applied Mathematics, 104(2): 193–212, May 17, 1999. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042799000576.

Stadlober:1990:RUA

[Sta90] Ernst Stadlober. The ratio of uniforms approach for generating discrete random variates. Journal of Computational and Applied Mathematics, 31(1): 181–189, July 24, 1990. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0377042790903495.

Starke:1993:FWP

[Sta93]  Gerhard Starke.  Fejér–

Walsh points for rational functions and their use in the ADI iterative method. Journal of Computational and Applied Mathematics, 46(1–2):129–141, June 14, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279390291I.

Stahl:1997:CDP

[Sta97] Herbert Stahl. The convergence of diagonal Padé approximants and the Padé Conjecture. Journal of Computational and Applied Mathematics, 86(1):287– 296, November 28, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042797001623.

Stahl:1998:SPP

[Sta98] Herbert Stahl. Spurious poles in Padé approximation. Journal of Computational and Applied Mathematics, 99(1–2):511–527, November 16, 1998. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042798001800.

Stewart:1990:ASI

[Ste90a]  Kris Stewart.  Avoiding

stability-induced inefficiencies in BDF methods. Journal of Computational and Applied Mathematics, 29(3): 357–367, March 15, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
037704279090017T.

Stewart:1990:MSS

[Ste90b] Kris Stewart. A model for stability of the semi-implicit backward differentiation formulas. Journal of Computational and Applied Mathematics, 33(3):245–259, December 31, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042705800020.

Stenberg:1995:STA

[Ste95] Rolf Stenberg. On some techniques for approximating boundary conditions in the finite element method. Journal of Computational and Applied Mathematics, 63(1–3):139–148, November 20, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042795000577.

Stenger:1997:MSM

[Ste97]  Frank Stenger.  Ma-

trices of Sinc methods. Journal of Computational and Applied Mathematics, 86(1):297–310, November 28, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042797001635.

Special issue dedicated to William B. Gragg (Monterey, CA, 1996).

Steinbach:1999:AFE

[Ste99] O. Steinbach. Adaptive finite element-boundary element solution of boundary value problems. Journal of Computational and Applied Mathematics, 106(2): 307–316, June 30, 1999. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042799000734.

Stolan:1995:ISA

[Sto95]  J. A. Stolan.  An im-

proved Siljak’s algorithm for solving polynomial equations converges quadratically to multiple zeros. Journal of Computational and Applied Mathematics, 64(3):247–268, December 20, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042794001146.

Stoyanova:1997:CFS

[Sto97] Srebra B. Stoyanova. Cubature formulae of the seventh degree of accuracy for the hypersphere. Journal of Computational and Applied Mathematics, 84(1):15–21, October 21, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042797000940.

Stone:1999:ZCA

[Sto99] James M. Stone. The ZEUS code for astrophysical magnetohydrodynamics: new extensions and applications. Journal of Computational and Applied Mathematics, 109(1–2):261–280, September 30, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042799002046.

Strom:1994:CBS

[Str94] Kyrre Strøm. On convolutions of B-splines. Journal of Computational and Applied Mathematics, 55(1):1– 29, October 31, 1994. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0377042794901821.

Struckmeier:1995:FGL

[Str95] J. Struckmeier. Fast generation of low-discrepancy sequences. Journal of Computational and Applied Mathematics, 61(1):29–41, July 20, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042794000545.

Sakurai:1991:HOI

[STS91] T. Sakurai, T. Torii, and H. Sugiura. A high-order iterative formula for simultaneous determination of zeros of a polynomial. Journal of Computational and Applied Mathematics, 38(1–3):387–397, December 23, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279190184L.

Stuller:1995:OMG

[Stu95]  Július Stuller.  Ordered

modified Gram–Schmidt orthogonalization revised. Journal of Computational and Applied Mathematics, 63(1–3):221–227, November 20, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042795000755.

Stys:1992:FDS

[Sty92] Tadeusz Stys. A finitedifference scheme overconvergent at certain mesh points. Journal of Computational and Applied Mathematics, 42(2):233–243, October 12, 1992. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279290077B.

Stynes:1995:FVM

[Sty95] Martin Stynes. Finite volume methods for convection– diffusion problems. Journal of Computational and Applied Mathematics, 63(1–3): 83–90, November 20, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042795000569.

Stys:1996:FDS

[Sty96] Tadeusz Styś. O(τ^{2} + h^{4})

finite difference scheme for decoupled system of two quasilinear parabolic equations. Journal of Computational and Applied Mathematics, 66(1–2):485–496, January 31, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042795001794.

Seymour:1999:ISP

[SU99] Chris Seymour and Keith Unsworth. Interactive shape preserving interpolation by curvature continuous rational cubic splines. Journal of Computational and Applied Mathematics, 102(1):87– 117, February 15, 1999. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042798002106.

Sun:1996:SOA

[Sun96] Zhi-Zhong Sun. A secondorder accurate finite difference scheme for a class of nonlocal parabolic equations with natural boundary conditions. Journal of Computational and Applied Mathematics, 76(1–2):137–146, December 17, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042796000970.

Suri:1991:RVF

[Sur91] Manil Suri. On the robustness of the h- and pversions of the finite-element method. Journal of Computational and Applied Mathematics, 35(1–3):303–310, June 26, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042791902178.

Scobelev:1992:SAD

[SV92a] B. Yu. Scobelev and E. V.

Vorozhtsov. Stability analysis of difference schemes by catastrophe theory methods. Journal of Computational and Applied Mathematics, 44(1):1–20, December 9, 1992. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042792900508.

Spigler:1992:LGA

[SV92b] Renato Spigler and Marco Vianello. Liouville–Green approximations for a class of linear oscillatory difference equations of the second order. Journal of Computational and Applied Mathematics, 41(1–2):105–116, August 20, 1992. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279290241O.

Szabados:1992:SMC

[SV92c] J. Szabados and P. Vértosi.

A survey on mean convergence of interpolatory processes. Journal of Computational and Applied Mathematics, 43(1–2):3–18, November 25, 1992. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/037704279290256W.

Sansigre:1995:LFS

[SV95a] Gabriela Sansigre and Galliano Valent. A large family of semi-classical polynomials: the perturbed Chebyshev. Journal of Computational and Applied Mathematics, 57(1–2):271–281, February 20, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042793E0251G.

Shenton:1995:CFA

[SV95b] L. R. Shenton and H. D. Vinod. Closed forms for asymptotic bias and variance in autoregressive models with unit roots. Journal of Computational and Applied Mathematics, 61(2): 231–243, July 31, 1995. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/037704279400069D.

Sinap:1996:OMP

[SV96] Ann Sinap and Walter Van Assche. Orthogonal matrix polynomials and applications. Journal of Computational and Applied Mathematics, 66(1–2):27–52, January 31, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279500193X.

Shampine:1995:CLE

[SW95a] L. F. Shampine and A. Witt.

Control of local error stabilizes integrations. Journal of Computational and Applied Mathematics, 62 (3):333–351, September 29, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042794001081.

Shampine:1995:SSS

[SW95b] L. F. Shampine and A. Witt.

A simple step size selection algorithm for ODE codes. Journal of Computational and Applied Mathematics, 58(3):345–354, April 20, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
037704279400007N.

Schaback:1996:ORF

[SW96] Robert Schaback and Z. Wu.

Operators on radial functions. Journal of Computational and Applied Mathematics, 73(1–2):257–270, October 5, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042796000477.

Simos:1997:FDM

[SW97] T. E. Simos and P. S. Williams. A finite-difference method for the numerical solution of the Schrödinger equation. Journal of Computational and Applied Mathematics, 79(2):189– 205, March 17, 1997. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042796001562.

Swann:1996:EEU

[Swa96]  Howard Swann.  Error

estimates using the cell discretization method for some parabolic problems. Journal of Computational and Applied Mathematics, 66(1–2):497–514, January 31, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042795001891.

Swann:1997:EEU

[Swa97]  Howard Swann.  Error

estimates using the cell discretization method for steady-state convection– diffusion equations. Journal of Computational and Applied Mathematics, 82(1– 2):389–405, September 15, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042797000514.

Shaw:1996:DST

[SWW96] S. Shaw, M. K. Warby, and J. R. Whiteman. Discrete schemes for treating hereditary problems of viscoelasticity and applications. Journal of Computational and Applied Mathematics, 74(1–2):313–329, November 5, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042796000295.

Shi:1998:SGQ

[SWY98] Xiquan Shi, Tianjun Wang, and Baocai Yin. Splines on generalized quasi-cross-cut partitions. Journal of Computational and Applied Mathematics, 96(2):139– 147, September 15, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042798000922.

Shivakumar:1999:DPM

[SX99] P. N. Shivakumar and Jungong Xue. On the double points of a Mathieu equation. Journal of Computational and Applied Mathematics, 107(1):111–125, July 15, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042799000849.

Shen:1990:SNS

[SY90] Y.-Q. Shen and T. J. Ypma. Solving nonlinear systems of equations with only one nonlinear variable. Journal of Computational and Applied Mathematics, 30(2): 235–246, May 28, 1990. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/037704279090031T.

Syam:1999:NPE

[Sya99] Muhammed I. Syam. A new proof of the Euler– Maclaurin expansion for quadrature over implicitly defined curves. Journal of Computational and Applied Mathematics, 104(1):19–25, April 15, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042798001988.

Sydora:1999:LNE

[Syd99]  R. D. Sydora.  Low-

noise electromagnetic and relativistic particle-in-cell plasma simulation models. Journal of Computational and Applied Mathematics, 109(1–2):243–259, September 30, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042799001612.

Selting:1997:NSA

[SZ97] P. A. Selting and Q. Zheng.

Numerical stability analysis of oscillating integrated circuits. Journal of Computational and Applied Mathematics, 82(1–2):367–378, September 15, 1997. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042797000873.

Sun:1998:TDS

[SZ98] Zhi-Zhong Sun and Qi-Ding Zhu. On Tsertsvadze’s difference scheme for the Kuramoto-tsuzuki equation. Journal of Computational and Applied Mathematics, 98(2):289–304, October 19, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042798001356.

Szafraniec:1993:LST

[Sza93] Franciszek Hugon Szafraniec.

A (little) step towards orthogonality of analytic polynomials. Journal of Computational and Applied Mathematics, 49(1–3):255–261, December 31, 1993. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042793901588.

Szafraniec:1995:MLS

[Sza95]  F. H. Szafraniec.  A

method of localizing the spectra of sequences of orthogonal polynomials. Journal of Computational and Applied Mathematics, 65(1–3):387–394, December 29, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042795001263.

Tan:1995:HTA

[TA95] Roger C. E. Tan and Alan L. Andrew. Henrici’s transformation and its application to the computation of derivatives of eigensystems. Journal of Computational and Applied Mathematics, 59(1):1– 8, April 28, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279400013Q.

Tagliani:1998:ITS

[Tag98]  Aldo Tagliani.  Inverse

two-sided Laplace transform for probability density functions. Journal of Computational and Applied Mathematics, 90(2):157– 170, April 17, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042798000132.

Taksar:1992:SPN

[Tak92] M. I. Taksar. Skorohod problems with nonsmooth boundary conditions. Journal of Computational and Applied Mathematics, 40(2): 233–251, June 24, 1992. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279290108A.

Tsuchiya:1997:PEE

[TB97] Takuya Tsuchiya and Ivo Babuska. A priori error estimates of finite element solutions of parametrized strongly nonlinear boundary value problems. Journal of Computational and Applied Mathematics, 79(1):41–66, March 3, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042796001495.

Tellez:1999:RBD

[TB99] Manuel Antonio Aguirre Téllez and Ana Lucía Barrenechea. A relation between the k-th derivative of the Dirac delta in (P ± i0) and the residue of distributions (P ± i0)^{λ}. Journal of Computational and Applied Mathematics, 108(1–2):31– 40, August 15, 1999. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042799000989.

Trivellato:1999:FVM

[TBF99] F. Trivellato, E. Bertolazzi, and B. Firmani. Finite volume modelling of free surface draining vortices. Journal of Computational and Applied Mathematics, 103(1):175–185, March 15, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042798002519.

Temme:1990:UAA

[TD90] N. M. Temme and A. B. Olde Daalhuis. Uniform asymptotic approximation of Fermi– Dirac integrals. Journal of Computational and Applied Mathematics, 31(3):383– 387, October 1, 1990. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042790900382.

Taseli:1997:NACa

[TD97] H. Taşeli and M. Demiralp. A new approach to the classical Stokes flow problem: Part I. Methodology and first-order analytical results. Journal of Computational and Applied Mathematics, 78(2):213– 232, February 17, 1997. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042796001410.

Taseli:1997:NACb

[TE97] H. Taşeli and R. Eid. A new approach to the classical Stokes flow problem: Part II. Series solutions and higher-order applications. Journal of Computational and Applied Mathematics, 78(2):233–254, February 17, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704279600146X.

Tellez:1998:D

[Tél98] Manuel Aguirre Téllez.

The distribution δ^{(k)}(P ± i0 − m^{2}). Journal of Computational and Applied Mathematics, 88(2):339– 348, March 2, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042797002227.

Temme:1992:AII

[Tem92] N. M. Temme. Asymptotic inversion of the incomplete beta function. Journal of Computational and Applied Mathematics, 41(1–2):145– 157, August 20, 1992. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/037704279290244R.

Temme:1995:UAE

[Tem95] N. M. Temme. Uniform asymptotic expansions of integrals: a selection of problems. Journal of Computational and Applied Mathematics, 65(1–3):395–417, December 29, 1995. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0377042795001271.

Tabata:1991:FEA

[TF91] Masahisa Tabata and Shoichi Fujima. Finite-element analysis of high Reynolds number flow past a circular cylinder. Journal of Computational and Applied Mathematics, 38(1–3):411–424, December 23, 1991. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279190186N.

Taksar:1991:PAC

[TG91] Michael I. Taksar and Winfried K. Grassmann. Probabilistic approach to computational algorithms for finding stationary distributions of Markov chains. Journal of Computational and Applied Mathematics, 36(2):131– 136, August 27, 1991. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279190021B.

Torrano:1993:MPB

[TG93] E. Torrano and R. Guadalupe.

On the moment problem in the bounded case. Journal of Computational and Applied Mathematics, 49(1–3):263–269, December 31, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042793901599.

Thoolen:1994:AMC

[TH94] P. M. C. Thoolen and P. W. Hemker. Approximate methods for n-component solute transport and ionexchange. Journal of Computational and Applied Mathematics, 53(2):275– 290, August 19, 1994. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0377042794900493.

Thompson:1990:SCD

[Tho90]  S. Thompson.  Stepsize

control for delay differential equations using continuously imbedded Runge– Kutta methods of Sarafyan. Journal of Computational and Applied Mathematics, 31(2):267–275, August 31, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279090169Z.

Thron:1990:OTE

[Thr90] W. J. Thron. Order and type of entire functions arising from separately convergent continued fractions. Journal of Computational and Applied Mathematics, 32(1–2):273–279, November 26, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042790904375.

Thron:1999:ITE

[Thr99] W. J. Thron. Improved truncation error bounds for limit periodic continued fractions with additional assumptions on its elements. Journal of Computational and Applied Mathematics, 105(1–2):467– 476, May 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042799000394.

Thukral:1999:FPT

[Thu99] R. Thukral. A family of Padé-type approximants for accelerating the convergence of sequences. Journal of Computational and Applied Mathematics, 102(2):287– 302, February 26, 1999. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042798002295.

Tichy:1990:RPC

[Tic90] Robert F. Tichy. Random points in the cube and on the sphere with applications to numerical analysis. Journal of Computational and Applied Mathematics, 31(1): 191–197, July 24, 1990. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042790903509.

Timus:1993:SFR

[Tim93] D. M. Timus. Some further results on the Hubbell rectangular source integral.

Journal of Computational and Applied Mathematics, 45(3):331–336, May 3, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
037704279390050L.

Tiwari:1998:ADD

[TK98] S. Tiwari and A. Klar. An adaptive domain decomposition procedure for Boltzmann and Euler equations. Journal of Computational and Applied Mathematics, 90(2):223–237, April 17, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042798000272.

Troff:1991:NMT

[TLL91] B. Troff, T. H. Lê, and Ta Phuoc Loc. A numerical method for the threedimensional unsteady incompressible Navier–Stokes equations. Journal of Computational and Applied Mathematics, 35(1–3):311– 318, June 26, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042791902189.

Twizell:1992:NMC

[TM92] E. H. Twizell and S. A. Matar. Numerical methods for computing the eigenvalues of linear fourthorder boundary-value problems. Journal of Computational and Applied Mathematics, 40(1):115–125, June 12, 1992. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
037704279290046Z.

Torres:1999:NSF

[TOCH99] Carlos R. Torres, José Ochoa, JoséE. Castillo, and Hideshi Hanazaki. Numerical simulation of flow past a sphere in vertical motion within a stratified fluid. Journal of Computational and Applied Mathematics, 103(1):67–76, March 15, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042798002416.

Tokarzewski:1996:PPA

[Tok96a] Stanisław Tokarzewski. Npoint Padé approximants to real-valued Stieltjes series with nonzero radii of convergence. Journal of Computational and Applied Mathematics, 75(2):259– 280, November 28, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0377042796000684.

Tokarzewski:1996:TPP

[Tok96b] Stanisław Tokarzewski. Twopoint Padé approximants for the expansions of Stieltjes functions in real domain. Journal of Computational and Applied Mathematics, 67(1):59–72, February 20, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042794001227.

Totik:1998:OPR

[Tot98] Vilmos Totik. Orthogonal polynomials with respect to varying weights. Journal of Computational and Applied Mathematics, 99(1–2):373–385, November 16, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704279800171X.

Teghem:1995:ELP

[TPA95] J. Teghem, M. Pirlot, and C. Antoniadis. Embedding of linear programming in a simulated annealing algorithm for solving a mixed integer production planning problem. Journal of Computational and Applied Mathematics, 64(1–2):91–102, November 30, 1995. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042795000097.

Tracogna:1996:ITS

[Tra96] S. Tracogna. Implementation of two-step Runge– Kutta methods for ordinary differential equations. Journal of Computational and Applied Mathematics, 76(1–2):113–136, December 17, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042796000933.

Tinoco-Ruiz:1999:ACG

[TRBS99] J. G. Tinoco-Ruiz and P. Barrera-Sánchez. Area control in generating smooth and convex grids over general plane regions. Journal of Computational and Applied Mathematics, 103(1): 19–32, March 15, 1999. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S037704279900237X.

Trench:1991:EWC

[Tre91] William F. Trench. Explicit weighting coefficients for predicting ARMA time series from the finite past. Journal of Computational and Applied Mathematics, 34(2):251–262, April 4, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        037704279190047N.

Trench:1992:ABS

[Tre92a] William F. Trench. Asymptotic behavior of solutions of a linear second-order difference equation. Journal of Computational and Applied Mathematics, 41(1–2):95– 103, August 20, 1992. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/037704279290240X.

Trench:1992:MIC

[Tre92b] William F. Trench. Mild integrability conditions for global solutions of an elliptic equation. Journal of Computational and Applied Mathematics, 39(1):38–48, February 28, 1992. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279290220R.

Trench:1999:ICI

[Tre99] William F. Trench. Invertibly convergent infinite products of matrices. Journal of Computational and Applied Mathematics, 101(1–2):255–263, January 15, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042798001915.

Trimeche:1993:RTD

[Tri93]  Khalifa Trimèche.  The

Radon transform and its dual associated with partial differential operators and applications to polynomials on the unit disk. Journal of Computational and Applied Mathematics, 49(1–3):271–280, December 31, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279390160D.

Thomas:1997:FHE

[TS97] R. M. Thomas and T. E.

Simos. A family of hybrid exponentially fitted predictor–corrector methods for the numerical integration of the radial Schrödinger equation. Journal of Computational and Applied Mathematics, 87(2):215–226, December 23, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704279700188X.

Tsao:1996:MAQ

[Tsa96] Min Tsao. A measure of accuracy for quadrature methods for Fredholm equations of the first kind. Journal of Computational and Applied Mathematics, 69(2): 243–249, May 23, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0377042795000364.

Tseng:1995:LCI

[Tse95] Paul Tseng. On linear convergence of iterative methods for the variational inequality problem. Journal of Computational and Applied Mathematics, 60(1–2):237– 252, June 20, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279400094H.

Tsitouras:1998:HOZ

[Tsi98] Ch. Tsitouras. High-order zero-dissipative Runge–Kutta– Nyström methods. Journal of Computational and Applied Mathematics, 95(1–2): 157–161, August 28, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042798000818.

Thomas:1996:FNT

[TSM96] R. M. Thomas, T. E.

Simos, and G. V. Mitsou. A family of Numerovtype exponentially fitted predictor–corrector methods for the numerical integration of the radial Schrödinger equation. Journal of Computational and Applied Mathematics, 67(2):255– 270, March 29, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279400126X.

Usman:1998:ESP

[UH98] Anila Usman and George Hall. Equilibrium states for predictor–corrector methods. Journal of Computational and Applied Mathematics, 89(2):275–308, March 16, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042797002392.

Uryasev:1994:DPF

[Ury94] Stanislav Uryas’ev. Derivatives of probability functions and integrals over sets given by inequalities. Journal of Computational and Applied Mathematics, 56(1–2):197–223, December 20, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042794903883.

Ushijima:1991:CAL

[Ush91]  Teruo Ushijima.  Com-

putational aspect of linear water wave problem. Journal of Computational and Applied Mathematics, 38(1–3):425–445, December 23, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279190187O.

Usmani:1992:UQS

[Usm92] Riaz A. Usmani. The use of quartic splines in the numerical solution of a fourthorder boundary value problem. Journal of Computational and Applied Mathematics, 44(2):187–200, December 17, 1992. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279290010U.

Uytterhoeven:1996:BRD

[Uyt96] G. Uytterhoeven. Book review: Digital video processing: A. Murat Tekalp Prentice-Hall signal processing series, Prentice Hall PTR, 1995, xxii + 526 pages, ISBN 0-13-190075-

7. Journal of Computational and Applied Mathematics, 66(1–2):N5–N6, January 31, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042796804742.

Valent:1993:OPQ

[Val93] Galliano Valent. Orthogonal polynomials for a quartic birth and death process. Journal of Computational and Applied Mathematics, 49(1–3):281–288, December 31, 1993. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042793901614.

Valent:1995:ASC

[Val95] Galliano Valent. Associated Stieltjes–Carlitz polynomials and a generalization of Heun’s differential equation. Journal of Computational and Applied Mathematics, 57(1–2):293–307, February 20, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042793E0253I.

Valent:1996:ESS

[Val96] Galliano Valent. Exact solutions of some quadratic and quartic birth and death processes and related orthogonal polynomials. Journal of Computational and Applied Mathematics, 67(1):103– 127, February 20, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0377042794001243.

VanHuffel:1990:PSV

[Van90a] Sabine Van Huffel. Partial singular value decomposition algorithm. Journal of Computational and Applied Mathematics, 33 (1):105–112, December 17, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042790902607.

VanHuffel:1990:PTL

[Van90b] Sabine Van Huffel. Partial total least squares algorithm. Journal of Computational and Applied Mathematics, 33(1):113–121, December 17, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279090261W.

VanAssche:1991:OPA

[Van91a] Walter Van Assche. Orthogonal polynomials, associated polynomials and functions of the second kind. Journal of Computational and Applied Mathematics, 37(1–3):237–249, November 18, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279190121Y.

VanGestel:1991:BRP

[Van91b] E. Van Gestel. Book review:

Programming in Martin-Löf’s Type Theory: an Introduction: Bengt Nordström, Kent Petersson and Jan M. Smith Intl. Series of Monographs on Computer Science, Vol. 7, Oxford Science Publications, Oxford, 1990, 231 pages. Journal of Computational and Applied Mathematics, 34(2): N4, April 4, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279190052L.

VanAssche:1993:CFT

[Van93a] Walter Van Assche. Christoffel functions and Turán determinants on several intervals. Journal of Computational and Applied Mathematics, 48(1–2):207–223, October 29, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042793903234.

VanAssche:1993:OP

[Van93b] Walter Van Assche. Open problems. Journal of Computational and Applied Mathematics, 48(1–2):225– 243, October 29, 1993. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0377042793903245.

VanIseghem:1993:GFR

[Van93c] Jeannette Van Iseghem. Generating function, recurrence relations, differential relations. Journal of Computational and Applied Mathematics, 49(1–3):299– 303, December 31, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042793901636.

VanRossum:1993:PSP

[Van93d] H. Van Rossum. Polynomial sequences with prescribed power sums of zeros. Journal of Computational and Applied Mathematics, 49(1–3):317–327, December 31, 1993. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0377042793901658.

VanBarel:1990:NARa

[VB90a] Marc Van Barel and Adhemar Bultheel. A new approach to the rational interpolation problem. Journal of Computational and Applied Mathematics, 32(1–2):281–289, November 26, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042790904386.

VanBarel:1990:NARb

[VB90b] Marc Van Barel and Adhemar Bultheel. A new approach to the rational interpolation problem: the vector case. Journal of Computational and Applied Mathematics, 33(3):331–346, December 31, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042705800081.

VanMelkebeek:1993:BOS

[VB93] Dieter Van Melkebeek and Adhemar Bultheel. Block orthogonal systems for symmetric P -forms. Journal of Computational and Applied Mathematics, 49(1–3):305– 315, December 31, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042793901647.

VanBarel:1994:DLL

[VB94] Marc Van Barel and Adhemar Bultheel. Discrete linearized least-squares rational approximation on the unit circle. Journal of Computational and Applied Mathematics, 50(1–3):545– 563, May 20, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042794903271.

VanBarel:1997:LAM

[VB97] Marc Van Barel and Adhemar Bultheel. Lookahead methods for block Hankel systems. Journal of Computational and Applied Mathematics, 86 (1):311–333, November 28, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042797001647.

VanDenBroeck:1991:DSF

[VBK91] Julien Van Den Broeck, Fernand Broeckx, and Leonard Kaufman. Decomposing stochastic frontier efficiency into secular and organisational efficiency. Journal of Computational and Applied Mathematics, 37(1–3):251–264, November 18, 1991. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279190122Z.

Verschelde:1994:SHC

[VC94] Jan Verschelde and Ronald Cools. Symmetric homotopy construction. Journal of Computational and Applied Mathematics, 50(1–3):575– 592, May 20, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042794903298.

VandenBerghe:1991:CNL

[VD91] G. Vanden Berghe and H. De Meyer. On a correction of Numerov-like eigenvalue approximations for Sturm–Liouville problems. Journal of Computational and Applied Mathematics, 37(1–3):179–186, November 18, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042791901162.

VandenBerghe:1994:FEE

[VD94] G. Vanden Berghe and H. De Meyer. A finiteelement estimate with trigonometric hat functions for Sturm–Liouville eigenvalues. Journal of Computational and Applied Mathematics, 53(3):389–396, August 30, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042794900663.

vanDoorn:1995:SNR

[vD95] Erik A. van Doorn. Some new results for chainsequence polynomials. Journal of Computational and Applied Mathematics, 57(1– 2):309–317, February 20, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042793E0254J.

vandeFliert:1995:NAC

[vdFvGdRdV95] B. W. van de Fliert, E. van Groesen, R. de Roo, and R. W. de Vries. Numerical algorithm for the calculation of nonsymmetric dipolar and rotating monopolar vortex structures. Journal of Computational and Applied Mathematics, 62(1):1– 25, August 31, 1995. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0377042794000774.

vanderHouwen:1997:PLS

[vdHM97] P. J. van der Houwen and E. Messina. Parallel linear system solvers for Runge– Kutta–Nyström methods. Journal of Computational and Applied Mathematics, 82(1–2):407–422, September 15, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042797000502.

vanderHouwen:1999:PAM

[vdHM99] P. J. van der Houwen and E. Messina. Parallel Adams methods. Journal of Computational and Applied Mathematics, 101(1–2):153– 165, January 15, 1999. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042798002143.

vanderHouwen:1998:UAF

[vdHS98] P. J. van der Houwen and B. P. Sommeijer. The use of approximate factorization in stiff ODE solvers. Journal of Computational and Applied Mathematics, 100(1):11–21, November 30, 1998. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042798001253.

vanderHouwen:1996:PPC

[vdHSdS96] P. J. van der Houwen, B. P. Sommeijer, and J. J. B. de Swart. Parallel predictor–corrector methods. Journal of Computational and Applied Mathematics, 66(1–2):53–71, January 31, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042795001581.

vanderHouwen:1995:PIA

[vdHSvdV95] P. J. van der Houwen, B. P. Sommeijer, and W. A. van der Veen. Parallel iteration across the steps of high-order Runge–Kutta methods for nonstiff initial value problems. Journal of Computational and Applied Mathematics, 60(3): 309–329, June 30, 1995. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0377042794000475.

VanKeer:1996:NMM

[VDM96] R. Van Keer, L. R. Dupré, and J. A. A. Melkebeek. On a numerical method for 2D magnetic field computations in a lamination with enforced total flux. Journal of Computational and Applied Mathematics, 72(1): 179–191, July 30, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0377042795002715.

vanderPloeg:1996:NGI

[vdPBW96] A. van der Ploeg, E. F. F.

Botta, and F. W. Wubs. Nested grids ILU-decomposition (NGILU). Journal of Computational and Applied Mathematics, 66(1–2):515– 526, January 31, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042795001824.

VandenBerghe:1990:CMN

[VDV90] G. Vanden Berghe, H. De Meyer, and J. Vanthournout. On a class of modified Newton–Cotes quadrature formulae based upon mixedtype interpolation. Journal of Computational and Applied Mathematics, 31(3): 331–349, October 1, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
037704279090034W.

vanderVen:1997:EMT

[vdVNTV97] H. van der Ven, B. E.

Niemann-Tuitman, and A. E. P. Veldman. An explicit multi-time-stepping algorithm for aerodynamic flows. Journal of Computational and Applied Mathematics, 82(1–2):423–431, September 15, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704279700054X.

vanderVorst:1993:SCB

[vdVV93] H. A. van der Vorst and C. Vuik. The superlinear convergence behaviour of GMRES. Journal of Computational and Applied Mathematics, 48(3):327– 341, November 30, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        037704279390028A.

Vecchio:1996:HSP

[Vec96] Antonia Vecchio. Highly stable parallel Volterra Runge– Kutta methods. Journal of Computational and Applied Mathematics, 71(2): 225–236, July 27, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0377042795002359.

Vekemans:1997:AP

[Vek97]  D. Vekemans.  Algo-

rithm for the E-prediction. Journal of Computational and Applied Mathematics, 85(2):181–202, November 12, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042797000095.

Venturino:1992:GMS

[Ven92]  Ezio Venturino.  The

Galerkin method for singular integral equations revisited. Journal of Computational and Applied Mathematics, 40(1):91–103, June 12, 1992. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279290044X.

Venturino:1993:UAS

[Ven93] Ezio Venturino. An unconventional algorithm for singular integral equations. Journal of Computational and Applied Mathematics, 49(1–3):329–337, December 31, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042793901669.

Venturino:1996:LIP

[Ven96] Ezio Venturino. On the (0.4)

lacunary interpolation problem and some related questions. Journal of Computational and Applied Mathematics, 76(1–2):287–300, December 17, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042796001148.

Verfurth:1991:P

[Ver91] R. Verfürth. Preface. Journal of Computational and Applied Mathematics, 36(1): 1, August 20, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042791902237.

Vermiglio:1993:MHO

[Ver93]  R. Vermiglio.  Multi-

step high-order interpolants of Runge–Kutta methods. Journal of Computational and Applied Mathematics, 45(1–2):75–88, April 8, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
037704279390266E.

Verfurth:1994:PEE

[Ver94a] R. Verfürth. A posteriori error estimation and adaptive mesh-refinement techniques. Journal of Computational and Applied Mathematics, 50(1–3):67– 83, May 20, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042794902909.

Verschelde:1994:BRH

[Ver94b] J. Verschelde. Book review: Handbook of convex geometry: P. M. Gruber and J. M. Wills (eds.), Convex Sets and Related Geometric Topics Geometry Volume A, 1993, 816 pages, Price: Dfl. 295.00 (US $168.50), ISBN 0-444-89596-5 Hardbound Volume B, 1993, 780 pages, Price: Dfl. 285.00 (US $162.75), ISBN 0-444-89597-5 Hardbound Two-Volume Set, Price: Dfl. 540.00 (US$308.50), ISBN 0-444-89598-1, North-Holland, Elsevier Science Publishers. Journal of Computational and Applied Mathematics, 54(3):N2–N3, October 20, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042794902615.

Verschelde:1996:BRB

[Ver96] J. Verschelde. Book review:

Basic algebraic geometry: Igor R. Shafarevich Second, revised and expanded edition, translated by Miles Reid, Springer-Verlag, 1994, Volume 1: Varieties in projective space, ISBN 3-540-54812-2, Softcover, DM 68. Volume 2: Schemes and complex manifolds, ISBN 3-540-57554-5, Softcover, DM

68. Journal of Computational and Applied Mathematics, 66(1–2):N3–N4, January 31, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042796804717.

Verlinden:1997:AGL

[Ver97] Pierre Verlinden. Acceleration of Gauss–Legendre quadrature for an integrand with an endpoint singularity. Journal of Computational and Applied Mathematics, 77(1–2):277–287, January 13, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042796001318.

vanGijzen:1995:PPG

[vG95] M. B. van Gijzen. A polynomial preconditioner for the GMRES algorithm. Journal of Computational and Applied Mathematics, 59(1): 91–107, April 28, 1995. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/037704279400015S.

Vozovoi:1996:MLF

[VIA96] L. Vozovoi, M. Israeli, and A. Averbuch. Multidomain local Fourier method for PDEs in complex geometries. Journal of Computational and Applied Mathematics, 66(1–2):543–555, January 31, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042795001859.

VanDooren:1996:CAD

[VJ96] R. Van Dooren and H. Janssen.

A continuation algorithm for discovering new chaotic motions in forced Duffing systems. Journal of Computational and Applied Mathematics, 66(1–2):527–541, January 31, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279500162X.

Vecchia:1994:RAU

[VM94] Biancamaria Della Vecchia and Giuseppe Mastroianni. Rational approximation in uniform and weighted L ^{1} norm. Journal of Computational and Applied Mathematics, 53(3):333–339, August 30, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042794900612.

Voss:1999:MIR

[VM99] D. A. Voss and P. H. Muir.

Mono-implicit Runge–Kutta schemes for the parallel solution of initial value ODEs. Journal of Computational and Applied Mathematics, 102(2):235–252, February 26, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042798002210.

Vecchia:1994:SAH

[VMV94] Biancamaria Della Vecchia, Giuseppe Mastroianni, and Péter Vértesi. Simultaneous approximation by Hermite interpolation of higher order. Journal of Computational and Applied Mathematics, 50(1–3):233–240, May 20, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042794903034.

vanNooyen:1993:IAV

[vN93] R. R. P. van Nooyen. An improved accuracy version of the mixed finite-element method for a second-order elliptic equation. Journal of Computational and Applied Mathematics, 47(1):11–33, June 28, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279390087R.

Voevudsky:1995:CRS

[Voe95] Valery P. Voevudsky. Convergence of regularized spline approximants to solutions of initial and boundary value problems for ODE. Journal of Computational and Applied Mathematics, 58(1):55–66, March 20, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042793E0267P.

Vogel:1994:SAS

[Vog94] Silvia Vogel. A stochastic approach to stability in stochastic programming. Journal of Computational and Applied Mathematics, 56(1–2):65–96, December 20, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042794903808.

Voit:1993:FHT

[Voi93]  Michael Voit.  A for-

mula of Hilb’s type for orthogonal polynomials. Journal of Computational and Applied Mathematics, 49(1–3):339–348, December 31, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279390167A.

Voit:1995:LTR

[Voi95] Michael Voit. Limit theorems for random walks on the double coset spaces U(n)||U(n − 1) for n → ∞. Journal of Computational and Applied Mathematics, 65(1–3):449–459, December 29, 1995. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0377042795001298.

Volk:1990:MDI

[Vol90] Wolfgang Volk. Making the difference interpolation method for splines more stable. Journal of Computational and Applied Mathematics, 33(1):53–59, December 17, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279090255X.

VanderJeugt:1997:TSF

[VPR97] J. Van der Jeugt, S. N. Pitre, and K. Srinivasa Rao. Transformation and summation formulas for double hypergeometric series. Journal of Computational and Applied Mathematics, 83(2):185–193, October 7, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042797000964.

Verschelde:1991:BRN

[VR91] J. Verschelde and D. Roose. Book review: Numerical Continuation Methods: an Introduction: E. L. Allgower, K. Georg, R. L. Graham, J. Stoer, R. Varga (Eds.) Springer-Verlag, 1990, Springer Series in Computational Mathematics. Volume 13. Approx. 400 pages, 37 figures. Hardcover, DM 128,-ISBN 3-540-12760-7. Journal of Computational and Applied Mathematics, 34(2):N2–N3, April 4, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279190050T.

Vrba:1995:MPB

[Vrb95] P. Vrba. Modelling of particle beams. Journal of Computational and Applied Mathematics, 63(1–3):457– 463, November 20, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042795000704.

VanDerHouwen:1990:PIH

[VS90] P. J. Van Der Houwen and B. P. Sommeijer. Parallel iteration of high-order Runge–Kutta methods with stepsize control. Journal of Computational and Applied Mathematics, 29(1):111– 127, January 10, 1990. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279090200J.

VanDoorn:1993:RWP

[VS93] Erik A. Van Doorn and Pauline Schrijner. Random walk polynomials and random walk measures. Journal of Computational and Applied Mathematics, 49(1–3):289–296, December 31, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042793901625.

Vuik:1995:NIG

[Vui95]  C. Vuik.  New insights

in GMRES-like methods with variable preconditioners. Journal of Computational and Applied Mathematics, 61(2):189–204, July 31, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        037704279400067B.

VanderStracten:1992:MRF

[VV92] M. Vander Stracten and H. Van de Vel. Multiple root-finding methods. Journal of Computational and Applied Mathematics, 40(1): 105–114, June 12, 1992. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/037704279290045Y.

VanBerkel:1994:CIE

[VV94] C. A. M. Van Berkel and S. J. L. Van Eijndhoven. Convolution integral equations with Gegenbauer function kernel. Journal of Computational and Applied Mathematics, 50(1–3):565– 574, May 20, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279490328X.

Valent:1995:ISW

[VV95] Galliano Valent and Walter Van Assche. The impact of Stieltjes’ work on continued fractions and orthogonal polynomials: additional material. Journal of Computational and Applied Mathematics, 65(1–3):419–447, December 29, 1995. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279500128X.

VanBarel:1996:IBL

[VV96] M. Van Barel and Z. Vavrín.

Inversion of a block Löwner matrix. Journal of Computational and Applied Mathematics, 69(2):261–284, May 23, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042795000372.

Vermolen:1998:NMC

[VV98] Fred Vermolen and Kees Vuik. A numerical method to compute the dissolution of second phases in ternary alloys. Journal of Computational and Applied Mathematics, 93(2): 123–143, July 13, 1998. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042798000764.

vanVeldhuizen:1993:DSR

[vVBZ93] M. van Veldhuizen, A. F. Bervoets, and W. Zijl. Dichotomy of a special recurrence relation from the earth sciences. Journal of Computational and Applied Mathematics, 47(1):61–66, June 28, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279390090X.

VandenBerghe:1992:FDF

[VVD92] G. Vanden Berghe, M. Van Daele, and H. De Meyer. A five-diagonal finite-difference method based on mixedtype interpolation for computing eigenvalues of fourthorder two-point boundaryvalue problems. Journal of Computational and Applied Mathematics, 41(3):359– 372, August 28, 1992. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279290142K.

VanDaele:1994:SAS

[VVD94] M. Van Daele, G. Vanden Berghe, and H. De Meyer. A smooth approximation for the solution of a fourth-order boundary value problem based on nonpolynomial splines. Journal of Computational and Applied Mathematics, 51(3): 383–394, June 30, 1994. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279200112M.

VanHecke:1997:MIR

[VVVD97] T. Van Hecke, M. Van Daele, G. Vanden Berghe, and H. De Meyer. A monoimplicit Runge–Kutta–Nyström modification of the Numerov method. Journal of Computational and Applied Mathematics, 78(1):161– 177, February 3, 1997. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042796001392.

VanDaele:1999:DCM

[VVVD99a] M. Van Daele, T. Van Hecke, G. Vanden Berghe, and H. De Meyer. Deferred correction with monoimplicit Runge–Kutta methods for first-order IVPs. Journal of Computational and Applied Mathematics, 111(1–2):37–47, November 15, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042799001302.

VanHecke:1999:CTS

[VVVD99b] T. Van Hecke, G. Vanden Berghe, M. Van Daele, and H. De Meyer. Comment on: A two-stage fourth-order “almost” P - stable method for oscillatory problems. Journal of Computational and Applied Mathematics, 103(2):297– 300, March 31, 1999. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042798002660.

See [CAZ98].

Vanmaele:1994:EFE

[VZ94] M. Vanmaele and A. Zenísek. External finite-element approximations of eigenfunctions in the case of multiple eigenvalues. Journal of Computational and Applied Mathematics, 50(1–3): 51–66, May 20, 1994. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042794902895.

Wong:1998:EIT

[WA98] Patricia J. Y. Wong and Ravi P. Agarwal. On eigenvalue intervals and twin eigenfunctions of higherorder boundary value problems. Journal of Computational and Applied Mathematics, 88(1):15–43, February 23, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042797002021.

Waadeland:1991:BVT

[Waa91] Haakon Waadeland. Boundary version of a twin region convergence theorem for continued fractions. Journal of Computational and Applied Mathematics, 36(3):361–369, September 24, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279190016D.

Wambsganss:1999:GLN

[Wam99] Joachim Wambsganss. Gravitational lensing: numerical simulations with a hierarchical tree code. Journal of Computational and Applied Mathematics, 109(1–2):353– 372, September 30, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042799001648.

Wang:1994:NCE

[Wan94]  Junping Wang.  New

convergence estimates for multilevel algorithms for finite-element approximations. Journal of Computational and Applied Mathematics, 50(1–3):593–604, May 20, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042794903301.

Wang:1997:PEE

[Wan97] Song Wang. An a posteriori error estimate for finite element approximations of a singularly perturbed advection–diffusion problem. Journal of Computational and Applied Mathematics, 87(2):227–242, December 23, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704279700191X.

Washburn:1995:DPB

[Was95]  Alan Washburn.  Dy-

namic programming and the backpacker’s linear search problem. Journal of Computational and Applied Mathematics, 60(3):

357–365, June 30, 1995. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042794000383.

Watkins:1997:UOP

[Wat97] David S. Watkins. Unitary orthogonalization processes. Journal of Computational and Applied Mathematics, 86(1):335–345, November 28, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042797001659.

Wazwaz:1991:ASD

[Waz91] Abdul-Majid Wazwaz. On the asymptotic solutions of a differential equation with multiple singular points. Journal of Computational and Applied Mathematics, 34(1):65–74, February 10, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279190148D.

Wille:1994:SCC

[WB94] David R. Willé and Christopher T. H. Baker. Stepsize control and continuity consistency for statedependent delay-differential equations. Journal of Computational and Applied Mathematics, 53(2):163– 170, August 19, 1994. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0377042794900434.

Willemans:1994:SFU

[WD94] Karin Willemans and Paul Dierckx. Surface fitting using convex Powell–Sabin splines. Journal of Computational and Applied Mathematics, 56(3):263–282, December 30, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042794900825.

Werner:1999:CSA

[WD99] Klaus Werner and Stefan Dreizler. The classical stellar atmosphere problem. Journal of Computational and Applied Mathematics, 109(1–2):65–93, September 30, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042799001545.

Wegert:1990:IMS

[Weg90a]  Elias Wegert.  An it-

erative method for solving nonlinear Riemann– Hilbert problems. Journal of Computational and Applied Mathematics, 29(3):311– 327, March 15, 1990. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279090014Q.

Wegmann:1990:DVN

[Weg90b]  Rudolf Wegmann.  Dis-

cretized versions of Newton type iterative methods for conformal mapping. Journal of Computational and Applied Mathematics, 29(2):207–224, February 15, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042790903587.

Wegert:1993:IMD

[Weg93a]  Elias Wegert.  Iterative

methods for discrete nonlinear Riemann–Hilbert problems. Journal of Computational and Applied Mathematics, 46(1–2):143–163, June 14, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279390292J.

Wegmann:1993:EPH

[Weg93b] Rudolf Wegmann. Extremal problems for harmonic mappings from the unit disc to convex regions. Journal of Computational and Applied Mathematics, 46(1–2):165– 181, June 14, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279390293K.

Wegmann:1996:FCM

[Weg96]  Rudolf Wegmann.  Fast

conformal mapping of an ellipse to a simply connected region. Journal of Computational and Applied Mathematics, 72(1): 101–126, July 30, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0377042795002472.

Wegenkittl:1998:THS

[Weg98] Stefan Wegenkittl. Are there hyperbolas in the scatter plots of inversive congruential pseudorandom numbers? Journal of Computational and Applied Mathematics, 95(1–2):117–125, August 28, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042798000776.

Weisz:1996:IMS

[Wei96] J. Weisz. On an iterative method for the solution of discretized elliptic problems with imperfect contact condition. Journal of Computational and Applied Mathematics, 72(2):319–333, August 13, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042796000039.

Weili:1997:SPN

[Wei97] Zhao Weili. Singular perturbations for nonlinear Robin problems. Journal of Computational and Applied Mathematics, 81(1): 59–74, June 1, 1997. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042797000101.

Weniger:1990:SSD

[Wen90] Ernst Joachim Weniger.

On the summation of some divergent hypergeometric series and related perturbation expansions. Journal of Computational and Applied Mathematics, 32(1–2):291–300, November 26, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042790904397.

Wenz:1997:NFE

[Wen97] Hans-Jörg Wenz. Note on fast evaluation of best L_{2} approximation polynomials on simplices. Journal of Computational and Applied Mathematics, 79(2):319– 322, March 17, 1997. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042797000058.

Wendland:1999:SPD

[Wen99]  Holger Wendland.  On

the smoothness of positive definite and radial functions. Journal of Computational and Applied Mathematics, 101(1–2):177–188, January 15, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042798002180.

Wern:1995:FES

[Wer95]  Harald Wern.  Finite-

element solutions for mechanical drilling methods: a new integral formalism. Journal of Computational and Applied Mathematics, 63(1–3):365–372, November 20, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042795000895.

Westbrook:1990:OPB

[Wes90] D. R. Westbrook. The obstacle problem for beams and plates. Journal of Computational and Applied Mathematics, 30(3):295– 311, July 10, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042790902814.

Wilders:1997:DIS

[WF97] P. Wilders and G. Fotia. Distributed implicit stepping with unstructured finite volumes for 2Dtransport. Journal of Computational and Applied Mathematics, 82(1–2):433– 446, September 15, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042797000526.

Witwit:1999:NCA

[WGK99] M. R. M. Witwit, N. A.

Gordon, and J. P. Killingbeck. Numerical computation and analysis of the Titchmarsh–Weyl m_{α}(λ) function for some simple potentials. Journal of Computational and Applied Mathematics, 106(1):131– 143, June 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042799000618.

Wick:1990:PMS

[Wic90] Joachim Wick. A point method for a semi-conductor equation. Journal of Computational and Applied Mathematics, 31(1):199– 203, July 24, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279090351Y.

Wimp:1990:PPP

[Wim90] Jet Wimp. Pollaczek polynomials and Padé approximants: some closed-form expressions. Journal of Computational and Applied Mathematics, 32(1–2):301– 310, November 26, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        037704279090440B.

Wimp:1991:CTA

[Wim91] Jet Wimp. Current trends in asymptotics: some problems and some solutions. Journal of Computational and Applied Mathematics, 35(1–3): 53–79, June 26, 1991. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/037704279190197R.

Witwit:1995:SIP

[Wit95] M. R. M. Witwit. Some inner product techniques for computing eigenvalues for three-dimensional anharmonic oscillators with quartic and sextic perturbations. Journal of Computational and Applied Mathematics, 62(3):285–299, September 29, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279400110X.

Witwit:1996:ELD

[Wit96a] M. R. M. Witwit. Energy levels for a double-well potential in three-dimensional system using Hill determinant approach. Journal of Computational and Applied Mathematics, 67(1):129– 140, February 20, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042794001197.

Witwit:1996:ELN

[Wit96b] M. R. M. Witwit. Energy levels of a nonpolynomial oscillator using finite difference technique. Journal of Computational and Applied Mathematics, 69(2): 331–343, May 23, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042795001328.

Witwit:1996:QTD

[Wit96c] M. R. M. Witwit. Quantum theory of a double-well potential: Energy levels for symmetric and nonsymmetric double-well potentials in a three-dimensional system. Journal of Computational and Applied Mathematics, 75(2):197–213, November 28, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042796000490.

Wimp:1998:SPS

[WK98] Jet Wimp and Harry Kiesel. Some properties of the Sobolev–Laguerre polynomials. Journal of Computational and Applied Mathematics, 99(1–2):401–412, November 16, 1998. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042798001733.

Wong:1992:AES

[WL92] R. Wong and H. Li. Asymptotic expansions for secondorder linear difference equations. Journal of Computational and Applied Mathematics, 41(1–2):65–94, August 20, 1992. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279290239T.

Watson:1992:PPU

[WM92] Layne T. Watson and Alexander P. Morgan. Polynomial programming using multi- homogeneous polynomial continuation. Journal of Computational and Applied Mathematics, 43(3):373–382, December 2, 1992. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279290022P.

Walton:1994:AQB

[WM94] D. J. Walton and D. S.

Meek. Approximation of quadratic Bézier curves by arc splines. Journal of Computational and Applied Mathematics, 54(1):107– 120, September 20, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042794903980.

Walton:1996:APC

[WM96a] D. J. Walton and D. S.

Meek. Approximation of a planar cubic Bézier spiral by circular arcs. Journal of Computational and Applied Mathematics, 75(1):47–56, November 12, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042796000568.

Walton:1996:PCB

[WM96b] D. J. Walton and D. S. Meek. A planar cubic Bézier spiral. Journal of Computational and Applied Mathematics, 72(1):85–100, July 30, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042795002464.

Wimp:1997:CJF

[WMC97] Jet Wimp, P. McCabe, and J. N. L. Connor. Computation of Jacobi functions of the second kind for use in nearside-farside scattering theory. Journal of Computational and Applied Mathematics, 82(1–2):447–464, September 15, 1997. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042797000836.

Wolfe:1994:IAC

[Wol94] M. A. Wolfe. An interval algorithm for constrained global optimization. Journal of Computational and Applied Mathematics, 50(1–3): 605–612, May 20, 1994. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279490331X.

Wood:1991:NUI

[Woo91] A. S. Wood. A note on the use of the isotherm migration method. Journal of Computational and Applied Mathematics, 36 (3):371–384, September 24, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
037704279190017E.

Wright:1993:PTB

[Wri93] K. Wright. Parallel treatment of block-bidiagonal matrices in the solution of ordinary differential boundary value problems. Journal of Computational and Applied Mathematics, 45(1–2): 191–200, April 8, 1993. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279390274F.

Wimp:1992:AIR

[WRW92] Jet Wimp, Chris Rorres, and Richard F. Wayland, Jr. Acoustic impulse responses for nonuniform media. Journal of Computational and Applied Mathematics, 42(1):89–107, September 8, 1992. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279290165T.

Wang:1997:EEB

[WS97] Morten M. T. Wang and Tony W. H. Sheu. An element-by-element BICGSTAB iterative method for threedimensional steady Navier– Stokes equations. Journal of Computational and Applied Mathematics, 79(1):147– 165, March 3, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042796001720.

Wu:1990:MSM

[Wu90] C. H. Wu. A multicolour SOR method for the finiteelement method. Journal of Computational and Applied Mathematics, 30(3): 283–294, July 10, 1990. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/037704279090280D.

Wong:1997:AEG

[WZ97] R. Wong and J.-M. Zhang. Asymptotic expansions of the generalized Bessel polynomials. Journal of Computational and Applied Mathematics, 85(1):87–112, November 6, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042797001313.

Xie:1996:SCP

[XC96] Sheng-Li Xie and Sui Sun Cheng. Stability criteria for parabolic type partial difference equations. Journal of Computational and Applied Mathematics, 75(1):57–66, November 12, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042796000581.

Xieping-Gao:1996:QSS

[XGSSKF96] Xieping-Gao, Shi-Shu, and Kaixin-Fu. Quartic splineon-spline interpolation. Journal of Computational and Applied Mathematics, 71(2): 213–223, July 27, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042795002235.

Xia:1996:NMC

[Xia96] Linhua Xia. New methods of computation for coupled channel equations. Journal of Computational and Applied Mathematics, 75(2):281–293, November 28, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042796000805.

Xiong:1992:TDE

[XK92] Zonghou Xiong and Andreas Kirsch. Three-dimensional earth conductivity inversion. Journal of Computational and Applied Mathematics, 42(1):109–121, September 8, 1992. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279290166U.

Xu:1998:SEM

[XM98] Chuanju Xu and Yvon Maday. A spectral element method for the timedependent two-dimensional Euler equations: applications to flow simulations. Journal of Computational and Applied Mathematics, 91(1):63–85, April 24, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042798000223.

Xu:1999:CPA

[Xu99] Jian-Jun Xu. Convergence of partially asynchronous block quasi-Newton methods for nonlinear systems of equations. Journal of Computational and Applied Mathematics, 103(2):307– 321, March 31, 1999. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042798002684.

Xiang:1998:ETS

[XZ98] Kaili Xiang and Jianjun Zhang. Explicit two-step high-accuracy hybrid methods with minimal phaselag for y′′ = f(x, y) and their application to the onedimensional Schrödinger equation. Journal of Computational and Applied Mathematics, 95(1–2):1–11, August 28, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704279800051X.

Ye:1995:DMS

[YA95] Xiu Ye and Gary Anderson. The derivation of minimal support basis functions for the discrete divergence operator. Journal of Computational and Applied Mathematics, 61(1): 105–116, July 20, 1995. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042794000567.

Yakubovich:1997:IIT

[Yak97] S. B. Yakubovich. Index integral transformations of Titchmarsh type. Journal of Computational and Applied Mathematics, 85(1):169– 179, November 6, 1997. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S037704279700143X.

Yamamoto:1991:P

[Yam91] Tetsuro Yamamoto. Preface. Journal of Computational and Applied Mathematics, 38(1–3):ix, December 23, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279190156E.

Yamamoto:1995:P

[Yam95] Tetsuro Yamamoto. Preface. Journal of Computational and Applied Mathematics, 60(1–2):1, June 20, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042795900977.

Yamagishi:1996:GCD

[Yam96] Y. Yamagishi. Global convergence of the Durand– Kerner method applied to the equation z^{3} = 0. Journal of Computational and Applied Mathematics, 70(1): 67–73, June 14, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042795001476.

Yamagishi:1997:SSO

[Yam97] Y. Yamagishi. Stable spiral orbits of SOR Durand– Kerner’s method applied to the equation x^{d} = 0. Journal of Computational and Applied Mathematics, 82(1–2):465–475, September 15, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042797000885.

Yang:1990:MMD

[Yan90] Daoqi Yang. Mixed methods with dynamic finite-element spaces for miscible displacement in porous media. Journal of Computational and Applied Mathematics, 30(3): 313–328, July 10, 1990. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042790902825.

Yang:1992:CMM

[Yan92] Daoqi Yang. A characteristic mixed method with dynamic finite-element space for convection-dominated diffusion problems. Journal of Computational and Applied Mathematics, 43(3): 343–353, December 2, 1992. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        037704279290020X.

Yang:1995:DDD

[Yan95]  Daoqi Yang.  Different

domain decompositions at different times for capturing moving local phenomena. Journal of Computational and Applied Mathematics, 59(1):39–48, April 28, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279400018V.

Yanez:1999:DOG

[Yán99]  Eduardo Yánez.  De-

formable orthogonal grids: lemniscates. Journal of Computational and Applied Mathematics, 104(1):41–48, April 15, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042798002489.

Yamamoto:1990:ENT

[YC90] Tetsuro Yamamoto and Xiaojun Chen. An existence and nonexistence theorem for solutions of nonlinear systems and its application to algebraic equations. Journal of Computational and Applied Mathematics, 30(1):

87–97, April 10, 1990. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279090008N.

Yang:1996:NMA

[YD96] Shi-Nine Yang and Wei-Chang Du. Numerical methods for approximating digitized curves by piecewise circular arcs. Journal of Computational and Applied Mathematics, 66(1–2):557– 569, January 31, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042795001913.

Yeyios:1992:TSA

[Yey92] A. K. Yeyios. On two sequences of algorithms for approximating square roots. Journal of Computational and Applied Mathematics, 40(1):63–72, June 12, 1992. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
037704279290042V.

Ye:1995:CNB

[YH95] Xiu Ye and Charles A. Hall.

The construction of a null basis for a discrete divergence operator. Journal of Computational and Applied Mathematics, 58(2):117– 133, March 30, 1995. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0377042793E0257M.

Yuan:1996:PCG

[YI96] J. Y. Yuan and A. N. Iusem. Preconditioned conjugate gradient method for generalized least squares problems. Journal of Computational and Applied Mathematics, 71(2):287– 297, July 27, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042795002391.

Young:1996:NCP

[YK96] David M. Young and David R. Kincaid. A new class of parallel alternatingtype iterative methods. Journal of Computational and Applied Mathematics, 74(1–2):331–344, November 5, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042796000301.

Yang:1997:LSF

[YL97] Suh-Yuh Yang and Jinn-Liang Liu. Least-squares finite element methods for the elasticity problem. Journal of Computational and Applied Mathematics, 87(1):39–60, December 18, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704279700174X.

Yalamov:1999:FAI

[YM99] P. Y. Yalamov and M. Mitrouli.

A fast algorithm for index of annihilation computations. Journal of Computational and Applied Mathematics, 108(1–2):99–111, August 15, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704279900103X.

Yumin:1998:ABS

[YMW98] Cheng Yumin, Peng Miaojuan, and Zhang Wu. On the asymptotic behavior of the solutions of a class of second order nonlinear differential equations. Journal of Computational and Applied Mathematics, 98(1):63–79, October 5, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042798001150.

Yamamoto:1995:NVS

[YN95] Nobito Yamamoto and Mitsuhiro T. Nakao. Numerical verifications for solutions to elliptic equations using residual iterations with a higher order finite element. Journal of Computational and Applied Mathematics, 60(1–2):271–279, June 20, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279400096J.

Yamatani:1998:REM

[YO98] Katsu Yamatani and Kohzaburo Ohnaka. A reliable estimation method of a dipole for three-dimensional Poisson equation. Journal of Computational and Applied Mathematics, 95(1–2):139– 151, August 28, 1998. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S037704279800082X.

Yao:1999:ILI

[YR99] Chengfu Yao and Jon G. Rokne. An integral linear interpolation approach to the design of incremental line algorithms. Journal of Computational and Applied Mathematics, 102(1):3–19, February 15, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042798002052.

Yuan:1996:NMG

[Yua96] Jin Yun Yuan. Numerical methods for generalized least squares problems. Journal of Computational and Applied Mathematics, 66(1–2):571–584, January 31, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042795001670.

Yun:1998:BIF

[Yun98] Jae Heon Yun. Block incomplete factorization preconditioners for a symmetric block-tridiagonal mmatrix. Journal of Computational and Applied Mathematics, 94(2):133–152, August 3, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042798000788.

Yildiz:1998:OCP

[YY98] B. Yildiz and G. Yagubov.

On an optimal control problem. Journal of Computational and Applied Mathematics, 88(2):275– 287, March 2, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042797001982.

Yang:1996:SQP

[YZY96] Boting Yang, Kecun Zhang, and Zhaoyong You. A successive quadratic programming method that uses new corrections for search directions. Journal of Computational and Applied Mathematics, 71(1):15–31, July 10, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042795002162.

Zhang:1999:CPQ

[ZC99] Caiming Zhang and Fuhua (Frank) Cheng. Constructing parametric quadratic curves. Journal of Computational and Applied Mathematics, 102(1):21–36, February 15, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042798002064.

Zhang:1999:CSP

[ZCG99] Guang Zhang, Sui Sun Cheng, and Ying Gao. Classification schemes for positive solutions of a secondorder nonlinear difference equation. Journal of Computational and Applied Mathematics, 101(1–2):39– 51, January 15, 1999. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042798001897.

Zarzo:1994:SPS

[ZD94] A. Zarzo and J. S. Dehesa.

Spectral properties of solutions of hypergeometrictype differential equations. Journal of Computational and Applied Mathematics, 50(1–3):613–623, May 20, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042794903328.

Zarzo:1990:NSR

[ZDR90] A. Zarzo, J. S. Dehesa, and A. Ronveaux. Newton sum rules of zeros of semiclassical orthogonal polynomials. Journal of Computational and Applied Mathematics, 33(1):85–96, December 17, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042790902582.

Zeilberger:1990:HSA

[Zei90] Doron Zeilberger. A holonomic systems approach to special functions identities. Journal of Computational and Applied Mathematics, 32(3):321–368, December 10, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/037704279090042X.

Zeilberger:1992:GCG

[Zei92] Doron Zeilberger. Gauss’s _{2} F _{1} (1) cannot be generalized to _{2} F _{1} ( x ). Journal of Computational and Applied Mathematics, 39(3): 379–382, May 8, 1992. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/037704279290211F.

Zellner:1995:BNB

[Zel95] Arnold Zellner. Bayesian and non-Bayesian approaches to statistical inference and decision-making. Journal of Computational and Applied Mathematics, 64(1–2):3–10, November 30, 1995. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/037704279500002X.

Zeng:1995:SNC

[Zen95] Jiang Zeng. The q-Stirling numbers, continued fractions and the q-Charlier and q-Laguerre polynomials. Journal of Computational and Applied Mathematics, 57(3):413–424, February 28, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042793E02114.

Zvan:1998:PMA

[ZFV98] R. Zvan, P. A. Forsyth, and K. R. Vetzal. Penalty methods for American options with stochastic volatility. Journal of Computational and Applied Mathematics, 91(2):199–218, May 4, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042798000375.

Zhang:1991:PSP

[Zha91] Jian-Wu Zhang. A parametric study of pinched orthotropic cylindrical shells. Journal of Computational and Applied Mathematics, 35(1–3):319–330, June 26, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
037704279190219A.

Zhang:1992:IMG

[Zha92] Jian Guo Zhang. An iterative method of global convergence without derivatives in the class of smooth functions. Journal of Computational and Applied Mathematics, 43(3):273–289, December 2, 1992. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279290015P.

Zhang:1994:SPV

[Zha94] Wen Zhang. The starting procedure in variablestepsize variable-order PECE codes. Journal of Computational and Applied Mathematics, 53(1):73–86, July 29, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
037704279200121O.

Zhan:1995:BIS

[Zha95]  Yinwei Zhan.  A bi-

variate C ^{3} interpolation scheme. Journal of Computational and Applied Mathematics, 61(2):179– 188, July 31, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279400066A.

Zhang:1996:CEM

[Zha96a] Jun Zhang. A cost-effective multigrid projection operator. Journal of Computational and Applied Mathematics, 76(1–2):325–333, December 17, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042796001173.

Zhang:1996:SNC

[Zha96b] Jun Zhang. Symbolic and numerical computation on Bessel functions of complex argument and large magnitude. Journal of Computational and Applied Mathematics, 75(1):99–118, November 12, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042796000635.

Zhang:1998:MLM

[Zha98]  Jun Zhang.  Multi-level

minimal residual smoothing: a family of general purpose multigrid acceleration techniques. Journal of Computational and Applied Mathematics, 100(1):41–51, November 30, 1998. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042798001332.

Zhengda:1993:NKT

[Zhe93] Huang Zhengda. A note on the Kantorovich theorem for Newton iteration. Journal of Computational and Applied Mathematics, 47(2): 211–217, ???? 1993. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279390004U.

Zhedanov:1997:RST

[Zhe97] Alexei Zhedanov. Rational spectral transformations and orthogonal polynomials. Journal of Computational and Applied Mathematics, 85(1):67–86, November 6, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042797001301.

Zhedanov:1998:CLB

[Zhe98] Alexei Zhedanov. The “classical” Laurent biorthogonal polynomials. Journal of Computational and Applied Mathematics, 98(1):121– 147, October 5, 1998. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042798001186.

Zhedanov:1999:MCK

[Zhe99]  Alexei Zhedanov.  A

method of constructing Krall’s polynomials. Journal of Computational and Applied Mathematics, 107(1): 1–20, July 15, 1999. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042799000709.

Zhou:1994:RAR

[Zho94] S. P. Zhou. Rational approximation rate for the Müntz system {x_{n}^{λ}} with λ_{n} ↘ 0. Journal of Computational and Applied Mathematics, 53(1):11–19, July 29, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0377042792001302.

Zhou:1997:ECM

[Zho97] Ping Zhou. Explicit construction of multivariate Padé approximants. Journal of Computational and Applied Mathematics, 79(1):1– 17, March 3, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042796000957.

Zhu:1997:BIP

[Zhu97] Ping Zhu. On Birkhoff interpolation by polynomials in several variables. Journal of Computational and Applied Mathematics, 85(2):263–270, November 12, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042797001222.

Zielinski:1990:APN

[Zie90] Ryszard Zieliński. An aperiodic pseudorandom number generator. Journal of Computational and Applied Mathematics, 31(1): 205–210, July 24, 1990. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/037704279090352Z.

Zingano:1999:NSU

[Zin99] Paulo R. Zingano. Nonlinear L^{2}-stability under large disturbances. Journal of Computational and Applied Mathematics, 103(1):207– 219, March 15, 1999. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042798002532.

Zitnan:1994:RRM

[Zit94] Peter Zitnan. The Rayleigh– Ritz method still competitive. Journal of Computational and Applied Mathematics, 54(3):297–306, October 20, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042794902526.

Zisis:1990:TDS

[ZL90] V. A. Zisis and E. G. Ladopoulos. Two-dimensional singular integral equations exact solutions. Journal of Computational and Applied Mathematics, 31(2):227– 232, August 31, 1990. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279090165V.

Zhang:1996:CSI

[ZL96] Jianzhong Zhang and Zhenhong Liu. Calculating some inverse linear programming problems. Journal of Computational and Applied Mathematics, 72(2):261– 273, August 13, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042795002774.

Zhang:1999:FSI

[ZL99] Jianzhong Zhang and Zhenhong Liu. A further study on inverse linear programming problems. Journal of Computational and Applied Mathematics, 106(2): 345–359, June 30, 1999. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042799000801.

Zhang:1996:EBA

[ZLQ96] J. M. Zhang, X. C. Li, and C. K. Qu. Error bounds for asymptotic solutions of second-order linear difference equations. Journal of Computational and Applied Mathematics, 71(2): 191–212, July 27, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0377042795002189.

Zhang:1997:CGP

[ZM97] Zhimin Zhang and Clyde F. Martin. Convergence and Gibbs’ phenomenon in cubic spline interpolation of discontinuous functions. Journal of Computational and Applied Mathematics, 87(2):359–371, December 23, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042797001994.

Zhong:1995:PAC

[ZP95] Chen Zhong and Fei Pusheng. A parallel algorithm for constrained optimization problems. Journal of Computational and Applied Mathematics, 61(2):225– 230, July 31, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0377042794000648.

Zarzo:1993:FOD

[ZRG93] A. Zarzo, A. Ronveaux, and E. Godoy. Fourt-order differential equation satisfied by the associated of any order of all classical orthogonal polynomials. A study of their distribution of zeros. Journal of Computational and Applied Mathematics, 49(1–3):349–359, December 31, 1993. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/037704279390168B.

Zafer:1998:TSE

[ZT98] A. Zafer and H. Taseli. Twosided eigenvalue bounds for the spherically symmetric states of the Schrödinger equation. Journal of Computational and Applied Mathematics, 95(1–2):83– 100, August 28, 1998. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042798000739.

Zhu:1999:NMV

[ZT99] Gong-Qin Zhu and Jie-Qing Tan. A note on matrix-valued rational interpolants. Journal of Computational and Applied Mathematics, 110(1):129–140, October 15, 1999. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042799002034.

Zenisek:1995:ABH

[ZV95] Alexander Zenísek and Michèle Vanmaele. Applicability of the Bramble– Hilbert lemma in interpolation problems of narrow quadrilateral isoparametric finite elements. Journal of Computational and Applied Mathematics, 63(1–3):109– 122, November 20, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042795000534.

Zhou:1997:NST

[ZW97] S. Zhou and D. R. Westbrook. Numerical solutions of the thermistor equations. Journal of Computational and Applied Mathematics, 79(1):101– 118, March 3, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042796001665.

Zhao-Yong:1997:CNB

[ZYCL97] You Zhao-Yong and Wang Chuan-Long. A concept of nonlinear block diagonal dominance. Journal of Computational and Applied Mathematics, 83(1):1– 10, September 30, 1997. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0377042797000678.

Zhaoyong:1991:BAC

[ZZ91] You Zhaoyong and Xu Zongben. Ball algorithms for constructing solutions of nonlinear operator equations. Journal of Computational and Applied Mathematics, 38(1–3):457–471, December 23, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/037704279190189Q.

Zhang:1994:PQN

[ZZ94] Jianzhong Zhang and Detong Zhu. A projective quasi-Newton method for nonlinear optimization. Journal of Computational and Applied Mathematics, 53(3):291– 307, August 30, 1994. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0377042794900582.

Zahle:1996:FDW

[ZZ96] M. Zähle and H. Ziezold.

Fractional derivatives of Weierstrass-type functions. Journal of Computational and Applied Mathematics, 76(1–2):265–275, December 17, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042796001100.

Zhou:1997:SFE

[ZZ97] Lei Zhou and Tian-Xiao Zhou. Stabilized finite element methods for the velocity-pressure-stress formulation of incompressible flows nonlinear model. Journal of Computational and Applied Mathematics, 81(1): 19–28, June 1, 1997. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042797000022.